A Complete Bibliography of Publications in Journal of Computational Physics: 2010–2014

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 March 2023 Version 1.75

Title word cross-reference

(1 + 1) [HPA14a]. (k) [NS11a]. 0 [MVCFM13]. 1 [BGR11, BCL13, DK14, FNGV14, GM10a, HvB11, MMR12, MZB12, MMT14, MHGNM12, MGN14b, Ols12, RCQL11, UJ11, ZMD^{+}11, vTCL^{+}10]. 1 + 1 [DBAV14]. 100 [YM14]. 12 [KM14b]. 2 [AW14, AH10b, BBAP14, BCL14, BSS11, BHM14, BLY10, CK11, Cha14, CL11a, CLAL14, Del11, DFD14, DCL11, GW13b, GC11, GHM14, HBM13, HLW13, Hu13, JMGN13, KBS14, LVFK14, LT12, LS13d, MAPP13, MMR12, MDTC13, MHHGN14, MV13, MBGW13, MLGN12, Ols12, SY12a, SCH13b, SM12c, SAA14, TR14, TGGT13, UWB12, WR11, YCY11, Zha10a, ZOWZ10]. 2.5 [KL12a, MBMV13]. 3 [AST13, ABMT14, BLM^{+}14, BWC11, BS14a, BGHM13, BLJ11, BD14b, CP14, CCW12, CC12, CH14, CGF^{+}13, DDL13, FBM14, FW12, FCG10, FGLB14, GW13b, Ge10, GCZ13, GMC13, GD10, GL12a, GLC^{+}11, GKS^{+}11, HRT11, HLYS11, JM10, KVM12, KKS13, LH10a, LP13a, LTSH10, LEH12, LLP14, LMBB11, LDA10, LL10b, LZZR10, MMM11, MVCFM13, MF13, MT14b, Nic10, NJ12, PPS12, PWB^{+}14, PBdGP14, RKGM14, SKN13a, SY14a, SCGE13, SF14, SCC14, SYE10, TGT^{+}10, TQW14, TDL^{+}14, TRSMM^{+}13, TGGT13, WYYK10, WG11, Xu11, Zha10a, ZYF^{+}10]. 3D [BPM13]. 4 [CS14c]. 6 [KM14b]. [1 + 1] [Hen13]. _{3} [Ber14]. _{540} [WPKK12]. A [NDBG14]. α [DR11, PLT14]. αβ [KRF12]. Bi [SRN13]. C^{0} [DSZ13, HLLW11, ZTG^{+}13]. C^{1} [ZTG^{+}13]. CFL > 1 [MHGNM12, MHHGN14]. d [TE14]. d^{2}/dx^{2} [LR11]. D_{4} [LB14]. δf [CPCU13]. e [KBSV14]. ℓ_{1} [YK13, PHD14]. ϵ [ECDB14]. η [Fan14]. F = 2 [Wan11b]. H [ZA14, BHL12a, CL10, VSK10]. hp [VSK10, CYK^{+}14]. k [CFW^{+}13, ECDB14, HPS13, HHL11, TE14, WPK14]. k · p [And13]. L^{2} [HLY12]. L^{∞} [CCKCG10]. L^{p} [XY14]. μ(I) [CM14]. N [BGZ12, CDSP13, Don14a, TPA^{+}13]. O(N log^{2} N) [WWS10, WW11]. ω [HHL11]. p [ABK13, BGR11, CN12, CVN13, KRT14, MHA10, MMD^{+}11, PVLT10, TBR13, VSK10]. P 1_{DG} [CH11b]. P 2 [CH11b]. P_{1} [Ols11]. P_{N} [OLFS13]. q [MSTL14]. QR [SYL12]. r [BBPC14, VW14a]. rz [BLA^{+}14, KBS14]. S [BWMG14, YS12b]. S_{2} [BMR14]. S_{N} [RGK12, BMR14, CBTR13, TR14]. SP_{N} [ZRM13]. SU(2) [CB11]. T [IFD^{+}12]. τ [FdVV12]. z [VW14a].

-adaptive [BHL12a, BBPC14, CN12, CVN13, CL10, TBR13]. -body [BGZ12, CDSP13, TPA^{+}13]. -Cylindrical [KBS14, BLA^{+}14]. -D [HLYS11, KKS13, CC12, DDL13, DK14, FCG10, FGLB14, LVFK14, MMR12]. -D-NSCBC [CGF^{+}13]. -DG-approximation [RGK12]. -eigenvalue [CFW^{+}13, WPK14]. -enriched [MMD^{+}11]. -estimation [FdVV12]. -matrix [IFD^{+}12, YS12b]. -minimization [PHD14]. -model [BWMG14]. -multigrid [BGR11, MHA10, PVLT10]. -Navier [KRT14]. -particles [PLT14]. -phase [Don14a]. -refinement [MSTL14]. -stability [CCKCG10]. -stable [NDBG14]. -th [HPS13]. -theorem [ZA14]. -version [ABK13].

1 [FNGV14, RS13b, SG12a]. 1A [XDZ^{+}14].

2 [RS14b, SG12b]. 228 [HY11]. 229 [Dad12, Dav15, NHT^{+}13]. 230 [Gho17, HSK^{+}15]. 231 [MPPP12b, TK15]. 233 [SS14b]. 237 [OBNN14]. 241 [YKWS13a]. 242 [CFCA13a, LWD13a]. 252 [XMDG14]. 259 [PS15]. 262 [EH15, XS15]. 265 [SYV17]. 270 [ZJS15]. 2F [LL10b].

3-node [LR11]. 3d [LH10b]. 3rd [PGL11]. 3rd- [PGL11].

4OC [Sen13]. 4th [PGL11]. 4th-order [PGL11].

5th [Ler14].

6th [Boe11].

7th [Ler14]. 7th-order [Ler14].

9347-9372 [NHT^{+}13].

ab-initio [SD14a]. Absorbing [ALS^{+}14, HGRB14, SG10b, SG10a, ZW10b, AMP14, BDT10, GD10, MBL^{+}14, RVbZ10, RCCC14, SG12a, SG12b, XS13, ZT13a, BGH10]. absorption [Den11b, FJ12, SCS14]. AC [HDD14, KBSV14]. accelerate [Yao14]. Accelerated [Hut14, KG12b, SCGE13, VDP11, VBSK10, WP14, BD14a, FTF14, GK13, GLC^{+}11, GD13, IFD^{+}12, KSP13, MdVBS13, NVS13, PPB14, QA11, RCQL11, RBCK10, SBB13, STC10b, Tak14, WK13b, Fat10]. Accelerating [JLS^{+}14, KSS14, NS14a, ZVNM10, ARR13, SS13a]. Acceleration [KSPP10, YM14, CFW^{+}13, EMSH14, FWP13, FWLP14, HNMM10, PMMB14, TR14, WPK14, YXD^{+}14, WTK14]. accelerator [CBCM^{+}11, LCNK10]. accelerators [TAAY14, VGCMG11]. accommodation [LH13]. Accuracies [XL11]. Accuracy [DB13, JCT11, LTC14, MZ11d, SK12a, BK11, BS14a, BSS14, BKV12, CFKL11, CJP^{+}14, CBGK13, CR14, DMP14, GHS13, GH10, GWCA14, HZ14, Ide12, KS11b, KG11, KM14b, LB12, LP14a, LT11b, LS13d, MB10a, Nic10, SRSV11, Wal13, YWL14, PZW14]. Accurate [BR12a, CF11b, GRS14, LL12a, MC12, MJ13a, MGN14b, SG12a, SG12b, SCC14, ZYHW13, AN13, BDA12, Boy11, Bre12b, CY12, DLC11, DKC14, DV14, GM12a, GM10b, GHM14, GH11b, GL12b, GVV13, HOK10, HK11, HF10, KS13a, Le 14b, Lee13, LKLG14, LdWF12, LT10b, LN10b, LLP10, MF14, MB12, MCA^{+}13, MH10, ML13, MB10b, NMZC13, NDBG14, NPV14, NT14, Per13, RE13, RWMG11, RMMD14, SHMS13, Sco12, Shu14, SBB10, Son11, SdAW14, WBC13, WYT14, XW14, YS13, ZK14, ZF14]. Accurately [Liu11a]. ACE [VBSK10]. acetylcholinesterase [Ham10]. achieve [VSK10]. acid [EHA14]. Acoustic [Lee12b, WdHX10, AH13a, CLG13, CVI^{+}13, FT11, FvdB12, GB14, GH11a, HVD13, HW10, KGG10, KNR11, LS13d, LM14b, MKB^{+}12, SM11b, WSBG10]. acoustics [KVM12, MBD13, NPC11a]. acquired [PMM^{+}11]. across [GGH14]. action [Wan11a, Wan13a]. activation [RJC13]. active [ASGW11, MdVBS13]. Ad [Ano12-32, BL10, LB10]. adaptation [ABD14, BDA12, CXLC12, Del11, FL11a, HKL10, KF13a, KK10b, LH10c, LDA10, MM10, MVG^{+}12, TG14b, YD12]. adapted [DO11b, HRBK10]. Adapting [KSB12, EAN11]. adaption [CSB14]. Adaptive [AHOT14, ABK13, BLOdV14, BCK11, BUH11, BS11b, CMG^{+}13, CJN^{+}13, DO11a, DMMGM14, HJ11, HZ11a, HHA14, KMHJ10, LTCN13, Lee14b, LLYE12a, MAPP13, MMD^{+}11, RVFK14, SR13b, SQC11, SYL12, SK12b, TvdHS^{+}12, WG11, YCLK12, YMLM^{+}14, ZHS10, AL10, AS14, Atz10, BHL12a, BDMP13, BBGP13, BK12b, BDLGC13, BBPC14, CN12, CVN13, CL10, CNR10, CGG12, CF10, CXZ10, CY14, CEL14, CKR14, Cou11, DMHL14, DDF14, DV13b, DZHB13, EQYF13, ETW10, FHL10, Fuj11, GG10, GM14a, HLT10a, HHL11, HG11, HLDP14, HHK13, HB13, HL12b, HWA10, Hu13, ID10, IDNG13, IG14, JU10, JAX11, JLY10, JHF10, JLC14, JK10, JG13, KVM12, KWSJ11, KMvM^{+}12, KA12, KG14, KSD12, KL12b, KL12c, LSM14, Lap11, LBT11, LHY13, LZ13, LLZ14, LX14, LN10c, MZ10, MS14b, MZ11c, Min13, MTG11, MG14b, MN14]. adaptive [MIKG12, MNL^{+}13, MJ13b, MBGW13, Mur10, OS12, PZW14, PZG14, PHRG13, PMM^{+}11, PRA13, PGLK14, PB13b, PWB^{+}14, Phi14, PHSA14, Qua11, RSS12, RE13, RWX11, RXLS14, RPA13, RS10b, SSHM10, SSB^{+}12, SFWP10, SWR11, SD14c, TE14, TA12, TDVG13, TAAY14, TBR13, UKP14, VRB10, Wan11a, Wan10, WR11, WS11, YLW10, YJK12, ZMQ13, ZN14a, ZN14b, ZDSP11, VV13, NT14]. adaptive-resolution [RSS12]. Adaptively [XZWW12]. Adaptivity [CHZ13b, AM12, DH10, DFVY14, DBM^{+}12, GBB^{+}13]. added [LJG14]. added-mass [LJG14]. addition [Far11]. additive [HJZ14, HWHW10]. ADE [FL13]. ADER [BD14b, BBD14, BMD^{+}13, BK14b, CTJT13, DZHB13, MMT14, MT14a, NF12, Nor13, Nor14]. ADER-DT [Nor14]. ADER-Taylor [NF12]. ADER-WENO [DZHB13]. adhesion [GWG14, RUL^{+}13]. ADI [BdF10, FF12, KAA12, SY11c, SN13b, ZFH14, Zha14]. adiabatic [BKH12]. Adjoint [CJNS12, HHL11, HLSO13, IWG13, KJC13, SR13d, ZPGO10, AS14, BSCML13, BW14, DRB14, GHM13, SBS11, SM10, VS14, Wan13b, YDN10, dPSS12]. Adjoint-based [HHL11, IWG13, VS14, YDN10]. adjustable [DGAH10]. adjustable-damping [DGAH10]. adjusting [Bal12a]. admits [ZA14]. adsorption [FCH13]. Advanced [SJB11]. Advances [LMLC12]. advancing [RSD10]. Advection [Dai13, AV13, Ald10, BR14, CD13b, DF10a, HAQ12, JW11, JZ13, KKAS12, KLMJ12, KWJR14, KJWR14, LN10a, LGP10, LSV10, MKGV12, MMP11, MMA11, MMA13, MT14a, NRS10, Nis10, Nis14b, QS11a, RSS12, Sou14, TLG10, WSYS12, WYS12b, WCVJ13, ZOL^{+}11, vdVR12a, vdVR12b]. advection-diffusion [KKAS12, LN10a, LSV10, MKGV12, Nis14b, RSS12, WCVJ13]. advection-diffusion-reaction [JZ13, MT14a]. advection-dispersion [CD13b, JW11, MMP11]. advection-dominated [TLG10]. advection-electrodiffusion [LGP10]. advection-propagation [Ald10]. advective [LDE^{+}13, SM12a, SM12b, VTTK13]. advective-reactive [VTTK13]. AEGIS [ZKV10]. aero [Whi10]. aero-optics [Whi10]. aeroacoustics [CR14, Fal13, KKS13, SK13, XS11, LLH10]. aerodynamic [KWS^{+}11, Lan14, LH10b, RWMG11, ZPGO10]. aeroelasticity [DYYA14]. affinity [JJM13]. Affordable [MT13]. after [GSTB13]. against [GMXG14, OAXL13]. agent [JJM13]. Agglomeration [Lan14, BBC^{+}12]. aggregate [BRS13]. aggregation [OS11a]. air [BCD14, CKS14, KED^{+}12, Nag14a, Nag14b]. air-water [BCD14]. air/water [CKS14]. aircraft [JWNL11]. airfoil [BBD11, DYYA14, JM12]. al [Dad12]. ALADINS [VV13]. ALDM [HAD10]. ALE [BKC^{+}11, FL14, GMB10, JLZ13, KPH13, KGSS10, LMQ^{+}11, MSB14, SYY14b, VKL^{+}13, WKL^{+}14, WK14, WC14b]. ALE-based [LMQ^{+}11]. ALE-DGSEM [WK14]. Algebraic [Gui12, WHT14, CK11, GW10, SSG11, SX14, TCW^{+}10, XIC11, Yao14, VV13]. Algebraically [KBSV14]. Algorithm [YK12, ARR13, AGBC14, ADMA13, AW14, AT13b, BK12a, BFR13, BHS13, BHS14a, BHS14b, BWC11, BILM14, BBAP14, BAG12, BKW10, BP10, CCB13, CdCNH14, CB12, CY12, Cha14, CRS14, CCB11, CCB12, CC13, CC12, Dad12, DBO^{+}10, Don14a, Don14b, DFW^{+}10, EWH14, FHL10, GMT10, Gen11, Gho17, GL13, GV13, GVH14, GJ14b, HAMA14, HSK^{+}15, HS13b, HLZ13, HF10, HWHW10, IFD^{+}12, Ian11, JNSA12, JTH10, KM11, KM13, KSPP10, KO12, KL14b, KS12, Le 14a, LKT^{+}12, sLwG10, LD12, LWD13a, LWD13b, LLP14, LG14, LXSR12, LBZ^{+}10, LX13b, LvZB10, LTL11, MGJZ10, MW12, MMM13, MCJ14, Men12, MBLM10, Mét11, MDM^{+}12, MW14a, MWY14, NS11a, Nic10, NS14b, OM12, PZW14, PZG14, PVLT10, PGL11, PE10, PT12, RCW11, RKM11, RH13, RBNS13, RS13a, SBT14, SNS10, SM12a, SNBN13, SSHT14]. algorithm [SAHS14, ST13, Shy10, SX14, SC10, SFWP10, SKt10, SRN13, SYE10, TBMH11, TRSMM^{+}13, TY11b, TR12, TSR13, Tyl14, VDD10b, VRBZ11, WFL10, WBC13, WZ10, XIC11, YCY11, Zad11, ZJL^{+}14, ZL13, ZWH^{+}12, vdVR12a, vdVR12b, FT11]. Algorithmic [Nor13]. Algorithms [DRW11, AKP^{+}12, BHBM10, CWL13, CW13, CJNS12, CHZ13a, CHHL11, CBTR13, Del14, DR14, DAEB13, ES13b, Gio10, GCW14, GKS^{+}11, HB13, JEU11, JGZL14, KN14, KIG11, LGE^{+}13, LWS14, MPWK13, NS11b, NPV14, PP14, SMP^{+}13, SSR^{+}13, TvdHS^{+}12, Tyg10]. alignments [SHFB12]. All-at-once [SW13]. All-electron [SL13a]. all-speed [CDT13, CDK12, sLwG10]. Allen [BSS12, GLWW14, RNT12]. Allmaras [CDB13]. allowing [MKB^{+}12]. almost [Pie12]. along [Gri13]. Alternate [Ols12]. Alternating [LX13a, LXL13, LP14b, nZzS11, Cui12, WW11, WD14]. alternating-direction [WW11, WD14]. alternative [Boo11, SPM^{+}13]. aluminium [BGK12]. Ambipolar [PMS11, PvDGvM11]. amino [EHA14]. amino-acid [EHA14]. Ampére [BCW13, BFO14, CCZ14a, CSC14, FO11, ZPF10]. amplitudes [LQ11]. AMR [CNSFD13, CN10, MV13, SRN10]. analogs [Ber14]. analogy [Cam13]. Analysis [Del10, DBHBB13, HdCNT12, KKvZB14, KG14, LH14a, LMS11, Man12, PGL11, PB13a, SS12, SRSV11, SSFM12, SKt10, TPA^{+}13, XS13, YBK13, ABKF11, APY14, APRS14, BHS14a, BHS14b, BA10, BAB^{+}13, BLA^{+}14, BW14, BB12b, BP13a, BRSS11, CWS13, CTG12, CM13, DRB14, DWL10, Den11b, Don10, DZL13, DB12, DEKBF14, EHA14, EC13, GB14, GLSC12, GCLS11, GVH14, GZQ13, Ham10, HPS13, HBAP10, HLSO13, Hig13, HVT12, IH13, IWG13, KWJR14, KJC13, KGG10, KWHD12, LSK11, LCK11, LJ13a, LD14, LDE^{+}13, LLS11, MZ11a, MS10a, MS14b, MZ11d, PGLK14, PVDW14, QS11b, RVbZ10, RB10, Rot13, SK12a, SM12b, SHFB12, SSR^{+}13, Soa13, SLC^{+}11, SG13, TNK10, TCL11, UJ11, VS11, VCJ11, WXZ10, WHB14, WIJ13, WH11, WG11, WL11, XMS12, XMDG13, XMDG14, ZKV10, vdVR12a, CRS12]. analysis-suitable [XMDG13, XMDG14]. analytic [DDL13, LW13b, RBNS13, TR12, WLW14]. Analytical [JR14, CC13, CSKP10, DFD14, GVH14, LVFK14, PBB14, XX13]. analytical-based [CSKP10]. analyze [HDPF11]. analyzing [HVD13]. anastomosis [GDY10]. anatomic [LS13a]. Anderson [WTK14]. aneurysm [GFK13]. angle [BS11b, DL10c, Gon12]. angular [BM10a]. animated [LS13a]. Anisotropic [MVG^{+}12, ABD14, AT13a, BWC11, BGH10, BDA12, CdCNH14, Cha14, Cou11, CJN^{+}13, DLNN12, EZ10, GH14, GW13b, HKL10, JS13b, LH10b, LH10c, LDA10, LHV13, MB12, MMP11, MTSG12, MN12, MBL^{+}14, OJ11, OK10, SS11, SG10b, SG10a, SG12a, SG12b, SH11, WBC13, WGD12, WG14, YD12, YN10, YB14, vEKdB14]. Anisotropy [LLMY14, MB13b, RR14, SG12a, SG12b, SRSV11]. anisotropy-free [RR14]. annihilation [SR13a]. annular [KTT12, Mit13]. anomalies [BD14c, OAXL13]. anomalous [MAD13, WV14]. ANOVA [LLZ14, YCLK12]. ANOVA-based [LLZ14]. antenna [ZLL10]. Anti [SHA11, SHA12, KL10b]. Anti-diffusion [SHA11, SHA12]. anti-diffusive [KL10b]. any [GPS10]. aortic [LS13a]. aperiodic [PMSM12, PS14b]. apertures [BL13]. applicability [HGSK^{+}13]. applicable [Sen13]. Application [AC11, Bal10, Bal12b, Bal14, BDLGC13, CYK^{+}14, CTP13, CRS14, CZ11a, FGE11, GH12a, GP12a, GK10, Ham10, KBRM12, KWS^{+}11, KMA13, KBGRM10, Kol11, KL14b, LHH14, MS10b, Mar10b, MPWK13, MM12, MRS14b, MHGNM12, MGN10b, MGN12b, NHE14, NMZC13, PPS12, RTS13, SdAW14, TDVG13, Val14, VS11, VBCM10, ZYLK13, BDA14, BC10a, BG10, BDTW10, CFCA13a, CFCA13b, CL14b, CAH14, CDLL12, GLSC12, GJ14a, GKG^{+}13, Hen13, HHL14, ISZ12, JLY10, KKS13, KST13, KKvZB14, KG14, LYE10, LTSH10, LH14a, LG14, LHV13, MJ13b, OBNN13, OBNN14, OD14, PRH^{+}14, PL12, QS11b, SSS11, SF14, VDD10a, WC14b, ZFJ11, ZTL13, DFNNRdlA14, MGN14b, TVEC12, VSC^{+}11]. Applications [BZ12, LRK11, ASGW11, AH13b, BGN10, BW13, BCW13, CMO14, CS12b, CZ13, DDF14, DMT^{+}11, DMM^{+}13, DMR14, hGzSwZ14, GCBH10, GKN^{+}11, JMW12, JKQ13, KL14a, KIG11, LGG^{+}13, MGN12a, NIN^{+}14, NB13, OM12, QL11, SS12, SLH^{+}11, SST10, TDL^{+}14, WZ10, XQX13, YS12b, ZLH12, CCFCM11, FR13]. applied [AAI10, BLOdV14, BDKV14, BJ12, CFW^{+}13, CRT10, CG13b, Del10, DOR10, DG12, EWH14, FH11, FY13, GVV13, KKC^{+}13, Lan14, LVFK14, LLS11, LL10b, NLGL14, NDBG14, NKF14, NVS13, SK12a, Stu10, TD12b, VWF13, VG14, WR11, ZHW10]. Applying [KS11c, PSL20]. Approach [TK12, TK15, AC13, ABL^{+}11, ABMT12, ACB^{+}13, BDL11, BFR13, BW13, BPS14, BR11, BCL14, BCL^{+}11, Boo11, CGC13, CMS10a, CGG12, CP14, CHKT13, CPX13, CW14a, CKR14, CST10, CRS12, Don14b, DL10b, EBGW13, ETW10, FM13, FLM12, FMT11b, GLLP14, GPCE13, GCE14, GZQ13, GMXG14, HD13, HMHO13, IAM^{+}12, JEX10, JYHT10, KKL10, KBTD12, KKvZB14, KMA13, Leu11, LWF11, LM13, LHPH11, LGH10b, MB13a, MCD14a, MVS13, MP13, Mön13, MKB^{+}13, MLBK14, MN14, NTV14, Nis10, OK14, OLPM14, PHRG13, PGLT10, PRA13, PHD14, PB13b, dlLPC10, QA11, RUL^{+}13, RWX11, RKGM14, RTKS10, RDHK12, SBC^{+}14, SSB13, Sam14, SMF^{+}13, SCIE14, Sco12, SD14b, SAHS14, SX14, ST11a, SGV13, SIT^{+}11, TGO13, Vai14, WZ13, WLW13a, YVL^{+}11, YKWS13a, YKWS13b, ZT13a, ZF12, ZCD10, ZJT10]. approaches [KMvM^{+}12, LCY11, ZL10]. Approximate [Abr10, EAG14, MS11, PR14a, hGzSwZ14, NGV14, Ren11, Rie11, SGL14, UN13]. Approximated [GL14]. Approximating [RSS10]. Approximation [ABPSM11, BGK12, Chi12, CH11a, Gra14, PBC11b, RPA13, AHA10b, AR14, AM12, Alm14, APC11, BGGGS11, BCP13, BHBM10, DDS12, DST14, DO11b, FP12, hGzS13, GHDS10, GS11a, Hua14, KL12a, KL14a, Kaz10, Lee14b, LRBT14, LLP10, ML14b, RGK12, RK14, SBN12, SCC14, SBO11, TR11, VP10, WMI10, Whi10, WK14, Wol12, YR13, ZN10, ZK13, ZL10, Zho14b, vHtTBC11]. approximations [AKP^{+}12, AC11, BS14b, BWMG14, Boy11, BHY14, EZ10, Gri12a, HW11, Hig13, HLYS11, JTXZ12, KDS14, KAFB11, KJ11, MDTC11, Mat14, MK11, Pir10, Pir11, VTTK13, ZT10a]. APU [CRK12]. aquarium [Nag14a, Nag14b]. aqueous [DXB^{+}13]. Arbitrarily [GCH14, MTD14, NS11a, RTKS10, Zha10b]. arbitrarily-oriented [MTD14]. arbitrarily-structured [RTKS10]. Arbitrary [DSZ13, GT12, GIQ11, KS14, LX12, OS14, ARR11, AEM13, BHS14c, BBG^{+}11, BR12a, BRSS11, BWMG14, CR13, DvW14, DFD14, ES13a, FK14, GW13b, GPY14, GH12a, GT11, HZS12, KBS14, KTT13, KS12, LK14, Li13, LS13c, LS10b, LSVW10, LS13d, LMS^{+}10, LLN^{+}10, MBMV13, MVS13, MGN14a, MGN14b, OS11b, Par12b, Pir12, RVB10, STC10a, SdAW14, TTD11, TAD14, TR14, VMS12, Wan10, WYC13, ZS14, ZJL^{+}14, BD14b]. arbitrary-geometry [GPY14]. arbitrary-Lagrangian [LMS^{+}10, BD14b]. Arc [KYP^{+}14]. architectures [HRBK10]. area [KG12a]. argon [KM14b]. arising [CHHL11, HY10, JHDC10, TCW^{+}10]. arithmetic [FMT11b]. ARM [GKG^{+}13]. ARM-based [GKG^{+}13]. Arnoldi [CHHL11]. Arnoldi-type [CHHL11]. Array [WZN14, KNR11]. Array-representation [WZN14]. arrays [Hel11a]. arterial [ADSF13, LL14a, XHF13]. Artificial [BHY14, OA10, TE11, Vai11, AOCR12, BD10a, Coo13, Deg11, hGzSnZ12, HZ11a, KSL10, Kaw13, KL12c, LS10b, OAKR14, OL13, RSS13, VW14a, Vre11, Wen10]. ash [MPS^{+}14]. aspect [BLR13b]. aspects [FJ12, Kwo14]. Asselin [LT14]. assemblies [LZC^{+}13]. assembly [KA13, RS14a, WT12a]. Assessment [JLB^{+}10, KSL10, MWRZ13, SOG14a]. assignment [GRT12]. Assimilating [BH13c]. assimilation [CS14c, GLS10, GCLS11, GHM13, Kaz10, NMX12, QA11]. assimilations [ACCM12]. associated [And10, SP11]. Association [AM13]. AstroBEAR [CNSFD13]. AstroGK [NHT^{+}13, NHT^{+}10]. Astrophysical [NHT^{+}10, NHT^{+}13, KSB14]. astrophysics [KHHK12]. Asymptotic [Boy11, CDK12, CLS13, Den11b, GJLY13, MN12, OA10, OLFS13, AC11, BCM12, BS10a, CdCNH14, DLNN12, DMR14, FJ10, JY11, Mie13, OK14, RLJ14, VBGM13, WLW13a, WW14, CM13, CLM13, DDN^{+}10]. Asymptotic-numerical [CLS13]. Asymptotic-Preserving [CDK12, GJLY13, MN12, CdCNH14, DLNN12, FJ10, JY11, RLJ14, DDN^{+}10].

Asynchronous [DA14, MK14a, OK12, SK13, Unf13]. atmosphere [IBO14]. atmospheres [FMM^{+}10]. Atmospheric [HMOH11, CP13b, KED^{+}12, KL12a, KG12a, KG14, KSD12, LA10, NMZC13, NMCZ14, NNS11, OBNN13, OBNN14, SSW13, SKW14, Ste14, UJ12, YMLM^{+}14]. atom [MMM11]. atomic [KJC13, MPPP12b, MPPP12a, RJC13]. Atomistic [GWK10, BLR13a, Gri12b, KMHJ10, ZJT10]. atoms [Ich13, JLS^{+}14, JMKK12, SSR14]. attenuation [BCL14]. Augmented [Ion13, MGN12a, SKCC10, SP14b, VSC^{+}11, BHK^{+}10, BV14, LDM14, NRS10]. AUSM [KS13a]. AUSM-family [KS13a]. Automated [DRB14, Sog14b]. Automatic [Dri10, GBC11, PB13b, ABJ^{+}12, DMR14, Kau10]. Auxiliary [SKN13a]. avalanches [FNGV14]. Average [CGM^{+}12]. average/point [XIIX14]. averaged [HAH14, SH14, YK12]. averages [Gil13, KSS14]. averaging [LM13, ECD14]. avoiding [HP13, Nor14]. aware [BHS14c]. axi [NCKN10]. axi-symmetric [NCKN10]. Axial [JK11, LK14]. axially [HK14]. Axisymmetric [BLA^{+}14, MA14, AW14, CMS10a, CRT10, DFD14, FGLB14, HKL14, HHL14, SSB13, ZKV10, ZGSZ10]. azimuthal [BdCB11].

B [BSV14, EH13]. B-spline [BSV14]. B-splines [EH13]. back [YZ12]. backflow [BC14]. backgrounds [FvdB12]. backward [LP13a, LQ10a, WZG^{+}13]. backward-facing [WZG^{+}13]. Bad [MMP11]. Baer [FP12, PZTW12, TT10]. bag [CB13]. balance [BKW10, CLS13, GGG^{+}13, LPE^{+}11, MAK14, MGN13, PWK11, Tok11a, VWF13]. balanced [BF12, CF10, DLGP13, DLM13b, DLM13c, FMT11a, FMM^{+}10, KM14a, LGLX14, MW14a, MPT13, MGN14b, TR12, WYS^{+}11, Xin14, Yok14, DMMGM14]. balanced-force [MW14a]. band [And13, BG10, BS14b, LDM14]. band-limited [BS14b]. Bandgap [MNF^{+}10, HCH^{+}10]. bands [BSS11]. bar [JXD14, RG12]. barrier [RJC13, Tok10]. base [MRS14b, VMS12]. Based [ABH^{+}14, ARR11, ARR13, AL10, AHOT14, AS12, BK12a, BCZ13, BD10a, BOD11, BBC^{+}12, BNM14, BWC13, BMK11, BMK12, BA10, BN14a, BERB^{+}12, BYCC14, BHK^{+}10, BS11b, BRSS11, BRS13, BRP14, BBD14, BGBR12, BDKV14, BCW13, CCKCG10, CNR11, CDF11, CK10, CM14, CXZ10, CP10, CBW10, CLZ12, CCL^{+}14, CLSX14, CY14, CHHS11, CGMQ14, CGS13, CKR14, CSKP10, CDS13, Cou11, Dad12, DV13b, DLNN12, Del11, DLC11, Den11a, DP12, DFD14, DKR12, DS11, DCL11, DLM13b, EH14, EH15, EL11, FGZ12, FGR12, FL13, FH11, FL11a, GDHM11, GC11, GG14, GL13, GKG^{+}13, GBB^{+}13, GBS^{+}14, GCL13, GHM13, GTSC14, GZQ13, HHS^{+}13, HPS13, HHL11, HBAP10, HK14, HHGC14, HMM^{+}10, HLSO13, HHK13, HH12b, HH13, HLZ14, HM13, HKL10, HMOH11, IAM^{+}12, ISZ12, IWG13, ITO^{+}10, JJM13, JTH10, JSX10, JWYH10]. based [KS11a, KF13a, KIF^{+}14, KKO13, KO12, LEM11, LGC13, LQ10a, LLZ11, sLwG10, LLX11b, LJ13a, LLZ14, LXSR12, LT11b, LSVW10, LS13d, LBM14, LMQ^{+}11, LMS^{+}10, LL12b, LL14b, LXL^{+}12, LXS^{+}13, MDTC13, MS14a, MT13, MSD12, MYM12, MBD13, Mön13, MF13, MMJ14, MGN10a, NTV14, NN12, OBT12, OK13, Par10, PBM11, PB13b, Per13, PT12, dlLPC10, RV10, RH13, RWX11, Ren11, RRCD14, SSS11, SP14a, SQC11, SS10c, SCB10, Soa13, SK12b, SGL14, sSLxRyZ14, SEBG12, SS10e, Tak14, TNK10, TD11a, UKP14, UGkM13, VBVD10, VB13, VS14, VWPF11, VWF13, VBSK10, Vol14, WZ14, WZ11, WLL13, WT13, WX14, WK13b, WZ10, WC14b, WG14, XZWW12, XS15, XPO14, XW10, XX13, YMS12, YDN10, YSW^{+}13, YF14, YSW14, YD12, YCH^{+}13, Yok13, YL14b, YYB12, ZVNM10, ZHPS11, ZHS10, ZGSZ10, ZD11b, ZMD14]. based [ZLL10, ZL13, vydCNS^{+}10, WP14]. bases [KL14a]. Basic [YS11a, ZWL^{+}12a, BL10]. Basics [DL13a, CMO14]. basin [LMQ^{+}11]. Basis [MPR12, TG14b, AB12, AHOT14, BFE12, Boy10, BY11, Boy11, BX12, CXLF14, CST10, EGW11, FHMS11, FL10, GHS12, HBM13, HHGC14, LH10a, LB14, LLYE12a, LLYE12b, NB14, PGLK14, PSDF13, Pir12, PH13, RA10a, RPA13, Sal14, ST13, SPM^{+}13, TCL11, ZBS12a, ZBS12b, ZD11b]. Basset [vHtTBC11]. bathymetries [RF14a]. Bayesian [AGS10, BZKL13, BH13c, ECDB14, ECD14, EM12, EWH14, HM13, JZ10, Jin12, KL14a, LMM14, RLPM12, Val14]. BCYCLIC [HPLS10]. BDF [Don10]. BDF-like [Don10]. BDF2 [EMK11]. beach [OBT12]. beam [AAI10, AETT11, BLQ14, GV13, JWYH10, Whi10]. beams [BM10a, LQ10a, Lip12, MdVBS13, QY10, WS14]. bed [CVC12, MGN10a, RB10, RB13]. Behavior [SYL10, MMP11, YKWS13a, YKWS13b]. behaviour [Ain14]. belief [CZ13]. BEM [CES14]. BEM-FEM [CES14]. benchmark [VCS14, BLM^{+}14]. Benchmarks [TSL^{+}14]. bending [FLM12, iT14]. Benjamin [BX12]. Bessel [BY11]. between [BHM11, BWC11, BCLR10, Deg11, FQV13, GKF11, GDGP11, KL10b, MDM^{+}12, MBGW13, Nag14a, Nag14b, Rem11, Rie10, SM12c, iT14, ZTG^{+}13]. Beyond [KVBP^{+}14, LE12]. beyonds [ARF12]. BGK [CL10, DL13a, DMR14, PP12, RHXQ14, XX13]. BGK/hydrodynamics [CL10]. Bhatnagar [MWRZ13]. Bhatnagar-Gross-Krook [MWRZ13]. bi [CHZ13b, CHZ13a, CSK14, DB12, HT10a, NB14]. bi-harmonic [NB14]. bi-Maxwellian [HT10a]. bi-orthogonal [CHZ13b, CHZ13a, CSK14]. bi-periodic [DB12]. biased [SST10]. biasing [BK12b]. bicharacteristic [DLM13b]. bicharacteristic-based [DLM13b]. bidirectional [KKC^{+}13]. bidomain [CEK14, Pie12]. BIE [HVD13]. bifurcation [GBC11, WIJ13]. bifurcations [CM13]. biharmonic [JRTY11, JKQ13]. bilayer [SBA13]. Billion [MMM11, ARF12]. Billion-atom [MMM11]. bimolecular [NMV13]. bimolecular-reactive [NMV13]. Binary [HH12b, HH13, CDS13, GZZ14, PBC^{+}11a, Sza12]. biological [GK10, TDL^{+}14]. bioluminescence [HQL^{+}10]. biomechanics [KVC^{+}13]. biomembranes [BNP10, ES10, LSM14, PRA13, RPA13]. biomolecular [LHMZ10, WRH10, Xie14a]. biomolecules [GK13]. biopolymer [CMBW13]. biorthogonal [dlLPCP12]. Biot [BCL13, FCG10]. blended [WA12]. blending [SGL14]. Bloch [ÁCB12, JWYH10, NLGL14]. Block [BMP14, Fat10, He13, HB13, CST12, Gao13, HPLS10, MA13, PJCS14, SSS10, SY11c, Zho10]. Block-spectral [He13]. Block-structured [HB13]. blood [BC10a, BCG^{+}13, MMT14, MPT13, WC14b, ZZK^{+}14, ZIOF10]. BlueGene [ARF12]. BlueGene/P [ARF12]. bluff [RBCK10]. blunt [NHE14]. Board [Ano12f, Ano12g, Ano12h, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano11z, Ano12b, Ano12c, Ano12d, Ano12e, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano13i, Ano13j, Ano13k, Ano13l, Ano14-34]. bodies [BHS13, BBGP13, BD12, BM11, De 10, MCA^{+}13, Mel11]. body [BGZ12, CDSP13, HSWZ13, HCS11, LKCY11, LY13a, MDM^{+}12, PB13b, RBCK10, SD14a, SCGE13, TPA^{+}13, WYYK10, WY11, XHF13, YL11]. boiling [TSL^{+}14]. Boltzmann [FRP14, MVS13, dv23, AJ14, BSDM12, CL10, CRS14, CLZ12, CHKT13, CGR11, CL13, CSB14, CLAL14, DD12, Del13, Del14, FY13, FSL10, GH14, GK13, GG14, GBB^{+}13, Gra14, GWCA14, HE14, HG11, HvB11, HW14a, HHL14, HW14b, HCM10, IF10, JNSA12, JYHT10, KH13, KMA13, LMLC12, LRT13, LL10a, LMK13, LP14a, LZC^{+}13, LZV12, LVZK14, MD13, MMR12, MB10a, MS10b, MS11, MS14a, Mel11, MZ11d, MZ14, ML11b, MTG11, MM11, MVG11, MHGM14, NSA11, PPB14, Pri14, RK14, RLJ14, RJG10, SMP^{+}13, SHFB12, SSC13, SSX13, SS14a, TLZ^{+}11, TRL14, VS11, VMS11, VMS12, VS14, VM10, WK10, WZG^{+}13, WZT^{+}14, WC11, WS10c, WS11, WWS^{+}13a, Xie14a, XS13, XM14, YYY^{+}14, YZ12, YN10, YN14, ZGSZ10, Zho14a, ZF14, vTCL^{+}10]. Boltzmann-Immersed [FRP14]. bond [NWLS12]. bone [AKL^{+}13, SR13b]. boom [AL10]. boosted [VGCMG11, YXD^{+}14]. BOR [TD11a]. Borehole [PMM^{+}11]. Boris [TAAY14]. Bose [AD14, BCW10, BCZ13, CCW11, CCC12, DH10, FT12, JCC14, LZZC14, MTZ14, ST13, Wan11b, Wan14, WC14a]. boson [BD11a]. bounce [YZ12]. bounce-back [YZ12]. bound [AM12, MB13b, VKL^{+}13]. boundaries [BWC13, BUH11, CN10, GPY14, Gri13, GT11, JKR14, MGLS12, MK12, MHHGN14, NS14b, SS10a, UGF14, YS11a, YS11b, ZN14a]. Boundary [ABH^{+}14, BB12a, BGH10, CF11a, CEK14, CTG12, CG12, FCYF11, FCY13, LMK13, MK12, SVK14, WY11, ABD14, AV10, AHA12, AMP14, AST13, ALS^{+}14, ADSF13, AKMD14, BF14, BN11, BN13, BN14a, BS10a, BR10b, BRS10, Bre12b, BHY14, CKS14, CCFCM11, CT10, CT13, CHC11, CFKL11, CB10, CLS10, CVI^{+}13, CGC^{+}12, CCG11, DDF14, DK14, DP12, Don14b, DKC14, Du11b, DGF14, DWZ10, EO11, ES13a, FDHP11, FL14, FY13, FCN^{+}13, hGzSnZ12, GCA13, GW11, Gen13, GK13, GHM14, GGH14, GS11b, GH10, HPA14a, HZ11a, HKM12, HG11, HLW13, HL12b, HDD14, HF11, HDPF11, HKL14, HCS11, HCS12, HW14a, HF10, IF10, JYHT10, JMW12, JC10, JCT11, JLC14, JM10, Kau10, KS11c, KSR^{+}14, KF12, KJ11, KT14, KLP10, KL10a, KLPS14, KTT13, KG12b, Le 14b, LRL10, LGP10, LKCY11, LY13a, LFK11]. boundary [LVFK14, Li10a, LKTL10, LL14a, LRK11, Liu11b, LX13b, LH14b, LL12b, LL14b, LYD10, Man12, MBS13, MMSI13, MBMV13, MZ14, MJ14, MST14, MP13, MBD13, MSTL14, MLBS14, NTV14, NMZC13, NMCZ14, NZ14, Nic11, NEE12, PSL20, PBT14, PGR10, PHRG13, PR11, PNPF10, PC13, PGA11, RVB10, RPV13, RzSZ13, RKGM14, RVbZ10, SG10b, SG10a, SG12a, SG12b, SHMS13, SM11b, SM11c, SSC13, SF12, SR11, SC11, SP11, SDH12, SN14, Tak14, TS10, TS11, TWSN12, TG14a, hT10b, TLZ^{+}11, TE11, TH14b, TA13, UGF14, Vai11, Vai14, VRB10, VM10, VW14b, WZ11, WZ10, WS10c, WYC13, Xu11, XMDG13, XMDG14, YYY^{+}14, YL11, YLW10, YCY11, YS12a, YHL13, YZ12, YW13, YHM12, ZT13a, ZF12, fZlC11, Zha10b, ZIOF10, ZzS11, ZW10b, ZLH12, ZF14, vDS12, FRP14, HGRB14, MDM^{+}12]. boundary-lattice [TLZ^{+}11, WS10c, ZF14]. boundary-layer [AKMD14, DWZ10, MMSI13]. boundary-phase [SSC13]. boundary-SAMR [KSR^{+}14]. boundary-thermal [HW14a, JYHT10]. boundary-value [HLW13, PR11]. boundary/level [CB10]. boundary/level-set [CB10]. bounded [BLQ14, CES14, EEK14, KH13, LSBJ14, VXB14]. Boussinesq [CW13, EEK14, KDS14, OBT12, PDZ^{+}14, RF14a, WZZZ14]. Boussinesq-type [EEK14, KDS14]. box [ZzS11, RP13]. box-type [ZzS11]. brain [GFK13]. branch [GBC11]. breakdown [CRS14]. breaking [IBO14, KDS14, SMS14]. breakthrough [CM13]. breakup [KBB13, SY12c]. breathing [YCH^{+}13]. bridging [HDPF11, LB13]. bright [BTX13]. Brinkman [CCLV10, MWY14]. brittle [AJV11, LYE10, SLC^{+}11]. broad [AF13]. broadband [AH13b, DG12]. Brownian [LKP14, PAG11]. Brunt [HdBK12]. Brusselator [LRBT14]. bubble [CP12, HRBK10, HY09, HY11, HHL14, MBGW13]. bubble-stabilized [HY09, HY11]. bubbles [APF13, KL11b]. Buckley [MJM13]. build [vTCL^{+}10]. bulk [BP10, CL14b]. Buneman [DST14, TAAY14]. buoyancy [EMS11]. Burgers’ [ZOWZ10]. burning [MRD13].

C [Pri14, Gas11, Gri13, RTKS10, WPKK12]. C-grid [Gas11, Gri13]. C-grids [RTKS10]. CAA [GYL12]. CABARET [SK13]. CAD [XMDG14, MPR12, XMDG13]. Cahn [HGW11, BSSW12, BSS12, BSB14, GLWW14, GGT13, GXX14, GLL14, HHK13, LDE^{+}13, RNT12, SJK11, WG11, ZW10a, ZTG^{+}13]. CaIrO [Ber14]. calculating [And13, ELM14, Tit12]. Calculation [MBS13, NVS13, IKO^{+}14, LH13, LLYE12a, MAPP13, MKB^{+}12, SS10c, Son11, SL13b, Xie14a, YTYJ12]. calculations [BHL12a, BG10, BV13, BHK^{+}10, CBPS13, CNSFD13, DR14, DSHP11, EH14, EH15, Fat10, GD10, GBB^{+}13, ITO^{+}10, JFC^{+}13, KM14b, LE13, LKTL10, NIN^{+}14, PA12, PA14, SV10, XS15]. calculator [VED13]. calculus [CT14, LCY11, RMMD14]. Calderon [KST13, NN12]. calibration [LMM14]. californica [Ham10]. Camassa [SCY11]. canal [vTCL^{+}10]. cancellation [SSS11]. canonical [HCS^{+}10, KSS14]. capability [AKMD14]. Capacitance [HP13]. capillaries [SHFB12]. capillary [MSB14]. capsule [FBM14, iT14]. capsules [HCS12, LT10a, LW11]. capture [DvW14]. capturing [BD10a, BCG^{+}13, Coo13, IST^{+}12, IXX14, JH12, LZZR10, Mar10b, OD13, RWX11, SPF10, Shu14, Xu11, YKWS13a, YKWS13b, ZJL^{+}14]. Caputo [hGzSwZ14]. cardiac [ABK12, ABK13, CYK^{+}14, KVC^{+}13, MNKS13]. cardinal [Boy11]. cardiovascular [BDM13, MVCFM13]. Carle [XX12]. Carlo [DWL10, Gho17, HH12b, HH13, AB13, AJG^{+}13, AKP^{+}12, BDL11, BHM11, BHM13, BP13a, Boo11, CRS14, CGP10, CBGK13, DD12, Den11b, DTU12, DBO^{+}10, EMSH14, GCA13, Gen11, GKN^{+}11, HJBB14, JNSA12, KSS14, LC13, LWZ^{+}13, MMM11, MSS^{+}14, MOV12, MSS12, MVG11, MJM13, MMJ14, NKSR12, NVS13, PBC^{+}11a, QA11, RLJ14, RSFS13, RRD^{+}14, SBC^{+}14, SJB11, Sch13a, SR13a, SD14a, SD14b, SSFM12, SSR^{+}13, SSK^{+}12, TSR13, WPP10, WK13b, WP14, ZKZ10, ZKKF10]. Carlo/deterministic [BDL11]. Cartesian [WYC13, GITW12, TD12a, BCL14, BCM^{+}11, CCY13, CR13, CCG11, DjY14, EQYF13, GIMT14, KWSJ11, KSU13, KSR^{+}14, LKTL10, MMR12, MBLM10, MTG11, MG14b, PHRG13, PMSH10, PB13b, SKU13, SY14a, SC11, UKP14, VBSK10, WY10, WY11, XFC^{+}14, ZN14a, ZN14b]. Cartesian-based [UKP14]. Cartesian-grid [WY11]. Cartesian-grids [WY10]. Cartesian/immersed [WYC13]. case [DL13b, FNGV14, FJ12, LX14, Pri14, RA10b, RS13b, RS14b, SMF^{+}13, TA13, YHT11, YT12]. cases [NL10, PE10, ZWL^{+}12b]. casting [CNR11]. catching [JXD14]. catching-up [JXD14]. Catmull [LW11]. Cattaneo [GMA10]. causal [AiINT14, TiI11, AAT^{+}14]. Causal-Path [AAT^{+}14]. cavitation [PS14a, PS15]. cavities [Du11b, MBMV13]. cavity [CLAL14, Du11a, DSZ13, LW13a, MAPP13, PPS12, RTS13, TRSMM^{+}13]. CC [Sen13]. CDG [LGLX14]. CDG-FE [LGLX14]. CE [CWL10]. CE/SE [CWL10]. Cell [AW14, HCS^{+}10, ADMA13, BILM14, CCB13, CY12, CCB11, CCB12, CC13, CCMG^{+}13, CCL^{+}14, CS10, CTJT13, CDLL12, CWZ^{+}11, DTG^{+}14, DOR10, DL12, DGAH10, DWZ10, FE11, Fuj11, GMB10, GHP13, GH12b, HT14, HLZ14, KGKS13, KD10, KHRD11, KHJ^{+}13, KSB14, Lap11, LY13a, Li12, MB12, MAB^{+}13, ML11a, MTB10, MT10, MKB^{+}13, PJCS14, SSB13, SHMS13, SF14, Son11, SD14c, VR14, WAM14, WZ10, YWY10, ZL10, DDN^{+}10, DC13, ILM^{+}13, MHL^{+}14, RCW11]. Cell-centered [AW14, CY12, CS10, CDLL12, DL12, FE11, GMB10, GH12b, KD10, MB12, MAB^{+}13, MTB10, MT10, MKB^{+}13, PJCS14, SSB13]. Cell-level [HCS^{+}10]. cell-local [WAM14]. cells [BHS14c, DFD14, EBGW13, Hel11a, KBSV14, Mac10, ZIOF10]. CENO [IG14]. center [LP12a, Men12, WP10]. center-difference [WP10]. centered [AW14, CY12, CS10, CDLL12, CG13b, DL12, FE11, GMB10, GH12b, HPOM10, KD10, MB12, MAB^{+}13, MTB10, MT10, MKB^{+}13, PJCS14, PLN11, SSB13]. centered-difference [HPOM10]. Central [LXY11, CLQX13, HWA10, IG14, LX12, LZZS13, RHXQ14, SIDG13, TK12, TK15, YR13, Zie11]. central-upwind [HWA10]. centred [AGBC14, AST12]. centres [DO11a]. certain [GPS10, Vai11]. CESE [BYCC14, QY12, RCQL11]. CFD [AQ11, EH15, XS15, ADT14, EH14, KS11a, KS11b, MT13, XDZ^{+}14]. CFD-DEM [ADT14]. CFL [CCGP13, TD11b, XCL14]. Chain [LC13, RMMD14, GW13a]. Chain-Rule [GW13a]. chains [KSS14, LHK12]. challenges [KYP^{+}14]. Chandrasekhar [Tan11]. change [AHOT14, DMHL14, HK13, HW14a, LYD10, SN13a]. changed [SKCC10]. changes [BLM11, NW10]. Channel [VT13, BPQO13, LTC14, OAK11, VKBK11]. channeling [SBS^{+}13]. channels [MGN14b]. chaos [AGS10, BS11b, GvdSVK10, KL14a, LBR14, PHD14, PGLK14, RWX11, RLPM12, TG14b, BW14, CSK13]. chaotic [Wan13b, WHB14]. characteristic [BDKV14, HKM12, LL14a, MBD13, PC13, SGL14]. characteristic-based [MBD13]. Characteristics [CBCM^{+}11, KS11a, Tal13]. Characterization [JAX11]. characterizations [BCZ13]. characterized [KG13a]. Characterizing [KIG11]. charge [CCB13, CCB11, CC13, DXB^{+}13, KHRD11, LHPH11, TAAY14]. charge- [CCB13, CC13]. charge-conservative [LHPH11]. charge-conserving [CCB11, KHRD11]. charged [LTC13, Zoh13]. Chebyshev [AV13, And10, BDLGC13, BY11, DAEB13, DBHBB13, HP13, HBZL14, LR11, LTSH10, LB14, MSD12, TDL10, Zho10, ZCS14]. Chebyshev-filtered [ZCS14]. checkerboards [Hel11b]. chemical [DR11, ERS14, JU10, KF13b, LRS10, Nag14a, Nag14b, NS14a, PPWZ11, Pud11, YR13]. chemically [KSR^{+}14, LI14, SRN10]. chemistry [MMSI13, RXLS14, YYB12, ZDSP11]. Cherenkov [GV14]. Choice [SF12, WC11]. Chow [ZZB14]. Christov [BX12]. CIP [FS10b]. circuits [Cam13, OBTC^{+}13]. circulant [LS13b]. Circular [YSW^{+}13]. circulation [BDTW10, MCJ14]. clamped [LHM14]. Clark [LW11]. class [ADB14, BJLR14, BDMV11, CCLS10, CP13b, FJ10, hGzS13, HDHZ13, JY11, LZZS13, Mar10b, NL10, PA14, QL11, RT14, SRL^{+}11, Tok11b, ZWL^{+}12a, ZWL^{+}12b, ZMD^{+}11]. classes [GPS10]. Classical [MTSG12, NVS13, RS13b, RS14b]. cleaning [TP12, YVL^{+}11]. clearance [Mit13]. climate [DMP14, JWNL11, OMCO14]. cloaking [LHY13]. cloaks [FGE11, LHY12]. Closest [MBR11]. closure [BHS14c, RB13, VTTK13, WABI11, TBMH11]. closure-independent [RB13]. closures [AHOT14, FCYF11, IAM^{+}12, MT13, Ols12]. cloud [RWMG11, Alm14]. CLSVOF [KBB13, Yok13]. cluster [GKG^{+}13, KEGM10, KO12, RJC13]. clustering [JJM13, Wit13]. clusters [BGK12, GHH14, LBRA^{+}13, Ran11, Ran12a]. co [CRG13, SLI10]. co-located [SLI10]. co-volume [CRG13]. CO5BOLD [FSL^{+}12]. coagulation [DRW11, WK13b, ZKZ10, ZKKF10]. coalescence [DLM^{+}13a, KBB13]. coalescence/breakup [KBB13]. Coarse [ALGB^{+}13, ZKKF10, BdCB11, Ich13, KKPV12, LG14, PJCS14, PA14, SS13a, ZZK^{+}14].

Coarse-grained [ALGB^{+}13, ZKKF10, LG14, PA14, ZZK^{+}14]. coarse-graining [Ich13]. coarse-grid [SS13a]. coarsening [NT14]. coastal [TQW14]. coated [KL11b]. cochlear [BS11a, SBWF14]. Code [BHL12b, BHB11, BGZ12, CDSP13, GLB^{+}11, IH13, JNW^{+}14, KSB14, Lap11, MA14, NHT^{+}10, NHT^{+}13, TGT^{+}10, TP14, TBNT11, YXD^{+}14, YS12b, ZKV10]. code-independent [IH13]. codes [CCMG^{+}13, GV14, HT14, IH13, KGKS13]. coefficient [DS12, HWW10, MGJZ10, Mar13, ZT10b]. coefficients [AGS10, CCG11, DV13a, GKN^{+}11, GP13, HPV13, OK14, PBC^{+}11a, SKCC10, XZWW12]. coexistence [LC13]. Coherent [MVG^{+}12, AMDR10, NRS10, YL14a, RS10b]. cold [TD12a]. Collaborating [XDZ^{+}14]. collective [LJ13a]. Collision [HH12b, HH13, HCS^{+}10, KM14b, KWH10, KS13b, LH13, LHK12, MZ14, PG11, WMK13]. collisional [HT14, LWF11]. collisionless [AHS14]. collisions [AiINT14, BP13a, CDS13, DCC^{+}13, GH14, HT10a, IAM^{+}12, LE13, RRCD14, RRD^{+}14, SSX13]. collocated [RP13, TLO^{+}14]. collocation [AA13, BdF10, CN12, CCC12, DYYA14, DBAV14, Dri10, FKM11, FF12, FK10, GRS14, HP13, JNX13, JCC14, Lee12b, LTSH10, LTT10, Mar13, NHE14, RVFK14, RMMD14, SAM10, ST13, SPM^{+}13, WZZ14, WI13b, WI13a, WTYC13, YZX14, ZN10, ZTS^{+}10]. colloidal [Kea14, LZC^{+}13]. Combination [DLLW14, HPS13, LP13a]. combinations [VDD^{+}13]. combine [BJLR14]. combined [ABD14, CSK13, DB12, DF10a, LZZC14, LLP14]. Combining [DDL13, KO14, CHKT13, HSN13, LSM14]. combusting [GG10]. combustion [BHB11, DLM^{+}13a, DKR12, Ger12, ISZ12, LI14]. Comment [Dad12, PSL20, XS15, ZJS15, dv23, AK10, EH15, RDHK12]. Comments [Gho17, HSK^{+}15]. commercial [JWNL11]. commitment [KH11]. common [Opp13]. communication [CRS14, NNS11, Nor14, OBT11, SSFM12]. communication-avoiding [Nor14]. commutator [AF11]. commutator-free [AF11]. Compact [Cui12, FKM11, HCD11, HM10, LN10c, MDTC11, Mur10, PJCS14, RzSZ13, TGGT13, WV14, BSS14, Boe11, DEH11, hGzS11, Ge10, GHM12, GCL13, GXS14, HE14, Kim13, LT12, LJ13b, LZZS13, MDTC13, NMT^{+}12, PDSS10, RBS12, SBS11, SZ11, SCY11, SY14b, TY11a, Tyl14, VW14b, WGX13, Wen10, YS13, ZFH14]. compact-WENO [GXS14]. compactly [Vai11]. companion [Boy13]. comparative [JHDC10, KGSS10]. Comparing [BY11]. Comparison [BHM11, BBC^{+}13, BM13a, CHHS11, DSHP11, LLX11a, MCDT12, MBGW13, iT14, BTX13, Boy13, GNGAS10, LKT^{+}12, LCP13, MOSW12, Pri14, PR14b, Rie10, SMP^{+}13, TDU14, WPK14, ZTG^{+}13, vRLPK11]. Comparisons [GDY10, BMD^{+}13, CB13, HL12a, JNW^{+}14, SC10]. Compatible [KBS14, CS12a, FGN14, KAB^{+}14, MK13, OS14, PRH^{+}14, TF10, dCRCS10]. Complement [Rod12, FMT11b, LYD10, SR13d]. complete [LHH14]. completely [NW12, YD11]. Complex [Lip12, RVbZ10, STD14, ADSF13, ABH^{+}14, BH13a, BUH11, BP13b, CKS14, CB10, CZ11b, CLS10, CGC^{+}12, DBO^{+}10, Guy13, HS13a, JTT14, JK11, KWS^{+}11, KIG11, LH13, LL14a, LN10c, LGE^{+}13, LH14b, MGN11, NJ12, PBdGP14, PHSA14, RF14a, SHCY13, SSM14, SM11b, SZ11, SJK11, SCC14, SY11c, SHZF12, UKP14, VBVD10, VS11, WYYK10, XDZ^{+}14, YZ12, ZN14a]. Complex-plane [STD14]. complexity [Pie12, PS14b]. complicated [CLZ12, Mar10a]. component [BD10b, HK11, MZ11a, MTZ14, PvDGvM11, PvDtTB^{+}11, ZKZ10]. components [ASK^{+}13, FK14, LRT13, PBC^{+}14, SBT14, dv23]. composite [BHS12, Du11a, JXD14, Mar13, TRL14]. composition [JWNL11, LBR14, WPP10]. composition-resolved [JWNL11]. Compositional [PM14, DFVY14, MF10, MF13]. comprehensive [LRK11]. compressibility [AOCR12, Deg11, OAKR14, OA10, PPTMDK13, TM10]. Compressible [JSA14, AB11, BHS13, BOD11, BDKV14, CVN13, CP10, CS10, CS14a, CS14b, CRT10, CC14, CGC^{+}12, CGF^{+}13, Del11, Del10, DL12, DL10b, FG12, FL11a, GITW12, GIMT14, GVV13, GKF11, GF13, GCL13, GDGP11, GXS14, HLT10a, HHA14, HMHO13, HK11, HK13, HAS13, JKW11, JLB^{+}10, JH12, KKP10, KSL10, KL10b, KG14, KK10c, KGK13, LMFH11, LGC13, LLX11a, LZ13, LLN^{+}10, LXS^{+}13, MWP14, MS10a, MF14, MESV10, MDM^{+}12, MJ13b, NMT^{+}12, NKF14, OL13, PZW14, PZG14, PC13, PL12, PR14b, QLW14, RGMC13, dCRCS10, RS10b, SM13, SZ10, SSHM10, SS10c, SPF10, Shu14, Shy10, SX14, SKW14, SHA12, SRL^{+}11, TS11, TBMH11, TT10, UKP14, VL14, WFK11, WWX^{+}10, WS10b, WP10, XMS12, XS13, XW14, YCLK12, YSW^{+}13, YSW14, ZHW10, ZS10b, ZS11, ZS12, dPSS12, vV12]. compression [HdCNT12, JU10, Opp13, OK10, RJG10]. Compressive [DvW14, ZDSP11]. Compton [DWL10]. Compton-scattering [DWL10]. Comput [ASS17, CFCA13a, Dav15, Gho17, HY11, LWD13a, NHT^{+}13, OBNN14, PS15, SYV17, SS14b, TK15, XMDG14, YKWS13a, ZJS15]. computable [AM12]. Computation [BS14a, CCLV10, DTYY14, GCLS11, LJ13b, MPS^{+}14, MRS14b, RW13, AK10, BM13a, CJLC11, CB10, DV13a, DKW10, ES10, EAB^{+}12, ES14, FN10, Fuj11, GT12, Gen13, GHM12, GP13, Gür14, KED^{+}12, LRS10, LL12a, LH10d, LHV13, MTD14, MLGN12, PBdGP14, PM14, RS13b, RS14b, Tan11, TVEC12, Wan13b, WC14b, XW14, ZD11a, ZKKF10]. Computational [ASXZ14, AS12, APRS14, Dad12, EH15, GHS13, HSK^{+}15, KEG^{+}12, KYP^{+}14, Try12, XS15, BSV14, CFCA13a, CFCA13b, CCB13, CP10, CGS13, CR14, EBGW13, Fal13, HHGC14, LCP13, OD14, OK13, PMS11, PSM11, PSM13, PMS14, PT12, SAHS14, SR13b, SC13, SS13c, Ton14, XHF13, XS11, XMDG13, XMDG14, LLH10]. Computationally [WG11, GL13, MJ13a]. computationally-efficient [GL13]. computations [AST13, BK12b, BX12, CRK12, CG13a, GP12b, GGG^{+}13, HLLW11, KS13a, LEH10, LEH12, LWD13a, LWD13b, LvZB10, OS11b, PB11, RWX11, SS13a, SHCY13, SLT^{+}14, SFWP10, SCB10, YS11b, ZS14]. compute [CP14, GD14, LFW11, STD14]. computed [LTC13]. Computer [HVT12]. computers [HOK10]. Computing [Ber14, Gil13, LSM14, LD12, WJC13, AD14, AC11, BCW10, BD11a, BCZ13, BTX13, BM10a, CDSP13, CF11b, CC12, CH14, DH10, Don11, FN14, FT12, GSTB13, GMG12, HF11, KST13, Kol11, LKP14, Leu11, LLX11b, xLqW13, LTC14, MWP14, MP13, MTZ14, NHE14, PMMB14, SSB13, Wan11b, Wan14, WC14a, YL14a]. concentrated [YM10]. concept [BF13, ZPGO10]. concepts [LRS10]. condensates [AD14, BCW10, BCZ13, CCW11, DH10, FT12, JCC14, MTZ14, ST13, Wan11b, Wan14, WC14a]. condensation [CCC12, HSN13, LZZC14]. condensed [MFM12, SNBN13]. condensed-phase [SNBN13]. condition [AV10, AHA12, Don14b, DKC14, GS11b, HZ11a, JCT11, JLC14, Lee12a, MZ14, PBT14, RPV13, SSC13, SF12, TDL10, UGF14, VM10, Xu11, XCL14]. Conditional [YF11, Oli14, TBMH11]. conditioned [CBPS13, WZZ14]. Conditions [TD11b, AMP14, ALS^{+}14, BDT10, BB13, BN11, BN13, BN14a, BHY14, CCGP13, CT10, CVI^{+}13, CG12, CGC^{+}12, DK14, Du11b, EC13, FY13, FCN^{+}13, hGzSnZ12, GCA13, GHM13, HPA14a, HKM12, HG11, HF11, HDPF11, IF10, JM10, KSS14, KMHJ10, LDS11, LFK11, LVFK14, LMK13, LKTL10, LL14a, LMS11, Liu11b, MNR11, MJ14, MLBS14, Nic11, PGR10, PHRG13, PBB14, PL10, PRL10, PC13, PGA11, RzSZ13, SG10b, SG10a, SG12a, SG12b, She13, SR11, Stu10, SP11, TS10, hT10b, TE11, Vai11, Vai14, VW14b, YZ12, ZT13a, ZzS11, ZW10b, BGH10]. conductance [SHFB12]. conducting [KSR^{+}14]. conduction [BD12, BDLGC13, Coo13, FAY^{+}13, KK10a, LCK11, MM11, MSTL13]. conduction-radiation [MSTL13]. conductivities [APRS14]. conductivity [BLM11, Hel11a, Hel11b]. conductors [Liu11a]. cone [HPA14b]. configurations [KSB12, RARO13, VBVD10]. confined [Jar12, LVZK14, MLBS14]. Confinement [WCM^{+}14, TTR^{+}12]. Conformal [MNS13, BY11, CWZ^{+}11, DZL13, PL10]. conforming [CRG13, EH13, KM13, KSR^{+}14, NKS10]. congestion [DHN11, DH13]. conical [HZ11a]. conjugate [KKvZB14, LN10b, NTV14, SD14c, Vol14, ZVNM10, MMM13]. connected [LM14b]. Connecting [PBB14]. connections [BCLR10]. Connectivity [WWZ13, SFWP10]. Connectivity-free [WWZ13]. conquer [BRP14]. consensus [YWS13]. Conservation [DMM^{+}13, AGBC14, BMS12, Bal12b, Bal14, BK14b, CCD11b, CCK13, CS10, CB10, CDL11, DMT^{+}11, DF10b, DZLD14, EFT13, Fan14, FD11, FCN^{+}13, FC13, GAV13, GPP11, HHS^{+}13, HTHG14, HAH14, IDNG13, JS13a, KPH13, Kri10, KL12c, LJX^{+}14, LRLL11, LR12, MN11, MSS12, MNS13, MLM12, MV13, NF12, Par12a, PYK10, QL11, QL12, RSS13, RTKS10, SK13, SM11c, SABH11, SHZ13, SLI10, TS10, TWSN12, hT10b, UGkM13, XL11, XLD^{+}11, XCL14, YHL13, ZS10a, ZWL^{+}12a, ZWL^{+}12b, ZLCW14]. conservation-flux [SK13]. Conservative [AINF13, CMS10b, JM12, KS14, QS11a, Ros10, YF14, ANIF14, AK10, AHA10a, AH10a, BOD11, BBG^{+}11, BRSS11, BRS13, BD14b, CDT13, CL14b, CS14b, CL11b, DB13, Era13, EN14a, EN14b, EAN11, FCN^{+}13, GH14, GW13a, GF13, HLM11, HSK^{+}15, HAS13, KR10, KAFB11, LSM14, LNU10, LP13b, LGF11, LSVW10, LHPH11, MKCB14, MD13, MD14, MDH^{+}10, MTB10, Mig14, MG14b, MDM^{+}12, Mor10, MK14b, NL12b, NS14b, OAK11, OS11b, OD13, OD14, PvDtTB^{+}11, Pir10, QC10, SN12, SN13a, SJK11, SS10d, TF10, TWMM14, WYS12b, WGX13, WXY14, Zad11, ZSW10, ZHY13, ZN14a, WY10]. conserved [LY13b, ZM10]. conserving [CCB13, CCB11, CC13, CCZ14a, CCZ14b, CSX14, HHL14, KBS14, KJ11, KHRD11, LTL^{+}14, MRS14a, ML11a, OS14, San13, TS12, Zer10, vV12]. considerations [CVIA10, LB10]. consistency [BH13a, FQV13, HLSO13, HZ14, PP14]. consistent [BNM14, BN13, BN14a, CDT13, CdCD12, DR11, GK14, KKLR10, LGE^{+}13, NL12b, OS14, PS14a, PS15, SKW14, XJ12, ZN14a]. consolidation [FCG10]. constant [BMK11, CWC11, DS12, LP12b, PM14, fZlC11, ZWL10]. constants [Ber14]. constitutive [LXM14, VP10]. Constrained [BHS14c, TP12, ZD12, CLSX14, CRT14, HMM14, HMS10, HRT11, HA14, IX10, KS13b, KBSV14, KMSS10, SAM10]. constraint [ACCM12, DHN11, McD14b, NMV13, PvDGvM11, WSS12, XCL14]. constraints [BH13a, BSS12, Kau10, MM12, ZD12]. constructed [WYS12a]. Constructing [XMDG13, XMDG14, SW11, ZF12]. Construction [ALR11, TDU14, WYS^{+}11, Cou11, KSR^{+}14, Kol11, LLY11, MF12, MW12, SRN10, Vai14, WZ14]. contact [BSSW12, DL10c, GW12, GW14, Han11, HGW11, JGZL14, Joh11, MKB^{+}13, RB14, SN12, SSC13, SYY14a, SYY14b, SSB^{+}12, SS13c, Vol14, XR14]. containing [YKWS13a, YKWS13b]. contamination [DL10b]. content [Try12]. Contents [Ano10a, Ano11a, Ano11b, Ano12a, Ano12x, Ano13g, Ano13h, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13t, Ano13-41, Ano13-42, Ano14a, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano10z, Ano10-27, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano10-32, Ano10-33, Ano10-34, Ano10-35, Ano10-36, Ano10-37, Ano10-38, Ano10-39, Ano10-40, Ano10-41, Ano10-42, Ano10-43, Ano10-44, Ano10-45, Ano10-46, Ano10-47, Ano10-48, Ano10-49, Ano10-50, Ano10-51, Ano10-52, Ano10-53, Ano10-54, Ano10-55, Ano10-56, Ano10-57, Ano10-58, Ano10-59, Ano10-60, Ano10-61, Ano10-62, Ano10-63]. contents [Ano10-64, Ano10-65, Ano10-66, Ano10-67, Ano10-68, Ano10-69, Ano10-70, Ano10-71, Ano10-72, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-33, Ano11-34, Ano11-35, Ano11-36, Ano11-37, Ano11-38, Ano11-39, Ano11-40, Ano11-41, Ano11-42, Ano11-43, Ano11-44, Ano11-45, Ano11-46, Ano11-47, Ano11-48, Ano11-49, Ano11-50, Ano11-51, Ano11-52, Ano11-53, Ano11-54, Ano11-55, Ano11-56, Ano11-57, Ano11-58, Ano11-59, Ano11-60, Ano11-61, Ano11-62, Ano11-63, Ano11-64, Ano11-65, Ano11-66, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano12-36, Ano12-37, Ano12-38, Ano12-39, Ano12-40, Ano12-41, Ano12-42, Ano12-43, Ano12-44, Ano12-45, Ano12-46, Ano12-47, Ano12-48, Ano12-49, Ano12-50, Ano12-51, Ano12-52, Ano12-53, Ano12-54, Ano12-55]. contents [Ano12-56, Ano12-57, Ano12-58, Ano12-59, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13u, Ano13v, Ano13w, Ano13x, Ano13y, Ano13z, Ano13-27, Ano13-28, Ano13-29, Ano13-30, Ano13-33, Ano13-34, Ano13-35, Ano13-36, Ano13-37, Ano13-38, Ano13-39, Ano13-40, Ano13-43, Ano13-44, Ano13-45, Ano13-46, Ano13-47, Ano13-48, Ano13-49, Ano13-50, Ano13-51, Ano13-52]. continuation [CF10, CCW11, HL12b, JCC14, SHZ13, ST13, WC14a, WIJ13]. continuation-multigrid [CF10]. continuation-WENO [SHZ13]. continuation/WENO [SABH11]. Continued [Ano13g, Ano13h, Ano13t, Ano10a, Ano10z, Ano10-27, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano10-32, Ano10-33, Ano10-34, Ano10-35, Ano10-36, Ano10-37, Ano10-38, Ano10-39, Ano10-40, Ano10-41, Ano10-42, Ano10-43, Ano10-44, Ano10-45, Ano10-46, Ano10-47, Ano10-48, Ano11a, Ano11b, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-33, Ano11-34, Ano11-35, Ano11-36, Ano11-37, Ano11-38, Ano11-39, Ano11-40, Ano11-41, Ano11-42, Ano12a, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13u, Ano13v, Ano13w, Ano13x, Ano13y, Ano13z, Ano13-27, Ano13-28, Ano13-29, Ano13-30]. continuity [BDM13, CXLF14, PvDtTB^{+}11]. Continuous [KG12a, BD11b, CEP12, CdCD12, DTYY14, DKA12, GAvdVB14, IST^{+}12, IXX14, PBT14, RE13, YL14a]. continuous/discontinuous [CdCD12]. Continuum [Mit13, FCH13, Fuj13, KMHJ10, LE13, NP12, PLBR13, SBA13, iT14, VCS14, XH10, XYL12, Yok14, ZJT10]. Continuum-kinetic-microscopic [Mit13]. continuum-molecular [PLBR13]. continuum-particle [SBA13]. contour [SM12a, SM12b]. contour-advective [SM12a, SM12b]. contrails [JWNL11]. contrast [CEL14, EGW11, EGG12, GPCE13, GCE14, GK14, Hel11a]. control [BVM14, BS14a, CEK14, DLLW14, DBM^{+}12, HA14, LY13a, MRS14b, SW13, VS14, fZlC11, ZJW12]. controllable [SRL^{+}11, sSLxRyZ14]. controlled [CLW10, KLB^{+}14, NMV13, SCS14, WH11]. Controlling [CN10, TE14]. convected [CES14, GCH14]. Convecting [BN14b, BPQO13]. convection [AH10a, BKC^{+}11, BD11b, CCD11a, CL14b, DMHL14, EBGW13, FSL^{+}12, GNGAS10, GBNS14, GC11, GMXG14, HY10, HW14b, JN12, KTT12, KHHK12, LK14, MMV^{+}13a, MHA10, MK14b, Ste14, TCW^{+}10, YN10, YN14, ZFH14, ZZS13, ZQ11]. convection-diffusion [AH10a, BKC^{+}11, CL14b, HY10, HW14b, JN12, LK14, MHA10, YN14, ZFH14, ZZS13, ZQ11]. convection-dominated [BD11b]. convective [BGH10, CVC12, FL11b, GDY10, KFOS14, Mor10, Pir10]. Convergence [CCY13, GBS^{+}14, BK12a, BMCK10, BV13, CSK13, GAV13, LTS^{+}10, LJ13a, NEE12, RBNS^{+}12, RBNS13, SPM^{+}13, WZZ14, WTK14]. convergent [BCP13, BHM14, BD14c, STD14]. converging [LTL11, Sto13]. Convex [VDD^{+}13, BLQ14, DFD14, GLWW14, SJ10, SBO11, hT10b]. convolution [BK14a, HW10]. convolutionary [ZD10]. cooling [ASGW11, MSTL14]. coordinate [MTZ14, OK10, YN14]. coordinates [AV10, BLA^{+}14, CS14b, CDHM11, FS10b, HE14, JMGN13, KBS14, Mig14, PG11, PPS12, Pir11, QZM^{+}12, SCH13b, VB13, YJK12]. copolymers [JHZ10]. core [CCD11a, Mit13, UJ12]. core-annular [Mit13]. corner [MAPP13]. corners [BRS10, Bre12a, GHM14]. correct [OJ11]. corrected [LBM10, LSVW10, OS11b, OJ11, VKL^{+}13, ZS14]. correction [BNM14, BD11b, CGMQ14, CSC14, DvW14, DF14, DS10, GW13a, IF10, LCP13, LMZ14, Min13, Par10, SRS10, SHFB12, SR13d, WW13, YWL14, MNR11]. corrector [CP13b]. correlated [CP13a]. correlations [VLM13]. Corrigendum [ASS17, CFCA13a, Dav15, MPPP12b, NHT^{+}13, OBNN14, PS15, SYV17, SS14b, TK15, XMDG14, YKWS13a]. corrosion [BBCH^{+}12]. cortical [AKL^{+}13]. cosmesis [GSTB13]. cost [GHS13]. costs [SSFM12]. Couette [BA10, GDW11, VT13]. Coulomb [BP13a, CDS13, DCC^{+}13, HT10a, Men12, PG11, RRCD14, RRD^{+}14]. counting [WLPG11]. Coupled [CES14, CHKT13, AMP14, BMP14, BSSW12, BGGW13, CWL13, CKS14, CP10, CS12b, CNST13, CL14b, CSW14, CH11a, FAY^{+}13, FCG10, GMG12, HVD13, HHK13, HY10, HCM10, KBTD12, KKvZB14, KSR^{+}14, KMHJ10, LFLV14, LZZR10, Min10b, MMH12, MGN10a, NKS10, RM14, SNBN13, SHCY13, SPB^{+}10, TWMM14, WK10, WXY13, WXY14, WX14, WSBG10, Zha14]. Coupling [AJRT11, AV10, AP12b, Li10a, MD13, MHA10, WKL^{+}14, BC10a, BLR13a, BK14b, BP10, CL10, DTG^{+}14, DSHP11, EGL^{+}13, EHA14, Her10, KR10, KN14, Kea14, KKC^{+}13, LP13a, LYE10, LDDBR13, MVCFM13, MKB^{+}12, MDM^{+}12, SKCC10, SCC14, SIT^{+}11, UPDB13, YM10, ZKM^{+}11, vTCL^{+}10, ABH^{+}14, YS11b]. covariance [GCLS11]. covariances [GLS10]. Covolume [JTXZ12]. Covolume-upwind [JTXZ12]. CPU [CCB12, DKH11, Pri14, XDZ^{+}14]. crack [SLC^{+}11]. cracks [LYE10]. Crank [ÇD12b, FF12, WXY13, YXZ13]. creating [SJ10]. criteria [DFVY14]. criterion [WFL10]. critical [MMM11, ZCD10]. Cross [JNW^{+}14, LH13, LRBT14, MBMV13, MF13]. Cross-code [JNW^{+}14]. cross-diffusion [LRBT14]. cross-flow [MF13]. cross-section [MBMV13]. crossing [SM12b]. crossings [HZ11a, SM12a, SM12b]. crowdedness [JJM13]. crowds [DH13]. crystal [BGGGS11, BHL^{+}13, BSS11, CTSM13, HCD11, LGH10b, YFL^{+}13, ZMQ13]. crystalline [STC10b]. crystals [CVI^{+}13, HCH^{+}10, KA13, MNF^{+}10, SBS^{+}13]. CSEM [BPM13]. CSF [Yok13]. CSLAM [HLM11, LNU10]. CTU [MT10]. cubed [BNT14, HLM11, IDNG13, LNU10, YC11]. cubed-sphere [BNT14, IDNG13, LNU10, YC11]. cubic [Del14, PLN11, Vai11, ZCD10]. CUDA [Pri14, HJBB14, JLS^{+}14]. current [BBF12, BM13a, Del13, JC10, KHRD11, SCB10]. current-based [SCB10]. current-dependent [Del13]. currents [BS14a]. curvature [ELM14, FS10a, Gür14, IKO^{+}14, LRS10, LH10d, OLC13, TVEC12, WJCG13]. curvatures [FLM12]. curve [Par12b, ZWJW10]. curve-like [Par12b]. curved [BWC11, BWC13, CDLL12, CKR14, Gen13, MBR11, NZ14, UGF14, VM10, VQVD11, Zha10b]. curves [Bre12a, MWP14]. Curvilinear [PDSS10, AV10, DKR14, FS10b, HE14, JRG13, NTV14, Pir11, SML13, TGRL13, YN14, YJK12, Zie11]. CUSP [SZH12]. cut [DWZ10, SHMS13, SD14c]. cut-cell [DWZ10, SHMS13, SD14c]. cutoff [LHV13]. CWENO [HAH14]. cycle [KWHD12, MNKS13, WHB14]. cycles [WIJ13]. cyclic [HPLS10]. cylinders [Lee12b, SSS10, STC10a]. Cylindrical [KBS14, MK13, BLA^{+}14, BdCB11, BY11, CS10, CS14b, DC13, GDGP11, JWV12, JNW^{+}14, KSB12, LHY12, LE12, Mig14, PG11, PPS12, RSK13, SSB13]. cylindrically [MTD14].

D [CGF^{+}13, Del11, DCL11, HLYS11, JM10, MMM11, MVCFM13, YCY11, ZOWZ10, ZJS15, KKS13, SYE10, AST13, ABMT14, AW14, AH10b, BGR11, BLM^{+}14, BWC11, BS14a, BBAP14, BGHM13, BCL13, BCL14, BLJ11, BD14b, BSS11, BHM14, BLY10, Cha14, CP14, CCW12, CL11a, CC12, CS14c, CLAL14, CH14, DDL13, DK14, DFD14, FBM14, FNGV14, FW12, FCG10, FGLB14, GW13b, Ge10, GC11, GCZ13, GM10a, GHM14, GD10, GL12a, GLC^{+}11, GKS^{+}11, HPA14a, HBM13, HRT11, HLW13, HvB11, Hu13, JMGN13, KL12a, KVM12, KBS14, LH10a, LP13a, LVFK14, LTSH10, LT12, LEH12, LLP14, LMBB11, LS13d, LDA10, LL10b, LZZR10, MAPP13, MMR12, MDTC13, MBMV13, MVCFM13, MZB12, MMT14, MF13, MHGNM12, MHHGN14, MV13, MBGW13, MT14b, MLGN12, MGN14b, Nic10, NJ12, Ols12, PPS12, PWB^{+}14, PBdGP14]. D [RCQL11, RKGM14, SKN13a, SY12a, SY14a, SCGE13, SF14, SCC14, SCH13b, SM12c, SAA14, TGT^{+}10, TQW14, TDL^{+}14, TRSMM^{+}13, TR14, TGGT13, UJ11, UWB12, WYYK10, WR11, WG11, Xu11, Zha10a, ZYF^{+}10, ZMD^{+}11, vTCL^{+}10]. D-lattice [BBAP14]. D-TLM [TRSMM^{+}13]. D-Var [CS14c]. D/ [MVCFM13]. DAE [MM12]. Damage [AJV11]. damped [MBMV13, MNS13]. damping [BK12a, DGAH10]. Darcy [DFVY14, EZ10, HPV13, KR10, LLMY14, NJ12, SW10, WX14]. Darcy-flux [EZ10]. dark [BTX13]. Data [AC11, BND^{+}12, QA11, RSFS13, AR14, BERB^{+}12, BGG12, CS14c, CF11c, GLS10, GCLS11, GHM13, HHA14, HMM^{+}10, HVT12, IK11, JLY10, JSX10, KS11a, Kaz10, KMSS10, NMX12, SS10a, Vai11, WMI10, WY11]. data-flow [BERB^{+}12]. Data-free [BND^{+}12]. datasets [AA13, HdCNT12]. Davidson [And13, HWHW10, Zho10]. DDM [DL10b, LD14]. DDR [JS13a]. Deborah [Kwo14]. debris [MGN12b]. Debye [NZ14]. decay [MS14b]. decomposed [SSFM12]. decomposition [AST13, AS12, AETT11, APY14, BPM13, BAG12, CBKK12, DLNN12, DMR14, DEKBF14, DAA11, ETW10, GPCE13, GCE14, GK10, Gri12b, GD13, HdCNT12, JC13, JWYH10, LTSH10, LTT10, LDE^{+}13, PL10, PRL10, RHB12, RHGT10, RSFS13, SYL12, SSB^{+}12, SCH13b, SKN13b, Sto13, Stu10, SS14c, VHG13, WABI11, WC14b, Xie14a, YCLK12]. decomposition-based [JWYH10]. deconvolution [KRF12]. Decoupled [OD12]. decoupling [NDV^{+}11]. Defect [BNM14, Coa12]. defects [Del14]. deferred [CGMQ14]. defined [GM10b, Sza12, YW13, YHM12]. definite [GM12a, RMSF11]. definition [MCLG10]. deflated [MMM13]. deflation [Vol14]. deformable [KF13a, MGN14a, SL13b, WKL^{+}14, XHF13]. deformation [BWS10, LT10a, LCB12, RA10a, ZL13]. deformational [NL10]. deformations [BBD11, LMQ^{+}11, NW10, TDL^{+}14]. deformed [XZWW12]. Deforming [BHS12, SYV17, AINF13, ANIF14, BBGP13, GCvRK11, LMZ14, MN11, RC12, SYV14]. Degasperis [YS13]. degenerate [CGC13, TSR13]. degree [KL11a]. DEIM [SN13b, XFB^{+}14]. delivery [RUL^{+}13]. Delta [ZT10a]. DEM-SPH [SSY13]. demagnetizing [GHDS10]. DEMOCRITUS [Lap11]. dendrite [GMG12, GMXG14]. dendrites [KBTD12]. dendritic [BGN10, RR14]. dense [CM14, GLLP14, JNW^{+}14, VP14]. Density [LE12, AHS14, AGSG11, AF13, ABPSM11, BGK12, BW14, BBF12, CCKCG10, DS12, FGR12, GW14, Gil13, GW10, GK14, GLL14, ITO^{+}10, JYHT10, JFC^{+}13, KSS14, KG13a, KL12b, LP13b, LRT13, LMGS13, LLYE12a, LLYE12b, LSE13, LHB10, Mor10, MIKG12, MNL^{+}13, MLGN12, Par10, SL13a, SD14b, SHM12, SBO11, SP14b, TBNT11, VK12, Yok13, Yok14, ZWL10, dv23, vydCNS^{+}10, Fat10]. density-driven [KL12b]. density-scaled [Yok13, Yok14]. dependencies [RE13]. dependent [AME14, BN12, Brü13, CdCNH14, CEK14, CHZ13b, CHZ13a, Del13, DTU12, Dur14, FGLB14, FWP13, FWLP14, Gen11, GvdSVK10, Gho17, GH12a, GS11b, HPA14b, HBM13, HBAP10, LFK11, LTL^{+}14, LSW14, MM10, OLFS13, RA10b, SG10b, SD14a, SLGB12, SSR14, SW13, Tal13, WD13a, WC11, ZD10]. deposition [KHRD11, Zoh13, vDS12]. depth [BH13a]. derefinement [BLOdV14]. Derivation [CHZ13a, SVS11, OLFS13]. derivative [CNR11, ÇD12b, CJLC11, DBZ14, hGzSwZ14, JR12, Kol11, Par12b, Pir10, SW11, SAM10, SBS11, WXY13, YMS12]. derivative-free [SAM10]. derivatives [BCL13, CTJT13, Gon12, HLY12, KSS14, Kaz10, LA10, Mar10a, Mat14, MW14b, ZY13]. derive [KBRM12]. deriving [LGE^{+}13]. descent [SBT14, WJC13]. describing [ÁCB12]. description [Jar12, MGN12b]. Descriptors [KIG11]. design [ASGW11, BF13, DRB14, EAN11, HM13, KJC13, LA10, LGC13, SRN13, WFL10, YDN10, ZLL10]. designed [OK10]. detailed [ZDSP11]. Detecting [Ran12a, DEH11, KL14b, Ran12b]. Detection [CDL11, GBC11, KS11a, LLSW14, MCLG10, SM12a, VR14]. deteriorating [PT12]. determination [PBC^{+}11a, XMS12, YLW10]. Determining [KWJR14, KJWR14]. Deterministic [AJ14, WWS^{+}13a, AC13, BDL11, BM14, Dav10, Dav15, GLC^{+}11]. detonation [JLY10]. detonations [WZSN12]. developed [SNS10]. Developing [LHY12, LS10b]. Development [FMT11b, HMOH11, KIF^{+}14, SSC13, SML13, SLH^{+}11, GHS13, SCY11]. developments [CMO14]. device [JLC14, LS10a, MJ14, SO13, SRN13]. devices [BS14a, CBKK12, CW10, CGR11, PAG11, TK14]. devising [EEK14]. DFLUX [GMC13]. DG [CLQX13, CKR14, HLY11, JC10, JXD14, RGK12, YK12, ZWL^{+}12a, ZWL^{+}12b, ZQ11, vDS12]. DG/FV [ZWL^{+}12a, ZWL^{+}12b]. DGFEM [BD10a]. DGM [FH11]. DGSEM [WK14]. Diagonal [Mat14, MAC14]. Diagonal-norm [Mat14, MAC14]. diagonalisable [BLTO11]. diagonalization [BHK^{+}10, CBPS13, ZCS14]. diagonally [NMCZ14]. diagonally-implicit [NMCZ14]. diatomic [LYXZ14]. diblock [JHZ10]. Dichotomy [FT11]. dielectric [CPJ11, LTC14, PVDW14, TK14, YD11]. dielectric/electrolyte [LTC14]. dielectrics [BWC11, Liu11a, WBC13]. dielectrophoresis [HDD14]. Difference [BMS12, MOSW12, SYV17, AINF13, ANIF14, AV10, AN13, BDMP13, BHL^{+}13, BN13, BN14a, BSS14, Boe11, BHM14, BHY14, CCW12, CH10, CRT14, CWZ^{+}11, CGP10, CR13, CVI^{+}13, CR14, DEH11, DMT^{+}11, DMM^{+}13, DB13, DA14, DV14, EO11, EAN11, FH11, FCYF11, FCN^{+}13, FC13, FL11b, FO11, FSL10, hGzS11, hGzSnZ12, GYL12, Gao13, hGzS13, Ge10, GMA10, GM10b, GHM12, Gri13, GZZ14, GQ13, HPA14a, HPA14b, HE14, Hic12, HZ14, HPOM10, HJK14, HY10, JM10, Kim13, Kol11, LT12, LX14, LCP13, LMS11, LMBB11, LMS14, LC14, LS13d, LW14, LBS10, Mat14, MTT13, MTB10, Mor10, MK14b, OBT11, Par13, PGLT10, PNPF10, Pri14, QS11a, RzSZ13, SRSV11, SBS11, SQC11, SZ11, SYV14, Son11, SC12, SIT^{+}11, SRL^{+}11, sSLxRyZ14, TH14a, TY11a, Tsu10, Tyl14, VW14b, WWS10, WW11, WSYS12, WXY13, WD13a]. difference [WGX13, WV14, WD14, WXY14, WP10, WAM14, XQX13, XQXC14, YS13, YWL14, ZS12, ZY13, nZzSlL14, ZW10b, vEKdB14]. differenced [LKM13]. differences [BN12, BdCB11, BFE12, FLB^{+}12, MW14b, UKP14]. differencing [HH12a, MOV12, NMT^{+}12, TK12, TK15]. different [BBC^{+}13, CBGK13, Far11, KSS14, LQ10b, RBS12, SMP^{+}13, SHFB12, SC10, XH13]. Differential [CBW10, PZ14, dCRCS10, AP12a, AB13, ARR11, BBM14, BJ12, CCLS10, CX13, CK11, CHZ13b, CHZ13a, DAEB13, DA14, Dri10, FVR13, Gno14, GM14a, Gri10, Har11, JTXZ12, Kat14, Ler14, LLZ11, LCY11, LS13c, LM13, MZ10, MDTC11, MDTC13, MGN10b, NP12, Pir12, RH13, RHB12, Ton14, TS14, WZ13, WW14, WH11, XH14, YMS12, YXZ13, YK13]. differential-algebraic [CK11]. differentially [CLS10, ZKZ10]. differentiation [hGzSwZ14, VS11]. diffraction [BL13, DB12]. Diffuse [AELV14, FG12, MZ14, ZCD10, Ala14, DjY14, GK14, GGT13, Min13, TSLV11, TFP13, DK14]. Diffuse-interface [ZCD10, GK14, Min13, TSLV11]. diffusing [SBA13]. Diffusion [ERT10, JS13a, KLB^{+}14, Mar10b, AV13, AH10a, AT13a, APY14, Atz10, BHM11, BHM13, BLM^{+}14, BH13b, BCK11, BBAP14, BS10a, BKC^{+}11, BHY14, BTY14, CYK^{+}14, ÇD12b, CGC13, CY12, Cha14, CL14b, CGP10, CM10, Cui12, DTU12, DBO^{+}10, DKW10, DSHP11, EMK11, EMSH14, FAY^{+}13, FHL10, FF12, FCY13, FWLP14, hGzS11, hGzSnZ12, hGzS13, GW13b, GC11, GM14a, HPS13, HL11, HLDP14, HY10, HW14b, HCM10, JC13, JZ13, JN12, KKAS12, LN10a, LK14, LS13b, LP12b, LH10c, LXL13, LRBT14, LYJ14, LXSR12, LSV10, LRN13, LM14a, LCWN12, LHMZ10, LHV13, MB12, MKGV12, MHA10, MSS^{+}14, MN12, Mie13, MAD13, MT14a, MY13, NMV13, NGV14, NMZC13, NLGL14, Nis10, Nis14a, Nis14b, Ols11, OMA14, PBC^{+}11a, PS12, PHRG13, PMS11, PBB14, PvDGvM11, PWB^{+}14, PH13, RSS12, RzSZ13, RF14b, RCCC14]. diffusion [RBK13, Sch13a, SR13a, SB14, SH11, SY11a, SY12b, SVS11, She13, SHA11, SHA12, Sou14, SAA14, TCW^{+}10, TR14, VDP11, VDD^{+}13, VWF13, VBSK10, Vre11, WWS10, WW11, WD13a, WD13b, WV14, WD14, WCVJ13, WDGY10, WGD12, WG14, YZX14, YN10, YN14, ZFH14, nZzS11, ZZS13, nZzSlL14, ZzS11, ZOL^{+}11, ZT10b, ZQ11, vEKdB14]. diffusion-advection-reaction [AV13]. diffusion-based [LXSR12]. Diffusion-controlled [KLB^{+}14, NMV13]. diffusion-reaction [DBO^{+}10, LHMZ10, PBB14]. diffusion-wave [BHY14, LXL13, WV14]. Diffusional [DCVM11, ZM10]. diffusive [AS12, BCD14, FF13, KL10b, KHHK12, MSTL13, NMCZ14, PAM10, RGK12, SAA14, ZKM^{+}11]. diffusivity [BR14, KSL10, KHGW13]. dihedral [Gon12]. dilute [LWF11]. Dimension [TPF11, EEK14, GPS10, MK14a, MMA13, NF12]. Dimensional [BNNW11, AH13a, Ald10, AK12, AN14, ADMA13, And13, ASK^{+}13, AKMD14, BDL11, Bal12b, BD14a, BG10, BILM14, BLM11, BYCC14, BCM^{+}11, BRS10, BBPC14, BHY14, BL13, BD14c, BLR14, BTY14, CWS13, Cap11, CBKK12, CNR10, CTG12, CRS14, CM14, CZ11a, CCB12, CS10, CS14b, CCY13, CCLV10, CLS13, CGC^{+}12, CTSM13, Cui12, DYYA14, DDF14, DB12, DKW10, Du11b, Fal13, FQV13, GMB10, GG10, Ger12, GMC13, GCFJ13, Gra14, GIQ11, GH12b, HdCNT12, Hen13, HKL14, HCH^{+}10, HCS12, IST^{+}12, IDNG13, JAX11, JLZ13, JGZL14, KS11a, KKL10, KSU13, Kau10, KK10b, KG12a, KLP10, KLPS14, KL11b, LP12a, Le 14b, Lee13, LJX^{+}14, LRLL11, LR12, LX13a, LXL13, LW13a, LLZ14, LRN13, LRK11, LW13b, LP11, LT11c, MZ10, MB12, MAB^{+}13, MNF^{+}10, MPT13, MGN10a, MGN11, Nor14, OD14, PB11, PYK10, PE10, QB14, RARO13]. dimensional [RVB10, RB10, RS13b, RW13, SCIE14, SHZ13, SYJ11, SSR14, SS14a, SMS14, Stu10, SWR11, SSY13, Tak14, TAI11, TA13, TP14, Val14, VMA14, VW14b, WCM^{+}14, Wan13a, WW11, WZ11, WZSN12, WD13a, WZN14, WD14, WLW14, WC11, WS14, WS10c, WS11, WC14b, WT14, XHF13, XJ14, YHT11, YT12, YZX14, Yin14, ZFH14, nZzS11, ZWL^{+}12a, ZWL^{+}12b, CDL11]. dimensions [BK14a, BR12a, CCG11, FK10, GH11a, GHS12, GL12b, GHL14, GD13, HR10, HWST12, JZ11, JRTY11, Le 14a, LB12, MB11, MSS12, MGN12a, RNT12, TDU14, TTD11, UGkM13, VRBZ11, WGX13, Wol12]. dimer [ZD12]. diminishing [KHHK12]. diode [MJK13]. diodes [JCT11]. dipolar [BCW10, JCC14, LZZC14]. dipole [ZLL10]. dipoles [MTD14]. Dirac [ASS17, ASS13, Alm14, AME14, BHM14, FGLB14, HPA14a, HPA14b, LSE13, XST13]. Direct [CDT13, KSP13, NVS13, OP10, TRL14, ZN14b, BHB11, BVM14, BLTO11, BD14b, Bre12a, CLS10, CGF^{+}13, DD12, DGAH10, EHA14, GCA13, GG14, GB13, GHL14, GHM13, GH10, HT12, JNSA12, JMW12, KM13, KF12, KGK13, LFL11, Mar13, MMSI13, PPWZ11, SBB13, SY12a, SYJ11, SC11, SSK^{+}12, VRB10, WWS10, WYT14, YHT11, YT12, YS12a, YHL13, dPSS12]. direct-coupling [EHA14]. direct-forcing [JMW12, VRB10]. direction [BLTO11, Cui12, LX13a, LXL13, WW11, WD14, nZzS11]. Directional [MF12, OL13, MSD12, TY11b, JS13a]. directions [CK10, PMSM12]. Directly [LR13, CW14b, YO11]. Dirichlet [AV10, Deg11, FN14, FN10, JRTY11, LL12b, MLBS14, Ost10]. Dirichlet-to-Neumann [AV10]. dis [GAvdVB14]. discharge [DBM^{+}12, KED^{+}12, LA10]. discharges [LEH10, LEH12, LE12]. discontinuities [BH13b, BD12, HCP10, JAX11, Joh11, WI13b, YKWS13a, YKWS13b]. discontinuity [VR14]. Discontinuous [BSCML13, BR12b, Cha13, GP12b, LI14, MMIW14, MK11, PDZ^{+}14, SH14, TR14, YWS13, AJRT11, AJ14, AWK^{+}11, AAT^{+}14, AT13b, APRS14, BNT14, BGR11, BBC^{+}12, BR11, BGHM13, BD11b, CJLC11, CZ11a, CAH14, CLX14, CCZ14a, CCZ14b, CW14b, CSX14, CCY13, CDB13, CCG11, CdCD12, DFFL10, DHH11, DL10a, DZL13, DKA12, DZLD14, DV14, FPN^{+}13, FT12, FH11, FMT11a, FP12, FP14, GDHM11, GAW^{+}14, GLL^{+}11, GZQ13, GXX14, HGMM12, Hig13, HvB11, JNX13, JZ11, KG12a, KKO13, Kol11, KG14, KRT14, Kuz14, LH10a, LXM14, LT11a, LP12b, LXY11, LX12, LH14a, LZZC14, LLP14, LSW14, LLYE12a, LLH10, LP14b, LLN^{+}10, LXL^{+}12, LXS^{+}13, MRS14a, MHA10, MN11, MESV10, MMD^{+}11, ML11b, MBGW13, MGN13, NPC11b, NPC11a, NPC11c, NP12, NvdVB13, OK14, OHF12, OS11a, OD13, Per13, QS11b, RHXQ14, RGMC13]. discontinuous [RC12, RCvdV13, RS11, SHR13, SSHT14, SGC13, SHCY13, SLT^{+}14, SBvdV11, SSL13, SD14c, TE11, TD11b, TD12b, TBR13, VMA14, VR14, WXZ10, WS10b, WZSN12, WK13a, WJTY10, WSW13, WSBG10, XXS10, XZWW12, Xin14, XLD^{+}11, XCL14, YP13, YO11, YMLM^{+}14, YHL13, YWL14, ZS10b, ZS11, ZZS13, ZT13b, ZYS14, ZS13, ZZSQ13, ZMD^{+}11, vdVR12a, vdVR12b, FW12, MOSW12, NDBG14, WKH10]. discontinuous-Galerkin [AWK^{+}11]. Discontinuous-Galerkin- [MOSW12]. discontinuous/continuous [DKA12]. discovery [BCM12, KKPV12]. discrepancy [LMM14]. Discrete [GAV13, KST13, AS14, ACB^{+}13, BCHM14, BPS14, BSCML13, BWMG14, BBM14, BM14, CFW^{+}13, CT10, CST10, GB14, GH11a, GD10, GL12a, GLC^{+}11, GXX14, GLL14, JWNL11, KPH13, KH11, LMFH11, Ler13, Ler14, LTSH10, MJ14, MRS14b, MM11, MVG11, SKW14, ST11a, SV10, TAAS13, TDL10, iT14, WZ11, YMS10, YMS12, ZYHW13, Zie11, SBN12]. discrete-fracture [ZYHW13]. Discretely [FCN^{+}13]. Discretisation [PAM10, CYK^{+}14, CVC12, SS10e]. discretisations [BLTO11, TA12, VSK10]. Discretization [Dav10, Dav15, KD10, SRS10, BdCB11, BSB14, BR10b, BMR14, CZ11a, Den11b, DHH11, DLM13c, EQYF13, FD11, Ge10, GG14, GHM14, LE13, LRN13, MMR12, MIJ10, MTG11, MG14b, Nor13, PRH^{+}14, PGR10, PvDtTB^{+}11, PMSM12, PJCS14, RF14a, SH14, SG14, SBT13, SSL13, SDH12, TD12a, TLO^{+}14, VW14a, VMA14, VKBK11, WAM14, YHM12]. Discretizations [SYV17, AB12, AJ14, AMP14, BBC^{+}12, BN13, BN14a, BRP14, BSV14, Gas11, GDHM11, JN12, MN11, SVK14, SYV14, SD14c, TD12b, WABI11, vdVR12a, vdVR12b, vV12]. discretize [TGO13]. discretized [Han11, LTC10, LTSH10, Mar13, NDBG14]. disjoint [Kim13]. disk [BY11, ZL13]. disks [YTYJ12]. dislocation [FTF14]. dispersal [LY14]. disperse [GJLY13]. Dispersion [KWJR14, TCL11, BSS14, CWS13, CD13b, CCW12, DBHBB13, GB14, GJ11, Gür14, HPA14a, Hig13, JW11, LS13d, MPS^{+}14, MMP11, NYM13, RSD10, SRL^{+}11, sSLxRyZ14, WA12, XS11]. dispersion-relation-based [LS13d]. Dispersive [Ain14, BRS11, BCL^{+}11, DKM11, HLY11, NZ14, OJ11, PDZ^{+}14, SB14, WXZ10]. dispersively [YS13]. displacement [BCG^{+}13, MMH12, ZL13]. displacements [ES13a]. dissection [SY14a]. dissimilar [WQCS13]. dissipation [BJLR14, CGM^{+}12, DBHBB13, GJ11, HAD10, HJK14, HK11, JXD14, KKP10, LFL11, LGC13, NYM13, Rie10, SRL^{+}11, sSLxRyZ14, TiI11, XS11, XW14, ZGSZ10]. dissipation-rate [JXD14]. dissipative [FR13, GPY14, LH10a, LTL^{+}14, LSW14, LEHA10, MGLS12, Pir11, RPV13, SCY11, VW14a, WYS^{+}11, WLL13, LFK11, MDPTK13, PAG11]. dissolution [YL10]. distance [ERT10, RS13a, WYS12a, ZT10a]. distances [SC12]. Distinguishing [OJ11]. distorted [WDGY10]. distributed [FB11, GSTB13, Kat14]. Distribution [MGJZ10, ALR11, BND^{+}12, CZL10, Cou11, HLT10b, JC10, LSK11, RA10b, SHR13, VQRD11, VQVD11, YSW14]. distributions [JW12]. disturbance [RZ10]. disturbances [VMS12]. Divergence [Kaw13, Nis12, AHA10b, BK12b, EH13, GW13a, GCvRK11, KAA12, KG13c, Li10b, LXY11, LX12, McD14b, NCV10, TS12, TP12, YVL^{+}11, YB14]. divergence-conforming [EH13]. divergence-free [KG13c, Li10b, LXY11, LX12, NCV10, YB14]. Divergence-free-preserving [Kaw13]. divergent [RH13]. divertor [DRB14]. divide [BRP14]. DNA [KLB^{+}14]. DNS [SSHM10, SRL^{+}11, VdMS^{+}14, YYB12]. Do [MW14b]. does [Den11a]. Domain [BPM13, ILM^{+}13, OK10, SKN13b, AWK^{+}11, AS12, AF13, APY14, BHB11, BILM14, BCB12, BAG12, BHY14, BL11, CES14, CBKK12, CW14a, CL11a, CWZ^{+}11, CVI^{+}13, DYYA14, DDF14, DMR14, DFFL10, DL10a, DZL13, DBHBB13, GB14, GLLP14, hGzSnZ12, hGzS13, GVH14, GP12b, GK10, GT11, HS13a, HR10, HBZL14, HJK14, HDD12, JC13, KKS13, LMLC12, LB12, LCNK10, LHY12, LTT10, LS13d, MBLM10, MN14, NZ14, NB14, PL10, PRL10, RVB10, RHB12, RHGT10, SBWF14, SABH11, SY11b, SF12, SSB^{+}12, SCH13b, SSFM12, SFWP10, Soa13, Sto13, Stu10, SS14c, TH14a, TD12a, VHG13, WJC13, WC14b, XMDG13, XMDG14, XH14, YD11]. domains [AJRT11, AAI10, AB11, AEM13, AT13b, AM11, BLQ14, BvBZ^{+}10, BUH11, BKC^{+}11, BR10b, BRS10, BL10, CCFCM11, CS13, CGS13, CR13, DKC14, DRZ14, Era13, FC13, GPY14, GM12b, GHM14, HG11, HWST12, JPC^{+}14, JKR14, JK11, KTT12, KF13a, KTT13, LP12a, LK14, LMZ14, LN10c, LB10, MKCB14, NS14b, PHSA14, PZ10, RC12, SJK11, SR11, TDVG13, WS14, Yin14, Zen11, ZWH^{+}12]. dominated [BD11b, KWSJ11, LFRM10, TLG10, vdVR12a, vdVR12b]. doping [CW10, CGR11]. Dory [VCS14]. dot [HWHW10]. dots [PBB14]. Double [EZ10, VG14, CBGK13, KHHK12, HGRB14]. double-diffusive [KHHK12]. Double-families [EZ10]. doubly [BS10a, CH14]. doubly-periodic [CH14]. DPD [MDPTK13, PPTMDK13]. drag [Min10b, SL13b]. drift [BH13b, MSS^{+}14, PMS11]. drift-diffusion-Poisson [BH13b]. driven [AF11, CLAL14, EMS11, GCFJ13, KWSJ11, KFOS14, KL12b, LS13a, LL12a, MWRZ13, RTS13, SGC13]. drop [BB13, DL10c, LL10a, YFL^{+}13]. Droplet [JW12, KBB13, LZ10, SL13b, Yok13]. droplet-laden [KBB13]. droplets [RWMG11, Zoh13]. drops [CL13]. Drude [HZS12, LBZ^{+}10]. drug [RUL^{+}13]. drum [ZWL10]. dry [DLM13b, LL10a, MHHGN14, MBGW13, YMLM^{+}14]. drying [CTP13]. DSMC [GWK10, Mac10, PMF13, SGV13, SG13, Sza12]. DSMC/Navier [SGV13]. DSMC/NEMD [GWK10]. DT [Nor14]. Dual [HZ14, BLRR11, BN13, BN14a, DjY14, KHGW13, PRH^{+}14, XJ12, XJ14]. dual-mesh [XJ12]. dual-primal [XJ14]. dual-resolution [DjY14]. Duality [BN14a]. duct [HY09, HY10, HY11, LT12]. due [BR14, CWL10, HH13, Mit13]. during [CRS14, IBO14, JNW^{+}14]. dust [Min10b]. Dynamic [PWB^{+}14, RVB10, RXLS14, CTSM13, Dad12, DDM10, Don10, GIQ11, HLZ14, JTH10, LDS11, SC10, SS13b, SS14b, Wit13, BM13b]. dynamical [APC11, CTG12, Chi12, GLM14, NT14, SP13, UJ12, Wan13b, YL14a, YL14b]. dynamically [CHZ13b, CHZ13a, CSK14, MN11, MMD^{+}11, RWX11, ULS13, YJK12, CSK13]. Dynamically-Orthogonal [CSK13]. Dynamics [MDPTK13, STL^{+}10, Abr10, AHA10a, AAI10, AGBC14, ARF12, AEM13, AM10, ACB^{+}13, BCW10, BD11a, BB12a, BHS14c, BLJ11, BDLGC13, CCT11, CP12, CD12a, CFCA13a, CFCA13b, CWC11, Chi12, Dav10, Dav15, DXB^{+}13, DKR14, DLM^{+}13a, Don11, Fal13, FBM14, FTF14, GPY14, GCLS11, GW11, GL13, HKL14, HCS^{+}10, iIT14, JZ11, JFC^{+}13, KH11, KL10a, KIF^{+}14, Kol11, KMHJ10, KSB14, KMSS10, KAG14, LSM14, LB13, LKLG14, LJX^{+}14, LSK11, LYE10, LJ13a, LTL^{+}14, LEHA10, LZ10, MCJ14, MK14a, ML14a, MB13b, MTZ14, NIN^{+}14, NNS11, ORHH13, OLC13, PTM13, PL12, PZ10, RTS13, RPV13, RTKS10, SB14, SSC13, SSB^{+}12, SSW13, SKW14, STC10b, TQW14, VMA14, Wan11b, WLL13, WZT^{+}14, WYT14, ZD11a, ZD12, ZYF^{+}10, ZKM^{+}11, dPSS12, LFK11, PAG11]. dynamics-based [LJ13a]. dynamo [LLJ10].

E-CUSP [SZH12]. Eady [BCW13]. Earth [ASXZ14, RKGM14, TRSMM^{+}13]. EAST [KYP^{+}14]. easy [SdAW14, Yao14]. easy-to-implement [SdAW14]. Eca [EH15, XS15]. Eccentered [PMM^{+}11]. eccentric [GDW11]. Eddy [CP13a, ADSF13, BGGW13, BA10, BM13a, DV13b, GW10, HA11, KSL10, KSB12, LFL11, MS14a, McD14b, OL13, OAK11, PGLT10, dlLPC10, dlLPCP12, RG13, SZ10, SFCF10, Ste14, TRL14, TGO13, VKBK11, WDP^{+}14, Den11a, MDH^{+}10]. eddy- [TGO13]. EDG [BR12b]. edge [AAG14, AST12, BD13, CXZ10, Cou11, GIQ11, ICC^{+}10, KBSV14, LHY13, Soa13, SS10e, TGT^{+}10]. edge-based [Soa13, SS10e]. edged [HWW10]. Editor [HAD10]. Editorial [Ano12f, Ano12g, Ano12h, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano11y, Ano11z, Ano12b, Ano12c, Ano12d, Ano12e, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano13i, Ano13j, Ano13k, Ano13l, Ano14-34]. EDQNM [MS14b]. EEM [KIF^{+}14]. Effect [JC10, FGE11, JR14, KH13, SSR^{+}13, TAI11, WJCG13, YHL13]. Effective [CFCA13a, CFCA13b, BdCB11, CR14, Hel11a, Hel11b, JLZ13, KWJR14, KJWR14, LKP14]. Effects [Coo13, GLL^{+}11, Lee12a, SB14, BB12a, BPHK13, DDM10, Gen11, Gho17, JWNL11, KS11b, LJG14]. efficiencies [LCP13]. efficiency [ASB10, CM13, GKG^{+}13, LTC14, PZW14, QL12, SEBG12, YWL14]. Efficient [ADT14, ABK12, ABK13, BMD^{+}13, BCW10, BCZ13, BLRR11, BF12, BR10b, CNSFD13, CS12b, CHHL11, DHH11, FNGV14, GM12a, GRT12, GKS^{+}11, GD13, Ham10, KL11b, LdWF12, LSE13, LFRM10, LL14b, MOV12, Men12, MY13, Nic11, NDB12, PGR10, PMSM12, Ran11, RS13b, RS14b, SR13a, TWSN12, TCW^{+}10, WWS13b, dPSS12, ARR13, AD14, ABH^{+}14, BHL^{+}13, BBAP14, BK12b, BZKL13, BSCML13, CY12, CCB12, DL13a, DL13b, Don14a, DL10b, DEKBF14, FL13, FP14, GW14, GMT10, GL13, GHM14, GRS14, GJ14b, GXX14, HAMA14, HNS12, HLZ13, Kri10, Lee13, LLX11b, LMZ14, LD14, LH14b, LCWN12, LFW11, MJ13a, MF10, MTZ14, MJ13b, MMH12, NDV^{+}11, Nor14, NT14, OTV13, RA10a, RS14a, RWMG11, SBN12, SS11, SJB11, SLT^{+}14, STH11, SR11, Shu14, SSR14, SS13c, TY11a, TLZ^{+}11, Wan11b, WT12a, WD13a]. efficient [WG11, WJC13, YL11, YS12a, vHtTBC11]. efficiently [VSK10]. Eigen [GZQ13, ZK13, ZD11b]. eigen-based [ZD11b]. eigen-problems [ZK13]. Eigen-structure [GZQ13]. eigenfunctions [Hua14]. eigenmodes [HP13]. Eigensolutions [RSK13]. eigensolver [BWC13, JMKK12]. eigenspectra [EEK14]. eigenvalue [Alm14, BDLGC13, CBPS13, CFW^{+}13, CXZ10, CHHL11, GLSC12, GBB^{+}13, HAMA14, HCH^{+}10, Hua14, HWHW10, LLJ10, MBR11, MRS14b, Ren11, SKCC10, WPK14, Xie14b, fZlC11, Zho10]. eigenvalues [And10, And13, JS13b, Ost10]. eigenvectors [And10]. eight [And13]. eikonal [DGM13, Luo13]. Einstein [AD14, BCW10, BCZ13, CCW11, CCC12, DH10, FT12, JCC14, LZZC14, MTZ14, ST13, Wan11b, Wan14, WC14a]. Eisenbud [RSK13]. elastic [BHS14a, Ber14, BBGP13, BLR14, DKR14, DCVM11, DV14, EL11, FT11, FG12, Ham10, HPOM10, HCS12, JXD14, KK14, Kwo14, Le 14a, MAB^{+}13, ML14a, OLC13, OLPM14, SG12a, SG12b, SCH13b, SM12c, TLZ^{+}11, Vol14, WSBG10]. elastic-acoustic [WSBG10]. elastic-plastic [DKR14, FG12, HPOM10, MAB^{+}13, OLPM14]. elasticity [ABHS12, BD10b, LSZ13, NPV14, TDVG13, ZWH^{+}12]. elasto [SSB13]. elasto-plastic [SSB13]. elastodynamic [CB14, Hut14]. elastodynamics [FN14, Ide12, KN14, NPC11a]. Elastoplast [BR12b]. elastoplastic [SNBN13]. elastoplastic-solids [SNBN13]. Electric [KYP^{+}14, BLM11, Cam13, EAG14, FHMS11, KED^{+}12, OBTC^{+}13, SCB10]. electrical [ABK12, ABK13, GJ14a, LHH14]. electrocardiographic [WKMJ13]. electrochemical [BP10, EBGW13, TCW^{+}10]. electrode [LHH14]. electrodiffusion [LGP10]. electrodynamic [YS12b]. electrohydrodynamic [LHPH11, RDHK12, VDD10a]. electrohydrodynamics [BBGP14]. electrokinetic [BDH13, LBK10, WK10]. electrolocation [BBGP14]. electrolyte [LTC14]. electrolytes [BDH13, Liu13]. electromagnetic [AWK^{+}11, BWC13, BLM11, CNR11, Du11a, Du11b, DSZ13, Fuj11, GHS12, HNS12, LTC13, Le 14b, Li10a, LHY12, LHY13, LLSW14, Mar10a, MTD14, Nic10, Par12b, PRL10, RKGM14, Soa13, TK14, TY11b, VHG13, ZLL10, Zho14b]. electromagnetics [BWC11, BCB12, BSV14, DFFL10, KAA12, KO14, OK13, PL10, WBC13]. electromagnetism [CV10, PT12]. Electron [PMS14, BGK12, CKD10, CM10, NIN^{+}14, SL13a, WS14]. electronic [BHL12a, BHK^{+}10, CBPS13, CW10, ITO^{+}10, Men12, Son11]. electronic-structure [ITO^{+}10]. electrons [BKH12, KHJ^{+}13, SBS^{+}13]. electroosmotic [LWCR10]. electrophysiology [CYK^{+}14]. electrostatic [CCB13, CCB11, CC13, CCL^{+}14, DXB^{+}13, LTC14, PA14, YD11]. electrostatics [Gen13, GK13, WRH10, Xie14a]. Element [ABMT14, ABH^{+}14, AB12, Ain14, ASS13, ASS17, AKL^{+}13, ABJ^{+}12, APC11, BGGGS11, BCP13, BPM13, BFR13, BHL12a, BSSW12, BWS10, BGN10, BW13, BFSVC14, BKC^{+}11, BCLR10, BTY14, BKV12, CN12, CYK^{+}14, CMS10a, CTG12, CT13, CXLF14, CLG13, CW14b, CH11a, CH11b, CT14, DH10, DB12, DSZ13, DKA12, DCVM11, DZLD14, EGW11, EGH13, EGL^{+}13, EN14a, EN14b, ES14, FPN^{+}13, FCG10, FK10, GT12, GDGP11, GTSC14, GQ13, HRK^{+}10, HR10, HPV13, HSN13, HGW11, HHK13, HvB11, HL12b, HY09, HY11, HZ11b, HLLW11, HLYS11, Hua14, Hut14, Hys12, Ian11, JNX13, JZ11, KR10, KGG10, KG13b, KEGM10, KBSV14, KG13c, KVBP^{+}14, KO14, LSM14, LRL10, LCNK10, LB13, LJX^{+}14, LH10c, Li10a, LHY12, LXL13, LHY13, LSZ13, LRBT14, LLMY14, LLS11, LHMZ10, MZ11b, MRS14a, MMV^{+}13a, MMV^{+}13b, MBMV13, ML11b, MK11, MVCFM13]. element [MSTL13, MSTL14, MIKG12, MNL^{+}13, MWY14, NKS10, ORHH13, OS12, PMM^{+}11, PR11, PBC11b, QZM^{+}12, RSS10, RKGM14, RCCW10, RHGT10, SF11, SL13a, Sog14b, SLC^{+}11, SK12b, SBvdV11, SJH12, SS13b, SS14b, SDH12, TF10, TD11a, TCL11, TTD11, VSK10, WA12, WKMJ13, WZZZ14, WYC13, XH13, XJ14, YL11, ZM10, ZK14, ZW10a, ZOWZ10, ZOJR11, ZYHW13, ZYT13, ZKM^{+}11]. element-finite [GQ13, RKGM14]. element-free [ZOWZ10]. element/Fourier [SJH12]. elements [ABK12, ABK13, BMS12, BD13, BSB14, BL10, CXZ10, CKR14, CS12c, CMBW13, DMHL14, DKR14, DL10a, DTYY14, ES10, FGZ12, FGE11, HBZL14, HMM^{+}10, HZS12, JMFO13, JRG13, MKGV12, MTS^{+}11, NLGL14, OS11b, RBK13, SCGE13, TEY12, VQVD11, ZTG^{+}13, ZD11b]. elimination [Opp13]. ellipsoidal [MWRZ13]. ellipsoidal-statistical [MWRZ13]. ellipsoids [GL13]. Elliptic [SG12a, AB12, BvBZ^{+}10, BL11, CK10, CS13, CV10, CR13, DLNN12, FGE11, FO11, GM10b, GKN^{+}11, Gri10, HPS13, HWST12, HWW10, HSWZ13, Hua14, JTXZ12, Kau10, LB12, LMS11, LC14, MDTC11, MMM13, Mar13, MWW^{+}13, OK14, OS11a, PG11, PPB14, SS11, SG12b, SY12a, SY14a, SD14c, XZW11, XZWW12, YK13, LB10]. elliptical [Lee12b, LR13]. ellipticity [TDL10]. Elsvr [Ano12-31]. embedded [BF14, CCG11, FL14, JM10, KT14, LKTL10, MP13, SK12b, SdAW14, UKP14, VRB10, MDM^{+}12]. embedded-boundary [VRB10]. Emden [xLqW13]. emission [Den11b, LGP14, RKD^{+}14]. empirical [CEGG14]. employing [GV13]. enabled [HJBB14]. enclosed [LT10a, LW11]. encode [ACG14]. endocytosis [TH14b]. endotoxemia [SMF^{+}13]. energetic [ABL^{+}11, BM10a]. Energy [BHL^{+}13, CCZ14a, CCZ14b, GKG^{+}13, HJK14, HLLW11, LTL^{+}14, MRS14a, MGN13, RHB12, San13, WZZZ14, WCVJ13, YHL13, AM10, BK12b, CWL13, CGM^{+}12, CCB13, CTG12, CCB11, CC13, CSX14, CJP^{+}14, DP12, FCYF11, FMT11a, Fus13, GZZ14, GLL14, HAD10, HCP10, JYHT10, KBS14, KJ11, LH13, LY13b, LLYE12a, LY14, LW14, ML11a, MHGM14, MGN14b, OS14, PS14a, PS15, RTKS10, SLI10, SO13, SEDW13, TBNT11, WGX13, WWS13b, vV12]. energy- [CCB11]. energy-based [DP12]. energy-conserved [LY13b]. Energy-conserving [CCZ14a, CCZ14b, LTL^{+}14, San13, CCB13, CC13]. Energy-dissipation [HJK14]. Energy-preserving [WZZZ14, CWL13, WWS13b]. energy-stable [GZZ14]. enforce [NMV13]. enforcement [Nic11]. Enforcing [NV14]. engineers [PZ14]. Enhanced [HSD11, LLW14, GCBH13, KGK13, LJX^{+}14, LFRM10, RF14a, XM14, ZMG11]. Enhancement [KG11, KG13a, Zho11, KRF12]. enhancements [BK11]. Enhancing [HGSK^{+}13]. EnKF [CZL10]. ENO [SIDG13, WI13a]. ENO-type [WI13a]. enriched [CJ14, MMD^{+}11, MSTL14, Sog14b]. Enriching [Try12]. enrichment [MSTL13]. Ensemble [GLM14, EWH13, Gil13, HMM14, KSS14]. ENSO [OMCO14]. enstrophy [KJ11]. enthalpy [BKKD14, CLZ12, KBTD12]. enthalpy-based [CLZ12]. entirely [BGZ12]. entrainment [Mit13]. Entropy [CVIA10, GPP11, LY14, AHOT14, DL12, FC13, FLM12, LGE^{+}13, MT13, RRCD14, RPA13, Rot13]. entropy-based [AHOT14]. Entropy/energy [LY14]. entry [CRS14]. envelope [CBCM^{+}11]. EPIRK [Tok11b]. epistemic [CPX13, JEX10]. EPS [SW11]. Equation [MM11, AH13a, Ain14, AJ14, ASS13, ASS17, AME14, BK14a, BH12, BD11a, BTX13, BHL^{+}13, BFO14, BR14, BB12b, BKH12, BAG12, BY11, BX12, BHY14, CCT11, CW13, CD13b, CES14, ÇD12b, CdCNH14, CB12, CT13, CGG12, CEP12, CRS14, CCW12, CLX14, CH10, CSX14, CGS13, Coa12, CRS12, Cui12, DDL13, DO11a, DOR10, DGM13, DL13a, Dur14, DV14, EMSH14, FKM11, FGZ12, FHMS11, FGLB14, FK14, FE11, FO11, GH14, Ge10, GCZ13, GMC13, GMA10, GG14, GM12b, GBB^{+}13, GH12a, GAV13, GL12b, GHJ14, GT11, GQ13, HM11, HPA14a, HPA14b, HZ11a, HPV13, HMS10, HRCW13, HLDP14, HvB11, HWW10, HW14b, JWV12, JNSA12, JC13, JRTY11, JW11, JKQ13, JWYH10, JY11, JM10, KKS13, KS11c, KJC13, KBGRM10, KL10b, KK10c, KQ11, KTT13, LH10a, LP12a]. equation [LKT^{+}12, LQ10a, LCY11, LX13a, LXL13, LMK13, xLqW13, LW13a, LLP14, LX14, LYJ14, LM14a, LDE^{+}13, Liu13, LS13d, MMR12, Mar10b, MKGV12, MHA10, MD14, MP13, MTG11, ML13, MYM12, MTS^{+}11, MAD13, MZB12, MVG11, NLGL14, NN12, Nis10, NCV10, PSL20, PRH^{+}14, PSM11, PBB14, PCF^{+}13, PS14a, PS15, PP12, PLN11, Pri14, QY10, QC10, RzSZ13, RLJ14, RVbZ10, RJG10, RS13b, RS14b, RNT12, SOG14a, Sam14, SSS10, SB14, SF14, SY12b, SCY11, SYL12, SJK11, SSX13, SSR14, ST11a, SJH12, SC12, Sto13, SYE10, Stu10, Tak14, Tal13, TG14a, TAT14, TGGT13, VW14b, WGX13, WV14, WZZZ14, WP10, WMYYG10, WJTY10, WG11, WDGY10, WWS^{+}13a, Xie14a, XQXC14, XST13, XX13, YXZ13, YZX14, Yao14, YHL13, YK12, YN10, YN14, YS13, ZN10, ZT13a, ZFH14, ZW10a, ZOWZ10, nZzS11, ZTG^{+}13, nZzSlL14]. equation [Zha10b, ZzS11, ZTL13, ZPF10, ZGSZ10, ZCS14, Zho14b, vTCL^{+}10]. equations [AK14, AP12a, AB13, ARR11, AIX10, AB11, AK12, AHOT14, AMDR10, AMP14, ÁCB12, AN14, AN13, AAT^{+}14, ALS^{+}14, AM11, APC11, BK11, BLQ14, BLRR11, BGR11, BGH10, BNNW11, Bér11, BN11, BN14a, BF12, BGHM13, Boe11, BdCB11, BKC^{+}11, BWMG14, BR12b, BRVE12, BDLGC13, BCLR10, BHL12b, BR10b, BRS10, Bre12a, BMR14, BTY14, BJ12, CVN13, CWS13, CCGP13, CCLS10, CX13, Cap11, CVC12, CGC13, CEK14, CJNS12, Cha13, Cha14, CS12b, CRG13, CS13, CL14b, CAH14, CXLF14, CHZ13b, CHZ13a, CLQX13, CW14b, CS14b, CLS10, CCY13, CRT10, CDK12, CT14, CG13b, CCG11, CMS10b, CLM13, CTSM13, CST12, Dad12, DLNN12, DCC^{+}13, DL13c, DAEB13, DBAV14, DHH11, DA14, DZL13, Dri10, DCL11, DAA11, DLM13c, EAG14, EH13, FPN^{+}13, FAY^{+}13, FL11a, FJ10, FR13, FMT11a, FM13]. equations [FW11, FVR13, FCY13, hGzS11, hGzSnZ12, hGzS13, Gas11, GL14, GM10b, GHM14, GG13, Gno14, GPS10, GM14a, Gri10, Gri12a, GLWW14, GMG12, GLL14, GXS14, HH12a, HR10, Har11, HHL11, HL12a, HKM12, HK14, HRT11, HOK10, HHGC14, HM10, HJK14, HJZ14, HDD12, HY10, HLT10b, Hu13, HAS13, HLY11, HLYS11, HF10, JMFO13, JU10, JKW11, JM12, JTH10, JCT11, JRTY11, JZ13, JLC14, JN12, JTXZ12, KL12a, KM14a, KKP10, Kat14, KDS14, KJ11, KHZ10, KG14, Kri14, KRT14, KMSS10, LN10a, LKM13, LT11a, Le 14b, LRL10, LTCN13, LK14, LS13b, Ler14, LLZ11, LTSH10, LXY11, LX12, LP14a, LY13b, LD14, LSW14, LS13c, LSV10, LRN13, LC14, LGE^{+}13, LP14b, LW14, LBM14, LM13, LDA10, LHMZ10, LL12b, LHV13, LL14b, LLN^{+}10, LXL^{+}12, Luo13, LQB14, MZ10, MA14, MMM13, MBS13, Mar10b, MMV^{+}13b, MIJ10]. equations [MMP11, MS14b, MN12, MAK14, MST14, MBA14, MMD^{+}11, MG14b, MT14a, MHGNM12, MY13, MWY14, MGN10b, MGN13, NMZC13, NPC11b, NPC11c, NP12, NOT14, NDBG14, NEE12, NvdVB13, OK14, OA10, OLFS13, OR10, Ols11, OBT12, OS12, PBT14, PDZ^{+}14, PS12, PSM13, PMS14, PVLT10, PGL11, PPB14, PWK11, PE10, PvDtTB^{+}11, PT12, PIN14, Pir12, PH13, Pir11, PMSM12, Pud11, PR14b, QY12, RBS12, RTS13, RP13, RH13, RHB12, RHXQ14, RGMC13, RCvdV13, RF14a, RS11, RM12, Sal14, SK12a, San13, SVK14, SS11, SH14, SM13, Sen13, SLGB12, Sha13, SZH12, STH11, SY11a, SVS11, She13, SML13, SWX12, SF12, SR11, SY11c, SR13c, Sou14, SN13b, SP11, ST11b, TAI11, TA12, TDVG13, TWMM14, TY11a, TCL11, Tok11a, TT10, Ton14, TE11, TS14]. equations [TLO^{+}14, TEY12, TBR13, ULS13, UJvL10, UN13, VV13, VK12, VTTK13, VWF13, VMA14, VQRD11, VBCM10, Vre14, WA12, WCM^{+}14, WZ13, WWS10, WXZ10, WPP10, WW11, WSYS12, WXY13, WW13, WD13a, WD13b, WD14, WW14, WXY14, WX14, WMYYG10, WH11, WLW13b, WAM14, WP14, Wit10, WG14, WYT14, XZW11, XZWW12, XFB^{+}14, XIIX14, Xin14, XH13, XM14, XH14, XJ14, YVL^{+}11, YO11, YC11, YWS13, YK13, YF14, YHM12, YF11, ZD10, Zen11, ZS10b, ZS11, ZS12, ZRM13, ZZS13, ZYLK13, ZYS14, ZW10b, ZT10b, ZLH12, ZTS^{+}10, ZQ13, vDS12, vV12, RB10]. equiaxed [KBTD12]. equidistribution [CDF11]. equilibrated [NB14]. equilibrium [BBF12, CVIA10, IAM^{+}12, JYHT10, MM11, MF13, PWB^{+}14, RPC14, SGV13, VP10, WYS^{+}11, Xin14, ZCD10, ZLCW14]. equivalence [Boy11, CSK14, SSM14]. Equivalent [CM10, Ler14, Wen10]. ergodic [YL14a]. Ericksen [CTSM13]. erodible [MGN10a]. Erratum [HY11, LWD13a]. Error [APY14, LH10b, MM10, SHFB12, SB14, WXZ10, WW13, WH11, AM12, Cou11, DFVY14, DBM^{+}12, ETW10, FdVV12, FVR13, GLS10, GCLS11, GBB^{+}13, HHL11, HLDP14, Hic12, HA14, HT12, HKL10, LLS11, NMT^{+}12, RKM11, SBT13, SEBG12, XY14]. Error-controlled [WH11]. errors [AH13b, CN10, Dav10, Dav15, GHM10a, GHM10b, JH12, LH10d, PSDF13, SM12b, SS14a, VKBK11, XL11]. Esirkepov [GV13]. essentially [CCD11b, CRT14, DB13, Fan14, FSTY14, FCYF11, HKLY13, HWA10, IG14, JN12, LQ10b, VKL^{+}13, YF14, ZS13]. essentially-bound-preserving [VKL^{+}13]. estimates [DFVY14, ECDB14, HLDP14, HKL10, SEBG12]. estimating [GMT10]. Estimation [Cam13, BJ12, EH14, EH15, EWH13, ETW10, FLM12, FdVV12, FVR13, GHM10a, GHM10b, HMM14, HHL11, Hic12, HA14, KG13b, LA10, LH10b, MPS^{+}14, MM10, MOV12, XS15, vydCNS^{+}10]. estimators [KS13b]. EULAG [SC13]. Euler [AK14, Bal10, Bal12b, BMD^{+}13, BDA14, BGR11, BN14a, BYCC14, CCGP13, Cap11, CD13a, CS14b, CRT10, CDK12, DHN11, DDS12, Del10, DCL11, GITW12, GXS14, HKM12, HDD12, HLT10b, Hu13, HAS13, JKW11, KM14a, KG14, Kri14, LTCN13, LX13a, LKTL10, LDA10, LXL^{+}12, MMV^{+}13b, MIJ10, NvdVB13, PIN14, Pir11, PR14b, SLGB12, SYY14a, TAI11, TE11, WCM^{+}14, Wit10, WYT14, YWS13, ZS10b, ZS11, ZS12, ZYLK13, ZW10b, vV12]. Eulerian [BD14b, AH10a, Atz11, BHS14c, BD10b, BDMP13, BBG^{+}11, BR12a, BRSS11, CBW10, DLM^{+}13a, ES13a, GT12, GIQ11, Her10, HPOM10, HAQ12, KBS14, KS12, KS14, LQ10a, Leu11, Li13, LHB10, LSVW10, LMS^{+}10, MRD13, MVS13, OS11b, OLPM14, OS14, Ric13, Sam14, SNBN13, SZF12, SX14, SB12, SIT^{+}11, VBCM10, Vre11, Wan10, WYC13, WYT14, YHT11, YT12, YL14a, YL14b, ZS14]. eutectic [BKKD14]. evaluating [ZJT10]. Evaluation [LX10, AINF13, BG13, BHBM10, GW13a, Gon10, Gon12, KBGO13, LLL10, NCKN10, QL12, SBWF14, VBSK10, dPSS12]. evaporating [VLM13]. evaporation [KLMJ12, MB13a, PS14a, PS15]. even [DAEB13, LS13d]. even-order [DAEB13, LS13d]. Event [BERB^{+}12]. Event-based [BERB^{+}12]. evolution [AJV11, BDKV14, CSKP10, DCVM11, GL14, GMG12, Hen13, HAD10, JWNL11, LX13a, LP14b, LMQ^{+}11, Ols11, RKM11, RSS10, Vai11, VK12, WT14, YMLM^{+}14]. evolutionary [APY14, CCLS10, JN12, LY14]. evolving [BF14, BD12, DKW10, LLZ11]. EVP [LKT^{+}12]. Ewald [LT11b, WLL13]. Ewald-based [LT11b]. Exact [RPC14, hT10b, VTTK13, BBAP14, FGR12, HTHG14, LTL11, MF14, SLI10]. Exactly [Xin14, DLGP13, GM11a, LXY11, LX12]. exceeding [YM14]. exceptional [CDLL12]. exchange [ABS^{+}14, BHBM10, Men12, Nag14a, Nag14b, PMSM12]. exchangers [PBdGP14]. excitation [VK12, WS14]. excited [CCW11]. exhibiting [NW10, SCY11]. Exner [MGN10a]. Exner-based [MGN10a]. expansion [BS11b, DV13a, KBGO13, MTS^{+}11, PSDF13, RJC13, VBSK10, ZD11b]. expansions [AGS10, BY11, FN14, LdWF12, MH10, PHD14, RJG10, Tyg10, Nic11]. Expectation [GJ14a, AETT11]. expectation-maximization [AETT11]. experimental [HM13]. experiments [BSDM12, KYP^{+}14, MBGW13, SEBG12]. Explicit [AB11, GDHM11, GMA10, RA10b, SWX12, WLW13b, ZL10, AV13, Abr10, AK12, BRVE12, CCMG^{+}13, DV13b, DKH11, Fal13, FP14, GHP13, Gui12, HK14, HW14c, KK10a, KHHK12, Kuz14, LCB12, MBA14, MHGNM12, MHHGN14, NNS11, PP12, PP14, RGMC13, SK12a, SLT^{+}14, SSHM10, Sou14, ZY13, LKLG14, DE14]. Explicit-Implicit-Null [DE14]. Explicit/Vertically [WLW13b]. Explorer [SBS^{+}13]. Exploring [CCW11, JJM13, JEU11, LBRA^{+}13, JHZ10]. explosives [SNBN13]. exponent [Leu11, CRK12]. Exponential [GBNS14, HBM13, LP14a, TS14, ZBS12a, ZBS12b, AF11, CTP13, CK11, LT12, RT14, TLG10, Tok11b]. Exponential-Krylov [TS14]. Exponentially [ZK14]. expressed [SBS11]. expressions [YYY^{+}14]. extended [BGM12, DRZ14, DCVM11, FK14, GDW11, GHM10a, KDS14, KBSV14, LL10b, MCJ14, Rot13, SF11, SWX12, VHWF13, YJK12, ZYS14].

Extending [CG13a]. Extension [KBB13, LD12, LR12, Liu11b, CGP10, MHHGN14, SC13, YMS10, ZT10a]. extensions [AR14, KS13b]. exterior [CT14, GH11a, GHJ14, KST13, SSM14, SY11b]. External [IH13]. extracted [ELM14]. extraction [MM12]. extrapolated [FF12]. Extrapolation [KSS14, FVR13, MCD14a]. extreme [Hel11a, SV13]. extremely [Gui12]. extremum [GW13b, SY11a]. extremum-preserving [GW13b].

face [PLN11]. face-centered-cubic [PLN11]. faceted [NW12]. facies [CZL10]. facing [PBC^{+}14, WZG^{+}13]. factor [CZ11a, GP12b, JZ13, LN10c, WZN14, ZOL^{+}11]. Factored [LQ11]. factorization [WdHX10, BNNW11]. factors [YM14]. Failsafe [KMSS10]. failure [LX10, LLX11b]. FaIMS [CB12]. families [EZ10]. family [CG13b, FE11, HPV13, KTT12, KS13a, Sen13]. fan [DG12]. far [HW11, IK11, LLSW14, Nic11]. far-field [HW11, LLSW14, Nic11]. Farley [DST14]. Fast [BK14a, BLQ14, BS11a, BRS13, BSB14, BBPC14, DLC11, EHA14, EMS11, EFT13, EAB^{+}12, FO11, GM10b, GG13, JC13, JHF10, JFC^{+}13, JMKK12, KQ11, LSK11, LLL10, LLY11, LvZB10, LQB14, MCLG10, MSD12, PG11, QY10, RGMC13, SBWF14, Tyg10, WD14, AAI10, AD14, ABS^{+}14, BWC13, BKH12, Bre12a, BLR14, CF11a, CB12, CCK13, Chi12, CC12, CKD10, CH14, DH10, DV13a, DGM13, DF14, GB13, Gon12, GHL14, HM11, HSN13, HW10, HW11, KM13, LBT11, LWD13a, LWD13b, LGG^{+}13, LT10b, LT11b, LC14, LCB12, Luo13, MCD14a, MST14, Mét11, MY13, NN12, PCF^{+}13, PVDW14, PA14, RE13, SY12a, SY14a, SH11, SV10, Tak14, TiI11, TY11b, VRBZ11, WT12a, WD13a, WWS^{+}13a, ZHPS11, Boy10, CB14, DDL13]. fast-electron [CKD10]. fast-multipole [Tak14]. fast-rotating [DH10]. Faster [SBB10, Gon10]. faucet [YFL^{+}13]. faults [LMQ^{+}11]. FBI [LQ10a]. FBI-transform-based [LQ10a]. FC [AB11, BL10, LB10]. FCC [PLN11]. FD [BMCK10, BMK11, BMK12, BFE12, CPJ11, FH11]. FDF [WPP10]. FDTD [FL13, GV14, LY13b, LBZ^{+}10, MK13, PLN11, Rem11, TD11a, WBC13]. FDV [RCCW10]. FE [LGLX14, SGS13, SPB^{+}10]. FE-FV [SGS13]. Feature [KWSJ11, BRP14]. Feature-driven [KWSJ11]. feature-preserving [BRP14]. features [LEH10]. featuring [LMQ^{+}11]. FEM [WKL^{+}14, AM12, BNP10, BSS11, CES14, DSHP11, HVD13, Hys12, LHH14, NOT14, NvdVB13, YCY11]. FEM/BIE [HVD13]. femur [AKL^{+}13]. Fermi [CB11, Liu13]. FETI [BH13b]. FFT [CP14, ES14, PVDW14, Vol14, WLL13, ZVNM10]. FFT-based [Vol14, ZVNM10]. Fibonacci [Rem11]. Fickian [STD^{+}10]. Fictitious [CCFCM11, AF13, BK12a, BL11, GLLP14, HS13a, LB12, NB14, YBK13]. fidelity [AME14]. Field [CGM^{+}12, AH10a, BR10a, BHL^{+}13, BKH12, BLR13a, CNR10, CL11b, DS12, Don14b, DFW^{+}10, EAB^{+}12, FHMS11, GW12, GW14, GH11b, GRS14, GWG14, GMG12, GMXG14, HD13, HW11, HH13, HLLW11, IK11, JR14, KED^{+}12, LXY11, LLSW14, LMBB11, LZ10, LVZK14, Nic11, PRA13, PB14a, RPA13, SMP^{+}13, SSC13, TNK10, TK14, YJK12, ZMQ13, ZYF^{+}10, Zho14b, ZWJW10]. field-lattice [SSC13]. field-wise [BLR13a]. fields [AETT11, BG10, BRSS11, BRS13, Brü13, CRK12, Far11, HW14c, KST13, KWS^{+}11, KSK13, LTC10, LTC13, LBS10, Mar10a, MTD14, PSDF13, STD14, SSR14, ZM10]. filaments [TLZ^{+}11, YFL^{+}13]. Filler [Ano12-32, Ano12-34, Ano12-35, Ano12-31, Ano12-33]. filling [MWP14]. film [BWS10, MB11]. Filon [WLW13a, WW14]. Filon-type [WLW13a, WW14]. filter [HMM14, LT14, LM13, Ran12b, WYS^{+}11]. Filtered [AGSG11, MH10, ZCS14]. Filtering [BGM12, GM11a, BH13a, DEH11, DV13b, Den11a, GHM10a, GHM10b, Har11, JMV^{+}12, SCIE14, Sti12]. filters [BGM12, BH13c, Fal13, GLM14, Kim13, MS12, MTAC12]. find [Boy13, LD12]. finding [And10]. fine [PJCS14]. fine-coarse [PJCS14]. finger [HZ11b]. fingering [SGC13, SHCY13]. Finite [BGGGS11, BWS10, BdCB11, CH10, CRT14, CR13, CC14, DKM11, FGE11, KLL13, LJX^{+}14, LRBT14, LW13b, Pri14, RCCW10, SGW10, SYV17, SLC^{+}11, SS13b, SS14b, TL13, WLW14, WT14, ZM10, nZzSlL14, vEKdB14, AB12, AINF13, ANIF14, AV10, Ain14, AIX10, ASS13, ASS17, ABK12, ABK13, AKL^{+}13, ABS^{+}14, ABJ^{+}12, APC11, BCP13, BPM13, BFR13, BHL12a, BSSW12, BGN10, BDMP13, BHL^{+}13, BW13, BN12, BN13, BN14a, BFSVC14, Boe11, BKC^{+}11, BFE12, BSB14, BD14b, BBD14, BCLR10, BHL12b, BHM14, BHY14, BTY14, BKV12, CCKCG10, CVC12, CMS10a, CCW12, CHKT13, CLSX14, CXLF14, CLG13, CW14b, CDL11, CWZ^{+}11, CH11a, CDHM11, CVI^{+}13, CMG^{+}13, CJ14, CH11b, CS12c, CT14, CR14, CMBW13, DH10, DMHL14, DEH11, DM13, DMT^{+}11, DMM^{+}13, DKR14, DB13, DA14, DB12, DCL11, DSZ13, DKA12, DCVM11, DZHB13, DZLD14]. finite [EGW11, EGH13, EGL^{+}13, ES10, EO11, EAN11, ES14, FGZ12, FPN^{+}13, FGR12, FH11, FCG10, FCYF11, FCN^{+}13, FC13, FLB^{+}12, FL11b, FE11, FO11, FSL10, FMM^{+}10, GT12, hGzS11, hGzSnZ12, Gao13, hGzS13, GDHM11, GMA10, Gri13, GZZ14, GDGP11, GQ13, GXS14, HRK^{+}10, HJ11, HKJ11, HPA14a, HPA14b, HR10, HPV13, HGW11, HRBK10, HE14, HMM^{+}10, Hic12, HZ14, HHK13, HZS12, HvB11, HM10, HJK14, HL12b, HY09, HY11, HZ11b, Hu13, HLLW11, HLYS11, HAQ12, Hua14, Hut14, Ian11, IX10, IDNG13, JMW12, JZ11, JRG13, JM10, JTXZ12, KCG14, KR10, KDS14, Kim13, Kol11, KEGM10, KVBP^{+}14, KO14, LSM14, LRL10, LCNK10, LB13, LdWF12, Leu11, LH10c, Li10a, LHY12, LR12, LXL13, LHY13, LX14, LSZ13, LLMY14, LSV10, LMS11, LMBB11, LMS14, LS13d, LW14, LLS11, LTL11, MZ11b, MAPP13, MF14]. finite [MB12, MMV^{+}13a, MMV^{+}13b, MMSI13, ML14a, Mat14, MOV12, MTB10, Mig14, MSS12, ML13, MVCFM13, MSTL13, MSTL14, MTS^{+}11, MMT14, Mor10, MK14b, MIKG12, MNL^{+}13, MWY14, NLGL14, Nis14a, Nis14b, NPV14, NKS10, NNS11, Nor14, ORHH13, OBT11, OS11b, PDSS10, PMM^{+}11, PR11, PvDtTB^{+}11, PNPF10, PS14b, PBC11b, QS11a, RSK13, RPV13, RSS10, RKGM14, RHGT10, RBK13, SSB13, SF11, SL13a, SRSV11, SCGE13, SQC11, SZ11, SY11a, SY12b, SS10c, SG14, SY11c, SYV14, Sog14b, SK12b, SBvdV11, Son11, SJH12, SC12, SR13d, SIT^{+}11, SRL^{+}11, SDH12, sSLxRyZ14, SIDG13, TH14a, TD11a, TCL11, TEY12, UKP14, UJvL10, UJ11, UJ12, VDP11, VWPF11, VWF13, VSK10, WA12, WWS10, WW11, WSYS12, WD13a, WKMJ13, WZZZ14, WD14, WAM14, WZ10, WYC13, WG14, XIIX14, XQX13, XQXC14, XLD^{+}11, XH13, XJ14, YVL^{+}11]. finite [YS13, ZW10a, ZS12, ZYHW13, ZHY13, ZY13, ZKM^{+}11, AGBC14, ABMT14, CRK12, Den11a, GP12a, LHMZ10, SGW10]. Finite-Difference [SYV17, CR13, vEKdB14, ANIF14, BDMP13, BHL^{+}13, BHM14, CWZ^{+}11, CVI^{+}13, CR14, DA14, FCYF11, FCN^{+}13, Gri13, GZZ14, HE14, Hic12, HZ14, HJK14, JM10, Kim13, Kol11, LS13d, OBT11, PNPF10, SYV14, TH14a, ZY13]. finite-differences [BFE12]. finite-element [GT12, HvB11, KEGM10, LHY12, MIKG12, MNL^{+}13, PMM^{+}11, RSS10, WYC13]. finite-rate [MMSI13]. finite-size [RPV13]. finite-spectral [WA12]. Finite-Volume [KLL13, CC14, CDHM11, CJ14, DCL11, FE11, IDNG13, MAPP13, Nis14a, Nis14b, NNS11, Nor14, PDSS10, PNPF10, PP14, SY11c, SR13d, SIDG13, UJvL10, UJ11, UJ12, VWPF11, VWF13]. finite/boundary [LRL10]. First [ABL^{+}11, Nis14a, Nis14b, BBAP14, CH10, DBZ14, JRTY11, KVBP^{+}14, LPE^{+}11, Nis10, VK12, DBO^{+}10, MSS^{+}14]. First- [Nis14a]. First-order [ABL^{+}11, KVBP^{+}14, LPE^{+}11, Nis10, VK12]. First-Passage [DBO^{+}10, MSS^{+}14]. FIS [DR11]. FIS- [DR11]. fish [BI11]. fish-like [BI11]. fitted [HSWZ13, NHE14]. fitting [HLW13, PB11, PPWZ11, RZ10]. five [Boy11, FK14, KL10b, KK10c, MDTC13]. five-equation [FK14, KL10b, KK10c]. five-point [MDTC13]. FIVER [FGR12]. fix [Rie11]. fixed [AR14, BHM11, BHM13, BOD11, RF14b]. fixed-source [RF14b]. flamelet [ISZ12]. flamelet-based [ISZ12]. flames [Par10]. flash [PM14]. flashing [LKP14]. Flatland [BM10a]. flexibility [BBC^{+}12]. Flexible [HL11, Cha14, CPX13, CNST13, FRP14, GRS14, HCS11, NL12a]. flexural [FGE11]. flight [NL12a, XW14]. flights [JWNL11]. floating [CKS14, LLdSC11]. flooding [KCG14]. floods [MGN12b]. Flow [DRW11, APF13, ALR11, AC13, ABL^{+}11, ABKF11, AP11, AP12b, Ala14, AST12, ACCM12, AT13b, APC11, BVM14, BHS13, BHS14a, BHS14b, BCHM14, BC10a, BW13, BGGW13, BB13, BA10, Bea12, BPQO13, BERB^{+}12, BDH13, BSS14, BdCB11, BS10b, BOK12, BM14, BCG^{+}13, CD10, CES14, CEP12, CJLC11, CFKL11, CZ13, CSW14, CY14, CS14a, CEL14, CH14, CdCD12, DKC14, EMS11, EWH13, EWH14, EQYF13, Era13, FMT11b, GB14, GLSC12, GDW11, GW10, GKF11, Gri12b, GK14, HS13a, HSD11, HSK^{+}15, HMHO13, HAD10, HK11, HY09, HY10, HY11, HLLW11, HCS12, Ion13, JK10, JMGN13, JK11, KVM12, KS11a, KST10, KR10, KWS^{+}11, KM11, KS11c, KSR^{+}14, KK10c, KG13c, KL11b, KG12b, Kwo14, LW11, LP13b, LDM14, LH10b, LBCL13, LQ10a, LXF10, LLX11a, LL12a, LT12, LMZ14, LLMY14, LLP10, LW13b, LDE^{+}13]. flow [LYXZ14, LvZB10, LBR14, MZ11b, MWP14, MB10a, MM10, MB11, MRS14b, MKB^{+}12, MDM^{+}12, MMT14, MF13, MJM13, MPT13, MMJ14, MT14b, MGN10a, MGN11, MGN12a, MGN14b, NL10, NS14a, NJ12, OP10, PSL20, PZTW12, PZW14, PZG14, PBWK11, PBM11, PMSH10, PC13, PPWZ11, QS11a, RBB12, RG13, RS10a, RXLS14, RHGT10, dCRCS10, RM14, SBB13, SS13a, SGS13, SSHT14, SSS10, SHFB12, SM11b, SSG11, SHM12, Shy10, SX14, SFWP10, SKN13b, SHA11, SHA12, SSL13, SL13b, SSY13, Tit12, TT10, TL13, UKP14, Unf13, UGF14, VMS12, VT13, Vre11, WPKK12, WFL10, WS10b, WSS12, WK13a, WHT14, WLW14, Wil13, WTSZ10, WC14b, Xu11, XW14, YMLM^{+}14, Zad11, ZM14, ZS14, ZYHW13, ZHY13, ZHS10, ZLH12, dPSS12]. flow-induced [SM11b]. flow/acoustics [KVM12]. flowfield [SZ11]. flowfields [RCCW10]. flowing [ZIOF10]. Flows [LLdSC11, ABMT12, ABMT14, AF13, ABPSM11, ABS^{+}14, BGGGS11, Bal10, Bal12b, BMD^{+}13, BDA14, BNM14, BDA12, BFSVC14, BCD14, BPS14, BR11, BLR13b, BD11b, BBD11, Bre12b, BDKV14, CIJS14, CCKCG10, CD13a, CFCA13a, CFCA13b, CS12a, CK10, CP14, CM14, CP10, CL14b, CS10, CB10, CZ11b, CVIA10, CL11b, CJP^{+}14, CRT10, CSKP10, CC14, CCLV10, CJN^{+}13, CGC^{+}12, CGF^{+}13, CDB13, CTSM13, DLDL10, DM13, Del11, DLZ^{+}11, DFVY14, DFNNRdlA14, DF14, DMMGM14, DS10, DS12, Don14a, Don14b, DL10b, DLM13b, EGG12, FGR12, FF13, FK14, FSL10, Fus13, GT12, GG10, GPCE13, GCE14, GJ14b, GITW12, GJLY13, GVV13, GCL13, GK10, GIQ11, GDGP11, GH12b, GLL14, HRK^{+}10, HLT10a, HHA14, HKM12, Hig13, HF11, HHH13, HK13, HKL14, IAM^{+}12, IG14, JTT14, JYHT10, JH12, KKAS12, KWSJ11, KKL10]. flows [KH13, KSL10, KM13, KF12, KT14, KG13a, KKO13, KRT14, KSD12, KBB13, LP13b, LNS13, LMFH11, Lee12a, LSBJ14, LGC13, Li10b, sLwG10, LZ13, Li13, LMGS13, LXSR12, LHB10, LFRM10, LWCR10, LWF11, LZV12, LVZK14, LDDBR13, LHPH11, LBK10, LXS^{+}13, LI14, MB13a, MS10a, MAB^{+}13, MMV^{+}13b, MCA^{+}13, McD14b, MGN14a, MZ11d, MWRZ13, MDH^{+}10, ML11b, Min13, MBD13, MW14a, MSB14, MLBS14, Mor10, MN14, MLGN12, MGN12b, NHE14, NL12b, NCKN10, NMT^{+}12, OAK11, OD13, PLBR13, PL14, PPS12, PP14, PB14b, QLW14, Qua11, RK14, RC12, RPC14, RB13, RBCK10, SRN10, SM11a, SBB13, SHR13, SF11, SHCY13, STH11, SFCF10, SYJ11, SSHM10, SS10c, SPF10, Shu14, SW10, SC13, SCB10, SBvdV11, SS13b, SS14b, SM10, SKR14, SST10, STC10b, Sza12, TS11, TQW14, TSL^{+}14, TSLV11, TK12]. flows [TK15, TBMH11, TRL14, TVEC12, TAD14, Tyl14, ULS13, VDD10a, VP14, VRBZ11, VBCM10, VSC^{+}11, VdMS^{+}14, VED13, VKBK11, VWP11, Wan13a, WK10, WYS^{+}11, WZ11, WWZ13, WZG^{+}13, WP10, WS10c, WS11, XQX13, XH10, XYL12, XR14, YDN10, YP13, YCY11, YS11a, YS11b, YCLK12, YSW^{+}13, YSW14, Yin14, Yok13, YYB12, ZA14, ZHW10, ZL10, ZFJ11, ZN14a, ZN14b, ZGSZ10, Zho11, ZF14, ZDSP11, vRLPK11, vVVK10, vdVR12a, vdVR12b]. Fluctuating [Kea14, SBA13, UPDB13]. fluctuation [NWLS12, XM14]. fluctuation-enhanced [XM14]. fluctuations [Atz10, Atz11, LHB10, PHSA14]. Fluid [BLR13a, BCG^{+}13, CCL^{+}14, DD12, LS13a, TDL^{+}14, AC13, AELV14, AF13, ACCM12, Atz11, BK12a, BHS12, BC10a, BOD11, BW13, BBGP13, BS10b, CFCA13a, CFCA13b, CNR10, CW14a, CS10, CLG13, CZ11b, CHHL11, CM10, Dad12, DDM10, Deg11, DM13, DF14, DLM^{+}13a, DL10b, DGF14, EC13, ETW10, Fal13, FL14, FRP14, FG12, FP14, GM12a, GCBH13, GJLY13, GVV13, GKF11, GZZ14, GBC11, HLT10a, HSK^{+}15, HMHO13, Hig13, HK13, HCS11, JSA14, JTH10, KFOS14, KMA13, KA12, KK10c, LT10a, LP13a, LZ13, LC13, LWF11, LW13b, LJG14, LHPH11, LGH10b, LZZR10, MD13, MF14, MK14a, MNKS13, Mit13, MMH12, NL12a, NPV14, NKF14, OL13, OP10, OD14, OM12, PPTMDK13, PMSH10, PHSA14, PZ10, QLW14, RBB12, Ric13, RHGT10, RMSF11, RDHK12, STH11, SC10, SC11, STL^{+}10, SBvdV11, SKt10, SIT^{+}11, TGT^{+}10]. fluid [TQW14, TPF11, TVEC12, TP14, ULS13, VDD10a, VRB10, VLM13, VF10, WFL10, WYS12a, WYS12b, WKL^{+}14, WLW14, WZT^{+}14, WY11, WTSZ10, WC14b, XL11, YS11b, YS12a, YBK13, Zad11, ZA14, ZBS12a, ZF12, ZL10, ZJL^{+}14, HS13b, MHL^{+}14, WY10]. fluid-body [WY11]. fluid-deformable [WKL^{+}14]. fluid-flexible [HCS11]. Fluid-Kinetic [MHL^{+}14]. fluid-particle [LWF11, SC11]. fluid-solid [CHHL11, ETW10, MNKS13, WZT^{+}14]. Fluid-structure [BCG^{+}13, LS13a, TDL^{+}14, Atz11, BK12a, BW13, BBGP13, Deg11, EC13, FL14, FP14, GCBH13, GKF11, LJG14, MF14, MMH12, NL12a, NPV14, OM12, PHSA14, SIT^{+}11, VRB10, WC14b, YS12a, YBK13, ZF12]. fluid-structure-interaction [Ric13]. fluid-surfactant [GZZ14]. fluid/solid [GM12a]. fluids [GF13, Hys12, KGG10, KL10b, LRT13, LGE^{+}13, PRA13, PS14a, PS15, UWB12, YFL^{+}13, ZYF^{+}10, dv23]. fluorescence [MB13b]. fluorophores [MB13b]. flux [APY14, BNT14, CVC12, Cha13, DFNNRdlA14, EZ10, Gao13, GMC13, GD10, GL12a, HLM11, KS13a, Kri14, KMSS10, LBM10, LSVW10, ML14b, OS11b, Par13, PM14, SBN12, SK13, VKL^{+}13, VCJ11, WSW13, WCVJ13, XQX13, ZS14]. flux-approximation [ML14b]. flux-corrected [LBM10, LSVW10, OS11b, VKL^{+}13]. flux-form [BNT14, HLM11]. fluxes [Cam13, PAM10]. FMM [GD13]. foam [KLP10, KLPS14]. focused [LX13b]. focusing [BRS11, MdVBS13]. Fokas [AN14]. Fokker [TTR^{+}12, TBNT11, Wol12, CCT11, GJ14b, JY11, LE13, Mar10b, PLT14, SJH12]. Foldy [HSZ10, HLZ13]. Force [SCS14, CN10, Dai13, HH12b, HH13, Kea14, LKCY11, LY13a, MWRZ13, MW14a, PBT14, PA12, YTYJ12, YM10, Yok14, vHtTBC11, SST10]. Force-controlled [SCS14]. force-coupling [Kea14, YM10]. force-driven [MWRZ13]. forced [CL13, TPA^{+}13]. forces [AT13b, HH13, MD13, SSHT14]. forcing [BL11, CLS10, GH10, Hig13, JMW12, KF12, LDS11, OBT11, VRB10, YS12a]. forecasting [Fal13]. form [ANIF14, BH12, BNT14, BN12, FCN^{+}13, HLM11, JM12, Mor10, MK14b, MV13, RCCW10, SBS11]. format [BGK12]. formation [DH13, SS13b, SS14b, Tok10]. forming [AVKP11]. forms [BBM14, PRH^{+}14, PZ14, RMMD14, YMS12, ZPF10]. formula [hGzSwZ14]. formulae [DFD14, NN12]. formulas [BMCK10]. Formulation [BRSS11, DEKBF14, RS13b, APF13, AHA10b, AHA13, AGBC14, AIX10, BLRR11, BLA^{+}14, BD12, CN12, CVN13, CFW^{+}13, CB14, CY12, CBW10, CF11c, CGP10, ES13a, ES13b, FL14, Fus13, GW13a, GIQ11, HSWZ13, HSZ10, HLZ13, JKQ13, KFOS14, KK10c, LRL10, LKT^{+}12, LYD10, LHM14, LBS10, MBMV13, NOT14, Nis12, OR10, OLC13, QS11b, Ric13, RMSF11, SHMS13, SZF12, SPB^{+}10, SYL10, SCB10, SB12, SP14b, TY11a, VMA14, Wan10, WZ11, XIIX14, Yok14, ZWL^{+}12a, BD10a]. formulations [BCLR10, FCN^{+}13, GCW14, MHA10, WX14, ZRM13, ZZB14]. Forward [Wan13b, HVT12, LP13a, MB10b, XW14]. forward-reverse [HVT12]. four [YJK12]. four-field [YJK12]. Fourier [AR14, APRS14, BPM13, BLTO11, BY11, CK10, FW12, GBS^{+}14, GZQ13, HW10, HDD12, HG10, JHF10, KSK13, LR11, PMSM12, PS14b, QZM^{+}12, SABH11, SHZ13, SJH12, ZHPS11]. Fourier-based [CK10, GBS^{+}14]. Fourier-continuation [SABH11]. Fourier-continuation/WENO [SABH11]. Fourier-series-based [ZHPS11]. Fourier-Wavelet [KSK13]. Fourth [WGX13, BYCC14, CdCD12, DAEB13, FCYF11, Ge10, HM10, Li10b, LY13b, Mat14, SVK14, YXZ13]. Fourth-order [WGX13, BYCC14, CdCD12, FCYF11, Ge10, HM10, Li10b, LY13b, SVK14]. FPGA [LWZ^{+}13]. fraction [QB14]. Fractional [ZK13, AKP^{+}12, BCL13, BCL14, BLY10, BHY14, BTY14, CX13, CD13b, ÇD12b, Cui12, FCY13, hGzS11, hGzSnZ12, hGzS13, hGzSwZ14, GMA10, HAMA14, JTT14, JPC^{+}14, JR12, Kat14, Kha13, LS13b, LCY11, LX13a, LXL13, LMGS13, LYJ14, MAD13, MY13, OM12, PS12, PH13, RzSZ13, Sou14, VW14b, WWS10, WW11, WXY13, WD13a, WD13b, WV14, WD14, WXY14, XH14, YZX14, YKWS13a, YKWS13b, ZK14, ZFH14, nZzS11, nZzSlL14, ZzS11]. fractional-step [AKP^{+}12, OM12]. fractions [FS10a]. Fracture [BJLR14, AJV11, BPS14, CWL10, HKJ11, ZYHW13, SBN12]. Fracture-Matrix [SBN12]. fractured [MF13, NJ12, YL10, ZYHW13]. fragmentation [KSU13, SKU13]. frame [DS11, GK14, JSX10, VGCMG11, YXD^{+}14, ZS14, ZJT10]. frame-indifferent [GK14]. frame-invariant [ZS14]. frames [BPQO13, BN14b]. framework [ACG14, BHS14c, BBGP13, CS14c, CT14, DBZ14, GHP13, HQL^{+}10, LLYE12a, LS10b, MRS14b, NMV13, OK12, OD14, SMP^{+}13, SKW14, SS10e, Unf13, VBVD10, VBSK10, WKMJ13, WC14b, XJ12, YS12a, YD12, Yok13]. FRC [KST13]. Free [AM10, BK12b, LH13, ABKF11, AF11, BCD14, BDTW10, BND^{+}12, CTP13, CSW14, CW14a, DDF14, FDHP11, GAvdVB14, GCvRK11, GC11, GCZ13, GL12a, GVV13, Han11, HMHO13, KM11, Kaw13, KG13a, KG13c, LTS^{+}10, LPE^{+}11, LKT^{+}12, Li10b, LXY11, LX12, LSZ13, LXSR12, LW14, LFLV14, LvZB10, LYD10, MCLG10, MIKG12, NCV10, OM12, PMMB14, PZ10, RR14, RB10, RB13, RLPM12, RM14, SAM10, SF11, SR13c, SST10, SBO11, SP14b, WWZ13, WIJ13, WY10, XMS12, YW13, Yok13, YB14, ZBS12a, ZBS12b, ZOWZ10, ZOJR11, ZCS14]. free-boundary [LYD10]. free-surface [ABKF11, BCD14, GVV13, HMHO13, LXSR12, MCLG10, PZ10, RB10, RB13, SF11, SST10, WY10]. freely [GMXG14]. Frensley [JCT11]. frequencies [CB12]. Frequency [HR10, AWK^{+}11, AETT11, AH10b, BCM12, CTG12, DTU12, EL11, GH11a, GGG^{+}13, HdBK12, LJ13b, LQB14, MJK13, MN14, SSS11, SF12, SKt10, TY11b]. frequency-dependent [DTU12]. frequency-domain [AWK^{+}11, SF12]. frequency-space [SSS11]. Friction [SPS14, CVC12, MGN12b]. Friedrichs [CCK13, LQ11]. frog [HT14, SWX12]. Front [SYJ11, AK14, BW13, LW11, MFM12, Mar10b, NW10, RZ10, RR14, TVEC12, UGkM13, WWZ13, Wit10]. front-tracking [LW11, RZ10, RR14, TVEC12, UGkM13]. front-tracking/ghost-fluid [TVEC12]. Frontiers [ASXZ14]. fronts [DLM13b, LEH10]. FSI [BHS13, BHS14a, BHS14b]. fuel [VDD10a]. Full [KED^{+}12, AMP14, FE11, GB14, LLJ10, SIT^{+}11, XHF13]. full-body [XHF13]. fullerene [WPKK12]. Fully [BBGP14, GM10a, GHP13, GF13, Hen13, LDA10, ASB10, AH10a, BZKL13, BCL^{+}11, Bre12b, CDSP13, CNR10, CCL^{+}14, FCG10, GH11a, GMC13, GXX14, HL12b, JZ13, KK10a, Kau10, LGF11, LL10b, MA14, MdVBS13, Mor10, Ric13, SNS10, SHCY13, SPB^{+}10, SFCF10, SCS14, TQW14, WC14b, YB14, ZBS12b, LEH12]. fully-coupled [SHCY13, SPB^{+}10]. fully-developed [SNS10]. fully-discrete [GXX14]. fully-implicit [SPB^{+}10]. fully-nonlinear [SCS14]. Function [MNR11, AGSG11, BFE12, BCM^{+}11, Boy10, BD14c, CRS14, DL13c, FWLP14, Gil13, Gon10, Gür14, HLW13, IST^{+}12, IXX14, JC10, JK11, KNR11, LK14, LSK11, LH10d, MCLG10, NKF14, ST13, TVE14, VK12, WZ11, WYS12a, YSW^{+}13, YSW14, ZT10a, ZMG11]. function-based [YSW^{+}13, YSW14]. Functional [Fat10, LGE^{+}13, BN12, CGS13, FN10, HZ14, ITO^{+}10, JFC^{+}13, LLYE12a, LLYE12b, LSE13, LBRA^{+}13, MM10, MIKG12, MNL^{+}13, SL13a, SSS11, SD14b, SBO11, SP14b, YWY10]. functionally [HBZL14]. functionals [RSS10]. Functions [PDC13, BY11, Boy11, BX12, CEK14, CCC12, CST10, EGW11, ERT10, Fat10, FL10, FB11, Gno14, GWG14, HBM13, HHGC14, JNX13, LD12, LB14, MK13, Min10a, NB14, Ost10, Pir12, PH13, RA10a, RPA13, Sal14, SPM^{+}13, TCL11, Val14, VT13, WZZ14, ZBS12a, ZBS12b, ZJW12, ZPGO10, MPR12]. fundamental [AEM13, BOK12, SYL12]. Further [DMM^{+}13]. fusion [BS14a, PLT14, TTR^{+}12, WCM^{+}14]. FV [SGS13, ZWL^{+}12a, ZWL^{+}12b].

Galerkin [BR12b, FW12, AJRT11, AJ14, AWK^{+}11, AAT^{+}14, BNT14, BGR11, BBC^{+}12, BR11, BGHM13, BSCML13, BTY14, Cha13, CZ11a, CS12b, CAH14, CLX14, CCZ14a, CCZ14b, CW14b, CSX14, CCY13, CJ14, CDB13, CdCD12, DAEB13, DFFL10, DHH11, DL10a, DZL13, DSZ13, DZLD14, FPN^{+}13, FT12, FH11, FP12, FP14, GAvdVB14, GH11a, GDHM11, GP12b, GZQ13, GXX14, HGMM12, Hig13, HvB11, Ian11, JMFO13, JZ11, KG12a, KKO13, KG14, KRT14, Kuz14, LH10a, LXM14, LHH14, LLJ10, LXY11, LX12, LXL13, LH14a, LZZC14, LLP14, LSW14, LLYE12a, LLMY14, LLH10, LP14b, Liv10, LLN^{+}10, LXL^{+}12, LXS^{+}13, LI14, MRS14a, MHA10, MN11, MOSW12, MST14, MESV10, MMD^{+}11, MMIW14, ML11b, MK11, MWW^{+}13, MWY14, MBGW13, NPC11b, NPC11a, NPC11c, NP12, NDBG14, NvdVB13, OHF12, OS11a, OD13, PDZ^{+}14, Per13, PIN14, QS11b, RV10]. Galerkin [RHXQ14, RGMC13, RC12, RCvdV13, RS11, SH14, SGC13, SHCY13, SLT^{+}14, SBvdV11, SSL13, SD14c, TG14a, TE11, TD11b, TD12b, TLNE10, TBR13, VMA14, VR14, WXZ10, WS10b, WT12a, WZSN12, WK13a, WKH10, WSBG10, WT14, XXS10, Xin14, XLD^{+}11, XCL14, YO11, YWS13, YMLM^{+}14, YHL13, YWL14, ZS10b, ZOWZ10, ZS11, ZOJR11, ZHY13, ZZS13, ZT13b, ZYS14, ZS13, ZT10b, ZZSQ13, ZMD^{+}11, vdVR12a, vdVR12b]. Galerkin-enriched [CJ14]. Galilean [Del14, WZT^{+}14]. gamma [RKD^{+}14]. gamma-ray [RKD^{+}14]. Gas [CJLC11, BRS12, BSDM12, BDLGC13, CXLC12, CGF^{+}13, DLDL10, GJ14b, GH12b, HMHO13, JZ11, KM11, Kol11, KGK13, KMSS10, LNS13, LXF10, LLX11a, LYXZ14, MZ11d, Mie13, Min10b, MHGM14, Nag14a, Nag14b, PZTW12, PLBR13, PL14, PL12, RTS13, RPC14, SS10c, Sza12, TK12, TK15, Tit12, TA13, UN13, VMA14, WZG^{+}13, WYT14, XH10, XX13, YSW^{+}13, YSW14, ZFJ11, vVVK10]. gas-dust [Min10b]. Gas-kinetic [CJLC11, KGK13, LXF10, LLX11a, LYXZ14, XH10, XX13, YSW^{+}13, YSW14]. gas-liquid-like [TK12, TK15]. gas-particle [WZG^{+}13]. gas-solid [BSDM12]. gaseous [WZSN12, YTYJ12]. gases [CL14a, KM14b, LXM14, PBC^{+}11a]. gauge [CB11]. Gauss [Boy10, DBAV14, TDL10, SAHS14, ZHS10]. Gaussian [AETT11, BLQ14, BHM11, BZ12, BZKL13, Boy10, HW14c, JWYH10, KSK13, LQ10a, QY10, RS13b, RS14b]. GCL [SYV17, AINF13, SYV14]. Geant4 [KLB^{+}14]. Geant4-DNA [KLB^{+}14]. Gegenbauer [DV13a]. gel [ALGB^{+}13]. GENE [GLB^{+}11]. general [AB11, BDTW10, BDKV14, BL10, DAA11, FE11, GCW14, KK10b, LFW11, LB10, MT14a, PNPF10, SM11a, SY12a, SRS10, VMA14, WYS12a, WG14, ZT13a, vV12]. generalised [LXSR12, Sal14]. generalised-Laguerre [Sal14]. Generalization [LBZ^{+}10, STD14]. generalizations [CHZ13b]. Generalized [AEM13, DXB^{+}13, EGH13, EGL^{+}13, HDHZ13, HSZ10, LBS10, PSM11, Pir10, PC13, RB13, SZ11, YBK13, ZZB14, ARR11, AHA12, CBPS13, DBZ14, Fal13, GvdSVK10, HE14, HHGC14, HLZ13, KJC13, LLX11a, Liu11b, LBR14, MMD^{+}11, MCDT12, OJW10, PBdGP14, Pir11, QLW14, SHZF12, XXS10, YHL13]. generated [ERT10, FLB^{+}12, FL11b, JH12]. generation [GL13, HW14c, Kau10, LX13b, MZ11a, PAG11, SGV13, SWR11, WTYC13, YB14, ZJW12]. generator [NW10]. generic [LEH10]. Gentile [Gho17]. Genuinely [Fal13, BBD14]. geodesic [PB13a, PB14a, Pud11]. geodynamics [ZLH12]. Geometric [ANIF14, DMT^{+}11, AMP14, ACB^{+}13, BNP10, CST10, ES10, HTHG14, MHA10, MN11, MYM12, MNS13, OD14, ST11a, XFC^{+}14, DMM^{+}13]. Geometrical [JRG13, BDM13, CKR14, LLZ11]. geometrically [GRS14, JK11, OS11b]. geometrically-conservative [OS11b]. geometries [CB10, CLS10, GCvRK11, HS13a, HBAP10, KG12b, MMR12, MTG11, PMSM12, PHSA14, SSB13, SM11b, SY11c, UKP14, ZYT13]. Geometry [SPS14, BCM^{+}11, CBW10, CS10, DOR10, DC13, FGLB14, GPY14, LRS10, LFRM10, SHCY13, UKP14]. geophysical [MG14a, TH14a, UJ11]. geophysics [TQW14]. geospherical [SS10e]. GFD [WYYK10]. ghost [CR13, HK13, KK14, KMA13, LY13a, VDD10a, XL11, ZF12, ZL10]. ghost-cell [LY13a]. ghost-fluid [TVEC12]. ghost-point [CR13]. Ginzburg [RSS10, SV13, TAT14]. glass [ARF12, MSTL14]. GLM [MTB10, MT10]. Global [BR10a, CLSX14, Gno14, SG13, CBGK13, DTG^{+}14, EGG12, GCE14, GLB^{+}11, IX10, JNM12, JMV^{+}12, JMGN13, LR13, MS10a, MBS13, Sti12, WH11, GPCE13]. globular [ORHH13]. gluon [AiINT14]. GMsFEM [CEL14, EGH13]. GNAT [CFCA13a, CFCA13b]. goal [BDA12, GBB^{+}13, LDA10, RKM11, WR11]. goal-based [GBB^{+}13]. goal-oriented [BDA12, LDA10, WR11]. Godunov [Del10, DOR10, LT11a, MMP11, MT10, MLBK14, PR14a, QL11, SS10b]. Godunov-like [MLBK14]. Godunov-type [LT11a]. good [CW13, CB10]. Gordon [BB12b, VBSK10, VW14b]. governing [PSM13]. GPU [BD14a, BGZ12, CCB12, CS13, CRK12, DKH11, FTF14, GLC^{+}11, HMOH11, Hut14, IFD^{+}12, JFC^{+}13, JLS^{+}14, KSP13, KEGM10, KO12, KA13, KAG14, MWP14, MdVBS13, NWLS12, NVS13, PMMB14, PAG11, Pri14, QA11, RCQL11, RBCK10, SBB13, Sch13a, STC10b, WK13b, Wei12, XDZ^{+}14]. GPU-accelerated [BD14a, IFD^{+}12, NVS13, STC10b, WK13b]. GPU-based [HMOH11, KO12]. GPUs [SSK^{+}12, APGL12, AJG^{+}13, BFE12, CB11, KAG14, LLL10, NS14a, SBWF14]. Grad [PCF^{+}13]. graded [HBZL14, HG11, MTG11]. gradient [BV14, Cam13, CJP^{+}14, GW12, IF10, LDM14, MS10b, MOV12, Mét11, NRS10, RSS10, SCC14, Wan14, XY14, ZMG11, Fat10, MMM13]. gradient-augmented [LDM14, NRS10]. gradients [HH13, MS11, Pir12, Vol14, ZVNM10]. Graetz [PBdGP14]. graft [GDY10]. grain [JR14]. grained [ALGB^{+}13, LG14, PA14, ZZK^{+}14, ZKKF10]. graining [Ich13, KKPV12]. Gram [CG13a]. granular [BRS12, BM13a, CM14, FR13, JMGN13, SSB^{+}12, VP10, VP14]. granulation [BKW10]. graphene [BHM14]. graphic [KHRD11, HJBB14]. graphical [GHH14, WZ13, WZ14]. graphics [KSPP10, LSK11, LLL10, Wit13, SSK^{+}12]. gravitation [KM14a]. gravitational [BGZ12, YTYJ12]. gravity [GCFJ13, NDV^{+}11, SGC13]. gravity-driven [GCFJ13, SGC13]. grazing [GH14]. greedy [SYL12, YLW10]. Green [PDZ^{+}14, BCL^{+}11, BD14c, CRS14, HAD10, JC10, JK11, KNR11, LGH10a, LK14, LD12, LGLX14, PE10, TVE14]. Green’s [CB14, MK13, PDC13, VT13]. grey [MSTL13, WL11]. Grid [HH12a, PGLK14, AP11, ABMT12, BCL14, Boy11, CP14, CDS13, CCG11, DjY14, EH14, EH15, Gas11, Gri13, HLM11, HG11, HSWZ13, IX10, IKO^{+}14, KED^{+}12, KSU13, Kau10, KSR^{+}14, KMvM^{+}12, LMLC12, LNU10, LLZ11, Li12, LRK11, MK13, MBLM10, MLBK14, MMJ14, NTV14, PVLT10, PB13a, PMSH10, PMF13, PJCS14, PLN11, Pud11, RBS12, RS14a, SS13a, SZF12, SBS11, Shy10, SWR11, SEBG12, TH14a, TQW14, TS12, WDP^{+}14, WY11, WTYC13, XS15]. grid-adaptive [KMvM^{+}12]. grid-based [CDS13, MMJ14]. grid-compatible [MK13]. grid-conforming [KSR^{+}14]. grid-independent [PMF13]. grid-particle [CP14]. grid-staggering [RBS12]. gridded [HW11]. Grids [SYV17, AINF13, ANIF14, ABHS12, BHS12, BCHM14, BOD11, BNM14, BV14, BM14, CCKCG10, CRG13, CLSX14, CCY13, CR13, DMT^{+}11, DMM^{+}13, DCL11, EZ10, EQYF13, Era13, EN14a, EN14b, EAB^{+}12, FGR12, FB11, FS10a, FE11, GC11, GCZ13, GH12a, GP12a, HRBK10, HW10, HM10, HLT10b, iIT14, IDNG13, JAX11, KVM12, KM13, Lan14, LRLL11, LR12, LB14, LX14, LKTL10, LTT10, LT11c, LLN^{+}10, LXL^{+}12, LXS^{+}13, LZZR10, MB12, MAB^{+}13, MZ11c, MTG11, MG14b, MK14b, NDV^{+}11, OD12, PB11, PRH^{+}14, PHRG13, PYK10, PB13a, PB14a, PAM10, RWX11, RTKS10, SKU13, SGS13, SSS10, SML13, SSHM10, SC11, SFWP10, SR13c, SYV14, Son11, SV10, TF10, TDVG13, TRL14, TD11b, TLO^{+}14, Tyl14, UJ11, VMA14, WYS12a, WY10, WG14, XIIX14, XDZ^{+}14, YMS10, Zie11, vV12]. Gross [MWRZ13, LX14, Sal14, TA12]. Ground [LLSW14, AD14, BCW10, BD11a, BCZ13, CCW11, Don11, FT12, Wan14, ZD11a]. group [AK12, CMO14, KBRM12, LB14]. grouping [Ran11]. growing [GMXG14]. growth [BGN10, CCFCM11, JR14, KBTD12, LGH10b]. GRP [HLT10a, LZ13, WYT14, YHT11, YT12]. Grüneisen [WP10]. GSM [WKL^{+}14]. GSM/ALE [WKL^{+}14]. GSPH [PR14a]. Guest [VCS14]. guide [FLB^{+}12]. gyrokinetic [BKH12, CB13, CRS12, GLB^{+}11, HdCNT12, JMV^{+}12, KWH10, KHJ^{+}13]. gyrokinetics [NHT^{+}10, NHT^{+}13].

H [TP14]. H-VLPL [TP14]. Hadwiger [CLW10]. haemodynamic [GCBH10]. haemodynamics [FQV13, NPV14]. Hagedorn [Zho14b]. half [CB14, Tyl14]. half-space [CB14]. half-staggered [Tyl14]. Hall [DTG^{+}14]. halo [BS14a]. Hamilton [CW14b, LP14b, YO11, ZQ13]. Hamiltonian [AEM13, DF10b, GCW14, HDHZ13, MNS13, NvdVB13, SWX12, WWS13b, ZTS^{+}10]. Hancock [BD14a]. Handling [ZL10, AHS14]. Hard [IAM^{+}12, De 10, FCH13, GL13]. Hard-sphere [IAM^{+}12, FCH13]. harmonic [DLC11, EAG14, GGG^{+}13, JMFO13, LLP14, MTS^{+}11, Mön13, NPC11b, NB14, PL10, RM12, SG10a, SF14, SP11, TEY12, Tyg10, XMDG13, XMDG14, XJ14, ZMD^{+}11]. harmonics [GS11a, LdWF12, MH10, MM11, Ols12, RARO13, RJG10, Sal14]. Harris [VCS14]. Hartmann [HY11, HY09]. Hartree [BD11a]. hash [BGG12]. Havriliak [CPJ11]. hazard [MPS^{+}14]. HDG [WK13a]. heart [LS13a]. heat [BD12, BDLGC13, CdCNH14, Coo13, CCG11, ETW10, GP12a, HS13a, HK11, JYHT10, KK10a, KKvZB14, KSR^{+}14, LN10a, LCK11, LN10b, MB13a, Mar10b, MM11, NTV14, PGR10, PBdGP14, SD14c, ZYT13]. heat-conducting [KSR^{+}14]. heating [KS13a]. heavy [AiINT14]. heavy-ion [AiINT14]. height [BCM^{+}11, LH10d]. heights [FS10a]. Hele [GXX14]. helical [Tan11]. helically [LP12a]. helicity [OR10]. helicopter [XW14]. Helmholtz [BAG12, CES14, CB12, CF11b, CCW12, CLX14, CC12, CGS13, Coa12, DDL13, FKM11, GD13, HM11, KQ11, KO14, LLL10, LW13a, LM14b, LQB14, NN12, RVbZ10, Sto13, Stu10, TGGT13, VBVD10, VG14, WdHX10, ZT13a, Zha10b]. Helmholtz-type [KO14]. hemodynamics [ADSF13, BC14, CG12, XHF13]. Hermite [CAH14, KBRM12, LXL^{+}12, RBK13, SP14a, YF14, Yin14, ZQ13, vDS12]. Hermite-based [SP14a]. Hermite-discontinuous [CAH14]. Hermitian [And10]. heterogeneity [LLMY14]. Heterogeneous [LR13, AB12, AST12, AT13a, BCL14, CTP13, CL11a, CSX14, DM13, Dur14, EGG12, LMFH11, LH10c, Liu11a, LW13b, MZ11b, MJM13, MMJ14, SKN13a, SGS13, STD^{+}10, TLG10, WT13, WHT14, WLW14, WGD12]. heterostructures [MZB12]. Heuristic [WFL10, Phi14]. HEVI [WLW13b]. hexagonal [Gas11]. hexahedra [BD13]. hexahedral [FGZ12, JLZ13]. hierarchic [MMD^{+}11, SL13a]. Hierarchical [AKP^{+}12, JZ10, Kuz14, Sog14b, ZKM^{+}11, BH13c, HKJ11, HKL10, LLY11, SY14a, SSB^{+}12, XLD^{+}11, LXS^{+}13]. hierarchy [MT13]. High [AQ11, AL10, AME14, AMP14, AF11, AN13, BGH10, BD13, BL10, CYK^{+}14, CCD11b, CJP^{+}14, CDHM11, DLGP13, DKR14, Fan14, FT11, FC13, FMM^{+}10, HRBK10, HOK10, HTHG14, IG14, KM14b, KEGM10, LGLX14, LL14a, LB10, MTT13, MTB10, Mig14, MSB14, NPC11a, NCKN10, OAKR14, PPWZ11, QB14, SHCY13, SFCF10, SYV17, SIDG13, TR11, UJvL10, VQVD11, WSYS12, Whi10, WD11, Wit13, XQXC14, XW14, Zha10b, AB12, AINF13, ANIF14, ALR11, Ain14, AETT11, ABK12, ABK13, AH10b, ABJ^{+}12, BHB11, BMS12, Bal12a, BNT14, BDMP13, BCM12, BSS14, BS10a, BLR13b, BK14b, BKW10, BL13, BLR14, CCKCG10, CX13, Cap11, CDSP13, CTG12, CWL10, CZ11a, CXLF14, CGMQ14, CEL14, CDL11, CLAL14, DEH11, DMT^{+}11, DMM^{+}13, DL13b, DAEB13, DSZ13, DWZ10, DV14, EGW11, EGG12, EL11, FW12, FK10]. high [FP14, Fuj11, GH11a, Gao13, GNGAS10, GBNS14, Ger12, GMC13, GPCE13, GCE14, GCL13, GH12b, GCH14, GXS14, HRK^{+}10, HH12a, HRCW13, HE14, HY09, HY11, HLT10b, HAS13, HMOH11, IX10, IKO^{+}14, JNSA12, JAX11, JZ11, JLB^{+}10, KLMJ12, KSU13, KF13a, Kaw13, KKvZB14, KG13a, KG13b, Kwo14, LTC10, LFL11, LP13b, Le 14a, LSBJ14, LGC13, Ler13, LLZ11, LXF10, LR12, LLZ14, LP14a, LN10b, LM14a, LLS11, LL14b, LQ11, LQB14, MZ10, MKGV12, MN11, MIJ10, MZ11d, MZ14, MJK13, MAD13, MPT13, NHE14, NMCZ14, NPC11c, Nic10, NMT^{+}12, OS11a, PDSS10, Par10, PMM^{+}11, Par12a, PGLT10, PVLT10, PCF^{+}13, Per13, QLW14, QC10, QS11a, RZ10, RKD^{+}14, RS11, SRN10, SKU13, SRSV11, SM11b, SQC11, SZH12, SCY11, SR11, Shy10, SYV14, Sou14, SPM^{+}13, SDH12, SD14c, SY14b, TS11, TWSN12, TK12, TK15]. high [Tit12, TY11b, Tyl14, TBNT11, UKP14, UJ12, Val14, VLM13, VWPF11, VWF13, VCJ11, VSK10, Wan11a, WMI10, WYS^{+}11, WZSN12, WW13, WZN14, WSBG10, WP14, XPO14, XDZ^{+}14, YL11, YHM12, YYB12, ZT13a, ZS10a, ZS10b, ZS11, ZS12, ZD11b, ZDSP11, vL11, vRLPK11]. High-accuracy [KM14b]. high-contrast [CEL14, EGW11, EGG12, GPCE13, GCE14]. high-convergence [SPM^{+}13]. high-dimensional [JAX11, LLZ14, MZ10, Val14, WZN14]. High-fidelity [AME14]. high-frequency [AH10b, MJK13, TY11b]. high-intensity [RKD^{+}14]. High-order [AL10, AF11, AN13, BGH10, BD13, BL10, CYK^{+}14, CDHM11, FC13, HOK10, IG14, KEGM10, LB10, MTB10, Mig14, MSB14, NPC11a, PPWZ11, SHCY13, SFCF10, SIDG13, UJvL10, VQVD11, Whi10, AB12, AINF13, ANIF14, ABK13, ABJ^{+}12, BHB11, BMS12, BNT14, BS10a, BL13, BLR14, Cap11, CDL11, DEH11, DMT^{+}11, DMM^{+}13, DSZ13, DWZ10, FP14, GNGAS10, GBNS14, Ger12, GMC13, GCL13, HH12a, HE14, HLT10b, HAS13, JNSA12, JZ11, Kaw13, LFL11, LSBJ14, Ler13, LXF10, LN10b, LM14a, LLS11, LQ11, MKGV12, MN11, MIJ10, MZ11d, MZ14, MAD13, MPT13, NMCZ14, NPC11c, NMT^{+}12, OS11a, PDSS10, PMM^{+}11, PGLT10, PVLT10, PCF^{+}13, Per13, RZ10, RS11, SRN10, SM11b, SCY11, SDH12, SD14c, TK12, TK15, Tit12, Tyl14, UKP14, UJ12, VWPF11, VWF13, VCJ11, VSK10, Wan11a, WYS^{+}11, WSBG10, WP14]. high-order [XDZ^{+}14, YL11, YHM12, YYB12, ZD11b, ZDSP11]. High-performance [FT11, OAKR14, WD11, Wit13, Fuj11, HMOH11]. high-precision [IKO^{+}14]. High-resolution [LL14a, JLB^{+}10, Par12a, Shy10, WD11, vL11]. high-speed [NHE14]. high-velocity [BDMP13, CWL10]. High-volume [QB14]. Higher [Dai13, DCC^{+}13, MF13, MIKG12, MNL^{+}13, SGS13, BX12, FL13, FVR13, Gen13, HAMA14, HdCNT12, Lee14b, LT14, MF10, Mön13, SK12b, SWR11, ZF12, vdVR12a, vdVR12b]. Higher-order [DCC^{+}13, MF13, MIKG12, MNL^{+}13, FL13, Gen13, HAMA14, LT14, MF10, Mön13, SK12b, ZF12]. Highly [ARR13, CLM13, DLNN12, EZ10, GHM14, HT14, LTCN13, LEHA10, LWS14, MN12, PVDW14, SS11, WFK11, WLW13a, WW14, WJC13]. Hilliard [GXX14, GLL14, HHK13, BSSW12, BSB14, GLWW14, GGT13, HGW11, LDE^{+}13, SJK11, WG11, ZW10a, ZTG^{+}13]. Hilliard/Navier [HHK13]. histogramming [LSK11]. historical [vL11]. History [CZL10, Dai13, vHtTBC11]. HLL [BBD14, Cap11, Guy13, MGN12a, Rie10, SLH^{+}11]. HLLC [Bal12b, BDA14, KB14, MGN12a, TT10]. HLLC-type [TT10]. HLLD [OLPM14]. HLLE [Bal10]. HOC [GC11, GCZ13]. Hodgkin [AN13]. Hoekstra [XS15]. Holm [SCY11]. homo [JJM13]. homo-oligomerization [JJM13]. homogeneity [iIT14]. homogeneous [AJ14, CSKP10, LE12, OTV13, STH11, SVS11, She13, TG14b]. homogenization [AB12, ZVNM10]. homologous [EHA14]. homotopy [HHS^{+}13]. Hoover [CWC11]. horizontal [EEK14]. Horizontally [WLW13b]. host [BRS12]. hot [CM10]. hp [Alm14, BSS11, DSHP11, LSW14, WZZ14, vdVR12a, vdVR12b]. hp-Cloud [Alm14]. hp-convergence [WZZ14]. hp-discontinuous [LSW14]. hp-FEM [BSS11, DSHP11]. hp-Multigrid [vdVR12a, vdVR12b]. HPC [SF14].

Hugoniot [Guy13]. human [AKL^{+}13, SMF^{+}13]. Huxley [AN13]. Huygens [Bér11, LQB14]. Hybrid [CBPS13, CP14, EWH14, GQ13, HDD14, LBC14, LQ10b, MM11, PLBR13, SBA13, WS14, ALR11, AP11, AP12b, AHS14, BDL11, BW13, BFSVC14, BBG^{+}11, BS10b, BLR13a, CDSP13, CCB12, CAH14, CPCU13, DTU12, FT12, GG13, GHM12, GWK10, HPOM10, KVM12, KK10b, Kha13, KS12, KSB14, LRL10, LEH10, LEH12, LDDBR13, MIJ10, Min10b, MSTL13, NW10, OK12, PP14, PAM10, RKGM14, RRCD14, RCCC14, SP13, SGW10, SM13, SS10b, SABH11, SHZ13, SMS14, SAA14, SGV13, sSLxRyZ14, TAT14, TP14, VLM13, WWX^{+}10, WSS12, WP10, WYC13, XJ12, YK12, ZWL^{+}12a, ZWL^{+}12b, ZDSP11, HPV13, TL13]. hybrid-kinetic [KSB14]. Hybrid-Mixed [HPV13]. Hybridizable [NPC11b, NP12, WK13a, CLX14, LLP14, NPC11a, NPC11c, RC12]. hydro [KMvM^{+}12]. hydrocarbon [BGBR12]. hydrocodes [LS10b]. Hydrodynamic [SBC^{+}14, AP11, AP12b, ÁCB12, CS12a, HP13, HZS12, LBZ^{+}10, MS10a, QY12, RUL^{+}13, SB12, TG14a, TLZ^{+}11, VKBK11]. hydrodynamic-kinetic [AP11, AP12b]. hydrodynamics [AHA12, AHA13, AiINT14, Bal12a, BLOdV14, BHS14c, BAB^{+}13, BLA^{+}14, BPHK13, CL10, DH13, GM10a, JGZL14, KK10a, KKLR10, KBS14, LXSR12, MKB^{+}13, MLBK14, PTM13, PBT14, Pri12, Ros10, Sco12, SS10b, SF14, SR13c, TiI11, TAAS13, UPDB13, VKL^{+}13, WT14, YHT11, YT12, Zha10a, ZT13b]. hydroplaning [VSC^{+}11]. hydrostatic [ABPSM11, BF12, NNS11, UJ12]. Hyperbolic [MMT14, AK12, AAG14, BMS12, Bal14, BR11, BK14b, BD14b, CCD11b, CCK13, CAH14, CH10, DZLD14, EFT13, Fan14, FPN^{+}13, Gra14, Guy13, HHS^{+}13, HAH14, IDNG13, JS13a, KL12c, LRLL11, LR12, LB10, MA14, MCDT12, NKF14, PYK10, QL11, QL12, RHB12, TP12, TLNE10, YVL^{+}11, Zen11, ZLCW14]. Hyperboloidal [Zen11]. hyperelasticity [FGN14, KD10]. HYPERS [OK12]. hypersingular [Le 14b]. hypersonic [DWZ10, KS13a, MMSI13, PPWZ11, SAHS14, SG13]. hysteresis [DL10c]. HZETRN [SBB10, SBT13].

IASI [HMOH11]. IBC [Ano10a, Ano10z, Ano10-27, Ano10-28, Ano10-29, Ano10-30, Ano10-31, Ano10-32, Ano10-33, Ano10-34, Ano10-35, Ano10-36, Ano10-37, Ano10-38, Ano10-39, Ano10-40, Ano10-41, Ano10-42, Ano10-43, Ano10-44, Ano10-45, Ano10-46, Ano10-47, Ano10-48, Ano11b, Ano11-27, Ano11-28, Ano11-29, Ano11-30, Ano11-31, Ano11-32, Ano11-33, Ano11-34, Ano11-35, Ano11-36, Ano11-37, Ano11-38, Ano11-39, Ano11-40, Ano11-41, Ano11-42, Ano12a, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano12y, Ano12z, Ano12-27, Ano12-28, Ano12-29, Ano12-30, Ano12-31, Ano12-32, Ano12-33, Ano12-34, Ano12-35, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano13x, Ano13y, Ano13z, Ano13-27, Ano13-28, Ano13-29, Ano13-30, Ano13-31, Ano13-32]. IBM [KBTD12]. IC [DZL13]. ice [CMG^{+}13, JG13, LTS^{+}10, LPE^{+}11, LKT^{+}12, LKLG14, LJX^{+}14, LFLV14, PTM13]. ice-ocean [LFLV14]. ice-sheet [PTM13]. ice-shelf [PTM13]. icosahedral [iIT14, IX10, PB13a, Pud11]. ideal [BD14a, CLQX13, CRT14, HRT11, LXY11, LX12, SZH12, SIDG13, YVL^{+}11]. IDeC [NS11a]. identical [CBGK13]. Identification [AGS10, BLM11, Kaz10, BCB12, BBF12, PM14]. identities [AINF13]. IDO [OAK11]. II [HH13, AP12b, ABK13, BHS14b, CHZ13b, DL13b, HH12a, JPC^{+}14, KJWR14, LR12, LEH12, LB10, Nis10, PZG14, PRA13, SV13, SYY14b, YS11b, ZWL^{+}12b, vdVR12b]. III [NL12b, Tyg10, UGkM13]. ill [CBPS13, JZ10]. ill-conditioned [CBPS13]. ill-posed [JZ10]. ILU [EL11, GM12b]. image [DXB^{+}13, YCH^{+}13]. image-based [YCH^{+}13]. imaging [DR14, HMM^{+}10, NLGL14, Par12b]. imbalances [SSR^{+}13]. IMC [LGP14]. IMEX [LKLG14, GBNS14, KKLR10, KG14, Par10, WLW13b]. Immersed [FRP14, PNPF10, ABD14, ADSF13, BBGP14, BF13, Bre12b, CKS14, CHC11, CFKL11, CB10, CLS10, DP12, DGF14, FDHP11, GGH14, GF13, GH10, HL12b, HDD14, HKL14, HCS11, HCS12, HW14a, JYHT10, JMW12, KKP10, KSR^{+}14, KF12, KLP10, KL10a, KLPS14, LT10a, LGP10, LKCY11, LY13a, LL12a, LSZ13, LH14b, MRD13, MVS13, MDH^{+}10, MZB12, NTV14, SM11b, SM11c, SSC13, TLZ^{+}11, TH14b, WZ11, WS10c, WYC13, Xu11, YS12a, YFL^{+}13, ZF12, ZF14, CF11a, GCBH10, GCBH13]. Immersed-boundary [PNPF10]. immersion [WY11]. immiscibility [TFP13]. immiscible [Hys12, LRT13, dv23]. impact [BDMP13, BN13, CWL10, KSU13, LL10a, SKU13, SDH12, Zoh13]. imparts [BRVE12]. impedance [GJ14a, IK11, Lee14a, LHH14, SP11]. implement [SdAW14]. Implementation [AM12, HE14, LTC10, LPE^{+}11, MF10, WLL13, BMD^{+}13, BB13, BILM14, BERB^{+}12, BDMV11, CFCA13a, CFCA13b, CNSFD13, CCB12, CDB13, FDHP11, FL13, GM12b, MTAC12, SLT^{+}14, TWSN12, TPA^{+}13, VWP11, WFK11, WD13a]. implementations [Ket10, NWLS12, OAKR14, Pri14]. Implementing [VSK10]. Implicit [DE14, PVLT10, PP14, SS12, Tit12, VF10, WLW13b, AV13, ASB10, BK11, Bea12, BSB14, CCD11a, CCB13, CCB11, CZ11a, CCB12, CC13, CCL^{+}14, CPCU13, CP13b, CGP10, Cui12, DTG^{+}14, DDF14, Den11a, DFFL10, DHH11, DGAH10, Fal13, FP14, Gen11, GMA10, GM10a, Gho17, GJ11, GMG12, HS13a, HL12b, HA11, Ian11, Jar12, JKW11, JSA14, JZ13, KK10a, KKvZB14, KGKS13, KG11, KTT13, KHHK12, Kuz14, Lan14, LXM14, LGP10, LY13a, LX13a, LXL13, LGP14, LL10b, MF14, MVCFM13, MBD13, MT14a, MTAC12, MN14, NYM13, NMCZ14, NPC11a, NPC11c, NDBG14, OM12, PGLT10, PGL11, PWB^{+}14, PP12, RS10a, RGMC13, RPC14, SS10a, SZF12, SPB^{+}10, SH11, SSHM10, SR13c, SK12b, SKt10, SN13b, TDU14, TBR13, WXY14, WMYYG10, WG11, WC14b, YVL^{+}11, YC11, nZzS11, ZOL^{+}11, ZYS14, LKLG14, Den11b]. Implicit [SPS14]. implicit-Euler [LX13a]. implicit-explicit [AV13, FP14, KHHK12, RGMC13, SSHM10, LKLG14]. implicit-macroscopically [PP12]. implicit/explicit [KK10a]. implicit/implicit [SKt10]. implicitly [LKM13, YW13]. implosion [JNW^{+}14]. Importance [DKH11]. imposition [UGF14]. imprecise [DMP14]. imprinting [PB13a]. improve [CS14c, CM13, DMP14, XCL14]. Improved [AA13, BV13, CWL10, GP13, iIT14, MS12, MGN11, Rot13, SM12c, Stu10, WW14, ADMA13, HKLY13, HCS11, KF12, PA14, SM11c, SY12b, SCC14, SLH^{+}11, TR12, VXB14, WZZZ14, WTK14, WS10c, YZ12, YYB12, ZJW12]. Improvement [Ren11]. Improvements [PGA11, Nor13]. Improving [GHM10a, LTS^{+}10, FSTY14, GHM10b]. impulse [TR12]. impulsive [Jin12, LXSR12]. in-order [KH11]. Including [CLG13, Gen11, Gho17, BBC^{+}13, BD14c, DLM13b, Ger12, KM11, MGN12a]. inclusions [Par12b, SBA13, SLC^{+}11]. Incompressible [BHS14a, BHS14b, LXSR12, ZMD14, APF13, BMP14, BNNW11, Boe11, BD11b, CIJS14, CCKCG10, CEP12, CP10, CXLF14, CB10, CLS10, CL11b, CJN^{+}13, CDB13, CdCD12, DLZ^{+}11, DP12, DF14, DS10, DS12, Don14a, Don14b, DKC14, EQYF13, FW12, FSL10, GLL14, HRK^{+}10, HOK10, HF11, HHH13, JM12, KKAS12, KGG10, KKO13, KRT14, LDM14, LSBJ14, LVFK14, LMGS13, LLP10, MBS13, MDH^{+}10, ML11b, Min13, MSB14, NPC11c, NL12b, NvdVB13, OA10, OP10, OM12, OS12, PMSH10, QS11a, RP13, RC12, RCvdV13, RHGT10, SS13a, SK12a, San13, SVK14, SHCY13, SLI10, STH11, SFCF10, SR11, SY11c, SHA11, SSL13, TY11a, TVEC12, Tyl14, VBCM10, Vre11, Vre14, WZ11, WSS12, WS10c, WTSZ10, WS11, XIIX14, XQX13, YCLK12, YSW14, Yin14, ZA14, ZBS12a, ZBS12b, ZHS10, ZWH^{+}12, VV13]. incompressible/compressible [CP10]. Incorporating [CPJ11, YYY^{+}14]. Incorporation [Mel11]. incorrect [OJ11]. increase [Ide12]. incremental [LEM11]. indefinite [RVbZ10]. independent [IH13, MSTL13, PMF13, RWX11, RB13, Sog14b, ST11a, ZHPS11]. index [CK11, MM12]. indicator [FSTY14, HKLY13, VR14]. indicators [LQ10b]. indifferent [GK14]. Individual [AKL^{+}13]. Individual-specific [AKL^{+}13]. Induced [SCB10, SM11b]. induction [RKGM14]. inductionless [BMP14, PBC11b]. inelastic [SSX13]. inequalities [BSS12, BSB14]. Inertial [NvdVB13, UPDB13, Dai13, Le 12, OTV13, TTR^{+}12, WCM^{+}14]. Inexact [NPV14]. inextensible [AELV14, FBM14, KL10a, RVB10]. inference [BND^{+}12, EM12, JZ10, Val14]. infiltration [GCFJ13]. infinite [BS10a, HVD13, YD11, Yin14]. infinitesimally [YTYJ12]. inflammation [SMF^{+}13]. Inflow [GHM13, GCA13, JCT11, JLC14, LDS11]. Influence [LTC13, DOR10, HDPF11, Liu11b]. influence-matrix [Liu11b]. information [FF13, RKM11, ZFJ11]. Infrared [HMOH11]. inhomogeneous [CKD10, LP14a, LRBT14, LQB14, PVDW14, SFCF10, YB14, ZTL13]. initial [GHM13, LWS14, LN14, MBS13, PBB14, RKM11, SN14, hT10b, Vai11, WLW13a]. initial-boundary-value [SN14]. initialization [FB11, MD14]. initio [SD14a]. injection [VDD10a]. inner [Zho10]. inner-outer [Zho10]. input [MZ11a, WZ14]. inputs [DO11b, KL14a]. insect [NL12a]. Insights [VCJ11]. insoluble [XYL12, XR14]. Inspection [Gas11]. inspired [RP13]. instabilities [CDT13, KL12b, SGC13, SHCY13]. instability [CF11b, DST14, DWZ10, DE14, GV14, HDPF11, Kwo14, MMSI13, MJ13b, VGCMG11, VCS14]. instationary [vVVK10]. integral [AH13a, AST13, AN14, BG10, BR10b, BRS10, Bre12a, BG12, BL13, CC12, CGMQ14, CGS13, Dri10, FHMS11, Gen13, GK13, GMC13, GHM14, GL12b, HK14, HHGC14, JM12, JRTY11, JKQ13, KS11c, KQ11, KTT13, KG12b, KNR11, Le 14b, Li10a, LL12b, LL14b, MST14, PSL20, QA11, RVB10, SRS10, SC11, Tak14, WJC13, XW10, YW13, YHM12, ZN10, ZIOF10]. integrally [Dri10]. integrals [BG13, KM14b, Men12]. integrate [Mar10a]. integrated [MDTC11, MDTC13]. integrated-RBF [MDTC11]. Integration [SGW10, AVKP11, Ala14, AMP14, CZ11a, CP13b, DCC^{+}13, GL14, GAW^{+}14, Gui12, HCP10, Ide12, JLZ13, JZ13, KK10a, LKLG14, LCK11, LN10c, MF12, MYM12, MNS13, NT14, RH13, SdAW14, TQW14, WZN14, Web14, YVL^{+}11, ZJS15, ZOL^{+}11]. integrations [SKW14, ZHS10]. integrator [AV13, BM10b, CTP13, HT14, LHK12, TLG10, TAAY14]. integrators [BRVE12, CK11, CMO14, CHHS11, GNGAS10, GBNS14, HDHZ13, KHZ10, ML14a, OBTC^{+}13, WWS13b]. integrin [JJM13]. integrin-ligand [JJM13]. integro [Dri10, YXZ13]. integro-differential [Dri10, YXZ13]. intensity [RKD^{+}14]. inter [RE13]. inter-scale [RE13]. interacting [AD14, OTV13, VSC^{+}11]. interaction [BK12a, BC10a, BW13, BBGP13, BCG^{+}13, CKS14, Deg11, DG12, DGAH10, DEKBF14, FL14, FRP14, FG12, FP14, GCBH13, HCS11, IBO14, KVM12, LS13a, LJG14, MMH12, NL12a, NPV14, OM12, PB11, Ric13, RMSF11, SZ11, STC10a, SM12c, TLZ^{+}11, TDL^{+}14, Unf13, VRB10, WY11, WYC13, WC14b, ZF12]. interactions [Atz11, CHC11, CKD10, DXB^{+}13, GM12a, GKF11, HRBK10, Lap11, PHSA14, RUL^{+}13, RZ10, RKD^{+}14, RWMG11, WKL^{+}14, YL11, YS12a, YBK13]. Interactive [BGBR12]. Interconnect [DL10a]. interconnecting [XJ14]. Interface [DK14, FS10a, SS12, AELV14, Ala14, Ald10, BB12a, BHS14c, BH13b, BB13, Bea12, BR12a, BF13, CN12, CDT13, CW10, Deg11, DjY14, DFD14, EC13, ELM14, FG12, FK14, GMB10, Gao13, GW11, GK14, GGT13, HHA14, Her10, HS13b, HDD14, HSWZ13, IST^{+}12, IXX14, IKO^{+}14, JSA14, JLZ13, JH12, KK14, KSU13, KKP10, KT14, KTT13, KGSS10, LP13b, LL12a, LSZ13, LFRM10, LH10d, LZZR10, MW12, MNR11, MRD13, MDH^{+}10, Min13, MZB12, MWW^{+}13, NZ14, NKF14, OD13, Qua11, Sam14, SN13a, SM11c, She13, SKCC10, SCC14, SPF10, Shu14, SX14, SHA11, SHA12, Sog14b, SS13b, SS14b, SdAW14, SD14c, TSLV11, TFP13, WS10b, WK13a, XZWW12, Xu11, ZCD10, Zha10b, ZLH12]. interface-aware [BHS14c]. interface-capturing [JH12]. interface-enriched [Sog14b]. interface-immersed [HDD14]. interface-tracking [BR12a]. interfaces [BF14, BD10b, BWC11, BvBZ^{+}10, BDH13, DvW14, GGH14, GL12b, GHL14, HWST12, HLZ14, HWW10, ID10, KL10b, NZ14, PS14a, PS15, PJCS14, UGkM13, XZW11, ZL10, ZJL^{+}14, Zha10b]. interfacial [AT13b, BS10b, CDT13, CL14b, CZ11b, Fuj13, NCKN10, SSHT14, WZT^{+}14, ZYF^{+}10]. Interferometer [HMOH11]. Interior [HLY11, DL10a, EGL^{+}13, JMKK12, LRL10]. Interlocked [Mét11]. intermediate [PDC13]. internal [Bal14, BL11, CGJ11, JYHT10, PLBR13, VED13]. interpolated [CLS10]. interpolating [dlLPCP12]. Interpolation [GM11b, GWCA14, LB14, WG14, BR10a, BBC^{+}13, BRSS11, BY11, CEGG14, CJM13, DSHP11, FS10b, KBRM12, KM14b, sLwG10, LSVW10, LCB12, LBS10, Mar10a, MSD12, PDSS10, dCRCS10, Soa13, Tak14, XY14, ZZB14]. Interpolation-based [WG14, Tak14]. interpolations [LTC10]. interpretation [ABPSM11]. interpretations [ANIF14]. interrogation [LFRM10]. interval [LJ13b]. intrinsic [Atz10, OLC13]. Introducing [ZA14]. introduction [CMO14]. Intrusive [TLNE10, AC13, NV14, PL12]. invariance [Del14, SS14a, WC11]. invariant [BN14b, LJ13b, WZT^{+}14, ZS14]. invariants [SGL14]. Inverse [DR14, IK11, Lee14a, TS10, WKMJ13, AS14, CNR11, CLN14, CB12, CXZ10, CV10, EWH14, FY13, GJ14a, IWG13, IJZ13, JR12, Jin12, KL11a, LLZ14, LLW14, Mét11, MB10b, TWSN12, WdHX10, WK13b, YLW10, ZM14, HG10]. inversion [LWD13a, LWD13b, OBNN13, OBNN14, WLPG11]. investigation [CIJS14, CG13b, LRT13, OLFS13, Ste14, dv23]. investigations [Ian11, PPTMDK13]. inviscid [ALR11, CCLS10, CF11b, LLX11a, NHE14, SLI10, TS11, YSW^{+}13]. invisibility [FGE11]. involving [DLC11, DS12, FK14, GBB^{+}13, TDL^{+}14]. ion [AiINT14, BGGW13, BHBM10, HH12b, HH13, LTC14, MTSG12, PMS14, Tok11a, ZCW11]. ionic [CEK14]. ionization [CCMG^{+}13, KM11, MPPP12b, MPPP12a]. ionized [Gio10, KA12, PMS11, PSM11, PSM13, PMS14]. ionosphere [TRSMM^{+}13]. ionospheric [DST14]. ions [LH13]. irregular [AAI10, AT13b, BvBZ^{+}10, GM12b, HG11, HWST12, JMGN13, MTG11, SR11, TDVG13, Tsu10, WTYC13, ZYT13, ZWH^{+}12]. irregularly [BH13c]. ischemia [WKMJ13]. Ising [KO12, LWZ^{+}13, MMM11]. iso [GRS14]. Isogeometric [ABKF11, BAB^{+}13, BLA^{+}14, BSV14, EH13, LDE^{+}13, BGGW13, BA10, CRS12]. isoparametric [VQVD11]. isoperimetric [WG11]. isothermal [LGE^{+}13, PL14]. Isotropic [SG14, TAAS13, BCL14, GHM12, MS14b, OBT11, OTV13, WWX^{+}10]. Issue [Ano13-41, Ano13-42, AQ11, Ano10-49, Ano10-50, Ano10-51, Ano10-52, Ano10-53, Ano10-54, Ano10-55, Ano10-56, Ano10-57, Ano10-58, Ano10-59, Ano10-60, Ano10-61, Ano10-62, Ano10-63, Ano10-64, Ano10-65, Ano10-66, Ano10-67, Ano10-68, Ano10-69, Ano10-70, Ano10-71, Ano10-72, Ano11-43, Ano11-44, Ano11-45, Ano11-46, Ano11-47, Ano11-48, Ano11-49, Ano11-50, Ano11-51, Ano11-52, Ano11-53, Ano11-54, Ano11-55, Ano11-56, Ano11-57, Ano11-58, Ano11-59, Ano11-60, Ano11-61, Ano11-62, Ano11-63, Ano11-64, Ano11-65, Ano11-66, Ano12-36, Ano12-37, Ano12-38, Ano12-39, Ano12-40, Ano12-41, Ano12-42, Ano12-43, Ano12-44, Ano12-45, Ano12-46, Ano12-47, Ano12-48, Ano12-49, Ano12-50, Ano12-51, Ano12-52, Ano12-53, Ano12-54, Ano12-55, Ano12-56, Ano12-57, Ano12-58, Ano12-59, Ano13-33, Ano13-34, Ano13-35, Ano13-36, Ano13-37, Ano13-38, Ano13-39, Ano13-40, Ano13-43]. issue [Ano13-44, Ano13-45, Ano13-46, Ano13-47, Ano13-48, Ano13-49, Ano13-50, Ano13-51, Ano13-52]. issues [MA13, NMT^{+}12, VLM13]. Itô [HHGC14]. iteration [ADMA13, BK11, CXZ10, Par12b, ZCS14]. Iterative [BHK^{+}10, CJ14, SSS11, ZRM13, AH13a, AK12, CBPS13, CGJ11, DL10b, EMS11, EWH13, HJ11, HKJ11, HL12a, JMW12, JHDC10, JMKK12, LSE13, LYJ14, LGP14, LTL11, MP13, PL12, RT14, Ren11, RVbZ10, Sha13, SS14c, SV10, Tok11b, WD13b, WTK14, XH13, YD11, YM14, Yao14]. Ito [ZYLK13]. IV [ZMD^{+}11].

J [ASS17, CFCA13a, Dav15, Gho17, HY11, LWD13a, NHT^{+}13, OBNN14, PS15, SYV17, SS14b, TK15, XMDG14, YKWS13a, ZJS15]. Jacobi [And13, CW14b, DBAV14, HWHW10, LP14b, YO11, YM14, ZQ13]. Jacobian [LFLV14, BDTW10, CTP13, GL12a, Han11, LTS^{+}10, LPE^{+}11, LKT^{+}12, LvZB10, PMMB14, SV13]. Jacobian-free [LFLV14, BDTW10, CTP13, GL12a, Han11, LTS^{+}10, LPE^{+}11, LKT^{+}12, LvZB10, PMMB14]. Jacobians [RPC14]. Jameson [AGBC14]. Jameson-Schmidt-Turkel [AGBC14]. JCOMP [MPPP12b]. Jenny [Dad12]. jet [SSS11]. jet-noise [SSS11]. jets [CMS10a, SY12c]. JKD [BCL13]. joint [BND^{+}12, VK12]. Jones [KM14b]. Journal [Dad12, EH15, HSK^{+}15, XS15, Try12]. JST [AGBC14]. jump [KMHJ10, MNR11, ZL10]. jumps [FGR12, FMT11b, MP13]. junction [Li13]. JVIE [PVDW14].

KADATH [Gra10]. Kalman [BGM12, GLM14, HMM14]. Kantorovich [CDF11]. Kapila [KK10c]. Karhunen [PSDF13]. Kelvin [CF11b, Gri13]. Kendall [Tan11]. Kernel [MZ11a, HD13, HK14, YXZ13, YW13, ZHPS11]. kernel-free [YW13]. kernel-independent [ZHPS11]. kernel-split [HK14]. kernels [MSD12]. kind [JRTY11, JKQ13, KS11c, PSL20]. kinematic [LLJ10, MMHA^{+}12, WT12b, WSW13]. kinematic-wave [WT12b]. kinematics [NW12]. Kinetic [ABPSM11, DBO^{+}10, MSS^{+}14, ZGSZ10, AP11, AP12b, AHOT14, AKP^{+}12, BWMG14, Cha13, CJLC11, CXLC12, DDM10, DL13a, DL13b, ES13b, FJ10, FR13, GCA13, GJLY13, GH12b, HAD10, KA12, KHJ^{+}13, KGK13, KSB14, LE13, LXF10, LLX11a, LYXZ14, MMM11, MOV12, Mie13, Mit13, NKSR12, OAXL13, PZTW12, PLT14, SBC^{+}14, SR13a, SLI10, Sza12, TDU14, VCS14, XH10, XX13, YSW^{+}13, YSW14, YF11, ZA14, ZKV10, Lip12, MHL^{+}14]. kinetic-fluid [DDM10, GJLY13, KA12]. kinetic-moment [GCA13]. kinetics [DR11, LRS10, NS14a]. Klein [BB12b, VBSK10, VW14b]. Kleiser [Liu11b]. knife [AV13]. Knudsen [FF13]. Kohn [LSE13, FGZ12, LLYE12a, LLYE12b, MNL^{+}13, SL13a, SBO11, ZCS14]. Kolgan [Rod12, vL11]. Korteweg [LGE^{+}13, YHL13]. Krause [YWS13]. Kronecker [GHDS10]. Krook [MWRZ13]. Krylov [YJK12, AD14, BDTW10, CFW^{+}13, CGG12, CZ11a, CCL^{+}14, DRZ14, GL12a, Han11, JZ13, LTS^{+}10, LPE^{+}11, LKT^{+}12, LFLV14, LvZB10, LL10b, MS10a, PMMB14, SEDW13, TS14, ZD10, vVVK10]. Krylov-based [CCL^{+}14]. krypton [KM14b]. Kullback [BK12b]. Kutta [MBA14, WLW13b, AV13, BMD^{+}13, BGR11, BR11, FP12, Ian11, KKvZB14, Ket10, KHHK12, Lan14, LP14a, LLH10, NYM13, NMCZ14, NDB12, RT14, RHXQ14, RA10b, SK12a, San13, SLT^{+}14, SK12b, Tok11b, TD11b, TD12b, WS10b, XCL14, ZT13b, ZS13, ZZSQ13, vdVR12b]. Kutta-Legendre [MBA14].

L [EH15, XS15]. LABSWE [Zho11]. lacunae [PT12]. laden [Bre12b, CD13a, KF12, KBB13, MB11, VdMS^{+}14, ZF14]. lagged [SS10a]. Lagrange [BCD14, CD13a, RHGT10, Wal13]. Lagrange-remap [BCD14]. Lagrangian [HAQ12, Ald10, AH10a, Atz11, BHS14c, BAB^{+}13, BLA^{+}14, BBG^{+}11, BR12a, BRSS11, BD14b, BBD14, CK11, CdCNH14, CS12a, CEP12, CS10, CS14a, CS14b, CGMQ14, CDLL12, CSC14, CMS10b, Del11, DL12, DKR14, DF10a, ES13a, GT12, GIQ11, GH12b, GQ13, HLM11, Her10, IL11, Ion13, JZ11, JGZL14, KBS14, KD10, KS12, KS14, LNU10, LGF11, Li13, LHB10, LS10b, LSVW10, LMS^{+}10, MAB^{+}13, MRD13, MVS13, MVG^{+}12, MKB^{+}13, MLBK14, MV13, MHGM14, OS11b, OS14, PRA13, PB14a, QC10, QS11a, QS11b, RSS12, RS11, SSB13, SZF12, Sco12, SYL10, SB12, SP14b, TS12, TBR13, VW14a, VXB14, VMA14, VBCM10, VSC^{+}11, VdMS^{+}14, Wan10, WPP10, WYS12b, WD11, WYC13, XQXC14, ZS14, Zer10]. Laguerre [Sal14]. Lamb [GP13, GD13]. laminar [LH10b, RK14, TAD14, Tyl14, vVVK10]. Landau [JY11, RSS10, SV13, TAT14]. Lane [xLqW13]. Langevin [BRVE12, DCC^{+}13, HT10a, KJC13]. Langevin-like [BRVE12]. Langmuir [MCJ14]. LANGUAGE ED [Ano12-31, Ano12-33, Ano13-31, Ano13-32]. Laplace [BR10b, FCY13, HLW13, SF14, SYL12]. Laplacian [Ost10, TAAS13, TDL10]. lapse [BLM11]. Large [BA10, Hel11a, LT10a, OBT11, OAK11, SZ10, SSK^{+}12, And10, ADSF13, BGGW13, BPS14, DV13b, DS12, Du11a, DSZ13, ES13a, FGR12, Fat10, FMT11b, GLSC12, GW10, HA11, JLS^{+}14, JK10, KSL10, KSB12, KEGM10, LFL11, LG14, LS10a, LBRA^{+}13, MS10a, MS14a, McD14b, MHGNM12, MMH12, NW10, NNS11, OAKR14, OL13, PGLT10, dlLPC10, dlLPCP12, QL11, QL12, RG13, SFCF10, TDL^{+}14, TRL14, VKBK11, WDP^{+}14, WIJ13, XDZ^{+}14, YB14, Zho10, CP13a, Den11a, MDH^{+}10, MHHGN14]. large-displacement [MMH12]. Large-eddy [OAK11, BGGW13, DV13b, GW10, KSL10, KSB12, MS14a, McD14b, OL13, dlLPC10, dlLPCP12, RG13, SFCF10, TRL14, Den11a, MDH^{+}10]. Large-scale [OBT11, SSK^{+}12, LG14, LS10a, MS10a, OAKR14, XDZ^{+}14]. large-sized [JLS^{+}14]. laser [CCMG^{+}13, CKD10, CBCM^{+}11, DGAH10, MdVBS13, RKD^{+}14, SSR14, VGCMG11, YXD^{+}14]. laser-matter [RKD^{+}14]. laser-plasma [CKD10, CBCM^{+}11, DGAH10]. late [OMCO14]. lateral [KJ11]. Lattice [CSB14, CLAL14, LL10a, LVZK14, MB10a, SMP^{+}13, WZG^{+}13, YN14, Zho14a, APGL12, BBAP14, BSDM12, CB11, CLZ12, CHKT13, FSL10, Gil13, Gra14, GWCA14, HE14, HW14a, HHL14, HW14b, HCM10, IF10, JYHT10, KH13, KMA13, KJC13, LMLC12, LRT13, LMK13, LZC^{+}13, LZV12, MD13, MS14a, MZ11d, MZ14, ML11b, NSA11, NKSR12, PPB14, RK14, RS10a, SSC13, SS14a, TAAS13, TLZ^{+}11, TRL14, VMS11, VMS12, VS14, VM10, WK10, WZT^{+}14, WC11, WS10c, WS11, XS13, YYY^{+}14, YZ12, YN10, ZGSZ10, ZF14, dv23, vTCL^{+}10, CL13, Del13, Del14, FRP14, Mel11, MVS13, SHFB12, VS11]. Lattice-Boltzmann [CLAL14, MB10a, GWCA14, MD13]. lattices [CCC12, Del14, GM10b, KGKS13]. law [DMT^{+}11, GLL14, HLLW11, MN11, DMM^{+}13, PSM11]. laws [BMS12, Bal12b, Bal14, BK14b, CCD11b, CCK13, CDL11, CLS13, DZLD14, EFT13, Fan14, FCN^{+}13, FC13, GAV13, GPP11, HHS^{+}13, HAH14, IDNG13, JS13a, Kri10, KL12c, LRLL11, LR12, MSS12, MNS13, MLM12, NF12, Par12a, PYK10, QL11, QL12, RSS13, SABH11, SHZ13, TS10, TWSN12, hT10b, XLD^{+}11, XCL14, ZS10a, ZWL^{+}12a, ZWL^{+}12b, ZLCW14]. Lax [CCK13, FY13, GL14, HSZ10, HLZ13, LQ11, TS10, TWSN12]. layer [AKMD14, BBCH^{+}12, BS10a, CCD11a, CVI^{+}13, DLC11, DWZ10, Dur14, FL13, KKS13, KBGO13, LH10a, LBT11, LMQ^{+}11, MMSI13, MBL^{+}14, MMIW14, NMZC13, NMCZ14, SF12, SDH12, WQCS13]. layered [AN14, ABS^{+}14, CC12, Dur14, HNS12, Li10a, LTC14, STD14]. layered-media [AN14]. Layers [MJ14, BDT10, FL14, HJK14, KAA12, LN10a, Man12, MSTL14, RVbZ10, RS10b, SSM14, XS13, Zen11, vDS12]. LBM [XS11, XMS12]. LCP [CVN13]. leads [BK11]. leak [GF13]. leak-proof [GF13]. leaky [YFL^{+}13]. Leap [HT14, SWX12]. leap-frog [SWX12]. leaping [FWP13, YR13]. learning [Phi14, RKM11]. Least [ADB14, CD13b, CXLF14, OS12, ABL^{+}11, BS11b, HGW11, HMM^{+}10, HY09, HY11, JMFO13, NOT14, PR11, WHB14]. Least-Squares [CD13b, OS12, HGW11, HMM^{+}10, HY09, HY11, NOT14, PR11]. least-squares/finite [HGW11]. left [LS13a, MNKS13]. Legendre [MBA14, KGG10, Ols12, QZM^{+}12]. Leibler [BK12b]. length [CG13a, Cou11, JR14, KS13b, Qua11]. Lennard [KM14b]. LES/PDF [VWP11]. LES/RANS [XJ12]. Leslie [CTSM13]. less [IK11, LD12, OBT11]. less-than [LD12]. Letter [HAD10]. Level [CKS14, CLW10, CV10, ILM^{+}13, RB14, ZLL10, ABD14, BOD11, BW13, BH13b, BILM14, BV14, BGG12, CCFCM11, CP12, CZL10, CWL10, CF11c, CKR14, DCVM11, ELM14, FB11, HZ11a, HMS10, HCS^{+}10, HQL^{+}10, Hys12, JCC14, JS13b, KAFB11, LSM14, LP13a, LDM14, LWF11, LZZR10, MCLG10, MD14, Min10a, MSS12, MHGM14, NRS10, NW10, OP10, OD13, PHRG13, SOG14a, SM11a, Sam14, SN12, Sha13, SJ10, SKR14, SS14c, VDD10a, VB13, WABI11, WSS12, WJTY10, WZ10, XYL12, XR14, YYY^{+}14, YL14b, ZT10a, fZlC11, ZWL10, ID10]. Level-set [CLW10, ZLL10, BOD11, BGG12, CB10, MCLG10, TVEC12, XYL12, XR14, ID10]. level-sets [ABD14]. Leveraging [WTK14]. Leverett [MJM13]. Levin [STD14]. lid [MAPP13, RTS13]. lid-cavity [MAPP13]. Lie [CMO14, CHHS11].

Lie-symmetry-based [CHHS11]. lift [SL13b]. lifting [CN12]. ligand [JJM13]. light [BHS13, IFD^{+}12]. light-scattering [IFD^{+}12]. like [BI11, BRVE12, CR14, Don10, FDHP11, LL14a, LZZS13, MLBK14, Par12b, TK12, TK15]. likelihood [MOV12]. limit [CJNS12, CTG12, DDS12, GH14, KWHD12, Mie13, RS13b, RS14b, SLI10, VP10, WHB14, WIJ13, WJTY10, YR13]. limitations [BCM^{+}11]. Limited [KG12a, AA13, BS14b, Nor14]. Limited-area [KG12a]. limiter [DFNNRdlA14, ICC^{+}10, WP10, XQX13, ZS14, ZT13b, ZS13]. limiters [GD10, GL12a, GTSC14, LRLL11, LR12, ZQ11, ZZSQ13]. limiting [DZLD14, Ger12, KKL10, KMSS10, Kuz14, MMD^{+}11, PYK10]. limits [ARF12, DWL10]. line [BSSW12, Coa12, GW12, GW14, HGW11, Per13, RB14, SN12, SSC13, XR14]. line-based [Per13]. Linear [ABHS12, LZZS13, AC13, Abr10, AHOT14, AH13b, BD10b, BRS11, BGM12, Cap11, CH10, DOR10, DEKBF14, FPN^{+}13, FN14, FVR13, GJ11, GGT13, Guy13, HR10, Ide12, JHDC10, JTXZ12, KJ11, KNR11, LLW14, LJ13b, LM13, MN12, Mie13, NL10, NvdVB13, PDZ^{+}14, Pie12, Ren11, RLPM12, SBWF14, SV13, SHZF12, TDVG13, VDP11, WL11, XFB^{+}14, ZKV10, ZWH^{+}12]. linearised [MBS13]. Linearity [WDGY10, WGD12]. linearity-preserving [WGD12]. linearization [FAY^{+}13, PR11]. Linearized [CD10, Chi12, HSD11, HDD12, MMR12, RBS12, SLGB12, SGL14, dPSS12]. Linearly [GJ11, BLJ11, CNST13, PSF^{+}14, WXY14]. liners [GB14]. lines [SS13c]. Link [AOCR12, OAKR14]. Link-wise [AOCR12, OAKR14]. linked [CVC12]. linking [WQCS13, ZKKF10]. Liouville [CT13, HAMA14, JR12, ZK13]. lipid [SM11a, SBA13]. Liquid [LNS13, BGGGS11, BDH13, CTSM13, DLDL10, HMHO13, HW14a, JLS^{+}14, KLMJ12, LT10a, LW11, MM12, SY12c, SSY13, STC10b, TK12, TK15, VDD10a, YFL^{+}13]. liquid-gas [DLDL10, HMHO13, LNS13]. liquid-liquid [MM12]. liquid/liquid [BDH13]. liquids [ZHW10]. load [SSR^{+}13]. Lobatto [DBAV14]. Local [AAT^{+}14, AH10b, BM14, CWL13, CW13, EGG12, FL10, GS11b, HLDP14, KRT14, XXS10, YDN10, ZQ11, BZ12, BS10a, BSS12, BGG12, CSX14, DTG^{+}14, EGW11, EAG14, ELM14, FLM12, GDHM11, GPCE13, GP12a, Gri12a, GZQ13, GXX14, HRBK10, JMGN13, Kri10, LGP10, LJX^{+}14, LLYE12a, LLYE12b, LR13, LBM14, MDTC11, NRS10, RV10, SPM^{+}13, SEBG12, WAM14, WTYC13, WT14, YO11, YR13, ZHS10, JWV12, SYJ11]. Local-global [EGG12]. Local-in-time [YDN10]. localization [DDM10, WKMJ13]. Localized [LK14, NT14, CSKP10, DD12, KSL10, MD14]. Locally [BCHM14, CWZ^{+}11, DFFL10, ZHY13, AELV14, EAN11, FJ12, MT14a, RVB10, SKU13]. located [SLI10]. locking [LSZ13, MMHA^{+}12]. locking-free [LSZ13]. locomotors [KST10]. Loève [PSDF13]. Logan [BY11]. logging [PMM^{+}11]. Long [AVKP11, LBR14, PA12, CGJ11, GAV13, LCNK10, MS14b, SC12]. Long-range [PA12]. Long-time [AVKP11, LBR14, MS14b]. longitudinal [BCG^{+}13]. loop [ADMA13, MRS14b]. Lorentz [VGCMG11, YXD^{+}14]. Lorentz-boosted [VGCMG11]. lossy [SVS11, She13, VBSK10]. Low [DST14, KHGW13, NGV14, Val14, AHS14, And13, APC11, BFSVC14, BCD14, CJNS12, CJLC11, DLDL10, Del10, DKA12, GKG^{+}13, GW10, GVV13, HGSK^{+}13, HKM12, HK11, KVM12, KKP10, Lee12a, MCA^{+}13, McD14b, MKB^{+}12, MBD13, Mor10, NYM13, NL12b, NDB12, NNS11, Ran11, Ran12a, Ran12b, Rie10, Rie11, SRN10, SM11b, SCB10, SKt10, SRL^{+}11, VSK10, WYS^{+}11, WAM14, XS11, XW14, YYB12]. low- [VSK10]. Low-diffusion [NGV14]. Low-diffusivity [KHGW13]. low-dispersion [NYM13, XS11]. low-dissipation [HK11, NYM13, XS11]. low-frequency [SKt10]. low-Mach [BFSVC14, GVV13, McD14b, Mor10, Rie11, SRN10, SM11b]. low-order [WAM14]. low-power [GKG^{+}13]. Low-rank [DST14, Val14, Ran11, Ran12a, Ran12b]. low-storage [NDB12]. LS [CB10]. LU [BMP14, PVLT10, PGL11]. Lubrication [LHV13]. Lubrication-type [LHV13]. lumpectomy [GSTB13]. lumped [AST12, DL10a]. lung [Mit13, YCH^{+}13]. LVAD [MNKS13]. LWR [BLM^{+}14, SSR^{+}13]. Lyapunov [CRK12, CG13a, Leu11].

M [EH15, XS15]. M1 [OHF12]. Mach [APC11, BFSVC14, CJNS12, DLDL10, Del10, GW10, GVV13, HGSK^{+}13, HKM12, Lee12a, McD14b, MKB^{+}12, MBD13, Mor10, Rie10, Rie11, SRN10, SM11b, YYB12]. machines [JFC^{+}13, MMM13]. macro [DLNN12]. macromolecular [LBRA^{+}13]. macromolecules [ORHH13, XFC^{+}14]. macroscale [SS13c]. macroscopic [IF10]. macroscopic-gradient [IF10]. macroscopically [PP12]. MADS [SRN13]. Madsen [OBT12]. magnetic [BM13a, EAG14, FMM^{+}10, GLL^{+}11, HH13, KST13, Kaw13, LXY11, LRN13, NLGL14, NL12b, SCB10, Web14, ZJS15]. magnetic-to-electric [EAG14]. magnetically [Jar12]. magnetized [MMA13, TD12a, VCS14]. magneto [SR13c, TG14a]. magneto-hydrodynamic [TG14a]. magneto-hydrodynamics [SR13c]. magnetodynamic [GLL^{+}11]. magnetoelastic [BWS10]. magnetohydrodynamic [Bal10, Bal12b, FM13, HRT11, Jar12, LT12, MLBS14, SCB10, VKBK11]. magnetohydrodynamical [BMD^{+}13]. magnetohydrodynamics [ASB10, BCP13, Bal12a, BD14a, CRT14, DTG^{+}14, Del13, DFNNRdlA14, Kaw13, KMvM^{+}12, KB14, Lee13, MTSG12, Pri12, SZH12, TP12, WFK11, WKH10, YVL^{+}11, YJK12, ZYS14, Zie11]. magnetostatic [LMBB11]. major [CM13]. MaMEC [vV12]. management [TE14]. manifold [LRS10]. manifolds [LS13c]. Manufactured [WCM^{+}14, BHL12b, GAW^{+}14, SHM12]. many [AJG^{+}13, EHA14, KKPV12, SD14a]. many-body [SD14a]. many-particle [AJG^{+}13]. map [LBR14, MTAC12]. mapped [CCB13, CDHM11, Shy10]. mapping [BY11, He13, Phi14]. maps [EM12, MPS^{+}14]. marching [sLwG10, MCD14a, MY14, RS13a]. marine [BPM13, CMG^{+}13, JG13, SF14]. marker [DKW10]. marker-particle [DKW10]. Markov [KSS14, LC13]. Mass [PvDtTB^{+}11, vV12, APF13, APGL12, ABS^{+}14, BK12a, BNT14, BRS13, GGH14, GK14, HSK^{+}15, HHL14, IL11, KJ11, LJX^{+}14, LSVW10, LJG14, MGJZ10, PP14, SN13a, SM11c, SLI10, SWR11, TS12, Zad11, Zer10]. mass-conservative [HSK^{+}15, SN13a, Zad11]. mass-conserving [HHL14, TS12]. Massive [ARF12]. Massively [AJG^{+}13, NIN^{+}14, HOK10, ITO^{+}10, MMM13]. Massively-parallel [NIN^{+}14, ITO^{+}10]. MAST [AST12]. MAST-edge [AST12]. master [HLDP14, JU10]. Matched [MJ14, CVI^{+}13, Dur14, FL13, GW11, HJK14, KKS13, KAA12, LH10a, LN10a, NZ14, SSM14, SF12, Zha10b, ZLH12]. matching [CZL10, KIG11, SCIE14]. Material [ZMG11, AT13b, BP13b, CW10, CS14a, CJP^{+}14, FGR12, FAY^{+}13, FK14, GMB10, GL12b, GHL14, JLZ13, JH12, KSU13, KBSV14, KGSS10, KS12, KS14, LX13b, MGJZ10, MMHA^{+}12, SSHT14, SC10, Wan10, XZW11, ZJT10]. materials [AJV11, ASGW11, BM13a, DLC11, DFW^{+}10, GG13, HBZL14, SSB^{+}12, ZBS12b]. Mathematical [SY11b, BBGP13]. Matlab [KA13]. matrices [AM13, And10, DS12, SY14a, SHZF12, ZT10b]. Matrix [Opp13, RJG10, WIJ13, Boy13, Brü13, DB12, HP13, HHGC14, HPLS10, HWW10, IFD^{+}12, Kim13, LTSH10, LWD13a, LWD13b, LLY11, Liu11b, OM12, SR13c, WT12a, YS12b, SBN12]. Matrix-free [WIJ13, OM12]. matrix-vector [LLY11]. matter [RARO13, RKD^{+}14]. matters [LG14]. maximization [AETT11, EHA14]. Maximum [ZZS13, FLM12, MT13, NMV13, RPA13, Rot13, VL14, XQX13, XQXC14, ZS10a, ZY13]. maximum-entropy [FLM12, MT13, RPA13]. Maximum-principle-satisfying [ZZS13, ZS10a]. Maxwell [Wol12, AAT^{+}14, AM11, Bér11, CWS13, CXZ10, CAH14, CCZ14b, CCY13, DDS12, DLC11, DZL13, DTYY14, EAG14, GG13, HR10, HJK14, HJZ14, HLY11, HLYS11, JMFO13, KHZ10, LRL10, LLP14, LY13b, LSW14, LBM14, NPC11b, PT12, PMSM12, RM12, SML13, SF12, SP11, ST11b, SS10d, TWMM14, TCL11, TEY12, WA12, WXZ10, XJ14, ZD10]. Maxwellian [HT10a]. MBN [SBS^{+}13]. McCormack [Sza12]. MCore [UJ12]. MD [MD13]. MDCT [YCH^{+}13]. mean [BDM13, FS10a, Gil13, BW14]. means [KL10b]. measurement [LLSW14]. measurements [BPM13, PMM^{+}11, SCIE14]. mechanical [GP12b, MPT13]. mechanics [ALS^{+}14, ABHS12, GBC11, JHDC10, MNKS13, NP12, NKS10, OLPM14, iT14]. mechanism [BRS11, MZ14, Rie10]. Mechanisms [RBNS^{+}12, BJLR14, Nag14a, Nag14b]. media [AN14, AST12, BHM13, BCL14, CTP13, CZ13, CSW14, CL11a, CC12, CSX14, Coa12, DM13, DFVY14, DL10b, DCVM11, Dur14, DV14, EGG12, FR13, FJ12, GPCE13, GCE14, GD10, HSK^{+}15, HZ11b, HLY11, JS13b, KFOS14, LMFH11, LP12b, LTC14, LW13b, LQB14, LE12, MZ11b, MdVBS13, MSTL13, MTD14, MF13, MJM13, MMJ14, NZ14, OJ11, OP10, OK13, PM14, RBB12, RCCC14, dCRCS10, STD14, SKN13a, SG10b, SG10a, SG12a, SG12b, SGS13, SGC13, SHCY13, SVS11, She13, SKN13b, SLC^{+}11, SRS14, STD^{+}10, TLG10, TL13, WXZ10, WT13, WHT14, WLW14, WSBG10, YD11, YL10, Zad11, ZM14, ZVNM10, ZYHW13, ZHY13, Zha10b, ZTL13]. mediated [ZM10]. medicine [SR13b]. medium [BRS12, CB12, CW14a, GCFJ13, IJZ13, Li10a, RM14, WS10b, WK14, XL11, ZYT13]. melt [KBTD12]. membrane [FLM12, JJM13, iT14]. membranes [PRA13, SBA13, WJCG13]. memory [BH13a, NS11b]. Merging [Lip12]. Mesh [DMMGM14, KS11b, ABD14, AIX10, ABMT14, AW14, BWC11, BWC13, BCK11, BDA12, BUH11, BBPC14, BCW13, CL10, CDF11, CGG12, CXLC12, CY14, CMG^{+}13, DH10, Del11, DF10b, DZHB13, ETW10, FL11a, Fuj11, GWCA14, HHL11, HB13, HKL10, JLY10, KWSJ11, KF13a, KG13a, KD10, KG14, LGP10, Lee13, LH10b, LH10c, LN10c, LX13b, LBM14, LDA10, LLS11, LCB12, MAPP13, MPR12, MVG^{+}12, MMH12, NL12b, NW10, PB13b, PWB^{+}14, Phi14, Qua11, RA10a, SVK14, SZF12, SQC11, SYY14a, SYY14b, SSW13, Sog14b, SBO11, SS10e, VRB10, Wan10, WR11, WG11, XZWW12, XJ12, YCY11, YD12, ZBS12b, ZJW12, ZL13, Lip12]. mesh-free [KG13a, SBO11, ZBS12b]. mesh-independent [Sog14b]. mesh-motion [CDF11]. meshed [BYCC14]. mesher [AKMD14]. meshes [ALR11, AGSG11, ABJ^{+}12, BMD^{+}13, BDA14, Bal14, BD14b, BBD14, BC10b, CCB13, Cha14, CZ11a, CRT10, CF11c, CDLL12, CBTR13, DvW14, DLM13c, Far11, FW12, GW13b, GBB^{+}13, GIQ11, IXX14, JLZ13, KWS^{+}11, KS11b, KK10b, KAFB11, KSD12, LRL10, LZ13, LS10b, LSV10, LMBB11, LM14a, MF12, MN11, MIJ10, MK11, NS11a, Per13, RS14a, SY12a, Sco12, SY11a, SY12b, TD12a, TGRL13, TTD11, TAD14, TR14, WR11, WQCS13, WAM14, WDGY10, WGD12, ZS10b, ZZS13, nZzSlL14, ZJL^{+}14, ZQ13, ZZSQ13]. meshfree [GM14a, PRA13, RPA13]. meshing [Cou11, CJN^{+}13, DDF14, KHGW13, Qua11, SPS14]. Meshkov [MJ13b]. meshless [DO11a, HCD11, HBM13, KBGRM10, MWP14, OJW10, RR14, SY11b, SPM^{+}13, SY14b, ZTL13]. meshsizes [Ge10]. mesoscale [DR11, KL12a]. mesoscopic [AVKP11]. metal [BWC13, JLS^{+}14]. metamaterials [HLYS11, LH14a, OK10, OK13]. metastable [ZHW10]. Method [CD13b, ILM^{+}13, KLL13, NDV^{+}11, SYJ11, VS11, ZSW10, ZKKF10, AB12, AB13, ABD14, AHA10a, AP11, AP12b, AS14, ASS13, ASS17, ÁCB12, AHS14, And13, AAG14, AAT^{+}14, ADSF13, AF13, AH10a, AEM13, AOCR12, AM11, BPM13, BFR13, BSSW12, BLOdV14, BD10b, BOD11, BDMP13, BGGW13, BH13b, BB13, BMK11, BMK12, BA10, BvBZ^{+}10, BFSVC14, BC14, BBGP13, BKKD14, BK12b, BYCC14, BUH11, BSCML13, BB12b, BV14, Boe11, BS10b, BLR13a, BLR13b, BM10a, BCM^{+}11, BR12b, BOK12, BDLGC13, BM11, BHL12b, BG12, Bre12b, BDKV14, Brü13, BGG12, BLR14, BTY14, BL11, CL10, CCD11a, CL14a, CKS14, CP12, CNR11, CLN14, CFCA13a, CFCA13b, CP13a, CPJ11, ÇD12b, CGM^{+}12, Cha13, CDT13, CTG12, CT13, CEP12, CF10, CP10, CWL10, CFKL11, CNST13, CS13, CZ13, CLSX14, CAH14]. method [CXLF14, CSW14, CLX14, CHZ13b, CHZ13a, CTJT13, CPCU13, CW14b, CB10, CZ11b, CLS10, CL11b, CGJ11, CGMQ14, CV10, CGS13, CRT10, CDL11, CDLL12, CKR14, CH11a, CSKP10, CSB14, CH14, CJ14, CCG11, CdCD12, Cui12, DYYA14, DR11, DH10, DMHL14, DLDL10, DV13a, DDN^{+}10, DLNN12, DM13, DRB14, DLC11, DLLW14, DvW14, DWL10, DTU12, DGM13, DP12, DFNNRdlA14, DMR14, DL13c, DjY14, DFD14, DF14, DBAV14, DFFL10, DKR12, DZL13, DB12, DBHBB13, DF10a, DGAH10, DKW10, DGF14, DWZ10, DF10b, DE14, DZLD14, EMK11, EGL^{+}13, EWH13, EAN11, ELM14, EMSH14, FDHP11, FN14, FHMS11, FGR12, FT12, Fat10, FRP14, FF13, FL13, FF12, FH11, FY13, FR13, FGLB14, FK10, FP12, FP14, FSL10, FCY13, FS10b, GAvdVB14, GH14, GT12, GH11a, GHS12, GG10, Gao13, GLSC12, GPY14, Ge10]. method [GC11, GCZ13, GW11, GG14, GD10, GL12a, GLC^{+}11, GGH14, GP12b, GM14a, Gra14, GW10, Gri12a, GK10, GWK10, GTSC14, GP13, GPP11, GD14, GXX14, GLL14, HM11, HRK^{+}10, HS13a, HP13, HGRB14, HJ11, HKJ11, HJBB14, HHA14, HR10, Han11, HHS^{+}13, HD13, HBM13, HGW11, HNS12, HL12a, HLS14, HGMM12, HE14, HWST12, HRT11, HW10, HW11, HHGC14, HvB11, HM10, HVT12, HLZ14, HJK14, HL12b, HDD12, HWW10, HK11, HK13, HY09, HY11, HZ11b, Hu13, HAS13, HKL14, HKL10, HCS11, HCS12, HW14a, HHL14, Hys12, ID10, Ich13, iIT14, IX10, IST^{+}12, IXX14, IJZ13, IKO^{+}14, IG14, JMFO13, JNM12, JSA14, JYHT10, JMW12, JS13b, JZ11, JLZ13, JHZ10, JW11, JXD14, JWYH10, JZ10, Jin12, JK10, JM10, JG13, JK11, KCG14, KK14, KED^{+}12, KKLR10, KKAS12, KLMJ12, KVM12, KH13, KPH13, KR10]. method [KSU13, KKP10, Kaw13, Kea14, KAFB11, KF12, KT14, KG11, KG13a, KMA13, KLP10, KL10a, KLPS14, KBGRM10, KBGO13, KSB12, KHJ^{+}13, KMHJ10, KBSV14, KG13c, KL11a, Kri14, KTT13, KG12b, KL12b, KL12c, KBB13, LBC14, LH10a, LDS11, LP12a, LW11, LXM14, LP13a, LT11a, LB12, LRL10, LGP10, LKCY11, Lee12b, LY13a, LB13, Lee14a, LK14, LDM14, LBCL13, LTS^{+}10, LPE^{+}11, LKT^{+}12, LSBJ14, LGF11, LLZ11, LH10c, Li10b, LYE10, sLwG10, LTSH10, LLX11b, LCK11, LHY12, LX13a, LXL13, LHY13, Li13, LMK13, LMGS13, LC13, LJ13a, LH14a, LZZC14, LLP14, LX14, LCP13, LGG^{+}13, LSW14, LCE^{+}10, LSE13, LSZ13, LSV10, LMS11, LMBB11, LM14a, LMS14, LRK11, LC14, LLH10, Liu11b, Liu11a, LZV12, LW13b, LS13d, LW14, LCWN12, LM13, LGP14, LMS^{+}10, LLS11, LHV13, LCB12, LBK10, LLN^{+}10, LXL^{+}12, LXS^{+}13, Luo13, LFW11, LZZR10, LI14, MGJZ10]. method [MZ11b, MWP14, MRS14a, MF14, MS10b, MS14a, MK14a, ML11a, MHL^{+}14, MMV^{+}13a, MMV^{+}13b, Mar13, MMSI13, MMHA^{+}12, MSS^{+}14, MIJ10, MD14, MRD13, MTT13, Mel11, MS14b, MZ11c, MST14, MBL^{+}14, MDH^{+}10, MBA14, MESV10, ML11b, MTZ14, Min13, MVCFM13, MZB12, MDM^{+}12, MLBS14, MJ13b, ML14b, NSA11, NTV14, NHE14, NRS10, NPC11c, NLGL14, Nic11, NJ12, NB14, NKS10, NCV10, OAKR14, OA10, OHF12, OD14, PBT14, PS12, PZW14, PZG14, PMM^{+}11, PPB14, PRA13, PL10, PMSH10, PRL10, PPS12, PB13b, Per13, PIN14, Pir12, PH13, PMSM12, PBC11b, PA14, PB14b, Pri14, QY12, QC10, QZM^{+}12, RSK13, RBB12, RVB10, RCQL11, RTS13, RK14, RP13, RG13, RSS12, RVFK14, RS10a, RSS13, RLJ14, RHXQ14, Ren11, RGMC13, RR14, RC12, RCvdV13, RWMG11, RPA13, RMMD14, RM12, STD14]. method [Sal14, SS13a, SW11, SBN12, SAM10, SS11, SN12, SN13a, SF11, SGS13, SM13, SR13a, SGC13, SHCY13, SS10b, SD14a, SM11b, SM11c, Sha13, SSC13, SF14, SH11, SYY14a, SYY14b, STH11, SYL12, ST13, SJK11, SR11, SSB^{+}12, SS10c, SKCC10, SCC14, SPF10, SCH13b, SBA13, SS14a, SC11, SY11c, SY12c, SLH^{+}11, SHA11, Sog14b, SBvdV11, Son11, Sou14, SM10, SM12c, SMS14, SAA14, Sto13, Stu10, SSK^{+}12, SS14c, SdAW14, SL13b, SSY13, SD14c, SY14b, Tak14, TNK10, TK14, Tal13, TM10, TQW14, TPF11, TG14a, TF10, TSLV11, TDVG13, TVE14, TLZ^{+}11, Tit12, TCL11, TL13, TRL14, TE11, TVEC12, TH14b, TBR13, UWB12, VRB10, VP14, VB13, VHG13, VMS11, VMS12, VS14, VM10, VDP11, VHWF13, VdMS^{+}14, Vre14, Wan11a, Wan13a, WWS10, WXZ10, WS10b, Wan10, WW11, Wan11b]. method [WZ11, WT12a, WSS12, WYS12a, WWZ13, WW13, WD13a, WK13a, WLL13, WZG^{+}13, WD13b, WQCS13, WZN14, WZZZ14, WLW14, WDP^{+}14, Wan14, WC14a, WP10, WMYYG10, WRH10, WJTY10, WK13b, WY10, WY11, WC11, WSBG10, WZ10, WS10c, WTSZ10, WS11, WLPG11, WJC13, WWS^{+}13a, WTYC13, WYC13, WT14, XZW11, XZWW12, XFB^{+}14, Xie14b, XIIX14, XPO14, Xu11, XYL12, XMDG13, XMDG14, XW14, XR14, XCL14, XJ14, YMS10, YMS12, YYY^{+}14, YVL^{+}11, YDN10, YL11, YO11, YD11, YXZ13, YWS13, YK13, YM14, YZX14, Yao14, YMLM^{+}14, YM10, YHL13, YK12, YZ12, YW13, Yok13, YN14, YL14a, YL14b, YL10, YS13, YB14, YF11, ZM10, ZBS12b, ZFH14, ZW10a, Zha10a, ZOWZ10, fZlC11, ZHPS11, ZMG11, ZFJ11, ZOJR11, ZYHW13, ZYT13, ZLCW14, ZKZ10, ZIOF10, ZTL13, ZMD14, ZW10b, Zho10, ZLH12, ZL13, ZCS14, ZF14, ZTS^{+}10, ZWJW10, ZQ11, ZZSQ13, ZKM^{+}11, vHtTBC11, ABMT14, CF11a]. Method [Chi12, CM13, DDL13, GCBH10, GCBH13, HG10, MBR11, MNR11, MOSW12, MVS13, SS12, BMS12]. methodology [AS12, FW11, GCA13, LGE^{+}13, VDD10a]. Methods [KSK13, AJRT11, AP12a, AQ11, AVKP11, AV10, AL10, AK12, AJV11, AWK^{+}11, ABHS12, APY14, Atz10, Atz11, ABJ^{+}12, BMD^{+}13, BCW10, BD11a, BCZ13, BTX13, BLQ14, BR10a, BC10a, BGN10, BBCH^{+}12, BBG^{+}11, BDTW10, BdF10, BGHM13, BHK^{+}10, BP13a, BRSS11, BY11, Boy13, BF13, BHBM10, BCW13, BSV14, BC10b, BKV12, CP14, CHC11, CF11b, CCW11, CZ11a, CS12b, CCC12, CHKT13, CCK13, CY14, CLQX13, CCZ14a, CCZ14b, CH10, CSK14, CSX14, CCY13, CR13, CDHM11, Coo13, CCLV10, DL12, DLGP13, DG12, DAEB13, Don10, Don11, DL10a, DSZ13, DSHP11, DCVM11, DV14, EGW11, EGH13, EFT13, FKM11, FDHP11, FPN^{+}13, FL14, FL11b, Fuj13, GPCE13, GCE14, GM10b, GBS^{+}14, GH11b, GRS14, GKN^{+}11, GL12b, GJ11, GBC11, GD13, GQ13, GZQ13, GH10, GWCA14, HPV13, HCD11, HMS10, HL12a, HRBK10]. methods [HLSO13, HTHG14, HLW13, HDD14, HLY11, IF10, JWNL11, JC13, Jar12, JKW11, JPC^{+}14, JLZ13, JC10, JZ13, JZ14, JLB^{+}10, KL14a, KG12a, Ket10, KAB^{+}14, KA13, KL11b, KQ11, KGSS10, KS12, KS14, KHHK12, Kuz14, LSM14, Lan14, LH13, LKP14, LOG12, Li10a, LXY11, LX12, xLqW13, LLW14, LGLX14, LYJ14, LLMY14, LT11b, LSVW10, LN10c, Liu13, LP14b, LQB14, MC12, MMP11, MNF^{+}10, MJ13a, MMIW14, Mig14, MF10, MSS12, MK14b, MIKG12, MNL^{+}13, MLM12, MWW^{+}13, MWY14, NV14, NPC11b, NPC11a, NP12, NN12, NF12, OJW10, OS11a, OS14, PDZ^{+}14, PVLT10, PWK11, PB14a, PMMB14, PNPF10, PS14b, PHSA14, PGA11, PPWZ11, QS11a, RT14, Ran11, RSS10, RHB12, RBCK10, RS10b, SK12a, San13, SM12b, SJB11, SHMS13, SRS10, SRSV11, SPB^{+}10, SWX12, SY11b, SSFM12, SC10, Soa13, SK12b]. methods [SC12, SR13d, SdAW14, ST11b, TSL^{+}14, Tok11b, TD11b, TS14, TLNE10, UJvL10, UJ11, UJ12, VGCMG11, VLM13, VWPF11, VWF13, VBCM10, VSC^{+}11, VG14, VSK10, Wal13, WK10, WPP10, WSYS12, WD14, WW14, WX14, WKH10, WPK14, XXS10, Xin14, XL11, XLD^{+}11, XS13, XST13, XH13, XH14, YC11, YF14, YKWS13a, YKWS13b, Yok13, YBK13, YWL14, ZT10a, ZK14, ZF12, ZWL^{+}12a, ZWL^{+}12b, ZHY13, nZzSlL14, Zha10b, ZOL^{+}11, ZT13b, ZHS10, ZS13, ZT10b, ZWL10, ZMD^{+}11, vDS12, vRLPK11, ABH^{+}14]. Metric [Cou11, Del11, SYV17, XY14, AINF13, LLS11, SYV14]. Metric-based [Del11]. metrics [ANIF14]. Metropolis [HJBB14]. Metzner [GDW11]. MgSiO [Ber14]. MHD [HY11, BMP14, BDA14, CNSFD13, CLQX13, CH11a, HY09, HY10, Li10b, LXY11, LX12, LT12, LL10b, MGLS12, MTB10, MT10, NL12b, PBC11b, SPB^{+}10, SC13, SIDG13, TG14a, XPO14, ZN14a, ZN14b]. MIB [NZ14, XZW11]. micro [BLR13b, DLNN12]. micro/nano [BLR13b]. microchannel [LVZK14, PL14, Tit12]. microchannels [WK10]. microfilaments [LG14]. Microlocal [Vai14]. micromagnetic [ES14]. micromagnetics [GHDS10, KVBP^{+}14]. Micromechanical [CMBW13]. micron [LL10a]. micron-scale [LL10a]. microphysical [OBNN13, OBNN14]. microscale [LXM14]. microscopic [Mit13]. Microscopically [PP12]. microstructural [JR14]. microstructured [GG13]. microvascular [ASGW11, ASK^{+}13]. Mie [WP10]. migration [TCW^{+}10]. migration-diffusion-convection-reaction [TCW^{+}10]. million [Liu11a]. millions [JMKK12]. Mimetic [BBM14, KTT12, LRN13, LMS14, KG13c, LS10b, LMS11, LMBB11, LM14a]. Minimal [JNX13, LRS10, TR11, WA12]. Minimising [GBB^{+}13]. Minimization [Oli14, BK12b, CLW10, CGS13, PR11, PHD14, Xie14a, YK13, YWY10]. minimized [HT12, sSLxRyZ14]. Minimizing [AH13b, CCW12, Kol11]. Minimum [CJM13, Ket10, Wan11a, Wan13a]. mining [AC11]. mirage [FGE11]. Mitigating [MMHA^{+}12]. mitigation [VGCMG11]. Mixed [CS12c, DL13c, Hys12, KG13c, AGBC14, AM12, BLRR11, CCB12, FCG10, FMT11b, GH11b, HLYS11, JM10, MZ11b, MMP11, MBA14, SM13, SKt10, TCL11, TTD11, HPV13]. mixed-dual [BLRR11]. mixed-element [TTD11]. mixed-precision [CCB12]. mixing [CS12a, JNW^{+}14, MJ13a, RS10b, VWP11]. mixture [MFM12, PS14a, PS15]. mixture-energy-consistent [PS14a, PS15]. mixtures [DGF14, Gio10, LC13, Liu11a, LR13, PvDtTB^{+}11, VP10, VP14]. mKdV [GAV13]. MLPG [LCK11, WTSZ10]. MLPG R [SM12c]. mobile [RB13]. mobility [CGC13, JJM13]. modal [BGHM13, PMSM12, PS14b]. Mode [GPCE13, GCE14, GKS^{+}11, KG12a, LL12b, LL14b]. Model [ECD14, HH13, And13, AC11, ABPSM11, BNT14, BHS14c, BBCH^{+}12, BLM^{+}14, BND^{+}12, BDH13, BCL13, BSDM12, BWMG14, BCL^{+}11, BKW10, CFCA13a, CFCA13b, CPJ11, CEK14, CM14, CF10, CBW10, CSW14, CGR11, CLG13, CH10, CGJ11, CF11c, CDS13, CBCM^{+}11, CDB13, Dad12, DTG^{+}14, DDS12, DXB^{+}13, DFW^{+}10, DRZ14, DKA12, DEKBF14, ECDB14, EGG12, EWH14, FG12, FCG10, FCH13, FK14, Fuj11, Fuj13, GW12, GW14, GDW11, GHM10a, GHM10b, GM11a, GHS13, GHP13, GVV13, GK14, GGT13, HKJ11, HHA14, HHL11, HVD13, HZS12, HMOH11, HW14b, HCM10, IX10, JJM13, JTH10, JG13, KL14a, KG12a, KRF12, KGG10, KL10b, KO12, KK10c, KBB13, LGH10a, LMFH11, LBT11, LHH14, LKT^{+}12, LJX^{+}14, LGLX14, LZC^{+}13, LWZ^{+}13, LTC14, LRBT14, LMM14, LBZ^{+}10, LFLV14, MZ10, MZ11a, MS14a, MWRZ13, MPWK13, Mit13]. model [MSTL14, MMT14, MVG11, MGN10a, NL12a, NWLS12, OHF12, ORHH13, PZTW12, PTM13, PBC^{+}11a, Par10, PBC^{+}14, PS14a, PS15, Pie12, RVFK14, RJC13, RBNS13, RCCC14, RG12, SBWF14, SM11a, SN13a, SMP^{+}13, SH14, Sch13a, SO13, SC13, SSW13, SRS14, SKR14, SN13b, SS13c, SV10, Sza12, TBMH11, TFP13, TS12, VP10, Vre11, WZ13, WZ14, WT12a, WMK13, WSW13, XHF13, XFB^{+}14, YP13, YCH^{+}13, Yok13, Yok14, YN10, ZA14, ZFJ11, ZYHW13, ZMQ13, Zho14a, DK14, HH12b]. model-order [DRZ14]. Model-Scenario [ECD14]. modeled [KRT14]. Modeling [ASXZ14, BI11, BM11, CW10, ES10, IBO14, KL12a, KST10, LZV12, LBK10, MGLS12, OLC13, VLM13, VKBK11, WK10, Wan10, YFL^{+}13, YXD^{+}14, ZHW10, AH13a, AL10, AAG14, ACCM12, BWS10, BC10a, BKKD14, BGBR12, BBD11, BHBM10, CRS14, CCMG^{+}13, CW14a, CL11a, CMG^{+}13, DST14, DKR12, DL10a, DGAH10, FT11, GSTB13, GJLY13, HNMM10, HGW11, HSK^{+}15, HBZL14, HH12b, HH13, HLZ14, HHH13, ICC^{+}10, KBTD12, KLB^{+}14, KEGM10, KKC^{+}13, Kwo14, LRT13, LB13, LLZ11, LYE10, LHY12, LLZ14, LG14, Lip12, LZ10, LVZK14, LP11, LM14b, LMQ^{+}11, MB13a, MCA^{+}13, MBMV13, MBL^{+}14, MF13, MGN14b, NZ14, NKF14, OD12, OP10, PLT14, PM14, RG13, RKGM14, SG10b, SG10a, SMF^{+}13, SB12, SKt10, TH14a, TTR^{+}12, Tok10, TvdHS^{+}12, VGCMG11, VHWF13, XFC^{+}14, XJ12, Zad11, ZM14, ZZK^{+}14, ZKM^{+}11, dv23]. Modelling [KSD12, RBB12, RKD^{+}14, STC10a, ERS14, FGE11, GM11b, HCD11, JR14, JW12, Nag14a, Nag14b, PLBR13, SMS14, VBVD10, vDS12]. models [AVKP11, AP12b, AELV14, Ala14, AA13, BF14, BDTW10, BS11a, BN14b, BJLR14, BM13a, BDMV11, CIJS14, CD10, CVC12, CNR10, CLSX14, CJP^{+}14, CP13b, CB13, DLM^{+}13a, EWH13, EEK14, FY13, FvdB12, FQV13, FP12, GHM10b, GH11b, GRS14, GP12a, Gri13, Gri10, Gri12a, Ham10, HMM14, HBAP10, HSD11, HLLW11, ISZ12, IWG13, KH13, KVC^{+}13, LXM14, Le 12, LTS^{+}10, LX10, LC13, LY14, LE12, MZ11d, MZ14, MJ13a, Mie13, MPPP12b, MPPP12a, Min13, MY14, NMX12, NB13, PBWK11, PR11, PRA13, PBM11, RV10, RE13, RPA13, SLGB12, SS14a, TQW14, TPF11, Tau12, TAT14, iT14, WABI11, WT12b, WMYYG10, Wei12, YWS13]. moderate [MCA^{+}13]. moderately [NS14a]. modern [YS12b]. modes [PBdGP14, RWX11, Tan11]. modification [LGP14]. modifications [HT14, MRS14b]. Modified [CWC11, SOG14a, SB12, BHL^{+}13, CG13b, Gür14, HG10, HW14b, JKQ13, KKS13, KQ11, LCK11, MAD13, SC10, WV14, ZMG11]. modular [MVCFM13]. modulated [IBO14]. modulation [ÁCB12]. MOF [JLZ13]. moist [MMV^{+}13a, Ste14]. Molecular [ARF12, LWCR10, BB12a, CD12a, CWC11, CLW10, Dad12, Dav10, Dav15, DXB^{+}13, GW11, Gen13, GL13, JTH10, JFC^{+}13, KSS14, KM11, KH11, KIF^{+}14, KAG14, LH13, LB13, LSK11, LYXZ14, MB13b, PLBR13, PA12, PA14, STC10b, VWP11, ZKM^{+}11]. molecule [HL11]. molecules [Son11]. Moment [TBMH11, AIX10, AHOT14, CLSX14, DKR12, GCA13, HPS13, IAM^{+}12, IX10, JSA14, KLMJ12, MMR12, MT13, MMA11, MMA13, NF12, PA12, UWB12, VLM13, VWPF11, VWF13, XIIX14, Yok13, HS13b, WP14]. Moment-Based [WP14]. Moment-of-Fluid [HS13b]. moment-preserving [MMR12]. moment/multi [VLM13]. moment/multi-fluid [VLM13]. moments [VHWF13, YF11, ZKKF10]. Momentum [IF10, BNT14, CWL13, CD12a, LPE^{+}11, LKT^{+}12, sLwG10, LSVW10, SLI10, SSR14, Tok11a, VKL^{+}13, vV12]. momentum-preserving [CWL13]. momentum-space [SSR14]. Monge [BFO14, BCW13, CDF11, CSC14, FO11, ZPF10]. mono [BM10a, Gra14]. mono-dimensional [Gra14]. mono-energetic [BM10a]. monolithic [APF13, BC10a, GM12a, KN14, NKS10, RMSF11]. monolithically [MMH12]. monolithically-coupled [MMH12]. Monoslope [BC10b]. monotone [LSV10, ML13, SY12b, WG14]. Monotonic [AT13a, ZSW10]. Monotonicity [BKV12, LP13b, LMS11]. Monte [Den11b, DBO^{+}10, Gho17, HJBB14, LC13, MSS^{+}14, WPP10, AB13, AJG^{+}13, AKP^{+}12, BDL11, BHM11, BHM13, BP13a, Boo11, CRS14, CGP10, CBGK13, DD12, DWL10, DTU12, EMSH14, GCA13, Gen11, GKN^{+}11, HH12b, HH13, JNSA12, KSS14, LWZ^{+}13, MMM11, MOV12, MSS12, MVG11, MJM13, MMJ14, NKSR12, NVS13, PBC^{+}11a, QA11, RLJ14, RSFS13, RRD^{+}14, SBC^{+}14, SJB11, Sch13a, SR13a, SD14a, SD14b, SSFM12, SSR^{+}13, SSK^{+}12, TSR13, WK13b, WP14, XX12, ZKZ10, ZKKF10]. Monte-Carlo [DD12, ZKKF10]. MOOD [CDL11]. Moreau [HHK13]. morphodynamic [Zho14a]. morphologies [KIG11, ZCD10]. Morrison [CP14]. Morse [ZZK^{+}14]. mortar [Han11]. mortar-discretized [Han11]. MOSFET [JC10]. motility [WZ10]. motion [Bea12, BBGP14, CDF11, Dad12, ERT10, HGW11, JTH10, KBTD12, LLZ11, MFM12, RUL^{+}13]. motions [LBRA^{+}13]. motors [LKP14]. movement [Ich13]. mover [CC13]. movers [DC13]. Moving [SYV17, TA13, AINF13, ANIF14, ABMT14, BSSW12, BBCH^{+}12, BKC^{+}11, BCW13, CCFCM11, CXLC12, CB10, DDF14, EQYF13, ES13a, FRP14, GW12, GW14, GGH14, HZ11b, HF10, KKAS12, KG11, KSD12, LCNK10, LKCY11, LY13a, LMZ14, LH14b, MKCB14, MRD13, MVS13, MZ11c, MK11, MK12, MK14b, NS14b, PZ10, Qua11, SHMS13, SYV14, SS13c, TS11, VSC^{+}11, WYYK10, WK14, WZ10, Xin14, XR14, YCY11, ZBS12a, ADB14]. moving-body [LKCY11]. Moving-Least-Squares [ADB14]. MPFA [FE11]. MPI [KEGM10, LS10a]. MPI-only [LS10a]. MPS [KG13a, TM10]. MRT [XMS12]. MRT-LBM [XMS12]. mud [MGN12b]. mud/debris [MGN12b]. Mullins [BGN10]. Multi [APGL12, ASGW11, ASK^{+}13, BILM14, BZ12, BZKL13, ERS14, FK10, Ger12, ILM^{+}13, IDNG13, KG13b, MMA11, MMA13, MSS12, NF12, PYK10, SABH11, SHZ13, XHF13, AHA10b, AIX10, AKL^{+}13, ABS^{+}14, BDL11, BD10b, BGGW13, CCD11a, CWS13, Cap11, CHKT13, CNST13, CLSX14, CY14, CS14a, CB13, DM13, DLM^{+}13a, DBHBB13, DBM^{+}12, EN14a, EN14b, Fal13, FGR12, FK14, GMB10, GG13, GCW14, GH12b, HHA14, HLM11, He13, Her10, HJZ14, HDD12, HK11, HDHZ13, HQL^{+}10, IX10, IST^{+}12, JNX13, JS13b, JHF10, JLZ13, JFC^{+}13, JH12, KED^{+}12, KKL10, KSU13, KN14, KBSV14, KGSS10, KS12, KS14, LMLC12, LNU10, LB12, LRLL11, LR12, LS10a, LX13b, MF14, MK14a, MS12, MHGM14, Nor14, PBWK11, PvDGvM11, PvDtTB^{+}11, RARO13, STD14, SBN12, SCGE13, SQC11, SWX12, SPF10, SY11c, TVE14, TRL14, UWB12, VF10, WS10b]. multi [XZW11, XIIX14, XL11, XH14, Yok13, ZN14b, ZLH12, CDL11, ILM^{+}13]. multi-block [SY11c]. multi-body [SCGE13]. multi-component [BD10b, HK11, PvDGvM11, PvDtTB^{+}11]. Multi-dimensional [Ger12, IDNG13, PYK10, SHZ13, BDL11, Cap11, Fal13, GH12b, IST^{+}12, KKL10, LRLL11, LR12, Nor14, RARO13, CDL11]. multi-dimensions [LB12, MSS12]. Multi-domain [SABH11, BILM14, DBHBB13, HDD12, LMLC12, XH14]. Multi-element [FK10, KG13b, JNX13]. multi-energy [MHGM14]. multi-flow [ZLH12]. multi-fluid [DLM^{+}13a, VLM13, VF10]. multi-GPU [JFC^{+}13]. multi-ion [BGGW13]. multi-layer [CCD11a]. multi-layered [ABS^{+}14, STD14]. Multi-level [BILM14, MSS12, JS13b]. Multi-mass [APGL12]. multi-material [CS14a, FGR12, FK14, GMB10, JLZ13, JH12, KSU13, KBSV14, KGSS10, KS12, KS14, LX13b, XZW11]. multi-medium [WS10b, XL11]. Multi-moment [MMA11, MMA13, NF12, AIX10, CLSX14, IX10, UWB12, XIIX14, Yok13]. Multi-output [BZ12, BZKL13]. multi-particle [GG13]. multi-phase [AHA10b, CY14, HHA14, HQL^{+}10, MF14, PBWK11, SPF10, ZN14b]. multi-physicochemical [CHKT13]. Multi-physics [ASGW11, ASK^{+}13, CNST13]. multi-point [SBN12]. multi-resolution [DM13, HHA14, TRL14]. Multi-scale [XHF13, AKL^{+}13, CHKT13, DBM^{+}12, He13, Her10, MK14a, SQC11]. Multi-scenario [ERS14]. multi-socket [LS10a]. multi-symplectic [CWS13, GCW14, HJZ14, HDHZ13, SWX12]. multi-target [MS12]. multi-terminal [TVE14]. multi-time-step [KN14]. multi-tracer [EN14a, EN14b, HLM11, LNU10]. multi-water-bag [CB13]. multi-wavelet [JHF10]. multiblock [ADSF13]. Multibody [LOG12]. multicanonical [SJB11]. Multicomponent [CGF^{+}13, Gio10, BP13a, CS12a, CC14, KCG14, LI14, NMT^{+}12, SMP^{+}13, Shu14, STL^{+}10, WP10]. multicore [HRBK10, LS10a, LBRA^{+}13]. Multidimensional [Bal10, BDA14, Bal14, BBD14, BKW10, CDF11, GV13, KSK13, MCD14a, MPWK13, RS14b]. multidomain [BDLGC13]. multifluid [MJ13b]. multifluids [MESV10]. Multifractal [RG13]. Multifrequency [She13, SVS11]. Multigrid [Ge10, GC11, PS12, PPB14, SEDW13, vdVR12a, vdVR12b, ASB10, BHL^{+}13, BGR11, BWC13, BM10a, CF10, CR13, GCZ13, GW10, GT11, GMG12, GMXG14, HSN13, JG13, Lan14, MHA10, NOT14, OS11a, PVLT10, SSG11, TCW^{+}10, WMYYG10, Xie14b]. multigrid-based [BWC13]. Multigroup [CGP10, RF14b, Ols11]. multilayer [ABPSM11, BOK12, DLM13b, FNKdLB13, SYL12]. multilayered [DB12]. Multilevel [KKPV12, KLL13, MJM13, RRD^{+}14, vdVR12a, AB13, CMS10a, CLX14, LLW14, MIJ10, MBLM10, MSD12, MMJ14, YC11]. multimaterial [BHS14c, BBG^{+}11, SKR14]. multimaterials [GIMT14]. multimesh [DSHP11]. multiobjective [ZM14]. multiparticle [HCS^{+}10]. multiphase [BDH13, BHL12b, CG13b, DFVY14, FP12, Fus13, HHL14, JSA14, KG13a, KG12b, LDM14, LHB10, MT14b, OD13, Qua11, SYJ11, Shy10, VDD10a, Vre11, WWZ13]. multiphysical [Mön13]. multiphysics [BP13b, CNSFD13, TQW14]. Multiple [YN10, ZFJ11, AP12a, AV10, ABH^{+}14, BFE12, CB12, GHS12, GCvRK11, GL12b, GHL14, GS11b, GGG^{+}13, HNMM10, HW14b, KSPP10, KL11b, LRT13, Lee12b, Li12, xLqW13, LW13a, Lip12, NMX12, Qua11, RM14, STC10a, SSK^{+}12, VBVD10, WW14, dv23]. multiple-cell [Li12]. multiple-frequency [GGG^{+}13]. Multiple-relaxation-time [YN10, HW14b]. multiple-scatterer [VBVD10]. multiple-time-scale [HNMM10]. multiple-time-step [RM14]. multiplication [LLY11]. multipliers [RHGT10]. multiplying [RF14b]. multipoint [APY14]. Multipole [DDL13, HW10, HW11, Lee12b, LOG12, LGG^{+}13, LC14, NN12, PA14, Tak14, ZHPS11, CB14]. multiprocessors [NS11b]. multiquadric [BMK11, BMK12]. multirate [BM10b, SLT^{+}14]. multiresolution [ACG14, dlLPC10, PA12, RCW11, RS10b]. Multiscale [CEGG14, EGW11, GSTB13, GW11, HHH13, JTT14, KF13b, KLL13, PL14, Ste14, TL13, WPKK12, WMK13, XFC^{+}14, AB12, BLQ14, BGGW13, BA10, BILM14, BDM13, BLR13a, BLR13b, BM13a, BSS11, Chi12, CJ14, DDM10, DM13, EGG12, EGH13, EGL^{+}13, ETW10, GW10, GFK13, GWK10, HJ11, HKJ11, HLSO13, JK10, KL12b, KKC^{+}13, LB13, LYE10, LDDBR13, LTL11, MZ11b, MCJ14, MKGV12, MMV^{+}13a, MMV^{+}13b, ML14b, OK12, OK13, PLBR13, RUL^{+}13, RG13, SKN13a, SMF^{+}13, Sco12, SC13, VBCM10, WZ13, WHT14, YCH^{+}13, ZOWZ10, ZOJR11, ZYHW13, ZHS10, HPV13]. multiscale-multigrid [GW10]. Multiscale/fractional [JTT14]. multislope [BC10b, CRT10]. multistep [CK11, LBM10]. multisymplectic [KHZ10, ST11b]. multitude [EQYF13]. multivariate [DKR12, WMI10]. Multiwavelet [VR14]. MUSCL [BD14a, BC10b, CCKCG10, CRT10].

N [Gho17]. Naghdi [BCL^{+}11, LGH10a, LGLX14, PDZ^{+}14, PE10]. nano [BLR13b, CW10, HZS12, JC10, KKPV12, KA13]. nano-crystals [KA13]. nano-electronic [CW10]. nano-pattern [KKPV12]. nano-plasmonic [HZS12]. nano-scale [JC10]. nanocarrier [RUL^{+}13]. nanochannels [LWCR10, ZKM^{+}11]. nanometers [NIN^{+}14]. nanoparticles [LH13]. nanostructures [NIN^{+}14, TVE14, XX12]. nanowire [RSK13]. narrow [LDM14]. narrow-band [LDM14]. NASA [KYP^{+}14]. Natural [KTT12, ZYT13, DMHL14, EBGW13]. natures [SCY11]. Navier [GLL14, HHK13, KRF12, VV13, AB11, ABPSM11, BK11, BSSW12, Bea12, BNNW11, BN11, BN14a, BGHM13, Boe11, BR12b, BCLR10, CVN13, CMS10a, CJNS12, Cha13, CXLF14, CLS10, CDK12, CST12, DLZ^{+}11, DHH11, DAA11, EH13, FW12, FL11a, HGW11, HL12a, HOK10, HM10, HKL14, HF10, JM12, KKP10, KF13a, KRT14, LXF10, LGE^{+}13, LLN^{+}10, MBS13, NPC11c, NOT14, NDBG14, NEE12, OA10, OR10, OS12, PBT14, PVLT10, PGL11, PNPF10, PGA11, RP13, RHXQ14, RGMC13, RCvdV13, RPC14, SK12a, San13, SVK14, SH14, SSHT14, SM13, Sen13, SSS10, Sha13, SLI10, SR11, SY11c, SGV13, TWMM14, TY11a, TLO^{+}14, UN13, VQRD11, VT13, Vre14, WW13, XFB^{+}14, XIIX14, XH13, ZW10b, ZLH12]. near [Alb10, BdCB11, Boy11, DDN^{+}10, GDY10, Mac10, RKM11, Vre14]. near-equivalence [Boy11]. near-initial-information [RKM11]. near-wall [GDY10]. nearly [KGG10, ML11b, ZA14, ZJW12, ZWH^{+}12]. nearly-orthogonal [ZJW12]. necessary [PBWK11]. Nédélec [HZS12, JMFO13]. need [PM14]. Negami [CPJ11]. negative [DR11, MMR12, MM12, NMV13]. neighbor [Mac10]. nematic [BGGGS11, CTSM13]. NEMD [GWK10]. neon [KM14b]. Nernst [LW14, LHMZ10, MG14b, ZCW11]. Nested [PMSH10, EWH14, SY14a]. network [BPS14, CCT11, Ham10, XHF13, vTCL^{+}10]. networks [AM13, BK14b, CMBW13, FLM12, Gui12, KL14b, MAK14, NJ12]. Neumann [AV10, Deg11, FN14, FN10, HG11, LL12b, MP13, RzSZ13, SSC13, TDL10, VCJ11, VW14b, ZzS11]. Neumann-to-Dirichlet [LL12b]. neuronal [CCT11]. Neutral [KM11, GH12b, WAM14]. neutrality [CRS12]. neutron [BLM^{+}14, DSHP11, JC13, Tal13, WPK14]. Newton [BDTW10, CGG12, CCL^{+}14, DMHL14, GL12a, Han11, HL12b, JG13, KL11a, LTS^{+}10, LPE^{+}11, LKT^{+}12, LFLV14, LvZB10, LL10b, NOT14, PMMB14, YJK12, vVVK10]. Newtonian [BNM14, KRT14, LXM14, PTM13, STH11, ZYF^{+}10]. Nicolson [ÇD12b, FF12, WXY13, YXZ13]. Nicolson/quasi [YXZ13]. Nicolson/quasi-wavelets [YXZ13]. nine [WDGY10]. nine-point [WDGY10]. Nitrogen [MPPP12b, MPPP12a]. Nitsche [AM11]. No [RPV13, SR11, VM10, Vre14]. No-slip [RPV13, SR11, VM10, Vre14]. nodal [DBZ14, DTYY14, LH14a, OS11b, ZS14]. nodal-continuous [DTYY14]. node [AT13b, BvBZ^{+}10, FL10, HWST12, LR11, SSHT14, TG14a, WZ10, ZWH^{+}12]. nodes [MGJZ10]. Noise [TAT14, AV13, AP11, DG12, HJZ14, Jin12, SSS11, VS14, vydCNS^{+}10]. Noiseless [PSF^{+}14]. nominally [MB12, MAB^{+}13]. Non [AM12, DZL13, ES14, FPN^{+}13, GGG^{+}13, LSW14, PL10, PL12, XFB^{+}14, AK10, AC13, AW14, ABJ^{+}12, BD10b, BNM14, BRS11, BSS12, BD14b, BAG12, BCG^{+}13, Cap11, CES14, CCD11b, CF11a, CRG13, CVIA10, CRT14, DR11, DB13, DO11b, DEKBF14, FSTY14, FCYF11, FVR13, GG10, GCvRK11, GP12a, Guy13, HKLY13, HG11, HWA10, IAM^{+}12, IG14, JC13, JN12, KIF^{+}14, KRT14, KG12b, LH10a, LXM14, LRL10, LQ10b, LWF11, LT11c, MMR12, MC12, MBLM10, MN12, MM12, MTG11, MM11, NMV13, NKS10, NNS11, PTM13, PDZ^{+}14, PL14, PT12, PWB^{+}14, Pir11, SBWF14, SG12b, SSM14, SBS11, SML13, STD^{+}10, SMS14, SJ10, SGV13, UJ12, Vai11, VDP11, WYS^{+}11, WLL13, WS14, YF14, ZN10, nZzSlL14]. non-adapted [DO11b]. non-breaking [SMS14]. non-collisional [LWF11]. non-compactly [Vai11]. Non-conformal [DZL13, PL10]. non-conforming [NKS10]. non-conservative [AK10, BD14b, YF14]. non-deteriorating [PT12]. non-dispersive [BRS11]. Non-dissipative [LSW14, Pir11]. non-divergence [GCvRK11]. non-elliptic [SG12b]. non-equilibrium [CVIA10, IAM^{+}12, MM11, PWB^{+}14, SGV13, WYS^{+}11]. non-equivalence [SSM14]. non-Fickian [STD^{+}10]. non-graded [HG11, MTG11]. non-hydrostatic [NNS11, UJ12]. non-isothermal [PL14]. Non-linear [FPN^{+}13, XFB^{+}14, AC13, BD10b, Cap11, DEKBF14, FVR13, Guy13, MN12, PDZ^{+}14, SBWF14, VDP11]. non-local [BSS12, GP12a]. non-negative [DR11, MMR12, MM12, NMV13]. non-Newtonian [BNM14, KRT14, LXM14, PTM13]. non-orthogonal [CRG13]. non-oscillatory [CCD11b, CRT14, DB13, FSTY14, FCYF11, HKLY13, HWA10, IG14, JN12, LQ10b, MC12, YF14]. non-overlapping [BAG12, JC13]. non-periodic [KG12b, LRL10, WS14]. non-premixed [GG10]. non-reactive [KIF^{+}14]. non-smooth [ZN10]. non-spherical [SJ10]. non-staggered [SML13]. non-stiff [CF11a]. Non-uniform [AM12, ES14, GGG^{+}13, AW14, ABJ^{+}12, CES14, LH10a, LT11c, MBLM10, SBS11, WLL13, nZzSlL14]. non-zero [BCG^{+}13]. nonconforming [HR10, ZW10a]. nonconservative [CG13b]. Nonequilibrium [CD12a, GWK10, PPWZ11, SAHS14, ZFJ11]. nonhomogeneous [BHM13]. nonhydrostatic [KG12a, MMV^{+}13b, SSW13]. Nonlinear [APRS14, CFW^{+}13, EMK11, LDN^{+}10, Shu14, WFK11, YP13, AMDR10, AH13b, APC11, BF14, BD11a, BTX13, BHL^{+}13, BBC^{+}13, BS11a, BCL^{+}11, BDLGC13, BGM12, CCT11, CWL13, CEGG14, CFCA13a, CFCA13b, CRS14, CZ11b, CGJ11, CV10, CF11c, CSKP10, CLS13, DBAV14, Don10, EO11, FD11, FF12, FCN^{+}13, FC13, FLB^{+}12, GAvdVB14, GJ14a, GCLS11, GL14, Gno14, GHP13, GJ11, GPP11, HMM14, HDHZ13, HM13, JZ13, KK10a, Kri10, LTCN13, LLZ14, LBZ^{+}10, LM14b, LL10b, MTG11, MSS12, MAD13, MY13, MN14, NMZC13, NMCZ14, NMT^{+}12, PR11, PMMB14, RSS13, SK13, SKN13b, SCS14, SN13b, SC12, TAT14, TLNE10, VK12, VMS12, VED13, WABI11, WXY13, WGX13, WT13, WXY14, WL11, XST13, YL11, YD11, ZN10, Zha14]. nonlinearity [BX12, DYYA14, Vai11]. nonlinearly [CCB12, CC13, JXD14]. nonlocal [BS11a, CM10, GLWW14, HZS12, WRH10]. nonnegative [BMR14, LHV13]. nonoscillatory [Fan14, ZS13]. nonparametric [CZ13]. Nonreflecting [CT10, JKR14, GS11b, Vai14]. nonreflective [Man12]. nonsmooth [YC11]. nonstandard [EO11]. Nonuniform [LLH10, FS10a, GC11, GCZ13, TD12a, YMS10]. norm [CJM13, HPA14a, MA13, Mat14, MAC14, XY14, ZY13, ZMD^{+}11]. normal [BDM13, ELM14, HLY12, IKO^{+}14, Vol14]. nose [NHE14, CWC11]. Note [Ano13-53, KGKS13, Try12, ZT10b, CFKL11, Don11, Gon12, SSL13, iT14]. Novel [FL13, ADSF13, BF13, GMXG14, JNM12, JMW12, LZZR10, TAAY14, YCY11]. NRxx [CL14a]. NSCBC [CGF^{+}13]. nuclear [NB13, SSFM12]. nucleation [LCE^{+}10, SJB11]. nucleus [ZCD10]. null [HD13, DE14]. null-field [HD13]. number [APC11, BFSVC14, CJNS12, DLDL10, Del10, FF13, FK14, GW10, HRK^{+}10, HGSK^{+}13, HKM12, Kwo14, Lee12a, LHB10, MKB^{+}12, MBD13, Mor10, NL12b, PAG11, Rie10, Rie11, SRN10, SM11b, SCB10, TGGT13, YYB12]. number-density [LHB10]. numbers [CBGK13, CLAL14, HY09, HY11, MCA^{+}13, MGN14a, WW13, vRLPK11]. Numerical [AP12a, ÁCB12, ABHS12, BD11a, BTX13, BBCH^{+}12, BFO14, BRS12, BJLR14, BX12, BKW10, BLY10, CCD11a, CWS13, CCMG^{+}13, CM13, CM10, CH11b, CB13, CG13b, CTSM13, DHN11, DDS12, DL10c, FJ12, GDW11, GPS10, GV13, GVH14, GAW^{+}14, Gür14, Har11, HL12a, HBZL14, HZS12, HH12b, HH13, HWW10, JWV12, JWNL11, JEX10, JPC^{+}14, JZ14, Kat14, KDS14, KJWR14, Kha13, KLP10, KLPS14, Kwo14, LKP14, LCY11, Liu13, LP11, LM14b, LT11c, MB13a, MYM12, NMT^{+}12, OLPM14, PPTMDK13, PBdGP14, SEBG12, Tal13, Tan11, Tok10, Tok11a, ULS13, VGCMG11, VWP11, Wol12, XST13, YJK12, ZYLK13, Zho14b, Zoh13, ARR13, AR14, AME14, Atz10, ABPSM11, BHB11, BVM14, BCW10, BSSW12, BCZ13, BLM^{+}14, Bér11, BS11a, BBGP13, BR14, BN14b, BRSS11, BS10b, BG13, BHBM10, BCW13, CCT11, CIJS14, CCFCM11, CX13]. numerical [CGM^{+}12, CGC13, CDT13, CM14, CWL10, CHKT13, CPX13, CSW14, CL11a, CVIA10, CH10, CSK14, CH14, CLS13, CS12c, CGC^{+}12, CGF^{+}13, Dai13, DLDL10, DLGP13, DBAV14, DST14, DFW^{+}10, DWZ10, DBM^{+}12, DE14, DLM13c, EH14, EH15, EL11, Era13, EC13, FBM14, FNGV14, FGLB14, FW11, Gao13, hGzSwZ14, GMT10, GM10b, GKG^{+}13, GV14, GHM13, Gui12, GLL14, HGW11, HLS14, HVD13, HM10, Hys12, JS13a, JTT14, JLB^{+}10, JH12, KL12a, KKP10, KSP13, KL10b, KAB^{+}14, KK10c, KRT14, KL11b, LTC13, LFL11, Le 14b, LGH10a, LBCL13, LP12b, LTS^{+}10, LCE^{+}10, LHB10, LW13b, LWS14, LMQ^{+}11, LHV13, LFW11, MDPTK13, Man12, MMSI13, MB11, MdVBS13, MTT13, MS14b, MVCFM13, MWY14, MPT13, MGN13, MGN14b, Nag14a, Nag14b, NMV13, OD12, OA10, OLFS13, OBT12, Par10, PBB14, PE10]. numerical [PS14a, PS15, PPWZ11, Pri14, Pud11, RTS13, RZ10, RHB12, RB10, RM12, SOG14a, SBB13, SNS10, SRSV11, SB14, STH11, SY11b, SJK11, SCS14, SL13b, ST11b, TiI11, TS10, TSL^{+}14, Tau12, TA12, TDVG13, TTR^{+}12, TvdHS^{+}12, TE11, TR11, VBVD10, VP10, VP14, VLM13, WFK11, WWX^{+}10, Wan11b, WLW14, WC14a, WRH10, WJC13, WWS^{+}13a, XS15, XPO14, Yok13, YL10, ZK13, ZVNM10, ZYHW13, ZPF10]. Numerically [GKF11, KHJ^{+}13, LS13c]. Nunziato [FP12, PZTW12, TT10]. NURBS [CXLF14]. Nyström [BG12, HK14, YHM12].

OBC [Ano10-49, Ano10-50, Ano10-51, Ano10-52, Ano10-53, Ano10-54, Ano10-55, Ano10-56, Ano10-57, Ano10-58, Ano10-59, Ano10-60, Ano10-61, Ano10-62, Ano10-63, Ano10-64, Ano10-65, Ano10-66, Ano10-67, Ano10-68, Ano10-69, Ano10-70, Ano10-71, Ano10-72, Ano11-43, Ano11-44, Ano11-45, Ano11-46, Ano11-47, Ano11-48, Ano11-49, Ano11-50, Ano11-51, Ano11-52, Ano11-53, Ano11-54, Ano11-55, Ano11-56, Ano11-57, Ano11-58, Ano11-59, Ano11-60, Ano11-61, Ano11-62, Ano11-63, Ano11-64, Ano11-65, Ano11-66, Ano12-36, Ano12-37, Ano12-38, Ano12-39, Ano12-40, Ano12-41, Ano12-42, Ano12-43, Ano12-44, Ano12-45, Ano12-46, Ano12-47, Ano12-48, Ano12-49, Ano12-50, Ano12-51, Ano12-52, Ano12-53, Ano12-54, Ano12-55, Ano12-56, Ano12-57, Ano12-58, Ano12-59, Ano13-33, Ano13-34, Ano13-35, Ano13-36, Ano13-37, Ano13-38, Ano13-39, Ano13-40, Ano13-41, Ano13-42, Ano13-43, Ano13-44]. OBC [Ano13-45, Ano13-46, Ano13-47, Ano13-48, Ano13-49, Ano13-50, Ano13-51, Ano13-52]. Object [YS12b, Lap11, LH14b]. Object-oriented [YS12b]. object-plasma [Lap11]. objective [SSS11]. objects [FRP14, PVDW14, Xu11]. observations [BLM11, GM11a]. observed [BM13b, BH13c, Har11]. obstacle [CHC11, KL11a, Le 14a, Lee14a]. obstacles [KKAS12]. Obtaining [CBGK13]. ocean [BDTW10, IBO14, Li12, LFLV14, OMCO14, SP14a, TQW14, ULS13, YK12]. ocean-atmosphere [IBO14]. oceanic [Hig13]. octree [BGZ12, PHRG13, PB13b]. octrees [Mur10]. ODE [BS14b, NV14, TPF11]. ODEs [RT14, ZK14]. Off [NKSR12, MMIW14, PB13b, YFL^{+}13]. off-body [PB13b]. Off-lattice [NKSR12]. Ohm [PSM11]. oil [OD12]. oligomerization [JJM13]. once [SW13]. One [EEK14, KS12, PRL10, YHT11, ACG14, Ain14, BH13b, BLTO11, CCB12, CW14a, CLS13, DRB14, FQV13, GDHM11, HH12a, Liu11a, LT11c, MM12, MPT13, MGN12a, NF12, Par12b, RB10, RS13b, SSR14, Stu10, TAI11, TA13, ZWL^{+}12a]. One-dimensional [YHT11, CCB12, CLS13, FQV13, LT11c, MPT13, RB10, SSR14, Stu10, TA13, ZWL^{+}12a]. one-domain [CW14a]. one-level [BH13b]. one-shot [DRB14]. one-sided [HH12a]. One-step [KS12, GDHM11, Par12b]. one-time [ACG14]. one-way [Ain14]. only [LS10a]. Ono [BX12]. opacities [Gen11, Gho17]. Open [DK14, BS10a, LL12a, MGLS12, MRS14b, MLM12, PGA11, SS10a]. open-loop [MRS14b]. OpenMP [ADT14]. operating [GMXG14]. operational [HHGC14]. Operator [Wal13, ZOL^{+}11, Zha14, BR11, BB12b, CC12, ETW10, FN14, FN10, HLDP14, KWH10, PG11, RXLS14, SSL13, WdHX10, WYS12b]. operator-splitting [WYS12b]. operators [BBC^{+}13, BG12, DBZ14, EAG14, FN14, FD11, Hua14, LR11, LB12, LL12b, MA13, Mat14, MAC14, Pir10, SW11, TAAS13, SRS10]. Optical [BFR13, CCC12, DR14]. optically [RGK12]. optics [CKR14, GHP13, GHH14, LBM14, Whi10, WJC13]. Optimal [BH13a, BGR11, BMK11, BMK12, BP13b, CSX14, CDL11, EC13, MAC14, NMCZ14, RSD10, SWR11, TD12b, XS11, AAG14, BVM14, BFR13, BD13, BAG12, DLLW14, EM12, FKM11, GLS10, HM13, IL11, Kaz10, LR11, LLJ10, MBS13, Oli14, PSDF13, SV13, VSK10, ZMD^{+}11, BFO14]. optimality [WFL10]. optimally [NRS10, WA12]. Optimised [GM11b]. Optimization [BRP14, GTSC14, LSVW10, MMM13, Ost10, vdVR12b, AJV11, ASK^{+}13, BHS14c, BPS14, BRSS11, BRS13, BKV12, CS14c, DLZ^{+}11, DLLW14, DR14, HA14, LD12, Man12, MNF^{+}10, Mét11, Mön13, SAM10, SSS11, SM10, Sti12, TNK10, TDVG13, WS10a, WdHX10, WKMJ13, YYY^{+}14, YDN10, YD12, fZlC11, ZLL10, ZWL10, ZPGO10]. Optimization-based [BRP14, GTSC14, LSVW10, BRSS11, BRS13, Mön13, YD12]. optimize [TK14]. Optimized [LLYE12b, ZY13, BSS14, CR14, GYL12, NDB12, TH14a, VG14]. optimizing [And13]. options [RA10a]. orbital [MIKG12, SP14b]. orbital-free [MIKG12, SP14b]. orbitals [Men12]. Order [SYV17, AB12, AINF13, ANIF14, ALR11, AQ11, ABL^{+}11, AM12, Ain14, AL10, AMP14, AF11, AN13, ACCM12, ABK12, ABK13, AT13b, ABJ^{+}12, BHB11, BMS12, Bal12a, BH12, BNT14, BWC11, BGH10, BvBZ^{+}10, BD13, BYCC14, BS10a, Boe11, BWMG14, BK14b, BBD11, BX12, Bre12b, BL10, BL13, BLR14, CIJS14, CYK^{+}14, CX13, Cap11, CD10, CCD11b, CS12b, CPCU13, CS14b, CH10, CSX14, CGMQ14, CDL11, CDHM11, CdCD12, Dai13, DEH11, DMT^{+}11, DMM^{+}13, DLGP13, DFNNRdlA14, DL13b, DCC^{+}13, DKR14, DAEB13, DRZ14, DCL11, DSZ13, DWZ10, DKA12, DV14, EO11, EAN11, Fan14, FPN^{+}13, FL13, FW12, FCYF11, FC13, FVR13, FP14, FMM^{+}10, Gao13, GNGAS10, GBNS14, Ge10, Gen13, GL14, Ger12, GMC13, GM12a, GH11b, GITW12, Gri12a, GCL13, GLWW14, GH12b, GCH14, GXS14, HH12a, HAMA14, HdCNT12, HBAP10]. order [HSD11, HRBK10, HRCW13, HE14, HWST12, HOK10, HTHG14, HM10, HLT10b, HAS13, HAH14, Hua14, Ide12, IX10, IG14, JNSA12, JZ11, JLZ13, KK10a, KKLR10, KLMJ12, KVM12, KKS13, KF13a, Kat14, Kaw13, KKvZB14, KSR^{+}14, KAFB11, KH11, KEGM10, KVBP^{+}14, LH10a, LTC10, LFL11, Le 14a, Lee14b, LPE^{+}11, LKLG14, LSBJ14, LGC13, Ler13, Ler14, LLZ11, Li10b, LXF10, LX12, LR12, LP14a, LT14, LGLX14, LY13b, LN10b, LM14a, LS13d, LWS14, LJG14, LM13, LLS11, LL14b, LQ11, Luo13, LB10, MDTC11, MDTC13, MF14, MB12, MAB^{+}13, MKGV12, MN11, MIJ10, MTT13, MZ11d, MZ14, MTB10, MT10, Mig14, MF10, MTG11, ML13, MTS^{+}11, MAD13, Mön13, MSB14, MF13, MIKG12, MNL^{+}13, MWW^{+}13, MPT13, NMCZ14, NPC11a, NPC11c, NDBG14, Nis10, Nis14a, Nis14b, NCKN10, NMT^{+}12, OS11a, Ols11]. order [PDSS10, PMM^{+}11, PGLT10, PVLT10, PGL11, PCF^{+}13, PL10, PRL10, Per13, PPWZ11, QLW14, QC10, QS11a, RV10, RZ10, RA10b, RS11, SRN10, SVK14, SGS13, SSHT14, SHCY13, SM11b, SLGB12, SZ10, SQC11, SZH12, SFCF10, SCY11, SR11, SYV14, SK12b, Sou14, SN13b, SDH12, SD14c, sSLxRyZ14, SY14b, SIDG13, TS11, TWSN12, TAI11, TK12, TK15, TDVG13, Tit12, TR11, TGGT13, Tyl14, UKP14, UJvL10, UJ12, VK12, VLM13, VWPF11, VWF13, VQRD11, VCJ11, VSK10, VQVD11, Wan11a, WMI10, WPP10, WYS^{+}11, WSYS12, WZSN12, WGX13, WLW13a, Whi10, WAM14, WSBG10, WP14, Wit10, WG14, WYT14, XQXC14, XDZ^{+}14, XW14, YL11, YXZ13, YHM12, YYB12, ZT13a, ZF12, ZS10a, ZS10b, ZS11, ZS12, ZZS13, Zha10b, ZTL13, ZD11b, ZCW11, ZF14, ZWH^{+}12, ZDSP11, vHtTBC11, vdVR12a, vdVR12b]. ordered [LCE^{+}10]. ordering [SC10]. ordinary [CX13, LCY11, RH13, TS14, WH11]. ordinates [CFW^{+}13, GD10, GL12a, GLC^{+}11, LTSH10, MM11]. ordinates-spherical [MM11]. organized [DH13]. organizing [Phi14, RSS12]. oriented [BDA12, ID10, LDA10, MB13b, MTD14, VB13, WR11, YS12b]. origin [BdCB11]. Original [HVD13]. Orthogonal [WTYC13, YZX14, AV13, CRG13, CHZ13b, CHZ13a, CSK14, FF12, Gri12b, Liv10, Pir12, ULS13, WABI11, ZJW12, Zie11, CSK13]. orthogonal-curvilinear [Zie11]. orthotropic [Ber14]. oscillating [HVT12]. oscillation [FDHP11]. oscillation-free [FDHP11]. oscillations [ÁCB12, KWHD12, Le 12, LKCY11, LY13a, SM11c, WHB14]. oscillator [BV13, MJK13]. oscillatory [CCD11b, CRT14, CLM13, DB13, FSTY14, FCYF11, HKLY13, HWA10, IG14, JN12, LQ10b, LWS14, MC12, MSD12, WLW13a, WW14, WJC13, WWS13b, YF14]. Oseen [Sha13]. Osher [LTCN13]. Osher-type [LTCN13]. Ostrovsky [MYM12]. other [YS11b]. outcome [GSTB13]. outer [Zho10]. outflow [Don14b, DKC14, GCA13, LFK11]. Output [FL11a, Hic12, KF13a, BN12, BZ12, BZKL13]. Output-based [FL11a, KF13a]. outputs [MM10]. overdamped [MDPTK13]. overfilled [Du11b]. overlapping [ABHS12, BAG12, GK10, JC13, OD12, RHB12, SFWP10]. overset [PB13b, RS14a, SSS10, TQW14, WDP^{+}14]. oversight [vL11]. oxide [BBCH^{+}12].

P [vL11, ARF12, NOT14]. P1 [Hua14]. Pacific [OMCO14]. package [KAG14]. packages [DZL13]. packet [RS13b, RS14b]. packets [Zho14b]. packing [LJ13a]. packs [SJ10]. paddle [OBT12]. Painlevé [FW11]. pair [CH11b, RKD^{+}14]. pairing [PMF13]. pairs [GL14]. panel [HK14]. panel-based [HK14]. parabolic [ARR11, BdF10, LD14, LM13, LHV13, LB10, MST14, MBA14, NS14b, Whi10, YWY10]. parabolized [UN13]. paradigm [BZKL13]. Parallel [BLTO11, CBTR13, FB11, Gen13, GL12a, GFK13, JMV^{+}12, KK10b, KMvM^{+}12, KH11, NS11b, RK14, RHGT10, SFWP10, TRSMM^{+}13, YC11, ARR11, ARR13, AAI10, ADT14, AJG^{+}13, AKP^{+}12, BC10a, BLRR11, BPOS12, CDSP13, CC12, CBGK13, DGM13, DZL13, FT11, Fuj11, GG10, GM12b, GMG12, GMXG14, HOK10, Her10, HPLS10, HH13, HWHW10, ITO^{+}10, LS10a, LC14, LFLV14, MMM13, MMM11, NIN^{+}14, OTV13, PZG14, PBdGP14, SLT^{+}14, SSB^{+}12, SYE10, TDU14, TPA^{+}13, Wan13a, GRT12]. Parallel-vector [NS11b]. parallelepipeds [KST13]. Parallelization [SNS10, CNSFD13, Kim13]. parallelized [CS13, RBNS^{+}12, RBNS13]. Parameter [AR14, BMK11, BMK12, ECDB14, EWH13, GHM10a, GHM10b, HBAP10, LA10, MPS^{+}14, ST13, WC14a, ZD10]. parameter-dependent [ZD10]. parameterization [CZL10, XMDG13, XMDG14]. parameterize [KM14b]. Parameterizing [ZZK^{+}14]. parameters [BND^{+}12, GMT10, IWG13, OBNN13, OBNN14, XMS12]. Parametric [BNP10, AKMD14, BGN10]. Parametrization [ZM14]. parametrized [XQX13]. Parareal [BLM^{+}14, BERB^{+}12, RBNS^{+}12, RBNS13, SNS10]. paraxial [AC11]. part [SM12a, SM12b, AP11, AP12b, Bal14, BHS14a, BHS14b, BD10a, DL13a, DL13b, DEKBF14, FNGV14, HR10, JWNL11, JPC^{+}14, JKR14, KWJR14, KJWR14, LB14, NL12b, PZW14, PZG14, PRA13, RPA13, RS13b, RS14b, SG10b, SG10a, SG12a, SG12b, SYY14a, SYY14b, YS11a, YS11b, ZMD^{+}11, vdVR12a, vdVR12b]. partial [AP12a, ARR11, CCLS10, CHZ13b, CHZ13a, DA14, FVR13, GM11a, GM14a, Gri10, JHF10, JTXZ12, LLZ11, LS13c, LM13, MZ10, MGN10b, NP12, Pir12, RH13, RHB12, WZ13, XH14, YXZ13]. Partially [Bea12, JKW11, Gio10, Har11]. participating [SRS14]. Particle [DGAH10, ILM^{+}13, KHRD11, Lap12, LFK11, Lip12, MDPTK13, PAG11, AHA12, AHA13, AP11, ADMA13, AJG^{+}13, AF13, AW14, BR10a, BLOdV14, BILM14, BBC^{+}13, BBGP14, BM10a, BKW10, BPHK13, Bre12b, CD13a, CCB13, CP14, CWL10, CCB11, CCB12, CC13, CCMG^{+}13, CCL^{+}14, CWZ^{+}11, CBGK13, DTG^{+}14, DRW11, DC13, DG12, DKW10, DF10b, ES13b, FCY13, Fuj11, Fuj13, GPY14, GG13, GM11b, GLC^{+}11, GHP13, GJ14b, GP13, HT14, Her10, HMM^{+}10, HLZ14, KKPV12, KF12, KGKS13, KIG11, KG11, KG13a, KWH10, KHJ^{+}13, KSB14, LBC14, LTC13, Lap11, LLZ11, LC13, LTL^{+}14, LXSR12, LT11b, LEHA10, LWF11, Mac10, MC12, ML11a, Mar10a, MS12, MB11, MPWK13, NS11b, OBNN13, OBNN14, PTM13, PBT14, PWK11, PMF13, PP14, Pri12, RPV13, RSS12, RSFS13, RBCK10, Ros10, SRS10, SBA13, SC11]. particle [SY12c, STD^{+}10, SGV13, Tau12, TE14, TAAY14, TP12, TVEC12, UPDB13, VB13, VdMS^{+}14, WPP10, WLL13, WZG^{+}13, WK13b, WAM14, Zha10a, ZF14, ZKKF10, vDS12, DDN^{+}10, DC13, Lip12, MHL^{+}14]. particle-fluid [AF13]. particle-grid [AP11]. Particle-in-cell [DGAH10, KHRD11, ADMA13, CCB13, CCB11, CCB12, CC13, CCMG^{+}13, CCL^{+}14, CWZ^{+}11, DTG^{+}14, Fuj11, GHP13, HT14, KGKS13, KHJ^{+}13, KSB14, ML11a, DDN^{+}10, DC13, MHL^{+}14]. particle-laden [Bre12b, CD13a, MB11, VdMS^{+}14, ZF14]. particle-LES [GM11b]. particle-mesh [DF10b, Lip12]. particle/finite [PP14]. particle/finite-volume [PP14]. particle/level [TVEC12]. particle/level-set [TVEC12]. particles [ARF12, AC11, Dai13, DLM^{+}13a, GHS12, ID10, LLdSC11, LTC10, LZC^{+}13, Liu11a, LR13, MD13, MGN14a, OTV13, PLT14, PSF^{+}14, TE14, vydCNS^{+}10]. particular [LGE^{+}13]. Particulate [PB14b, AF13, KM13, Zoh13]. partition [Gil13, OJW10, SAHS14, YL14a]. partitioned [AV13, BK12a, BHS14a, BHS14b, NPV14, RS14a]. parts [BN12, DBZ14, Hic12, MA13, Mat14, NL13, SN14]. passage [BBAP14, DBO^{+}10, MSS^{+}14]. passing [GBS^{+}14]. passive [GHM12, LBC14, VXB14]. past [SSS10, WPKK12, Wil13]. patch [BRVE12]. patches [GK10, RCW11]. path [DH13, QA11, WJC13, AAT^{+}14]. Patient [KVC^{+}13, IWG13]. Patient-specific [KVC^{+}13, IWG13]. pattern [AVKP11, KKPV12, NKSR12, SAM10]. pattern-forming [AVKP11]. patterns [HDPF11, JHZ10, LCWN12, PMMB14, PMSM12, VSC^{+}11]. PC [Liu11a, NV14]. PCG [GM12b]. PCKF [LLZ14]. PDE [BD10a, CF11c, HA14, IH13, SK12b, SPM^{+}13]. PDE-based [BD10a]. PDE-constrained [HA14]. PDEs [BFE12, BL11, CEGG14, CGM^{+}12, DO11b, DE14, FL11b, GKG^{+}13, GCW14, GKN^{+}11, GJ11, HDHZ13, JPC^{+}14, KL14a, Kha13, LJ13b, LHV13, LB10, Mar13, MNS13, SKW14, SS14c, Vai11, VK12]. pdf [vDS12, MJ13a, VTTK13, VWP11, WPP10]. PDF/FDF [WPP10]. peer [JKW11]. Pegasus [KSB14]. Penalization [ICC^{+}10, ABD14, BDKV14, CP14, KKAS12, MLBS14, PBC^{+}14]. penalty [AAG14, CN12, DL10a, EGL^{+}13, HT12, HLY11, HCS11, HCS12, KL10a, LRL10, VSC^{+}11, VdMS^{+}14, XH14]. penetration [BDMP13, KSU13, SKU13]. pentadiagonal [Kim13]. peptides [ALGB^{+}13]. percolating [KL14b]. Perfectly [KAA12, LN10a, MJ14, CVI^{+}13, Dur14, FL13, HJK14, KKS13, LH10a, SSM14, SF12]. perfectly-matched-layer [CVI^{+}13]. Performance [LS10a, Wei12, CS14c, DKH11, FT11, Fuj11, GM12b, GKG^{+}13, HMOH11, JEU11, KG13a, OAKR14, SSR^{+}13, VSK10, WD11, Wit13, ZRM13]. Peridynamic [OMA14, BD12, KFOS14, WT12a]. periodic [BG10, BD14c, CK10, Coa12, CH14, DB12, FJ12, GL13, GB13, GHL14, GLWW14, GD14, HNS12, HVD13, JTT14, JWYH10, KL14b, KG12b, LRL10, LBCL13, LT10b, MN14, Nic10, NN12, PS14b, WS14, ZVNM10, vRLPK11]. peripheral [GDY10]. peristaltic [WS10a]. permeabilities [dCRCS10]. permeability [GLL^{+}11]. permeable [GGH14]. perpendicular [HH13]. perspective [HZ14]. perspectives [MG14a]. perturbation [Ham10, KO14, LKTL10]. perturbations [MBS13, Wan13a]. perturbative [BX12]. perturbed [FJ12, Kha13]. Petrov [AWK^{+}11, FPN^{+}13, ZMD^{+}11]. PGD [DAA11]. Phase [GMXG14, JR14, LC13, LZ10, TK14, AHA10b, ABL^{+}11, ALGB^{+}13, Ala14, ABMT12, ABMT14, BHL^{+}13, BW13, CNR10, CSW14, CY14, CL11b, CKR14, DMHL14, DS12, Don14a, Don14b, DFW^{+}10, ES10, FGR12, GT12, GW12, GW14, GH11b, GRS14, GJLY13, GK14, GWG14, GMG12, HHA14, HK13, HLLW11, HQL^{+}10, HW14a, IK11, KCG14, KT14, LP13b, LQ10a, Li13, LVZK14, LHPH11, LYD10, MB13a, MF14, MFM12, MW14a, MF13, MJM13, MMJ14, PZTW12, PBWK11, PRA13, PBM11, PS14a, PS15, PM14, PZ10, RBB12, RDHK12, RB13, RPA13, dCRCS10, SN13a, SF11, SMP^{+}13, SGS13, SNBN13, SSC13, SPF10, SX14, SHA11, SHA12, TNK10, TSLV11, TT10, TL13, VP10, VP14, VBCM10, VSC^{+}11, WSS12, XYL12, XR14, YK12, ZM10, ZHW10, ZYHW13, ZMQ13, ZHY13, ZN14b, ZYF^{+}10, ZWJW10]. phase-averaged [YK12]. phase-change [DMHL14]. Phase-coexistence [LC13]. Phase-field [LZ10, CNR10, DS12, DFW^{+}10, GH11b, GRS14, GMG12, LVZK14, PRA13, RPA13, ZYF^{+}10]. phase-less [IK11]. phase-space [CKR14]. phases [LCE^{+}10]. phenomena [BP13b, HZ11b, IBO14, JXD14, LL14a, RBK13]. phonon [NSA11]. phonons [MM11]. photoelectron [WLPG11]. photoionization [GH12a]. photon [RKD^{+}14, WLPG11]. photon-photoelectron-counting [WLPG11]. photonic [BSS11, HCD11, HCH^{+}10, MNF^{+}10]. Phys [ASS17, CFCA13a, Dav15, Gho17, HY11, LWD13a, NHT^{+}13, OBNN14, PS15, SYV17, SS14b, TK15, XMDG14, YKWS13a, ZJS15]. Physalis [Liu11a, GP13]. physical [BBC^{+}12, FQV13, Hen13, WJC13]. Physically [DR11, BNM14, BGBR12]. physically-based [BGBR12]. physicochemical [CHKT13]. Physics [ASXZ14, Dad12, EH15, HSK^{+}15, KAB^{+}14, PRH^{+}14, XS15, ASGW11, ASK^{+}13, CNST13, Gra10, HMM14, KEG^{+}12, MMIW14, NB14, Ton14, TBNT11, Try12]. Physics-compatible [KAB^{+}14, PRH^{+}14]. PIC [AW14, CN10, GV13, GVH14, GV14, Lap11, TP14, YXD^{+}14]. PIC/fluid [TP14]. piecewise [Gon10, HLY12, KJ11, fZlC11, ZWL10]. piezoelectric [CVI^{+}13]. pinch [YFL^{+}13]. pinch-off [YFL^{+}13]. PIROCK [AV13]. piston [ZYLK13]. Pitaevskii [LX14, Sal14, TA12]. pitching [JM12]. PKD [MOSW12]. planar [Bre12a, CSKP10, DFD14, LSZ13, Tit12, LEH10]. Planck [JY11, CCT11, GJ14b, LE13, LW14, LHMZ10, Mar10b, MG14b, PLT14, SJH12, TTR^{+}12, TBNT11, Wol12, ZCW11]. plane [BHK^{+}10, DB12, DKW10, JFC^{+}13, KED^{+}12, RKGM14, STD14, VT13]. plasma [AiINT14, AAG14, BBF12, CWZ^{+}11, CKD10, CBCM^{+}11, DCC^{+}13, DST14, DGAH10, ES13b, GKS^{+}11, Gür14, HH12b, HH13, ILM^{+}13, ICC^{+}10, KM11, KEG^{+}12, KSB14, LA10, Lap11, Lip12, MHL^{+}14, MMIW14, MMA13, PBC^{+}14, PLT14, RBNS^{+}12, SNS10, SKt10, TGT^{+}10, TTR^{+}12, TD12a, Unf13, ZZK^{+}14, ZKV10, vydCNS^{+}10]. plasma/flow [Unf13]. plasmadynamics [PMS11, PSM11, PSM13, PMS14]. plasmas [AHS14, BP13a, CDS13, CM10, DMR14, GP12a, HdBK12, HT10a, Jar12, KST13, KA12, KHJ^{+}13, Tok11a, TSR13, VHG13, VCS14]. plasmonic [HZS12]. plastic [DKR14, FG12, HPOM10, LTS^{+}10, LKT^{+}12, MAB^{+}13, OLPM14, SSB13]. plasticity [BJLR14, RG12]. plate [LHM14]. platelet [WMK13]. plates [FGE11]. platforms [CDSP13]. PLIC [MZ11c]. plus [KK10a, RV10, Wal13]. PML [OJ11]. POD [RV10, SN13b, TR12]. POD/DEIM [SN13b]. Point [MBR11, XLD^{+}11, Alb10, AEM13, BGGGS11, BCM12, CR13, DLC11, Her10, KED^{+}12, KNR11, LQ11, MGJZ10, MDTC13, MMHA^{+}12, ML14b, RA10a, SBN12, Soa13, UPDB13, WDGY10, ZMG11, ZD12]. point-based [DLC11]. point-source [LQ11]. point-to-plane [KED^{+}12]. point-value [XIIX14]. Point-wise [XLD^{+}11]. points [CDLL12, DS11, Li13]. pointwise [KG13c]. Poiseuille [MWRZ13]. Poisson [Don11, ZD11a, ARR13, AAI10, BH13b, BLTO11, BY11, BHM14, CRS14, CGR11, CCG11, CLM13, DO11a, DDN^{+}10, DMR14, Ge10, GCZ13, GK13, GM12b, GCH14, GT11, HGMM12, HRCW13, HG11, HT12, JWV12, JM10, KTT13, Liu13, LW14, LHMZ10, MRS14a, MS10b, MS11, MNR11, MJK13, MP13, MTG11, MG14b, MY13, PRH^{+}14, PGR10, PSF^{+}14, QS11b, RS11, VCS14, Xie14a, XM14, ZCW11]. Poisson-BGK [DMR14]. polar [QZM^{+}12, SCH13b]. pollution [JWNL11]. polyatomic [CL14a, Son11]. polycrystalline [DFW^{+}10]. polycrystals [Ber14]. polydisperse [BDMV11, FNKdLB13, KLMJ12, VLM13]. polygonal [BR10b, LS10b, LSV10, NS11a, SY11a, SY12b, TR14, WGD12, WJC13]. polyhedra [SdAW14]. polyhedral [LMBB11, LM14a]. polymer [KCG14, KL14b, LHK12]. polymeric [ASK^{+}13]. polymers [SJB11]. polymorphs [Ber14]. Polynomial [HG10, BS11b, BWMG14, Boy13, GvdSVK10, HWHW10, KL14a, LBR14, PHD14, PGLK14, RWX11, RLPM12, SF14, WTYC13, CSK13, JWV12]. polynomials [BY11, CJM13, DV13a, Gon10, KBRM12, LB14, Liv10, MOSW12, Ols12, Rem11]. population [BKW10, MAK14, PWK11, VWF13]. pore [HW14c, MB10a, OP10, RBB12]. pore-level [OP10]. pore-scale [MB10a]. Pores [SW10]. poro [SCH13b]. poro-elastic [SCH13b]. poroelastic [BCL13]. porous [AST12, BCL14, CTP13, CZ13, CSW14, CW14a, CL11a, DM13, DFVY14, DL10b, GPCE13, GCE14, GCFJ13, HSK^{+}15, HZ11b, KTT12, KFOS14, LW13b, MZ11b, MF13, MJM13, MMJ14, OP10, PM14, RBB12, dCRCS10, RM14, SKN13a, SGS13, SGC13, SHCY13, SKN13b, TLG10, TL13, WT13, WHT14, WLW14, YL10, Zad11, ZM14, ZHY13]. posed [JZ10]. posedness [RNT12]. positive [EZ10, GM12a, RMSF11]. Positivity [CCGP13, CLQX13, CS14a, HAS13, Par12a, Par13, QS11b, ZS11, ZS12, Bal12a, BKC^{+}11, GXS14, MKCB14, NV14, RS11, ZS10b, ZYS14]. Positivity-preserving [CCGP13, CLQX13, CS14a, HAS13, Par12a, Par13, ZS11, ZS12, BKC^{+}11, GXS14, RS11, ZS10b, ZYS14]. positrons [SBS^{+}13]. post [AM12, AM10, Ide12, ZL13]. post-processing [AM12, AM10, Ide12]. post-smoothing [ZL13]. posterior [Lee14b]. posteriori [DFVY14, DZLD14, ETW10, HKL10]. posteriors [AGS10]. postprocessing [BKV12]. Potential [GCBH10, GCBH13, BK12b, CLN14, CLW10, ES14, Gon10, HCP10, KJ11, LLL10, LTC14, PSM11, PA14, RTKS10, Wei12, ZZK^{+}14]. potentials [AJRT11, Gon12, JWYH10, KIF^{+}14, KBGO13, LT10b, MTT13, STD14, VBSK10, WJTY10]. Power [CM13, GKG^{+}13]. practical [CCGP13, Yok13]. Prager [LGG^{+}13]. pre [ZL13]. pre-displacement [ZL13]. precipitates [DCVM11]. precise [LCK11]. precision [CDSP13, CCB12, CBGK13, FMT11b, IKO^{+}14]. Preconditioned [Fat10, HKM12, LYJ14, AD14, CCW12, HWHW10, MS10a, MP13, WD13b, XJ14]. preconditioner [BNNW11, DDL13, Du11a, LR11, LS13b, MMH12, SEDW13, SS14c, TEY12, VG14, Vol14]. preconditioners [ABJ^{+}12, BMP14, EL11, MY13, RF14b]. Preconditioning [BGHM13, BSS12, HCH^{+}10, NN12, Pie12, BMR14, CBPS13, CCL^{+}14, CST12, FAY^{+}13, Lee12a, RVbZ10, Shu14]. predict [GSTB13]. Predicting [RKM11, CLS10, RJC13, ZCD10]. prediction [BCW13, CS12c, DMP14, GHM10a, KVM12]. Predictive [ECD14]. predictor [CP13b]. predictor-corrector [CP13b]. predictors [GDHM11]. Preface [AQ11, LPK13]. Preliminary [Ste14]. premixed [GG10, Par10, TBMH11]. prescribed [MP13]. presence [Dai13, SSX13]. preservation [CS10, FF13, PP14, TFP13, ZFJ11]. Preserving [CGM^{+}12, CDK12, Bal12a, BSS14, BRP14, BKC^{+}11, BKV12, CWL13, CW13, CCGP13, CdCNH14, CLQX13, CS14a, CS14b, DLNN12, DMR14, FJ10, Fus13, GW13b, GCW14, GJLY13, GXS14, HPA14a, HAS13, HLLW11, Hua14, JY11, Kaw13, KBS14, KAA12, LP13b, MKCB14, MMR12, MN12, Mie13, OHF12, Par12a, Par13, QS11b, RSD10, RLJ14, RS11, SY11a, SML13, TLO^{+}14, VW14a, VKL^{+}13, VBGM13, WZZZ14, WDGY10, WGD12, WWS13b, XQX13, XQXC14, ZS10b, ZS11, ZS12, ZYS14, CLM13, DDN^{+}10]. Pressure [Hig13, HF11, MKB^{+}12, SS10c, BD11b, CIJS14, CP10, CFKL11, DF14, FE11, KED^{+}12, KFOS14, KS13a, LLP10, MTSG12, Min13, Par10, SM11c, SC11, SSL13, SR13d, TM10, Vre14, XPO14, ZGSZ10]. pressure-based [CP10, XPO14]. pressure-correction [BD11b, DF14, SR13d]. Pressure-velocity [MKB^{+}12]. pressureless [YWS13]. pressures [TK12, TK15]. Preventing [JH12]. prevention [NDV^{+}11, SM12b]. primal [XJ14]. primary [KPH13]. Principal [SBT14, Hua14, MZ11a]. principle [Kol11, SY11a, XQX13, XQXC14, ZS10a, ZZS13]. principles [NMV13, TD11a]. priori [FVR13, LLS11]. priors [LMM14]. prisms [BD13]. Probabilistic [Gri10, MPPP12b, MPPP12a, FK10, LTT10, LM13, MPS^{+}14, WZ13, WZ14]. Probability [BW14, LX10, LLX11b, VK12]. problem [Alm14, AS12, AH10a, BMP14, Bal14, BGN10, BB13, BFO14, Boo11, BLR14, BCW13, CFW^{+}13, CLN14, CB12, CRS14, CXZ10, CH10, CV10, CH11a, GW12, GW14, HLS14, IL11, JRTY11, JR12, KL11a, Lee14a, LLJ10, LW13a, LN10b, Mét11, Mön13, MCDT12, PBC11b, RSK13, Ren11, SV13, Sti12, hT10b, TPA^{+}13, VG14, Vol14, WPK14, WG11, WLPG11, Xie14b, XL11, YLW10, YD11, YWY10, ZYLK13, RB10]. Problems [YKWS13a, YKWS13b, AB12, AV13, ALR11, AQ11, AC13, AS14, AT13a, APY14, AH10b, AT13b, ABH^{+}14, BF14, BHS12, BHS14a, BHS14b, BSSW12, BD10b, BOD11, BDMP13, BvBZ^{+}10, BN12, BRS11, BdF10, BDM13, BOK12, BDLGC13, BDKV14, BLY10, BL13, CBPS13, CCFCM11, CNR11, CYK^{+}14, CTG12, CCK13, CHHL11, CEL14, CR13, CdCD12, DMHL14, DDF14, Deg11, DLC11, EGW11, EMS11, EWH14, Era13, EO11, ES13a, FGE11, FGR12, FL14, FJ10, GYL12, GW13b, GLSC12, GC11, GJ14a, GLS10, GCLS11, GB13, GBS^{+}14, GHH14, GL12b, GLL^{+}11, Guy13, HP13, HAMA14, Han11, HHS^{+}13, HPS13, HNMM10, HBM13, HWST12, Hen13, HLW13, HSWZ13, HY09, HY10, HY11, HCH^{+}10, HAQ12, Hua14, HF10, HWHW10, Ide12, JAX11, JZ10, Jin12, KK10a, KSU13, KS13b, KO14, Le 14b, LKCY11, LY13a, LH10c, LT12]. problems [LLW14, LSZ13, LMS11, LMBB11, LRK11, LWS14, LYD10, LN14, LHM14, MWP14, MKCB14, MBR11, MDTC11, MDTC13, MF14, MM10, MK14a, MNR11, MHHGN14, MWW^{+}13, MMH12, MV13, NTV14, NL10, NPV14, NW10, NB14, NKS10, NS14b, OS11a, PGR10, PHRG13, PR11, PvDGvM11, PRL10, QLW14, RCQL11, RA10b, Ric13, RF14b, SKU13, SY12a, SY14a, SSHT14, SR13a, SK13, SY11b, SSG11, SKCC10, SCC14, Shy10, SKN13b, SK12b, SAA14, SIT^{+}11, SDH12, SD14c, SN14, TH14a, TD12b, TA13, UJ11, VBVD10, VRB10, VS14, WYYK10, WLW13a, WY11, WCVJ13, WL11, WGD12, ZBS12a, ZBS12b, ZK13, ZF12, ZSW10, Zer10, fZlC11, ZOJR11, Zho10, ZQ11]. procedure [And10, EH14, EH15, FGN14, GW13a, GP13, Her10, KKC^{+}13, LCP13, LMZ14, SS10a, TS10, XS15, YWL14]. procedures [NEE12, SBB10]. Procesi [YS13]. process [BZ12, BZKL13, BHBM10, JLS^{+}14, KKL10, PYK10, SS13b, SS14b]. processes [BHM11, BHM13, BCK11, BBAP14, CHKT13, FHL10, HL11, KL12a, LP12b, LHMZ10, MSS^{+}14, RJC13, SJB11, ZM14, ZKKF10]. Processing [HJBB14, SSK^{+}12, AM12, AM10, DMP14, GHH14, Ide12, KHRD11, LSK11, LLL10]. processor [BGZ12]. processors [CBGK13, KSPP10, LBRA^{+}13, Wit13]. product [BGK12, DSZ13, GHDS10, OJW10]. production [RKD^{+}14]. products [AK10, BBM14]. profile [BBF12, LBK10]. profiles [CGR11, Ler14]. Programming [BVM14, MNF^{+}10]. Progress [LRT13, dv23, ISZ12]. project [Rod12]. projected [vVVK10]. projection [BB13, BFSVC14, DSHP11, GH10, HLY12, KG13b, PB14b, RV10, Ren11, SM11a, SS13a, SSL13, Vre14, Wan14]. projections [KST13, KMSS10]. prolate [Men12, WZZ14]. prolate-collocation [WZZ14]. proof [GF13]. propagating [LXSR12]. Propagation [Li12, AJRT11, Ald10, AF11, ABK12, ABK13, CZ13, CL11a, CH11b, DRZ14, DKM11, FH11, FJ12, FMM^{+}10, GJ14a, Gri13, HBZL14, JPC^{+}14, JKR14, KVM12, KEGM10, LH14a, LT11c, LBR14, RT14, RF14a, SCH13b, Soa13, SLC^{+}11, TK14, TH14a, TAT14, Tok11b, TD12b, WSBG10, ZMD^{+}11]. propellant [DLM^{+}13a]. propensity [FWLP14]. Proper [DAA11, Gri12b, WABI11]. Properties [LKM13, AR14, AT13b, BMCK10, BCM^{+}11, BP13b, CS10, CB10, DBHBB13, FD11, GCL13, HTHG14, LTC13, LKP14, LTL^{+}14, LR13, MPT13, OL13, SSHT14]. property [ANIF14, Mie13, MGN14b]. proposals [OJ11]. propulsion [BBGP14]. prosthesis [LS13a]. protein [EHA14, FCH13, SBA13]. prototype [HvB11]. Provably [GH11b]. proximal [AKL^{+}13]. PSE [CZ11b, SRS10]. Pseudo [MLM12, PAG11, GG14, GVH14, LZC^{+}13, Mac10, MLBS14, MMH12, PE10, SCH13b, VHG13, Zha14, vRLPK11]. Pseudo-random [PAG11]. pseudo-solid [LZC^{+}13, MMH12]. Pseudo-spectral [MLM12, GG14, GVH14, MLBS14, PE10, SCH13b, VHG13, vRLPK11]. pseudo-subcells [Mac10]. pseudo-time [Zha14]. pseudoanalytical [MTD14]. pseudolikelihood [EHA14]. Pseudospectral [HG10, xLqW13, Brü13, Don11, DBHBB13, GKS^{+}11, Hen13, HT12, HDD12, PPS12, Sal14, SW11, Yin14]. public [BHB11]. Publisher [Ano13-53]. PUFEM [EL11]. pulled [LEH10]. pumping [WS10a]. PuReMD [KAG14]. pursuit [SCIE14]. PVM [DFNNRdlA14]. pyramid [HHA14]. pyramids [BD13].

QCD [APGL12]. QSM [ZSW10]. Quadratic [BVM14, FKM11]. Quadrature [IAM^{+}12, KBGO13, BK14a, DKR12, SAHS14, VWPF11, VWF13, VHWF13, YF11, ZKKF10]. Quadrature-based [IAM^{+}12, VWPF11, VWF13]. quadratures [BRS10, NCKN10]. quadrilateral [LX13b, WAM14, WDGY10]. quadtree [LBT11, PHRG13]. quadtree/octree [PHRG13]. quality [GP12b, KS11b, LLS11]. Quantification [BR14, AC13, AA13, BZ12, BZKL13, CPX13, CNST13, CZ13, JEX10, JNX13, LTT10, MPPP12b, MPPP12a, SP13, SM12a, WT12b, WI13a]. quantile [KG13b]. quantitative [ZTG^{+}13]. quantum [AM13, AF11, ALS^{+}14, CBKK12, GHH14, HWHW10, JC10, JLC14, JHDC10, MJ14, SD14a, SO13, Wit13, XX12]. quantum-scale [CBKK12]. quark [AiINT14]. quark-gluon [AiINT14]. Quartic [ZSW10]. Quasi [FVR13, GKN^{+}11, Kim13, ADMA13, BG10, BAG12, BD14c, CLSX14, CRS12, EZ10, GB13, GLL14, NJ12, PT12, TAI11, TM10]. Quasi-a [FVR13]. Quasi-Compressibility [TM10]. Quasi-disjoint [Kim13]. quasi-incompressible [GLL14]. quasi-lacunae [PT12]. Quasi-Monte [GKN^{+}11]. quasi-neutrality [CRS12]. quasi-one-dimensional [TAI11]. quasi-optimal [BAG12]. quasi-periodic [BG10, BD14c, GB13]. quasi-positive [EZ10]. quasi-static [ADMA13]. quasi-uniform [CLSX14]. quasi-wavelets [YXZ13]. quasicrystals [JZ14]. quasidiffusion [WAM14]. quasigeostrophic [GM14b]. quasineutral [DDS12]. quasineutrality [DDN^{+}10]. quasistatic [SLC^{+}11]. QuickPIC [ADMA13].

R13 [RTS13]. racks [ARF12]. Radau [TDL10]. Radial [LSK11, ASS13, ASS17, BFE12, Boy10, BY11, Boy11, FL10, LB14, Pir12, PH13, RA10a, ST13, SPM^{+}13, MPR12]. Radiation [VHWF13, BDL11, Den11b, EMSH14, FAY^{+}13, GM10a, GD10, KKLR10, LKM13, LBS10, MSTL13, NB13, Ols11, Ols12, PWB^{+}14, RARO13, RGK12, SS10b, SAHS14, SVS11, She13, WR11]. Radiative [WP14, Boo11, CGG12, DTU12, EMK11, GL12a, GS11a, HMOH11, KBGRM10, KS13b, LTSH10, MH10, MSTL14, MTS^{+}11, OHF12, RCCC14, SRS14, WL11, ZYT13, ZTL13]. radiative-transfer [DTU12]. Random [DG12, GL13, PDC13, STD^{+}10, ARR11, ARR13, Ber14, CW10, CZ13, GKN^{+}11, HPS13, Hel11a, Hel11b, HW14c, KSK13, LJ13a, LTT10, MTAC12, MJM13, MMJ14, PBB14, PSDF13, PAG11, Ran12a, RWX11, SJ10, Wan13a, ZT10b]. Randomized [KSK13]. range [AF13, JSX10, PA12]. rank [DST14, HW11, Ran11, Ran12a, Ran12b, Val14]. Rankine [Guy13, ZMD14]. RANS [ECD14, GHS13, HHL11, WMYYG10, XJ12]. RANS-LES [GHS13]. rapid [VBSK10]. Rapidly [BD14c, JM12, STD14, Sto13]. rare [LLX11b]. rarefied [BCHM14, BM14, CVIA10, GJ14b, LXM14, MZ11d, RTS13, Tit12, TA13, XH10]. rate [HAD10, JW11, JXD14, MMSI13, MPPP12b, MPPP12a]. rates [SR13a]. ratio [BLR13b, LP13b]. rational [BX12, CHHL11, HLW13]. rational-function [HLW13]. ratios [AF13, CCKCG10, DS12, Hel11a, HK11, KG13a, LRT13, MOV12, Par10, dv23]. ray [RKD^{+}14]. Rayleigh [And10, CLAL14]. RBC [Ler13, Ler14]. RBF [BMCK10, BMK11, BMK12, BFE12, Boy11, DO11a, FLB^{+}12, FL11b, HCD11, KBGRM10, MDTC11]. RBF-FD [BFE12]. RBF-generated [FLB^{+}12, FL11b]. RBFs [Boy11, MDTC13]. re [CRS14, Den11b, FB11, HAH14, MD14]. re-averaged [HAH14]. re-emission [Den11b]. re-entry [CRS14]. re-initialization [FB11, MD14]. reacting [HK13, KSR^{+}14, LI14, MFM12, PZW14, PZG14, SRN10, VWP11, YYB12, vVVK10]. Reaction [Sch13a, AV13, Atz10, BCK11, BBAP14, CYK^{+}14, DL13c, DBO^{+}10, FHL10, FF12, FWLP14, Gui12, HL11, HLDP14, ISZ12, JZ13, LCWN12, LHMZ10, MSS^{+}14, MPPP12b, MPPP12a, MT14a, Nag14a, Nag14b, PBB14, SR13a, TCW^{+}10, ZLCW14, ZOL^{+}11]. Reaction-diffusion [Sch13a, Atz10, BCK11, BBAP14, CYK^{+}14, FF12, FWLP14, HL11, HLDP14, LCWN12, SR13a]. reaction-diffusion-advection [ZOL^{+}11]. reaction-drift-diffusion [MSS^{+}14]. reaction-subdiffusion [DL13c]. reactions [BSDM12, KLB^{+}14, KF13b, PDC13, Pud11]. reactive [BR11, Cam13, CGC^{+}12, CGF^{+}13, KIF^{+}14, KAG14, MKCB14, NV14, NMV13, NS14a, PP14, RXLS14, TLG10, VTTK13]. reactive-flow [NS14a]. reactor [MAK14, SSFM12]. real [BSCML13, BBF12, BB12b, CGF^{+}13, HPA14b, ITO^{+}10, NIN^{+}14, RPC14, UN13, VSC^{+}11]. real-space [HPA14b, ITO^{+}10, NIN^{+}14]. real-time [BSCML13, NIN^{+}14]. ReALE [LMS^{+}10]. realizability [OHF12]. realizability-preserving [OHF12]. Realizable [VWPF11, VWF13]. really [Den11a]. recognition [NKSR12]. reconnection [LMS^{+}10]. reconnection-based [LMS^{+}10]. reconstructed [LLN^{+}10, LXS^{+}13]. reconstructing [CF11c, KSS14]. Reconstruction [BBF12, AM10, BFR13, BNM14, BP13b, CLN14, CY12, Cha14, CTJT13, DFD14, DS11, GMB10, Gao13, GW13a, HD13, HS13b, HG10, Hu13, HQL^{+}10, HAH14, IST^{+}12, JSX10, JLZ13, KGSS10, LCP13, LMZ14, LXL^{+}12, LXS^{+}13, Mig14, NS11a, SS10a, SLH^{+}11, VCJ11, WCVJ13, XLD^{+}11, YF14, YWL14, SYJ11]. reconstruction-based [LXL^{+}12]. reconstructions [PB14a]. Recovering [CGR11, BKV12]. Recovery [HLY12]. rectangular [CS13, FS10a, NDV^{+}11, ZS10b]. Recursive [TVE14, BMP14, Ren11]. Redistribution [Chi12, BUH11, BBPC14]. reduce [BDTW10]. Reduced [AB12, BBD11, HBAP10, HW11, RV10, RA10a, SBT13, ACCM12, BH13c, CIJS14, CD10, CKR14, FPN^{+}13, FHMS11, GHS12, GL14, Gri12a, HSD11, KJC13, LP12a, SM11c, SLGB12, Sza12, TS12]. Reduced-order [BBD11, HBAP10, CD10, HSD11, SLGB12]. Reduced-rank [HW11]. reducing [AP11, LH10d]. Reduction [BF14, BdCB11, CFCA13a, CFCA13b, DRZ14, EGG12, MLM12, SN13b, TPF11, XFB^{+}14, YP13, ZD10, vydCNS^{+}10]. reference [BPQO13, BN14b, LGE^{+}13, LBS10]. refined [BCHM14, SKU13, UJ11, ZJL^{+}14]. refinement [BLOdV14, BCK11, CGG12, CY14, CN10, DZHB13, EH14, EH15, ETW10, FL10, Fuj11, GWCA14, HHL11, HB13, JLY10, KWSJ11, KG14, LMLC12, LGP10, LH10b, LN10c, LBM14, MSTL14, OS12, PWB^{+}14, Phi14, SQC11, SEBG12, VRB10, Wan10, WR11, XS15, DMMGM14]. reflection [LL14a, MZ14, WK14]. reflective [MK12]. reformulated [Dri10]. reformulation [MMT14]. Reformulations [MT14a]. refractive [SVS11, She13]. regeneration [YCY11]. regenerative [SR13b]. regime [LT11a, LQB14, Rie10]. regimes [AP11, LYXZ14, OMCO14, RGK12]. region [WT13, WTYC13]. regional [YCH^{+}13]. regions [AHS14, CY14, NDB12]. regression [BZ12, BS11b, HMM14]. regular [CF11c, HW10]. Regularization [ISZ12, BFR13, BC14, LBCL13, LDN^{+}10]. regularizations [CCLS10]. Regularized [OBNN13, OBNN14, BM10b, BM11, CGJ11, CCLV10, CH14, OLC13]. Regularizing [Alb10]. regulated [KF13b]. Regulator [JS13a]. reinitialization [HMS10, RB14]. reinitializing [Min10a]. related [Ber14, FJ10, KYP^{+}14, LWD13a, LWD13b, RSS10]. Relation [SC12, BSS14, GDGP11, LS13d, RSD10, Rem11]. Relations [PBM11]. Relationship [Oli14]. relative [MK14a, dCRCS10]. relativistic [AK14, AiINT14, ALS^{+}14, BD11a, DGAH10, GV13, GHH14, KMvM^{+}12, KB14, Mar10b, QY12, Ros10, SBS^{+}13, SS10d, TiI11, TP14, WT14, YHT11, YT12, ZT13b]. relativity [CH10]. relaxation [GLSC12, Gra14, HW14b, PBM11, PC13, XMS12, YM14, YN10, ZL13, Chi12]. Relaxed [BNNW11]. Relevance [SR13b, DF10b, WCM^{+}14]. Reliable [SLGB12]. remap [BCD14, BRSS11, BRS13, BRP14, KS14, LEM11, OS11b, OS14, Wal13, ZS14]. remap’-based [LEM11]. remapping [BBG^{+}11, KS12, LSVW10, VKL^{+}13]. remedy [OAXL13]. remeshed [RCW11]. remeshing [DDF14, MC12]. remodeling [SR13b]. remodelling [GDY10]. Removing [DE14]. renormalized [AH13a]. repair [MPR12]. Reparameterization [BJ12]. repeating [PMSM12]. Replay [NT14]. Reply [EH15]. representation [BG10, BV13, CCT11, CLG13, HT10a, JSA14, JHF10, KBRM12, KNR11, LLY11, MZ10, PL12, Val14, WZN14, XFC^{+}14]. representations [BK12b, RLPM12]. Reprint [EN14b, Nag14b]. reproducing [AHA10b]. rescaling [FR13]. reserving [JXD14]. reservoir [SGV13]. Residual [WP14, ALR11, BA10, DCL11, GCL13, HLSO13, HLT10b, LGC13, RA10b, RF14a, SHR13, VQRD11, VQVD11, XFB^{+}14]. residual-based [BA10, DCL11, GCL13, HLSO13, LGC13]. resistance [BB12a]. resistive [ASB10, BCP13, SPB^{+}10, SR13c, XPO14]. resistivity [Del13, Kaw13]. resolution [BLRR11, BdCB11, BLR13a, DM13, DjY14, DAA11, HHA14, JNM12, JLB^{+}10, KWJR14, KJWR14, KBSV14, LL14a, LZZS13, Par12a, RSS12, Shy10, TRL14, WSYS12, WD11, WI13b, XPO14, vL11]. resolve [Den11a]. resolved [BBGP14, Bre12b, HAD10, JWNL11, MdVBS13, VdMS^{+}14]. resolving [LR13, PDZ^{+}14, SYY14a, SYY14b, SW10, WLPG11]. resonance [NLGL14]. resonant [JCT11, LT11a, MJK13]. resonator [GP12b]. resonators [LM14b]. resp [CGM^{+}12]. respiratory [KKC^{+}13]. response [Abr10, SSB13, TR12, VK12]. response-excitation [VK12]. responses [LJ13b, LR13, RKGM14, Wan10]. restart [Zho10]. results [APRS14, CBGK13]. resummation [RH13]. retarded [AJRT11]. retrodictive [VDD10b]. reversal [BCM12, VMS11]. reverse [HVT12]. reversible [PDC13]. Review [Jar12, SN14, TTR^{+}12]. Revisit [XIC11]. revisited [CG12, FdVV12, OA10, PSDF13]. Reweighted [YK13]. reweighting [KSS14]. Reynolds [HRK^{+}10, MCA^{+}13, MGN14a, MY14, NGV14, NL12b, SH14, SCB10, WW13, vRLPK11]. Reynolds-averaged [SH14]. Reynolds-stress [NGV14]. rheology [CM14]. Rhie [ZZB14]. Richards [ML13]. Richtmyer [MJ13b]. ridge [BY11]. Riemann [AiINT14, Bal10, Bal12b, BDA14, Bal14, BBD14, Cap11, FGR12, Guy13, KB14, KMHJ10, LLX11a, MF14, MCDT12, MGN11, MGN12a, MLGN12, MGN12b, NGV14, OLPM14, PBM11, PR14a, QLW14, Rie11, RB10, SYY14a, SYY14b, SGL14, TWMM14, TT10, UN13, WKH10, WSW13, XL11]. Riemannian [LRS10]. Riesz [BTY14, ÇD12b, WXY13]. rigid [BHS13, BBGP13, De 10, MDM^{+}12, Xu11]. rigidity [FLM12]. rising [HHL14]. RK [XQX13]. RK-WENO [XQX13]. Robert [BY11, LT14]. Robin [BN11, Deg11, HG11, PBT14, PGR10, PHRG13]. Robust [AD14, CDF11, MH10, RS14a, TGRL13, WZSN12, AHS14, CF11a, CLX14, DKC14, ELM14, FMT11b, GMT10, HLT10b, KPH13, LCWN12, MT13, MBL^{+}14, MF10, Nor14, Par10, RGK12, RTS13, STD14, SYY14a, SYY14b, SdAW14, WFK11, WMYYG10, WI13a]. rod [OLC13]. Roe [LL14a, PBM11, PIN14, Rie10, Rie11, RB13]. role [JJM13, OMCO14, WKH10]. roles [PR11]. roots [Boy13]. Rotary [SPS14]. rotating [AD14, CCC12, CRG13, CT14, DH10, FW12, GNGAS10, GBNS14, HdBK12, JTT14, MTZ14, NvdVB13, PE10, PPS12, RBS12, SP14a, ST13]. rotation [LdWF12]. Rotational [FL10, WC11, DS10, SS14a]. rotationally [YHM12]. rotations [ES13a]. rotators [FGE11]. Rotne [LGG^{+}13]. rotor [XW14]. rough [HPV13, HLS14, LWCR10]. roughness [DWZ10]. rule [Ide12, GW13a]. Runge [MBA14, AV13, BMD^{+}13, BGR11, BR11, FP12, Ian11, KKvZB14, Ket10, KHHK12, Lan14, LP14a, LLH10, NYM13, NMCZ14, NDB12, RT14, RHXQ14, RA10b, SK12a, San13, SLT^{+}14, SK12b, Tok11b, TD11b, TD12b, WS10b, WLW13b, XCL14, ZT13b, ZS13, ZZSQ13, vdVR12b]. running [CDSP13]. runs [BGZ12]. runup [DKM11].

S [DL10b, LY13b, LD14]. S-DDM [LD14]. S-FDTD [LY13b]. SA [BMR14]. saddle [BGGGS11, ZD12]. saddle-point [BGGGS11]. Sadegh [HSK^{+}15]. sampled [CF11c]. Sampling [RLPM12, EWH14, Gil13, GPS10, GGG^{+}13, HCS^{+}10, LLW14, PDC13, Ran12a]. Sampling-free [RLPM12]. SAMR [KSR^{+}14]. SAT [LN14]. Satisfying [SYV17, LW14, SYV14, ZS10a, ZZS13]. saturated [AST12, HSK^{+}15, Zad11]. SBP [LN14, MAC14]. SBS [MdVBS13]. SBS-active [MdVBS13]. Scalable [BC10a, CST12, KG12a, SPB^{+}10, SS14c]. scalar [BRSS11, BRS13, Boe11, BHL12b, BBM14, BL13, CB12, CLG13, ES14, GMC13, GHM12, GHJ14, KKAS12, KHGW13, LBC14, MB13a, STD14, SG10b, SG10a, SC12, TH14a, VBCM10, ZS10a]. scalar-wave [TH14a]. scalars [VXB14]. Scale [HA11, Abr10, AST13, AKL^{+}13, BHS14c, BPS14, CBKK12, CHKT13, DBM^{+}12, Fat10, HNMM10, He13, Her10, JR14, JC10, JK10, KF13b, LL10a, LG14, LS10a, LHB10, MS10a, MB10a, MK14a, OAKR14, OBT11, RBB12, RG13, RE13, SQC11, SRS14, SSK^{+}12, XHF13, XDZ^{+}14]. scaled [Yok13, Yok14]. scales [AP12a, CG13a, Qua11, RWX11, WW14]. Scaling [PSDF13, ARF12, HCS^{+}10, SSM14]. scapula [SR13b]. scattered [WMI10]. scatterer [VBVD10]. scattering [AV10, AM13, AN14, AH10b, ABH^{+}14, BV13, Bre12a, BG13, BLR14, DWL10, DB12, Du11a, Du11b, DSZ13, EL11, FGE11, FvdB12, GH14, GH11a, GHS12, GB13, GG13, GD10, GL12b, GHL14, GS11b, HNS12, HLS14, HW10, IFD^{+}12, IJZ13, IK11, KL11a, KO14, Le 14a, Le 14b, Lee12b, Lee14a, Li10a, LW13a, LLW14, MTT13, Nic10, RCCC14, SM11b, TY11b, WdHX10]. Scenario [ECD14, ERS14]. scene [JSX10]. scheduled [YM14]. schema [CX13]. scheme [AK14, AiINT14, Ald10, AST12, AH13b, BCD14, BSS14, BKC^{+}11, BD14b, BD11b, BHM14, CN12, CWS13, CD12a, CS12a, CWL10, CJLC11, CCW12, CLZ12, CXLC12, CRG13, CL14b, CS10, CTJT13, CS14a, CS14b, CC14, CDK12, DOR10, DL13a, DL13b, DS10, DS12, DCL11, DLM13b, Era13, EN14a, EN14b, EO11, FSTY14, FH11, GMB10, hGzS11, hGzSnZ12, GW12, GYL12, Gao13, GW13b, GW14, Ge10, GC11, GCZ13, GITW12, GIMT14, Gra14, GVV13, GZZ14, GH12b, GCH14, GT11, GMG12, GXS14, HKLY13, HPA14a, HPA14b, HLT10a, HLM11, HK14, HJZ14, HWA10, HLT10b, HAQ12, HAH14, IL11, IDNG13, JCC14, KTT12, KSL10, KJ11, KD10, KL10b, KAA12, Kri10, KHGW13, KGK13, LNU10, LGH10a, LBM10, LBT11, LTCN13, Lee13, LP12b, LKLG14, LLJ10, LLX11a, LT12, LZ13, LY13b, LD14, LL14a, LEHA10, LJG14, LYXZ14, MKCB14, MMR12]. scheme [MB12, MAB^{+}13, MB11, MN12, Mie13, MT10, Min10b, MMA11, MMA13, ML13, MTS^{+}11, MAD13, MY14, MHGNM12, MHHGN14, MV13, NYM13, NKSR12, NRS10, NMZC13, NL12b, NMT^{+}12, Nor14, NT14, OAK11, OD13, PZTW12, PGL11, PLN11, QL12, RARO13, RP13, RzSZ13, Ren11, RRCD14, RS11, SK13, SBS11, SQC11, SZ11, SZH12, SY11a, SY12b, SCY11, SML13, SSHM10, SR13c, sSLxRyZ14, SBO11, SIDG13, SS10d, TiI11, TH14a, TK12, TK15, TY11a, VXB14, VBGM13, VW14b, WFK11, WA12, WMI10, WYYK10, WYS12b, WXY13, WZZ14, WXY14, WGD12, WYT14, XIC11, XH10, XW10, XX13, YHT11, YT12, YSW^{+}13, YSW14, YWY10, YZ12, YHM12, YYB12, Zer10, ZN14a, ZzS11, ZYS14, ZDSP11, Zie11, vL11, SGW10]. schemes [AINF13, ANIF14, ALR11, Ain14, AN13, ADB14, BK11, Bal12a, BMD^{+}13, BH12, BHL^{+}13, BBC^{+}13, BF12, BR14, BR12a, BK14b, BBD14, BDMV11, CCKCG10, CCGP13, CCD11b, CY12, CSK13, CRT14, CDLL12, CP13b, CSC14, CG13b, CMS10b, CLM13, CR14, Dai13, DEH11, Deg11, Del10, DMT^{+}11, DMM^{+}13, DB13, DA14, DZHB13, DKM11, EAN11, Fan14, FNGV14, FJ10, FCYF11, FC13, FMT11a, FE11, FMM^{+}10, GDHM11, Ger12, GMA10, GP12a, GJLY13, GCL13, GK14, GLWW14, GDGP11, GGT13, HHS^{+}13, Hic12, HY10, HAS13, HLLW11, JTT14, JY11, Joh11, JH12, KM14a, Kaw13, KKvZB14, KWJR14, KJWR14, Kim13, KS13a, Kol11, KVBP^{+}14, LTC13, LEM11, Ler13, Ler14, LQ10b, LR12, LZZS13, LY14, LQ11, MC12, Mar10b, MTB10, Mig14, Mön13, MMT14, MT14a, MSB14, Mor10, MPT13, MGN13]. schemes [Mur10, NMCZ14, NDB12, NDBG14, Nis14a, Nis14b, Nor13, OAXL13, PDSS10, Par10, Par12a, Par13, PP12, PP14, PR14b, QL11, RSD10, RBS12, RZ10, RXLS14, RA10b, Rie10, Sen13, SLT^{+}14, SZ10, SST10, SRL^{+}11, SN14, TD12b, TTD11, TAD14, Tsu10, TGGT13, ULS13, VDD^{+}13, VWPF11, VWF13, VQRD11, VCJ11, VR14, VQVD11, WPP10, WYS^{+}11, WZSN12, WGX13, WV14, WLW13b, WCVJ13, WG11, WDGY10, WG14, Xie14a, XQX13, XQXC14, XS11, XMS12, ZS10a, ZS10b, nZzS11, ZS11, ZS12, ZZS13, ZY13, Zha14, ZQ13, vEKdB14]. Schmidt [AGBC14, CG13a]. Schrödinger [AMDR10, BTX13, CWL13, DBAV14, Don11, GH12a, HZ11a, JWYH10, LQ10a, MJK13, MZB12, QY10, RS13b, RS14b, SSR14, ST11a, SYE10, WXY13, WGX13, WXY14, WJTY10, ZD11a, Zho14b]. Schrödinger-type [AMDR10]. Schumann [Liu11b]. Schur [FMT11b, LTSH10, LYD10]. Schur-decomposition [LTSH10]. Schwarz [HWHW10, JC13, LLP14, VG14, YJK12]. science [CJP^{+}14]. scientific [AC11]. scientists [PZ14]. SCIVAL BUTTERFLY [Ano12-32]. SCIVERSE APPS [Ano12-34]. SCIVERSE SCIENCEDIRECT [Ano12-35]. scrape [MMIW14]. scrape-off [MMIW14]. screens [BL13]. SE [CWL10]. sea [LTS^{+}10, LKT^{+}12, LKLG14, LFLV14]. seakeeping [SCGE13]. search [RS13a, Rot13, SAM10, ZD12]. Searching [Mac10, Phi14]. Second [CS14b, GLWW14, HH12a, JRTY11, JKQ13, ML13, Ols11, RA10b, Wit10, ZCW11, AT13b, BH12, BWC11, BvBZ^{+}10, Bre12b, CS12b, CPCU13, CH10, CSX14, DFNNRdlA14, EO11, GM12a, GH11b, GITW12, HWST12, Hua14, JLZ13, KK10a, KKLR10, KSS14, KKS13, KS11c, KSR^{+}14, KL11a, LKLG14, LWS14, LJG14, LM13, Luo13, MDTC11, MDTC13, MF14, MB12, MAB^{+}13, MT10, MTG11, MWW^{+}13, Nis14a, Nis14b, PSL20, PL10, PRL10, SSHT14, SBS11, TAI11, TDVG13, WPP10, WLW13a, WG14, ZZS13, ZTL13, ZF14, ZWH^{+}12, vHtTBC11]. second- [Nis14a]. second-kind [KS11c, PSL20]. Second-order [HH12a, ML13, ZCW11, BH12, BWC11, Bre12b, CS12b, CPCU13, CSX14, GM12a, GH11b, Hua14, JLZ13, KSR^{+}14, LKLG14, LWS14, LJG14, LM13, MDTC11, MDTC13, MF14, MB12, MAB^{+}13, MT10, MTG11, PL10, TDVG13, WPP10, WLW13a, WG14, ZF14, ZWH^{+}12]. secondary [KPH13, MK14b]. section [LH13, MBMV13]. sedimentary [LMQ^{+}11]. sedimentation [BDMV11, FNKdLB13]. segregated [SS10a]. seismic [KEGM10, LDN^{+}10, Mét11, MBL^{+}14, SCH13b]. Sekerka [BGN10]. Selection [KL14a, LMM14, AR14, EWH14, RA10a, WI13a]. Self [Bal12a, DH13, YTYJ12, Bal14, BBGP14, BDLGC13, CN10, ID10, KKLR10, KA13, LZC^{+}13, MdVBS13, Phi14, RSS12, SBS11]. self-adaptive [ID10]. self-adjoint [SBS11]. Self-adjusting [Bal12a]. self-assemblies [LZC^{+}13]. self-assembly [KA13]. self-consistent [KKLR10]. self-focusing [MdVBS13]. self-force [CN10]. Self-gravitational [YTYJ12]. Self-organized [DH13]. self-organizing [Phi14, RSS12]. self-propulsion [BBGP14]. self-similar [Bal14, BDLGC13]. Semi [Ald10, ABJ^{+}12, CK11, MBD13, PBB14, SZF12, AC13, BK11, BS10a, CCD11a, CdCNH14, CPCU13, CGMQ14, CP13b, CSC14, CMS10b, DHH11, GH12b, GQ13, HLM11, JSA14, KGKS13, KG11, LNU10, LGF11, LVFK14, MV13, NDV^{+}11, PB14a, QC10, QS11a, QS11b, RS10a, RS11, RS13b, RS14b, SH11, SKt10, TS12, TBR13, VXB14, WYS12b, WD11, XQXC14, Zer10, ZYS14, Zie11]. semi- [KGKS13]. Semi-analytical [PBB14, LVFK14]. Semi-automatic [ABJ^{+}12]. semi-classical [RS13b, RS14b]. semi-discrete [Zie11]. Semi-implicit [MBD13, SZF12, BK11, CCD11a, CPCU13, CP13b, DHH11, JSA14, KG11, RS10a, SH11, SKt10, TBR13, ZYS14]. semi-implicit/implicit [SKt10]. semi-infinite [BS10a]. semi-intrusive [AC13]. Semi-Lagrangian [Ald10, CK11, CdCNH14, CGMQ14, CSC14, CMS10b, GH12b, GQ13, HLM11, LNU10, LGF11, MV13, PB14a, QC10, QS11a, QS11b, RS11, TS12, TBR13, VXB14, WYS12b, WD11, XQXC14, Zer10]. semi-staggered [NDV^{+}11]. semiclassical [LX14, WJTY10]. Semiconductor [SRN13, ÁCB12, CGR11, LS10a, MZB12, SO13]. semiconductors [RJG10]. semidefinite [MNF^{+}10]. semilinear [ARR11, JZ13, RT14]. semirelativistic [MTSG12]. semitransparent [ZYT13]. sensitivities [TAI11]. Sensitivity [KWHD12, XMS12, BW14, CEP12, DRB14, HBAP10, KJC13, LA10, MRS14b, PGLK14, SG13, TNK10, VS11, Wan13b, WHB14, WH11, WG11]. sensor [DEH11]. separable [BZKL13]. separated [Val14]. separation [Abr10, Alb10, HA11, KF13b]. sEPIRK [RT14]. sequences [EHA14]. Sequential [NMX12, BVM14]. serial [LKT^{+}12]. Series [BY11, Boy10, CM13, Gno14, RH13, ZHPS11]. servers [RSFS13]. set [BOD11, BW13, BV14, BGG12, CKS14, CP12, CZL10, CWL10, CLW10, CB10, CV10, CF11c, CKR14, CST10, DCVM11, ELM14, FB11, GRT12, HMS10, HQL^{+}10, Hys12, KL12a, KAFB11, LSM14, LP13a, LDM14, LLYE12a, LLYE12b, LWF11, LZZR10, MCLG10, MD14, Min10a, NRS10, NW10, OD13, PHRG13, RB14, SOG14a, SM11a, Sam14, SN12, SZ10, SKR14, TVEC12, VDD10a, VB13, WSS12, WJTY10, WZ10, XYL12, XR14, YYY^{+}14, YL14b, ZT10a, fZlC11, ZLL10, ZWL10, ID10]. set-based [WZ10]. set-front [BW13]. set-volume [WSS12]. sets [ABD14, CCFCM11, SJ10, WMI10]. setting [RSS10]. seven [BY11]. several [BJLR14]. severely [DKC14]. severely-truncated [DKC14]. SGS [PVLT10, PGL11]. Shadowing [WHB14]. Shafranov [PCF^{+}13]. Shallow [FLB^{+}12, AIX10, ABS^{+}14, BNT14, BF12, CVC12, CRG13, CLSX14, CT14, DLGP13, DFNNRdlA14, DMMGM14, DKA12, DLM13b, DLM13c, FNKdLB13, FMT11a, Gas11, IX10, Ion13, JTT14, KJ11, LT11a, Le 12, LBT11, LGLX14, MHGNM12, MGN10b, MGN11, MGN12a, MLGN12, MGN12b, MGN13, OBT12, PE10, PBM11, Pud11, RBS12, RVFK14, SHR13, SST10, SN13b, TS12, TBR13, UJvL10, Xin14, YC11, Zho11, vTCL^{+}10, vV12, RB10, SP14a]. shallow-water [BF12, CT14, Le 12, SHR13, TS12, UJvL10]. Sham [FGZ12, LLYE12a, LLYE12b, LSE13, MNL^{+}13, SL13a, SBO11, ZCS14]. Shannon [GPS10]. Shape [TNK10, WS10a, AJV11, BMK11, BMK12, DRB14, HBAP10, KIG11, MGN14b, STC10a, SM10, TDVG13, WFL10, ZMG11, ZWL10, ZPGO10]. shaped [HZS12, Par12b]. shapes [SHFB12, SJ10, SDH12]. Shapiro [Fal13]. shared [NS11b]. shared-memory [NS11b]. Sharp [CCGP13, Bea12, CDT13, DvW14, FK14, HHA14, HWW10, KSU13, LP13b, LZZR10, SN13a, SM11c, TDVG13, ZJL^{+}14]. sharp-edged [HWW10]. sharp-interface [CDT13, SN13a, SM11c]. sharpening [NKF14, SX14, SHA12]. Shaw [GXX14]. shear [BKW10, CD12a, CLG13, CZ11b, CSKP10, GDY10, HCS12, LW11, LDE^{+}13, RG12, SH14, SL13b, Wan13a]. shear-stress [SH14]. Sheath [PSM13]. sheet [Alb10, LPE^{+}11, LJX^{+}14, PTM13]. sheets [CMG^{+}13, JG13]. shelf [PTM13]. shell [Brü13]. shells [BHS14b, GNGAS10, GBNS14, JNW^{+}14, LT10a, LW11]. Shepp [BY11]. Sherman [CP14]. shielding [NB13]. shift [OMCO14]. shifted [CH10, SHZF12]. ship [PMMB14]. Shock [BD10a, KS11a, KYP^{+}14, PB11, AH13b, CHC11, Coo13, DEH11, HRBK10, Ian11, JLB^{+}10, KS13a, Ler14, NHE14, NVS13, OAXL13, PPWZ11, RCQL11, RZ10, SBC^{+}14, Sco12, SZ11, YKWS13a, YKWS13b]. shock-bubble [HRBK10]. shock-capturing [YKWS13a, YKWS13b]. shock-detecting [DEH11]. shock-disturbance [RZ10]. shock-fitted [NHE14]. shock-fitting [PB11, PPWZ11, RZ10]. shock-stable [KS13a]. shock-structure [Ian11]. shock/complex [SZ11]. shock/obstacle [CHC11]. shocks [Gno14, Ler13, Ler14, SG13, TAI11]. Short [iT14, BDT10, Don11, Gon12, SSL13]. Shortley [JM10]. shot [DRB14]. shrinking [ZD12]. side [SMP^{+}13]. side-by-side [SMP^{+}13]. sided [HH12a, MY13]. Sign [Hua14, SKCC10]. sign-changed [SKCC10]. Sign-preserving [Hua14]. signals [BGM12, BH13c]. signed [ERT10]. silica [KIF^{+}14, MPWK13]. silica-water [KIF^{+}14]. silicon [KIF^{+}14]. silicon-water [KIF^{+}14]. similar [Bal14, BDLGC13]. similarity [Deg11]. simple [BD14a, CLZ12, DV13a, GM12b, GITW12, GIMT14, GT11, KST10, NKF14, Ran12b, TGO13, WYS12b, XIC11, YS12a, YSW14, Zer10, ZS13, KKO13]. Simplex [WI13a, SD14c, WI13b, YD12]. simplified [BLRR11, Don11, GHM14, OLFS13, Tau12]. simplifying [LEM11]. simulate [DRZ14, FRP14, Ich13, TQW14, WYYK10]. simulated [SHFB12]. Simulating [De 10, GWG14, HZ11b, KL10a, LP12b, MNKS13, NW12, RR14, ACB^{+}13, CD13a, CTP13, FMM^{+}10, HKL14, JCT11, KLMJ12, Liu11a, LHPH11, LHMZ10, MAB^{+}13, MCJ14, MRD13, RS10a, SY12c, SM12c, TK12, TK15, VDD10a, VRBZ11, Wei12, WS10c, ZA14]. Simulation [AMDR10, BCL13, CCKCG10, CMS10a, CKD10, FK14, HM13, LZC^{+}13, MFM12, MDH^{+}10, MLBS14, NVS13, PMM^{+}11, SKU13, SBS^{+}13, WYC13, YS11a, YS11b, YM10, Zha10a, ALGB^{+}13, ADSF13, AST12, ABK12, ABK13, AKL^{+}13, BVM14, BPM13, BHM13, BBCH^{+}12, BGGW13, BA10, BBC^{+}13, BI11, BCD14, BCL14, BPOS12, BJLR14, BHM14, BP10, CCD11a, CKS14, CCFCM11, CD12a, CD10, CBKK12, CDT13, CWL10, CW10, CWC11, CW14a, CPCU13, CWZ^{+}11, CVI^{+}13, CGF^{+}13, CBCM^{+}11, DR11, DLDL10, DV13b, DD12, DM13, DLM^{+}13a, DWZ10, DL10c, EBGW13, FHL10, FWLP14, Fus13, GLLP14, Gao13, GCA13, GCFJ13, GP12a, GW10, GHM13, GP13, HS13a, HSD11, HRBK10, HL11, HLDP14, HB13, HH12b, HH13, HA11, HCS12, HHL14, HWHW10, Hys12, JNSA12, JTT14, JYHT10, JEU11, JLS^{+}14, JK10, JMGN13, KH13, KSL10, KM13]. simulation [KF12, KSP13, KH11, KL10b, KO12, KHJ^{+}13, KA13, KHHK12, KAG14, LFL11, LA10, LCNK10, LSBJ14, LHY13, LMZ14, LWZ^{+}13, LFRM10, LZV12, LGH10b, MB10a, MS14a, MMSI13, McD14b, MTT13, MGLS12, MVS13, MJ13a, MF10, NT14, Oli14, OL13, OTV13, OAK11, OLPM14, PZW14, PZG14, PGLT10, PL14, dlLPC10, dlLPCP12, PPWZ11, RVB10, RG13, RZ10, RRD^{+}14, SBB13, SPS14, SZF12, SR13b, SZ10, SFCF10, SYJ11, SO13, SPF10, Shu14, SCH13b, SW10, SS13b, SS14b, Ste14, SSK^{+}12, SSY13, STC10b, TSL^{+}14, TE14, TD12a, TAAY14, TRSMM^{+}13, TRL14, Unf13, VDD10b, VBGM13, VSC^{+}11, WPKK12, WWX^{+}10, WDP^{+}14, WK13b, WS14, WTSZ10, XX12, YSW^{+}13, YSW14, YL10, YJK12, ZYHW13, ZN14b, ZMD14, Zoh13, vRLPK11, RS10b]. Simulation-based [HM13]. Simulations [FSL^{+}12, GCvRK11, MMV^{+}13a, MT14b, Qua11, RG12, ARR13, AAI10, AHS14, ADT14, AWK^{+}11, AJG^{+}13, AW14, BHB11, BCHM14, BB12a, BHM11, BNM14, BILM14, BPQO13, BPS14, BSS14, BSCML13, BLJ11, BLR13a, Bre12b, BM14, BLY10, CB11, CNR10, CCL^{+}14, CY14, CVIA10, CSK14, CGMQ14, CBGK13, CL13, CSB14, CLAL14, CGC^{+}12, CMBW13, Dav10, Dav15, DHN11, DD12, DC13, DXB^{+}13, DTU12, DS12, DKC14, DBM^{+}12, ECD14, ES13b, FBM14, FLB^{+}12, GB14, GV13, GLC^{+}11, GAV13, GK10, Gri12b, GFK13, GWK10, GH12b, GKS^{+}11, HAD10, HCS^{+}10, HCH^{+}10, Hut14, IFD^{+}12, ILM^{+}13, JLB^{+}10, JMV^{+}12, KSS14, KF13a, KM11, Kea14, KLP10, KIF^{+}14, KLPS14, KSB12, KWH10, KA12, KHRD11, KG14, KYP^{+}14, KF13b, KGSS10, KGK13, Lan14, Lap12, LL10a, LY13a, LH10b, LLX11a, LEH12, LC13, LS10a, LT11b]. simulations [LHB10, LBZ^{+}10, LWCR10, LYXZ14, LDDBR13, MHL^{+}14, MMM11, MdVBS13, MOV12, MGN14a, MJK13, MJ14, MMA11, MMA13, MVCFM13, MJ13b, MBGW13, NKSR12, NS14a, NS11b, OK12, OBT11, PPS12, PAG11, PSF^{+}14, PC13, PA12, PA14, QB14, RSS12, RXLS14, RBNS^{+}12, RBNS13, RRCD14, RPC14, RHGT10, RDHK12, RSFS13, RB10, RBCK10, SNS10, SBN12, SKN13a, SM12a, SM12b, SJB11, SNBN13, Sch13a, SD14a, SCGE13, SQC11, SSB^{+}12, SC11, SHA12, Sog14b, STD^{+}10, SG13, SSK^{+}12, SS13c, Tau12, TCW^{+}10, TDL^{+}14, VHG13, VF10, VKBK11, VCS14, WZG^{+}13, WP10, WRH10, WZT^{+}14, WY10, WY11, XDZ^{+}14, YCY11, YFL^{+}13, Yok13, YBK13, Zha14, ZYF^{+}10, ZJT10, vydCNS^{+}10, CP13a, Den11a]. simulator [AGSG11, LS10a]. simultaneous [RUL^{+}13]. Sinc [ZN10]. sine [CCC12]. Single [HPA14b, BB13, GCvRK11, HL11, KSPP10, KO12, LC13, LLSW14, PRH^{+}14, PBM11, SSS10]. single-block [SSS10]. single-cluster [KO12]. Single-cone [HPA14b]. single-drop [BB13]. single-particle [LC13]. single-phase [PBM11]. singular [AM11, BG12, BG13, DTYY14, FO11, HdCNT12, YXZ13, ZN10]. singularities [LQ11]. singularity [RM12]. sink [CD12a]. sintering [ZKKF10]. six [HdCNT12, PS14a, PS15]. six-dimensional [HdCNT12]. six-equation [PS14a, PS15]. sixteen [KM14b]. sixth [NDBG14, SZ10, sSLxRyZ14, TGGT13]. sixth-order [NDBG14]. Size [VLM13, JWNL11, JW12, NIN^{+}14, RPV13]. size- [JWNL11]. Size-velocity [VLM13]. sized [JLS^{+}14]. Skew [Mor10, MK14b]. Skew-symmetric [Mor10, MK14b]. skewed [Cha14]. skewness [DvW14]. skill [BGM12]. SLAP [Ald10]. Slater [Don11, ZD11a]. slender [BM11, FRP14]. sliding [BD10b, DL10c, FW12]. slip [RPV13, SR11, UGF14, VM10, Vre14]. slope [Kuz14, MMD^{+}11]. slope-limiting [MMD^{+}11]. slopes [JMGN13]. sloshing [KGG10]. Slow [HNMM10, Abr10, Chi12, LRS10]. Slow-time [HNMM10]. small [AST13, BBD11, FF13, GW13b, LKTL10, LHB10, MGJZ10, Wan13a]. small-perturbation [LKTL10]. small-scale [AST13, LHB10]. smallest [And10]. smart [EO11]. Smith [KSPP10]. smooth [BL10, GPY14, LB10, Mel11, RSS13, ZN10]. Smoothed [OS11a, PTM13, PBT14, Pri12, AHA12, AHA13, BLOdV14, BPHK13, LXSR12, LEHA10, LBK10, Ros10, Soa13, TP12, Zha10a]. Smoother [vdVR12a, vdVR12b]. smoothers [BGR11, Lan14]. smoothing [TM10, ZL13, ZWJW10]. smoothness [FSTY14, HKLY13, iIT14, ZJW12]. Sobolev [CJM13, MS10b, MS11, RSS10]. socket [LS10a]. soft [LG14].

Software [OK13, DKH11]. SOI [GM11b]. solenoidal [AH10a, SCIE14]. solid [AGBC14, ABHS12, BOD11, BN11, BSDM12, CWC11, CHHL11, DLM^{+}13a, DKW10, ETW10, FG12, GPY14, GM12a, GKF11, GF13, HW14a, KK14, KSR^{+}14, LZC^{+}13, MNKS13, MMH12, OLPM14, SSY13, WZT^{+}14, ZL10]. solid-fluid [ZL10]. solid-liquid [HW14a, SSY13]. solid-solid [KK14]. solid/fluid [BOD11]. solidification [BKKD14, JLS^{+}14, RR14]. solidifying [KBTD12]. solids [BHS14a, HPOM10, LYE10, ML14a, MRD13, MVS13, SSB13, SNBN13, VSC^{+}11]. solitary [BX12, SC12]. solitons [BTX13]. soluble [BS10b, CL14b, GT12, KT14, LZ10, MT14b, TSLV11]. solute [CVC12, GMG12, MKCB14]. Solution [BFE12, CGG12, GHH14, MZB12, RBS12, SHZF12, SC12, ZD10, AJ14, AS14, AK12, AME14, AT13a, ABH^{+}14, ABPSM11, BSSW12, BGR11, BFO14, Bér11, BS11a, BSB14, BHBM10, CX13, CD13b, CGC13, CRS14, CLS13, CTSM13, Dad12, Deg11, DBAV14, DHH11, DFW^{+}10, DTYY14, EL11, FT12, FM13, FWP13, GG10, GLS10, GDW11, GPS10, GKG^{+}13, GL12a, GHJ14, GCH14, HM11, HP13, HGW11, He13, HLS14, HKM12, HOK10, HW10, HZS12, HLW13, IL11, IDNG13, IG14, JWV12, JTH10, Kat14, KBGRM10, KG13c, Kri10, KO14, Le 14b, Lee13, LBCL13, LH10c, LW13a, LLP14, LRK11, LTL11, MZ10, MMV^{+}13b, MA13, MAK14, MM11, MSTL13, MY13, MVG11, MJ13b, OK14, OS12, PPB14, PE10, PMSM12, PLN11, Pud11, RE13, Sal14, SVS11, SCY11, SSX13, SR13c, SW13, Stu10, SP11]. solution [SS14c, Tal13, TWMM14, Tok11a, TE11, Wal13, WKMJ13, WLW13b, WG11, WS11, Xie14a, YC11, ZN10, ZBS12a, ZBS12b, ZYLK13, ZTL13, ZPF10, ZMD14]. solution-adaptive [GG10, IDNG13, IG14, MJ13b, WS11]. solutions [AEM13, AH10b, BOK12, BDLGC13, BHL12b, BKV12, CJP^{+}14, CM10, DXB^{+}13, EC13, FO11, Gno14, GAW^{+}14, HM10, JHDC10, KS11b, Kol11, KS13b, KG14, KRT14, xLqW13, LHMZ10, LHV13, MS11, MP13, MGN10b, PBB14, SRSV11, SYL10, SYL12, SHM12, SSR14, SPM^{+}13, TCL11, VK12, WCM^{+}14, WC14a, WTK14, WWS^{+}13a, XX13]. solvable [GM11a]. solvated [GK13]. solvation [CBW10, CLW10, DXB^{+}13, Zha14]. solve [BR11, CGJ11, GMG12, LLJ10, NLGL14]. Solver [MMJ14, AH13a, ARR11, AAI10, AiINT14, AB11, ABS^{+}14, Bal10, Bal12b, BDA14, BHL12a, BD14a, BLM^{+}14, BYCC14, BKH12, BLTO11, Bre12a, BL13, CCT11, Cap11, CMS10a, CK10, CCW12, CDB13, DDM10, FGZ12, FW12, FMT11b, GK13, GMC13, GM12a, GB13, GG13, Gra10, GHL14, Guy13, HGSK^{+}13, HSN13, HRCW13, HG11, HMHO13, HPOM10, HHK13, HPLS10, HT12, KB14, KKO13, LP13b, LXF10, LLX11a, LFLV14, LL12b, LL10b, Mar13, MGN14a, MB10b, MHGM14, MLGN12, NNS11, OR10, OLPM14, PCF^{+}13, PBM11, Rie11, RB13, SBB13, SV13, SY12a, SY14a, SABH11, SP14a, SLI10, SYY14a, SYY14b, SSG11, SGL14, TWMM14, TT10, UN13, VV13, VED13, VT13, Vol14, WT13, WHT14, WSW13, WP14, YS11a, Yao14, Yin14, YYB12, YXD^{+}14, ZVNM10, ZCW11]. Solver-based [MMJ14]. solvers [APGL12, AD14, BS14b, BBD14, BL10, EMS11, FO11, JS13a, LL14b, LB10, MCLG10, MCDT12, MGN11, MGN12a, NTV14, NGV14, OS11a, PG11, PBM11, PNPF10, PVDW14, PR14a, SHZ13, SF12, SHM12, SEDW13, TDU14, WKH10, YS11b, dPSS12, vVVK10]. Solving [AM11, BWMG14, CT13, JU10, LS13c, MBR11, SSS10, XM14, BHS12, BY11, BKW10, BL11, CEGG14, CTG12, CL14b, CW14b, CL11b, CCG11, DMHL14, DAEB13, ES13a, EMSH14, GCZ13, GM10b, Gri12a, GT11, HAMA14, HHS^{+}13, HHGC14, HWW10, HSWZ13, HAS13, JC13, JW11, KSU13, Kha13, LPE^{+}11, LKT^{+}12, LLZ11, LTSH10, LRLL11, LT12, LR12, LD14, LSE13, LL14a, LWS14, LBM14, LYD10, MMM13, MS14b, MM12, Mön13, NJ12, PDZ^{+}14, PVLT10, PGL11, PRL10, Pir12, QY12, RM12, SS11, SML13, SKCC10, ST11a, SYE10, SIT^{+}11, TDVG13, TY11a, VMA14, VBCM10, WLW13a, WPK14, XLD^{+}11, XCL14, XJ14, YVL^{+}11, YO11, YXZ13, YK12, ZMQ13, ZTG^{+}13, ZYS14, ZLH12, ZCS14]. Some [GCW14, LWS14, XH13, CHHS11, JS13a]. sonic [AL10, PMM^{+}11]. Sørensen [OBT12]. sound [KVM12, SM11b]. Sounding [HMOH11]. soundproof [SKW14]. Source [BCM12, BCB12, CD12a, CEK14, CM10, HQL^{+}10, LHB10, LQ11, LBS10, MPS^{+}14, MAD13, MCDT12, MGN10b, MGN12a, Nis12, NF12, RF14b, WSYS12, WKMJ13, YKWS13a, YKWS13b, ZS11, ZMD14]. source-and-sink [CD12a]. Sources [LKCY11, CB12, CG13b, FJ10, KNR11]. South [OMCO14]. Space [AS14, SBvdV11, AJ14, BBC^{+}12, BB12b, BL13, BTY14, CP13a, CB14, CRS14, CXLC12, CH10, CKR14, DBM^{+}12, DZHB13, FL11a, GAvdVB14, hGzS13, GG14, HPA14b, HBM13, HL12a, ITO^{+}10, JK10, LE13, Lap12, LSW14, LS13d, MWP14, MK14a, MdVBS13, MST14, MY13, NIN^{+}14, NB13, Phi14, QY12, RSS13, RC12, RCvdV13, Rot13, SHR13, SSS11, SAHS14, SSR14, TH14a, TA12, TAAY14, TvdHS^{+}12, TD12b, TTD11, UGkM13, WYYK10, WXY13, WD13a, WD13b, WD14, WXY14]. space-filling [MWP14]. space-fractional [MY13, WD13a, WD13b, WD14]. spaced [BH13c]. spaces [SL13a, TG14b]. spacetime [KMHJ10]. Spalart [CDB13]. spall [CWL10]. Sparse [GS11a, LP12a, And10, ABJ^{+}12, BGZ12, BK12b, BS11b, DO11b, FGR12, JAX11, KL14a, LWD13a, LWD13b, LTT10, PHD14, Per13, ZCS14]. sparsely [BM13b, BH13c]. Spatial [LM13, ANIF14, BHY14, BMR14, CZ11a, CTJT13, Den11b, DSHP11, FWLP14, HL12a, JEU11, LY13b, LS13c, MMR12, MLM12, SLH^{+}11, TDU14, WI13b, YB14, ZY13]. Spatially [Atz10, LEH10, LEH12, PHSA14, TA13, AJ14, GHM10a, MCJ14, SR13a, TGO13]. spatially-extended [MCJ14]. SPDEs [CSK13]. special [AQ11, QY12, Ros10, ZT13b]. special-relativistic [Ros10]. species [Boy11, Tok11a]. specific [AKL^{+}13, HK11, IWG13, KVC^{+}13]. Spectral [CD13b, CCC12, GHM12, GCL13, JCC14, JHZ10, KSK13, LT11b, WX14, AB11, AD14, BdF10, BY11, BL11, CCD11a, CYK^{+}14, CS12b, CS13, CKR14, CR14, Dri10, EGW11, EN14a, EN14b, FDHP11, FD11, GH14, GYL12, GG14, GVH14, Gra10, GTSC14, GKS^{+}11, HSN13, HNS12, He13, HLS14, HBZL14, JNM12, JZ11, KGG10, KG13b, KSB12, KG13c, LR11, LBC14, LP12a, Le 14a, LB12, LTSH10, LCP13, LZZS13, MA14, MRS14a, Mar10a, MKGV12, Mar13, ML11b, MTZ14, MK11, MLBS14, MLM12, MHGM14, NHE14, NCV10, Ost10, OS12, PGLT10, PGL11, PE10, PL12, QZM^{+}12, SCH13b, TF10, VHG13, VSK10, WA12, WWS^{+}13a, XH14, YK12, YWL14, YXD^{+}14, ZK14, ZIOF10, ZD11b, ZW10b, vRLPK11, BMS12, MOSW12]. Spectral-Difference-Method [MOSW12]. spectral-element [EN14a, EN14b, ML11b]. spectral-Galerkin [CS12b]. spectral-Lagrangian [MHGM14]. spectral-like [CR14, FDHP11, LZZS13]. spectral-particle [LBC14]. spectral-tau [LP12a]. spectral/ [CYK^{+}14, VSK10]. spectral/DG [CKR14, YK12]. Spectrally [HF10, Le 14b, LT10b, Boy11, RMMD14, GM11b]. Spectrally-accurate [HF10, Boy11]. spectrum [BD14c]. spectrum-including [BD14c]. speculation [KH11]. Speed [BMD^{+}13, CDT13, CJLC11, CDK12, KSU13, ML14a, NHE14, SKU13, sLwG10]. speeds [LNS13]. SPH [PR14a, AHA10a, AHA10b, ADB14, HF11, LVFK14, MCLG10, MCA^{+}13, PR14b, SSY13, VL14, ZMD14]. sphere [BNT14, DKA12, FCH13, FL10, FLB^{+}12, GM14a, GHJ14, HLM11, IAM^{+}12, IDNG13, JNM12, LNU10, LLJ10, LJ13a, NL10, NCV10, SP14a, TS12, UJvL10, WD11, YC11]. spheres [RS13a, Wil13]. spherical [Brü13, CCD11a, CLSX14, EN14a, EN14b, GNGAS10, GBNS14, GS11a, GDGP11, JNW^{+}14, LdWF12, Li12, MH10, Mel11, Mig14, MM11, MTS^{+}11, Ols12, PB13a, RARO13, RJG10, Sal14, SJ10, Tyg10]. spherically [Don11]. spheroidal [Men12, WZZ14]. spiciness [OMCO14]. spin [BCZ13, BDTW10, CCW11, Gil13, Wan11b, Wan14, WC14a, Wei12]. spin-1 [BCZ13, CCW11, WC14a]. spin-2 [Wan14]. spin-up [BDTW10]. Spinodal [LDE^{+}13]. splashing [Yok13]. Spline [ZSW10, BSV14, DL13c, FKM11, FF12, LTC10, YZX14]. splines [EH13]. split [BB12b, FGLB14, FCN^{+}13, HK14, Pir10, RT14]. split-step [FGLB14]. Splitting [KHZ10, BCL^{+}11, Cha13, CGMQ14, FGN14, GLWW14, HLDP14, HJK14, Kri14, LX14, LEHA10, Par13, PGA11, RXLS14, Sal14, VV13, Wal13, WPP10, WYS12b, ZOL^{+}11, Zha14]. sponge [Man12]. Spray [MW12, JW12, KLMJ12]. sprays [VLM13]. spring [iIT14]. Spurious [Le 12, TCL11, YKWS13a, YKWS13b, HP13, LKCY11, LY13a, Min13, SM11c]. SQP [BVM14]. square [BY11, CLAL14]. square-to-disk [BY11]. Squares [CD13b, WHB14, ABL^{+}11, CXLF14, Hel11a, HMM^{+}10, HY09, HY11, JMFO13, NOT14, OS12, PR11, ADB14]. squares/finite [HGW11]. squircle [LB14]. Stability [DWL10, LJ13a, Alm14, BK11, BB12b, BRVE12, CCKCG10, GLSC12, GV13, GVH14, HP13, KG11, KG13a, LXSR12, MS10a, MA13, NDB12, VED13, WFK11, WL11, ZKV10]. Stabilization [CST12, DL12, FL11b, TM10, Vre11, BC14, Gui12, HH12a, MKGV12, Ren11]. Stabilized [AB13, ASS13, ASS17, HRK^{+}10, Pir11, BCP13, CH11a, CdCD12, HY09, HY11, MMV^{+}13a, MMV^{+}13b, MBA14, PBC11b, SPB^{+}10, Sha13, WTSZ10, WGD12]. Stable [BN11, CVI^{+}13, DV14, FN10, GL12b, GK14, LLP10, MTD14, PVDW14, SSR14, XH14, BMS12, BHS13, BHS14a, BHS14b, BGN10, BHL^{+}13, BL10, CN12, DS10, EAN11, FCYF11, FC13, FMT11a, GW12, GH11b, GKF11, GZZ14, HNS12, KKS13, KB14, KS13a, KHJ^{+}13, Lee13, LGF11, LN10b, LY14, LJG14, LB10, MAD13, MY14, MWY14, NMZC13, NDBG14, RHB12, SHR13, Sco12, TD12a, WBC13, WCVJ13, ZFH14]. STAG [CB10]. stage [IJZ13]. stages [WDP^{+}14]. staggered [Boe11, CCY13, HM10, JGZL14, KS14, Lee13, LT11c, MLBK14, NS11a, NL12b, NDV^{+}11, OS14, PB14a, SVK14, SML13, TH14a, Tyl14, VW14a, VKL^{+}13]. staggered-grid [TH14a]. staggering [RBS12]. Standard [WR11, Del14]. standing [RW13]. stars [BD11a]. starting [Ton14]. state [BJ12, CEK14, CWC11, CCK13, Don11, EFT13, FT12, GMXG14, HHS^{+}13, LTCN13, LCWN12, NJ12, NEE12, PDC13, VED13, WD13b, Wan14, WP10, WTSZ10, ZD11a, ZLCW14]. state-dependent [CEK14]. states [AD14, BCW10, BD11a, BCZ13, CCW11, DH10, Gil13, KTT12, MK12]. static [ADMA13, DL10c, HSN13, RARO13]. stationary [BCHM14, BS14a, CT13, DMT^{+}11, DMM^{+}13, HL12a, OBT11]. Statistical [DF10b, WLPG11, AA13, BJ12, HMM14, LTC13, MWRZ13]. statistically [GM11a]. statistics [Lee14b]. Steady [Ler13, Ler14, ALR11, CCK13, EFT13, FSL10, GBC11, HHS^{+}13, HY09, HY10, HY11, HLT10b, Hu13, JK11, KTT12, KKO13, LT12, LCWN12, LDA10, LFW11, NJ12, NEE12, PVLT10, PGL11, Sha13, STH11, VED13, WD13b, WTSZ10, XH13]. steady-state [NEE12, VED13, WD13b]. steep [JMGN13]. steepening [SHA11]. steepest [WJC13]. steered [AM10]. Stefan [BGN10, LYD10, PGR10, PHRG13]. Stefan-type [PGR10, PHRG13]. stellar [FSL^{+}12]. stencil [GW13b, MDTC13, WI13a]. stencils [Cha14, DO11a, PJCS14, Tsu10]. step [AKP^{+}12, BBG^{+}11, CJN^{+}13, DWL10, FGLB14, GDHM11, JTT14, KN14, KS12, LCK11, LLH10, LDDBR13, MHGNM12, NNS11, OM12, Par12b, QL11, QL12, RB10, RM14, VL14, WZG^{+}13, YKWS13a, YKWS13b, MHHGN14]. Stephen [ZJS15]. steplength [BHM11, BHM13]. stepping [AB11, DS12, FP14, HRBK10, KA13, Kri10, LMGS13, MBA14, SK13, TAAY14, ZMQ13, AAT^{+}14]. steps [Lee14a, LP12b]. stepsize [CTP13]. stiff [AV13, AB13, CF11a, FJ10, Gui12, Har11, NS14a, WSYS12, YR13, YKWS13a, YKWS13b, YYB12, ZLCW14, ZOL^{+}11]. stiffly [Min10b]. Stochastic [Atz11, BM13b, FSL10, GM14b, HW14c, MAK14, PWK11, Sti12, AP12a, AB13, ACG14, AP11, AA13, Atz10, BCK11, BKW10, BH13c, Cam13, CHZ13b, CHZ13a, Dad12, Dav10, Dav15, DRW11, DO11b, EWH13, ERS14, FHL10, FM13, FWLP14, GHM10a, GHM10b, Gri10, Gri12a, Har11, HHGC14, HJZ14, JAX11, JNX13, JTH10, JEU11, KKPV12, KL14a, KG13b, KF13b, LHH14, LG14, MZ10, MZ11a, MZ11b, MPWK13, ORHH13, PBB14, PIN14, PHSA14, PL12, RWX11, RNT12, SAM10, SS14c, ULS13, Val14, VDD10b, VK12, WZ13, WZ14, WI13b, WI13a, YCLK12, YK13, YR13, ZYLK13, ZT10b]. Stokes [HGW11, KRF12, LGE^{+}13, PSL20, AB11, AT13b, ABPSM11, BK11, BSSW12, Bea12, BNNW11, BN11, BN14a, BGHM13, Boe11, BR12b, BOK12, BCLR10, CVN13, CMS10a, CJNS12, Cha13, CFKL11, CXLF14, CLS10, CDK12, CH14, CST12, DLZ^{+}11, DHH11, DAA11, EH13, FW12, FL11a, FMT11b, GLL14, HL12a, HOK10, HHK13, HM10, HKL14, HF10, JM12, JK11, KST10, KR10, KKP10, KF13a, KS11c, KG13c, KRT14, KL11b, LBCL13, LJX^{+}14, LXF10, LL12a, LT10b, LLN^{+}10, MBS13, NPC11c, NOT14, NDBG14, NCKN10, NEE12, OA10, OR10, OLC13, OS12, PBT14, PVLT10, PGL11, PNPF10, PGA11, RP13, RHXQ14, RGMC13, RCvdV13, RPC14, SK12a, San13, SVK14, SH14, SSHT14, SM13, Sen13, SSS10, Sha13, SLI10, SR11, SY11c, SGV13, SW13, TWMM14, TY11a, TLO^{+}14, UN13, UGF14, VV13, VQRD11, VT13]. Stokes [Vre14, WW13, WK13a, WX14, Wil13, XFB^{+}14, XIIX14, XH13, ZW10b, ZLH12]. Stokeslets [BM10b, BM11]. stopping [DFVY14]. storage [Ket10, NDB12, WT12a]. straight [Gri13]. Straightforward [LFL11]. strand [KWS^{+}11]. Strategies [MM12, BP10, Har11, RBS12, SKN13b]. strategy [BRP14, CD13a, DHH11, DBM^{+}12, HL12b, Par12b, PVLT10, PMF13, RE13, SPM^{+}13, ZMQ13]. stratified [ABPSM11, DFNNRdlA14, FMM^{+}10, MMV^{+}13b, MTD14]. Stratonovich [ZYLK13]. Stratonovich- [ZYLK13]. stray [EAB^{+}12]. stream [WZ11]. streamer [DBM^{+}12, KED^{+}12, LEH10, LEH12, LE12, VBGM13]. streamfunction [TY11a]. streaming [MZ14]. streaming-collision [MZ14]. Stress [ZM10, BDM13, GDY10, MY14, NGV14, SH14]. Stress-mediated [ZM10]. stressed [JR14]. stretch [OLPM14]. Stretched [FvdB12, RVbZ10]. stretching [RG12]. strictly [BMR14]. Strong [ACCM12, LJG14, NEE12, RM12, SSR14, TWMM14]. strongly [AD14, CGJ11, KR10, KKvZB14, MDPTK13, MMP11, ZTL13]. strongly-coupled [KKvZB14]. Structural [GCBH10, GCBH13, BHS14b, BGBR12, CLG13, DYYA14]. structure [Atz11, BGGGS11, BK12a, Bal14, BHS12, BC10a, BG10, BW13, BBGP13, BHK^{+}10, BGG12, BCG^{+}13, CBPS13, CW13, CW10, Deg11, DEKBF14, EHA14, EC13, FL14, FP14, GCBH13, GCW14, GKF11, GZQ13, HHA14, HNS12, Ian11, ITO^{+}10, JLY10, LS13a, LL14a, LJG14, MF14, MMH12, NL12a, NPV14, OM12, PHSA14, Ric13, RMSF11, Son11, SIT^{+}11, TDL^{+}14, VRB10, WKL^{+}14, WYC13, WC14b, YS12a, YBK13, ZF12, ZCD10]. structure-preserving [CW13, GCW14]. structured [BMD^{+}13, Bal14, CG12, EZ10, GH12a, HB13, PJCS14, RTKS10, SSS10, WYS12a, WQCS13, ZJW12, ZD11b]. structures [AMDR10, And13, BHL12a, CKS14, DB12, GHL14, GKF11, GF13, HZS12, HW14c, KJC13, KL14b, LRL10, LCNK10, LX13b, MB13b, MYM12, PS14b, SM12c, MVG^{+}12]. studies [BRS12, DMM^{+}13, EH14, EH15, SBC^{+}14, SY11b, VWP11, XS15]. Study [vTCL^{+}10, BK12a, BRSS11, BKW10, CSK13, CRS12, GKG^{+}13, GMXG14, JR14, JHDC10, KGSS10, LKT^{+}12, LCE^{+}10, MDPTK13, SF11, SMF^{+}13, SSC13, STC10b, TA12]. studying [CSKP10, DYYA14, LW11, YL14b]. Sturm [HAMA14, JR12, ZK13]. Sub [BK11, ABMT12, BHS14c, CTJT13, hGzS11, hGzSnZ12, hGzS13, MAD13, RzSZ13, WV14, YZX14, nZzS11, ZzS11]. sub-cell [CTJT13]. sub-diffusion [hGzS11, hGzSnZ12, hGzS13, MAD13, RzSZ13, WV14, YZX14, nZzS11, ZzS11]. sub-grid [ABMT12]. Sub-iteration [BK11]. sub-scale [BHS14c]. Subcell [WI13b, DZLD14, WSYS12]. subcells [Mac10]. subdiffusion [BLY10, DL13c]. subdivision [BBC^{+}13, LW11]. subdomains [AS12, SAA14]. subexpression [Opp13]. Subgrid [GM11b, SRS14, JWNL11, RG13, Sha13, TVEC12]. Subgrid-scale [SRS14, RG13]. subgridding [Bér11, TD11a]. subject [PBT14, PHSA14]. submatrices [Ran12b]. submitted [BBD11]. subscale [APC11]. subspace [BB13, DR14, DRZ14, MNF^{+}10, PB14b, ZD10, ZCS14]. substrate [DKW10, GWG14, LZC^{+}13]. substructuring [MP13]. subsurface [CD10, EWH13, EWH14, HSD11]. subzonal [DL12]. Sufficient [SP11]. suitable [KAFB11, XMDG13, XMDG14]. suite [BGM12]. suited [ARR11, MZ11c]. Summation [NL13, BN12, DBZ14, Hic12, LT10b, MA13, Mat14, MSD12, SN14, WLL13]. Summation-by-parts [NL13, BN12, DBZ14, Hic12, SN14]. summing [Boy10]. sums [GD14]. super [MBA14]. super-time-stepping [MBA14]. supercomputer [XDZ^{+}14]. supercomputers [KSP13]. Superconvergence [GZQ13, HLYS11, CCY13]. Superconvergent [BN12]. supercooled [ARF12]. supercritical [LX14, TK12, TK15]. superfast [WD13b]. superfast-preconditioned [WD13b]. superhydrophobic [CL13]. superlattice [Pri14]. superlattices [ÁCB12]. supermesh [MF12]. superparameterization [GM14b, MG14a]. Superresolution [BM13b]. supersonic [DKR12]. support [FE11, MNKS13, SY14b]. supported [Vai11]. Suppressing [GV14]. suprathermal [PLT14]. Surface [BPHK13, AHA10b, ABKF11, AST13, ABMT12, AKMD14, BCD14, BLJ11, BP10, CL14b, CF11c, DS11, DKW10, DWZ10, EAG14, ES10, GAvdVB14, GVV13, HLS14, HVD13, HMHO13, IK11, KS11c, Li12, LXSR12, LFRM10, MCLG10, PSL20, PDZ^{+}14, PZ10, RA10a, RB10, RB13, SF11, SZF12, SST10, SGV13, SP11, VB13, WY10, Yok13, Yok14, Zha10a]. surface-bulk [CL14b]. surface-tension [AHA10b]. surfaces [BR12a, BG12, CYK^{+}14, CL13, DKW10, Gen13, HK14, LW11, LL10a, LLZ11, MBR11, MKB^{+}13, NW12, Pir12, TR11, WWZ13, YW13, YHM12, ZBS12a]. surfactant [AHA10a, BS10b, CL14b, Fuj13, GZZ14, KL11b, XYL12, XR14]. surfactant-coated [KL11b]. surfactants [GT12, KT14, LZ10, MT14b, TSLV11]. surrogate [LX10, LLX11b, SAM10]. surrogate-based [LLX11b]. surrogates [Val14]. suspended [LLdSC11, RVB10]. suspensions [GLLP14, Kea14, QB14, YM10]. SVD [WYYK10]. swapping [GIQ11]. sweep [VG14]. sweeping [CCK13, DGM13, EFT13, LQ11, Luo13, LQB14, TEY12]. sweeps [CBTR13]. swimmers [GCvRK11]. swimming [BI11]. swirling [CRT10]. Swiss [AV13]. Swiss-knife [AV13]. switching [CEK14, GBC11, HSK^{+}15, SCY11, Zad11]. Symmetric [EGL^{+}13, Don11, GM12a, HK14, Mor10, MK14b, NCKN10, NCV10, PGR10, RMSF11, SHZF12, YHM12, HS13b]. Symmetrizing [LB14]. Symmetry [TLO^{+}14, VKL^{+}13, ANIF14, CS10, CS14b, CHHS11, GDGP11, KBS14, LB14, LE12, VW14a]. Symmetry- [VKL^{+}13]. Symmetry-preserving [TLO^{+}14, CS14b]. Symplectic [ST11b, Web14, ZJS15, ZTS^{+}10, CWS13, GCW14, HJZ14, HDHZ13, MLM12, SWX12]. synchronized [LSVW10, OS11b]. synchronous [MMM11]. synoptic [OMCO14]. synthesis [CP13a]. synthetic [EMSH14, TR14]. synthetic-acceleration [EMSH14]. System [ASXZ14, ABL^{+}11, ARF12, AAG14, ABPSM11, BSSW12, BH13b, BHM14, CWL13, CSW14, CCZ14a, CCZ14b, CS14c, DDN^{+}10, DHN11, Del10, DMR14, Don11, FNKdLB13, GZZ14, GCH14, GXX14, HGW11, HGMM12, HHK13, HY10, JLS^{+}14, MRS14a, MS10b, Nis10, QS11b, SV13, SS10d, Wol12, XXS10, ZD11a, ZJW12, ZN14a, ZN14b]. systematic [SF11, ZF12]. systems [AVKP11, AF11, Atz10, BH13a, BD14b, BGBR12, BM13b, BP10, Cam13, CK11, CS12b, CLZ12, CNST13, CAH14, Chi12, CH10, CHHL11, CDL11, CLS13, DLGP13, DBO^{+}10, EFT13, ERS14, FF12, FWLP14, GMT10, GHM10a, Gil13, Gra14, GLM14, HMHO13, HM13, JHDC10, KKPV12, KJC13, Kim13, KL14b, KL12c, LCWN12, MDPTK13, MS11, MMM11, MMD^{+}11, Min10b, MSS12, MCDT12, MLM12, NV14, NMV13, NMCZ14, NF12, NT14, Par12a, Phi14, PA12, RT14, RM14, SP13, SHZF12, TAT14, TCW^{+}10, TDL^{+}14, TLNE10, Wan13b, WZN14, WIJ13, Web14, WTK14, WWS13b, YR13, YN14, YL14a, YL14b, ZWL^{+}12a, ZJS15, ZOL^{+}11].

T [Ham10, BSV14]. T-spline [BSV14]. T3 [WKL^{+}14]. table [BGG12]. tails [TR12]. tally [RSFS13]. tangent [DKW10, NKF14]. tangent-plane [DKW10]. tangential [BC14]. tank [OBT12, SCS14]. target [DRB14, MS12]. targeted [RUL^{+}13]. tau [FWP13, LP12a, YR13]. tau-leaping [FWP13]. Taylor [JWV12, NF12, BA10, GDW11, HAD10]. TD [CPJ11]. tearing [XJ14]. technique [BPOS12, CBKK12, DRZ14, FS10b, GHM14, HP13, HPS13, HDPF11, IH13, JWV12, KK10a, KKS13, LB13, LMQ^{+}11, LH10d, LN14, MZ10, MS10a, NKF14, PB11, PBC^{+}14, RM14, SLH^{+}11, SHA12, Sog14b, sSLxRyZ14, UKP14, WS10b, YLW10]. techniques [ABD14, AC11, CWZ^{+}11, GAW^{+}14, HSN13, Kha13, LDN^{+}10, OK13, PRH^{+}14, PBWK11, PGLK14, SS14c]. temperature [CWC11, Gen11, Gho17, GRT12, HH13, KSS14, LTL^{+}14, PBC^{+}14, ZFJ11]. temperature-dependent [Gen11, Gho17, LTL^{+}14]. Tempering [GRT12]. Temporal [Ich13, RUL^{+}13]. ten [NIN^{+}14]. ten-nanometers [NIN^{+}14]. tension [AHA10b, AST13, ABMT12, BPHK13, LFRM10, SZF12, Zha10a]. tension-dominated [LFRM10]. tensor [BGK12, Brü13, CB14, Cou11, DSZ13, EAB^{+}12, FE11, GS11a, LGG^{+}13, LS10b, OLPM14, Wen10]. tensor-product [BGK12]. tensorial [VdMS^{+}14]. tensors [EZ10, GHDS10, NCV10, XY14]. term [LFL11, MAD13, Nis12, TGO13]. terminal [TVE14]. terms [MCDT12, Mor10, MGN10b, MGN12a, MGN12b, NF12, SZ10, WSYS12, YKWS13a, YKWS13b, ZS11, ZLCW14]. Test [GHM10b, GM11a, HH12b, HH13, LTC10, NL10, PE10, Tau12, ZMD^{+}11]. test-ion [HH12b, HH13]. test-particle [Tau12]. testing [KJWR14]. tests [WFK11]. tetragonal [ZCD10]. tetrahedral [BD14b, FB11, IXX14, LXL^{+}12, LXS^{+}13, LZZR10, Sco12]. textbook [ASB10]. th [HPS13]. their [BCZ13, BP13a, CS12b, Gon12, KST13, KM14b, Ler14, SM12b, She13, SC10]. theorem [ZA14]. Theoretical [QS11b, RB10, Gra10, LB10]. Theory [CSK14, FCN^{+}13, LG14, OBNN13, OBNN14, ZK13, Bre12a, BG13, CB11, ITO^{+}10, JFC^{+}13, KS11a, KBRM12, KO14, LLYE12a, LLYE12b, LSE13, MIKG12, MNL^{+}13, RG12, SL13a, SD14b, SBO11, SP14b, ZGSZ10, Fat10]. Thermal [WP14, Atz11, BB12a, Boo11, CLZ12, DSHP11, EMK11, EC13, EMSH14, FAY^{+}13, GNGAS10, GBNS14, GMG12, HH12b, HH13, HW14a, JYHT10, KMA13, LKM13, LMK13, LGP14, ML14a, OMA14, PHSA14, PPWZ11, WL11, ZGSZ10]. thermal-solute [GMG12]. thermalization [RRCD14]. thermally [BMP14, CH11a, CLAL14]. thermo [ML14a]. thermo-elastic [ML14a]. thermoacoustic [WIJ13]. thermocapillary [LZV12, LVZK14, MB13a]. thermochemistry [CGC^{+}12]. thermodynamic [GK14]. thermodynamically [FGN14, LGE^{+}13]. thermomechanical [Han11, LG14]. thermostat [CWC11, LG14]. thermostats [Dav10, Dav15]. thick [RGK12]. Thickness [CY14]. Thickness-based [CY14]. thin [ACB^{+}13, CY14, DLC11, FGE11, GF13, LT10a, LW11, LHM14, MB11, Par12b, SP14a, WJCG13, YTYJ12]. THINC [IST^{+}12, IXX14, SX14, XIC11]. Third [VQRD11, DAEB13, DCL11, HAH14, Mat14, Nis14a, Nis14b, WYT14]. third-order [DCL11, Nis14a, WYT14]. threads [ACB^{+}13]. Three [AH13a, CNR10, CGC^{+}12, DDF14, GCFJ13, HCS12, Kau10, dCRCS10, SSY13, Ald10, AN14, ADMA13, ASK^{+}13, AKMD14, BK14a, BLM11, BR12a, BBPC14, BL13, BLR14, CWS13, CTG12, CM14, CLSX14, CCY13, CCLV10, CCG11, DB12, DKW10, FQV13, GH11a, GHS12, GG10, GMC13, GD13, HWST12, HKL14, HCH^{+}10, IDNG13, JLZ13, JGZL14, JRTY11, KSU13, KK10b, KG12a, KLPS14, LP12a, Le 14a, Le 14b, Lee13, LJX^{+}14, Li13, LRK11, MMA13, MF13, OD14, RW13, SCIE14, SYJ11, SS14a, Tak14, TDU14, TTD11, TP14, VRBZ11, WCM^{+}14, WZ11, WD13a, WD14, WLW14, WC11, Wil13, WS10c, WC14b, XHF13, XJ14, ZFH14]. Three-dimensional [AH13a, CNR10, CGC^{+}12, DDF14, GCFJ13, HCS12, Kau10, SSY13, Ald10, ASK^{+}13, AKMD14, BLM11, BL13, BLR14, CM14, CCY13, CCLV10, DB12, DKW10, FQV13, GG10, GMC13, HKL14, HCH^{+}10, IDNG13, JLZ13, JGZL14, KSU13, KK10b, KG12a, KLPS14, LP12a, Le 14b, LJX^{+}14, LRK11, OD14, RW13, SCIE14, SYJ11, SS14a, Tak14, TP14, WCM^{+}14, WZ11, WD13a, WD14, WLW14, WC11, WS10c, WC14b, XHF13, XJ14, ZFH14]. Three-phase [dCRCS10, Li13, MF13]. thresholding [HW14c]. throughout [BD14c]. TianHe [XDZ^{+}14]. TianHe-1A [XDZ^{+}14]. Time [Ala14, BDA12, CL11a, CRK12, FWP13, GvdSVK10, GH12a, KA13, LFK11, LX14, LS13d, LDDBR13, MK14a, MSTL13, MHHGN14, NZ14, RH13, SG10b, SG10a, Soa13, ACG14, Abr10, AVKP11, AB11, AS14, AME14, AF11, AWK^{+}11, BLM^{+}14, BCM12, BCB12, BN12, BDTW10, BLM11, BBAP14, BSCML13, BBF12, BSB14, BM10b, Brü13, BHY14, CP13a, CdCNH14, CHZ13b, CHZ13a, CH10, CLS10, CGMQ14, CP13b, CWZ^{+}11, CVI^{+}13, CG13a, CJN^{+}13, Cui12, DYYA14, DLC11, DWL10, DCC^{+}13, DFFL10, DS12, DL10a, DZL13, DBM^{+}12, DZHB13, Dur14, EAG14, FL11a, FGLB14, FP14, FCY13, FWLP14, GB14, GAvdVB14, GNGAS10, GBNS14, GDHM11, GVH14, GBS^{+}14, GH11b, GAV13, GP12b, GS11b, GGG^{+}13, HPA14b, HNMM10, HBM13, HL12a, HBZL14, HRBK10, Hen13, HJK14, HDD12, HK11, HW14b, Ide12, JMFO13, KK10a]. time [KKS13, KN14, Kri10, LKM13, LCNK10, LP12b, LKLG14, Leu11, sLwG10, LCK11, LHY12, LMGS13, LLP14, LT14, LJ13b, LSW14, LLH10, LBR14, MF14, MM10, MdVBS13, MS14b, MBLM10, MST14, MBA14, Mön13, MY14, MHGNM12, NPC11b, NIN^{+}14, NL13, NNS11, Nor13, OLFS13, Ols11, PL10, PGA11, QY12, QL11, QL12, RSD10, RK14, RSS10, RSS13, RGMC13, RBNS^{+}12, RBNS13, RC12, RCvdV13, RA10b, RM12, RM14, SBWF14, Sal14, SNS10, SHR13, SD14a, SK13, SLGB12, SSR14, SBvdV11, ST11a, SW13, SP11, Tal13, TH14a, TA12, TD12a, TAAY14, TEY12, VV13, VMS11, VL14, Wal13, WD13a, WW14, WG11, XJ14, YVL^{+}11, YDN10, YCH^{+}13, YN10, ZD10, ZFH14, ZMQ13, nZzSlL14, Zha14, ZMD^{+}11, AAT^{+}14]. time-accurate [GH11b, HK11, MF14]. time-convolutionary [ZD10]. Time-dependent [GvdSVK10, LFK11, SG10b, AME14, BN12, Brü13, CdCNH14, CHZ13b, CHZ13a, Dur14, FGLB14, GS11b, HPA14b, LSW14, MM10, OLFS13, SD14a, SLGB12, SSR14, SW13, WD13a]. time-differenced [LKM13]. Time-domain [NZ14, Soa13, AWK^{+}11, CWZ^{+}11, CVI^{+}13, DL10a, GB14, GVH14, GP12b, HBZL14, HJK14, HDD12, LHY12, SBWF14, TD12a]. time-fractional [BHY14, ZFH14]. time-harmonic [DLC11, EAG14, LLP14, Mön13, NPC11b, PL10, RM12, TEY12, XJ14, ZMD^{+}11]. Time-independent [MSTL13]. time-invariant [LJ13b]. time-lapse [BLM11]. time-marching [sLwG10]. time-parallelized [RBNS^{+}12, RBNS13]. time-propagation [AF11]. time-reversal [VMS11]. Time-space [LS13d, HL12a, TH14a]. Time-splitting [LX14, PGA11, Sal14]. Time-step [LDDBR13, CJN^{+}13, DWL10, LLH10]. Time-stepping [KA13, AB11, DS12, HRBK10, Kri10, LMGS13, SK13, ZMQ13, AAT^{+}14]. time-varying [CLS10, YCH^{+}13]. timestepping [LGP10]. tip [GMXG14]. tire [VSC^{+}11]. TLM [TRSMM^{+}13]. TOKAM [TGT^{+}10]. TOKAM- [TGT^{+}10]. tokamak [AAG14, HdBK12, PBC^{+}14, BBF12, ICC^{+}10]. tokamaks [LL10b, MMIW14, TGT^{+}10]. tomography [BFR13, GJ14a, HQL^{+}10, LHH14, LDN^{+}10]. tool [YL14b]. tools [BFR13, PMM^{+}11]. topography [FMT11a, Ion13, JTT14, LT11a, MGN13, YC11]. Topological [Par12b, CNR11, NW10]. topologies [LS13c]. Topology [DLZ^{+}11, YYY^{+}14, CZ11b, DLLW14, TNK10, ZLL10]. toroidal [KHJ^{+}13]. toroidally [ZKV10]. Torrey [NLGL14]. torus [CJM13]. Total [KHHK12, LLYE12a, BDM13, KBS14, OS14, VMA14, YL14b, dCRCS10]. Total-variation-diminishing [KHHK12]. TR [EMK11]. TR/BDF2 [EMK11]. tracer [EN14a, EN14b, HLM11, LNU10, WD11]. tracers [GM11a]. track [KS13b]. tracking [APF13, AK14, Ald10, BW13, BR12a, FL14, Her10, ID10, KT14, LTC10, LW11, MS12, NW10, Qua11, RZ10, RR14, Sam14, SN13a, STD^{+}10, UGkM13, VF10, WWZ13, Wit10]. tracking-level [NW10]. tracking/ghost [TVEC12]. tracking/Lagrangian [Her10]. traction [PGA11]. tractions [KS11c, PSL20]. traffic [WSW13]. trains [IBO14]. trajectories [BBC^{+}13, LTC13, LRS10, LSK11, Mar10a]. Trajectory [HCP10, NVS13, CSC14]. transducers [HVD13]. transfer [BRS13, Boo11, CGG12, DTU12, ETW10, GL12a, GGH14, GS11a, HS13a, HMOH11, IL11, IF10, JYHT10, KKvZB14, LN10b, MB13a, MH10, NTV14, OHF12, PBdGP14, RCCC14, SRS14, SD14c, WL11, ZYT13, ZTL13, WP14]. transform [HW11, JHF10, LQ10a, RS13b, RS14b, Boy10]. Transformation [LBM14, OK13, GC11, GCZ13, JW11, KO14, MTZ14, PIN14, RG12, ZM10]. transformation-free [GC11, GCZ13]. transformations [MMD^{+}11, OK10, ZCD10]. transformed [FCY13, PSF^{+}14, Nic11]. transforms [QY10, STD14]. Transient [LCK11, MJK13, AJRT11, BLM^{+}14, BCL13, BD12, LP11, MSTL13, MGN10a, NHE14, NJ12, SK12b]. transition [JW11, KGK13, ZHW10]. transitional [TAD14]. transitions [CL13]. translocation [WMK13]. transmembrane [ALGB^{+}13]. transmission [JS13b, MTT13, PL10, PRL10, Stu10, WK14]. transonic [BBD11, SSG11]. transparent [Du11b, HPA14a, MJ14]. Transport [KS13b, MM11, AHA13, AAG14, AS12, BDL11, BLRR11, BGGW13, BRP14, Boe11, CTP13, CVC12, CdCNH14, CRT14, CKD10, CM10, CBTR13, DTU12, DR14, EN14a, EN14b, FL10, FLB^{+}12, Fuj13, GDY10, GMC13, Gio10, GPS10, GBB^{+}13, GLC^{+}11, GP12a, HLM11, HRT11, HH12b, HH13, JC10, JLC14, KFOS14, KBGRM10, KHGW13, LBC14, LKM13, LKP14, LNU10, LEM11, LP13a, LBM10, LBS10, MKCB14, MPS^{+}14, MMR12, MTS^{+}11, MB10b, MV13, MJM13, MMJ14, MGN11, NSA11, NL10, NGV14, Nor14, Oli14, Ols11, Ols12, OD14, PMS14, PB14a, Pri14, RARO13, RGK12, RF14b, RCCC14, RSFS13, RJG10, SGS13, SH14, SO13, SBB10, STD^{+}10, SAA14, SWR11, Tal13, TGT^{+}10, TLG10, Tok10, TR14, VTTK13, VXB14, VHWF13, VWP11, WPP10, WR11, WT13, WD11, WAM14, WPK14, WTK14, XX12, ZSW10, Zer10]. transport [ZCW11, ZKM^{+}11, vDS12]. Transport-constrained [KS13b]. transport-diffusion [DTU12, RCCC14, SAA14]. transport-velocity [AHA13]. Transportation [BFO14]. transversely [BCL14]. trapezoidal [Ide12]. traveltimes [LQ11]. tread [VSC^{+}11]. treat [SS10a]. treating [JWNL11, MKB^{+}13, RWMG11, WS10b]. treatment [BB12a, GF13, Joh11, KDS14, Kha13, KK10c, LH14b, Man12, MBD13, SVK14, SN12, SJH12, TS11, TWSN12, TR12]. treatments [JC10, OS14]. tree [CG12, LL14a]. tree-like [LL14a]. treecode [GK13]. treecode-accelerated [GK13]. trees [ARR11, ARR13, TE14]. Trefftz [Tsu10, VBVD10]. Trefftz-based [VBVD10]. triangle [KBRM12, SDH12]. triangles [WCVJ13]. triangular [BMS12, BBD14, CF11c, CH11b, FE11, Gas11, IXX14, LXM14, SV10, TD11b, WR11, ZZS13]. triangulated [BLJ11]. triangulation [Gen13]. triangulations [MOSW12]. tricubic [Mar10a]. tridiagonal [HPLS10]. trigonometric [Boy13, CJM13]. triple [Li13]. troubled [VR14]. troubled-cell [VR14]. truncate [ACG14]. truncated [DKC14]. Truncation [SS14a, FdVV12, FVR13, SEBG12]. Trust [WT13, MW14b]. Trust-region [WT13]. tsunami [BSCML13]. Tube [KYP^{+}14, LM14b, RCQL11]. tumor [CCFCM11]. tunneling [CCMG^{+}13, JCT11, MJK13]. turbomachinery [GBS^{+}14, WDP^{+}14]. turbulence [CP13a, CDB13, ECDB14, Ger12, GLB^{+}11, GHM12, GLM14, GM14b, HHL11, JLB^{+}10, KSP13, KGK13, MG14a, MS14b, OBT11, OTV13, dlLPC10, dlLPCP12, SNS10, SRL^{+}11, TGT^{+}10, Tau12, WWX^{+}10, WMYYG10, YB14]. Turbulent [BPQO13, BHB11, BVM14, BH13a, BGGW13, BA10, BGM12, BM13b, BH13c, CFCA13a, CFCA13b, FL14, Fus13, GG10, GHM10a, GM11a, GW10, HHH13, JK10, KH13, KSL10, LBC14, LTC13, LDS11, LSBJ14, MJ13a, Nag14a, Nag14b, OAK11, PP14, RG13, RBNS^{+}12, RBNS13, RS10b, SBB13, SFCF10, SRS14, TRL14, TAD14, Tyl14, VXB14, VWP11, YYB12, vDS12]. Turing [GMT10]. Turkel [AGBC14]. TVD [BR14, QL12]. twist [OLC13]. Two [BBG^{+}11, DTG^{+}14, Du11b, GK14, HY10, JGZL14, PZ10, RDHK12, WABI11, YT12, ZWL^{+}12b, ABL^{+}11, Ala14, AK12, ABMT12, ABMT14, Bal12b, BD14a, BG10, BW13, BILM14, BYCC14, BCM^{+}11, BRS10, BHY14, BD14c, BTY14, CBKK12, CTG12, CRS14, CSW14, CS10, CS14b, CZ11b, CL11b, CTSM13, Cui12, DYYA14, DF14, DS12, Don14b, DGF14, ES10, FGR12, GMB10, GT12, GJLY13, GL12b, GHL14, GVV13, GWG14, GIQ11, HZ11a, HR10, HLLW11, IJZ13, JCC14, JZ11, KCG14, KS11a, KT14, KLP10, KK10c, KL11b, LP12a, LP13b, LBT11, Lee14a, Li10a, LX13a, LXL13, LW13a, LRN13, LW13b, LP11, LHPH11, MB13a, MF14, MB12, MAB^{+}13, MFM12, MB11, MNF^{+}10, Men12, MTZ14, MW14a, MY13, ML14b, MJM13, MMJ14, MGN10a, MGN11, MGN12a, NKF14, PB11, PZTW12, PE10, PBM11, PS14a, PS15, QB14, RBB12, RVB10, RE13]. two [RB13, RNT12, Sam14, SF11, SGS13, Sha13, ST13, SYJ11, SX14, SHA11, SHA12, SBvdV11, SKt10, SMS14, SS14c, TSLV11, Tok11a, TT10, TL13, UGkM13, VP10, VP14, VMA14, VBCM10, VSC^{+}11, VW14b, Wan13a, WW11, WZ11, WSS12, WZSN12, WGX13, WQCS13, WC14a, WMYYG10, Wol12, WS14, WS11, WT14, XYL12, XR14, YS11b, YZX14, Yin14, ZHW10, nZzS11, ZYHW13, ZHY13, ZKZ10, ZHS10, ZWL10]. two- [SYJ11, WZ11]. two-center [LP12a, Men12]. two-component [MTZ14, ZKZ10]. two-density [ZWL10]. Two-dimensional [YT12, ZWL^{+}12b, Bal12b, BD14a, BG10, BILM14, BYCC14, BCM^{+}11, BRS10, BHY14, BD14c, BTY14, CBKK12, CRS14, CS10, CS14b, CTSM13, Cui12, Du11b, GMB10, GIQ11, KS11a, KLP10, KL11b, LX13a, LXL13, LW13a, LRN13, LW13b, LP11, MB12, MAB^{+}13, MNF^{+}10, MGN10a, MGN11, PB11, PE10, QB14, RVB10, VMA14, Wan13a, WW11, WZSN12, WS14, WS11, WT14, YZX14, Yin14, nZzS11]. two-dimensions [JZ11]. two-equation [WMYYG10]. two-fluid [CZ11b, DF14, DGF14, GVV13, KK10c, NKF14, SBvdV11, SKt10, YS11b]. two-layer [LBT11]. two-layered [Li10a]. Two-level [WABI11, HZ11a, JCC14, Sha13, SS14c]. two-parameter [ST13, WC14a]. Two-phase [GK14, PZ10, RDHK12, ABL^{+}11, Ala14, ABMT12, ABMT14, BW13, CSW14, CL11b, DS12, Don14b, FGR12, GT12, GJLY13, HLLW11, KCG14, KT14, LP13b, LHPH11, MB13a, MF14, MW14a, PZTW12, PBM11, PS14a, PS15, RBB12, RB13, SF11, SGS13, SX14, SHA11, SHA12, TSLV11, TT10, TL13, VP10, VP14, VBCM10, VSC^{+}11, WSS12, XYL12, XR14, ZHW10, ZYHW13, ZHY13]. two-point [ML14b]. two-scalar [MB13a]. two-scale [RE13]. two-sided [MY13]. two-stage [IJZ13]. Two-step [BBG^{+}11]. two-steps [Lee14a]. Two-way [DTG^{+}14, Sam14]. type [AMDR10, AM11, CHHL11, Del10, EEK14, FY13, FP12, HLT10b, KDS14, KO14, LT11a, LTCN13, LT14, LM13, LHV13, PGR10, PHRG13, RT14, SV13, TT10, Tok11b, WLW13a, WW14, WI13a, ZzS11, ZQ11, ZZSQ13]. type-II [SV13].

ultra [AK14, DL13a, DL13b, LHM14, SBS^{+}13]. ultra-relativistic [AK14, SBS^{+}13]. ultrarelativistic [AC11]. ultrascale [GK10]. unbounded [AJRT11, Boy11, BHY14, CK10, CGS13, DKC14, DRZ14, Era13, hGzSnZ12, hGzS13, HLS14, HRCW13, JPC^{+}14, JKR14, SY11b, Zen11]. uncertain [BND^{+}12, TLNE10]. Uncertainty [LTT10, SP13, WT12b, AC13, AA13, BZ12, BZKL13, CPX13, CNST13, CZ13, EH14, EH15, ERS14, JEX10, JNX13, KL14a, LBR14, MPPP12b, MPPP12a, PGLK14, WI13a, XS15]. unconditional [BRVE12]. Unconditionally [MY14, SHR13, BCP13, BL10, DS10, GH11b, LGF11, LB10, MAD13, TD12a, ZFH14]. undulatory [YS11a, YS11b]. unequal [Ge10]. unidimensional [OK12]. Unification [Nis10]. Unified [LVFK14, LYXZ14, SMP^{+}13, BBGP13, CXLC12, GHS13, Mie13, RTKS10, XH10]. uniform [AM12, AW14, ABJ^{+}12, Boy11, CD12a, CES14, CLSX14, ES14, GGG^{+}13, LH10a, LT11c, LFW11, MBLM10, MBGW13, SBS11, WLL13, nZzSlL14]. uniformly [Luo13]. uniqueness [SP11]. unit [Hel11a]. Units [SSK^{+}12, GHH14, KHRD11, LSK11, LLL10, HJBB14]. units-Radial [LSK11]. unity [OJW10]. Universal [BRS10]. Unlike [KWH10]. Unlike-particle [KWH10]. unorganized [DS11]. unresolved [GLM14]. unsplit [LP13a, Lee13, MT10, OJ11, OD14]. unstable [GCFJ13]. unstaggered [HRT11]. unsteady [BDA12, BGHM13, BSCML13, DLZ^{+}11, EH13, GCL13, IH13, KKvZB14, LDS11, LGC13, LLP10, LvZB10, MHHGN14, MGN11, MLGN12, MGN12b, PC13, Sen13, STH11, SFWP10, SM10, XW14, YDN10, YP13]. Unstructured [CVC12, ALR11, ABD14, AIX10, AGSG11, BDA14, BNM14, BYCC14, BV14, BD14b, BBD14, BC10b, CCKCG10, CZ11a, CRT10, CBTR13, DCL11, DLM13c, EZ10, Era13, FB11, GMB10, GCA13, GBB^{+}13, GP12a, HLT10b, IXX14, IKO^{+}14, JLZ13, KS11b, KDS14, KAFB11, KD10, Lan14, LRLL11, LR12, LZ13, LSV10, LM14a, MF12, MB12, MAB^{+}13, MIJ10, MGN14a, PB11, PYK10, PVLT10, Per13, RS14a, SGS13, SCGE13, SYY14a, SYY14b, SSHM10, SR13c, SSW13, Son11, SS10e, TF10, TLO^{+}14, TTD11, TAD14, VMA14, WR11, WAM14, XIIX14, ZN14a, ZN14b, ZQ13, ZZSQ13]. unstructured-mesh [SSW13]. unstructured-meshed [BYCC14]. untangling [TGRL13]. Untilted [SG12b]. update [RLPM12]. updates [MMH12]. upon [HKL10]. upscaling [DD12]. Upwind [BH12, RF14a, SST10, BF12, BSS14, HY10, HWA10, JTXZ12, MHGNM12, Mur10, PBWK11, Rie10, SHCY13, VQVD11]. Upwind-biased [SST10]. upwind-schemes [Rie10]. upwinding [TLNE10]. UQ [NV14]. Use [Ide12, CHC11, DMP14, GD10, HLW13, KM11, KGKS13, MD13, Mar10a, dlLPCP12, ZPGO10]. used [NB13]. uselessness [Boy10]. Using [Gon10, LA10, SJ10, ABD14, AH13a, AHA10b, AJ14, AA13, ADT14, And13, AETT11, ABK12, ABK13, AEM13, BGGGS11, BA10, BFO14, BN12, BS14b, BFE12, BSB14, BX12, BHL12b, BGG12, Cam13, CGM^{+}12, Cha13, CGG12, CP14, CCC12, CLG13, CSK13, CF11c, CH11a, CCLV10, DL12, DCC^{+}13, DjY14, DAEB13, DCVM11, EGW11, ES10, EQYF13, ELM14, ERT10, FT11, FMT11b, GMB10, hGzSnZ12, GW11, Gno14, GBB^{+}13, GLC^{+}11, GHH14, GWG14, GIQ11, GMXG14, HdCNT12, HBZL14, HMHO13, HZS12, Hut14, iIT14, IX10, JWV12, JMFO13, JYHT10, JLZ13, JLS^{+}14, Joh11, JMGN13, KVM12, KM13, KJ11, Kha13, KSPP10, KSP13, KJC13, KMHJ10, KHGW13, LRL10, LCNK10, Lee12b, LCK11, LC13, LZC^{+}13, Liu11a, LZV12, LHPH11, LBR14, LCB12, MB13a, MBR11, Mac10, MOV12, MTT13, MOSW12, MNF^{+}10, MTG11]. using [MTS^{+}11, MZB12, MDM^{+}12, MLBS14, MY13, NTV14, NS14a, NMT^{+}12, NB14, Nor13, Ols12, OLC13, OLPM14, PDSS10, PB11, PMM^{+}11, Phi14, PBdGP14, PBC11b, PM14, Qua11, QA11, RBB12, RCW11, RHB12, RA10a, RR14, RS13a, RDHK12, RBCK10, RS10b, RG12, Sal14, SNS10, SAM10, SY14a, SSS10, SCGE13, SZ10, SYL12, ST13, SHM12, SCH13b, SFWP10, SC12, SPM^{+}13, SRN13, SS14c, SDH12, TK12, TK15, TE14, TBMH11, TS12, VHWF13, VT13, Vol14, WK10, WWS^{+}13a, WTYC13, XFB^{+}14, YYY^{+}14, YM14, YM10, YXD^{+}14, ZD10, ZYHW13, ZY13, ZQ11, ZZSQ13, vydCNS^{+}10]. utilizing [SAHS14, UJ12].

V [NOT14, NOT14]. V-V-P [NOT14]. Väisälä [HdBK12]. Validation [BSDM12, GM10a, NB13, OAKR14, YK12]. validations [DEKBF14, LXSR12]. validity [JRG13]. valuations [CLW10]. value [EO11, HdCNT12, HLW13, LRK11, LWS14, LN14, PR11, SN14, hT10b, WLW13a, XIIX14]. valued [PSDF13]. values [FS10a]. valve [LS13a]. vanishing [KSB12]. vapor [CP12]. Var [CS14c]. variability [ECDB14]. Variable [MTS^{+}11, AF13, BMK12, CTP13, CGC13, GW14, GW10, GLL14, HK11, ISZ12, JNM12, KAFB11, LRT13, LMGS13, Mar13, Mor10, MLGN12, PIN14, SGS13, SHM12, TGGT13, dv23]. variable-density [AF13, GW10, SHM12]. variable-order [KAFB11]. variable-stepsize [CTP13]. variables [Abr10, LGE^{+}13, SKN13a, ZJT10]. variably [AST12, HSK^{+}15, Zad11]. Variance [Sza12]. Variance-reduced [Sza12]. variant [TAAY14]. variants [EMK11]. variation [FN10, KHHK12, YL14b, YB14]. Variational [ES13b, GAvdVB14, LHK12, MKGV12, ML14a, OBTC^{+}13, ZOWZ10, ZOJR11, ZWL10, ABL^{+}11, ACCM12, BGGW13, BA10, BSS12, BSB14, GLS10, GCLS11, GH11b, GW10, HLSO13, JZ10, Jin12, JK10, Kaz10, LHM14, MMV^{+}13a, MMV^{+}13b, NW10, RG13, Sco12, XMDG13, XMDG14, YMS10, YMS12, ZHS10, ZWJW10]. variations [PBC^{+}14]. Various [ZPF10, XL11, ZRM13]. varying [CLS10, MPT13, RCCW10, SR13a, TGO13, YCH^{+}13]. vascular [GDY10, IWG13]. Vector [CGM^{+}12, Cha13, CG13a, KSK13, LLY11, NS11b, PB14a, PSDF13, ZS14]. vector-valued [PSDF13]. vectorization [DKH11]. vectors [ELM14, NCV10, Ran11]. vehicles [CRS14]. velocimetry [HMM^{+}10]. velocities [Min13, NCKN10]. Velocity [OR10, AHA13, AJ14, BCHM14, BDMP13, BNM14, BD11b, BM14, CIJS14, CWL10, CXLC12, DS10, FR13, GG14, HCS^{+}10, JW12, LE13, Liu11b, McD14b, MKB^{+}12, MVG11, NS11a, SCIE14, SV10, VLM13]. velocity-correction [DS10]. velocity-pressure [CIJS14]. velocity-space [LE13]. Velocity-vorticity-helicity [OR10]. ventilation [YCH^{+}13]. ventricle [LS13a]. ventricular [MNKS13]. Verification [SHM12, BHL12b, GM10a, IH13]. versatile [CTG12]. version [ABK13, GLB^{+}11, HLM11, KKS13]. versions [MGN12a]. versus [CM10, GBNS14, MJ14]. vertex [AGBC14, CCKCG10]. vertex-based [CCKCG10]. Vertically [WLW13b]. very [ALR11, SC12]. vesicle [BLJ11, FBM14, GWG14, QB14, VRBZ11]. vesicle-substrate [GWG14]. vesicles [AELV14, HKL14, KL10a, RVB10, SM11a, STL^{+}10]. vessels [MPT13]. VFRoe [PBM11]. via [CT13, ECD14, FS10a, FLM12, GW13a, GD13, HLDP14, JZ10, KL14a, LFL11, Lee14a, LX10, LCP13, LMZ14, LTT10, MF12, Mar13, MTZ14, Min10b, MVG11, OK10, Ran12a, RSS10, RBNS13, RSFS13, VS11, WdHX10, YR13, YWL14, YJK12]. VIALS [YL14b]. vicinity [CM13]. violent [BCD14, SM12c, ZMD14]. virtual [AT13b, BvBZ^{+}10, HWST12, SSHT14, SJ10, ZWH^{+}12]. visco [AH13a]. visco-acoustic [AH13a]. viscoelastic [LP11, MMT14, ZYF^{+}10]. viscoelasticity [DP12]. viscoplastic [FNGV14, Ion13]. viscosities [XH13]. viscosity [BD10a, CD12a, CLG13, CSB14, Coo13, FMT11b, GW14, GPP11, KSB12, KL12c, LRT13, LS10b, RSS13, SGS13, TGO13, VW14a, Wen10, dv23]. Viscous [CCLS10, AiINT14, AELV14, ACB^{+}13, BJLR14, BDKV14, CP14, CF11b, CB10, CC14, CGF^{+}13, FSL10, IG14, LFL11, LTS^{+}10, LKT^{+}12, LMZ14, LLP10, MCA^{+}13, MESV10, Min10b, MN14, PRA13, PMSH10, RHGT10, SGC13, SHCY13, SYL10, SZ10, SS10c, STL^{+}10, SM10, SR13d, TGO13, UWB12, WS11, XW14, YS11a, YS11b, YFL^{+}13, YSW14, ZL10, ZGSZ10, ZDSP11]. viscous-plastic [LTS^{+}10, LKT^{+}12]. Vladimir [vL11]. Vlasov [ARR13, CCZ14a, CCZ14b, CGMQ14, CMS10b, CLM13, DDN^{+}10, DMR14, GCH14, GQ13, HGMM12, HB13, MRS14a, MMA11, MMA13, PSF^{+}14, QC10, QS11b, RS11, SS10d, TTR^{+}12, TBNT11, VCS14, Wol12, XQXC14, YF14]. VLPL [TP14]. VOF [OD14, CMS10a, DvW14, GMB10, KBTD12, MZ11c, VdMS^{+}14, XIC11]. VOF-IBM-enthalpy [KBTD12]. voids [SLC^{+}11]. volcanic [MPS^{+}14]. Volterra [HHGC14, ZN10]. Volume [AGBC14, KLL13, SGW10, SC10, AIX10, ABS^{+}14, BFSVC14, BD14b, BBD14, BHL12b, BDKV14, CCKCG10, CVC12, CHKT13, CRG13, CLSX14, CDL11, CDHM11, CC14, CMG^{+}13, CJ14, DM13, DCL11, DZHB13, DKM11, FGR12, FS10a, FE11, FMM^{+}10, GDHM11, GDGP11, GXS14, HJ11, HKJ11, HRBK10, HMHO13, HM10, Hu13, HAQ12, IX10, IDNG13, JMW12, JTXZ12, KCG14, KKAS12, KDS14, LP13a, LR12, LRBT14, LSV10, LHPH11, LGH10b, LTL11, LZZR10, MAPP13, MF14, MB12, Mig14, MSS12, ML13, MMT14, MLBS14, Nis14a, Nis14b, NNS11, Nor14, OD14, PDSS10, PRH^{+}14, PGL11, PvDtTB^{+}11, PNPF10, PM14, PP14, QB14, RSK13, RBB12, RDHK12, SSB13, SY11a, SY12b, SS10c, SG14, SY11c, SR13d, SIDG13, UJvL10, UJ11, UJ12, VDP11, VWPF11, VWF13, WSS12, WYS12a, WYS12b, WZZZ14, WZ10, WT14, WG14, XIIX14]. volume [XLD^{+}11, YVL^{+}11, ZHY13, ZJL^{+}14, Den11a, GP12a, TL13, WY10]. Volume-average [XIIX14]. Volume-average/point-value [XIIX14]. Volume-based [Den11a]. Volume-of-fluid [SC10, HMHO13, LP13a, LHPH11, OD14, RBB12, RDHK12, WYS12a, WYS12b, ZJL^{+}14, WY10]. volume/element [BFSVC14]. volumetric [HW10]. Voronoi [Son11, SS12]. Voronoi-cell [Son11]. Vortex [RS10b, Alb10, AEM13, CF10, DH10, GD13, KWSJ11, RBCK10, UWB12, vRLPK11, RCW11]. vortex-dominated [KWSJ11]. Vortex-In-Cell [RCW11]. vortical [CK10, vRLPK11]. vortices [CSKP10, LFW11, MAPP13]. Vorticity [TS12, DF10b, KJ11, OR10, RTKS10]. Vorticity-divergence [TS12]. voxelized [RS13a]. Vries [YHL13]. vs [GKG^{+}13, MMJ14, SMP^{+}13, ZL10].

wake [GBS^{+}14]. wakefield [VGCMG11, YXD^{+}14]. walk [PBB14, STD^{+}10]. Wall [GDY10, MS14a, RS13a, AHA12, BN11, KH13, KIF^{+}14, LSBJ14, LVFK14, NEE12, Vre14, ZPGO10]. wall-bounded [KH13, LSBJ14]. walls [KM11]. Walsh [Gno14]. WAMR [PZW14, PZG14]. Water [JPC^{+}14, JKR14, RB10, AIX10, ABS^{+}14, BNT14, BDT10, BCD14, BF12, CKS14, CVC12, CRG13, CLSX14, CT14, CB13, DLGP13, DMMGM14, DKA12, DLM13b, DLM13c, FNKdLB13, FMT11a, FLB^{+}12, GAvdVB14, Gas11, GCFJ13, IX10, JTT14, KJ11, KIF^{+}14, LT11a, Le 12, LBT11, LGLX14, MHGNM12, MGN10b, MGN12b, MGN13, Nag14a, Nag14b, OBT12, PDZ^{+}14, PE10, Pud11, RBS12, RVFK14, RW13, SHR13, STC10a, SN13b, TS12, TBR13, UJvL10, WPKK12, Xin14, YC11, ZOJR11, ZMD14, Zho11, vTCL^{+}10, vV12, SP14a]. Waterman [KSPP10]. Wave [ABH^{+}14, BCL14, FJ12, MGN12b, AJRT11, Ain14, AMP14, AETT11, BK14a, BH12, BLQ14, BGH10, BRS11, BHK^{+}10, BHY14, BLR14, CTG12, CL11a, CH10, CSX14, CH11b, DOR10, DB12, DRZ14, DEKBF14, Dur14, DV14, DKM11, EL11, Fat10, FH11, FMM^{+}10, GHJ14, Gri13, HH12a, HVD13, HBZL14, HDHZ13, IBO14, JPC^{+}14, JKR14, JFC^{+}13, JXD14, KKS13, KS11a, KDS14, KEGM10, LH10a, LP12a, Le 14b, LXL13, LH14a, LL14a, LS13d, LT11c, MBL^{+}14, Mön13, NDV^{+}11, NCV10, OBT12, PRL10, PMMB14, PLN11, RKGM14, RF14a, RS13b, RS14b, SBC^{+}14, Sam14, SRSV11, SM11b, SWX12, SCH13b, Soa13, STC10a, SCS14, SC12, Tak14, TK14, TH14a, TD12b, TGGT13, VBSK10, VED13, WT12b, WV14, WZZ14, WLW13b, WSW13, WSBG10, WK14, WYC13, YL11, YK12, ZOJR11, ZMD14, Zho14b, ZTS^{+}10, ZMD^{+}11]. wave-body [YL11]. wave-structure [DEKBF14, WYC13]. wave-wave [YL11]. Waveguide [LL12b, LL14b]. waveguides [MBMV13]. Wavelet [DV13b, DS11, EL11, JSX10, KSK13, vydCNS^{+}10, GM14a, HRBK10, JU10, JHF10, LX14, PZW14, PZG14, dlLPC10, RVFK14, RWX11, ZTS^{+}10]. wavelet-adapted [HRBK10]. wavelet-adaptive [LX14]. Wavelet-based [DV13b, vydCNS^{+}10, dlLPC10]. wavelets [dlLPCP12, YXZ13]. wavenumber [SML13]. wavenumber-preserving [SML13]. wavepacket [QY10]. waves [AH13b, BDT10, BCL13, BX12, CT10, CGJ11, CVI^{+}13, FGE11, FT11, GAvdVB14, GKS^{+}11, JLB^{+}10, Li12, LGLX14, LXSR12, LP11, LM14b, ML14a, MBMV13, MTT13, NVS13, NvdVB13, PDZ^{+}14, PS14a, PS15, RW13, SG10b, SG10a, SM12c, SMS14, SC12]. way [Ain14, Alm14, DTG^{+}14, PRL10, Sam14]. Weak [LLMY14, MWW^{+}13, MGN10b, NEE12, UGF14, WPP10, ACCM12, HSWZ13, LHM14, MWY14]. weakly [BCL^{+}11, BG12, HMHO13, KA12, PMS11, PSM11, PSM13, PMS14, RS10b, VL14, XMS12, XS13, YXZ13, ZN10]. weakly-compressible [XMS12, XS13]. weakly-ionized [PMS11, PSM11, PSM13, PMS14]. weather [BCW13, CS12c, DMP14, Fal13, Lap12, OMCO14, TvdHS^{+}12]. Webb [ZJS15]. Weighted [HMM^{+}10, NMT^{+}12, CCD11b, CRT14, DB13, Fan14, FSTY14, FCYF11, HKLY13, HWA10, LQ10b, MPWK13, PWK11, PHD14, SYL12, XW10, YF14, ZKZ10, ZS13, DRW11]. weighted-integral [XW10]. weighting [AW14]. weights [AH13b, TE14]. Welding [SPS14]. Well [DLM13b, FMT11a, KM14a, MPT13, BF12, DLGP13, DLM13c, FMM^{+}10, LGLX14, RNT12, WYS^{+}11, WZZ14, Xin14, DMMGM14]. Well-balanced [DLM13b, FMT11a, KM14a, MPT13, BF12, DLGP13, DLM13c, FMM^{+}10, LGLX14, WYS^{+}11, Xin14, DMMGM14]. well-conditioned [WZZ14]. well-posedness [RNT12]. Weller [JM10]. wells [OD12]. Wendroff [FY13, TS10, TWSN12]. WENO [ADB14, AH13b, BD14b, BBD14, BDMV11, CCD11b, CTJT13, CC14, DZHB13, Fan14, GXS14, HHS^{+}13, HPOM10, Hu13, HAQ12, HAH14, JZ13, Joh11, KHGW13, KGK13, LXL^{+}12, LXS^{+}13, Nor14, QC10, QS11a, SABH11, SHZ13, SQC11, SZH12, TTD11, TAD14, XQX13, XQXC14, ZS12, ZT13b, ZQ11, ZQ13, ZZSQ13]. WENO- [Fan14]. WENO-enhanced [KGK13]. WENO-limited [Nor14]. wet [DLM13b, MHHGN14]. wet/dry [DLM13b, MHHGN14]. wetting [BPHK13, CL13, SS13b, SS14b]. WFA [DRW11]. which [ZA14]. whistler [GKS^{+}11, WS14]. whistler-mode [GKS^{+}11]. White [GDW11]. widely [RCCW10]. Wigner [JC10, JCT11, JLC14, RSK13, SD14a, SD14b]. windkessel [IWG13]. windows [LCNK10]. wings [NL12a]. wireline [PMM^{+}11]. wise [AOCR12, BLR13a, OAKR14, XLD^{+}11]. within [KH13, LB14, RSK13, RG13, SBC^{+}14, XHF13, Yok14]. without [AV13, AM10, Del14, HSWZ13, MW12, Min13, PM14]. wood [CTP13, BD14c]. Woodbury [CP14]. work [ZGSZ10]. worldwide [JWNL11]. written [VMA14].

XFEM [KBSV14]. XFEM-AC [KBSV14]. XTOR [LL10b].

Yamakawa [LGG^{+}13]. Yee [BWC11, BWC13, KGKS13]. Yee-lattices [KGKS13]. Yosida [HHK13]. Yosida-based [HHK13].

Z [CCD11b]. Zadeh [HSK^{+}15]. Zakharov [XXS10]. Zernike [BY11, LB14]. zero [BCG^{+}13]. Zipper [WQCS13]. zone [RG12].

References

Alwan:2013:ISM

[AA13] Aravind Alwan and N. R. Aluru. Improved statistical models for limited datasets in uncertainty quantification using stochastic collocation. Journal of Computational Physics, 255(??):521–539, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

              tronic). URL http://www.sciencedirect.com/science/
              article/pii/S0021999113005627.

Angot:2014:OPM

[AAG14] Philippe Angot, Thomas Auphan, and Olivier Guès. An optimal penalty method for a hyperbolic system modeling the edge plasma transport in a tokamak. Journal of Computational Physics, 261(??):1–22, March 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

              tronic). URL http://www.sciencedirect.com/science/
              article/pii/S0021999113008450.

Adelmann:2010:FPP

[AAI10] A. Adelmann, P. Arbenz, and Y. Ineichen. A fast parallel Poisson solver on irregular domains applied to beam dynamics simulations. Journal of Computational Physics, 229(12):4554–4566, June 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001026.

Angulo:2014:CPL

[AAT^{+}14] L. D. Angulo, J. Alvarez, F. L. Teixeira, M. F. Pantoja, and S. G. Garcia. Causal-Path Local Time-Stepping in the discontinuous Galerkin method for Maxwell’s equations. Journal of Computational Physics, 256(??):678–695, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006153.

Albin:2011:SFS

[AB11] Nathan Albin and Oscar P. Bruno. A spectral FC solver for the compressible Navier–Stokes equations in general domains I: Explicit time-stepping. Journal of Computational Physics, 230(16):6248–6270, July 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002695.

Abdulle:2012:RBF

[AB12] Assyr Abdulle and Yun Bai. Reduced basis finite element heterogeneous multiscale method for high-order discretizations of elliptic homogenization problems. Journal of Computational Physics, 231(21):7014–7036, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001209.

Abdulle:2013:SMM

[AB13] Assyr Abdulle and Adrian Blumenthal. Stabilized multilevel Monte Carlo method for stiff stochastic differential equations. Journal of Computational Physics, 251(??):445–460, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004051.

Abgrall:2014:IBM

[ABD14] R. Abgrall, H. Beaugendre, and C. Dobrzynski. An immersed boundary method using unstructured anisotropic mesh adaptation combined with level-sets and penalization techniques. Journal of Computational Physics, 257(??):83–101, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005962.

Atak:2014:CBE

[ABH^{+}14] Onur Atak, Bart Bergen, Daan Huybrechs, Bert Pluymers, and Wim Desmet. Coupling of Boundary Element and Wave Based Methods for the efficient solution of complex multiple scattering problems. Journal of Computational Physics, 258(??):165–184, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007109.

Appelo:2012:NMS

[ABHS12] Daniel Appelö, Jeffrey W. Banks, William D. Henshaw, and Donald W. Schwendeman. Numerical methods for solid mechanics on overlapping grids: Linear elasticity. Journal of Computational Physics, 231(18):6012–6050, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001805.

Austin:2012:SAS

[ABJ^{+}12] Travis M. Austin, Marian Brezina, Ben Jamroz, Chetan Jhurani, Thomas A. Manteuffel, and John Ruge. Semiautomatic sparse preconditioners for high-order finite element methods on non-uniform meshes. Journal of Computational Physics, 231(14):4694–4708, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001556.

Arthurs:2012:ESC

[ABK12] Christopher J. Arthurs, Martin J. Bishop, and David Kay. Efficient simulation of cardiac electrical propagation using high order finite elements. Journal of Computational Physics, 231(10):3946–3962, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200068X.

Arthurs:2013:ESC

[ABK13] Christopher J. Arthurs, Martin J. Bishop, and David Kay. Efficient simulation of cardiac electrical propagation using high-order finite elements II: Adaptive p-version. Journal of Computational Physics, 253(??):443–470, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004841.

Akkerman:2011:IAF

[ABKF11] I. Akkerman, Y. Bazilevs, C. E. Kees, and M. W. Farthing. Isogeometric analysis of free-surface flow. Journal of Computational Physics, 230(11):4137–4152, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006595.

Adler:2011:FOS

[ABL^{+}11] J. H. Adler, J. Brannick, C. Liu, T. Manteuffel, and L. Zikatanov. First-order system least squares and the energetic variational approach for two-phase flow. Journal of Computational Physics, 230(17):6647–6663, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002993.

Aniszewski:2012:NAS

[ABMT12] W. Aniszewski, A. Boguslawski, M. Marek, and A. Tyliszczak. A new approach to sub-grid surface tension for LES of twophase flows. Journal of Computational Physics, 231(21): 7368–7397, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003890.

Anjos:2014:MMF

[ABMT14] G. R. Anjos, N. Borhani, N. Mangiavacchi, and J. R. Thome. A 3D moving mesh Finite Element Method for two-phase flows. Journal of Computational Physics, 270(??):366–377, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002551.

Audusse:2011:AHN

[ABPSM11] E. Audusse, M.-O. Bristeau, M. Pelanti, and J. Sainte-Marie. Approximation of the hydrostatic Navier–Stokes system for density stratified flows by a multilayer model: Kinetic interpretation and numerical solution. Journal of Computational Physics, 230(9):3453–3478, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000672.

Abramov:2010:ALR

[Abr10] Rafail V. Abramov. Approximate linear response for slow variables of dynamics with explicit time scale separation. Journal of Computational Physics, 229(20):7739–7746, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000344X.

Audusse:2014:FFV

[ABS^{+}14] Emmanuel Audusse, Fayssal Benkhaldoun, Saida Sari, Mohammed Seaid, and Pablo Tassi. A fast finite volume solver for multi-layered shallow water flows with mass exchange. Journal of Computational Physics, 272(??):23–45, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002897.

Assous:2011:DMT

[AC11] Franck Assous and Joël Chaskalovic. Data mining techniques for scientific computing: Application to asymptotic paraxial approximations to model ultrarelativistic particles. Journal of Computational Physics, 230(12):4811–4827, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001483.

Abgrall:2013:SID

[AC13] Rémi Abgrall and Pietro Marco Congedo. A semi-intrusive deterministic approach to uncertainty quantification in nonlinear fluid flow problems. Journal of Computational Physics, 235(??):828–845, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004263.

Alvaro:2012:NMH

[ÁCB12] M. Álvaro, M. Carretero, and L. L. Bonilla. Numerical method for hydrodynamic modulation equations describing Bloch oscillations in semiconductor superlattices. Journal of Computational Physics, 231(13):4499–4514, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001337.

Audoly:2013:DGA

[ACB^{+}13] B. Audoly, N. Clauvelin, P.-T. Brun, M. Bergou, E. Grinspun, and M. Wardetzky. A discrete geometric approach for simulating the dynamics of thin viscous threads. Journal of Computational Physics, 253(??):18–49, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004646.

Artana:2012:SWC

[ACCM12] G. Artana, A. Cammilleri, J. Carlier, and E. Mémin. Strong and weak constraint variational assimilations for reduced order fluid flow modeling. Journal of Computational Physics, 231(8):3264–3288, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000319.

Abgrall:2014:OTT

[ACG14] R. Abgrall, P. M. Congedo, and G. Geraci. A one-time truncate and encode multiresolution stochastic framework. Journal of Computational Physics, 257(??):19–56, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005342.

Antoine:2014:REP

[AD14] Xavier Antoine and Romain Duboscq. Robust and efficient preconditioned Krylov spectral solvers for computing the ground states of fast rotating and strongly interacting Bose–Einstein condensates. Journal of Computational Physics, 258(??):509–523, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007213.

Avesani:2014:NCM

[ADB14] Diego Avesani, Michael Dumbser, and Alberto Bellin. A new class of Moving-Least-Squares WENO–SPH schemes. Journal of Computational Physics, 270(??):278–299, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002289.

An:2013:IIL

[ADMA13] Weiming An, Viktor K. Decyk, Warren B. Mori, and Thomas M. Antonsen, Jr. An improved iteration loop for the three dimensional quasi-static particle-in-cell algorithm: QuickPIC. Journal of Computational Physics, 250 (??):165–177, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113003525.

Anupindi:2013:NMI

[ADSF13] Kameswararao Anupindi, Yann Delorme, Dinesh A. Shetty, and Steven H. Frankel. A novel multiblock immersed boundary method for large eddy simulation of complex arterial hemodynamics. Journal of Computational Physics, 254(??):200–218, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005238.

Amritkar:2014:EPC

[ADT14] Amit Amritkar, Surya Deb, and Danesh Tafti. Efficient parallel CFD-DEM simulations using OpenMP. Journal of Computational Physics, 256(??):501–519, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006128.

Aland:2014:DIM

[AELV14] Sebastian Aland, Sabine Egerer, John Lowengrub, and Axel Voigt. Diffuse interface models of locally inextensible vesicles in a viscous fluid. Journal of Computational Physics, 277(??):32–47, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005622.

Ashbee:2013:GHP

[AEM13] T. L. Ashbee, J. G. Esler, and N. R. McDonald. Generalized Hamiltonian point vortex dynamics on arbitrary domains using the method of fundamental solutions. Journal of Computational Physics, 246(??):289–303, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002222.

Ariel:2011:GBD

[AETT11] Gil Ariel, Björn Engquist, Nicolay M. Tanushev, and Richard Tsai. Gaussian beam decomposition of high frequency wave fields using expectation-maximization. Journal of Computational Physics, 230(6):2303–2321, March 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006820.

Alvermann:2011:HOC

[AF11] A. Alvermann and H. Fehske. High-order commutator-free exponential time-propagation of driven quantum systems. Journal of Computational Physics, 230(15):5930–5956, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002300.

Apte:2013:VDF

[AF13] Sourabh V. Apte and Justin R. Finn. A variable-density fictitious domain method for particulate flows with broad range of particle-fluid density ratios. Journal of Computational Physics, 243(??):109–129, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007528.

Aguirre:2014:VCF

[AGBC14] Miquel Aguirre, Antonio J. Gil, Javier Bonet, and Aurelio Arranz Carreño. A vertex centred Finite Volume Jameson-Schmidt-Turkel (JST) algorithm for a mixed conservation formulation in solid dynamics. Journal of Computational Physics, 259(??):672–699, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008115.

Arnst:2010:IBP

[AGS10] M. Arnst, R. Ghanem, and C. Soize. Identification of Bayesian posteriors for coefficients of chaos expansions. Journal of Computational Physics, 229(9):3134–3154, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109007190.

Ansari:2011:FDF

[AGSG11] N. Ansari, G. M. Goldin, M. R. H. Sheikhi, and P. Givi. Filtered density function simulator on unstructured meshes. Journal of Computational Physics, 230(19):7132–7150, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003196.

Arbogast:2010:FCE

[AH10a] Todd Arbogast and Chieh-Sen Huang. A fully conservative Eulerian–Lagrangian method for a convection-diffusion problem in a solenoidal field. Journal of Computational Physics, 229(9):3415–3427, May 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000240.

Asheim:2010:LSH

[AH10b] Andreas Asheim and Daan Huybrechs. Local solutions to high-frequency 2D scattering problems. Journal of Computational Physics, 229(14):5357–5372, July 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001531.

Abubakar:2013:TDV

[AH13a] A. Abubakar and T. M. Habashy. Three-dimensional viscoacoustic modeling using a renormalized integral equation iterative solver. Journal of Computational Physics, 249(??): 1–12, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911300257X.

Arshed:2013:MEL

[AH13b] Ghulam M. Arshed and Klaus A. Hoffmann. Minimizing errors from linear and nonlinear weights of WENO scheme for broadband applications with shock waves. Journal of Computational Physics, 246(??):58–77, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002155.

Adami:2010:CSM

[AHA10a] S. Adami, X. Y. Hu, and N. A. Adams. A conservative SPH method for surfactant dynamics. Journal of Computational Physics, 229(5):1909–1926, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006342.

Adami:2010:NST

[AHA10b] S. Adami, X. Y. Hu, and N. A. Adams. A new surfacetension formulation for multi-phase SPH using a reproducing divergence approximation. Journal of Computational Physics, 229(13):5011–5021, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001324.

Adami:2012:GWB

[AHA12] S. Adami, X. Y. Hu, and N. A. Adams. A generalized wall boundary condition for smoothed particle hydrodynamics. Journal of Computational Physics, 231(21):7057–7075, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200229X.

Adami:2013:TVF

[AHA13] S. Adami, X. Y. Hu, and N. A. Adams. A transportvelocity formulation for smoothed particle hydrodynamics. Journal of Computational Physics, 241(??):292–307, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300096X.

Alldredge:2014:ACB

[AHOT14] Graham W. Alldredge, Cory D. Hauck, Dianne P. O’Leary, and André L. Tits. Adaptive change of basis in entropybased moment closures for linear kinetic equations. Journal of Computational Physics, 258(??):489–508, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007250.

Amano:2014:RMH

[AHS14] Takanobu Amano, Katsuaki Higashimori, and Keisuke Shirakawa. A robust method for handling low density regions in hybrid simulations for collisionless plasmas. Journal of Computational Physics, 275(??):197–212, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004550.

Akamatsu:2014:NSC

[AiINT14] Yukinao Akamatsu, Shu ichiro Inutsuka, Chiho Nonaka, and Makoto Takamoto. A new scheme of causal viscous hydrodynamics for relativistic heavy-ion collisions: a Riemann solver for quark-gluon plasma. Journal of Computational Physics, 256(??):34–54, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005913.

Ainsworth:2014:DBH

[Ain14] Mark Ainsworth. Dispersive behaviour of high order finite element schemes for the one-way wave equation. Journal of Computational Physics, 259(??):1–10, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007493.

Abe:2013:CME

[AINF13] Yoshiaki Abe, Nobuyuki Iizuka, Taku Nonomura, and Kozo Fujii. Conservative metric evaluation for high-order finite difference schemes with the GCL identities on moving and deforming grids. Journal of Computational Physics, 232(1):14–21, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004779.

Akoh:2010:MMF

[AIX10] Ryosuke Akoh, Satoshi Ii, and Feng Xiao. A multimoment finite volume formulation for shallow water equations on unstructured mesh. Journal of Computational Physics, 229(12):4567–4590, June 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001038.

Alekseenko:2014:DSS

[AJ14] A. Alekseenko and E. Josyula. Deterministic solution of the spatially homogeneous Boltzmann equation using discontinuous Galerkin discretizations in the velocity space. Journal of Computational Physics, 272(??):170–188, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002186.

Anderson:2013:MPM

[AJG^{+}13] Joshua A. Anderson, Eric Jankowski, Thomas L. Grubb, Michael Engel, and Sharon C. Glotzer. Massively parallel Monte Carlo for many-particle simulations on GPUs. Journal of Computational Physics, 254(??):27–38, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004968.

Abboud:2011:CDG

[AJRT11] Toufic Abboud, Patrick Joly, Jerónimo Rodri’guez, and Isabelle Terrasse. Coupling discontinuous Galerkin methods and retarded potentials for transient wave propagation on unbounded domains. Journal of Computational Physics, 230(15):5877– 5907, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002245.

Allaire:2011:DFE

[AJV11] Grégoire Allaire, François Jouve, and Nicolas Van Goethem. Damage and fracture evolution in brittle materials by shape optimization methods. Journal of Computational Physics, 230(12):5010–5044, June 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001677.

Abgrall:2010:CCN

[AK10] Rémi Abgrall and Smadar Karni. A comment on the computation of non-conservative products. Journal of Computational Physics, 229(8):2759–2763, April 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006949.

Ali:2012:NEG

[AK12] Norhashidah Hj. Mohd. Ali and Lee Ming Kew. New explicit group iterative methods in the solution of two dimensional hyperbolic equations. Journal of Computational Physics, 231(20):6953–6968, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003403.

Abdelrahman:2014:NFT

[AK14] Mahmoud A. E. Abdelrahman and Matthias Kunik. A new front tracking scheme for the ultra-relativistic Euler equations. Journal of Computational Physics, 275(??):213–235, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004665.

Ascenzi:2013:ISM

[AKL^{+}13] Maria-Grazia Ascenzi, Neal P. Kawas, Andre Lutz, Dieter Kardas, Udo Nackenhorst, and Joyce H. Keyak. Individualspecific multi-scale finite element simulation of cortical bone of human proximal femur. Journal of Computational Physics, 244(??):298–311, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002835.

Aubry:2014:TDP

[AKMD14] R. Aubry, B. K. Karamete, E. L. Mestreau, and S. Dey. A three-dimensional parametric mesher with surface boundarylayer capability. Journal of Computational Physics, 270 (??):161–181, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114002447.

Arampatzis:2012:HFS

[AKP^{+}12] Giorgos Arampatzis, Markos A. Katsoulakis, Petr Plechác, Michela Taufer, and Lifan Xu. Hierarchical fractional-step approximations and parallel kinetic Monte Carlo algorithms. Journal of Computational Physics, 231(23):7795–7814, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003907.

Alauzet:2010:HOS

[AL10] F. Alauzet and A. Loseille. High-order sonic boom modeling based on adaptive methods. Journal of Computational Physics, 229(3):561–593, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005129.

Aland:2014:TID

[Ala14] Sebastian Aland. Time integration for diffuse interface models for two-phase flow. Journal of Computational Physics, 262(??):58–71, April 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000102.

Alben:2010:RVS

[Alb10] Silas Alben. Regularizing a vortex sheet near a separation point. Journal of Computational Physics, 229(13):5280–5298, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001634.

Aldredge:2010:SLA

[Ald10] R. C. Aldredge. Semi-Lagrangian advection-propagation (SLAP) scheme for three-dimensional interface tracking. Journal of Computational Physics, 229(12):4683–4702, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001166.

Al-Lehyani:2013:CGS

[ALGB^{+}13] Ibrahim H. Al-Lehyani, John M. A. Grime, Matthew Bano, Kim McKelvey, and Michael P. Allen. Coarse-grained simulation of transmembrane peptides in the gel phase. Journal of Computational Physics, 238(??):97–105, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007450.

Almanasreh:2014:HCA

[Alm14] Hasan Almanasreh. hp-Cloud approximation of the Dirac eigenvalue problem: the way of stability. Journal of Computational Physics, 272(??):487–506, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002332.

Abgrall:2011:CVH

[ALR11] R. Abgrall, A. Larat, and M. Ricchiuto. Construction of very high order residual distribution schemes for steady inviscid flow problems on hybrid unstructured meshes. Journal of Computational Physics, 230(11):4103–4136, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004286.

Antoine:2014:ABC

[ALS^{+}14] X. Antoine, E. Lorin, J. Sater, F. Fillion-Gourdeau, and A. D. Bandrauk. Absorbing boundary conditions for relativistic quantum mechanics equations. Journal of Computational Physics, 277(??):268–304, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005300.

Athenes:2010:FER

[AM10] Manuel Athènes and Mihai-Cosmin Marinica. Free energy reconstruction from steered dynamics without postprocessing. Journal of Computational Physics, 229(19):7129– 7146, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003177.

Assous:2011:SME

[AM11] F. Assous and M. Michaeli. Solving Maxwell’s equations in singular domains with a Nitsche type method. Journal of Computational Physics, 230(12):4922–4939, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001562.

Ainsworth:2012:NUO

[AM12] Mark Ainsworth and Xinhui Ma. Non-uniform order mixed FEM approximation: Implementation, post-processing, computable error bound and adaptivity. Journal of Computational Physics, 231(2):436–453, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005316.

Almeida:2013:ASM

[AM13] F. A. G. Almeida and A. M. S. Macêdo. Association of scattering matrices in quantum networks. Journal of Computational Physics, 243(??):1–13, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300171X.

Alonso-Mallo:2010:SCS

[AMDR10] I. Alonso-Mallo, A. Durán, and N. Reguera. Simulation of coherent structures in nonlinear Schrödinger-type equations. Journal of Computational Physics, 229(21):8180–8198, October 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004110.

Almquist:2014:HFN

[AME14] Martin Almquist, Ken Mattsson, and Tomas Edvinsson. Highfidelity numerical solution of the time-dependent Dirac equation. Journal of Computational Physics, 262(??):86–103, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008462.

Alonso-Mallo:2014:HOF

[AMP14] I. Alonso-Mallo and A. M. Portillo. High order full discretizations of coupled wave equations with absorbing boundary conditions and geometric integration. Journal of Computational Physics, 265(??):16–33, May 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000916.

Amsallem:2013:HOA

[AN13] David Amsallem and Jan Nordström. High-order accurate difference schemes for the Hodgkin–Huxley equations. Journal of Computational Physics, 252(??):573–590, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004658.

Ambrose:2014:FIE

[AN14] David M. Ambrose and David P. Nicholls. Fokas integral equations for three dimensional layered-media scattering. Journal of Computational Physics, 276(??):1–25, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005117.

Anderson:2010:RCP

[And10] Christopher R. Anderson. A Rayleigh–Chebyshev procedure for finding the smallest eigenvalues and associated eigenvectors of large sparse Hermitian matrices. Journal of Computational Physics, 229(19):7477–7487, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003451.

Andrzejewski:2013:OJD

[And13] Janusz Andrzejewski. On optimizing Jacobi–Davidson method for calculating eigenvalues in low dimensional structures using eight band k · p model. Journal of Computational Physics, 249(??):22–35, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002519.

Abe:2014:GIS

[ANIF14] Yoshiaki Abe, Taku Nonomura, Nobuyuki Iizuka, and Kozo Fujii. Geometric interpretations and spatial symmetry property of metrics in the conservative form for high-order finitedifference schemes on moving and deforming grids. Journal of Computational Physics, 260(??):163–203, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008188.

Anonymous:2010:CCI

[Ano10a] Anonymous. Contents continued from IBC. Journal of Computational Physics, 229(19):??, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000402X.

Anonymous:2010:EBa

[Ano10b]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(1):??, January 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005919.

Anonymous:2010:EBb

[Ano10c] Anonymous. Editorial Board. Journal of Computational Physics, 229(2):??, January 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006184.

Anonymous:2010:EBc

[Ano10d] Anonymous. Editorial Board. Journal of Computational Physics, 229(3):??, February 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006421.

Anonymous:2010:EBd

[Ano10e] Anonymous. Editorial Board. Journal of Computational Physics, 229(4):??, February 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006871.

Anonymous:2010:EBe

[Ano10f]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(5):??, March 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109007074.

Anonymous:2010:EBf

[Ano10g]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(6):??, March 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000082.

Anonymous:2010:EBg

[Ano10h]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(7):??, April 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000367.

Anonymous:2010:EBh

[Ano10i]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(8):??, April 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000063X.

Anonymous:2010:EBi

[Ano10j]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(9):??, May 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000938.

Anonymous:2010:EBj

[Ano10k]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(10):??, May 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001075.

Anonymous:2010:EBk

[Ano10l]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(11):??, June 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001373.

Anonymous:2010:EBl

[Ano10m]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(12):??, June 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001701.

Anonymous:2010:EBm

[Ano10n]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(13):??, July 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002081.

Anonymous:2010:EBn

[Ano10o]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(14):??, July 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002342.

Anonymous:2010:EBo

[Ano10p]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 229(15):??, August 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002433.

Anonymous:2010:EBp

[Ano10q] Anonymous. Editorial Board. Journal of Computational Physics, 229(16):??, August 10, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002603.

Anonymous:2010:EBq

[Ano10r] Anonymous. Editorial Board. Journal of Computational Physics, 229(17):??, August 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003025.

Anonymous:2010:EBr

[Ano10s] Anonymous. Editorial Board. Journal of Computational Physics, 229(18):??, September 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003542.

Anonymous:2010:EBs

[Ano10t] Anonymous. Editorial Board. Journal of Computational Physics, 229(19):??, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003803.

Anonymous:2010:EBt

[Ano10u] Anonymous. Editorial Board. Journal of Computational Physics, 229(20):??, October 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003888.

Anonymous:2010:EBu

[Ano10v] Anonymous. Editorial Board. Journal of Computational Physics, 229(21):??, October 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004584.

Anonymous:2010:EBv

[Ano10w] Anonymous. Editorial Board. Journal of Computational Physics, 229(22):??, November 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004821.

Anonymous:2010:EBw

[Ano10x] Anonymous. Editorial Board. Journal of Computational Physics, 229(23):??, November 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005152.

Anonymous:2010:EBx

[Ano10y] Anonymous. Editorial Board. Journal of Computational Physics, 229(24):??, December 10, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000536X.

Anonymous:2010:ICCa

[Ano10z] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(2):??, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006214.

Anonymous:2010:ICCb

[Ano10-27] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(3):??, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006457.

Anonymous:2010:ICCc

[Ano10-28] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(4):??, February 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006901.

Anonymous:2010:ICCd

[Ano10-29] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(5):??, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109007104.

Anonymous:2010:ICCe

[Ano10-30] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(6):??, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000112.

Anonymous:2010:ICCf

[Ano10-31] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(7):??, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000392.

Anonymous:2010:ICCg

[Ano10-32] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(8):??, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000665.

Anonymous:2010:ICCh

[Ano10-33] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(9):??, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000963.

Anonymous:2010:ICCi

[Ano10-34] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(10):??, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001105.

Anonymous:2010:ICCj

[Ano10-35] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(11):??, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001403.

Anonymous:2010:ICCk

[Ano10-36] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(12):??, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001737.

Anonymous:2010:ICCl

[Ano10-37] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(13):??, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002111.

Anonymous:2010:ICCm

[Ano10-38] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(14):??, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002378.

Anonymous:2010:ICCn

[Ano10-39] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(15):??, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002469.

Anonymous:2010:ICCo

[Ano10-40] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(16):??, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002639.

Anonymous:2010:ICCp

[Ano10-41] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(17):??, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003050.

Anonymous:2010:ICCq

[Ano10-42] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(18):??, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003578.

Anonymous:2010:ICCr

[Ano10-43] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(19):??, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003839.

Anonymous:2010:ICCs

[Ano10-44] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(20):??, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003918.

Anonymous:2010:ICCt

[Ano10-45] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(21):??, October 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004614.

Anonymous:2010:ICCu

[Ano10-46] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(22):??, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004857.

Anonymous:2010:ICCv

[Ano10-47] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(23):??, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005188.

Anonymous:2010:ICCw

[Ano10-48] Anonymous. IBC (contents continued). Journal of Computational Physics, 229(24):??, December 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005395.

Anonymous:2010:OICa

[Ano10-49] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(1):??, January 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005956.

Anonymous:2010:OICb

[Ano10-50] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(2):??, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006226.

Anonymous:2010:OICc

[Ano10-51] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(3):??, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006469.

Anonymous:2010:OICd

[Ano10-52] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(4):??, February 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006913.

Anonymous:2010:OICe

[Ano10-53] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(5):??, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109007116.

Anonymous:2010:OICf

[Ano10-54] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(6):??, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000124.

Anonymous:2010:OICg

[Ano10-55] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(7):??, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000409.

Anonymous:2010:OICh

[Ano10-56] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(8):??, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000677.

Anonymous:2010:OICi

[Ano10-57] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(9):??, May 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000975.

Anonymous:2010:OICj

[Ano10-58] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(10):??, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001117.

Anonymous:2010:OICk

[Ano10-59] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(11):??, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001415.

Anonymous:2010:OICl

[Ano10-60] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(12):??, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001749.

Anonymous:2010:OICm

[Ano10-61] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(13):??, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002123.

Anonymous:2010:OICn

[Ano10-62] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(14):??, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000238X.

Anonymous:2010:OICo

[Ano10-63] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(15):??, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002470.

Anonymous:2010:OICp

[Ano10-64] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(16):??, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002640.

Anonymous:2010:OICq

[Ano10-65] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(17):??, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003062.

Anonymous:2010:OICr

[Ano10-66] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(18):??, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000358X.

Anonymous:2010:OICs

[Ano10-67] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(19):??, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003840.

Anonymous:2010:OICt

[Ano10-68] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(20):??, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000392X.

Anonymous:2010:OICu

[Ano10-69] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(21):??, October 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004626.

Anonymous:2010:OICv

[Ano10-70] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(22):??, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004869.

Anonymous:2010:OICw

[Ano10-71] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(23):??, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000519X.

Anonymous:2010:OICx

[Ano10-72] Anonymous. OBC (issue contents). Journal of Computational Physics, 229(24):??, December 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005401.

Anonymous:2011:CC

[Ano11a] Anonymous. Contents continued. Journal of Computational Physics, 230(12):??, June 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002531.

Anonymous:2011:CCI

[Ano11b] Anonymous. Contents continued from IBC. Journal of Computational Physics, 230(4):??, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006972.

Anonymous:2011:EBa

[Ano11c]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(1):??, January 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005632.

Anonymous:2011:EBb

[Ano11d] Anonymous. Editorial Board. Journal of Computational Physics, 230(2):??, January 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005930.

Anonymous:2011:EBc

[Ano11e] Anonymous. Editorial Board. Journal of Computational Physics, 230(3):??, February 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006273.

Anonymous:2011:EBd

[Ano11f] Anonymous. Editorial Board. Journal of Computational Physics, 230(4):??, February 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006893.

Anonymous:2011:EBe

[Ano11g]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(5):??, March 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000222.

Anonymous:2011:EBf

[Ano11h]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(6):??, March 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000416.

Anonymous:2011:EBg

[Ano11i]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(7):??, April 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000702.

Anonymous:2011:EBh

[Ano11j]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(8):??, April 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000969.

Anonymous:2011:EBi

[Ano11k]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(9):??, May 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001380.

Anonymous:2011:EBj

[Ano11l]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(10):??, May 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001707.

Anonymous:2011:EBk

[Ano11m]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(11):??, May 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001926.

Anonymous:2011:EBl

[Ano11n]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(12):??, June 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002397.

Anonymous:2011:EBm

[Ano11o]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(13):??, June 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002634.

Anonymous:2011:EBn

[Ano11p]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(14):??, June 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002889.

Anonymous:2011:EBo

[Ano11q]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(15):??, July 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003056.

Anonymous:2011:EBp

[Ano11r]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(16):??, July 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003263.

Anonymous:2011:EBq

[Ano11s]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(17):??, July 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003639.

Anonymous:2011:EBr

[Ano11t]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 230(18):??, August 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003895.

Anonymous:2011:EBs

[Ano11u] Anonymous. Editorial Board. Journal of Computational Physics, 230(19):??, August 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004189.

Anonymous:2011:EBt

[Ano11v] Anonymous. Editorial Board. Journal of Computational Physics, 230(20):??, August 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004414.

Anonymous:2011:EBu

[Ano11w] Anonymous. Editorial Board. Journal of Computational Physics, 230(21):??, September 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004657.

Anonymous:2011:EBv

[Ano11x] Anonymous. Editorial Board. Journal of Computational Physics, 230(22):??, September 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004876.

Anonymous:2011:EBw

[Ano11y] Anonymous. Editorial Board. Journal of Computational Physics, 230(23):??, September 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100533X.

Anonymous:2011:EBx

[Ano11z] Anonymous. Editorial Board. Journal of Computational Physics, 230(24):??, October 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005559.

Anonymous:2011:ICCa

[Ano11-27] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(3):??, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006303.

Anonymous:2011:ICCb

[Ano11-28] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(4):??, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006923.

Anonymous:2011:ICCc

[Ano11-29] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(5):??, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000192.

Anonymous:2011:ICCd

[Ano11-30] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(7):??, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000738.

Anonymous:2011:ICCe

[Ano11-31] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(8):??, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000994.

Anonymous:2011:ICCf

[Ano11-32] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(9):??, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001410.

Anonymous:2011:ICCg

[Ano11-33] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(10):??, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001732.

Anonymous:2011:ICCh

[Ano11-34] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(11):??, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001951.

Anonymous:2011:ICCi

[Ano11-35] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(12):??, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002427.

Anonymous:2011:ICCj

[Ano11-36] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(13):??, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100266X.

Anonymous:2011:ICCk

[Ano11-37] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(14):??, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002919.

Anonymous:2011:ICCl

[Ano11-38] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(15):??, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003081.

Anonymous:2011:ICCm

[Ano11-39] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(17):??, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003706.

Anonymous:2011:ICCn

[Ano11-40] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(19):??, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004268.

Anonymous:2011:ICCo

[Ano11-41] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(22):??, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004979.

Anonymous:2011:ICCp

[Ano11-42] Anonymous. IBC (contents continued). Journal of Computational Physics, 230(23):??, September 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005407.

Anonymous:2011:OICa

[Ano11-43] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(1):??, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000567X.

Anonymous:2011:OICb

[Ano11-44] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(2):??, January 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005978.

Anonymous:2011:OICc

[Ano11-45] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(3):??, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006315.

Anonymous:2011:OICd

[Ano11-46] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(4):??, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006935.

Anonymous:2011:OICe

[Ano11-47] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(5):??, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000209.

Anonymous:2011:OICf

[Ano11-48] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(6):??, March 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000453.

Anonymous:2011:OICg

[Ano11-49] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(7):??, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100074X.

Anonymous:2011:OICh

[Ano11-50] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(8):??, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001008.

Anonymous:2011:OICi

[Ano11-51] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(9):??, May 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001422.

Anonymous:2011:OICj

[Ano11-52] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(10):??, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001744.

Anonymous:2011:OICk

[Ano11-53] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(11):??, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001963.

Anonymous:2011:OICl

[Ano11-54] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(12):??, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002439.

Anonymous:2011:OICm

[Ano11-55] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(13):??, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002671.

Anonymous:2011:OICn

[Ano11-56] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(14):??, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002920.

Anonymous:2011:OICo

[Ano11-57] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(15):??, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003093.

Anonymous:2011:OICp

[Ano11-58] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(16):??, July 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003305.

Anonymous:2011:OICq

[Ano11-59] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(17):??, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003718.

Anonymous:2011:OICr

[Ano11-60] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(18):??, August 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003962.

Anonymous:2011:OICs

[Ano11-61] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(19):??, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100427X.

Anonymous:2011:OICt

[Ano11-62] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(20):??, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004487.

Anonymous:2011:OICu

[Ano11-63] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(21):??, September 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004748.

Anonymous:2011:OICv

[Ano11-64] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(22):??, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004980.

Anonymous:2011:OICw

[Ano11-65] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(23):??, September 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005419.

Anonymous:2011:OICx

[Ano11-66] Anonymous. OBC (issue contents). Journal of Computational Physics, 230(24):??, October 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005808.

Anonymous:2012:CCI

[Ano12a] Anonymous. Contents continued from IBC. Journal of Computational Physics, 231(4):??, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007157.

Anonymous:2012:EBa

[Ano12b]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(1):??, January 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100605X.

Anonymous:2012:EBb

[Ano12c] Anonymous. Editorial Board. Journal of Computational Physics, 231(2):??, January 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006358.

Anonymous:2012:EBc

[Ano12d] Anonymous. Editorial Board. Journal of Computational Physics, 231(3):??, February 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006607.

Anonymous:2012:EBd

[Ano12e] Anonymous. Editorial Board. Journal of Computational Physics, 231(4):??, February 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111007091.

Anonymous:2012:EBe

[Ano12f]  Anonymous.  Editorial board.  Journal of Computa-

tional Physics, 231(5):??, March 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111007388.

Anonymous:2012:EBf

[Ano12g]  Anonymous.  Editorial board.  Journal of Computa-

tional Physics, 231(6):??, March 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000162.

Anonymous:2012:EBg

[Ano12h]  Anonymous.  Editorial board.  Journal of Computa-

tional Physics, 231(7):??, April 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000502.

Anonymous:2012:EBh

[Ano12i]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(8):??, April 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000769.

Anonymous:2012:EBi

[Ano12j]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(9):??, May 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000964.

Anonymous:2012:EBj

[Ano12k]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(10):??, May 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911200109X.

Anonymous:2012:EBk

[Ano12l]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(11):??, June 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001258.

Anonymous:2012:EBl

[Ano12m]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(12):??, June 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001611.

Anonymous:2012:EBm

[Ano12n]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 231(13):??, May 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001908.

Anonymous:2012:ICCa

[Ano12o] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(2):??, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006413.

Anonymous:2012:ICCb

[Ano12p] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(3):??, February 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006668.

Anonymous:2012:ICCc

[Ano12q] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(4):??, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007133.

Anonymous:2012:ICCd

[Ano12r] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(6):??, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000216.

Anonymous:2012:ICCe

[Ano12s] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(7):??, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000563.

Anonymous:2012:ICCf

[Ano12t] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(8):??, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000824.

Anonymous:2012:ICCg

[Ano12u] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(9):??, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911200099X.

Anonymous:2012:ICCh

[Ano12v] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(11):??, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001283.

Anonymous:2012:ICCi

[Ano12w] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(14):??, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002410.

Anonymous:2012:ICCj

[Ano12x] Anonymous. IBC contents continued. Journal of Computational Physics, 231(16):??, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002938.

Anonymous:2012:ICCk

[Ano12y] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(17):??, July 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003282.

Anonymous:2012:ICCl

[Ano12z] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(18):??, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003531.

Anonymous:2012:ICCm

[Ano12-27] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(19):??, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003816.

Anonymous:2012:ICCn

[Ano12-28] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(20):??, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004032.

Anonymous:2012:ICCo

[Ano12-29] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(21):??, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004469.

Anonymous:2012:ICCp

[Ano12-30] Anonymous. IBC (contents continued). Journal of Computational Physics, 231(24):??, October 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005281.

Anonymous:2012:IEF

[Ano12-31] Anonymous. IBC elsvr filler: LANGUAGE ED. Journal of Computational Physics, 231(13):??, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002112.

Anonymous:2012:IFA

[Ano12-32] Anonymous. IBC filler ad: SCIVAL BUTTERFLY. Journal of Computational Physics, 231(15):??, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002719.

Anonymous:2012:IFL

[Ano12-33] Anonymous. IBC filler: LANGUAGE ED. Journal of Computational Physics, 231(10):??, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200112X.

Anonymous:2012:IFSa

[Ano12-34] Anonymous. IBC filler: SCIVERSE APPS. Journal of Computational Physics, 231(22):??, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004822.

Anonymous:2012:IFSb

[Ano12-35] Anonymous. IBC filler: SCIVERSE SCIENCEDIRECT. Journal of Computational Physics, 231(23):??, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004986.

Anonymous:2012:OICa

[Ano12-36] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(1):??, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005973.

Anonymous:2012:OICb

[Ano12-37] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(2):??, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006425.

Anonymous:2012:OICc

[Ano12-38] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(3):??, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100667X.

Anonymous:2012:OICd

[Ano12-39] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(4):??, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007145.

Anonymous:2012:OICe

[Ano12-40] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(5):??, March 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111007431.

Anonymous:2012:OICf

[Ano12-41] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(6):??, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000228.

Anonymous:2012:OICg

[Ano12-42] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(7):??, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000575.

Anonymous:2012:OICh

[Ano12-43] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(8):??, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000836.

Anonymous:2012:OICi

[Ano12-44] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(9):??, May 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001003.

Anonymous:2012:OICj

[Ano12-45] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(10):??, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001131.

Anonymous:2012:OICk

[Ano12-46] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(11):??, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001295.

Anonymous:2012:OICl

[Ano12-47] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(12):??, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001659.

Anonymous:2012:OICm

[Ano12-48] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(13):??, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001970.

Anonymous:2012:OICn

[Ano12-49] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(14):??, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002422.

Anonymous:2012:OICo

[Ano12-50] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(15):??, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002720.

Anonymous:2012:OICp

[Ano12-51] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(16):??, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911200294X.

Anonymous:2012:OICq

[Ano12-52] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(17):??, July 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003294.

Anonymous:2012:OICr

[Ano12-53] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(18):??, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003543.

Anonymous:2012:OICs

[Ano12-54] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(19):??, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003828.

Anonymous:2012:OICt

[Ano12-55] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(20):??, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004044.

Anonymous:2012:OICu

[Ano12-56] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(21):??, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004470.

Anonymous:2012:OICv

[Ano12-57] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(22):??, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004834.

Anonymous:2012:OICw

[Ano12-58] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(23):??, October 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004998.

Anonymous:2012:OICx

[Ano12-59] Anonymous. OBC (issue contents). Journal of Computational Physics, 231(24):??, October 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005293.

Anonymous:2013:Ca

[Ano13a] Anonymous. Contents. Journal of Computational Physics, 252(??):??, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005184.

Anonymous:2013:Cb

[Ano13b] Anonymous. Contents. Journal of Computational Physics, 252(??):??, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005196.

Anonymous:2013:Cc

[Ano13c] Anonymous. Contents. Journal of Computational Physics, 253(??):??, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005470.

Anonymous:2013:Cd

[Ano13d] Anonymous. Contents. Journal of Computational Physics, 253(??):??, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005482.

Anonymous:2013:Ce

[Ano13e] Anonymous. Contents. Journal of Computational Physics, 254(??):??, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005834.

Anonymous:2013:Cf

[Ano13f] Anonymous. Contents. Journal of Computational Physics, 255(??):??, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006360.

Anonymous:2013:CCa

[Ano13g] Anonymous. Contents continued. Journal of Computational Physics, 235(??):iv, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000247.

Anonymous:2013:CCb

[Ano13h] Anonymous. Contents continued. Journal of Computational Physics, 242(??):iv, June 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003021.

Anonymous:2013:EBa

[Ano13i]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 232(1):??, January 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112005608.

Anonymous:2013:EBb

[Ano13j]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 244(??):??, July 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002787.

Anonymous:2013:EBc

[Ano13k]  Anonymous.  Editorial Board.  Journal of Computa-

tional Physics, 246(??):??, August 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003720.

Anonymous:2013:EBd

[Ano13l] Anonymous. Editorial Board. Journal of Computational Physics, 251(??):??, October 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005020.

Anonymous:2013:ICCa

[Ano13m] Anonymous. IBC (contents continued). Journal of Computational Physics, 232(1):??, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112005633.

Anonymous:2013:ICCb

[Ano13n] Anonymous. IBC (contents continued). Journal of Computational Physics, 233(??):??, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006158.

Anonymous:2013:ICCc

[Ano13o] Anonymous. IBC (contents continued). Journal of Computational Physics, 234(??):??, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006778.

Anonymous:2013:ICCd

[Ano13p] Anonymous. IBC (contents continued). Journal of Computational Physics, 235(??):??, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000156.

Anonymous:2013:ICCe

[Ano13q] Anonymous. IBC (contents continued). Journal of Computational Physics, 236(??):??, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000624.

Anonymous:2013:ICCf

[Ano13r] Anonymous. IBC (contents continued). Journal of Computational Physics, 237(??):??, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000831.

Anonymous:2013:ICCg

[Ano13s] Anonymous. IBC (contents continued). Journal of Computational Physics, 238(??):??, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001071.

Anonymous:2013:ICCh

[Ano13t] Anonymous. IBC (contents continued). Journal of Computational Physics, 241(??):??, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002313.

Anonymous:2013:ICCi

[Ano13u] Anonymous. IBC (contents continued). Journal of Computational Physics, 242(??):??, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003008.

Anonymous:2013:ICCj

[Ano13v] Anonymous. IBC (contents continued). Journal of Computational Physics, 243(??):??, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003112.

Anonymous:2013:ICCk

[Ano13w] Anonymous. IBC (contents continued). Journal of Computational Physics, 244(??):??, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002817.

Anonymous:2013:ICCl

[Ano13x] Anonymous. IBC (contents continued). Journal of Computational Physics, 245(??):??, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003574.

Anonymous:2013:ICCm

[Ano13y] Anonymous. IBC (contents continued). Journal of Computational Physics, 246(??):??, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003756.

Anonymous:2013:ICCn

[Ano13z] Anonymous. IBC (contents continued). Journal of Computational Physics, 247(??):??, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003835.

Anonymous:2013:ICCo

[Ano13-27] Anonymous. IBC (contents continued). Journal of Computational Physics, 248(??):??, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003914.

Anonymous:2013:ICCp

[Ano13-28] Anonymous. IBC (contents continued). Journal of Computational Physics, 249(??):??, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004233.

Anonymous:2013:ICCq

[Ano13-29] Anonymous. IBC (contents continued). Journal of Computational Physics, 250(??):??, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300452X.

Anonymous:2013:ICCr

[Ano13-30] Anonymous. IBC (contents continued). Journal of Computational Physics, 251(??):??, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005056.

Anonymous:2013:ILa

[Ano13-31] Anonymous. IBC LANGUAGE ED. Journal of Computational Physics, 239(??):??, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911300137X.

Anonymous:2013:ILb

[Ano13-32] Anonymous. IBC: LANGUAGE ED. Journal of Computational Physics, 240(??):??, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001666.

Anonymous:2013:OICa

[Ano13-33] Anonymous. OBC (issue contents). Journal of Computational Physics, 232(1):??, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112005645.

Anonymous:2013:OICb

[Ano13-34] Anonymous. OBC (issue contents). Journal of Computational Physics, 233(??):??, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200616X.

Anonymous:2013:OICc

[Ano13-35] Anonymous. OBC (issue contents). Journal of Computational Physics, 234(??):??, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200678X.

Anonymous:2013:OICd

[Ano13-36] Anonymous. OBC (issue contents). Journal of Computational Physics, 235(??):??, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000168.

Anonymous:2013:OICe

[Ano13-37] Anonymous. OBC (issue contents). Journal of Computational Physics, 236(??):??, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000636.

Anonymous:2013:OICf

[Ano13-38] Anonymous. OBC (issue contents). Journal of Computational Physics, 237(??):??, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000843.

Anonymous:2013:OICg

[Ano13-39] Anonymous. OBC (issue contents). Journal of Computational Physics, 238(??):??, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001083.

Anonymous:2013:OICh

[Ano13-40] Anonymous. OBC (issue contents). Journal of Computational Physics, 239(??):??, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001381.

Anonymous:2013:OICi

[Ano13-41] Anonymous. OBC (issue contents). Journal of Computational Physics, 240(??):??, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001678.

Anonymous:2013:OICj

[Ano13-42] Anonymous. OBC (issue contents). Journal of Computational Physics, 241(??):??, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002325.

Anonymous:2013:OICk

[Ano13-43] Anonymous. OBC (issue contents). Journal of Computational Physics, 242(??):??, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911300301X.

Anonymous:2013:OICl

[Ano13-44] Anonymous. OBC (issue contents). Journal of Computational Physics, 243(??):??, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003124.

Anonymous:2013:OICm

[Ano13-45] Anonymous. OBC (issue contents). Journal of Computational Physics, 244(??):??, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002829.

Anonymous:2013:OICn

[Ano13-46] Anonymous. OBC (issue contents). Journal of Computational Physics, 245(??):??, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003586.

Anonymous:2013:OICo

[Ano13-47] Anonymous. OBC (issue contents). Journal of Computational Physics, 246(??):??, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003768.

Anonymous:2013:OICp

[Ano13-48] Anonymous. OBC (issue contents). Journal of Computational Physics, 247(??):??, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003847.

Anonymous:2013:OICq

[Ano13-49] Anonymous. OBC (issue contents). Journal of Computational Physics, 248(??):??, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003926.

Anonymous:2013:OICr

[Ano13-50] Anonymous. OBC (issue contents). Journal of Computational Physics, 249(??):??, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004245.

Anonymous:2013:OICs

[Ano13-51] Anonymous. OBC (issue contents). Journal of Computational Physics, 250(??):??, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113004531.

Anonymous:2013:OICt

[Ano13-52] Anonymous. OBC (issue contents). Journal of Computational Physics, 251(??):??, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005068.

Anonymous:2013:PN

[Ano13-53] Anonymous. Publisher note. Journal of Computational Physics, 236(??):??, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113000648.

Apology for printing errors in [ASK^{+}13]; the online version is correct.

Anonymous:2014:Ca

[Ano14a] Anonymous. Contents. Journal of Computational Physics, 256(??):??, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006888.

Anonymous:2014:Cb

[Ano14b] Anonymous. Contents. Journal of Computational Physics, 257(??):??, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007456.

Anonymous:2014:Cc

[Ano14c] Anonymous. Contents. Journal of Computational Physics, 257(??):??, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007699.

Anonymous:2014:Cd

[Ano14d] Anonymous. Contents. Journal of Computational Physics, 258(??):??, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113008280.

Anonymous:2014:Ce

[Ano14e] Anonymous. Contents. Journal of Computational Physics, 259(??):??, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114000072.

Anonymous:2014:Cf

[Ano14f] Anonymous. Contents. Journal of Computational Physics, 260(??):??, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114000382.

Anonymous:2014:Cg

[Ano14g] Anonymous. Contents. Journal of Computational Physics, 261(??):??, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114000734.

Anonymous:2014:Ch

[Ano14h] Anonymous. Contents. Journal of Computational Physics, 262(??):??, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114000886.

Anonymous:2014:Ci

[Ano14i] Anonymous. Contents. Journal of Computational Physics, 263(??):??, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114001211.

Anonymous:2014:Cj

[Ano14j] Anonymous. Contents. Journal of Computational Physics, 264(??):??, May 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114001272.

Anonymous:2014:Ck

[Ano14k] Anonymous. Contents. Journal of Computational Physics, 265(??):??, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114001582.

Anonymous:2014:Cl

[Ano14l] Anonymous. Contents. Journal of Computational Physics, 266(??):??, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114001946.

Anonymous:2014:Cm

[Ano14m] Anonymous. Contents. Journal of Computational Physics, 267(??):??, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114002162.

Anonymous:2014:Cn

[Ano14n] Anonymous. Contents. Journal of Computational Physics, 268(??):??, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114002691.

Anonymous:2014:Co

[Ano14o] Anonymous. Contents. Journal of Computational Physics, 269(??):??, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114003064.

Anonymous:2014:Cp

[Ano14p] Anonymous. Contents. Journal of Computational Physics, 270(??):??, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114003507.

Anonymous:2014:Cq

[Ano14q] Anonymous. Contents. Journal of Computational Physics, 271(??):??, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114003234.

Anonymous:2014:Cr

[Ano14r] Anonymous. Contents. Journal of Computational Physics, 272(??):??, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911400388X.

Anonymous:2014:Cs

[Ano14s] Anonymous. Contents. Journal of Computational Physics, 273(??):IBC, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114004604.

Anonymous:2014:Ct

[Ano14t] Anonymous. Contents. Journal of Computational Physics, 273(??):OBC, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114004616.

Anonymous:2014:Cu

[Ano14u] Anonymous. Contents. Journal of Computational Physics, 274(??):II, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005075.

Anonymous:2014:Cv

[Ano14v] Anonymous. Contents. Journal of Computational Physics, 274(??):IBC, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005002.

Anonymous:2014:Cw

[Ano14w] Anonymous. Contents. Journal of Computational Physics, 274(??):OBC, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005014.

Anonymous:2014:Cx

[Ano14x] Anonymous. Contents. Journal of Computational Physics, 275(??):IBC, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005671.

Anonymous:2014:Cy

[Ano14y] Anonymous. Contents. Journal of Computational Physics, 275(??):OBC, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005683.

Anonymous:2014:Cz

[Ano14z] Anonymous. Contents. Journal of Computational Physics, 276(??):IBC, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911400597X.

Anonymous:2014:Caa

[Ano14-27] Anonymous. Contents. Journal of Computational Physics, 276(??):OBC, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005981.

Anonymous:2014:Cab

[Ano14-28] Anonymous. Contents. Journal of Computational Physics, 277(??):IBC, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006135.

Anonymous:2014:Cac

[Ano14-29] Anonymous. Contents. Journal of Computational Physics, 277(??):OBC, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006147.

Anonymous:2014:Cad

[Ano14-30] Anonymous. Contents. Journal of Computational Physics, 278(??):IBC, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911400638X.

Anonymous:2014:Cae

[Ano14-31] Anonymous. Contents. Journal of Computational Physics, 278(??):OBC, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006391.

Anonymous:2014:Caf

[Ano14-32] Anonymous. Contents. Journal of Computational Physics, 279(??):IBC, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006858.

Anonymous:2014:Cag

[Ano14-33] Anonymous. Contents. Journal of Computational Physics, 279(??):OBC, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911400686X.

Anonymous:2014:EB

[Ano14-34] Anonymous. Editorial Board. Journal of Computational Physics, 276(??):ii, November 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005944.

Asinari:2012:LWA

[AOCR12] Pietro Asinari, Taku Ohwada, Eliodoro Chiavazzo, and Antonio F. Di Rienzo. Link-wise artificial compressibility method. Journal of Computational Physics, 231(15):5109–5143, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002082.

Alaia:2011:HMH

[AP11] Alessandro Alaia and Gabriella Puppo. A hybrid method for hydrodynamic-kinetic flow Part I: a particle-grid method for reducing stochastic noise in kinetic regimes. Journal of Computational Physics, 230(14):5660–5683, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002087.

Abdulle:2012:NMS

[AP12a] A. Abdulle and G. A. Pavliotis. Numerical methods for stochastic partial differential equations with multiple scales. Journal of Computational Physics, 231(6):2482–2497, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006954.

Alaia:2012:HMH

[AP12b] Alessandro Alaia and Gabriella Puppo. A hybrid method for hydrodynamic-kinetic flow — Part II — coupling of hydrodynamic and kinetic models. Journal of Computational Physics, 231(16):5217–5242, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001234.

Avila:2011:FED

[APC11] Matias Avila, Javier Principe, and Ramon Codina. A finite element dynamical nonlinear subscale approximation for the low Mach number flow equations. Journal of Computational Physics, 230(22):7988–8009, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004104.

Aanjaneya:2013:MMT

[APF13] Mridul Aanjaneya, Saket Patkar, and Ronald Fedkiw. A monolithic mass tracking formulation for bubbles in incompressible flow. Journal of Computational Physics, 247(??):17–61, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300226X.

Alexandru:2012:MMS

[APGL12] A. Alexandru, C. Pelissier, B. Gamari, and F. X. Lee. Multimass solvers for lattice QCD on GPUs. Journal of Computational Physics, 231(4):1866–1878, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006450.

Astala:2014:NFA

[APRS14] Kari Astala, Lassi Päivärinta, Juan Manuel Reyes, and Samuli Siltanen. Nonlinear Fourier analysis for discontinuous conductivities: Computational results. Journal of Computational Physics, 276(??):74–91, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005257.

Arraras:2014:EAM

[APY14] A. Arrarás, L. Portero, and I. Yotov. Error analysis of multipoint flux domain decomposition methods for evolutionary diffusion problems. Journal of Computational Physics, 257(??):1321–1351, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005512.

Abgrall:2011:PSI

[AQ11] Rémi Abgrall and Jianxian Qiu. Preface to the special issue “High order methods for CFD problems”. Journal of Computational Physics, 230(11):4101–4102, May 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001471.

Adcock:2014:PSN

[AR14] Ben Adcock and Joseph Ruan. Parameter selection and numerical approximation properties of Fourier extensions from fixed data. Journal of Computational Physics, 273(??):453– 471, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114003994.

Allsopp:2012:MDB

[ARF12] Nicholas Allsopp, Giancarlo Ruocco, and Andrea Fratalocchi. Molecular dynamics beyonds the limits: Massive scaling on 72 racks of a BlueGene/P and supercooled glass dynamics of a 1 billion particles system. Journal of Computational Physics, 231(8):3432–3445, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200040X.

Acebron:2011:NPS

[ARR11] Juan A. Acebrón and Ángel Rodríguez-Rozas. A new parallel solver suited for arbitrary semilinear parabolic partial differential equations based on generalized random trees. Journal of Computational Physics, 230(21):7891–7909, September 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004165.

Acebron:2013:HEN

[ARR13] Juan A. Acebrón and Ángel Rodríguez-Rozas. Highly efficient numerical algorithm based on random trees for accelerating parallel Vlasov–Poisson simulations. Journal of Computational Physics, 250(??):224–245, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003653.

Anistratov:2012:CTM

[AS12] Dmitriy Y. Anistratov and Nicholas D. Stehle. Computational transport methodology based on decomposition of a problem domain into transport and diffusive subdomains. Journal of Computational Physics, 231(24):8009–8028, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003154.

Alexe:2014:STA

[AS14] Mihai Alexe and Adrian Sandu. Space–time adaptive solution of inverse problems with the discrete adjoint method. Journal of Computational Physics, 270(??):21–39, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002290.

Adams:2010:TTM

[ASB10] Mark F. Adams, Ravi Samtaney, and Achi Brandt. Toward textbook multigrid efficiency for fully implicit resistive magnetohydrodynamics. Journal of Computational Physics, 229(18):6208–6219, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002007.

Aragon:2011:MPD

[ASGW11] Alejandro M. Aragón, Kyle J. Smith, Philippe H. Geubelle, and Scott R. White. Multi-physics design of microvascular materials for active cooling applications. Journal of Computational Physics, 230(13):5178–5198, June 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001550.

Aragon:2013:MPO

[ASK^{+}13] Alejandro M. Aragón, Rajat Saksena, Brian D. Kozola, Philippe H. Geubelle, Kenneth T. Christensen, and Scott R. White. Multi-physics optimization of three-dimensional microvascular polymeric components. Journal of Computational Physics, 233(??):132–147, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004214. The printed version con-

tains errors introduced during journal production. The online version is correct; see [Ano13-53].

Almanasreh:2013:SFE

[ASS13] Hasan Almanasreh, Sten Salomonson, and Nils Svanstedt. Stabilized finite element method for the radial Dirac equation. Journal of Computational Physics, 236(??):426–442, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007000. See corrigendum

[ASS17].

Almanasreh:2017:CSF

[ASS17] Hasan Almanasreh, Sten Salomonson, and Nils Svanstedt. Corrigendum to “Stabilized finite element method for the radial Dirac equation” [J. Comput. Phys. 236 (2013) 426– 442]. Journal of Computational Physics, 340(??):712, July 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999117301432.

Arico:2012:MEC

[AST12] Costanza Aricò, Marco Sinagra, and Tullio Tucciarelli. The MAST-edge centred lumped scheme for the flow simulation in variably saturated heterogeneous porous media. Journal of Computational Physics, 231(4):1387–1425, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006103.

Ambrose:2013:SSD

[AST13] David M. Ambrose, Michael Siegel, and Svetlana Tlupova. A small-scale decomposition for 3D boundary integral computations with surface tension. Journal of Computational Physics, 247(??):168–191, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002234.

Abgrall:2014:FCP

[ASXZ14] Rémi Abgrall, Piotr Smolarkiewicz, Feng Xiao, and Stéphane Zaleski. Frontiers in computational physics: Modeling the Earth system. Journal of Computational Physics, 271(??):1, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002563.

Arico:2013:MSH

[AT13a] Costanza Aricò and Tullio Tucciarelli. Monotonic solution of heterogeneous anisotropic diffusion problems. Journal of Computational Physics, 252(??):219–249, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004415.

Assencio:2013:SOV

[AT13b] Diego C. Assêncio and Joseph M. Teran. A second order virtual node algorithm for Stokes flow problems with interfacial forces, discontinuous material properties and irregular domains. Journal of Computational Physics, 250(??):77–105, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003355.

Atzberger:2010:SAS

[Atz10] Paul J. Atzberger. Spatially adaptive stochastic numerical methods for intrinsic fluctuations in reaction-diffusion systems. Journal of Computational Physics, 229(9):3474–3501, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000276.

Atzberger:2011:SEL

[Atz11] Paul J. Atzberger. Stochastic Eulerian Lagrangian methods for fluid-structure interactions with thermal fluctuations. Journal of Computational Physics, 230(8):2821–2837, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110007023.

Acosta:2010:CDN

[AV10] Sebastian Acosta and Vianey Villamizar. Coupling of Dirichletto-Neumann boundary condition and finite difference methods in curvilinear coordinates for multiple scattering. Journal of Computational Physics, 229(15):5498–5517, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001877.

Abdulle:2013:PSK

[AV13] Assyr Abdulle and Gilles Vilmart. PIROCK: a Swissknife partitioned implicit-explicit orthogonal Runge–Kutta Chebyshev integrator for stiff diffusion-advection-reaction problems with or without noise. Journal of Computational Physics, 242(??):869–888, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001198.

Abukhdeir:2011:LTI

[AVKP11] Nasser Mohieddin Abukhdeir, Dionisios G. Vlachos, Markos Katsoulakis, and Michael Plexousakis. Long-time integration methods for mesoscopic models of pattern-forming systems. Journal of Computational Physics, 230(14):5704–5715, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002130.

Araki:2014:CCP

[AW14] Samuel J. Araki and Richard E. Wirz. Cell-centered particle weighting algorithm for PIC simulations in a nonuniform 2D axisymmetric mesh. Journal of Computational Physics, 272(??):218–226, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003003.

Anderson:2011:PGD

[AWK^{+}11] W. Kyle Anderson, Li Wang, Sagar Kapadia, Craig Tanis, and Bruce Hilbert. Petrov–Galerkin and discontinuous-Galerkin methods for time-domain and frequency-domain electromagnetic simulations. Journal of Computational Physics, 230(23):8360–8385, September 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003986.

Bazilevs:2010:LES

[BA10] Y. Bazilevs and I. Akkerman. Large eddy simulation of turbulent Taylor–Couette flow using isogeometric analysis and the residual-based variational multiscale method. Journal of Computational Physics, 229(9):3402–3414, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000239.

Bazilevs:2013:IAL

[BAB^{+}13] Y. Bazilevs, I. Akkerman, D. J. Benson, G. Scovazzi, and M. J. Shashkov. Isogeometric analysis of Lagrangian hydrodynamics. Journal of Computational Physics, 243(??):224–243, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300140X.

Boubendir:2012:QON

[BAG12] Y. Boubendir, X. Antoine, and C. Geuzaine. A quasioptimal non-overlapping domain decomposition algorithm for the Helmholtz equation. Journal of Computational Physics, 231(2):262–280, January 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004797.

Balsara:2010:MHR

[Bal10] Dinshaw S. Balsara. Multidimensional HLLE Riemann solver: Application to Euler and magnetohydrodynamic flows. Journal of Computational Physics, 229(6):1970–1993, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006378.

Balsara:2012:SAP

[Bal12a] Dinshaw S. Balsara. Self-adjusting, positivity preserving high order schemes for hydrodynamics and magnetohydrodynamics. Journal of Computational Physics, 231(22):7504– 7517, September 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112000629.

Balsara:2012:TDH

[Bal12b] Dinshaw S. Balsara. A two-dimensional HLLC Riemann solver for conservation laws: Application to Euler and magnetohydrodynamic flows. Journal of Computational Physics, 231(22):7476–7503, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007467.

Balsara:2014:MRP

[Bal14] Dinshaw S. Balsara. Multidimensional Riemann problem with self-similar internal structure. Part I — application to hyperbolic conservation laws on structured meshes. Journal of Computational Physics, 277(??):163–200, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005464.

Barisik:2012:BTE

[BB12a] Murat Barisik and Ali Beskok. Boundary treatment effects on molecular dynamics simulations of interface thermal resistance. Journal of Computational Physics, 231(23):7881–7892, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200410X.

Blumenthal:2012:SAR

[BB12b] Frederick Blumenthal and Heiko Bauke. A stability analysis of a real space split operator method for the Klein–Gordon equation. Journal of Computational Physics, 231(2):454–464, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005444.

Baumler:2013:SPM

[BB13] Kathrin Bäumler and Eberhard Bänsch. A subspace projection method for the implementation of interface conditions in a single-drop flow problem. Journal of Computational Physics, 252(??):438–457, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004543.

Bezzola:2014:EEF

[BBAP14] Andri Bezzola, Benjamin B. Bales, Richard C. Alkire, and Linda R. Petzold. An exact and efficient first passage time algorithm for reaction-diffusion processes on a 2D-lattice. Journal of Computational Physics, 256(??):183–197, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005974.

Bassi:2012:FAB

[BBC^{+}12] F. Bassi, L. Botti, A. Colombo, D. A. Di Pietro, and P. Tesini. On the flexibility of agglomeration based physical space discontinuous Galerkin discretizations. Journal of Computational Physics, 231(1):45–65, January 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005055.

Bensiali:2013:CDI

[BBC^{+}13] Bouchra Bensiali, Kowsik Bodi, Guido Ciraolo, Philippe Ghendrih, and Jacques Liandrat. Comparison of different interpolation operators including nonlinear subdivision schemes in the simulation of particle trajectories. Journal of Computational Physics, 236(??):346–366, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200705X.

Bataillon:2012:NMS

[BBCH^{+}12] C. Bataillon, F. Bouchon, C. Chainais-Hillairet, J. Fuhrmann, E. Hoarau, and R. Touzani. Numerical methods for the simulation of a corrosion model with moving oxide layer. Journal of Computational Physics, 231(18):6213–6231, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003087.

Bourguet:2011:ROM

[BBD11] Rémi Bourguet, Marianna Braza, and Alain Dervieux. Reduced-order modeling of transonic flows around an airfoil submitted to small deformations. Journal of Computational Physics, 230(1):159–184, January 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005139.

Boscheri:2014:LAW

[BBD14] Walter Boscheri, Dinshaw S. Balsara, and Michael Dumbser. Lagrangian ADER–WENO finite volume schemes on unstructured triangular meshes based on genuinely multidimensional HLL Riemann solvers. Journal of Computational Physics, 267(??):112–138, June 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001430.

Blum:2012:REP

[BBF12] J. Blum, C. Boulbe, and B. Faugeras. Reconstruction of the equilibrium of the plasma in a Tokamak and identification of the current density profile in real time. Journal of Computational Physics, 231(3):960–980, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002294.

Berndt:2011:TSH

[BBG^{+}11] Markus Berndt, Jérôme Breil, Stéphane Galera, Milan Kucharik, Pierre-Henri Maire, and Mikhail Shashkov. Twostep hybrid conservative remapping for multimaterial arbitrary Lagrangian–Eulerian methods. Journal of Computational Physics, 230(17):6664–6687, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003007.

Bhalla:2013:UMF

[BBGP13] Amneet Pal Singh Bhalla, Rahul Bale, Boyce E. Griffith, and Neelesh A. Patankar. A unified mathematical framework and an adaptive numerical method for fluid-structure interaction with rigid, deforming, and elastic bodies. Journal of Computational Physics, 250(??):446–476, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003173.

Bhalla:2014:FRI

[BBGP14] Amneet Pal Singh Bhalla, Rahul Bale, Boyce E. Griffith, and Neelesh A. Patankar. Fully resolved immersed electrohydrodynamics for particle motion, electrolocation, and selfpropulsion. Journal of Computational Physics, 256(??):88–108, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005871.

Brezzi:2014:MSP

[BBM14] F. Brezzi, A. Buffa, and G. Manzini. Mimetic scalar products of discrete differential forms. Journal of Computational Physics, 257(??):1228–1259, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300555X.

Browne:2014:FTD

[BBPC14] P. A. Browne, C. J. Budd, C. Piccolo, and M. Cullen. Fast three dimensional r-adaptive mesh redistribution. Journal of Computational Physics, 275(??):174–196, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004161.

Barker:2010:SPM

[BC10a] Andrew T. Barker and Xiao-Chuan Cai. Scalable parallel methods for monolithic coupling in fluid–structure interaction with application to blood flow modeling. Journal of Computational Physics, 229(3):642–659, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005464.

Buffard:2010:MMM

[BC10b] Thierry Buffard and Stéphane Clain. Monoslope and multislope MUSCL methods for unstructured meshes. Journal of Computational Physics, 229(10):3745–3776, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000495.

Bertoglio:2014:TRM

[BC14] Cristóbal Bertoglio and Alfonso Caiazzo. A tangential regularization method for backflow stabilization in hemodynamics. Journal of Computational Physics, 261(??):162–171, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000126.

Benoit:2012:SIT

[BCB12] J. Benoit, C. Chauvière, and P. Bonnet. Source identification in time domain electromagnetics. Journal of Computational Physics, 231(8):3446–3456, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000411.

Bernard-Champmartin:2014:LDL

[BCD14] Aude Bernard-Champmartin and Florian De Vuyst. A low diffusive Lagrange-remap scheme for the simulation of violent air-water free-surface flows. Journal of Computational Physics, 274(??):19–49, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003957.

Bukac:2013:FSI

[BCG^{+}13] Martina Bukac, Suncica Canić, Roland Glowinski, Josip Tambaca, and Annalisa Quaini. Fluid-structure interaction in blood flow capturing non-zero longitudinal structure displacement. Journal of Computational Physics, 235(??):515–541, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004901.

Baranger:2014:LRD

[BCHM14] C. Baranger, J. Claudel, N. Hérouard, and L. Mieussens. Locally refined discrete velocity grids for stationary rarefied flow simulations. Journal of Computational Physics, 257 (??):572–593, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006827.

Bayati:2011:AMR

[BCK11] Basil Bayati, Philippe Chatelain, and Petros Koumoutsakos. Adaptive mesh refinement for stochastic reaction-diffusion processes. Journal of Computational Physics, 230(1):13–26, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004900.

Bonneton:2011:SAF

[BCL^{+}11] P. Bonneton, F. Chazel, D. Lannes, F. Marche, and M. Tissier. A splitting approach for the fully nonlinear and weakly dispersive Green–Naghdi model. Journal of Computational Physics, 230(4):1479–1498, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006236.

Blanc:2013:BJM

[BCL13] Emilie Blanc, Guillaume Chiavassa, and Bruno Lombard. Biot–JKD model: Simulation of 1D transient poroelastic waves with fractional derivatives. Journal of Computational Physics, 237(??):1–20, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007346.

Blanc:2014:WSH

[BCL14] Emilie Blanc, Guillaume Chiavassa, and Bruno Lombard. Wave simulation in 2D heterogeneous transversely isotropic porous media with fractional attenuation: a Cartesian grid approach. Journal of Computational Physics, 275(??):118–142, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400480X.

Bowers:2010:NCB

[BCLR10] Abigail L. Bowers, Benjamin R. Cousins, Alexander Linke, and Leo G. Rebholz. New connections between finite element formulations of the Navier–Stokes equations. Journal of Computational Physics, 229(24):9020–9025, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004912.

Bornia:2011:PLH

[BCM^{+}11] G. Bornia, A. Cervone, S. Manservisi, R. Scardovelli, and S. Zaleski. On the properties and limitations of the height function method in two-dimensional Cartesian geometry. Journal of Computational Physics, 230(4):851–862, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006443.

Benamou:2012:SPD

[BCM12] Jean-David Benamou, Francis Collino, and Simon Marmorat. Source point discovery through high frequency asymptotic time reversal. Journal of Computational Physics, 231(14):

4643–4661, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112001532.

Badia:2013:UCS

[BCP13] Santiago Badia, Ramon Codina, and Ramon Planas. On an unconditionally convergent stabilized finite element approximation of resistive magnetohydrodynamics. Journal of Computational Physics, 234(??):399–416, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005761.

Bao:2010:ENM

[BCW10] Weizhu Bao, Yongyong Cai, and Hanquan Wang. Efficient numerical methods for computing ground states and dynamics of dipolar Bose–Einstein condensates. Journal of Computational Physics, 229(20):7874–7892, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003670.

Budd:2013:MAB

[BCW13] C. J. Budd, M. J. P. Cullen, and E. J. Walsh. Monge– Ampére based moving mesh methods for numerical weather prediction, with applications to the Eady problem. Journal of Computational Physics, 236(??):247–270, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006912.

Bao:2013:ENM

[BCZ13] Weizhu Bao, I-Liang Chern, and Yanzhi Zhang. Efficient numerical methods for computing ground states of spin-1 Bose– Einstein condensates based on their characterizations. Journal of Computational Physics, 253(??):189–208, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300466X.

Barter:2010:SCP

[BD10a] Garrett E. Barter and David L. Darmofal. Shock capturing with PDE-based artificial viscosity for DGFEM: Part I. Formulation. Journal of Computational Physics, 229(5):1810–1827, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006299.

Barton:2010:EMM

[BD10b] Philip T. Barton and Dimitris Drikakis. An Eulerian method for multi-component problems in non-linear elasticity with sliding interfaces. Journal of Computational Physics, 229(15):5518–5540, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001889.

Bao:2011:NMC

[BD11a] Weizhu Bao and Xuanchun Dong. Numerical methods for computing ground states and dynamics of nonlinear relativistic Hartree equation for boson stars. Journal of Computational Physics, 230(13):5449–5469, June 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002129.

Botti:2011:PCS

[BD11b] Lorenzo Botti and Daniele A. Di Pietro. A pressure-correction scheme for convection-dominated incompressible flows with discontinuous velocity and continuous pressure. Journal of Computational Physics, 230(3):572–585, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005486.

Bobaru:2012:PFT

[BD12]  Florin Bobaru and Monchai Duangpanya.  A peridy-

namic formulation for transient heat conduction in bodies with evolving discontinuities. Journal of Computational Physics, 231(7):2764–2785, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007297.

Bergot:2013:HOO

[BD13] Morgane Bergot and Marc Duruflé. High-order optimal edge elements for pyramids, prisms and hexahedra. Journal of Computational Physics, 232(1):189–213, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004391.

Bard:2014:SGA

[BD14a] Christopher M. Bard and John C. Dorelli. A simple GPU-accelerated two-dimensional MUSCL–Hancock solver for ideal magnetohydrodynamics. Journal of Computational Physics, 259(??):444–460, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300805X.

Boscheri:2014:DAL

[BD14b] Walter Boscheri and Michael Dumbser. A direct Arbitrary-Lagrangian–Eulerian ADER–WENO finite volume scheme on unstructured tetrahedral meshes for conservative and nonconservative hyperbolic systems in 3D. Journal of Computational Physics, 275(??):484–523, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400477X.

Bruno:2014:RCT

[BD14c] Oscar P. Bruno and Bérangère Delourme. Rapidly convergent two-dimensional quasi-periodic Green function throughout the spectrum-including Wood anomalies. Journal of Computational Physics, 262(??):262–290, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008553.

Belme:2012:TAA

[BDA12] A. Belme, A. Dervieux, and F. Alauzet. Time accurate anisotropic goal-oriented mesh adaptation for unsteady flows. Journal of Computational Physics, 231(19):6323–6348, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002276.

Balsara:2014:MHR

[BDA14] Dinshaw S. Balsara, Michael Dumbser, and Remi Abgrall. Multidimensional HLLC Riemann solver for unstructured meshes — with application to Euler and MHD flows. Journal of Computational Physics, 261(??):172–208, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300836X.

Bogey:2011:FDC

[BdCB11] Christophe Bogey, Nicolas de Cacqueray, and Christophe Bailly. Finite differences for coarse azimuthal discretization and for reduction of effective resolution near origin of cylindrical flow equations. Journal of Computational Physics, 230(4):1134–1146, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005875.

Bialecki:2010:ASC

[BdF10] B. Bialecki and J. de Frutos. ADI spectral collocation methods for parabolic problems. Journal of Computational Physics, 229(13):5182–5193, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000152X.

Berry:2013:MEF

[BDH13] J. D. Berry, M. R. Davidson, and D. J. E. Harvie. A multiphase electrokinetic flow model for electrolytes with liquid/liquid interfaces. Journal of Computational Physics, 251(??):209–222, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003665.

Brown-Dymkoski:2014:CBV

[BDKV14] Eric Brown-Dymkoski, Nurlybek Kasimov, and Oleg V. Vasilyev. A characteristic based volume penalization method for general evolution problems applied to compressible viscous flows. Journal of Computational Physics, 262(??):344–357, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000151.

Bal:2011:HMC

[BDL11] Guillaume Bal, Anthony B. Davis, and Ian Langmore. A hybrid (Monte Carlo/deterministic) approach for multidimensional radiation transport. Journal of Computational Physics, 230(20):7723–7735, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004025.

Boudesocque-Dubois:2013:AMC

[BDLGC13] Carine Boudesocque-Dubois, Virginie Lombard, Serge Gauthier, and Jean-Marie Clarisse. An adaptive multidomain Chebyshev method for nonlinear eigenvalue problems: Application to self-similar solutions of gas dynamics equations with nonlinear heat conduction. Journal of Computational Physics, 235(??):723–741, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006663.

Blanco:2013:CMT

[BDM13] Pablo J. Blanco, Simone Deparis, and A. Cristiano I. Malossi. On the continuity of mean total normal stress in geometrical multiscale cardiovascular problems. Journal of Computational Physics, 251(??):136–155, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004038.

Barton:2013:EAF

[BDMP13] P. T. Barton, R. Deiterding, D. Meiron, and D. Pullin. Eulerian adaptive finite-difference method for high-velocity impact and penetration problems. Journal of Computational Physics, 240(??):76–99, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000442.

Burger:2011:IWS

[BDMV11] Raimund Bürger, Rosa Donat, Pep Mulet, and Carlos A. Vega. On the implementation of WENO schemes for a class of polydisperse sedimentation models. Journal of Computational Physics, 230(6):2322–2344, March 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006832.

Barucq:2010:NAL

[BDT10] H. Barucq, J. Diaz, and M. Tlemcani. New absorbing layers conditions for short water waves. Journal of Computational Physics, 229(1):58–72, January 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004732.

Bernsen:2010:AJF

[BDTW10] Erik Bernsen, Henk A. Dijkstra, Jonas Thies, and Fred W. Wubs. The application of Jacobian-free Newton–Krylov methods to reduce the spin-up time of ocean general circulation models. Journal of Computational Physics, 229(21):8167–8179, October 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003980.

Beale:2012:PIM

[Bea12] J. Thomas Beale. Partially implicit motion of a sharp interface in Navier–Stokes flow. Journal of Computational Physics, 231(18):6159–6172, July 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002604.

Berenger:2011:HSN

[Bér11] Jean-Pierre Bérenger. The Huygens subgridding for the numerical solution of the Maxwell equations. Journal of Computational Physics, 230(14):5635–5659, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002075.

Berryman:2014:CEC

[Ber14] James G. Berryman. Computing elastic constants for random polycrystals of orthotropic MgSiO_{3}, related polymorphs, and CaIrO_{3} analogs. Journal of Computational Physics, 271(??):379–396, August 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113004683.

Berry:2012:EBP

[BERB^{+}12] L. A. Berry, W. Elwasif, J. M. Reynolds-Barredo, D. Samaddar, R. Sanchez, and D. E. Newman. Event-based parareal: a data-flow based implementation of parareal. Journal of Computational Physics, 231(17):5945–5954, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002586.

Berthon:2012:EWB

[BF12] Christophe Berthon and Françoise Foucher. Efficient wellbalanced hydrostatic upwind schemes for shallow-water equations. Journal of Computational Physics, 231(15):4993–5015, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001453.

Brehm:2013:NCD

[BF13] C. Brehm and H. F. Fasel. A novel concept for the design of immersed interface methods. Journal of Computational Physics, 242(??):234–267, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000715.

Balajewicz:2014:RNE

[BF14] Maciej Balajewicz and Charbel Farhat. Reduction of nonlinear embedded boundary models for problems with evolving interfaces. Journal of Computational Physics, 274(??):489–504, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004458.

Bollig:2012:SPU

[BFE12] Evan F. Bollig, Natasha Flyer, and Gordon Erlebacher. Solution to PDEs using radial basis function finite-differences (RBF-FD) on multiple GPUs. Journal of Computational Physics, 231(21):7133–7151, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003452.

Benamou:2014:NSO

[BFO14] Jean-David Benamou, Brittany D. Froese, and Adam M. Oberman. Numerical solution of the Optimal Transportation problem using the Monge–Ampère equation. Journal of Computational Physics, 260(??):107–126, March 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008140.

Balima:2013:OTR

[BFR13] O. Balima, Y. Favennec, and D. Rousse. Optical tomography reconstruction algorithm with the finite element method: an optimal approach with regularization tools. Journal of Computational Physics, 251(??):461–479, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003379.

Bermudez:2014:PHF

[BFSVC14] A. Bermúdez, J. L. Ferrín, L. Saavedra, and M. E. Vázquez-Cendón. A projection hybrid finite volume/element method for low-Mach number flows. Journal of Computational Physics, 271(??):360–378, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006438.

Barnett:2010:NIR

[BG10] Alex Barnett and Leslie Greengard. A new integral representation for quasi-periodic fields and its application to twodimensional band structure calculations. Journal of Computational Physics, 229(19):6898–6914, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002901.

Bremer:2012:NMW

[BG12] James Bremer and Zydrunas Gimbutas. A Nyström method for weakly singular integral operators on surfaces. Journal of Computational Physics, 231(14):4885–4903, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200174X.

Bremer:2013:NES

[BG13] James Bremer and Zydrunas Gimbutas. On the numerical evaluation of the singular integrals of scattering theory. Journal of Computational Physics, 251(??):327–343, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004166.

Bosson:2012:IPB

[BGBR12] Mael Bosson, Sergei Grudinin, Xavier Bouju, and Stephane Redon. Interactive physically-based structural modeling of hydrocarbon systems. Journal of Computational Physics, 231 (6):2581–2598, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111007042.

Brun:2012:LLS

[BGG12] Emmanuel Brun, Arthur Guittet, and Frédéric Gibou. A local level-set method using a hash table data structure. Journal of Computational Physics, 231(6):2528–2536, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006991.

Badia:2011:FEA

[BGGGS11] Santiago Badia, Francisco Guillén-González, and Juan Vicente Gutiérrez-Santacreu. Finite element approximation of nematic liquid crystal flows using a saddle-point structure. Journal of Computational Physics, 230(4):1686–1706, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006480.

Bauer:2013:IVM

[BGGW13] Georg Bauer, Peter Gamnitzer, Volker Gravemeier, and Wolfgang A. Wall. An isogeometric variational multiscale method for large-eddy simulation of coupled multiion transport in turbulent flow. Journal of Computational Physics, 251(??):194–208, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003689.

Becache:2010:HOA

[BGH10] Eliane Bécache, Dan Givoli, and Thomas Hagstrom. Highorder Absorbing Boundary Conditions for anisotropic and convective wave equations. Journal of Computational Physics, 229(4):1099–1129, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005579.

Birken:2013:PMD

[BGHM13] Philipp Birken, Gregor Gassner, Mark Haas, and Claus-Dieter Munz. Preconditioning for modal discontinuous Galerkin methods for unsteady 3D Navier–Stokes equations. Journal of Computational Physics, 240(??):20–35, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000284.

Blesgen:2012:AED

[BGK12] T. Blesgen, V. Gavini, and V. Khoromskaia. Approximation of the electron density of aluminium clusters in tensor-product format. Journal of Computational Physics, 231(6):2551–2564, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111007078.

Branicki:2012:FST

[BGM12] M. Branicki, B. Gershgorin, and A. J. Majda. Filtering skill for turbulent signals for a suite of nonlinear and linear extended Kalman filters. Journal of Computational Physics, 231(4):1462–1498, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006292.

Barrett:2010:SPF

[BGN10] John W. Barrett, Harald Garcke, and Robert Nürnberg. On stable parametric finite element methods for the Stefan problem and the Mullins–Sekerka problem with applications to dendritic growth. Journal of Computational Physics, 229(18):6270–6299, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002263.

Bassi:2011:ORK

[BGR11] F. Bassi, A. Ghidoni, and S. Rebay. Optimal Runge– Kutta smoothers for the p-multigrid discontinuous Galerkin solution of the 1D Euler equations. Journal of Computational Physics, 230(11):4153–4175, May 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002068.

Bedorf:2012:SOG

[BGZ12] Jeroen Bédorf, Evghenii Gaburov, and Simon Portegies Zwart. A sparse octree gravitational N-body code that runs entirely on the GPU processor. Journal of Computational Physics, 231(7):2825–2839, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007364.

Banks:2012:USW

[BH12] Jeffrey W. Banks and William D. Henshaw. Upwind schemes for the wave equation in second-order form. Journal of Computational Physics, 231(17):5854–5889, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002367.

Bakunova:2013:OFC

[BH13a] Eugenia S. Bakunova and John Harlim. Optimal filtering of complex turbulent systems with memory depth through consistency constraints. Journal of Computational Physics, 237(??):320–343, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007085.

Baumgartner:2013:OLF

[BH13b] Stefan Baumgartner and Clemens Heitzinger. A onelevel FETI method for the drift-diffusion-Poisson system with discontinuities at an interface. Journal of Computational Physics, 243(??):74–86, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001708.

Brown:2013:AIS

[BH13c] Kristen A. Brown and John Harlim. Assimilating irregularly spaced sparsely observed turbulent signals with hierarchical Bayesian reduced stochastic filters. Journal of Computational Physics, 235(??):143–160, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006651.

Babkovskaia:2011:HOP

[BHB11] N. Babkovskaia, N. E. L. Haugen, and A. Brandenburg. A high-order public domain code for direct numerical simulations of turbulent combustion. Journal of Computational Physics, 230(1):1–12, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004754.

Bu:2010:ESA

[BHBM10] Sunyoung Bu, Jingfang Huang, Treavor H. Boyer, and Cass T. Miller. An evaluation of solution algorithms and numerical approximation methods for modeling an ion exchange process. Journal of Computational Physics, 229(13):4996–5010, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001312.

Blaha:2010:IDA

[BHK^{+}10] P. Blaha, H. Hofstätter, O. Koch, R. Laskowski, and K. Schwarz. Iterative diagonalization in augmented plane wave based methods in electronic structure calculations. Journal of Computational Physics, 229(2):453–460, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005282.

Bao:2012:AFE

[BHL12a] Gang Bao, Guanghui Hu, and Di Liu. An h-adaptive finite element solver for the calculations of the electronic structures. Journal of Computational Physics, 231(14):4967–4979, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001738.

Brady:2012:CVF

[BHL12b] P. T. Brady, M. Herrmann, and J. M. Lopez. Code verification for finite volume multiphase scalar equations using the method of manufactured solutions. Journal of Computational Physics, 231(7):2924–2944, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000083.

Baskaran:2013:ESE

[BHL^{+}13] Arvind Baskaran, Zhengzheng Hu, John S. Lowengrub, Cheng Wang, Steven M. Wise, and Peng Zhou. Energy stable and efficient finite-difference nonlinear multigrid schemes for the modified phase field crystal equation. Journal of Computational Physics, 250(??):270–292, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002891.

Barlett:2011:CBF

[BHM11] V. Ruiz Barlett, M. Hoyuelos, and H. O. Mártin. Comparison between fixed and Gaussian steplength in Monte Carlo simulations for diffusion processes. Journal of Computational Physics, 230(10):3719–3726, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000660.

Barlett:2013:MCS

[BHM13] V. Ruiz Barlett, M. Hoyuelos, and H. O. Mártin. Monte Carlo simulation with fixed steplength for diffusion processes in nonhomogeneous media. Journal of Computational Physics, 239(??):51–56, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000028.

Brinkman:2014:CFD

[BHM14] D. Brinkman, C. Heitzinger, and P. A. Markowich. A convergent 2D finite-difference scheme for the Dirac–Poisson system and the simulation of graphene. Journal of Computational Physics, 257(??):318–332, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006700.

Banks:2012:DCG

[BHS12] Jeffrey W. Banks, William D. Henshaw, and Donald W. Schwendeman. Deforming composite grids for solving fluid structure problems. Journal of Computational Physics, 231 (9):3518–3547, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111007558.

Banks:2013:SFA

[BHS13] J. W. Banks, W. D. Henshaw, and B. Sjögreen. A stable FSI algorithm for light rigid bodies in compressible flow. Journal of Computational Physics, 245(??):399–430, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300185X.

Banks:2014:ANSa

[BHS14a] J. W. Banks, W. D. Henshaw, and D. W. Schwendeman. An analysis of a new stable partitioned algorithm for FSI problems. Part I: Incompressible flow and elastic solids. Journal of Computational Physics, 269(??):108–137, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400179X.

Banks:2014:ANSb

[BHS14b] J. W. Banks, W. D. Henshaw, and D. W. Schwendeman. An analysis of a new stable partitioned algorithm for FSI problems. Part II: Incompressible flow and structural shells. Journal of Computational Physics, 268(??):399–416, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001764.

Barlow:2014:COF

[BHS14c] Andrew Barlow, Ryan Hill, and Mikhail Shashkov. Constrained optimization framework for interface-aware sub-scale dynamics closure model for multimaterial cells in Lagrangian and arbitrary Lagrangian–Eulerian hydrodynamics. Journal of Computational Physics, 276(??):92–135, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005245.

Brunner:2014:ABC

[BHY14] Hermann Brunner, Houde Han, and Dongsheng Yin. Artificial boundary conditions and finite difference approximations for a time-fractional diffusion-wave equation on a twodimensional unbounded spatial domain. Journal of Computational Physics, 276(??):541–562, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005385.

Bergmann:2011:MSF

[BI11] M. Bergmann and A. Iollo. Modeling and simulation of fish-like swimming. Journal of Computational Physics, 230(2):329–348, January 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005115.

Beck:2014:MLM

[BILM14] A. Beck, M. E. Innocenti, G. Lapenta, and S. Markidis. Multi-level multi-domain algorithm implementation for twodimensional multiscale particle in cell simulations. Journal of Computational Physics, 271(??):430–443, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008152.

Butler:2012:RSS

[BJ12] T. Butler and M. Juntunen. Reparameterization for statistical state estimation applied to differential equations. Journal of Computational Physics, 231(6):2641–2654, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007315.

Bonnetier:2014:NSC

[BJLR14] Eric Bonnetier, Lukás Jakabcin, Stéphane Labbé, and Anne Replumaz. Numerical simulation of a class of models that combine several mechanisms of dissipation: Fracture, plasticity, viscous dissipation. Journal of Computational Physics, 271(??):397–414, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002319.

Baek:2011:SIL

[BK11] Hyoungsu Baek and George Em Karniadakis. Sub-iteration leads to accuracy and stability enhancements of semi-implicit schemes for the Navier–Stokes equations. Journal of Computational Physics, 230(12):4384–4402, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000295.

Baek:2012:CSN

[BK12a] Hyoungsu Baek and George Em Karniadakis. A convergence study of a new partitioned fluid-structure interaction algorithm based on fictitious mass and damping. Journal of Computational Physics, 231(2):629–652, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005687.

Bilionis:2012:FEC

[BK12b] I. Bilionis and P. S. Koutsourelakis. Free energy computations by minimization of Kullback–Leibler divergence: An efficient adaptive biasing potential method for sparse representations. Journal of Computational Physics, 231(9):3849–3870, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000630.

Banjai:2014:FCQ

[BK14a] L. Banjai and M. Kachanovska. Fast convolution quadrature for the wave equation in three dimensions. Journal of Computational Physics, 279(??):103–126, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006251.

Borsche:2014:ASH

[BK14b] Raul Borsche and Jochen Kall. ADER schemes and high order coupling on networks of hyperbolic conservation laws. Journal of Computational Physics, 273(??):658–670, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004070.

Boiarkine:2011:PPA

[BKC^{+}11] Oleg Boiarkine, Dmitri Kuzmin, Suncica Canić, Giovanna Guidoboni, and Andro Mikelić. A positivity-preserving ALE finite element scheme for convection-diffusion equations in moving domains. Journal of Computational Physics, 230 (8):2896–2914, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111000040.

Borchardt:2012:FSG

[BKH12] M. Borchardt, R. Kleiber, and W. Hackbusch. A fast solver for the gyrokinetic field equation with adiabatic electrons. Journal of Computational Physics, 231(18):6207–6212, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003002.

Bhattacharya:2014:EMM

[BKKD14] Anirban Bhattacharya, Apoorva Kiran, Shyamprasad Karagadde, and Pradip Dutta. An enthalpy method for modeling eutectic solidification. Journal of Computational Physics, 262(??):217–230, April 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400028X.

Burdakov:2012:MRA

[BKV12] Oleg Burdakov, Ivan Kapyrin, and Yuri Vassilevski. Monotonicity recovering and accuracy preserving optimization methods for postprocessing finite element solutions. Journal of Computational Physics, 231(8):3126–3142, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000095.

Braumann:2010:NSS

[BKW10] Andreas Braumann, Markus Kraft, and Wolfgang Wagner. Numerical study of a stochastic particle algorithm solving a multidimensional population balance model for high shear granulation. Journal of Computational Physics, 229(20):7672–7691, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003360.

Bruno:2010:HOU

[BL10] Oscar P. Bruno and Mark Lyon. High-order unconditionally stable FC–AD solvers for general smooth domains I. Basic elements. Journal of Computational Physics, 229(6):2009–2033, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006391.

Buffat:2011:SFD

[BL11] Marc Buffat and Lionel Le Penven. A spectral fictitious domain method with internal forcing for solving elliptic PDEs. Journal of Computational Physics, 230(7):2433–2450, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006686.

Bruno:2013:HOI

[BL13] Oscar P. Bruno and Stéphane K. Lintner. A high-order integral solver for scalar problems of diffraction by screens and apertures in three-dimensional space. Journal of Computational Physics, 252(??):250–274, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004464.

Bazilevs:2014:IAL

[BLA^{+}14] Y. Bazilevs, C. C. Long, I. Akkerman, D. J. Benson, and M. J. Shashkov. Isogeometric analysis of Lagrangian hydrodynamics: Axisymmetric formulation in the rz-cylindrical coordinates. Journal of Computational Physics, 262(??):244–261, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000205.

Boedec:2011:VDS

[BLJ11] G. Boedec, M. Leonetti, and M. Jaeger. 3D vesicle dynamics simulations with a linearly triangulated surface. Journal of Computational Physics, 230(4):1020–1034, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005772.

Berre:2011:ITD

[BLM11] Inga Berre, Martha Lien, and Trond Mannseth. Identification of three-dimensional electric conductivity changes from time-lapse electromagnetic observations. Journal of Computational Physics, 230(10):3915–3928, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001045.

Baudron:2014:PTN

[BLM^{+}14] Anne-Marie Baudron, Jean-Jacques Lautard, Yvon Maday, Mohamed Kamel Riahi, and Julien Salomon. Parareal in time 3D numerical solver for the LWR Benchmark neutron diffusion transient model. Journal of Computational Physics, 279(??):67–79, December 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006056.

Barcarolo:2014:APR

[BLOdV14] D. A. Barcarolo, D. Le Touzé, G. Oger, and F. de Vuyst. Adaptive particle refinement and derefinement applied to the smoothed particle hydrodynamics method. Journal of Computational Physics, 273(??):640–657, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004033.

Bao:2014:FMG

[BLQ14] Gang Bao, Jun Lai, and Jianliang Qian. Fast multiscale Gaussian beam methods for wave equations in bounded convex domains. Journal of Computational Physics, 261(??):36–64, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113008425.

Borg:2013:FSA

[BLR13a] Matthew K. Borg, Duncan A. Lockerby, and Jason M. Reese. Fluid simulations with atomistic resolution: a hybrid multiscale method with field-wise coupling. Journal of Computational Physics, 255(??):149–165, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005603.

Borg:2013:MMM

[BLR13b] Matthew K. Borg, Duncan A. Lockerby, and Jason M. Reese. A multiscale method for micro/nano flows of high aspect ratio. Journal of Computational Physics, 233(??):400–413, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005402.

Bu:2014:FHO

[BLR14]  Fanbin Bu, Junshan Lin, and Fernando Reitich.  A

fast and high-order method for the three-dimensional elastic wave scattering problem. Journal of Computational Physics, 258(??):856–870, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113007596.

Barrault:2011:EPR

[BLRR11] M. Barrault, B. Lathuilière, P. Ramet, and J. Roman. Efficient parallel resolution of the simplified transport equations in mixed-dual formulation. Journal of Computational Physics, 230(5):2004–2020, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006625.

Borrell:2011:PDP

[BLTO11] R. Borrell, O. Lehmkuhl, F. X. Trias, and A. Oliva. Parallel direct Poisson solver for discretisations with one Fourier diagonalisable direction. Journal of Computational Physics, 230(12):4723–4741, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100132X.

Brunner:2010:NSF

[BLY10] Hermann Brunner, Leevan Ling, and Masahiro Yamamoto. Numerical simulations of 2D fractional subdiffusion problems. Journal of Computational Physics, 229(18):6613–6622, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002767.

Borgers:2010:AMM

[BM10a] Christoph Börgers and Scott MacLachlan. An angular multigrid method for computing mono-energetic particle beams in Flatland. Journal of Computational Physics, 229(8):2914–2931, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109007025.

Bouzarth:2010:MTI

[BM10b] Elizabeth L. Bouzarth and Michael L. Minion. A multirate time integrator for regularized Stokeslets. Journal of Computational Physics, 229(11):4208–4224, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000793.

Bouzarth:2011:MSB

[BM11] Elizabeth L. Bouzarth and Michael L. Minion. Modeling slender bodies with the method of regularized Stokeslets. Journal of Computational Physics, 230(10):3929–3947, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001069.

Bottauscio:2013:CMM

[BM13a]  Oriano Bottauscio and Alessandra Manzin.  Compari-

son of multiscale models for eddy current computation in granular magnetic materials. Journal of Computational Physics, 253(??):1–17, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004671.

Branicki:2013:DSS

[BM13b] M. Branicki and A. J. Majda. Dynamic Stochastic Superresolution of sparsely observed turbulent systems. Journal of Computational Physics, 241(??):333–363, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007176.

Brull:2014:LDV

[BM14] S. Brull and L. Mieussens. Local discrete velocity grids for deterministic rarefied flow simulations. Journal of Computational Physics, 266(??):22–46, June 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001016.

Bayona:2010:RFF

[BMCK10] Victor Bayona, Miguel Moscoso, Manuel Carretero, and Manuel Kindelan. RBF–FD formulas and convergence properties. Journal of Computational Physics, 229(22):8281–8295, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003773.

Balsara:2013:EIA

[BMD^{+}13] Dinshaw S. Balsara, Chad Meyer, Michael Dumbser, Huijing Du, and Zhiliang Xu. Efficient implementation of ADER schemes for Euler and magnetohydrodynamical flows on structured meshes — speed comparisons with Runge–Kutta methods. Journal of Computational Physics, 235(??):934–969, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002860.

Bayona:2011:OCS

[BMK11] Victor Bayona, Miguel Moscoso, and Manuel Kindelan. Optimal constant shape parameter for multiquadric based RBF– FD method. Journal of Computational Physics, 230(19): 7384–7399, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111003548.

Bayona:2012:OVS

[BMK12] Victor Bayona, Miguel Moscoso, and Manuel Kindelan. Optimal variable shape parameter for multiquadric based RBF–FD method. Journal of Computational Physics, 231(6):2466–2481, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006929.

Badia:2014:BRL

[BMP14] Santiago Badia, Alberto F. Martín, and Ramon Planas. Block recursive LU preconditioners for the thermally coupled incompressible inductionless MHD problem. Journal of Computational Physics, 274(??):562–591, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004355.

Bruss:2014:SPE

[BMR14] Don E. Bruss, Jim E. Morel, and Jean C. Ragusa. S_{2} SA preconditioning for the S_{n} equations with strictly nonnegative spatial discretization. Journal of Computational Physics, 273(??):706–719, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003684.

Balan:2012:SHO

[BMS12] Aravind Balan, Georg May, and Joachim Schöberl. A stable high-order Spectral Difference method for hyperbolic conservation laws on triangular elements. Journal of Computational Physics, 231(5):2359–2375, March 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006978.

Berg:2011:SRS

[BN11] Jens Berg and Jan Nordström. Stable Robin solid wall boundary conditions for the Navier–Stokes equations. Journal of Computational Physics, 230(19):7519–7532, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004001.

Berg:2012:SFO

[BN12] Jens Berg and Jan Nordström. Superconvergent functional output for time-dependent problems using finite differences on summation-by-parts form. Journal of Computational Physics, 231(20):6846–6860, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003476.

Berg:2013:IBC

[BN13] Jens Berg and Jan Nordström. On the impact of boundary conditions on dual consistent finite difference discretizations. Journal of Computational Physics, 236(??):41–55, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006997.

Berg:2014:DBB

[BN14a] Jens Berg and Jan Nordström. Duality based boundary conditions and dual consistent finite difference discretizations of the Navier–Stokes and Euler equations. Journal of Computational Physics, 259(??):135–153, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007936.

Bihlo:2014:CRF

[BN14b] Alexander Bihlo and Jean-Christophe Nave. Convecting reference frames and invariant numerical models. Journal of Computational Physics, 272(??):656–663, September 1, 2014.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003118.

Berry:2012:DFI

[BND^{+}12] Robert D. Berry, Habib N. Najm, Bert J. Debusschere, Youssef M. Marzouk, and Helgi Adalsteinsson. Data-free inference of the joint distribution of uncertain model parameters. Journal of Computational Physics, 231(5):2180–2198, March 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006310.

Basumatary:2014:DCB

[BNM14] Mantulal Basumatary, Ganesh Natarajan, and Subhash C. Mishra. Defect correction based velocity reconstruction for physically consistent simulations of non-Newtonian flows on unstructured grids. Journal of Computational Physics, 272 (??):227–244, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114002964.

Benzi:2011:RDF

[BNNW11] Michele Benzi, Michael Ng, Qiang Niu, and Zhen Wang. A Relaxed Dimensional Factorization preconditioner for the incompressible Navier–Stokes equations. Journal of Computational Physics, 230(16):6185–6202, July 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002257.

Bonito:2010:PFG

[BNP10] Andrea Bonito, Ricardo H. Nochetto, and M. Sebastian Pauletti. Parametric FEM for geometric biomembranes. Journal of Computational Physics, 229(9):3171–3188, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000033.

Bao:2014:MMF

[BNT14] Lei Bao, Ramachandran D. Nair, and Henry M. Tufo. A mass and momentum flux-form high-order discontinuous Galerkin shallow water model on the cubed-sphere. Journal of Computational Physics, 271(??):224–243, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300795X.

Barton:2011:CLS

[BOD11] P. T. Barton, B. Obadia, and D. Drikakis. A conservative levelset based method for compressible solid/fluid problems on fixed grids. Journal of Computational Physics, 230(21):7867–7890, September 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100413X.

Boersma:2011:OSC

[Boe11] Bendiks Jan Boersma. A 6th order staggered compact finite difference method for the incompressible Navier– Stokes and scalar transport equations. Journal of Computational Physics, 230(12):4940–4954, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001574.

Boselli:2012:MMF

[BOK12] F. Boselli, D. Obrist, and L. Kleiser. A multilayer method of fundamental solutions for Stokes flow problems. Journal of Computational Physics, 231(18):6139–6158, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002653.

Booth:2011:AMC

[Boo11] Thomas E. Booth. An alternative Monte Carlo approach to the thermal radiative transfer problem. Journal of Computational Physics, 230(4):1516–1527, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006339.

Boyd:2010:UFG

[Boy10] John P. Boyd. The uselessness of the Fast Gauss Transform for summing Gaussian radial basis function series. Journal of Computational Physics, 229(4):1311–1326, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005841.

Boyd:2011:NEF

[Boy11] John P. Boyd. The near-equivalence of five species of spectrally-accurate radial basis functions (RBFs): Asymptotic approximations to the RBF cardinal functions on a uniform, unbounded grid. Journal of Computational Physics, 230(4):1304–1318, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006054.

Boyd:2013:CCM

[Boy13] John P. Boyd. A comparison of companion matrix methods to find roots of a trigonometric polynomial. Journal of Computational Physics, 246(??):96–112, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002003.

Buoni:2010:ASE

[BP10] Matthew Buoni and Linda Petzold. An algorithm for simulation of electrochemical systems with surface–bulk coupling strategies. Journal of Computational Physics, 229(2):379–398, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005245.

Bobylev:2013:MCM

[BP13a] A. V. Bobylev and I. F. Potapenko. Monte Carlo methods and their analysis for Coulomb collisions in multicomponent plasmas. Journal of Computational Physics, 246(??):123–144, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002027.

Bukshtynov:2013:ORM

[BP13b] Vladislav Bukshtynov and Bartosz Protas. Optimal reconstruction of material properties in complex multiphysics phenomena. Journal of Computational Physics, 242(??):889–914, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001538.

Breinlinger:2013:STW

[BPHK13] Thomas Breinlinger, Pit Polfer, Adham Hashibon, and Torsten Kraft. Surface tension and wetting effects with smoothed particle hydrodynamics. Journal of Computational Physics, 243(??):14–27, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001575.

Bakr:2013:DDF

[BPM13] Shaaban A. Bakr, David Pardo, and Trond Mannseth. Domain decomposition Fourier finite element method for the simulation of 3D marine CSEM measurements. Journal of Computational Physics, 255(??):456–470, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005858.

Blanco-Pillado:2012:NPS

[BPOS12] Jose J. Blanco-Pillado, Ken D. Olum, and Benjamin Shlaer. A new parallel simulation technique. Journal of Computational Physics, 231(1):98–108, January 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005171.

Bernardini:2013:TCF

[BPQO13] M. Bernardini, S. Pirozzoli, M. Quadrio, and P. Orlandi. Turbulent channel flow simulations in convecting reference frames. Journal of Computational Physics, 232(1):1–6, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004408.

Berrone:2014:OAL

[BPS14] Stefano Berrone, Sandra Pieraccini, and Stefano Scialò. An optimization approach for large scale simulations of discrete fracture network flows. Journal of Computational Physics, 256(??):838–853, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006426.

Barba:2010:GFI

[BR10a] L. A. Barba and Louis F. Rossi. Global field interpolation for particle methods. Journal of Computational Physics, 229(4):1292–1310, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900583X.

Bremer:2010:EDL

[BR10b] James Bremer and Vladimir Rokhlin. Efficient discretization of Laplace boundary integral equations on polygonal domains. Journal of Computational Physics, 229(7):2507–2525, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006718.

Billet:2011:RKD

[BR11] G. Billet and J. Ryan. A Runge–Kutta discontinuous Galerkin approach to solve reactive flows: The hyperbolic operator. Journal of Computational Physics, 230(4):1064–1083, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005814.

Bjontegaard:2012:AIT

[BR12a] Tormod Bjøntegaard and Einar M. Rønquist. Accurate interface-tracking of surfaces in three dimensions for arbitrary Lagrangian–Eulerian schemes. Journal of Computational Physics, 231(19):6514–6531, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003178.

Borrel:2012:EDG

[BR12b] M. Borrel and J. Ryan. The Elastoplast Discontinuous Galerkin (EDG) method for the Navier–Stokes equations. Journal of Computational Physics, 231(1):1–22, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004773.

Bidadi:2014:QND

[BR14] Shreyas Bidadi and Sarma L. Rani. Quantification of numerical diffusivity due to TVD schemes in the advection equation. Journal of Computational Physics, 261(??):65–82, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008103.

Bremer:2012:FDS

[Bre12a] James Bremer. A fast direct solver for the integral equations of scattering theory on planar curves with corners. Journal of Computational Physics, 231(4):1879–1899, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100670X.

Breugem:2012:SOA

[Bre12b] Wim-Paul Breugem. A second-order accurate immersed boundary method for fully resolved simulations of particle-laden flows. Journal of Computational Physics, 231(13):4469–4498, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001374.

Bochev:2014:OBR

[BRP14] Pavel Bochev, Denis Ridzal, and Kara Peterson. Optimizationbased remap and transport: a divide and conquer strategy for feature-preserving discretizations. Journal of Computational Physics, 257(??):1113–1139, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002416.

Bremer:2010:UQB

[BRS10] James Bremer, Vladimir Rokhlin, and Ian Sammis. Universal quadratures for boundary integral equations on twodimensional domains with corners. Journal of Computational Physics, 229(22):8259–8280, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003633.

Bhumkar:2011:LFM

[BRS11] Yogesh G. Bhumkar, Manoj K. Rajpoot, and Tapan K. Sengupta. A linear focusing mechanism for dispersive and non-dispersive wave problems. Journal of Computational Physics, 230(4):1652–1675, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006418.

Bisi:2012:NSG

[BRS12] M. Bisi, S. Rjasanow, and G. Spiga. Numerical studies of a granular gas in a host medium. Journal of Computational Physics, 231(4):1339–1359, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006085.

Bochev:2013:FOB

[BRS13] Pavel Bochev, Denis Ridzal, and Mikhail Shashkov. Fast optimization-based conservative remap of scalar fields through aggregate mass transfer. Journal of Computational Physics, 246(??):37–57, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113002180.

Bochev:2011:FAN

[BRSS11] Pavel Bochev, Denis Ridzal, Guglielmo Scovazzi, and Mikhail Shashkov. Formulation, analysis and numerical study of an optimization-based conservative interpolation (remap) of scalar fields for arbitrary Lagrangian–Eulerian methods. Journal of Computational Physics, 230(13):5199–5225, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001604.

Brugmann:2013:PMM

[Brü13] Bernd Brügmann. A pseudospectral matrix method for time-dependent tensor fields on a spherical shell. Journal of Computational Physics, 235(??):216–240, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006833.

Bou-Rabee:2012:PIU

[BRVE12] Nawaf Bou-Rabee and Eric Vanden-Eijnden. A patch that imparts unconditional stability to explicit integrators for Langevin-like equations. Journal of Computational Physics, 231(6):2565–2580, March 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007054.

Birk:2010:LHO

[BS10a] C. Birk and Ch. Song. A local high-order doubly asymptotic open boundary for diffusion in a semi-infinite layer. Journal of Computational Physics, 229(17):6156–6179, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002408.

Booty:2010:HNM

[BS10b] M. R. Booty and M. Siegel. A hybrid numerical method for interfacial fluid flow with soluble surfactant. Journal of Computational Physics, 229(10):3864–3883, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000550.

Bertaccini:2011:FNS

[BS11a] Daniele Bertaccini and Renata Sisto. Fast numerical solution of nonlinear nonlocal cochlear models. Journal of Computational Physics, 230(7):2575–2587, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007096.

Blatman:2011:ASP

[BS11b] Géraud Blatman and Bruno Sudret. Adaptive sparse polynomial chaos expansion based on least angle regression. Journal of Computational Physics, 230(6):2345–2367, March 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006856.

Bettini:2014:CSH

[BS14a] Paolo Bettini and Ruben Specogna. Computation of stationary 3D halo currents in fusion devices with accuracy control. Journal of Computational Physics, 273(??):100–117, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003404.

Beylkin:2014:OSU

[BS14b] G. Beylkin and K. Sandberg. ODE solvers using bandlimited approximations. Journal of Computational Physics, 265(??):156–171, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114000953.

Bosch:2014:FSC

[BSB14] Jessica Bosch, Martin Stoll, and Peter Benner. Fast solution of Cahn–Hilliard variational inequalities using implicit time discretization and finite elements. Journal of Computational Physics, 262(??):38–57, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000084.

Blaise:2013:DGU

[BSCML13] Sébastien Blaise, Amik St-Cyr, Dimitri Mavriplis, and Brian Lockwood. Discontinuous Galerkin unsteady discrete adjoint method for real-time efficient tsunami simulations. Journal of Computational Physics, 232(1):416–430, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004688.

Bohn:2012:VLB

[BSDM12] C. D. Bohn, S. A. Scott, J. S. Dennis, and C. R. Müller. Validation of a lattice Boltzmann model for gassolid reactions with experiments. Journal of Computational Physics, 231(16):5334–5350, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002021.

Brandsmeier:2011:MHF

[BSS11] Holger Brandsmeier, Kersten Schmidt, and Christoph Schwab. A multiscale hp-FEM for 2D photonic crystal bands. Journal of Computational Physics, 230(2):349–374, January 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005127.

Blank:2012:PAC

[BSS12] Luise Blank, Lavinia Sarbu, and Martin Stoll. Preconditioning for Allen–Cahn variational inequalities with nonlocal constraints. Journal of Computational Physics, 231 (16):5406–5420, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112002173.

Bhumkar:2014:DRP

[BSS14] Yogesh G. Bhumkar, Tony W. H. Sheu, and Tapan K. Sengupta. A dispersion relation preserving optimized upwind compact difference scheme for high accuracy flow simulations. Journal of Computational Physics, 278(??):378–399, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006081.

Bao:2012:FEM

[BSSW12] Kai Bao, Yi Shi, Shuyu Sun, and Xiao-Ping Wang. A finite element method for the numerical solution of the coupled Cahn–Hilliard and Navier–Stokes system for moving contact line problems. Journal of Computational Physics, 231(24):8083–8099, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004111.

Buffa:2014:IMC

[BSV14]  A. Buffa, G. Sangalli, and R. Vázquez.  Isogeomet-

ric methods for computational electromagnetics: B-spline and T-spline discretizations. Journal of Computational Physics, 257(??):1291–1320, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005536.

Bao:2013:NMC

[BTX13] Weizhu Bao, Qinglin Tang, and Zhiguo Xu. Numerical methods and comparison for computing dark and bright solitons in the nonlinear Schrödinger equation. Journal of Computational Physics, 235(??):423–445, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006602.

Bu:2014:GFE

[BTY14] Weiping Bu, Yifa Tang, and Jiye Yang. Galerkin finite element method for two-dimensional Riesz space fractional diffusion equations. Journal of Computational Physics, 276(??):26–38, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005166.

Bin:2011:AMR

[BUH11] Jonghoon Bin, Ali Uzun, and M. Yousuff Hussaini. Adaptive mesh redistribution method for domains with complex boundaries. Journal of Computational Physics, 230(8):3178–3204, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000398.

Bidasyuk:2013:ICS

[BV13] Y. Bidasyuk and W. Vanroose. Improved convergence of scattering calculations in the oscillator representation. Journal of Computational Physics, 234(??):60–78, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005499.

Bockmann:2014:GAL

[BV14] Arne Bøckmann and Magnus Vartdal. A gradient augmented level set method for unstructured grids. Journal of Computational Physics, 258(??):47–72, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007006.

Bedrossian:2010:SOV

[BvBZ^{+}10] Jacob Bedrossian, James H. von Brecht, Siwei Zhu, Eftychios Sifakis, and Joseph M. Teran. A second order virtual node method for elliptic problems with interfaces and irregular domains. Journal of Computational Physics, 229(18):6405–6426, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002548.

Badreddine:2014:SQP

[BVM14] Hassan Badreddine, Stefan Vandewalle, and Johan Meyers. Sequential Quadratic Programming (SQP) for optimal control in direct numerical simulation of turbulent flow. Journal of Computational Physics, 256(??):1–16, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005883.

Basting:2013:HLS

[BW13] Steffen Basting and Martin Weismann. A hybrid level setfront tracking finite element approach for fluid-structure interaction and two-phase flow applications. Journal of Computational Physics, 255(??):228–244, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005561.

Blonigan:2014:PDA

[BW14] Patrick J. Blonigan and Qiqi Wang. Probability density adjoint for sensitivity analysis of the Mean of Chaos. Journal of Computational Physics, 270(??):660–686, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002903.

Bauer:2011:SOE

[BWC11] Carl A. Bauer, Gregory R. Werner, and John R. Cary. A second-order 3D electromagnetics algorithm for curved interfaces between anisotropic dielectrics on a Yee mesh. Journal of Computational Physics, 230(5):2060–2075, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006698.

Bauer:2013:FMB

[BWC13] Carl A. Bauer, Gregory R. Werner, and John R. Cary. A fast multigrid-based electromagnetic eigensolver for curved metal boundaries on the Yee mesh. Journal of Computational Physics, 251(??):524–534, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004154.

Bond:2014:SDM

[BWMG14] D. M. Bond, V. Wheatley, M. N. Macrossan, and M. Goldsworthy. Solving the discrete S-model kinetic equations with arbitrary order polynomial approximations. Journal of Computational Physics, 259(??):175–198, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007882.

Barham:2010:FEM

[BWS10] Matthew I. Barham, Daniel A. White, and David J. Steigmann. Finite element modeling of the deformation of magnetoelastic film. Journal of Computational Physics, 229(18):6193–6207, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000183X.

Boyd:2012:NPC

[BX12] John P. Boyd and Zhengjie Xu. Numerical and perturbative computations of solitary waves of the Benjamin– Ono equation with higher order nonlinearity using Christov rational basis functions. Journal of Computational Physics, 231(4):1216–1229, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005870.

Boyd:2011:CSS

[BY11] John P. Boyd and Fu Yu. Comparing seven spectral methods for interpolation and for solving the Poisson equation in a disk: Zernike polynomials, Logan–Shepp ridge polynomials, Chebyshev–Fourier series, cylindrical Robert functions, Bessel–Fourier expansions, square-to-disk conformal mapping and radial basis functions. Journal of Computational Physics, 230(4):1408–1438, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006133.

Bilyeu:2014:TDF

[BYCC14] David L. Bilyeu, S.-T. John Yu, Yung-Yu Chen, and Jean-Luc Cambier. A two-dimensional fourth-order unstructuredmeshed Euler solver based on the CESE method. Journal of Computational Physics, 257(??):981–999, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300658X.

Bilionis:2012:MOL

[BZ12] Ilias Bilionis and Nicholas Zabaras. Multi-output local Gaussian process regression: Applications to uncertainty quantification. Journal of Computational Physics, 231(17):5718–5746, July 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002513.

Bilionis:2013:MOS

[BZKL13] Ilias Bilionis, Nicholas Zabaras, Bledar A. Konomi, and Guang Lin. Multi-output separable Gaussian process: Towards an efficient, fully Bayesian paradigm for uncertainty quantification. Journal of Computational Physics, 241(??):212–239, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000417.

Chen:2014:HHD

[CAH14] Xi (Ronald) Chen, Daniel Appelö, and Thomas Hagstrom. A hybrid hermite-discontinuous Galerkin method for hyperbolic systems with application to Maxwell’s equations. Journal of Computational Physics, 257(??):501–520, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006608.

Cameron:2013:ERF

[Cam13] M. K. Cameron. Estimation of reactive fluxes in gradient stochastic systems using an analogy with electric circuits. Journal of Computational Physics, 247(??):137–152, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002386.

Capdeville:2011:HOM

[Cap11] G. Capdeville. A high-order multi-dimensional HLL–Riemann solver for non-linear Euler equations. Journal of Computational Physics, 230(8):2915–2951, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000052.

Cheny:2010:LSM

[CB10] Yoann Cheny and Olivier Botella. The LS–STAG method: a new immersed boundary/level-set method for the computation of incompressible viscous flows in complex moving geometries with good conservation properties. Journal of Computational Physics, 229(4):1043–1076, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900552X.

Cardoso:2011:LGT

[CB11] Nuno Cardoso and Pedro Bicudo. SU(2) lattice gauge theory simulations on Fermi GPUs. Journal of Computational Physics, 230(10):3998–4010, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001124.

Chaillat:2012:FFA

[CB12] Stéphanie Chaillat and George Biros. FaIMS: a fast algorithm for the inverse medium problem with multiple frequencies and multiple sources for the scalar Helmholtz equation. Journal of Computational Physics, 231(12):4403–4421, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200085X.

Coulette:2013:NCG

[CB13] David Coulette and Nicolas Besse. Numerical comparisons of gyrokinetic multi-water-bag models. Journal of Computational Physics, 248(??):1–32, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002532.

Chaillat:2014:NFM

[CB14] Stéphanie Chaillat and Marc Bonnet. A new Fast Multipole formulation for the elastodynamic half-space Green’s tensor. Journal of Computational Physics, 258(??):787–808, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007602.

Cowan:2011:CEM

[CBCM^{+}11] Benjamin M. Cowan, David L. Bruhwiler, Estelle Cormier-Michel, Eric Esarey, Cameron G. R. Geddes, Peter Messmer, and Kevin M. Paul. Characteristics of an envelope model for laser-plasma accelerator simulation. Journal of Computational Physics, 230(1):61–86, January 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005036.

Cleveland:2013:OIR

[CBGK13] Mathew A. Cleveland, Thomas A. Brunner, Nicholas A. Gentile, and Jeffrey A. Keasler. Obtaining identical results with double precision global accuracy on different numbers of processors in parallel particle Monte Carlo simulations. Journal of Computational Physics, 251(??):223–236, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004075.

Cauley:2012:TDD

[CBKK12] Stephen Cauley, Venkataramanan Balakrishnan, Gerhard Klimeck, and Cheng-Kok Koh. A two-dimensional domain decomposition technique for the simulation of quantum-scale devices. Journal of Computational Physics, 231(4):1293–1313, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005997.

Cai:2013:HPI

[CBPS13] Yunfeng Cai, Zhaojun Bai, John E. Pask, and N. Sukumar. Hybrid preconditioning for iterative diagonalization of ill-conditioned generalized eigenvalue problems in electronic structure calculations. Journal of Computational Physics, 255(??):16–30, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004932.

Colomer:2013:PAT

[CBTR13] G. Colomer, R. Borrell, F. X. Trias, and I. Rodríguez. Parallel algorithms for S_{n} transport sweeps on unstructured meshes. Journal of Computational Physics, 232(1):118–135, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003683.

Chen:2010:DGB

[CBW10] Zhan Chen, Nathan A. Baker, and G. W. Wei. Differential geometry based solvation model I: Eulerian formulation. Journal of Computational Physics, 229(22):8231–8258, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003517.

Cho:2012:PFA

[CC12] Min Hyung Cho and Wei Cai. A parallel fast algorithm for computing the Helmholtz integral operator in 3-D layered media. Journal of Computational Physics, 231(17):5910–5925, July 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002641.

Chen:2013:APM

[CC13] G. Chen and L. Chacón. An analytical particle mover for the charge- and energy-conserving, nonlinearly implicit, electrostatic particle-in-cell algorithm. Journal of Computational Physics, 247(??):79–87, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002507.

Coralic:2014:FVW

[CC14] Vedran Coralic and Tim Colonius. Finite-volume WENO scheme for viscous compressible multicomponent flows. Journal of Computational Physics, 274(??):95–121, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004082.

Chen:2011:ECC

[CCB11] G. Chen, L. Chacón, and D. C. Barnes. An energy- and charge-conserving, implicit, electrostatic particle-in-cell algorithm. Journal of Computational Physics, 230(18):7018–7036, August 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003421.

Chen:2012:EMP

[CCB12] G. Chen, L. Chacón, and D. C. Barnes. An efficient mixedprecision, hybrid CPU–GPU implementation of a nonlinearly implicit one-dimensional particle-in-cell algorithm. Journal of Computational Physics, 231(16):5374–5388, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002227.

Chacon:2013:CEC

[CCB13] L. Chacón, G. Chen, and D. C. Barnes. A charge- and energy-conserving implicit, electrostatic particle-in-cell algorithm on mapped computational meshes. Journal of Computational Physics, 233(??):1–9, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004275.

Chen:2012:SCM

[CCC12] Huei-Shuang Chen, Shing-Lin Chang, and Cheng-Sheng Chien. Spectral collocation methods using sine functions for a rotating Bose–Einstein condensation in optical lattices. Journal of Computational Physics, 231(4):1553–1569, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006309.

Cai:2011:NSC

[CCD11a] Tao Cai, Kwing L. Chan, and Licai Deng. Numerical simulation of core convection by a multi-layer semiimplicit spherical spectral method. Journal of Computational Physics, 230(24):8698–8712, October 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005018.

Castro:2011:HOW

[CCD11b] Marcos Castro, Bruno Costa, and Wai Sun Don. High order weighted essentially non-oscillatory WENO–Z schemes for hyperbolic conservation laws. Journal of Computational Physics, 230(5):1766–1792, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006431.

Calzada:2011:FDL

[CCFCM11] M. Carmen Calzada, Gema Camacho, Enrique Fernández-Cara, and Mercedes Marín. Fictitious domains and level sets for moving boundary problems. Applications to the numerical simulation of tumor growth. Journal of Computational Physics, 230(4):1335–1358, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006078.

Crockett:2011:CGE

[CCG11] R. K. Crockett, P. Colella, and D. T. Graves. A Cartesian grid embedded boundary method for solving the Poisson and heat equations with discontinuous coefficients in three dimensions. Journal of Computational Physics, 230(7):2451–2469, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006819.

Calgaro:2013:PPS

[CCGP13] C. Calgaro, E. Creusé, T. Goudon, and Y. Penel. Positivitypreserving schemes for Euler equations: Sharp and practical CFL conditions. Journal of Computational Physics, 234 (??):417–438, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112005852.

Chen:2013:LFF

[CCK13] Weitao Chen, Ching-Shan Chou, and Chiu-Yen Kao. Lax– Friedrichs fast sweeping methods for steady state problems for hyperbolic conservation laws. Journal of Computational Physics, 234(??):452–471, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005979.

Calgaro:2010:SVB

[CCKCG10] Caterina Calgaro, Emile Chane-Kane, Emmanuel Creusé, and Thierry Goudon. L^{∞}-stability of vertex-based MUSCL finite volume schemes on unstructured grids: Simulation of incompressible flows with high density ratios. Journal of Computational Physics, 229(17):6027–6046, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002214.

Chen:2014:FPN

[CCL^{+}14] G. Chen, L. Chacón, C. A. Leibs, D. A. Knoll, and W. Taitano. Fluid preconditioning for Newton–Krylov-based, fully implicit, electrostatic particle-in-cell simulations. Journal of Computational Physics, 258(??):555–567, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007286.

Camassa:2010:VIR

[CCLS10] Roberto Camassa, Pao-Hsiung Chiu, Long Lee, and Tony W. H. Sheu. Viscous and inviscid regularizations in a class of evolutionary partial differential equations. Journal of Computational Physics, 229(19):6676–6687, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003165.

Cortez:2010:CTD

[CCLV10] Ricardo Cortez, Bree Cummins, Karin Leiderman, and Douglas Varela. Computation of three-dimensional Brinkman flows using regularized methods. Journal of Computational Physics, 229(20):7609–7624, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000327X.

Chen:2013:NML

[CCMG^{+}13] M. Chen, E. Cormier-Michel, C. G. R. Geddes, D. L. Bruhwiler, L. L. Yu, E. Esarey, C. B. Schroeder, and W. P. Leemans. Numerical modeling of laser tunneling ionization in explicit particle-in-cell codes. Journal of Computational Physics, 236(??):220–228, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007097.

Caceres:2011:NSN

[CCT11] María J. Cáceres, José A. Carrillo, and Louis Tao. A numerical solver for a nonlinear Fokker–Planck equation representation of neuronal network dynamics. Journal of Computational Physics, 230(4):1084–1099, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005838.

Chen:2011:EGS

[CCW11] Jen-Hao Chen, I-Liang Chern, and Weichung Wang. Exploring ground states and excited states of spin-1 Bose–Einstein condensates by continuation methods. Journal of Computational Physics, 230(6):2222–2236, March 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006637.

Chen:2012:DMF

[CCW12] Zhongying Chen, Dongsheng Cheng, and Tingting Wu. A dispersion minimizing finite difference scheme and preconditioned solver for the 3D Helmholtz equation. Journal of Computational Physics, 231(24):8152–8175, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004330.

Chung:2013:CSS

[CCY13] Eric T. Chung, Patrick Ciarlet, Jr., and Tang Fei Yu. Convergence and superconvergence of staggered discontinuous Galerkin methods for the three-dimensional Maxwell’s equations on Cartesian grids. Journal of Computational Physics, 235(??):14–31, February 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112006201.

Cheng:2014:ECDa

[CCZ14a] Yingda Cheng, Andrew J. Christlieb, and Xinghui Zhong. Energy-conserving discontinuous Galerkin methods for the Vlasov–Ampère system. Journal of Computational Physics, 256(??):630–655, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006189.

Cheng:2014:ECDb

[CCZ14b] Yingda Cheng, Andrew J. Christlieb, and Xinghui Zhong. Energy-conserving discontinuous Galerkin methods for the Vlasov–Maxwell system. Journal of Computational Physics, 279(??):145–173, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006172.

Cardoso:2010:LRO

[CD10] M. A. Cardoso and L. J. Durlofsky. Linearized reducedorder models for subsurface flow simulation. Journal of Computational Physics, 229(3):681–700, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900549X.

Cao:2012:NMD

[CD12a] Bing-Yang Cao and Ruo-Yu Dong. Nonequilibrium molecular dynamics simulation of shear viscosity by a uniform momentum source-and-sink scheme. Journal of Computational Physics, 231(16):5306–5316, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001982.

Celik:2012:CNM

[ÇD12b] Cem Çelik and Melda Duman. Crank–Nicolson method for the fractional diffusion equation with the Riesz fractional derivative. Journal of Computational Physics, 231(4):1743–1750, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006504.

Capecelatro:2013:ELS

[CD13a] Jesse Capecelatro and Olivier Desjardins. An Euler–Lagrange strategy for simulating particle-laden flows. Journal of Computational Physics, 238(??):1–31, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007462.

Carella:2013:LSS

[CD13b] Alfredo Raúl Carella and Carlos Alberto Dorao. Leastsquares spectral method for the solution of a fractional advection-dispersion equation. Journal of Computational Physics, 232(1):33–45, January 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002549.

Crivellini:2013:SAT

[CDB13] Andrea Crivellini, Valerio D’Alessandro, and Francesco Bassi. A Spalart–Allmaras turbulence model implementation in a discontinuous Galerkin solver for incompressible flows. Journal of Computational Physics, 241(??):388–415, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300034X.

Cruz:2012:CSC

[CdCD12] A. G. B. Cruz, E. G. Dutra do Carmo, and F. P. Duda. A consistent and stabilized continuous/discontinuous Galerkin method for fourth-order incompressible flow problems. Journal of Computational Physics, 231(16):5469–5488, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002264.

Chacon:2014:APS

[CdCNH14] L. Chacón, D. del Castillo-Negrete, and C. D. Hauck. An asymptotic-preserving semi-Lagrangian algorithm for the timedependent anisotropic heat transport equation. Journal of Computational Physics, 272(??):719–746, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003258.

Chacon:2011:RMM

[CDF11] L. Chacón, G. L. Delzanno, and J. M. Finn. Robust, multidimensional mesh-motion based on Monge–Kantorovich equidistribution. Journal of Computational Physics, 230(1):87–103, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005073.

Colella:2011:HOF

[CDHM11] P. Colella, M. R. Dorr, J. A. F. Hittinger, and D. F. Martin. High-order, finite-volume methods in mapped coordinates. Journal of Computational Physics, 230(8):2952–2976, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000064.

Cordier:2012:APA

[CDK12] Floraine Cordier, Pierre Degond, and Anela Kumbaro. An asymptotic-preserving all-speed scheme for the Euler and Navier–Stokes equations. Journal of Computational Physics, 231(17):5685–5704, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002069.

Clain:2011:HOF

[CDL11] S. Clain, S. Diot, and R. Loubère. A high-order finite volume method for systems of conservation laws — Multidimensional Optimal Order Detection (MOOD). Journal of Computational Physics, 230(10):4028–4050, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100115X.

Claisse:2012:NEP

[CDLL12] A. Claisse, B. Després, E. Labourasse, and F. Ledoux. A new exceptional points method with application to cell-centered Lagrangian schemes and curved meshes. Journal of Computational Physics, 231(11):4324–4354, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001064.

Cohen:2013:GBB

[CDS13] Bruce I. Cohen, Andris M. Dimits, and David J. Strozzi. A grid-based binary model for Coulomb collisions in plasmas. Journal of Computational Physics, 234(??):33–43, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005141.

Capuzzo-Dolcetta:2013:FPH

[CDSP13] R. Capuzzo-Dolcetta, M. Spera, and D. Punzo. A fully parallel, high precision, N-body code running on hybrid computing platforms. Journal of Computational Physics, 236(??):580–593, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006900.

Chang:2013:DNS

[CDT13] Chih-Hao Chang, Xiaolong Deng, and Theo G. Theofanous. Direct numerical simulation of interfacial instabilities: a consistent, conservative, all-speed, sharp-interface method. Journal of Computational Physics, 242(??):946–990, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000454.

Calo:2014:MEI

[CEGG14] Victor M. Calo, Yalchin Efendiev, Juan Galvis, and Mehdi Ghommem. Multiscale empirical interpolation for solving nonlinear PDEs. Journal of Computational Physics, 278 (??):204–220, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005452.

Chamakuri:2014:BCB

[CEK14] Nagaiah Chamakuri, Christian Engwer, and Karl Kunisch. Boundary control of bidomain equations with state-dependent switching source functions in the ionic model. Journal of Computational Physics, 273(??):227–242, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003635.

Chung:2014:AGH

[CEL14] Eric T. Chung, Yalchin Efendiev, and Guanglian Li. An adaptive GMsFEM for high-contrast flow problems. Journal of Computational Physics, 273(??):54–76, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400343X.

Charlot:2012:CLS

[CEP12] L. Charlot, S. Etienne, and D. Pelletier. A continuous Lagrangian sensitivity equation method for incompressible flow. Journal of Computational Physics, 231(18):5989–6011, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001404.

Casenave:2014:CBF

[CES14] Fabien Casenave, Alexandre Ern, and Guillaume Sylvand. Coupled BEM-FEM for the convected Helmholtz equation with non-uniform flow in a bounded domain. Journal of Computational Physics, 257(??):627–644, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300692X.

Chen:2010:ACM

[CF10] Ye Chen and Scott R. Fulton. An adaptive continuationmultigrid method for the balanced vortex model. Journal of Computational Physics, 229(6):2236–2248, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006615.

Ceniceros:2011:FRN

[CF11a] Hector D. Ceniceros and Jordan E. Fisher. A fast, robust, and non-stiff Immersed Boundary Method. Journal of Computational Physics, 230(12):5133–5153, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001884.

Chen:2011:AMC

[CF11b] Michael J. Chen and Lawrence K. Forbes. Accurate methods for computing inviscid and viscous Kelvin–Helmholtz instability. Journal of Computational Physics, 230(4):1499–1515, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000625X.

Claisse:2011:NPM

[CF11c] A. Claisse and P. Frey. A nonlinear PDE model for reconstructing a regular surface from sampled data using a level set formulation on triangular meshes. Journal of Computational Physics, 230(12):4636–4656, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001288.

Carlberg:2013:CGM

[CFCA13a] Kevin Carlberg, Charbel Farhat, Julien Cortial, and David Amsallem. Corrigendum to “The GNAT method for nonlinear model reduction: Effective implementation and application to computational fluid dynamics and turbulent flows” [J. Comput. Phys. 242 (2013) 623–647]. Journal of Computational Physics, 250(??):713, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003628. See [CFCA13b].

Carlberg:2013:GMN

[CFCA13b] Kevin Carlberg, Charbel Farhat, Julien Cortial, and David Amsallem. The GNAT method for nonlinear model reduction: Effective implementation and application to computational fluid dynamics and turbulent flows. Journal of Computational Physics, 242(??):623–647, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999113001472. See corrigendum [CFCA13a].

Chen:2011:NPA

[CFKL11] Kuan-Yu Chen, Ko-An Feng, Yongsam Kim, and Ming-Chih Lai. A note on pressure accuracy in immersed boundary method for Stokes flow. Journal of Computational Physics, 230(12):4377–4383, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001628.

Calef:2013:NKA

[CFW^{+}13] Matthew T. Calef, Erin D. Fichtl, James S. Warsa, Markus Berndt, and Neil N. Carlson. Nonlinear Krylov acceleration applied to a discrete ordinates formulation of the k-eigenvalue problem. Journal of Computational Physics, 238(??):188–209, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007553.

Cousins:2012:BCH

[CG12] W. Cousins and P. A. Gremaud. Boundary conditions for hemodynamics: the structured tree revisited. Journal of Computational Physics, 231(18):6086–6096, July 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002203.

Costa:2013:ELT

[CG13a] Anthony B. Costa and Jason R. Green. Extending the length and time scales of Gram–Schmidt Lyapunov vector computations. Journal of Computational Physics, 246(??):113–122, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002350.

Crochet:2013:NIM

[CG13b] M. W. Crochet and K. A. Gonthier. Numerical investigation of a modified family of centered schemes applied to multiphase equations with nonconservative sources. Journal of Computational Physics, 255(??):266–292, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300541X.

Coussement:2012:TDB

[CGC^{+}12] Axel Coussement, Olivier Gicquel, Jean Caudal, Benoît Fiorina, and Gérard Degrez. Three-dimensional boundary conditions for numerical simulations of reactive compressible flows with complex thermochemistry. Journal of Computational Physics, 231(17):5571–5611, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001593.

Ceniceros:2013:NAN

[CGC13] Hector D. Ceniceros and Carlos J. García-Cervera. A new approach for the numerical solution of diffusion equations with variable and degenerate mobility. Journal of Computational Physics, 246(??):1–10, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002143.

Coussement:2013:MRG

[CGF^{+}13] Axel Coussement, Olivier Gicquel, Benoît Fiorina, Gérard Degrez, and Nasser Darabiha. Multicomponent real gas 3-D-NSCBC for direct numerical simulation of reactive compressible viscous flows. Journal of Computational Physics, 245(??):259–280, July 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001149.

Charest:2012:SER

[CGG12] Marc R. J. Charest, Clinton P. T. Groth, and Ömer L. Gülder. Solution of the equation of radiative transfer using a Newton–Krylov approach and adaptive mesh refinement. Journal of Computational Physics, 231(8):3023–3040, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006711.

Choi:2011:IMS

[CGJ11] Wooyoung Choi, Arnaud Goullet, and Tae-Chang Jo. An iterative method to solve a regularized model for strongly nonlinear long internal waves. Journal of Computational Physics, 230(5):2021–2030, March 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006650.

Celledoni:2012:PER

[CGM^{+}12] E. Celledoni, V. Grimm, R. I. McLachlan, D. I. McLaren, D. O’Neale, B. Owren, and G. R. W. Quispel. Preserving energy resp. dissipation in numerical PDEs using the “Average Vector Field” method. Journal of Computational Physics, 231(20):6770–6789, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003373.

Christlieb:2014:HOT

[CGMQ14] Andrew Christlieb, Wei Guo, Maureen Morton, and Jing-Mei Qiu. A high order time splitting method based on integral deferred correction for semi-Lagrangian Vlasov simulations. Journal of Computational Physics, 267(??):7–27, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001326.

Cleveland:2010:EIM

[CGP10] Mathew A. Cleveland, Nick A. Gentile, and Todd S. Palmer. An extension of implicit Monte Carlo diffusion: Multigroup and the difference formulation. Journal of Computational Physics, 229(16):5707–5723, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001804.

Cheng:2011:RDP

[CGR11] Yingda Cheng, Irene M. Gamba, and Kui Ren. Recovering doping profiles in semiconductor devices with the Boltzmann–Poisson model. Journal of Computational Physics, 230(9):3391–3412, May 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000593.

Ciraolo:2013:CMH

[CGS13] Giulio Ciraolo, Francesco Gargano, and Vincenzo Sciacca. A computational method for the Helmholtz equation in unbounded domains based on the minimization of an integral functional. Journal of Computational Physics, 246(??):78–95, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002258.

Chirvasa:2010:FDM

[CH10] M. Chirvasa and S. Husa. Finite difference methods for second order in space, first order in time hyperbolic systems and the linear shifted wave equation as a model problem in numerical relativity. Journal of Computational Physics, 229(7):2675–2696, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006950.

Codina:2011:ATC

[CH11a] Ramon Codina and Noel Hernández. Approximation of the thermally coupled MHD problem using a stabilized finite element method. Journal of Computational Physics, 230(4):1281–1303, February 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006042.

Cotter:2011:NWP

[CH11b] C. J. Cotter and D. A. Ham. Numerical wave propagation for the triangular P 1_{DG}–P 2 finite element pair. Journal of Computational Physics, 230(8):2806–2820, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006984.

Cortez:2014:FNM

[CH14] Ricardo Cortez and Franz Hoffmann. A fast numerical method for computing doubly-periodic regularized Stokes flow in 3D. Journal of Computational Physics, 258(??):1–14, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007080.

Chandrashekar:2013:DGM

[Cha13] Praveen Chandrashekar. Discontinuous Galerkin method for Navier–Stokes equations using kinetic flux vector splitting. Journal of Computational Physics, 233(??):527–551, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005487.

Chang:2014:RAF

[Cha14] Lina Chang. A reconstruction algorithm with flexible stencils for anisotropic diffusion equations on 2D skewed meshes. Journal of Computational Physics, 256(??):484–500, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006177.

Chaudhuri:2011:UIB

[CHC11] Arnab Chaudhuri, Abdellah Hadjadj, and Ashwin Chinnayya. On the use of immersed boundary methods for shock/obstacle interactions. Journal of Computational Physics, 230(5):1731–1748, March 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006248.

Chou:2011:EAT

[CHHL11] So-Hsiang Chou, Tsung-Ming Huang, Wei-Qiang Huang, and Wen-Wei Lin. Efficient Arnoldi-type algorithms for rational eigenvalue problems arising in fluid-solid systems. Journal of Computational Physics, 230(5):2189–2206, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006868.

Chhay:2011:CSL

[CHHS11] M. Chhay, E. Hoarau, A. Hamdouni, and P. Sagaut. Comparison of some Lie-symmetry-based integrators. Journal of Computational Physics, 230(5):2174–2188, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006790.

Chiavazzo:2012:ASF

[Chi12] Eliodoro Chiavazzo. Approximation of slow and fast dynamics in multiscale dynamical systems by the linearized Relaxation Redistribution Method. Journal of Computational Physics, 231(4):1751–1765, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006498.

Chen:2013:CNA

[CHKT13] Li Chen, Ya-Ling He, Qinjun Kang, and Wen-Quan Tao. Coupled numerical approach combining finite volume and lattice Boltzmann methods for multi-scale multi-physicochemical processes. Journal of Computational Physics, 255(??):83–105, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300524X.

Cheng:2013:DBOb

[CHZ13a] Mulin Cheng, Thomas Y. Hou, and Zhiwen Zhang. A dynamically bi-orthogonal method for time-dependent stochastic partial differential equations I: Derivation and algorithms. Journal of Computational Physics, 242(??):843–868, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001526.

Cheng:2013:DBOa

[CHZ13b] Mulin Cheng, Thomas Y. Hou, and Zhiwen Zhang. A dynamically bi-orthogonal method for time-dependent stochastic partial differential equations II: Adaptivity and generalizations. Journal of Computational Physics, 242(??):753–776, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001393.

Caiazzo:2014:NIV

[CIJS14] Alfonso Caiazzo, Traian Iliescu, Volker John, and Swetlana Schyschlowa. A numerical investigation of velocitypressure reduced order models for incompressible flows. Journal of Computational Physics, 259(??):598–616, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008036.

Cortinovis:2014:IGE

[CJ14] Davide Cortinovis and Patrick Jenny. Iterative Galerkinenriched multiscale finite-volume method. Journal of Computational Physics, 277(??):248–267, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005737.

Chen:2011:GKS

[CJLC11] Songze Chen, Changqiu Jin, Cunbiao Li, and Qingdong Cai. Gas-kinetic scheme with discontinuous derivative for low speed flow computation. Journal of Computational Physics, 230(5):2045–2059, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006674.

Chandrasekaran:2013:MSN

[CJM13] S. Chandrasekaran, K. R. Jayaraman, and H. N. Mhaskar. Minimum Sobolev norm interpolation with trigonometric polynomials on the torus. Journal of Computational Physics, 249(??):96–112, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002192.

Coupez:2013:ATS

[CJN^{+}13] T. Coupez, G. Jannoun, N. Nassif, H. C. Nguyen, H. Digonnet, and E. Hachem. Adaptive time-step with anisotropic meshing for incompressible flows. Journal of Computational Physics, 241(??):195–211, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007413.

Chandler:2012:AAN

[CJNS12] Gary J. Chandler, Matthew P. Juniper, Joseph W. Nichols, and Peter J. Schmid. Adjoint algorithms for the Navier– Stokes equations in the low Mach number limit. Journal of Computational Physics, 231(4):1900–1916, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006681.

Christlieb:2014:HAS

[CJP^{+}14] Andrew Christlieb, Jaylan Jones, Keith Promislow, Brian Wetton, and Mark Willoughby. High accuracy solutions to energy gradient flows from material science models. Journal of Computational Physics, 257(??):193–215, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006633.

Chatelain:2010:FBE

[CK10] Philippe Chatelain and Petros Koumoutsakos. A Fourier-based elliptic solver for vortical flows with periodic and unbounded directions. Journal of Computational Physics, 229(7):2425–2431, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000021.

Celledoni:2011:SLM

[CK11] Elena Celledoni and Bawfeh Kingsley Kometa. Semi-Lagrangian multistep exponential integrators for index 2 differential-algebraic systems. Journal of Computational Physics, 230(9):3413–3429, May 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000611.

Cohen:2010:SLP

[CKD10] B. I. Cohen, A. J. Kemp, and L. Divol. Simulation of laser-plasma interactions and fast-electron transport in inhomogeneous plasma. Journal of Computational Physics, 229 (12):4591–4612, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911000104X.

Cockburn:2014:ASD

[CKR14] Bernardo Cockburn, Chiu-Yen Kao, and Fernando Reitich. An adaptive spectral/DG method for a reduced phase-space based level set approach to geometrical optics on curved elements. Journal of Computational Physics, 259(??):636–649, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113008176.

Calderer:2014:LSI

[CKS14] Antoni Calderer, Seokkoo Kang, and Fotis Sotiropoulos. Level set immersed boundary method for coupled simulation of air/ water interaction with complex floating structures. Journal of Computational Physics, 277(??):201–227, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005567.

Cai:2010:AMM

[CL10] Zhenning Cai and Ruo Li. An h-adaptive mesh method for Boltzmann–BGK/hydrodynamics coupling. Journal of Computational Physics, 229(5):1661–1680, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006123.

Chiavassa:2011:TDN

[CL11a] Guillaume Chiavassa and Bruno Lombard. Time domain numerical modeling of wave propagation in 2D heterogeneous porous media. Journal of Computational Physics, 230 (13):5288–5309, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111001811.

Chiu:2011:CPF

[CL11b] Pao-Hsiung Chiu and Yan-Ting Lin. A conservative phase field method for solving incompressible two-phase flows. Journal of Computational Physics, 230(1):185–204, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005243.

Connington:2013:LBS

[CL13] Kevin Connington and Taehun Lee. Lattice Boltzmann simulations of forced wetting transitions of drops on superhydrophobic surfaces. Journal of Computational Physics, 250 (??):601–615, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113003434.

Cai:2014:NXM

[CL14a] Zhenning Cai and Ruo Li. The NRxx method for polyatomic gases. Journal of Computational Physics, 267(??):63– 91, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001466.

Chen:2014:CSS

[CL14b] Kuan-Yu Chen and Ming-Chih Lai. A conservative scheme for solving coupled surface-bulk convection-diffusion equations with an application to interfacial flows with soluble surfactant. Journal of Computational Physics, 257(??):1–18, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006682.

Contrino:2014:LBS

[CLAL14] Dario Contrino, Pierre Lallemand, Pietro Asinari, and Li-Shi Luo. Lattice-Boltzmann simulations of the thermally driven 2D square cavity at high Rayleigh numbers. Journal of Computational Physics, 275(??):257–272, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004549.

Cheng:2013:IFS

[CLG13] Lei Cheng, Yizeng Li, and Karl Grosh. Including fluid shear viscosity in a structural acoustic finite element model using a scalar fluid representation. Journal of Computational Physics, 247(??):248–261, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002477.

Crouseilles:2013:APS

[CLM13] Nicolas Crouseilles, Mohammed Lemou, and Florian Méhats. Asymptotic Preserving schemes for highly oscillatory Vlasov– Poisson equations. Journal of Computational Physics, 248 (??):287–308, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113002878.

Canelas:2014:NRM

[CLN14] Alfredo Canelas, Antoine Laurain, and Antonio A. Novotny. A new reconstruction method for the inverse potential problem. Journal of Computational Physics, 268(??):417–431, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006967.

Cheng:2013:PPD

[CLQX13] Yue Cheng, Fengyan Li, Jianxian Qiu, and Liwei Xu. Positivity-preserving DG and central DG methods for ideal MHD equations. Journal of Computational Physics, 238 (??):255–280, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112007504.

Chiu:2010:DID

[CLS10] P. H. Chiu, R. K. Lin, and Tony W. H. Sheu. A differentially interpolated direct forcing immersed boundary method for predicting incompressible Navier–Stokes equations in time-varying complex geometries. Journal of Computational Physics, 229(12):4476–4500, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000860.

Costarelli:2013:ANS

[CLS13] Danilo Costarelli, Matteo Laurenzi, and Renato Spigler. Asymptotic-numerical solution of nonlinear systems of onedimensional balance laws. Journal of Computational Physics, 245(??):347–363, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113001496.

Chen:2014:GSW

[CLSX14] Chungang Chen, Xingliang Li, Xueshun Shen, and Feng Xiao. Global shallow water models based on multi-moment constrained finite volume method and three quasi-uniform spherical grids. Journal of Computational Physics, 271(??):191–223, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300702X.

Cheng:2010:LSM

[CLW10] Li-Tien Cheng, Bo Li, and Zhongming Wang. Levelset minimization of potential controlled Hadwiger valuations for molecular solvation. Journal of Computational Physics, 229(22):8497–8510, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004250.

Chen:2014:RMM

[CLX14] Huangxin Chen, Peipei Lu, and Xuejun Xu. A robust multilevel method for hybridizable discontinuous Galerkin method for the Helmholtz equation. Journal of Computational Physics, 264(??):133–151, May 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000801.

Chen:2012:SEB

[CLZ12] Sheng Chen, K. H. Luo, and Chuguang Zheng. A simple enthalpy-based lattice Boltzmann scheme for complicated thermal systems. Journal of Computational Physics, 231(24):8278–8294, October 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004652.

Colombant:2010:NFS

[CM10] Denis Colombant and Wallace Manheimer. Numerical fluid solutions for nonlocal electron transport in hot plasmas: Equivalent diffusion versus nonlocal source. Journal of Computational Physics, 229(11):4369–4381, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000902.

Cochelin:2013:PSA

[CM13] Bruno Cochelin and Marc Medale. Power series analysis as a major breakthrough to improve the efficiency of Asymptotic Numerical Method in the vicinity of bifurcations. Journal of Computational Physics, 236(??):594–607, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006936.

Chauchat:2014:TDN

[CM14] Julien Chauchat and Marc Médale. A three-dimensional numerical model for dense granular flows based on the μ(I) rheology. Journal of Computational Physics, 256(??):696–712, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006098.

Cyron:2013:MSB

[CMBW13] Christian J. Cyron, Kei W. Müller, Andreas R. Bausch, and Wolfgang A. Wall. Micromechanical simulations of biopolymer networks with finite elements. Journal of Computational Physics, 244(??):236–251, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006262.

Cornford:2013:AMF

[CMG^{+}13] Stephen L. Cornford, Daniel F. Martin, Daniel T. Graves, Douglas F. Ranken, Anne M. Le Brocq, Rupert M. Gladstone, Antony J. Payne, Esmond G. Ng, and William H. Lipscomb. Adaptive mesh, finite volume modeling of marine ice sheets. Journal of Computational Physics, 232(1):529–549, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005050.

Celledoni:2014:ILG

[CMO14] Elena Celledoni, Håkon Marthinsen, and Brynjulf Owren. An introduction to Lie group integrators — basics, new developments and applications. Journal of Computational Physics, 257(??):1040–1061, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000041.

Cervone:2010:SAJ

[CMS10a] A. Cervone, S. Manservisi, and R. Scardovelli. Simulation of axisymmetric jets with a finite element Navier–Stokes solver and a multilevel VOF approach. Journal of Computational Physics, 229(19):6853–6873, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000286X.

Crouseilles:2010:CSL

[CMS10b] Nicolas Crouseilles, Michel Mehrenberger, and Eric Sonnendrücker. Conservative semi-Lagrangian schemes for Vlasov equations. Journal of Computational Physics, 229(6):1927– 1953, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109006263.

Colella:2010:CSF

[CN10] Phillip Colella and Peter C. Norgaard. Controlling selfforce errors at refinement boundaries for AMR–PIC. Journal of Computational Physics, 229(4):947–957, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109003696.

Cagnone:2012:SIE

[CN12] J. S. Cagnone and S. K. Nadarajah. A stable interface element scheme for the p-adaptive lifting collocation penalty formulation. Journal of Computational Physics, 231(4):1615–1634, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006188.

Ceniceros:2010:TDF

[CNR10] Hector D. Ceniceros, Rudimar L. Nós, and Alexandre M. Roma. Three-dimensional, fully adaptive simulations of phasefield fluid models. Journal of Computational Physics, 229 (17):6135–6155, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110002329.

Canelas:2011:NMI

[CNR11] Alfredo Canelas, Antonio A. Novotny, and Jean R. Roche. A new method for inverse electromagnetic casting problems based on the topological derivative. Journal of Computational Physics, 230(9):3570–3588, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000830.

Carroll-Nellenback:2013:EPA

[CNSFD13] Jonathan J. Carroll-Nellenback, Brandon Shroyer, Adam Frank, and Chen Ding. Efficient parallelization for AMR MHD multiphysics calculations; implementation in AstroBEAR. Journal of Computational Physics, 236(??):461–476, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005931.

Chen:2013:FUQ

[CNST13] Xiao Chen, Brenda Ng, Yunwei Sun, and Charles Tong. A flexible uncertainty quantification method for linearly coupled multi-physics systems. Journal of Computational Physics, 248(??):383–401, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002581.

Coatleven:2012:HEP

[Coa12] Julien Coatléven. Helmholtz equation in periodic media with a line defect. Journal of Computational Physics, 231(4):1675–1704, February 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100622X.

Cook:2013:EHC

[Coo13] Andrew W. Cook. Effects of heat conduction on artificial viscosity methods for shock capturing. Journal of Computational Physics, 255(??):48–52, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005305.

Coupez:2011:MCL

[Cou11] T. Coupez. Metric construction by length distribution tensor and edge based error for anisotropic adaptive meshing. Journal of Computational Physics, 230(7):2391–2405, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000656X.

Chen:2010:CPB

[CP10] Z. J. Chen and A. J. Przekwas. A coupled pressure-based computational method for incompressible/compressible flows. Journal of Computational Physics, 229(24):9150–9165, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004766.

Can:2012:LSM

[CP12] Edip Can and Andrea Prosperetti. A level set method for vapor bubble dynamics. Journal of Computational Physics, 231(4):1533–1552, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006218.

Castro:2013:TSC

[CP13a] Hugo G. Castro and Rodrigo R. Paz. A time and space correlated turbulence synthesis method for Large Eddy Simulations. Journal of Computational Physics, 235(??):742–763, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006389.

Clancy:2013:CSI

[CP13b] Colm Clancy and Janusz A. Pudykiewicz. A class of semi-implicit predictor-corrector schemes for the time integration of atmospheric models. Journal of Computational Physics, 250(??):665–684, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004895.

Chatelin:2014:HGP

[CP14] Robin Chatelin and Philippe Poncet. Hybrid grid-particle methods and penalization: a Sherman–Morrison–Woodbury approach to compute 3D viscous flows using FFT. Journal of Computational Physics, 269(??):314–328, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002046.

Cheng:2013:SOS

[CPCU13] Jianhua Cheng, Scott E. Parker, Yang Chen, and Dmitri A. Uzdensky. A second-order semi-implicit δf method for hybrid simulation. Journal of Computational Physics, 245(??):364– 375, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001952.

Causley:2011:IHN

[CPJ11] Matthew F. Causley, Peter G. Petropoulos, and Shidong Jiang. Incorporating the Havriliak–Negami dielectric model in the FD–TD method. Journal of Computational Physics, 230(10):3884–3899, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111000945.

Chen:2013:FNA

[CPX13] Xiaoxiao Chen, Eun-Jae Park, and Dongbin Xiu. A flexible numerical approach for quantification of epistemic uncertainty. Journal of Computational Physics, 240(??):211–224, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000491.

Coco:2013:FDG

[CR13]  Armando Coco and Giovanni Russo.  Finite-difference

ghost-point multigrid methods on Cartesian grids for elliptic problems in arbitrary domains. Journal of Computational Physics, 241(??):464–501, May 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007292.

Cunha:2014:EAS

[CR14] G. Cunha and S. Redonnet. On the effective accuracy of spectral-like optimized finite-difference schemes for computational aeroacoustics. Journal of Computational Physics, 263(??):222–232, April 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000515.

Chen:2013:CVS

[CRG13] Qingshan Chen, Todd Ringler, and Max Gunzburger. A co-volume scheme for the rotating shallow water equations on conforming non-orthogonal grids. Journal of Computational Physics, 240(??):174–197, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000272.

Conti:2012:GAC

[CRK12] Christian Conti, Diego Rossinelli, and Petros Koumoutsakos. GPU and APU computations of Finite Time Lyapunov Exponent fields. Journal of Computational Physics, 231(5):2229–2244, March 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006322.

Crouseilles:2012:IAA

[CRS12] Nicolas Crouseilles, Ahmed Ratnani, and Eric Sonnendrücker. An Isogeometric Analysis approach for the study of the gyrokinetic quasi-neutrality equation. Journal of Computational Physics, 231(2):373–393, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005225.

Chatterjee:2014:NGF

[CRS14] Kausik Chatterjee, John R. Roadcap, and Surendra Singh. A new Green’s function Monte Carlo algorithm for the solution of the two-dimensional nonlinear Poisson–Boltzmann equation: Application to the modeling of the communication breakdown problem in space vehicles during re-entry. Journal of Computational Physics, 276(??):479–485, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400535X.

Clain:2010:MMM

[CRT10] S. Clain, D. Rochette, and R. Touzani. A multislope MUSCL method on unstructured meshes applied to compressible Euler equations for axisymmetric swirling flows. Journal of Computational Physics, 229(13):4884–4906, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001142.

Christlieb:2014:FDW

[CRT14] Andrew J. Christlieb, James A. Rossmanith, and Qi Tang. Finite difference weighted essentially non-oscillatory schemes with constrained transport for ideal magnetohydrodynamics. Journal of Computational Physics, 268(??):302–325, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001703.

Cheng:2010:CCL

[CS10] Juan Cheng and Chi-Wang Shu. A cell-centered Lagrangian scheme with the preservation of symmetry and conservation properties for compressible fluid flows in twodimensional cylindrical geometry. Journal of Computational Physics, 229(19):7191–7206, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003220.

Chang:2012:CLH

[CS12a] C. H. Chang and A. K. Stagg. A compatible Lagrangian hydrodynamic scheme for multicomponent flows with mixing. Journal of Computational Physics, 231(11):4279–4294, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000848.

Chen:2012:ESG

[CS12b] Feng Chen and Jie Shen. Efficient spectral-Galerkin methods for systems of coupled second-order equations and their applications. Journal of Computational Physics, 231(15):5016–5028, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001350.

Cotter:2012:MFE

[CS12c] C. J. Cotter and J. Shipton. Mixed finite elements for numerical weather prediction. Journal of Computational Physics, 231(21):7076–7091, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002628.

Chen:2013:GPS

[CS13] Feng Chen and Jie Shen. A GPU parallelized spectral method for elliptic equations in rectangular domains. Journal of Computational Physics, 250(??):555–564, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003975.

Cheng:2014:PPL

[CS14a] Juan Cheng and Chi-Wang Shu. Positivity-preserving Lagrangian scheme for multi-material compressible flow. Journal of Computational Physics, 257(??):143–168, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300661X.

Cheng:2014:SOS

[CS14b] Juan Cheng and Chi-Wang Shu. Second order symmetrypreserving conservative Lagrangian scheme for compressible Euler equations in two-dimensional cylindrical coordinates. Journal of Computational Physics, 272(??):245–265, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002940.

Cioaca:2014:OFI

[CS14c] Alexandru Cioaca and Adrian Sandu. An optimization framework to improve 4D-Var data assimilation system performance. Journal of Computational Physics, 275(??):377–389, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004859.

Conrad:2014:VAM

[CSB14] Daniel Conrad, Andreas Schneider, and Martin Böhle. A viscosity adaption method for lattice Boltzmann simulations. Journal of Computational Physics, 276(??):681–690, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005543.

Cossette:2014:MAT

[CSC14] Jean-François Cossette, Piotr K. Smolarkiewicz, and Paul Charbonneau. The Monge–Ampère trajectory correction for semi-Lagrangian schemes. Journal of Computational Physics, 274(??):208–229, October 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003623.

Choi:2013:CSS

[CSK13] Minseok Choi, Themistoklis P. Sapsis, and George Em Karniadakis. A convergence study for SPDEs using combined Polynomial Chaos and Dynamically-Orthogonal schemes. Journal of Computational Physics, 245(??):281–301, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001794.

Choi:2014:EDO

[CSK14] Minseok Choi, Themistoklis P. Sapsis, and George Em Karniadakis. On the equivalence of dynamically orthogonal and bi-orthogonal methods: Theory and numerical simulations. Journal of Computational Physics, 270(??):1–20, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400237X.

Cohen:2010:ABM

[CSKP10] J. Cohen, I. G. Shukhman, M. Karp, and J. Philip. An analytical-based method for studying the nonlinear evolution of localized vortices in planar homogeneous shear flows. Journal of Computational Physics, 229(20):7765–7773, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003505.

Codecasa:2010:NSB

[CST10] Lorenzo Codecasa, Ruben Specogna, and Francesco Trevisan. A new set of basis functions for the discrete geometric approach. Journal of Computational Physics, 229(19):7401–7410, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003384.

Cyr:2012:SSB

[CST12] Eric C. Cyr, John N. Shadid, and Raymond S. Tuminaro. Stabilization and scalable block preconditioning for the Navier–Stokes equations. Journal of Computational Physics, 231(2):345–363, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005195.

Chen:2014:NMM

[CSW14] Jie Chen, Shuyu Sun, and Xiao-Ping Wang. A numerical method for a model of two-phase flow in a coupled free flow and porous media system. Journal of Computational Physics, 268(??):1–16, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001697.

Chou:2014:OEC

[CSX14] Ching-Shan Chou, Chi-Wang Shu, and Yulong Xing. Optimal energy conserving local discontinuous Galerkin methods for second-order wave equation in heterogeneous media. Journal of Computational Physics, 272(??):88–107, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002721.

Carpio:2010:NBC

[CT10] A. Carpio and B. Tapiador. Nonreflecting boundary conditions for discrete waves. Journal of Computational Physics, 229(5):1879–1896, March 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006329.

Chappell:2013:SSL

[CT13] David J. Chappell and Gregor Tanner. Solving the stationary Liouville equation via a boundary element method. Journal of Computational Physics, 234(??):487–498, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005906.

Cotter:2014:FEE

[CT14] C. J. Cotter and J. Thuburn. A finite element exterior calculus framework for the rotating shallow-water equations. Journal of Computational Physics, 257(??):1506–1526, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006761.

Chappell:2012:BED

[CTG12] David J. Chappell, Gregor Tanner, and Stefano Giani. Boundary element dynamical energy analysis: a versatile method for solving two or three dimensional wave problems in the high frequency limit. Journal of Computational Physics, 231(18):6181–6191, July 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002847.

Cheng:2013:SCW

[CTJT13] Jun Bo Cheng, Eleuterio F. Toro, Song Jiang, and Weijun Tang. A sub-cell WENO reconstruction method for spatial derivatives in the ADER scheme. Journal of Computational Physics, 251(??):53–80, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004002.

Carr:2013:VSJ

[CTP13] E. J. Carr, I. W. Turner, and P. Perré. A variable-stepsize Jacobian-free exponential integrator for simulating transport in heterogeneous porous media: Application to wood drying. Journal of Computational Physics, 233(??):66–82, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200397X.

Cruz:2013:NSE

[CTSM13] Pedro A. Cruz, Murilo F. Tomé, Iain W. Stewart, and Sean McKee. Numerical solution of the Ericksen–Leslie dynamic equations for two-dimensional nematic liquid crystal flows. Journal of Computational Physics, 247(??):109–136, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002453.

Cui:2012:CAD

[Cui12] Mingrong Cui. Compact alternating direction implicit method for two-dimensional time fractional diffusion equation. Journal of Computational Physics, 231(6):2621–2633, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007200.

Cimrak:2010:LSM

[CV10] Ivan Cimrák and Roger Van Keer. Level set method for the inverse elliptic problem in nonlinear electromagnetism. Journal of Computational Physics, 229(24):9269–9283, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004948.

Cea:2012:UFV

[CVC12] L. Cea and M. E. Vázquez-Cendón. Unstructured finite volume discretisation of bed friction and convective flux in solute transport models linked to the shallow water equations. Journal of Computational Physics, 231(8):3317–3339, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000289.

Cooper:2013:SPM

[CVI^{+}13] J. D. Cooper, A. Valavanis, Z. Ikonić, P. Harrison, and J. E. Cunningham. Stable perfectly-matched-layer boundary conditions for finite-difference time-domain simulation of acoustic waves in piezoelectric crystals. Journal of Computational Physics, 253(??):239–246, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004920.

Chigullapalli:2010:ECN

[CVIA10] S. Chigullapalli, A. Venkattraman, M. S. Ivanov, and A. A. Alexeenko. Entropy considerations in numerical simulations of non-equilibrium rarefied flows. Journal of Computational Physics, 229(6):2139–2158, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006561.

Cagnone:2013:ALF

[CVN13] J. S. Cagnone, B. C. Vermeire, and S. Nadarajah. A padaptive LCP formulation for the compressible Navier–Stokes equations. Journal of Computational Physics, 233(??):324–338, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005396.

Chen:2010:MSE

[CW10] Duan Chen and Guo-Wei Wei. Modeling and simulation of electronic structure, material interface and random doping in nano-electronic devices. Journal of Computational Physics, 229(12):4431–4460, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000756.

Cai:2013:LSP

[CW13] Jiaxiang Cai and Yushun Wang. Local structure-preserving algorithms for the ‘good’ Boussinesq equation. Journal of Computational Physics, 239(??):72–89, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000399.

Chen:2014:ODA

[CW14a] Huangxin Chen and Xiao-Ping Wang. A one-domain approach for modeling and simulation of free fluid over a porous medium. Journal of Computational Physics, 259(??):650–671, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113008073.

Cheng:2014:NDG

[CW14b] Yingda Cheng and Zixuan Wang. A new discontinuous Galerkin finite element method for directly solving the Hamilton–Jacobi equations. Journal of Computational Physics, 268(??):134–153, July 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001673.

Chen:2011:MNH

[CWC11] Wen-Hwa Chen, Chun-Hung Wu, and Hsien-Chie Cheng. Modified Nosé–Hoover thermostat for solid state for constant temperature molecular dynamics simulation. Journal of Computational Physics, 230(16):6354–6366, July 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002786.

Chen:2010:ICS

[CWL10] Qianyi Chen, Jingtao Wang, and Kaixin Liu. Improved CE/SE scheme with particle level set method for numerical simulation of spall fracture due to high-velocity impact. Journal of Computational Physics, 229(19):7503–7519, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003487.

Cai:2013:LEP

[CWL13] Jiaxiang Cai, Yushun Wang, and Hua Liang. Local energypreserving and momentum-preserving algorithms for coupled nonlinear Schrödinger system. Journal of Computational Physics, 239(??):30–50, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000090.

Cai:2013:NDA

[CWS13] Wenjun Cai, Yushun Wang, and Yongzhong Song. Numerical dispersion analysis of a multi-symplectic scheme for the three dimensional Maxwell’s equations. Journal of Computational Physics, 234(??):330–352, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005888.

Clark:2011:LCF

[CWZ^{+}11] R. E. Clark, D. R. Welch, W. R. Zimmerman, C. L. Miller, T. C. Genoni, D. V. Rose, D. W. Price, P. N. Martin, D. J. Short, A. W. P. Jones, and J. R. Threadgold. Locally conformal finite-difference time-domain techniques for particle-in-cell plasma simulation. Journal of Computational Physics, 230(3):695–705, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005589.

Cao:2013:HOS

[CX13] Junying Cao and Chuanju Xu. A high order schema for the numerical solution of the fractional ordinary differential equations. Journal of Computational Physics, 238(??):154–168, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007449.

Chen:2012:UGK

[CXLC12] Songze Chen, Kun Xu, Cunbiao Lee, and Qingdong Cai. A unified gas kinetic scheme with moving mesh and velocity space adaptation. Journal of Computational Physics, 231 (20):6643–6664, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112002616.

Chen:2014:LSF

[CXLF14] De-Xiang Chen, Zi-Li Xu, Shi Liu, and Yong-Xin Feng. Least squares finite element method with high continuity NURBS basis for incompressible Navier–Stokes equations. Journal of Computational Physics, 260(??):204–221, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008383.

Chen:2010:AII

[CXZ10] Junqing Chen, Yifeng Xu, and Jun Zou. An adaptive inverse iteration for Maxwell eigenvalue problem based on edge elements. Journal of Computational Physics, 229(7):2649–2658, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006846.

Chang:2012:EAR

[CY12] Lina Chang and Guangwei Yuan. An efficient and accurate reconstruction algorithm for the formulation of cellcentered diffusion schemes. Journal of Computational Physics, 231(20):6935–6952, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003348.

Chen:2014:TBA

[CY14] Xiaodong Chen and Vigor Yang. Thickness-based adaptive mesh refinement methods for multi-phase flow simulations with thin regions. Journal of Computational Physics, 269(??):22–39, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001612.

Cantwell:2014:HOS

[CYK^{+}14] Chris D. Cantwell, Sergey Yakovlev, Robert M. Kirby, Nicholas S. Peters, and Spencer J. Sherwin. High-order spectral/hp element discretisation for reaction-diffusion problems on surfaces: Application to cardiac electrophysiology. Journal of Computational Physics, 257(??):813–829, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006955.

Chen:2011:KII

[CZ11a] Shanqin Chen and Yong-Tao Zhang. Krylov implicit integration factor methods for spatial discretization on high dimensional unstructured meshes: Application to discontinuous Galerkin methods. Journal of Computational Physics, 230(11):4336–4352, May 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000283.

Cheung:2011:NPM

[CZ11b] Lawrence C. Cheung and Tamer A. Zaki. A nonlinear PSE method for two-fluid shear flows with complex interfacial topology. Journal of Computational Physics, 230(17):6756–6777, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003111.

Chen:2013:NBP

[CZ13] Peng Chen and Nicholas Zabaras. A nonparametric belief propagation method for uncertainty quantification with applications to flow in random porous media. Journal of Computational Physics, 250(??):616–643, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003306.

Chang:2010:HMF

[CZL10] Haibin Chang, Dongxiao Zhang, and Zhiming Lu. History matching of facies distribution with the EnKF and level set parameterization. Journal of Computational Physics, 229 (20):8011–8030, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003748.

Donzis:2014:AFD

[DA14] Diego A. Donzis and Konduri Aditya. Asynchronous finitedifference schemes for partial differential equations. Journal of Computational Physics, 274(??):370–392, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004240.

Dumon:2011:PGD

[DAA11] A. Dumon, C. Allery, and A. Ammar. Proper general decomposition (PGD) for the resolution of Navier–Stokes equations. Journal of Computational Physics, 230(4):1387–1407, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006121.

Dadzie:2012:CSA

[Dad12] S. Kokou Dadzie. Comment on “A solution algorithm for the fluid dynamic equations based on a stochastic model for molecular motion, Jenny et al., Journal of Computational Physics, 229 (2010)”. Journal of Computational Physics, 231(21):7011–7013, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002811. See [JTH10].

Doha:2013:NAS

[DAEB13] E. H. Doha, W. M. Abd-Elhameed, and M. A. Bassuony. New algorithms for solving high even-order differential equations using third and fourth Chebyshev–Galerkin methods. Journal of Computational Physics, 236(??):563–579, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006857.

Daitche:2013:AIP

[Dai13] Anton Daitche. Advection of inertial particles in the presence of the history force: Higher order numerical schemes. Journal of Computational Physics, 254(??):93–106, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300497X.

Davidchack:2010:DEM

[Dav10] Ruslan L. Davidchack. Discretization errors in molecular dynamics simulations with deterministic and stochastic thermostats. Journal of Computational Physics, 229(24):9323– 9346, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110004985.

See corrigendum [Dav15].

Davidchack:2015:CDE

[Dav15] Ruslan L. Davidchack. Corrigendum to “Discretization errors in molecular dynamics simulations with deterministic and stochastic thermostats” [J. Comput. Phys. 229 (2010) 9323– 9346]. Journal of Computational Physics, 298(??):816, October 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999115004489. See [Dav10].

Dossou:2012:CTD

[DB12] Kokou B. Dossou and Lindsay C. Botten. A combined threedimensional finite element and scattering matrix method for the analysis of plane wave diffraction by bi-periodic, multilayered structures. Journal of Computational Physics, 231 (20):6969–6989, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003592.

Don:2013:AWE

[DB13] Wai-Sun Don and Rafael Borges. Accuracy of the weighted essentially non-oscillatory conservative finite difference schemes. Journal of Computational Physics, 250(??):347–372, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003501.

Doha:2014:JGL

[DBAV14] E. H. Doha, A. H. Bhrawy, M. A. Abdelkawy, and Robert A. Van Gorder. Jacobi–Gauss–Lobatto collocation method for the numerical solution of 1 + 1 nonlinear Schrödinger equations. Journal of Computational Physics, 261(??):244–255, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000229.

Dragna:2013:ADD

[DBHBB13] D. Dragna, C. Bogey, M. Hornikx, and P. Blanc-Benon. Analysis of the dissipation and dispersion properties of the multidomain Chebyshev pseudospectral method. Journal of Computational Physics, 255(??):31–47, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005287.

Duarte:2012:NNS

[DBM^{+}12] Max Duarte, Zdenek Bonaventura, Marc Massot, Anne Bourdon, Stéphane Descombes, and Thierry Dumont. A new numerical strategy with space–time adaptivity and error control for multi-scale streamer discharge simulations. Journal of Computational Physics, 231(3):1002–1019, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004062.

Donev:2010:FPK

[DBO^{+}10] Aleksandar Donev, Vasily V. Bulatov, Tomas Oppelstrup, George H. Gilmer, Babak Sadigh, and Malvin H. Kalos. A First-Passage Kinetic Monte Carlo algorithm for complex diffusion-reaction systems. Journal of Computational Physics, 229(9):3214–3236, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000057.

DelReyFernandez:2014:GFN

[DBZ14] David C. Del Rey Fernández, Pieter D. Boom, and David W. Zingg. A generalized framework for nodal first derivative summation-by-parts operators. Journal of Computational Physics, 266(??):214–239, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400076X.

Delzanno:2013:PMC

[DC13] G. L. Delzanno and E. Camporeale. On particle movers in cylindrical geometry for Particle-In-Cell simulations. Journal of Computational Physics, 253(??):259–277, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004798.

Dimits:2013:HOT

[DCC^{+}13] A. M. Dimits, B. I. Cohen, R. E. Caflisch, M. S. Rosin, and L. F. Ricketson. Higher-order time integration of Coulomb collisions in a plasma using Langevin equations. Journal of Computational Physics, 242(??):561–580, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000910.

Du:2011:TOF

[DCL11] Xi Du, Christophe Corre, and Alain Lerat. A thirdorder finite-volume residual-based scheme for the 2D Euler equations on unstructured grids. Journal of Computational Physics, 230(11):4201–4215, May 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100057X.

Roupert:2010:TPC

[dCRCS10] R. di Chiara Roupert, G. Chavent, and G. Schäfer. Threephase compressible flow in porous media: Total Differential Compatible interpolation of relative permeabilities. Journal of Computational Physics, 229(12):4762–4780, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001233.

Duddu:2011:DEP

[DCVM11] Ravindra Duddu, David L. Chopp, Peter Voorhees, and Brian Moran. Diffusional evolution of precipitates in elastic media using the extended finite element and the level set methods. Journal of Computational Physics, 230(4):1249–1264, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006029.

Degond:2012:FSL

[DD12] Pierre Degond and Giacomo Dimarco. Fluid simulations with localized Boltzmann upscaling by direct simulation Monte-Carlo. Journal of Computational Physics, 231(6):2414–2437, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006851.

Dapogny:2014:TDA

[DDF14] C. Dapogny, C. Dobrzynski, and P. Frey. Three-dimensional adaptive domain remeshing, implicit domain meshing, and applications to free and moving boundary problems. Journal of Computational Physics, 262(??):358–378, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000266.

Darbas:2013:CAP

[DDL13] M. Darbas, E. Darrigrand, and Y. Lafranche. Combining analytic preconditioner and Fast Multipole Method for the 3-D Helmholtz equation. Journal of Computational Physics, 236(??):289–316, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006948.

Degond:2010:MKF

[DDM10] Pierre Degond, Giacomo Dimarco, and Luc Mieussens. A multiscale kinetic-fluid solver with dynamic localization of kinetic effects. Journal of Computational Physics, 229(13):4907–4933, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001191.

Degond:2010:APP

[DDN^{+}10] Pierre Degond, Fabrice Deluzet, Laurent Navoret, An-Bang Sun, and Marie-Hélène Vignal. Asymptotic-Preserving Particle-In-Cell method for the Vlasov–Poisson system near quasineutrality. Journal of Computational Physics, 229(16): 5630–5652, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001774.

Degond:2012:NAE

[DDS12] P. Degond, F. Deluzet, and D. Savelief. Numerical approximation of the Euler–Maxwell model in the quasineutral limit.

Journal of Computational Physics, 231(4):1917–1946, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006577.

DeMichele:2010:SHR

[De 10] C. De Michele. Simulating hard rigid bodies. Journal of Computational Physics, 229(9):3276–3294, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000173.

Duchemin:2014:EIN

[DE14] Laurent Duchemin and Jens Eggers. The Explicit-Implicit-Null method: Removing the numerical instability of PDEs. Journal of Computational Physics, 263(??):37–52, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000400.

Degroote:2011:SBD

[Deg11] Joris Degroote. On the similarity between Dirichlet–Neumann with interface artificial compressibility and Robin–Neumann schemes for the solution of fluid-structure interaction problems. Journal of Computational Physics, 230(17):6399–6403, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003160.

Darian:2011:SDS

[DEH11] Hossein Mahmoodi Darian, Vahid Esfahanian, and Kazem Hejranfar. A shock-detecting sensor for filtering of highorder compact finite difference schemes. Journal of Computational Physics, 230(3):494–514, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005310.

Ducrozet:2014:NLW

[DEKBF14] Guillaume Ducrozet, Allan P. Engsig-Karup, Harry B. Bingham, and Pierre Ferrant. A non-linear wave decomposition model for efficient wave-structure interaction. Part A:

Formulation, validations and analysis. Journal of Computational Physics, 257(??):863–883, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006220.

Dellacherie:2010:AGT

[Del10] Stéphane Dellacherie. Analysis of Godunov type schemes applied to the compressible Euler system at low Mach number. Journal of Computational Physics, 229(4):978–1016, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005361.

DelPino:2011:MBM

[Del11] Stéphane Del Pino. Metric-based mesh adaptation for 2D Lagrangian compressible flows. Journal of Computational Physics, 230(5):1793–1821, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006455.

Dellar:2013:LBM

[Del13] Paul J. Dellar. Lattice Boltzmann magnetohydrodynamics with current-dependent resistivity. Journal of Computational Physics, 237(??):115–131, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007012.

Dellar:2014:LBA

[Del14] Paul J. Dellar. Lattice Boltzmann algorithms without cubic defects in Galilean invariance on standard lattices. Journal of Computational Physics, 259(??):270–283, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007833.

Denaro:2011:WDF

[Den11a] Filippo Maria Denaro. What does Finite Volume-based implicit filtering really resolve in Large-Eddy Simulations? Journal of Computational Physics, 230(10):3849–3883, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000933.

Densmore:2011:AAS

[Den11b] Jeffery D. Densmore. Asymptotic analysis of the spatial discretization of radiation absorption and re-emission in Implicit Monte Carlo. Journal of Computational Physics, 230(4):1116–1133, February 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005863.

Dritschel:2010:CLA

[DF10a]  David G. Dritschel and Jérôme Fontane.  The com-

bined Lagrangian advection method. Journal of Computational Physics, 229(14):5408–5417, July 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001671.

Dubinkina:2010:SRV

[DF10b] Svetlana Dubinkina and Jason Frank. Statistical relevance of vorticity conservation in the Hamiltonian particle-mesh method. Journal of Computational Physics, 229(7):2634–2648, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006834.

Dodd:2014:FPC

[DF14] Michael S. Dodd and Antonino Ferrante. A fast pressurecorrection method for incompressible two-fluid flows. Journal of Computational Physics, 273(??):416–434, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003702.

Diot:2014:IRM

[DFD14] S. Diot, M. M. François, and E. D. Dendy. An interface reconstruction method based on analytical formulae for 2D planar and axisymmetric arbitrary convex cells. Journal of Computational Physics, 275(??):53–64, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004781.

Dolean:2010:LID

[DFFL10] Victorita Dolean, Hassan Fahs, Loula Fezoui, and Stéphane Lanteri. Locally implicit discontinuous Galerkin method for time domain electromagnetics. Journal of Computational Physics, 229(2):512–526, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005300.

Diaz:2014:SOP

[DFNNRdlA14] M. J. Castro Díaz, E. D. Fernández-Nieto, G. Narbona-Reina, and M. de la Asunción. A second order PVM flux limiter method. Application to magnetohydrodynamics and shallow stratified flows. Journal of Computational Physics, 262(??):172–193, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400014X.

DiPietro:2014:PEE

[DFVY14] Daniele A. Di Pietro, Eric Flauraud, Martin Vohralík, and Soleiman Yousef. A posteriori error estimates, stopping criteria, and adaptivity for multiphase compositional Darcy flows in porous media. Journal of Computational Physics, 276(??):163–187, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114004811.

Dorr:2010:NAS

[DFW^{+}10] M. R. Dorr, J.-L. Fattebert, M. E. Wickett, J. F. Belak, and P. E. A. Turchi. A numerical algorithm for the solution of a phase-field model of polycrystalline materials. Journal of Computational Physics, 229(3):626–641, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005336.

Dieste:2012:RPM

[DG12] M. Dieste and G. Gabard. Random particle methods applied to broadband fan interaction noise. Journal of Computational Physics, 231(24):8133–8151, October 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004299.

Drouin:2010:PCM

[DGAH10] M. Drouin, L. Gremillet, J.-C. Adam, and A. Héron. Particlein-cell modeling of relativistic laser-plasma interaction with the adjustable-damping, direct implicit method. Journal of Computational Physics, 229(12):4781–4812, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001257.

Du:2014:IBM

[DGF14] Jian Du, Robert D. Guy, and Aaron L. Fogelson. An immersed boundary method for two-fluid mixtures. Journal of Computational Physics, 262(??):231–243, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000291.

Detrixhe:2013:PFS

[DGM13] Miles Detrixhe, Frédéric Gibou, and Chohong Min. A parallel fast sweeping method for the eikonal equation. Journal of Computational Physics, 237(??):46–55, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200722X.

Danaila:2010:FEM

[DH10] Ionut Danaila and Frédéric Hecht. A finite element method with mesh adaptivity for computing vortex states in fastrotating Bose–Einstein condensates. Journal of Computational Physics, 229(19):6946–6960, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002937.

Degond:2013:SOH

[DH13] Pierre Degond and Jiale Hua. Self-organized hydrodynamics with congestion and path formation in crowds. Journal of Computational Physics, 237(??):299–319, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007139.

Dolejsi:2011:ESS

[DHH11] V. Dolejsí, M. Holík, and J. Hozman. Efficient solution strategy for the semi-implicit discontinuous Galerkin discretization of the Navier–Stokes equations. Journal of Computational Physics, 230(11):4176–4200, May 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005851.

Degond:2011:NSE

[DHN11] Pierre Degond, Jiale Hua, and Laurent Navoret. Numerical simulations of the Euler system with congestion constraint. Journal of Computational Physics, 230(22):8057– 8088, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004153.

Ding:2014:DIM

[DjY14] Hang Ding and Cheng jun Yuan. On the diffuse interface method using a dual-resolution Cartesian grid. Journal of Computational Physics, 273(??):243–254, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003416.

Desmarais:2014:OBC

[DK14] J. L. Desmarais and J. G. M. Kuerten. Open boundary conditions for the Diffuse Interface Model in 1-D. Journal of Computational Physics, 263(??):393–418, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400059X.

Duben:2012:DCL

[DKA12] Peter D. Düben, Peter Korn, and Vadym Aizinger. A discontinuous/continuous low order finite element shallow water model on the sphere. Journal of Computational Physics, 231(6):2396–2413, March 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006735.

Dong:2014:RAO

[DKC14] S. Dong, G. E. Karniadakis, and C. Chryssostomidis. A robust and accurate outflow boundary condition for incompressible flow simulations on severely-truncated unbounded domains. Journal of Computational Physics, 261(??):83–105, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113008504.

Dickson:2011:IEV

[DKH11] Neil G. Dickson, Kamran Karimi, and Firas Hamze. Importance of explicit vectorization for CPU and GPU software performance. Journal of Computational Physics, 230 (13):5383–5398, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111002026.

Dutykh:2011:FVS

[DKM11] Denys Dutykh, Theodoros Katsaounis, and Dimitrios Mitsotakis. Finite volume schemes for dispersive wave propagation and runup. Journal of Computational Physics, 230 (8):3035–3061, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111000118.

Donde:2012:MQB

[DKR12] Pratik Donde, Heeseok Koo, and Venkat Raman. A multivariate quadrature based moment method for LES based modeling of supersonic combustion. Journal of Computational Physics, 231(17):5805–5821, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002136.

Dobrev:2014:HOC

[DKR14] Veselin A. Dobrev, Tzanio V. Kolev, and Robert N. Rieben. High order curvilinear finite elements for elasticplastic Lagrangian dynamics. Journal of Computational Physics, 257(??):1062–1080, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000466.

Du:2010:TPM

[DKW10] Ping Du, Mikhail Khenner, and Harris Wong. A tangentplane marker-particle method for the computation of threedimensional solid surfaces evolving by surface diffusion on a substrate. Journal of Computational Physics, 229(3):813–827, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005580.

Dosopoulos:2010:ILE

[DL10a] Stylianos Dosopoulos and Jin-Fa Lee. Interconnect and lumped elements modeling in interior penalty discontinuous Galerkin time-domain methods. Journal of Computational Physics, 229(22):8521–8536, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004298.

Du:2010:EDI

[DL10b] Chuanbin Du and Dong Liang. An efficient S–DDM iterative approach for compressible contamination fluid flows in porous media. Journal of Computational Physics, 229(12):4501–4521, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000999.

Dupont:2010:NSS

[DL10c] Jean-Baptiste Dupont and Dominique Legendre. Numerical simulation of static and sliding drop with contact angle hysteresis. Journal of Computational Physics, 229(7):2453–2478, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109004203.

Despres:2012:SCC

[DL12] Bruno Després and Emmanuel Labourasse. Stabilization of cell-centered compressible Lagrangian methods using subzonal entropy. Journal of Computational Physics, 231(20):6559–6595, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001994.

Dimarco:2013:TUEa

[DL13a] Giacomo Dimarco and Raphaël Loubere. Towards an ultra efficient kinetic scheme. Part I: Basics on the BGK equation. Journal of Computational Physics, 255(??):680–698, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006870.

Dimarco:2013:TUEb

[DL13b] Giacomo Dimarco and Raphaël Loubere. Towards an ultra efficient kinetic scheme. Part II: the high order case. Journal of Computational Physics, 255(??):699–719, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004907.

Ding:2013:MSF

[DL13c] Hengfei Ding and Changpin Li. Mixed spline function method for reaction-subdiffusion equations. Journal of Computational Physics, 242(??):103–123, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001241.

Demaldent:2011:FAP

[DLC11] Edouard Demaldent, David P. Levadoux, and Gary Cohen. Fast and accurate point-based method for time-harmonic Maxwell problems involving thin layer materials. Journal of Computational Physics, 230(14):5774–5786, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002221.

Daru:2010:NMS

[DLDL10] V. Daru, P. Le Quéré, M.-C. Duluc, and O. Le Maître. A numerical method for the simulation of low Mach number liquid-gas flows. Journal of Computational Physics, 229(23):8844–8867, November 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000450X.

Diaz:2013:HOE

[DLGP13] M. J. Castro Díaz, J. A. López-García, and Carlos Parés. High order exactly well-balanced numerical methods for shallow water systems. Journal of Computational Physics, 246 (??):242–264, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113002118.

Plata:2010:WBM

[dlLPC10] M. de la Llave Plata and R. S. Cant. A wavelet-based multiresolution approach to large-eddy simulation of turbulence. Journal of Computational Physics, 229(20):7715–7738, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003426.

Plata:2012:UBI

[dlLPCP12] Marta de la Llave Plata, Stewart Cant, and Robert Prosser. On the use of biorthogonal interpolating wavelets for largeeddy simulation of turbulence. Journal of Computational Physics, 231(20):6754–6769, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003397.

Deng:2014:CTO

[DLLW14] Yongbo Deng, Zhenyu Liu, Yongshun Liu, and Yihui Wu. Combination of topology optimization and optimal control method. Journal of Computational Physics, 257(??):374–399, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006475.

Doisneau:2013:EMF

[DLM^{+}13a] F. Doisneau, F. Laurent, A. Murrone, J. Dupays, and M. Massot. Eulerian multi-fluid models for the simulation of dynamics and coalescence of particles in solid propellant combustion. Journal of Computational Physics, 234(??):230–262, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005700.

Dudzinski:2013:WBB

[DLM13b] M. Dudzinski and M. Lukácová-Medvid’ová. Well-balanced bicharacteristic-based scheme for multilayer shallow water flows including wet/dry fronts. Journal of Computational Physics, 235(??):82–113, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006407.

Duran:2013:WBN

[DLM13c] A. Duran, Q. Liang, and F. Marche. On the wellbalanced numerical discretization of shallow water equations on unstructured meshes. Journal of Computational Physics, 235(??):565–586, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112006365.

Degond:2012:APM

[DLNN12] Pierre Degond, Alexei Lozinski, Jacek Narski, and Claudia Negulescu. An asymptotic-preserving method for highly anisotropic elliptic equations based on a micro–macro decomposition. Journal of Computational Physics, 231(7):2724–2740, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006966.

Deng:2011:TOU

[DLZ^{+}11] Yongbo Deng, Zhenyu Liu, Ping Zhang, Yongshun Liu, and Yihui Wu. Topology optimization of unsteady incompressible Navier–Stokes flows. Journal of Computational Physics, 230(17):6688–6708, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003019.

Dehkordi:2013:MRM

[DM13] Mehdi Mosharaf Dehkordi and Mehrdad Taghizadeh Manzari. A multi-resolution multiscale finite volume method for simulation of fluid flows in heterogeneous porous media. Journal of Computational Physics, 248(??):339–362, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002672.

Danaila:2014:NMA

[DMHL14] Ionut Danaila, Raluca Moglan, Frédéric Hecht, and Stéphane Le Masson. A Newton method with adaptive finite elements for solving phase-change problems with natural convection. Journal of Computational Physics, 274(??):826–840, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004434.

Deng:2013:FSG

[DMM^{+}13] Xiaogang Deng, Yaobing Min, Meiliang Mao, Huayong Liu, Guohua Tu, and Hanxin Zhang. Further studies on Geometric Conservation Law and applications to high-order finite difference schemes with stationary grids. Journal of Computational Physics, 239(??):90–111, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007255.

Donat:2014:WBA

[DMMGM14] Rosa Donat, M. Carmen Martí, Anna Martínez-Gavara, and Pep Mulet. Well-Balanced Adaptive Mesh Refinement for shallow water flows. Journal of Computational Physics, 257(??):937–953, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006463.

Duben:2014:UIP

[DMP14] Peter D. Düben, Hugh McNamara, and T. N. Palmer. The use of imprecise processing to improve accuracy in weather & climate prediction. Journal of Computational Physics, 271(??):2–18, August 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113007183.

Dimarco:2014:APA

[DMR14] Giacomo Dimarco, Luc Mieussens, and Vittorio Rispoli. An asymptotic preserving automatic domain decomposition method for the Vlasov–Poisson-BGK system with applications to plasmas. Journal of Computational Physics, 274 (??):122–139, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114004069.

Deng:2011:GCL

[DMT^{+}11] Xiaogang Deng, Meiliang Mao, Guohua Tu, Huayong Liu, and Hanxin Zhang. Geometric conservation law and applications to high-order finite difference schemes with stationary grids. Journal of Computational Physics, 230(4):1100–1115, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000584X.

Davydov:2011:AMC

[DO11a] Oleg Davydov and Dang Thi Oanh. Adaptive meshless centres and RBF stencils for Poisson equation. Journal of Computational Physics, 230(2):287–304, January 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004997.

Doostan:2011:NAS

[DO11b] Alireza Doostan and Houman Owhadi. A non-adapted sparse approximation of PDEs with stochastic inputs. Journal of Computational Physics, 230(8):3015–3034, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000106.

Dong:2010:BLM

[Don10] S. Dong. BDF-like methods for nonlinear dynamic analysis. Journal of Computational Physics, 229(8):3019–3045, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109007141.

Dong:2011:SNS

[Don11] Xuanchun Dong. A short note on simplified pseudospectral methods for computing ground state and dynamics of spherically symmetric Schrödinger–Poisson–Slater system. Journal of Computational Physics, 230(22):7917–7922, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004542.

Dong:2014:EAI

[Don14a] S. Dong. An efficient algorithm for incompressible N-phase flows. Journal of Computational Physics, 276(??):691–728, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005488.

Dong:2014:OBC

[Don14b] S. Dong. An outflow boundary condition and algorithm for incompressible two-phase flows with phase field approach. Journal of Computational Physics, 266(??):47–73, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001314.

Dellacherie:2010:ICG

[DOR10] Stéphane Dellacherie, Pascal Omnes, and Felix Rieper. The influence of cell geometry on the Godunov scheme applied to the linear wave equation. Journal of Computational Physics, 229(14):5315–5338, July 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001221.

Devendran:2012:IBE

[DP12] Dharshi Devendran and Charles S. Peskin. An immersed boundary energy-based method for incompressible viscoelasticity. Journal of Computational Physics, 231(14):4613–4642, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001210.

dePando:2012:EED

[dPSS12] Miguel Fosas de Pando, Denis Sipp, and Peter J. Schmid. Efficient evaluation of the direct and adjoint linearized dynamics from compressible flow solvers. Journal of Computational Physics, 231(23):7739–7755, October 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003725.

Dana:2011:PCS

[DR11] Saswati Dana and Soumyendu Raha. Physically consistent simulation of mesoscale chemical kinetics: The non-negative FIS-α method. Journal of Computational Physics, 230(24):8813–8834, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005158.

Ding:2014:ITC

[DR14] Tian Ding and Kui Ren. Inverse transport calculations in optical imaging with subspace optimization algorithms. Journal of Computational Physics, 273(??):212–226, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400360X.

Dekeyser:2014:ADT

[DRB14] W. Dekeyser, D. Reiter, and M. Baelmans. Automated divertor target design by adjoint shape sensitivity analysis and a one-shot method. Journal of Computational Physics, 278(??):117–132, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005774.

Driscoll:2010:ASC

[Dri10] Tobin A. Driscoll. Automatic spectral collocation for integral, integro-differential, and integrally reformulated differential equations. Journal of Computational Physics, 229 (17):5980–5998, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110002056.

DeVille:2011:WFA

[DRW11] R. E. L. DeVille, N. Riemer, and M. West. Weighted Flow Algorithms (WFA) for stochastic particle coagulation. Journal of Computational Physics, 230(23):8427–8451, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004554.

Druskin:2014:EKS

[DRZ14] Vladimir Druskin, Rob Remis, and Mikhail Zaslavsky. An extended Krylov subspace model-order reduction technique to simulate wave propagation in unbounded domains. Journal of Computational Physics, 272(??):608–618, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003271.

Dong:2010:USR

[DS10] S. Dong and J. Shen. An unconditionally stable rotational velocity-correction scheme for incompressible flows. Journal of Computational Physics, 229(19):7013–7029, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002986.

Dong:2011:WFB

[DS11] Bin Dong and Zuowei Shen. Wavelet frame based surface reconstruction from unorganized points. Journal of Computational Physics, 230(22):8247–8255, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004505.

Dong:2012:TSS

[DS12] S. Dong and J. Shen. A time-stepping scheme involving constant coefficient matrices for phase-field simulations of twophase incompressible flows with large density ratios. Journal of Computational Physics, 231(17):5788–5804, July 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002239.

Dubcova:2011:CMH

[DSHP11] Lenka Dubcova, Pavel Solin, Glen Hansen, and HyeongKae Park. Comparison of multimesh hp-FEM to interpolation and projection methods for spatial coupling of thermal and neutron diffusion calculations. Journal of Computational Physics, 230(4):1182–1197, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005905.

Dolgov:2014:LRA

[DST14] S. V. Dolgov, A. P. Smirnov, and E. E. Tyrtyshnikov. Lowrank approximation in the numerical modeling of the Farley– Buneman instability in ionospheric plasma. Journal of Computational Physics, 263(??):268–282, April 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000564.

Du:2013:AHO

[DSZ13] Kui Du, Weiwei Sun, and Xiaoping Zhang. Arbitrary high-order C^{0} tensor product Galerkin finite element methods for the electromagnetic scattering from a large cavity. Journal of Computational Physics, 242(??):181–195, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001253.

Daldorff:2014:TWC

[DTG^{+}14] Lars K. S. Daldorff, Gábor Tóth, Tamas I. Gombosi, Giovanni Lapenta, Jorge Amaya, Stefano Markidis, and Jeremiah U. Brackbill. Two-way coupling of a global Hall magnetohydrodynamics model with a local implicit particle-in-cell model. Journal of Computational Physics, 268(??):236–254, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400182X.

Densmore:2012:HTD

[DTU12] Jeffery D. Densmore, Kelly G. Thompson, and Todd J. Urbatsch. A hybrid transport-diffusion Monte Carlo method for frequency-dependent radiative-transfer simulations. Journal of Computational Physics, 231(20):6924–6934, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200335X.

Duan:2014:CMS

[DTYY14] Huo-Yuan Duan, Roger C. E. Tan, Suh-Yuh Yang, and Cheng-Shu You. Computation of Maxwell singular solution by nodal-continuous elements. Journal of Computational Physics, 268(??):63–83, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001727.

Du:2011:CPE

[Du11a] Kui Du. A composite preconditioner for the electromagnetic scattering from a large cavity. Journal of Computational Physics, 230(22):8089–8108, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100430X.

Du:2011:TTB

[Du11b] Kui Du. Two transparent boundary conditions for the electromagnetic scattering from two-dimensional overfilled cavities. Journal of Computational Physics, 230(15):5822–5835, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002178.

Duru:2014:PML

[Dur14] Kenneth Duru. A perfectly matched layer for the timedependent wave equation in heterogeneous and layered media. Journal of Computational Physics, 257(??):757–781, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006980.

DeMicheli:2013:EGP

[DV13a] Enrico De Micheli and Giovanni Alberto Viano. The expansion in Gegenbauer polynomials: a simple method for the fast computation of the Gegenbauer coefficients. Journal of Computational Physics, 239(??):112–122, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000387.

DeStefano:2013:WBA

[DV13b] Giuliano De Stefano and Oleg V. Vasilyev. Wavelet-based adaptive large-eddy simulation with explicit filtering. Journal of Computational Physics, 238(??):240–254, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200575X.

Duru:2014:SHO

[DV14] Kenneth Duru and Kristoffer Virta. Stable and high order accurate difference methods for the elastic wave equation in discontinuous media. Journal of Computational Physics, 279(??):37–62, December 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006226.

daSilvaInacio:2023:CPI

[dv23] Leonardo Afonso da Silva Inácio and André von Borries Lopes. Comment on “Progress and investigation on lattice Boltzmann modeling of multiple immiscible fluids or components with variable density and viscosity ratios”. Journal of Computational Physics, 474(??):??, February 1, 2023. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199912200883X. See [LRT13].

Denner:2014:CVM

[DvW14] Fabian Denner and Berend G. M. van Wachem. Compressive VOF method with skewness correction to capture sharp interfaces on arbitrary meshes. Journal of Computational Physics, 279(??):127–144, December 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006287.

Densmore:2010:SAT

[DWL10] Jeffery D. Densmore, James S. Warsa, and Robert B. Lowrie. Stability analysis and time-step limits for a Monte Carlo Compton-scattering method. Journal of Computational Physics, 229(10):3691–3705, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000458.

Duan:2010:HOC

[DWZ10] Le Duan, Xiaowen Wang, and Xiaolin Zhong. A highorder cut-cell method for numerical simulation of hypersonic boundary-layer instability with surface roughness. Journal of Computational Physics, 229(19):7207–7237, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003232.

Deng:2013:GIC

[DXB^{+}13] Shaozhong Deng, Changfeng Xue, Andriy Baumketner, Donald Jacobs, and Wei Cai. Generalized image charge solvation model for electrostatic interactions in molecular dynamics simulations of aqueous solutions. Journal of Computational Physics, 245(??):84–106, July 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002052.

Dai:2014:TDC

[DYYA14] Honghua Dai, Xiaokui Yue, Jianping Yuan, and Satya N. Atluri. A time domain collocation method for studying the aeroelasticity of a two dimensional airfoil with a structural nonlinearity. Journal of Computational Physics, 270(??):214–237, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002502.

Dumbser:2013:AWF

[DZHB13] Michael Dumbser, Olindo Zanotti, Arturo Hidalgo, and Dinshaw S. Balsara. ADER-WENO finite volume schemes with space–time adaptive mesh refinement. Journal of Computational Physics, 248(??):257–286, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002660.

Dosopoulos:2013:NCP

[DZL13] Stylianos Dosopoulos, Bo Zhao, and Jin-Fa Lee. Nonconformal and parallel discontinuous Galerkin time domain method for Maxwell’s equations: EM analysis of IC packages. Journal of Computational Physics, 238(??):48–70, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007309.

Dumbser:2014:PSL

[DZLD14] Michael Dumbser, Olindo Zanotti, Raphaël Loubère, and Steven Diot. A posteriori subcell limiting of the discontinuous Galerkin finite element method for hyperbolic conservation laws. Journal of Computational Physics, 278(??):47–75, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005555.

Exl:2012:FSF

[EAB^{+}12] L. Exl, W. Auzinger, S. Bance, M. Gusenbauer, F. Reichel, and T. Schrefl. Fast stray field computation on tensor grids. Journal of Computational Physics, 231(7):2840–2850, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007510.

ElBouajaji:2014:ALM

[EAG14] M. El Bouajaji, X. Antoine, and C. Geuzaine. Approximate local magnetic-to-electric surface operators for timeharmonic Maxwell’s equations. Journal of Computational Physics, 279(??):241–260, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114006470.

Eriksson:2011:SCM

[EAN11] Sofia Eriksson, Qaisar Abbas, and Jan Nordström. A stable and conservative method for locally adapting the design order of finite difference schemes. Journal of Computational Physics, 230(11):4216–4231, May 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006352.

Ehrl:2013:CAS

[EBGW13] Andreas Ehrl, Georg Bauer, Volker Gravemeier, and Wolfgang A. Wall. A computational approach for the simulation of natural convection in electrochemical cells. Journal of Computational Physics, 235(??):764–785, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005116.

Errera:2013:OSN

[EC13] Marc-Paul Errera and Sébastien Chemin. Optimal solutions of numerical interface conditions in fluid-structure thermal analysis. Journal of Computational Physics, 245(??):431–455, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001770.

Edeling:2014:PRS

[ECD14] W. N. Edeling, P. Cinnella, and R. P. Dwight. Predictive RANS simulations via Bayesian Model-Scenario Averaging. Journal of Computational Physics, 275(??):65–91, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004707.

Edeling:2014:BEP

[ECDB14] W. N. Edeling, P. Cinnella, R. P. Dwight, and H. Bijl. Bayesian estimates of parameter variability in the k–ϵ turbulence model. Journal of Computational Physics, 258(??):73–94, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007031.

Eskilsson:2014:DBT

[EEK14] Claes Eskilsson and Allan P. Engsig-Karup. On devising Boussinesq-type models with bounded eigenspectra: One horizontal dimension. Journal of Computational Physics, 271(??):261–280, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005925.

Engquist:2013:FSM

[EFT13] Björn Engquist, Brittany D. Froese, and Yen-Hsi Richard Tsai. Fast sweeping methods for hyperbolic systems of conservation laws at steady state. Journal of Computational Physics, 255(??):316–338, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005743.

Efendiev:2012:LGM

[EGG12] Yalchin Efendiev, Juan Galvis, and Eduardo Gildin. Localglobal multiscale model reduction for flows in high-contrast heterogeneous media. Journal of Computational Physics, 231(24):8100–8113, October 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004160.

Efendiev:2013:GMFa

[EGH13] Yalchin Efendiev, Juan Galvis, and Thomas Y. Hou. Generalized multiscale finite element methods (GMsFEM). Journal of Computational Physics, 251(??):116–135, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003392.

Efendiev:2013:GMFb

[EGL^{+}13] Y. Efendiev, J. Galvis, R. Lazarov, M. Moon, and M. Sarkis. Generalized multiscale finite element method. symmetric interior penalty coupling. Journal of Computational Physics, 255(??):1–15, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911300510X.

Efendiev:2011:MFE

[EGW11] Yalchin Efendiev, Juan Galvis, and Xiao-Hui Wu. Multiscale finite element methods for high-contrast problems using local spectral basis functions. Journal of Computational Physics, 230(4):937–955, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005292.

Evans:2013:IDC

[EH13] John A. Evans and Thomas J. R. Hughes. Isogeometric divergence-conforming B-splines for the unsteady Navier– Stokes equations. Journal of Computational Physics, 241 (??):141–167, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113000363.

Eca:2014:PEN

[EH14] L. Eça and M. Hoekstra. A procedure for the estimation of the numerical uncertainty of CFD calculations based on grid refinement studies. Journal of Computational Physics, 262(??):104– 130, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000278. See com-

ment [XS15] and reply [EH15].

Eca:2015:RCP

[EH15] L. Eça and M. Hoekstra. Reply to comment on “A procedure for the estimation of the numerical uncertainty of CFD calculations based on grid refinement studies” (L. Eça and M. Hoekstra, Journal of Computational Physics 262 (2014) 104–130). Journal of Computational Physics, 301(??):487–488, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911500577X. See [EH14,

XS15].

Ekeberg:2014:FPM

[EHA14] Magnus Ekeberg, Tuomo Hartonen, and Erik Aurell. Fast pseudolikelihood maximization for direct-coupling analysis of protein structure from many homologous amino-acid sequences. Journal of Computational Physics, 276(??):341–356, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005178.

ElKacimi:2011:WBI

[EL11] A. El Kacimi and O. Laghrouche. Wavelet based ILU preconditioners for the numerical solution by PUFEM of high frequency elastic wave scattering. Journal of Computational Physics, 230(8):3119–3134, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000301.

Ervik:2014:RMC

[ELM14] Åsmund Ervik, Karl Yngve Lervåg, and Svend Tollak Munkejord. A robust method for calculating interface curvature and normal vectors using an extracted local level set. Journal of Computational Physics, 257(??):259–277, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006712.

ElMoselhy:2012:BIO

[EM12] Tarek A. El Moselhy and Youssef M. Marzouk. Bayesian inference with optimal maps. Journal of Computational Physics, 231(23):7815–7850, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003956.

Edwards:2011:NVT

[EMK11] Jarrod D. Edwards, Jim E. Morel, and Dana A. Knoll. Nonlinear variants of the TR/BDF2 method for thermal radiative diffusion. Journal of Computational Physics, 230(4):1198–1214, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005917.

Elman:2011:FIS

[EMS11] Howard Elman, Milan Mihajlović, and David Silvester. Fast iterative solvers for buoyancy driven flow problems. Journal of Computational Physics, 230(10):3900–3914, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001033.

Evans:2014:MCS

[EMSH14] Thomas M. Evans, Scott W. Mosher, Stuart R. Slattery, and Steven P. Hamilton. A Monte Carlo synthetic-acceleration method for solving the thermal radiation diffusion equation. Journal of Computational Physics, 258(??):338–358, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007195.

Erath:2014:CMT

[EN14a] Christoph Erath and Ramachandran D. Nair. A conservative multi-tracer transport scheme for spectral-element spherical grids. Journal of Computational Physics, 256(??):118–134, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005949.

Erath:2014:RCM

[EN14b] Christoph Erath and Ramachandran D. Nair. Reprint of: a conservative multi-tracer transport scheme for spectralelement spherical grids. Journal of Computational Physics, 271(??):244–260, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114002629.

Erdogan:2011:SNF

[EO11] Utku Erdogan and Turgut Ozis. A smart nonstandard finite difference scheme for second order nonlinear boundary value problems. Journal of Computational Physics, 230(17):6464–6474, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002816.

English:2013:ADI

[EQYF13] R. Elliot English, Linhai Qiu, Yue Yu, and Ronald Fedkiw. An adaptive discretization of incompressible flow using a multitude of moving Cartesian grids. Journal of Computational Physics, 254(??):107–154, December 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005226.

Erath:2013:NCN

[Era13] Christoph Erath. A new conservative numerical scheme for flow problems on unstructured grids and unbounded domains. Journal of Computational Physics, 245(??):476–492, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002398.

Evans:2014:MSM

[ERS14] R. David Evans and Luis A. Ricardez-Sandoval. Multi-scenario modelling of uncertainty in stochastic chemical systems. Journal of Computational Physics, 273(??):374–392, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400374X.

EsedogAlu:2010:DGM

[ERT10] Selim EsedogÂl̄u, Steven Ruuth, and Richard Tsai. Diffusion generated motion using signed distance functions. Journal of Computational Physics, 229(4):1017–1042, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005476.

Elliott:2010:MCT

[ES10] Charles M. Elliott and Björn Stinner. Modeling and computation of two phase geometric biomembranes using surface finite elements. Journal of Computational Physics, 229(18):6585–6612, September 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002755.

Erzincanli:2013:ALE

[ES13a] Belkis Erzincanli and Mehmet Sahin. An arbitrary Lagrangian– Eulerian formulation for solving moving boundary problems with large displacements and rotations. Journal of Computational Physics, 255(??):660–679, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005767.

Evstatiev:2013:VFP

[ES13b] E. G. Evstatiev and B. A. Shadwick. Variational formulation of particle algorithms for kinetic plasma simulations. Journal of Computational Physics, 245(??):376–398, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001800.

Exl:2014:NUF

[ES14] L. Exl and T. Schrefl. Non-uniform FFT for the finite element computation of the micromagnetic scalar potential. Journal of Computational Physics, 270(??):490–505, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002769.

Estep:2010:PEE

[ETW10] Donald Estep, Simon Tavener, and Tim Wildey. A posteriori error estimation and adaptive mesh refinement for a multiscale operator decomposition approach to fluid-solid heat transfer. Journal of Computational Physics, 229(11):4143–4158, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000768.

Elsheikh:2013:ISE

[EWH13] Ahmed H. Elsheikh, Mary F. Wheeler, and Ibrahim Hoteit. An iterative stochastic ensemble method for parameter estimation of subsurface flow models. Journal of Computational Physics, 242(??):696–714, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001113.

Elsheikh:2014:HNS

[EWH14] Ahmed H. Elsheikh, Mary F. Wheeler, and Ibrahim Hoteit. Hybrid nested sampling algorithm for Bayesian model selection applied to inverse subsurface flow problems. Journal of Computational Physics, 258(??):319–337, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006669.

Edwards:2010:DFQ

[EZ10] Michael G. Edwards and Hongwen Zheng. Double-families of quasi-positive Darcy-flux approximations with highly anisotropic tensors on structured and unstructured grids. Journal of Computational Physics, 229(3):594–625, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005294.

Falissard:2013:GMD

[Fal13] F. Falissard. Genuinely multi-dimensional explicit and implicit generalized Shapiro filters for weather forecasting, computational fluid dynamics and aeroacoustics. Journal of Computational Physics, 253(??):344–367, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004713.

Fan:2014:HOW

[Fan14] Ping Fan. High order weighted essentially nonoscillatory WENO-η schemes for hyperbolic conservation laws. Journal of Computational Physics, 269(??):355–385, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002204.

Farrell:2011:AFD

[Far11] P. E. Farrell. The addition of fields on different meshes. Journal of Computational Physics, 230(9):3265–3269, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000532.

Fattebert:2010:ABP

[Fat10] J.-L. Fattebert. Accelerated Block Preconditioned Gradient method for large scale wave functions calculations in Density Functional Theory. Journal of Computational Physics, 229(2):441–452, January 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005270.

Feng:2013:LPR

[FAY^{+}13] Tao Feng, Hengbin An, Xijun Yu, Qin Li, and Rongpei Zhang. On linearization and preconditioning for radiation diffusion coupled to material thermal conduction equations. Journal of Computational Physics, 236(??):28–40, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006894.

Fortmeier:2011:PRI

[FB11] Oliver Fortmeier and H. Martin Bücker. Parallel reinitialization of level set functions on distributed unstructured tetrahedral grids. Journal of Computational Physics, 230(12):4437–4453, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111000878.

Farutin:2014:NSV

[FBM14] Alexander Farutin, Thierry Biben, and Chaouqi Misbah. 3D numerical simulations of vesicle and inextensible capsule dynamics. Journal of Computational Physics, 275(??):539–568, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004884.

Fisher:2013:HOE

[FC13] Travis C. Fisher and Mark H. Carpenter. High-order entropy stable finite difference schemes for nonlinear conservation laws: Finite domains. Journal of Computational Physics, 252(??):518–557, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004385.

Ferronato:2010:FCD

[FCG10] Massimiliano Ferronato, Nicola Castelletto, and Giuseppe Gambolati. A fully coupled 3-D mixed finite element model of Biot consolidation. Journal of Computational Physics, 229(12):4813–4830, June 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001282.

Finch:2013:CHS

[FCH13] Craig Finch, Thomas Clarke, and James J. Hickman. A continuum hard-sphere model of protein adsorption. Journal of Computational Physics, 244(??):212–222, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004196.

Fisher:2013:DCF

[FCN^{+}13] Travis C. Fisher, Mark H. Carpenter, Jan Nordström, Nail K. Yamaleev, and Charles Swanson. Discretely conservative finite-difference formulations for nonlinear conservation laws in split form: Theory and boundary conditions. Journal of Computational Physics, 234(??):353–375, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005712.

Fu:2013:BPM

[FCY13] Zhuo-Jia Fu, Wen Chen, and Hai-Tian Yang. Boundary particle method for Laplace transformed time fractional diffusion equations. Journal of Computational Physics, 235(??):52–66, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006195.

Fisher:2011:BCF

[FCYF11] Travis C. Fisher, Mark H. Carpenter, Nail K. Yamaleev, and Steven H. Frankel. Boundary closures for fourth-order energy stable weighted essentially non-oscillatory finite-difference schemes. Journal of Computational Physics, 230(10):3727– 3752, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111000684.

Fauconnier:2011:SCP

[FD11] D. Fauconnier and E. Dick. On the spectral and conservation properties of nonlinear discretization operators. Journal of Computational Physics, 230(12):4488–4518, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001148.

Fang:2011:TOF

[FDHP11] Jiannong Fang, Marc Diebold, Chad Higgins, and Marc B. Parlange. Towards oscillation-free implementation of the immersed boundary method with spectral-like methods. Journal of Computational Physics, 230(22):8179–8191, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004360.

Fraysse:2012:ETE

[FdVV12] F. Fraysse, J. de Vicente, and E. Valero. The estimation of truncation error by τ-estimation revisited. Journal of Computational Physics, 231(9):3457–3482, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006887.

Friis:2011:FMF

[FE11] Helmer A. Friis and Michael G. Edwards. A family of MPFA finite-volume schemes with full pressure support for the general tensor pressure equation on cell-centered triangular grids. Journal of Computational Physics, 230(1):205–231, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005061.

Fernandes:2012:AEC

[FF12] Ryan I. Fernandes and Graeme Fairweather. An ADI extrapolated Crank–Nicolson orthogonal spline collocation method for nonlinear reaction-diffusion systems. Journal of Computational Physics, 231(19):6248–6267, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001726.

Fei:2013:DIP

[FF13] Fei Fei and Jing Fan. A diffusive information preservation method for small Knudsen number flows. Journal of Computational Physics, 243(??):179–193, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001897.

Favrie:2012:DIM

[FG12] N. Favrie and S. L. Gavrilyuk. Diffuse interface model for compressible fluid — compressible elastic-plastic solid interaction. Journal of Computational Physics, 231(7):2695–2723, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006826.

Farhat:2011:FEM

[FGE11] M. Farhat, S. Guenneau, and S. Enoch. Finite elements modelling of scattering problems for flexural waves in thin plates: Application to elliptic invisibility cloaks, rotators and the mirage effect. Journal of Computational Physics, 230(6):2237–2245, March 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000673X.

Fillion-Gourdeau:2014:SSN

[FGLB14] François Fillion-Gourdeau, Emmanuel Lorin, and André D. Bandrauk. A split-step numerical method for the timedependent Dirac equation in 3-D axisymmetric geometry. Journal of Computational Physics, 272(??):559–587, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002630.

Favrie:2014:TCS

[FGN14] N. Favrie, S. Gavrilyuk, and S. Ndanou. A thermodynamically compatible splitting procedure in hyperelasticity. Journal of Computational Physics, 270(??):300–324, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002381.

Farhat:2012:FFV

[FGR12] Charbel Farhat, Jean-Frédéric Gerbeau, and Arthur Rallu. FIVER: a finite volume method based on exact twophase Riemann problems and sparse grids for multi-material flows with large density jumps. Journal of Computational Physics, 231(19):6360–6379, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002823.

Fang:2012:KSE

[FGZ12] Jun Fang, Xingyu Gao, and Aihui Zhou. A Kohn–Sham equation solver based on hexahedral finite elements. Journal of Computational Physics, 231(8):3166–3180, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000125.

Fernando:2011:DFF

[FH11] Anne M. Fernando and Fang Q. Hu. DGM–FD: a finite difference scheme based on the discontinuous Galerkin method applied to wave propagation. Journal of Computational Physics, 230(12):4871–4898, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001513.

Ferm:2010:AAS

[FHL10] Lars Ferm, Andreas Hellander, and Per Lötstedt. An adaptive algorithm for simulation of stochastic reaction–diffusion processes. Journal of Computational Physics, 229(2):343–360, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005221.

Fares:2011:RBM

[FHMS11] M. Fares, J. S. Hesthaven, Y. Maday, and B. Stamm. The reduced basis method for the electric field integral equation. Journal of Computational Physics, 230(14):5532–5555, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001665.

Filbet:2010:CAP

[FJ10] Francis Filbet and Shi Jin. A class of asymptotic-preserving schemes for kinetic equations and related problems with stiff sources. Journal of Computational Physics, 229(20):7625–7648, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003323.

Fliss:2012:WPL

[FJ12] Sonia Fliss and Patrick Joly. Wave propagation in locally perturbed periodic media (case with absorption): Numerical aspects. Journal of Computational Physics, 231(4):1244–1271, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006000.

Foo:2010:MEP

[FK10] Jasmine Foo and George Em Karniadakis. Multi-element probabilistic collocation method in high dimensions. Journal of Computational Physics, 229(5):1536–1557, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006044.

Friess:2014:SSI

[FK14] Marie Billaud Friess and Samuel Kokh. Simulation of sharp interface multi-material flows involving an arbitrary number of components through an extended five-equation model. Journal of Computational Physics, 273(??):488–519, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003581.

Fairweather:2011:COQ

[FKM11] Graeme Fairweather, Andreas Karageorghis, and Jon Maack. Compact optimal quadratic spline collocation methods for the Helmholtz equation. Journal of Computational Physics, 230 (8):2880–2895, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111000039.

Flyer:2010:RTS

[FL10] Natasha Flyer and Erik Lehto. Rotational transport on a sphere: Local node refinement with radial basis functions. Journal of Computational Physics, 229(6):1954–1969, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006354.

Fidkowski:2011:OBS

[FL11a] Krzysztof J. Fidkowski and Yuxing Luo. Output-based space– time mesh adaptation for the compressible Navier–Stokes equations. Journal of Computational Physics, 230(14):5753–5773, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100221X.

Fornberg:2011:SRG

[FL11b] Bengt Fornberg and Erik Lehto. Stabilization of RBFgenerated finite difference methods for convective PDEs. Journal of Computational Physics, 230(6):2270–2285, March 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006789.

Feng:2013:NEF

[FL13] Naixing Feng and Jianxiong Li. Novel and efficient FDTD implementation of higher-order perfectly matched layer based on ADE method. Journal of Computational Physics, 232 (1):318–326, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112004585.

Farhat:2014:AFE

[FL14] Charbel Farhat and Vinod K. Lakshminarayan. An ALE formulation of embedded boundary methods for tracking boundary layers in turbulent fluid-structure interaction problems. Journal of Computational Physics, 263(??):53–70, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400045X.

Flyer:2012:GRG

[FLB^{+}12] Natasha Flyer, Erik Lehto, Sébastien Blaise, Grady B. Wright, and Amik St-Cyr. A guide to RBF-generated finite differences for nonlinear transport: Shallow water simulations on a sphere. Journal of Computational Physics, 231(11):4078–4095, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000587.

Fraternali:2012:ECB

[FLM12] F. Fraternali, C. D. Lorenz, and G. Marcelli. On the estimation of the curvatures and bending rigidity of membrane networks via a local maximum-entropy approach. Journal of Computational Physics, 231(2):528–540, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005493.

Floriani:2013:SAS

[FM13] E. Floriani and R. Vilela Mendes. A stochastic approach to the solution of magnetohydrodynamic equations. Journal of Computational Physics, 242(??):777–789, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001587.

Fuchs:2010:HOW

[FMM^{+}10] F. G. Fuchs, A. D. McMurry, S. Mishra, N. H. Risebro, and K. Waagan. High order well-balanced finite volume schemes for simulating wave propagation in stratified magnetic atmospheres. Journal of Computational Physics, 229(11):4033–4058, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000616.

Fjordholm:2011:WBE

[FMT11a] Ulrik S. Fjordholm, Siddhartha Mishra, and Eitan Tadmor. Well-balanced and energy stable schemes for the shallow water equations with discontinuous topography. Journal of Computational Physics, 230(14):5587–5609, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002038.

Furuichi:2011:DSF

[FMT11b] Mikito Furuichi, Dave A. May, and Paul J. Tackley. Development of a Stokes flow solver robust to large viscosity jumps using a Schur complement approach with mixed precision arithmetic. Journal of Computational Physics, 230(24):8835–8851, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005274.

Fazioli:2010:SCF

[FN10] Carlo Fazioli and David P. Nicholls. Stable computation of the functional variation of the Dirichlet–Neumann operator. Journal of Computational Physics, 229(3):906–920, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005737.

Fang:2014:OEM

[FN14] Zheng Fang and David P. Nicholls. An operator expansions method for computing Dirichlet–Neumann operators in linear elastodynamics. Journal of Computational Physics, 272(??):266–278, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003076.

Fernandez-Nieto:2014:ENS

[FNGV14] Enrique D. Fernández-Nieto, José M. Gallardo, and Paul Vigneaux. Efficient numerical schemes for viscoplastic avalanches. Part 1: the 1D case. Journal of Computational Physics, 264(??):55–90, May 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000539.

Fernandez-Nieto:2013:MSW

[FNKdLB13] E. D. Fernández-Nieto, E. H. Koné, T. Morales de Luna, and R. Bürger. A multilayer shallow water system for polydisperse sedimentation. Journal of Computational Physics, 238(??):281–314, April 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007395.

Froese:2011:FFD

[FO11] B. D. Froese and A. M. Oberman. Fast finite difference solvers for singular solutions of the elliptic Monge–Ampère equation. Journal of Computational Physics, 230(3):818–834, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005760.

Franquet:2012:RKD

[FP12] Erwin Franquet and Vincent Perrier. Runge–Kutta discontinuous Galerkin method for the approximation of Baer and Nunziato type multiphase models. Journal of Computational Physics, 231(11):4096–4141, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200071X.

Froehle:2014:HOD

[FP14] Bradley Froehle and Per-Olof Persson. A high-order discontinuous Galerkin method for fluid-structure interaction with efficient implicit-explicit time stepping. Journal of Computational Physics, 272(??):455–470, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002216.

Fang:2013:NLP

[FPN^{+}13] F. Fang, C. C. Pain, I. M. Navon, A. H. Elsheikh, J. Du, and D. Xiao. Non-linear Petrov–Galerkin methods for reduced order hyperbolic equations and discontinuous finite element methods. Journal of Computational Physics, 234(??):540–559, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006006.

Formaggia:2013:PCB

[FQV13] Luca Formaggia, Alfio Quarteroni, and Christian Vergara. On the physical consistency between three-dimensional and one-dimensional models in haemodynamics. Journal of Computational Physics, 244(??):97–112, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004342.

Filbet:2013:RVM

[FR13] Francis Filbet and Thomas Rey. A rescaling velocity method for dissipative kinetic equations. Applications to granular media. Journal of Computational Physics, 248(??):177–199, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300288X.

Favier:2014:LBI

[FRP14] Julien Favier, Alistair Revell, and Alfredo Pinelli. A Lattice Boltzmann-Immersed Boundary method to simulate the fluid interaction with moving and slender flexible objects. Journal of Computational Physics, 261(??):145–161, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008607.

Francois:2010:ICV

[FS10a] Marianne M. François and Blair K. Swartz. Interface curvature via volume fractions, heights, and mean values on nonuniform rectangular grids. Journal of Computational Physics, 229(3):527–540, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005749.

Fujimatsu:2010:NIT

[FS10b] Nobuyoshi Fujimatsu and Kojiro Suzuki. New interpolation technique for the CIP method on curvilinear coordinates. Journal of Computational Physics, 229(16):5573–5596, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001580.

Fu:2010:SFD

[FSL10] S. C. Fu, R. M. C. So, and W. W. F. Leung. Stochastic finite difference lattice Boltzmann method for steady incompressible viscous flows. Journal of Computational Physics, 229(17):6084–6103, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002287.

Freytag:2012:SSC

[FSL^{+}12] B. Freytag, M. Steffen, H.-G. Ludwig, S. Wedemeyer-Böhm, W. Schaffenberger, and O. Steiner. Simulations of stellar convection with CO5BOLD. Journal of Computational Physics, 231(3):919–959, February 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005699.

Fan:2014:NSI

[FSTY14] Ping Fan, Yiqing Shen, Baolin Tian, and Chao Yang. A new smoothness indicator for improving the weighted essentially non-oscillatory scheme. Journal of Computational Physics, 269(??):329–354, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114002198.

Fatyanov:2011:HPM

[FT11] Alexey G. Fatyanov and Andrew V. Terekhov. Highperformance modeling acoustic and elastic waves using the parallel Dichotomy Algorithm. Journal of Computational Physics, 230(5):1992–2003, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006613.

Farhat:2012:HDG

[FT12] Charbel Farhat and Jari Toivanen. A hybrid discontinuous Galerkin method for computing the ground state solution of Bose–Einstein condensates. Journal of Computational Physics, 231(14):4709–4722, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001519.

Ferroni:2014:GAD

[FTF14] Francesco Ferroni, Edmund Tarleton, and Steven Fitzgerald. GPU accelerated dislocation dynamics. Journal of Computational Physics, 272(??):619–628, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003283.

Fujimoto:2011:NEP

[Fuj11] Keizo Fujimoto. A new electromagnetic particle-in-cell model with adaptive mesh refinement for high-performance parallel computation. Journal of Computational Physics, 230(23):8508–8526, September 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004608.

Fujioka:2013:CMI

[Fuj13] Hideki Fujioka. A continuum model of interfacial surfactant transport for particle methods. Journal of Computational Physics, 234(??):280–294, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005864.

Fuster:2013:EPF

[Fus13] D. Fuster. An energy preserving formulation for the simulation of multiphase turbulent flows. Journal of Computational Physics, 235(??):114–128, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006304.

Fokkema:2012:SBA

[FvdB12] Jacob T. Fokkema and Peter M. van den Berg. Stretched backgrounds for acoustic scattering models. Journal of Computational Physics, 231(4):1728–1742, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006449.

Fraysse:2013:QPT

[FVR13] F. Fraysse, E. Valero, and G. Rubio. Quasi-a priori truncation error estimation and higher order extrapolation for non-linear partial differential equations. Journal of Computational Physics, 253(??):389–404, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004919.

Fornberg:2011:NMP

[FW11] Bengt Fornberg and J. A. C. Weideman. A numerical methodology for the Painlevé equations. Journal of Computational Physics, 230(15):5957–5973, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002312.

Ferrer:2012:HOD

[FW12] Esteban Ferrer and Richard H. J. Willden. A high order Discontinuous Galerkin–Fourier incompressible 3D Navier–Stokes solver with rotating sliding meshes. Journal of Computational Physics, 231(21):7037–7056, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002215.

Fu:2014:TDP

[FWLP14] Jin Fu, Sheng Wu, Hong Li, and Linda R. Petzold. The time dependent propensity function for acceleration of spatial stochastic simulation of reaction-diffusion systems. Journal of Computational Physics, 274(??):524–549, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400432X.

Fu:2013:TDS

[FWP13] Jin Fu, Sheng Wu, and Linda R. Petzold. Time dependent solution for acceleration of tau-leaping. Journal of Computational Physics, 235(??):446–457, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006390.

Filbet:2013:ILW

[FY13] Francis Filbet and Chang Yang. An inverse Lax–Wendroff method for boundary conditions applied to Boltzmann type models. Journal of Computational Physics, 245(??):43–61, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001939.

Zhang:2011:BPC

[fZlC11] Zheng fang Zhang and Xiao liang Cheng. A boundary piecewise constant level set method for boundary control of eigenvalue optimization problems. Journal of Computational Physics, 230(2):458–473, January 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005498.

Gao:2013:BIF

[Gao13] Junhui Gao. A block interface flux reconstruction method for numerical simulation with high order finite difference scheme. Journal of Computational Physics, 241(??):1–17, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000338.

Gassmann:2011:IHT

[Gas11] Almut Gassmann. Inspection of hexagonal and triangular C-grid discretizations of the shallow water equations. Journal of Computational Physics, 230(7):2706–2721, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000325.

Gorria:2013:DCL

[GAV13] C. Gorria, M. A. Alejo, and L. Vega. Discrete conservation laws and the convergence of long time simulations of the mKdV equation. Journal of Computational Physics, 235(??):274–285, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006481.

Gagarina:2014:VST

[GAvdVB14] E. Gagarina, V. R. Ambati, J. J. W. van der Vegt, and O. Bokhove. Variational space–time (dis)continuous Galerkin method for nonlinear free surface water waves. Journal of Computational Physics, 275(??):459–483, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004422.

Grier:2014:NIT

[GAW^{+}14] Benjamin Grier, Edward Alyanak, Michael White, José Camberos, and Richard Figliola. Numerical integration techniques for discontinuous manufactured solutions. Journal of Computational Physics, 278(??):193–203, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005993.

Gillman:2013:FDS

[GB13] A. Gillman and A. Barnett. A fast direct solver for quasi-periodic scattering problems. Journal of Computational Physics, 248(??):309–322, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002647.

Gabard:2014:FDD

[GB14] G. Gabard and E. J. Brambley. A full discrete dispersion analysis of time-domain simulations of acoustic liners with flow. Journal of Computational Physics, 273(??):310–326, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003362.

Goffin:2013:MEE

[GBB^{+}13] Mark A. Goffin, Christopher M. J. Baker, Andrew G. Buchan, Christopher C. Pain, Matthew D. Eaton, and Paul N. Smith. Minimising the error in eigenvalue calculations involving the Boltzmann transport equation using goal-based adaptivity on unstructured meshes. Journal of Computational Physics, 242(??):726–752, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000089.

Guevel:2011:ADB

[GBC11] Y. Guevel, H. Boutyour, and J. M. Cadou. Automatic detection and branch switching methods for steady bifurcation in fluid mechanics. Journal of Computational Physics, 230(9):3614– 3629, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000866.

Garcia:2014:EVI

[GBNS14] Ferran Garcia, Luca Bonaventura, Marta Net, and Juan Sánchez. Exponential versus IMEX high-order time integrators for thermal convection in rotating spherical shells. Journal of Computational Physics, 264(??):41–54, May 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000606.

Gomar:2014:CFB

[GBS^{+}14] Adrien Gomar, Quentin Bouvy, Frédéric Sicot, Guillaume Dufour, Paola Cinnella, and Benjamin François. Convergence of Fourier-based time methods for turbomachinery wake passing problems. Journal of Computational Physics, 278(??):229–256, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400583X.

Ge:2011:MMB

[GC11] Yongbin Ge and Fujun Cao. Multigrid method based on the transformation-free HOC scheme on nonuniform grids for 2D convection diffusion problems. Journal of Computational Physics, 230(10):4051–4070, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001161.

Gatsonis:2013:UDS

[GCA13] Nikolaos A. Gatsonis, Ryan E. Chamberlin, and Sergey N. Averkin. An unstructured direct simulation Monte Carlo methodology with kinetic-moment inflow and outflow boundary conditions. Journal of Computational Physics, 233(??): 148–174, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911200455X.

Gil:2010:ISP

[GCBH10] A. J. Gil, A. Arranz Carreño, J. Bonet, and O. Hassan. The Immersed Structural Potential Method for haemodynamic applications. Journal of Computational Physics, 229(22):8613–8641, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004420.

Gil:2013:EIS

[GCBH13] A. J. Gil, A. Arranz Carreño, J. Bonet, and O. Hassan. An enhanced Immersed Structural Potential Method for fluid-structure interaction. Journal of Computational Physics, 250(??):178–205, October 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003422.

Ghommem:2014:MDM

[GCE14] Mehdi Ghommem, Victor M. Calo, and Yalchin Efendiev. Mode decomposition methods for flows in high-contrast porous media. A global approach. Journal of Computational Physics, 257(??):400–413, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006451.

Gomez:2013:TDS

[GCFJ13] Hector Gomez, Luis Cueto-Felgueroso, and Ruben Juanes. Three-dimensional simulation of unstable gravity-driven infiltration of water into a porous medium. Journal of Computational Physics, 238(??):217–239, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007498.

Guclu:2014:AHO

[GCH14] Yaman Güçlü, Andrew J. Christlieb, and William N. G. Hitchon. Arbitrarily high order convected scheme solution of the Vlasov–Poisson system. Journal of Computational Physics, 270(??):711–752, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002575.

Grimich:2013:SPH

[GCL13] K. Grimich, P. Cinnella, and A. Lerat. Spectral properties of high-order residual-based compact schemes for unsteady compressible flows. Journal of Computational Physics, 252(??):142–162, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113004282.

Gejadze:2011:CAE

[GCLS11] I. Yu. Gejadze, G. J. M. Copeland, F.-X. Le Dimet, and V. Shutyaev. Computation of the analysis error covariance in variational data assimilation problems with nonlinear dynamics. Journal of Computational Physics, 230(22):7923–7943, September 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001902.

Gazzola:2011:SSM

[GCvRK11] Mattia Gazzola, Philippe Chatelain, Wim M. van Rees, and Petros Koumoutsakos. Simulations of single and multiple swimmers with non-divergence free deforming geometries. Journal of Computational Physics, 230(19):7093–7114, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002737.

Gong:2014:SNS

[GCW14] Yuezheng Gong, Jiaxiang Cai, and Yushun Wang. Some new structure-preserving algorithms for general multi-symplectic formulations of Hamiltonian PDEs. Journal of Computational Physics, 279(??):80–102, December 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006275.

Ge:2013:TFH

[GCZ13] Yongbin Ge, Fujun Cao, and Jun Zhang. A transformationfree HOC scheme and multigrid method for solving the 3D Poisson equation on nonuniform grids. Journal of Computational Physics, 234(??):199–216, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005797.

Godoy:2010:UFL

[GD10] William F. Godoy and Paul E. DesJardin. On the use of flux limiters in the discrete ordinates method for 3D radiation calculations in absorbing and scattering media. Journal of Computational Physics, 229(9):3189–3213, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000045.

Gumerov:2013:EFA

[GD13]  Nail A. Gumerov and Ramani Duraiswami.  Efficient

FMM accelerated vortex methods in three dimensions via the Lamb–Helmholtz decomposition. Journal of Computational Physics, 240(??):310–328, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000521.

Gumerov:2014:MCP

[GD14] Nail A. Gumerov and Ramani Duraiswami. A method to compute periodic sums. Journal of Computational Physics, 272(??):307–326, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003088.

Guardone:2011:RBF

[GDGP11] A. Guardone, D. De Santis, G. Geraci, and M. Pasta. On the relation between finite element and finite volume schemes for compressible flows with cylindrical and spherical symmetry. Journal of Computational Physics, 230(3):680–694, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005577.

Gassner:2011:EOS

[GDHM11] Gregor Gassner, Michael Dumbser, Florian Hindenlang, and Claus-Dieter Munz. Explicit one-step time discretizations for discontinuous Galerkin and finite volume schemes based on local predictors. Journal of Computational Physics, 230(11):4232–4247, May 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005802.

Germann:2011:NSE

[GDW11] N. Germann, M. Dressler, and E. J. Windhab. Numerical solution of an extended White–Metzner model for eccentric Taylor–Couette flow. Journal of Computational Physics, 230(21):7853–7866, September 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004128.

Gambaruto:2010:WSS

[GDY10] A. M. Gambaruto, D. J. Doorly, and T. Yamaguchi. Wall shear stress and near-wall convective transport: Comparisons with vascular remodelling in a peripheral graft anastomosis. Journal of Computational Physics, 229(14):5339–5356, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001488.

Ge:2010:MMF

[Ge10] Yongbin Ge. Multigrid method and fourth-order compact difference discretization scheme with unequal meshsizes for 3D Poisson equation. Journal of Computational Physics, 229(18):6381–6391, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002512.

Gentile:2011:IET

[Gen11] N. A. Gentile. Including the effects of temperature-dependent opacities in the implicit Monte Carlo algorithm. Journal of Computational Physics, 230(12):5100–5114, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100180X. See comments

[Gho17].

Geng:2013:PHO

[Gen13] Weihua Geng. Parallel higher-order boundary integral electrostatics computation on molecular surfaces with curved triangulation. Journal of Computational Physics, 241(??):253–265, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000739.

Gerlinger:2012:MDL

[Ger12] Peter Gerlinger. Multi-dimensional limiting for high-order schemes including turbulence and combustion. Journal of Computational Physics, 231(5):2199–2228, March 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006243.

Gretarsson:2013:FCL

[GF13] Jón Tómas Grétarsson and Ron Fedkiw. Fully conservative leak-proof treatment of thin solid structures immersed in compressible fluids. Journal of Computational Physics, 245(??):160–204, July 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001289.

Grinberg:2013:PMS

[GFK13] Leopold Grinberg, Dmitry A. Fedosov, and George Em Karniadakis. Parallel multiscale simulations of a brain aneurysm. Journal of Computational Physics, 244(??):131–147, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200469X.

Gao:2010:PSA

[GG10] Xinfeng Gao and Clinton P. T. Groth. A parallel solutionadaptive method for three-dimensional turbulent non-premixed combusting flows. Journal of Computational Physics, 229 (9):3250–3275, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110000161.

Gimbutas:2013:FMP

[GG13] Z. Gimbutas and L. Greengard. Fast multi-particle scattering: a hybrid solver for the Maxwell equations in microstructured materials. Journal of Computational Physics, 232(1):22–32, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004184.

Ghiroldi:2014:DMB

[GG14] G. P. Ghiroldi and L. Gibelli. A direct method for the Boltzmann equation based on a pseudo-spectral velocity space discretization. Journal of Computational Physics, 258(??):568– 584, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007316.

Guedeney:2013:NUT

[GGG^{+}13] Thomas Guédeney, Adrien Gomar, François Gallard, Frédéric Sicot, Guillaume Dufour, and Guillaume Puigt. Non-uniform time sampling for multiple-frequency harmonic balance computations. Journal of Computational Physics, 236(??):317–345, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006869.

Gong:2014:IBM

[GGH14] Xiaobo Gong, Zhaoxin Gong, and Huaxiong Huang. An immersed boundary method for mass transfer across permeable moving interfaces. Journal of Computational Physics, 278 (??):148–168, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005798.

Guillen-Gonzalez:2013:LSC

[GGT13] F. Guillén-González and G. Tierra. On linear schemes for a Cahn–Hilliard diffuse interface model. Journal of Computational Physics, 234(??):140–171, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005517.

Guy:2010:ADF

[GH10] Robert D. Guy and David A. Hartenstine. On the accuracy of direct forcing immersed boundary methods with projection methods. Journal of Computational Physics, 229(7):2479–2496, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005798.

Ganesh:2011:FDG

[GH11a] M. Ganesh and S. C. Hawkins. A fully discrete Galerkin method for high frequency exterior acoustic scattering in three dimensions. Journal of Computational Physics, 230 (1):104–125, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110005085.

Gomez:2011:PUS

[GH11b] Hector Gomez and Thomas J. R. Hughes. Provably unconditionally stable, second-order time-accurate, mixed variational methods for phase-field models. Journal of Computational Physics, 230(13):5310–5327, June 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001847.

Gordon:2012:TDS

[GH12a] D. F. Gordon and B. Hafizi. Time dependent Schrödinger equation on arbitrary structured grids: Application to photoionization. Journal of Computational Physics, 231(19):6349–6359, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200263X.

Guclu:2012:HOC

[GH12b] Y. Güçlü and W. N. G. Hitchon. A high order cellcentered semi-Lagrangian scheme for multi-dimensional kinetic simulations of neutral gas flows. Journal of Computational Physics, 231(8):3289–3316, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000290.

Gamba:2014:CSM

[GH14] Irene M. Gamba and Jeffrey R. Haack. A conservative spectral method for the Boltzmann equation with anisotropic scattering and the grazing collisions limit. Journal of Computational Physics, 270(??):40–57, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002228.

Goncharov:2010:KPA

[GHDS10] A. V. Goncharov, G. Hrkac, J. S. Dean, and T. Schrefl. Kronecker product approximation of demagnetizing tensors for micromagnetics. Journal of Computational Physics, 229 (7):2544–2549, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109006743.

Gordon:2014:SRQ

[GHH14] D. F. Gordon, B. Hafizi, and M. H. Helle. Solution of relativistic quantum optics problems using clusters of graphical processing units. Journal of Computational Physics, 267(??):50–62, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400148X.

Greengard:2014:SSW

[GHJ14] Leslie Greengard, Thomas Hagstrom, and Shidong Jiang. The solution of the scalar wave equation in the exterior of a sphere. Journal of Computational Physics, 274(??):191–207, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003945.

Greengard:2014:FDS

[GHL14] Leslie Greengard, Kenneth L. Ho, and June-Yub Lee. A fast direct solver for scattering from periodic structures with multiple material interfaces in two dimensions. Journal of Computational Physics, 258(??):738–751, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007614.

Gershgorin:2010:IFP

[GHM10a] B. Gershgorin, J. Harlim, and A. J. Majda. Improving filtering and prediction of spatially extended turbulent systems with model errors through stochastic parameter estimation. Journal of Computational Physics, 229(1):32–57, January 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005142.

Gershgorin:2010:TMI

[GHM10b] B. Gershgorin, J. Harlim, and A. J. Majda. Test models for improving filtering with model errors through stochastic parameter estimation. Journal of Computational Physics, 229(1):1–31, January 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910900463X.

Gotoh:2012:SCD

[GHM12] T. Gotoh, S. Hatanaka, and H. Miura. Spectral compact difference hybrid computation of passive scalar in isotropic turbulence. Journal of Computational Physics, 231(21):7398–7414, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003701.

Gronskis:2013:IIC

[GHM13] A. Gronskis, D. Heitz, and E. Mémin. Inflow and initial conditions for direct numerical simulation based on adjoint data assimilation. Journal of Computational Physics, 242(??):480– 497, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001290.

Gillman:2014:STE

[GHM14] A. Gillman, S. Hao, and P. G. Martinsson. A simplified technique for the efficient and highly accurate discretization of boundary integral equations in 2D on domains with corners. Journal of Computational Physics, 256(??):214–219, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005937.

Ghosh:2017:CIE

[Gho17] Karabi Ghosh. Comments on “Including the effects of temperature-dependent opacities in the implicit Monte Carlo algorithm” by N. A. Gentile [J. Comput. Phys. 230 (2011) 5100–5114]. Journal of Computational Physics, 330(??):111– 113, February 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999116305915.

See [Gen11].

Gordon:2013:FEN

[GHP13] D. F. Gordon, M. H. Helle, and J. R. Peñano. Fully explicit nonlinear optics model in a particle-in-cell framework. Journal of Computational Physics, 250(??):388–402, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003458.

Ganesh:2012:RBM

[GHS12] M. Ganesh, J. S. Hesthaven, and B. Stamm. A reduced basis method for electromagnetic scattering by multiple particles in three dimensions. Journal of Computational Physics, 231(23):7756–7779, October 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003671.

Gopalan:2013:URM

[GHS13] Harish Gopalan, Stefan Heinz, and Michael K. Stöllinger. A unified RANS-LES model: Computational development, accuracy and cost. Journal of Computational Physics, 249(??):249–274, September 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003203.

Gillespie:2013:CPF

[Gil13] Dirk Gillespie. Computing the partition function, ensemble averages, and density of states for lattice spin systems by sampling the mean. Journal of Computational Physics, 250(??):1–12, October 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003252.

Gorsse:2014:SCS

[GIMT14] Yannick Gorsse, Angelo Iollo, Thomas Milcent, and Haysam Telib. A simple Cartesian scheme for compressible multimaterials. Journal of Computational Physics, 272(??):772–798, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003374.

Giovangigli:2010:MTA

[Gio10] Vincent Giovangigli. Multicomponent transport algorithms for partially ionized mixtures. Journal of Computational Physics, 229(11):4117–4142, June 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110000744.

Guardone:2011:ALE

[GIQ11] A. Guardone, D. Isola, and G. Quaranta. Arbitrary Lagrangian Eulerian formulation for two-dimensional flows using dynamic meshes with edge swapping. Journal of Computational Physics, 230(20):7706–7722, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003998.

Gorsse:2012:SSO

[GITW12] Yannick Gorsse, Angelo Iollo, Haysam Telib, and Lisl Weynans. A simple second order cartesian scheme for compressible Euler flows. Journal of Computational Physics, 231(23):7780–7794, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003877.

Grooms:2011:LIM

[GJ11] Ian Grooms and Keith Julien. Linearly implicit methods for nonlinear PDEs with linear dispersion and dissipation. Journal of Computational Physics, 230(9):3630–3650, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000891.

Gehre:2014:EPN

[GJ14a] Matthias Gehre and Bangti Jin. Expectation propagation for nonlinear inverse problems — with an application to electrical impedance tomography. Journal of Computational Physics, 259(??):513–535, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008097.

Gorji:2014:EPF

[GJ14b] M. Hossein Gorji and Patrick Jenny. An efficient particle Fokker–Planck algorithm for rarefied gas flows. Journal of Computational Physics, 262(??):325–343, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008541.

Goudon:2013:APS

[GJLY13] Thierry Goudon, Shi Jin, Jian-Guo Liu, and Bokai Yan. Asymptotic-preserving schemes for kinetic-fluid modeling of disperse two-phase flows. Journal of Computational Physics, 246(??):145–164, August 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002167.

Grinberg:2010:NDD

[GK10] L. Grinberg and G. E. Karniadakis. A new domain decomposition method with overlapping patches for ultrascale simulations: Application to biological flows. Journal of Computational Physics, 229(15):5541–5563, August 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001907.

Geng:2013:TAB

[GK13] Weihua Geng and Robert Krasny. A treecode-accelerated boundary integral Poisson–Boltzmann solver for electrostatics of solvated biomolecules. Journal of Computational Physics, 247(??):62–78, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002404.

Grun:2014:TPF

[GK14] G. Grün and F. Klingbeil. Two-phase flow with mass density contrast: Stable schemes for a thermodynamic consistent and frame-indifferent diffuse-interface model. Journal of Computational Physics, 257(??):708–725, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007043.

Gretarsson:2011:NSF

[GKF11] Jón Tómas Grétarsson, Nipun Kwatra, and Ronald Fedkiw. Numerically stable fluid-structure interactions between compressible flow and solid structures. Journal of Computational Physics, 230(8):3062–3084, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000131.

Goddeke:2013:EEV

[GKG^{+}13] Dominik Göddeke, Dimitri Komatitsch, Markus Geveler, Dirk Ribbrock, Nikola Rajovic, Nikola Puzovic, and Alex Ramirez. Energy efficiency vs. performance of the numerical solution of PDEs: an application study on a low-power ARM-based cluster. Journal of Computational Physics, 237(??):132–150, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007115.

Graham:2011:QMC

[GKN^{+}11] I. G. Graham, F. Y. Kuo, D. Nuyens, R. Scheichl, and I. H. Sloan. Quasi-Monte Carlo methods for elliptic PDEs with random coefficients and applications. Journal of Computational Physics, 230(10):3668–3694, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000489.

Gumerov:2011:ESP

[GKS^{+}11] Nail A. Gumerov, Alexey V. Karavaev, A. Surjalal Sharma, Xi Shao, and Konstantinos D. Papadopoulos. Efficient spectral and pseudospectral algorithms for 3D simulations of whistler-mode waves in a plasma. Journal of Computational Physics, 230(7):2605–2619, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007126.

Godoy:2012:PJF

[GL12a] William F. Godoy and Xu Liu. Parallel Jacobian-free Newton Krylov solution of the discrete ordinates method with flux limiters for 3D radiative transfer. Journal of Computational Physics, 231(11):4257–4278, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000897.

Greengard:2012:SAI

[GL12b] Leslie Greengard and June-Yub Lee. Stable and accurate integral equation methods for scattering problems with multiple material interfaces in two dimensions. Journal of Computational Physics, 231(6):2389–2395, March 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006905.

Ghossein:2013:RGP

[GL13] Elias Ghossein and Martin Lévesque. Random generation of periodic hard ellipsoids based on molecular dynamics: a computationally-efficient algorithm. Journal of Computational Physics, 253(??):471–490, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004750.

Gerbeau:2014:ALP

[GL14] Jean-Frédéric Gerbeau and Damiano Lombardi. Approximated Lax pairs for the reduced order integration of nonlinear evolution equations. Journal of Computational Physics, 265(??):246–269, May 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000928.

Gorler:2011:GVG

[GLB^{+}11] T. Görler, X. Lapillonne, S. Brunner, T. Dannert, F. Jenko, F. Merz, and D. Told. The global version of the gyrokinetic turbulence code GENE. Journal of Computational Physics, 230(18):7053–7071, August 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003457.

Gong:2011:GAS

[GLC^{+}11] Chunye Gong, Jie Liu, Lihua Chi, Haowei Huang, Jingyue Fang, and Zhenghu Gong. GPU accelerated simulations of 3D deterministic particle transport using discrete ordinates method. Journal of Computational Physics, 230(15):6010– 6022, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002348.

Guermond:2011:EDM

[GLL^{+}11] J.-L. Guermond, J. Léorat, F. Luddens, C. Nore, and A. Ribeiro. Effects of discontinuous magnetic permeability on magnetodynamic problems. Journal of Computational Physics, 230(16):6299–6319, July 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002749.

Guo:2014:NMQ

[GLL14] Z. Guo, P. Lin, and J. S. Lowengrub. A numerical method for the quasi-incompressible Cahn–Hilliard–Navier–Stokes equations for variable density flows with a discrete energy law. Journal of Computational Physics, 276(??):486–507, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005312.

Gallier:2014:FDA

[GLLP14] Stany Gallier, Elisabeth Lemaire, Laurent Lobry, and François Peters. A fictitious domain approach for the simulation of dense suspensions. Journal of Computational Physics, 256(??):367–387, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006207.

Grooms:2014:EKF

[GLM14] Ian Grooms, Yoonsang Lee, and Andrew J. Majda. Ensemble Kalman filters for dynamical systems with unresolved turbulence. Journal of Computational Physics, 273(??):435–452, September 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114004008.

Gejadze:2010:OSE

[GLS10] I. Yu. Gejadze, F.-X. Le Dimet, and V. Shutyaev. On optimal solution error covariances in variational data assimilation problems. Journal of Computational Physics, 229(6):2159–2178, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006573.

Garnaud:2012:RML

[GLSC12] X. Garnaud, L. Lesshafft, P. J. Schmid, and J.-M. Chomaz. A relaxation method for large eigenvalue problems, with an application to flow stability analysis. Journal of Computational Physics, 231(10):3912–3927, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000691.

Guan:2014:SOC

[GLWW14] Zhen Guan, John S. Lowengrub, Cheng Wang, and Steven M. Wise. Second order convex splitting schemes for periodic nonlocal Cahn–Hilliard and Allen–Cahn equations. Journal of Computational Physics, 277(??):48–71, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005476.

Ghosh:2010:FIR

[GM10a] Karabi Ghosh and S. V. G. Menon. Fully implicit 1D radiation hydrodynamics: Validation and verification. Journal of Computational Physics, 229(19):7488–7502, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003463.

Gillman:2010:FAN

[GM10b] A. Gillman and P. G. Martinsson. Fast and accurate numerical methods for solving elliptic difference equations defined on lattices. Journal of Computational Physics, 229(24):9026–9041, December 10, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004171.

Gershgorin:2011:FSE

[GM11a] B. Gershgorin and A. J. Majda. Filtering a statistically exactly solvable test model for turbulent tracers from partial observations. Journal of Computational Physics, 230(4):1602–1638, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000639X.

Gobert:2011:SMP

[GM11b] Christian Gobert and Michael Manhart. Subgrid modelling for particle-LES by Spectrally Optimised Interpolation (SOI). Journal of Computational Physics, 230(21):7796–7820, September 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004013.

Gibou:2012:ESP

[GM12a] Frédéric Gibou and Chohong Min. Efficient symmetric positive definite second-order accurate monolithic solver for fluid/solid interactions. Journal of Computational Physics, 231(8):3246–3263, April 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000307.

Gibou:2012:PSP

[GM12b] Frédéric Gibou and Chohong Min. On the performance of a simple parallel implementation of the ILU–PCG for the Poisson equation on irregular domains. Journal of Computational Physics, 231(14):4531–4536, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001325.

Goyal:2014:AMD

[GM14a] Kavita Goyal and Mani Mehra. An adaptive meshfree diffusion wavelet method for partial differential equations on the sphere. Journal of Computational Physics, 272(??):747–771, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003131.

Grooms:2014:SSQ

[GM14b] Ian Grooms and Andrew J. Majda. Stochastic superparameterization in quasigeostrophic turbulence. Journal of Computational Physics, 271(??):78–98, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006256.

Ghazizadeh:2010:EIF

[GMA10] H. R. Ghazizadeh, M. Maerefat, and A. Azimi. Explicit and implicit finite difference schemes for fractional Cattaneo equation. Journal of Computational Physics, 229(19):7042–7057, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003001.

Galera:2010:TDU

[GMB10] Stéphane Galera, Pierre-Henri Maire, and Jérôme Breil. A twodimensional unstructured cell-centered multi-material ALE scheme using VOF interface reconstruction. Journal of Computational Physics, 229(16):5755–5787, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001956.

Germaine:2013:HOF

[GMC13] Emmanuel Germaine, Laurent Mydlarski, and Luca Cortelezzi. 3DFLUX: a high-order fully three-dimensional flux integral solver for the scalar transport equation. Journal of Computational Physics, 240(??):121–144, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000405.

Guo:2012:IPM

[GMG12] Z. Guo, J. Mi, and P. S. Grant. An implicit parallel multigrid computing scheme to solve coupled thermal-solute phasefield equations for dendrite evolution. Journal of Computational Physics, 231(4):1781–1796, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006486.

Garvie:2010:ERN

[GMT10] Marcus R. Garvie, Philip K. Maini, and Catalin Trenchea. An efficient and robust numerical algorithm for estimating parameters in Turing systems. Journal of Computational Physics, 229(19):7058–7071, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003128.

Guo:2014:PFS

[GMXG14] Z. Guo, J. Mi, S. Xiong, and P. S. Grant. Phase field study of the tip operating state of a freely growing dendrite against convection using a novel parallel multigrid approach. Journal of Computational Physics, 257(??):278–297, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006694.

Garcia:2010:CHO

[GNGAS10] F. Garcia, M. Net, B. García-Archilla, and J. Sánchez. A comparison of high-order time integrators for thermal convection in rotating spherical shells. Journal of Computational Physics, 229(20):7997–8010, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003736.

Gnoffo:2014:GSS

[Gno14] Peter A. Gnoffo. Global series solutions of nonlinear differential equations with shocks using Walsh functions. Journal of Computational Physics, 258(??):650–688, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007304.

Gonnet:2010:UPP

[Gon10] Pedro Gonnet. Using piecewise polynomials for faster potential function evaluation. Journal of Computational Physics, 229(2):313–324, January 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005208.

Gonnet:2012:SNF

[Gon12] Pedro Gonnet. A short note on the fast evaluation of dihedral angle potentials and their derivatives. Journal of Computational Physics, 231(7):2691–2694, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007340.

Goudon:2012:FVS

[GP12a] Thierry Goudon and Martin Parisot. Finite Volume schemes on unstructured grids for non-local models: Application to the simulation of heat transport in plasmas. Journal of Computational Physics, 231(24):8188–8208, October 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004548.

Govindjee:2012:TDD

[GP12b] Sanjay Govindjee and Per-Olof Persson. A time-domain discontinuous Galerkin method for mechanical resonator quality factor computations. Journal of Computational Physics, 231(19):6380–6392, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911200304X.

Gudmundsson:2013:IPC

[GP13] K. Gudmundsson and A. Prosperetti. Improved procedure for the computation of Lamb’s coefficients in the PHYSALIS method for particle simulation. Journal of Computational Physics, 234(??):44–59, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005177.

Ghommem:2013:MDM

[GPCE13] Mehdi Ghommem, Michael Presho, Victor M. Calo, and Yalchin Efendiev. Mode decomposition methods for flows in high-contrast porous media. Global–local approach. Journal of Computational Physics, 253(??):226–238, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004634.

Guermond:2011:EVM

[GPP11] Jean-Luc Guermond, Richard Pasquetti, and Bojan Popov. Entropy viscosity method for nonlinear conservation laws. Journal of Computational Physics, 230(11):4248–4267, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006583.

Gobbi:2010:NSC

[GPS10] Romina Gobbi, Silvia Palpacelli, and Renato Spigler. Numerical solution of certain classes of transport equations in any dimension by Shannon sampling. Journal of Computational Physics, 229(9):3502–3522, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000288.

Gatsonis:2014:SDP

[GPY14] Nikolaos A. Gatsonis, Raffaele Potami, and Jun Yang. A smooth dissipative particle dynamics method for domains with arbitrary-geometry solid boundaries. Journal of Computational Physics, 256(??):441–464, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006037.

Guo:2013:HSL

[GQ13] Wei Guo and Jing-Mei Qiu. Hybrid semi-Lagrangian finite element-finite difference methods for the Vlasov equation. Journal of Computational Physics, 234(??):108–132, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005451.

Grandclement:2010:KSS

[Gra10] Philippe Grandclément. KADATH: a spectral solver for theoretical physics. Journal of Computational Physics, 229(9):

3334–3357, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110000203.

Graille:2014:AMD

[Gra14] B. Graille. Approximation of mono-dimensional hyperbolic systems: a lattice Boltzmann scheme as a relaxation method. Journal of Computational Physics, 266(??):74–88, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001375.

Grigoriu:2010:PMS

[Gri10] Mircea Grigoriu. Probabilistic models for stochastic elliptic partial differential equations. Journal of Computational Physics, 229(22):8406–8429, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000416X.

Grigoriu:2012:MSS

[Gri12a] M. Grigoriu. A method for solving stochastic equations by reduced order models and local approximations. Journal of Computational Physics, 231(19):6495–6513, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003208.

Grinberg:2012:POD

[Gri12b] L. Grinberg. Proper orthogonal decomposition of atomistic flow simulations. Journal of Computational Physics, 231 (16):5542–5556, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112002318.

Griffiths:2013:KWP

[Gri13] Stephen D. Griffiths. Kelvin wave propagation along straight boundaries in C-grid finite-difference models. Journal of Computational Physics, 255(??):639–659, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005846.

Gomez:2014:AEI

[GRS14] Hector Gomez, Alessandro Reali, and Giancarlo Sangalli. Accurate, efficient, and (iso)geometrically flexible collocation methods for phase-field models. Journal of Computational Physics, 262(??):153–171, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008528.

Guidetti:2012:EAT

[GRT12] M. Guidetti, V. Rolando, and R. Tripiccione. Efficient assignment of the temperature set for Parallel Tempering. Journal of Computational Physics, 231(4):1524–1532, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100619X.

Grella:2011:STS

[GS11a] K. Grella and Ch. Schwab. Sparse tensor spherical harmonics approximation in radiative transfer. Journal of Computational Physics, 230(23):8452–8473, September 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004566.

Grote:2011:LNB

[GS11b] Marcus J. Grote and Imbo Sim. Local nonreflecting boundary condition for time-dependent multiple scattering. Journal of Computational Physics, 230(8):3135–3154, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000350.

Garbey:2013:MMD

[GSTB13] M. Garbey, R. Salmon, D. Thanoon, and B. L. Bass. Multiscale modeling and distributed computing to predict cosmesis outcome after a lumpectomy. Journal of Computational Physics, 244(??):321–335, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004354.

Guillet:2011:SMS

[GT11] Thomas Guillet and Romain Teyssier. A simple multigrid scheme for solving the Poisson equation with arbitrary domain boundaries. Journal of Computational Physics, 230(12):4756– 4771, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001343.

Ganesan:2012:ALE

[GT12] Sashikumaar Ganesan and Lutz Tobiska. Arbitrary Lagrangian– Eulerian finite-element method for computation of twophase flows with soluble surfactants. Journal of Computational Physics, 231(9):3685–3702, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000393.

Guba:2014:OBL

[GTSC14] Oksana Guba, Mark Taylor, and Amik St-Cyr. Optimizationbased limiters for the spectral element method. Journal of Computational Physics, 267(??):176–195, June 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001491.

Guidry:2012:ASE

[Gui12] Mike Guidry. Algebraic stabilization of explicit numerical integration for extremely stiff reaction networks. Journal of Computational Physics, 231(16):5266–5288, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002070.

Gurcan:2014:NCM

[Gür14] Ö. D. Gürcan. Numerical computation of the modified plasma dispersion function with curvature. Journal of Computational Physics, 269(??):156–167, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001983.

Guy:2013:HRH

[Guy13] Capdeville Guy. A HLL–Rankine–Hugoniot Riemann solver for complex non-linear hyperbolic problems. Journal of Computational Physics, 251(??):156–193, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003641.

Godfrey:2013:NSR

[GV13] Brendan B. Godfrey and Jean-Luc Vay. Numerical stability of relativistic beam multidimensional PIC simulations employing the esirkepov algorithm. Journal of Computational Physics, 248(??):33–46, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002556.

Godfrey:2014:SNC

[GV14] Brendan B. Godfrey and Jean-Luc Vay. Suppressing the numerical Cherenkov instability in FDTD PIC codes. Journal of Computational Physics, 267(??):1–6, June 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001429.

Gerritsma:2010:TDG

[GvdSVK10] Marc Gerritsma, Jan-Bart van der Steen, Peter Vos, and George Karniadakis. Time-dependent generalized polynomial chaos. Journal of Computational Physics, 229(22):8333–8363, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004134.

Godfrey:2014:NSA

[GVH14] Brendan B. Godfrey, Jean-Luc Vay, and Irving Haber. Numerical stability analysis of the pseudo-spectral analytical time-domain PIC algorithm. Journal of Computational Physics, 258(??):689–704, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007298.

Grenier:2013:ALM

[GVV13] N. Grenier, J.-P. Vila, and P. Villedieu. An accurate low-Mach scheme for a compressible two-fluid model applied to freesurface flows. Journal of Computational Physics, 252(??):1–19, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004312.

Gravemeier:2010:AVM

[GW10] Volker Gravemeier and Wolfgang A. Wall. An algebraic variational multiscale-multigrid method for large-eddy simulation of turbulent variable-density flow at low Mach number. Journal of Computational Physics, 229(17):6047–6070, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002238.

Geng:2011:MMD

[GW11] Weihua Geng and G. W. Wei. Multiscale molecular dynamics using the matched interface and boundary method. Journal of Computational Physics, 230(2):435–457, January 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005346.

Gao:2012:GSS

[GW12] Min Gao and Xiao-Ping Wang. A gradient stable scheme for a phase field model for the moving contact line problem. Journal of Computational Physics, 231(4):1372–1386, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006152.

Gao:2013:CCP

[GW13a] Haiyang Gao and Z. J. Wang. A conservative correction procedure via reconstruction formulation with the Chain-Rule divergence evaluation. Journal of Computational Physics, 232(1):7–13, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004767.

Gao:2013:SSE

[GW13b]  Zhiming Gao and Jiming Wu.  A small stencil and

extremum-preserving scheme for anisotropic diffusion problems on arbitrary 2D and 3D meshes. Journal of Computational Physics, 250(??):308–331, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003446.

Gao:2014:ESP

[GW14] Min Gao and Xiao-Ping Wang. An efficient scheme for a phase field model for the moving contact line problem with variable density and viscosity. Journal of Computational Physics, 272(??):704–718, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003301.

Guzik:2014:IMA

[GWCA14] Stephen M. Guzik, Todd H. Weisgraber, Phillip Colella, and Berni J. Alder. Interpolation methods and the accuracy of lattice-Boltzmann mesh refinement. Journal of Computational Physics, 259(??):461–487, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008012.

Gu:2014:SVS

[GWG14] R. Gu, X. Wang, and M. Gunzburger. Simulating vesiclesubstrate adhesion using two phase field functions. Journal of Computational Physics, 275(??):626–641, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004902.

Gu:2010:AHD

[GWK10] Kai Gu, Charles B. Watkins, and Joel Koplik. Atomistic hybrid DSMC/NEMD method for nonequilibrium multiscale simulations. Journal of Computational Physics, 229(5):1381–1400, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005877.

Guo:2014:PPH

[GXS14] Yan Guo, Tao Xiong, and Yufeng Shi. A positivitypreserving high order finite volume compact-WENO scheme for compressible Euler equations. Journal of Computational Physics, 274(??):505–523, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004537.

Guo:2014:EFD

[GXX14] Ruihan Guo, Yinhua Xia, and Yan Xu. An efficient fully-discrete local discontinuous Galerkin method for the Cahn–Hilliard–Hele–Shaw system. Journal of Computational Physics, 264(??):23–40, May 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000643.

Gao:2012:OSD

[GYL12] Junhui Gao, Zhigang Yang, and Xiaodong Li. An optimized spectral difference scheme for CAA problems. Journal of Computational Physics, 231(14):4848–4866, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001817.

Guo:2013:SDG

[GZQ13] Wei Guo, Xinghui Zhong, and Jing-Mei Qiu. Superconvergence of discontinuous Galerkin and local discontinuous Galerkin methods: Eigen-structure analysis based on Fourier approach. Journal of Computational Physics, 235(??):458–485, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006213.

Gu:2014:ESF

[GZZ14] Shuting Gu, Hui Zhang, and Zhengru Zhang. An energystable finite-difference scheme for the binary fluid-surfactant system. Journal of Computational Physics, 270(??):416–431, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002472.

Hu:2011:SSI

[HA11] X. Y. Hu and N. A. Adams. Scale separation for implicit large eddy simulation. Journal of Computational Physics, 230(19):7240–7249, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003342.

Hicken:2014:PCO

[HA14] J. E. Hicken and J. J. Alonso. PDE-constrained optimization with error estimation and control. Journal of Computational Physics, 263(??):136–150, April 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008589.

Hickel:2010:LEE

[HAD10] Stefan Hickel, Nikolaus A. Adams, and J. Andrzej Domaradzki. Letter to the Editor: On the evolution of dissipation rate and resolved kinetic energy in ALDM simulations of the Taylor–Green flow. Journal of Computational Physics, 229(6):2422–2423, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006366.

Huang:2014:RAW

[HAH14] Chieh-Sen Huang, Todd Arbogast, and Chen-Hui Hung. A reaveraged WENO reconstruction and a third order CWENO scheme for hyperbolic conservation laws. Journal of Computational Physics, 262(??):291–312, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000114.

Hamacher:2010:EPA

[Ham10] K. Hamacher. Efficient perturbation analysis of elastic network models — application to acetylcholinesterase of T. californica. Journal of Computational Physics, 229(19):7309– 7316, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911000330X.

Hajji:2014:EAS

[HAMA14] Mohamed A. Hajji, Qasem M. Al-Mdallal, and Fathi M. Allan. An efficient algorithm for solving higher-order fractional Sturm–Liouville eigenvalue problems. Journal of Computational Physics, 272(??):550–558, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003179.

Hansen:2011:JFN

[Han11] Glen Hansen. A Jacobian-free Newton Krylov method for mortar-discretized thermomechanical contact problems. Journal of Computational Physics, 230(17):6546–6562, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002865.

Huang:2012:ELW

[HAQ12] Chieh-Sen Huang, Todd Arbogast, and Jianxian Qiu. An Eulerian–Lagrangian WENO finite volume scheme for advection problems. Journal of Computational Physics, 231 (11):4028–4052, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112000605.

Harlim:2011:NSF

[Har11] John Harlim. Numerical strategies for filtering partially observed stiff stochastic differential equations. Journal of Computational Physics, 230(3):744–762, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000570X.

Hu:2013:PPM

[HAS13] Xiangyu Y. Hu, Nikolaus A. Adams, and Chi-Wang Shu. Positivity-preserving method for high-order conservative schemes solving compressible Euler equations. Journal of Computational Physics, 242(??):169–180, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000557.

Hittinger:2013:BSA

[HB13] J. A. F. Hittinger and J. W. Banks. Block-structured adaptive mesh refinement algorithms for Vlasov simulation. Journal of Computational Physics, 241(??):118–140, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000740.

Hay:2010:ROM

[HBAP10] A. Hay, J. Borggaard, I. Akhtar, and D. Pelletier. Reducedorder models for parameter dependent geometries based on shape sensitivity analysis. Journal of Computational Physics, 229(4):1327–1352, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005853.

Hashemi:2013:EBF

[HBM13] S. H. Hashemi, B. Boroomand, and B. Movahedian. Exponential basis functions in space and time: a meshless method for 2D time dependent problems. Journal of Computational Physics, 241(??):526–545, May 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000776.

Hedayatrasa:2014:NMW

[HBZL14] Saeid Hedayatrasa, Tinh Quoc Bui, Chuanzeng Zhang, and Chee Wah Lim. Numerical modeling of wave propagation in functionally graded materials using time-domain spectral Chebyshev elements. Journal of Computational Physics, 258(??):381–404, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113007134.

Hart:2011:CRM

[HCD11] E. E. Hart, S. J. Cox, and K. Djidjeli. Compact RBF meshless methods for photonic crystal modelling. Journal of Computational Physics, 230(12):4910–4921, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001537.

Huang:2010:PBE

[HCH^{+}10] Tsung-Ming Huang, Wei-Jen Chang, Yin-Liang Huang, Wen-Wei Lin, Wei-Cheng Wang, and Weichung Wang. Preconditioning bandgap eigenvalue problems in three-dimensional photonic crystals simulations. Journal of Computational Physics, 229(23):8684–8703, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004407.

Huber:2010:LBM

[HCM10] Christian Huber, Bastien Chopard, and Michael Manga. A lattice Boltzmann model for coupled diffusion. Journal of Computational Physics, 229(20):7956–7976, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003712.

Hurd:2010:TIP

[HCP10] Patricia Hurd, Teresa Cusati, and Maurizio Persico. Trajectory integration with potential energy discontinuities. Journal of Computational Physics, 229(6):2109–2116, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006548.

Huang:2010:CLC

[HCS^{+}10] C. C. Huang, A. Chatterji, G. Sutmann, G. Gompper, and R. G. Winkler. Cell-level canonical sampling by velocity scaling for multiparticle collision dynamics simulations. Journal of Computational Physics, 229(1):168–177, January 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005166.

Huang:2011:IPI

[HCS11] Wei-Xi Huang, Cheong Bong Chang, and Hyung Jin Sung. An improved penalty immersed boundary method for fluid-flexible body interaction. Journal of Computational Physics, 230(12):5061–5079, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001781.

Huang:2012:TDS

[HCS12] Wei-Xi Huang, Cheong Bong Chang, and Hyung Jin Sung. Three-dimensional simulation of elastic capsules in shear flow by the penalty immersed boundary method. Journal of Computational Physics, 231(8):3340–3364, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000277.

Harness:2013:NFM

[HD13] Y. Harness and A. Ditkowski. The null-field method: a reconstruction kernel approach. Journal of Computational Physics, 248(??):127–146, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300260X.

Haverkort:2012:BVF

[HdBK12] J. W. Haverkort, H. J. de Blank, and B. Koren. The Brunt–Väisälä frequency of rotating tokamak plasmas. Journal of Computational Physics, 231(3):981–1001, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001598.

Hatch:2012:ACS

[HdCNT12] D. R. Hatch, D. del Castillo-Negrete, and P. W. Terry. Analysis and compression of six-dimensional gyrokinetic datasets using higher order singular value decomposition. Journal of Computational Physics, 231(11):4234–4256, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000861.

Hornikx:2012:MDF

[HDD12] Maarten Hornikx, Wim De Roeck, and Wim Desmet. A multi-domain Fourier pseudospectral time-domain method for the linearized Euler equations. Journal of Computational Physics, 231(14):4759–4774, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001568.

Hossan:2014:HII

[HDD14] Mohammad Robiul Hossan, Robert Dillon, and Prashanta Dutta. Hybrid immersed interface-immersed boundary methods for AC dielectrophoresis. Journal of Computational Physics, 270(??):640–659, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002757.

Hu:2013:GMS

[HDHZ13] Weipeng Hu, Zichen Deng, Songmei Han, and Wenrong Zhang. Generalized multi-symplectic integrators for a class of Hamiltonian nonlinear wave PDEs. Journal of Computational Physics, 235(??):394–406, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006353.

Hu:2011:BTA

[HDPF11] Heng Hu, Noureddine Damil, and Michel Potier-Ferry. A bridging technique to analyze the influence of boundary conditions on instability patterns. Journal of Computational Physics, 230(10):3753–3764, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000775.

He:2013:BSM

[He13] L. He. Block-spectral mapping for multi-scale solution. Journal of Computational Physics, 250(??):13–26, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003288.

Hejranfar:2014:IHO

[HE14] Kazem Hejranfar and Eslam Ezzatneshan. Implementation of a high-order compact finite-difference lattice Boltzmann method in generalized curvilinear coordinates. Journal of Computational Physics, 267(??):28–49, June 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001508.

Helsing:2011:ECA

[Hel11a] Johan Helsing. The effective conductivity of arrays of squares: Large random unit cells and extreme contrast ratios. Journal of Computational Physics, 230(20):7533–7547, August 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003433.

Helsing:2011:ECR

[Hel11b] Johan Helsing. The effective conductivity of random checkerboards. Journal of Computational Physics, 230(4):1171–1181, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005899.

Hennig:2013:FPT

[Hen13] Jörg Hennig. Fully pseudospectral time evolution and its application to [1 + 1] dimensional physical problems. Journal of Computational Physics, 235(??):322–333, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006432.

Herrmann:2010:PEI

[Her10] M. Herrmann. A parallel Eulerian interface tracking/ Lagrangian point particle multi-scale coupling procedure. Journal of Computational Physics, 229(3):745–759, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005543.

Husain:2010:SAA

[HF10] S. Z. Husain and J. M. Floryan. Spectrally-accurate algorithm for moving boundary problems for the Navier–Stokes equations. Journal of Computational Physics, 229(6):2287–2313, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006640.

Hosseini:2011:PBC

[HF11] S. Majid Hosseini and James J. Feng. Pressure boundary conditions for computing incompressible flows with SPH. Journal of Computational Physics, 230(19):7473–7487, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003779.

Hrycak:2010:PFR

[HG10] Tomasz Hrycak and Karlheinz Gröchenig. Pseudospectral Fourier reconstruction with the modified Inverse Polynomial Reconstruction Method. Journal of Computational Physics, 229(3):933–946, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005786.

Helgadottir:2011:PBS

[HG11] Ásdís Helgadóttir and Frédéric Gibou. A Poisson–Boltzmann solver on irregular domains with Neumann or Robin boundary conditions on non-graded adaptive grid. Journal of Computational Physics, 230(10):3830–3848, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000921.

Heath:2012:DGM

[HGMM12] R. E. Heath, I. M. Gamba, P. J. Morrison, and C. Michler. A discontinuous Galerkin method for the Vlasov–Poisson system. Journal of Computational Physics, 231(4):1140–1174, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005523.

Hagstrom:2014:DAB

[HGRB14] Thomas Hagstrom, Dan Givoli, Daniel Rabinovich, and Jacobo Bielak. The Double Absorbing Boundary method. Journal of Computational Physics, 259(??):220–241, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007870.

Happenhofer:2013:LMN

[HGSK^{+}13] Natalie Happenhofer, Hannes Grimm-Strele, Friedrich Kupka, Bernhard Löw-Baselli, and Herbert Muthsam. A low Mach number solver: Enhancing applicability. Journal of Computational Physics, 236(??):96–118, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006572.

He:2011:LSF

[HGW11] Qiaolin He, Roland Glowinski, and Xiao-Ping Wang. A least-squares/finite element method for the numerical solution of the Navier–Stokes–Cahn–Hilliard system modeling the motion of the contact line. Journal of Computational Physics, 230(12):4991–5009, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001653.

Gao:2011:CFD

[hGzS11] Guang hua Gao and Zhi zhong Sun. A compact finite difference scheme for the fractional sub-diffusion equations. Journal of Computational Physics, 230(3):586–595, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005516.

Gao:2013:FDA

[hGzS13] Guang hua Gao and Zhi zhong Sun. The finite difference approximation for a class of fractional sub-diffusion equations on a space unbounded domain. Journal of Computational Physics, 236(??):443–460, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006882.

Gao:2012:FDS

[hGzSnZ12] Guang hua Gao, Zhi zhong Sun, and Ya nan Zhang. A finite difference scheme for fractional sub-diffusion equations on an unbounded domain using artificial boundary conditions. Journal of Computational Physics, 231(7):2865–2879, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111007492.

Gao:2014:NFN

[hGzSwZ14] Guang hua Gao, Zhi zhong Sun, and Hong wei Zhang. A new fractional numerical differentiation formula to approximate the Caputo fractional derivative and its applications. Journal of Computational Physics, 259(??):33–50, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007791.

Hagstrom:2012:GSH

[HH12a] Thomas Hagstrom and George Hagstrom. Grid stabilization of high-order one-sided differencing II: Second-order wave equations. Journal of Computational Physics, 231(23):7907–7931, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004172.

Homma:2012:NMT

[HH12b] Yuki Homma and Akiyoshi Hatayama. Numerical modeling of thermal force in a plasma for test-ion transport simulation based on Monte Carlo Binary Collision Model. Journal of Computational Physics, 231(8):3211–3227, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000046.

Homma:2013:NMT

[HH13] Yuki Homma and Akiyoshi Hatayama. Numerical modeling of the thermal force in a plasma for test-ion transport simulation based on a Monte Carlo Binary Collision Model (II) — thermal forces due to temperature gradients parallel and perpendicular to the magnetic field. Journal of Computational Physics, 250(??):206–223, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003331.

Han:2014:AMR

[HHA14] L. H. Han, X. Y. Hu, and N. A. Adams. Adaptive multiresolution method for compressible multi-phase flows with sharp interface model and pyramid data structure. Journal of Computational Physics, 262(??):131–152, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000230.

Heydari:2014:CMS

[HHGC14] M. H. Heydari, M. R. Hooshmandasl, F. M. Maalek Ghaini, and C. Cattani. A computational method for solving stochastic Itô–Volterra integral equations based on stochastic operational matrix for generalized hat basis functions. Journal of Computational Physics, 270(??):402–415, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002514.

Hou:2013:MMI

[HHH13] T. Y. Hou, X. Hu, and F. Hussain. Multiscale modeling of incompressible turbulent flows. Journal of Computational Physics, 232(1):383–396, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004755.

Hintermuller:2013:AFE

[HHK13] M. Hintermüller, M. Hinze, and C. Kahle. An adaptive finite element Moreau–Yosida-based solver for a coupled Cahn–Hilliard/Navier–Stokes system. Journal of Computational Physics, 235(??):810–827, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005992.

Hartmann:2011:ABE

[HHL11] Ralf Hartmann, Joachim Held, and Tobias Leicht. Adjointbased error estimation and adaptive mesh refinement for the RANS and k–ω turbulence model equations. Journal of Computational Physics, 230(11):4268–4284, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005826.

Huang:2014:MCA

[HHL14] Haibo Huang, Jun-Jie Huang, and Xi-Yun Lu. A massconserving axisymmetric multiphase lattice Boltzmann method and its application in simulation of bubble rising. Journal of Computational Physics, 269(??):386–402, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002095.

Hao:2013:HMB

[HHS^{+}13] Wenrui Hao, Jonathan D. Hauenstein, Chi-Wang Shu, Andrew J. Sommese, Zhiliang Xu, and Yong-Tao Zhang. A homotopy method based on WENO schemes for solving steady state problems of hyperbolic conservation laws. Journal of Computational Physics, 250(??):332–346, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300332X.

Hicken:2012:OEE

[Hic12] J. E. Hicken. Output error estimation for summationby-parts finite-difference schemes. Journal of Computational Physics, 231(9):3828–3848, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000617.

Higdon:2013:PFD

[Hig13] Robert L. Higdon. Pressure forcing and dispersion analysis for discontinuous Galerkin approximations to oceanic fluid flows. Journal of Computational Physics, 249(??):36–66, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002933.

Hajibeygi:2011:AIM

[HJ11] Hadi Hajibeygi and Patrick Jenny. Adaptive iterative multiscale finite volume method. Journal of Computational Physics, 230(3):628–643, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005541.

Hall:2014:MMC

[HJBB14] Clifford Hall, Weixiao Ji, and Estela Blaisten-Barojas. The Metropolis Monte Carlo method with CUDA enabled Graphic Processing Units. Journal of Computational Physics, 258(??):871–879, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007626.

Hong:2014:EDS

[HJK14] Jialin Hong, Lihai Ji, and Linghua Kong. Energy-dissipation splitting finite-difference time-domain method for Maxwell equations with perfectly matched layers. Journal of Computational Physics, 269(??):201–214, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400206X.

Hong:2014:SMS

[HJZ14] Jialin Hong, Lihai Ji, and Liying Zhang. A stochastic multisymplectic scheme for stochastic Maxwell equations with additive noise. Journal of Computational Physics, 268(??):255– 268, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001818.

Houim:2011:LDT

[HK11] Ryan W. Houim and Kenneth K. Kuo. A low-dissipation and time-accurate method for compressible multi-component flow with variable specific heat ratios. Journal of Computational Physics, 230(23):8527–8553, September 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004621.

Houim:2013:GFM

[HK13] Ryan W. Houim and Kenneth K. Kuo. A ghost fluid method for compressible reacting flows with phase change. Journal of Computational Physics, 235(??):865–900, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005530.

Helsing:2014:EKS

[HK14] Johan Helsing and Anders Karlsson. An explicit kernelsplit panel-based Nyström scheme for integral equations on axially symmetric surfaces. Journal of Computational Physics, 272(??):686–703, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003295.

Hajibeygi:2011:HFM

[HKJ11] Hadi Hajibeygi, Dimitris Karvounis, and Patrick Jenny. A hierarchical fracture model for the iterative multiscale finite volume method. Journal of Computational Physics, 230 (24):8729–8743, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111005080.

Huang:2010:NAM

[HKL10] Weizhang Huang, Lennard Kamenski, and Jens Lang. A new anisotropic mesh adaptation method based upon hierarchical a posteriori error estimates. Journal of Computational Physics, 229(6):2179–2198, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006585.

Hu:2014:IBM

[HKL14] Wei-Fan Hu, Yongsam Kim, and Ming-Chih Lai. An immersed boundary method for simulating the dynamics of threedimensional axisymmetric vesicles in Navier–Stokes flows. Journal of Computational Physics, 257(??):670–686, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006943.

Ha:2013:IWE

[HKLY13] Youngsoo Ha, Chang Ho Kim, Yeon Ju Lee, and Jungho Yoon. An improved weighted essentially non-oscillatory scheme with a new smoothness indicator. Journal of Computational Physics, 232(1):68–86, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003233.

Hejranfar:2012:PCB

[HKM12] Kazem Hejranfar and Ramin Kamali-Moghadam. Preconditioned characteristic boundary conditions for solution of the preconditioned Euler equations at low Mach number flows. Journal of Computational Physics, 231(12):4384–4402, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000721.

Hellander:2011:FSM

[HL11] Stefan Hellander and Per Lötstedt. Flexible single molecule simulation of reaction-diffusion processes. Journal of Computational Physics, 230(10):3948–3965, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001094.

He:2012:NCT

[HL12a] Yinnian He and Jian Li. Numerical comparisons of timespace iterative method and spatial iterative methods for the stationary Navier–Stokes equations. Journal of Computational Physics, 231(20):6790–6800, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003130.

Hoppe:2012:ANC

[HL12b] R. H. W. Hoppe and C. Linsenmann. An adaptive Newton continuation strategy for the fully implicit finite element immersed boundary method. Journal of Computational Physics, 231(14):4676–4693, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001416.

Hellander:2014:LEE

[HLDP14] Andreas Hellander, Michael J. Lawson, Brian Drawert, and Linda Petzold. Local error estimates for adaptive simulation of the reaction-diffusion master equation via operator splitting. Journal of Computational Physics, 266(??):89–100, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000989.

Hua:2011:ELP

[HLLW11] Jinsong Hua, Ping Lin, Chun Liu, and Qi Wang. Energy law preserving C^{0} finite element schemes for phase field models in two-phase flow computations. Journal of Computational Physics, 230(19):7115–7131, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003172.

Harris:2011:FFV

[HLM11] Lucas M. Harris, Peter H. Lauritzen, and Rashmi Mittal. A flux-form version of the conservative semi-Lagrangian multitracer transport scheme (CSLAM) on the cubed sphere grid. Journal of Computational Physics, 230(4):1215–1237, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006005.

He:2014:NSM

[HLS14] Ying He, Peijun Li, and Jie Shen. A new spectral method for numerical solution of the unbounded rough surface scattering problem. Journal of Computational Physics, 275(??):608–625, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005191.

Hicken:2013:ACA

[HLSO13] J. E. Hicken, J. Li, O. Sahni, and A. A. Oberai. Adjoint consistency analysis of residual-based variational multiscale methods. Journal of Computational Physics, 255(??):396–406, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005366.

Han:2010:AGS

[HLT10a] Ee Han, Jiequan Li, and Huazhong Tang. An adaptive GRP scheme for compressible fluid flows. Journal of Computational Physics, 229(5):1448–1466, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005993.

Hu:2010:RHO

[HLT10b] Guanghui Hu, Ruo Li, and Tao Tang. A robust highorder residual distribution type scheme for steady Euler equations on unstructured grids. Journal of Computational Physics, 229(5):1681–1697, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006135.

Hochman:2013:URF

[HLW13] Amit Hochman, Yehuda Leviatan, and Jacob K. White. On the use of rational-function fitting methods for the solution of 2D Laplace boundary-value problems. Journal of Computational Physics, 238(??):337–358, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004615.

Huang:2011:IPD

[HLY11] Yunqing Huang, Jichun Li, and Wei Yang. Interior penalty DG methods for Maxwell’s equations in dispersive media. Journal of Computational Physics, 230(12):4559–4570, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001203.

Huang:2012:RND

[HLY12] Yunqing Huang, Hailiang Liu, and Nianyu Yi. Recovery of normal derivatives from the piecewise L^{2} projection. Journal of Computational Physics, 231(4):1230–1243, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005845.

Huang:2011:SMF

[HLYS11] Yunqing Huang, Jichun Li, Wei Yang, and Shuyu Sun. Superconvergence of mixed finite element approximations to 3-D Maxwell’s equations in metamaterials. Journal of Computational Physics, 230(22):8275–8289, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004530.

Huang:2013:EAG

[HLZ13] Kai Huang, Peijun Li, and Hongkai Zhao. An efficient algorithm for the generalized Foldy–Lax formulation. Journal of Computational Physics, 234(??):376–398, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005724.

Hon:2014:CBP

[HLZ14] Sean Y. Hon, Shingyu Leung, and Hongkai Zhao. A cell based particle method for modeling dynamic interfaces. Journal of Computational Physics, 272(??):279–306, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002952.

Hokpunna:2010:CFO

[HM10] Arpiruk Hokpunna and Michael Manhart. Compact fourthorder finite volume method for numerical solutions of Navier– Stokes equations on staggered grids. Journal of Computational Physics, 229(20):7545–7570, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003141.

Haber:2011:FMS

[HM11] Eldad Haber and Scott MacLachlan. A fast method for the solution of the Helmholtz equation. Journal of Computational Physics, 230(12):4403–4418, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000337.

Huan:2013:SBO

[HM13] Xun Huan and Youssef M. Marzouk. Simulation-based optimal Bayesian experimental design for nonlinear systems. Journal of Computational Physics, 232(1):288–317, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004597.

Heyns:2013:WCF

[HMHO13] Johan A. Heyns, Arnaud G. Malan, Thomas M. Harms, and Oliver F. Oxtoby. A weakly compressible free-surface flow solver for liquid-gas systems using the volume-of-fluid approach. Journal of Computational Physics, 240(??):145–157, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000533.

Heys:2010:WLS

[HMM^{+}10] J. J. Heys, T. A. Manteuffel, S. F. McCormick, M. Milano, J. Westerdale, and M. Belohlavek. Weighted least-squares finite elements based on particle imaging velocimetry data. Journal of Computational Physics, 229(1):107–118, January 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005014.

Harlim:2014:EKF

[HMM14] John Harlim, Adam Mahdi, and Andrew J. Majda. An ensemble Kalman filter for statistical estimation of physics constrained nonlinear regression models. Journal of Computational Physics, 257(??):782–812, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007018.

Huang:2011:DGB

[HMOH11] Bormin Huang, Jarno Mielikainen, Hyunjong Oh, and Hung-Lung Allen Huang. Development of a GPU-based highperformance radiative transfer model for the Infrared Atmospheric Sounding Interferometer (IASI). Journal of Computational Physics, 230(6):2207–2221, March 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000505X.

Hartmann:2010:CRE

[HMS10] Daniel Hartmann, Matthias Meinke, and Wolfgang Schröder. The constrained reinitialization equation for level set methods. Journal of Computational Physics, 229(5):1514–1535, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006032.

Haselbacher:2010:STA

[HNMM10] A. Haselbacher, F. M. Najjar, L. Massa, and R. D. Moser. Slow-time acceleration for modeling multiple-time-scale problems. Journal of Computational Physics, 229(2):325–342, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910900521X.

He:2012:ESS

[HNS12] Ying He, David P. Nicholls, and Jie Shen. An efficient and stable spectral method for electromagnetic scattering from a layered periodic structure. Journal of Computational Physics, 231(8):3007–3022, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006334.

Henniger:2010:HOA

[HOK10] R. Henniger, D. Obrist, and L. Kleiser. High-order accurate solution of the incompressible Navier–Stokes equations on massively parallel computers. Journal of Computational Physics, 229(10):3543–3572, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000306.

Hagan:2013:CMT

[HP13] Jonathan Hagan and Janis Priede. Capacitance matrix technique for avoiding spurious eigenmodes in the solution of hydrodynamic stability problems by Chebyshev collocation method. Journal of Computational Physics, 238(??):210–216, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007437.

Hammer:2014:DNP

[HPA14a] René Hammer, Walter Pötz, and Anton Arnold. A dispersion and norm preserving finite difference scheme with transparent boundary conditions for the Dirac equation in (1 + 1)D. Journal of Computational Physics, 256(??):728–747, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300627X.

Hammer:2014:SCR

[HPA14b] René Hammer, Walter Pötz, and Anton Arnold. Single-cone real-space finite difference scheme for the time-dependent Dirac equation. Journal of Computational Physics, 265(??):50–70, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000552.

Hirshman:2010:BPB

[HPLS10] S. P. Hirshman, K. S. Perumalla, V. E. Lynch, and R. Sanchez. BCYCLIC: a parallel block tridiagonal matrix cyclic solver. Journal of Computational Physics, 229(18):6392–6404, September 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002536.

Hill:2010:EHW

[HPOM10] D. J. Hill, D. Pullin, M. Ortiz, and D. Meiron. An Eulerian hybrid WENO centered-difference solver for elastic-plastic solids. Journal of Computational Physics, 229(24):9053–9072, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004675.

Harbrecht:2013:CTB

[HPS13] Helmut Harbrecht, Michael Peters, and Markus Siebenmorgen. Combination technique based k-th moment analysis of elliptic problems with random diffusion. Journal of Computational Physics, 252(??):128–141, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004373.

Harder:2013:FMH

[HPV13] Christopher Harder, Diego Paredes, and Frédéric Valentin. A family of Multiscale Hybrid-Mixed finite element methods for the Darcy equation with rough coefficients. Journal of Computational Physics, 245(??):107–130, July 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001976.

Huang:2010:MPL

[HQL^{+}10] Heyu Huang, Xiaochao Qu, Jimin Liang, Xiaowei He, Xueli Chen, Da’an Yang, and Jie Tian. A multi-phase level set framework for source reconstruction in bioluminescence tomography. Journal of Computational Physics, 229(13):5246–5256, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001609.

Hansbo:2010:LNF

[HR10] Peter Hansbo and Thomas Rylander. A linear nonconforming finite element method for Maxwell’s equations in two dimensions. Part I: Frequency domain. Journal of Computational Physics, 229(18):6534–6547, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002706.

Hejazialhosseini:2010:HOF

[HRBK10] Babak Hejazialhosseini, Diego Rossinelli, Michael Bergdorf, and Petros Koumoutsakos. High order finite volume methods on wavelet-adapted grids with local time-stepping on multicore architectures for the simulation of shock-bubble interactions. Journal of Computational Physics, 229(22):8364–8383, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004146.

Hejlesen:2013:HOS

[HRCW13] Mads Mølholm Hejlesen, Johannes Tophøj Rasmussen, Philippe Chatelain, and Jens Honoré Walther. A high order solver for the unbounded Poisson equation. Journal of Computational Physics, 252(??):458–467, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004324.

Hachem:2010:SFE

[HRK^{+}10] E. Hachem, B. Rivaux, T. Kloczko, H. Digonnet, and T. Coupez. Stabilized finite element method for incompressible flows with high Reynolds number. Journal of Computational Physics, 229(23):8643–8665, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004237.

Helzel:2011:UCT

[HRT11] Christiane Helzel, James A. Rossmanith, and Bertram Taetz. An unstaggered constrained transport method for the 3D ideal magnetohydrodynamic equations. Journal of Computational Physics, 230(10):3803–3829, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100091X.

Haeri:2013:NIF

[HS13a] S. Haeri and J. S. Shrimpton. A new implicit fictitious domain method for the simulation of flow in complex geometries with heat transfer. Journal of Computational Physics, 237(??):21–45, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007322.

Hill:2013:SMF

[HS13b] Ryan N. Hill and Mikhail Shashkov. The Symmetric Momentof-Fluid interface reconstruction algorithm. Journal of Computational Physics, 249(??):180–184, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003227.

He:2011:ELR

[HSD11] J. He, J. Sætrom, and L. J. Durlofsky. Enhanced linearized reduced-order models for subsurface flow simulation. Journal of Computational Physics, 230(23):8313–8341, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003561.

He:2015:CMC

[HSK^{+}15] Wenkui He, Haibing Shao, Olaf Kolditz, Wenqing Wang, and Thomas Kalbacher. Comments on “A mass-conservative switching algorithm for modeling fluid flow in variably saturated porous media, K. Sadegh Zadeh, Journal of Computational Physics, 230 (2011)”. Journal of Computational Physics, 295(??):815–820, August 15, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911500265X. See [Zad11].

Haupt:2013:FSE

[HSN13] L. Haupt, J. Stiller, and W. E. Nagel. A fast spectral element solver combining static condensation and multigrid techniques. Journal of Computational Physics, 255(??):384–395, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005251.

Hou:2013:WFS

[HSWZ13] Songming Hou, Peng Song, Liqun Wang, and Hongkai Zhao. A weak formulation for solving elliptic interface problems without body fitted grid. Journal of Computational Physics, 249(??):80–95, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002908.

Huang:2010:GFL

[HSZ10] Kai Huang, Knut Solna, and Hongkai Zhao. Generalized Foldy–Lax formulation. Journal of Computational Physics, 229(12):4544–4553, June 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001014.

Hellinger:2010:LRC

[HT10a] Petr Hellinger and Pavel M. Trávnícek. Langevin representation of Coulomb collisions for bi-Maxwellian plasmas. Journal of Computational Physics, 229(14):5432–5439, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001853.

Teng:2010:EBC

[hT10b] Zhen huan Teng. Exact boundary conditions for the initial value problem of convex conservation laws. Journal of Computational Physics, 229(10):3792–3801, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000513.

Horng:2012:EMP

[HT12] Tzyy-Leng Horng and Chun-Hao Teng. An error minimized pseudospectral penalty direct Poisson solver. Journal of Computational Physics, 231(6):2498–2509, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100698X.

Hanzlikova:2014:LFI

[HT14] N. Hanzlikova and M. M. Turner. Leap frog integrator modifications in highly collisional particle-in-cell codes. Journal of Computational Physics, 268(??):355–362, July 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001995.

Hiemstra:2014:HOG

[HTHG14] R. R. Hiemstra, D. Toshniwal, R. H. M. Huijsmans, and M. I. Gerritsma. High order geometric methods with exact conservation properties. Journal of Computational Physics, 257(??):1444–1471, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006414.

Hu:2013:AFV

[Hu13] Guanghui Hu. An adaptive finite volume method for 2D steady Euler equations with WENO reconstruction. Journal of Computational Physics, 252(??):591–605, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004786.

Huang:2014:SPP

[Hua14] Weizhang Huang. Sign-preserving of principal eigenfunctions in P1 finite element approximation of eigenvalue problems of second-order elliptic operators. Journal of Computational Physics, 274(??):230–244, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004197.

Huthwaite:2014:AFE

[Hut14]  Peter Huthwaite.  Accelerated finite element elastody-

namic simulations using the GPU. Journal of Computational Physics, 257(??):687–707, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006931.

Hoitinga:2011:DGF

[HvB11] W. Hoitinga and E. H. van Brummelen. A discontinuous Galerkin finite-element method for a 1D prototype of the Boltzmann equation. Journal of Computational Physics, 230(15):6115–6135, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002488.

Hecht:2013:OCF

[HVD13] Frédéric Hecht, Pascal Ventura, and Pierre Dufilié. Original coupled FEM/BIE numerical model for analyzing infinite periodic surface acoustic wave transducers. Journal of Computational Physics, 246(??):265–274, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002210.

Holland:2012:CDA

[HVT12] Bryan W. Holland, Shaghayegh Vafaei, and Bruno Tomberli. Computer data analysis of the oscillating forward-reverse method. Journal of Computational Physics, 231(11):4355– 4364, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001076.

Hesford:2010:FMM

[HW10] Andrew J. Hesford and Robert C. Waag. The fast multipole method and Fourier convolution for the solution of acoustic scattering on regular volumetric grids. Journal of Computational Physics, 229(21):8199–8210, October 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004183.

Hesford:2011:RRA

[HW11]  Andrew J. Hesford and Robert C. Waag.  Reduced-

rank approximations to the far-field transform in the gridded fast multipole method. Journal of Computational Physics, 230(10):3656–3667, May 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001057.

Huang:2014:IBT

[HW14a] Rongzong Huang and Huiying Wu. An immersed boundarythermal lattice Boltzmann method for solid-liquid phase change. Journal of Computational Physics, 277(??):305– 319, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005749.

Huang:2014:MMR

[HW14b] Rongzong Huang and Huiying Wu. A modified multiplerelaxation-time lattice Boltzmann model for convectiondiffusion equation. Journal of Computational Physics, 274 (??):50–63, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114004057.

Hyman:2014:SGE

[HW14c] Jeffrey D. Hyman and C. Larrabee Winter. Stochastic generation of explicit pore structures by thresholding Gaussian random fields. Journal of Computational Physics, 277(??):16–31, November 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005397.

Hu:2010:ACU

[HWA10] X. Y. Hu, Q. Wang, and N. A. Adams. An adaptive centralupwind weighted essentially non-oscillatory scheme. Journal of Computational Physics, 229(23):8952–8965, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004560.

Hwang:2010:PAS

[HWHW10] Feng-Nan Hwang, Zih-Hao Wei, Tsung-Ming Huang, and Weichung Wang. A parallel additive Schwarz preconditioned Jacobi–Davidson algorithm for polynomial eigenvalue problems in quantum dot simulation. Journal of Computational Physics, 229(8):2932–2947, April 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109007037.

Hellrung:2012:SOV

[HWST12] Jeffrey Lee Hellrung, Jr., Luming Wang, Eftychios Sifakis, and Joseph M. Teran. A second order virtual node method for elliptic problems with interfaces and irregular domains in three dimensions. Journal of Computational Physics, 231(4):2015–2048, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006784.

Hou:2010:NMS

[HWW10] Songming Hou, Wei Wang, and Liqun Wang. Numerical method for solving matrix coefficient elliptic equation with sharp-edged interfaces. Journal of Computational Physics, 229(19):7162–7179, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003207.

Hsieh:2009:BSL

[HY09] Po-Wen Hsieh and Suh-Yuh Yang. A bubble-stabilized leastsquares finite element method for steady MHD duct flow problems at high Hartmann numbers. Journal of Computational Physics, 228(22):8301–8320, December 1, 2009. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004410. See erratum [HY11].

Hsieh:2010:TNU

[HY10] Po-Wen Hsieh and Suh-Yuh Yang. Two new upwind difference schemes for a coupled system of convection-diffusion equations arising from the steady MHD duct flow problems.

Journal of Computational Physics, 229(24):9216–9234, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004894.

Hsieh:2011:EBS

[HY11] Po-Wen Hsieh and Suh-Yuh Yang. Erratum to “A bubblestabilized least-squares finite element method for steady MHD duct flow problems at high Hartmann numbers” [J. Comput. Phys. 228 (2009) 8301–8320]. Journal of Computational Physics, 230(2):474–475, January 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005474. See [HY09].

Hysing:2012:MEF

[Hys12] S. Hysing. Mixed element FEM level set method for numerical simulation of immiscible fluids. Journal of Computational Physics, 231(6):2449–2465, March 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006917.

Han:2011:AAB

[HZ11a] Houde Han and Zhiwen Zhang. Adaptive artificial boundary condition for the two-level Schrödinger equation with conical crossings. Journal of Computational Physics, 230(4): 1319–1334, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110006066.

Hu:2011:SFP

[HZ11b] Guanghui Hu and Paul Andries Zegeling. Simulating finger phenomena in porous media with a moving finite element method. Journal of Computational Physics, 230(8):3249–3263, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000568.

Hicken:2014:DCF

[HZ14] J. E. Hicken and D. W. Zingg. Dual consistency and functional accuracy: a finite-difference perspective. Journal of Computational Physics, 256(??):161–182, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005524.

Hiremath:2012:NSN

[HZS12] Kirankumar R. Hiremath, Lin Zschiedrich, and Frank Schmidt. Numerical solution of nonlocal hydrodynamic Drude model for arbitrary shaped nano-plasmonic structures using Nédélec finite elements. Journal of Computational Physics, 231(17):5890–5896, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002550.

Icardi:2012:QBM

[IAM^{+}12] M. Icardi, P. Asinari, D. L. Marchisio, S. Izquierdo, and R. O. Fox. Quadrature-based moment closures for non-equilibrium flows: Hard-sphere collisions and approach to equilibrium. Journal of Computational Physics, 231(21):7431–7449, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003853.

Iannelli:2011:IGF

[Ian11] Joe Iannelli. An implicit Galerkin finite element Runge– Kutta algorithm for shock-structure investigations. Journal of Computational Physics, 230(1):260–286, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005280.

Iafrati:2014:MOA

[IBO14] A. Iafrati, A. Babanin, and M. Onorato. Modeling of ocean-atmosphere interaction phenomena during the breaking of modulated wave trains. Journal of Computational Physics, 271(??):151–171, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300853X.

Isoardi:2010:PML

[ICC^{+}10] L. Isoardi, G. Chiavassa, G. Ciraolo, P. Haldenwang, E. Serre, Ph. Ghendrih, Y. Sarazin, F. Schwander, and P. Tamain. Penalization modeling of a limiter in the Tokamak edge plasma. Journal of Computational Physics, 229(6):2220–2235, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006603.

Ichinomiya:2013:TCG

[Ich13] Takashi Ichinomiya. Temporal coarse-graining method to simulate the movement of atoms. Journal of Computational Physics, 251(??):319–326, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004270.

Ianniello:2010:SAO

[ID10] S. Ianniello and A. Di Mascio. A self-adaptive oriented particles Level-Set method for tracking interfaces. Journal of Computational Physics, 229(4):1353–1380, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005865.

Idesman:2012:UPP

[Ide12] A. Idesman. Use of post-processing to increase the order of accuracy of the trapezoidal rule at time integration of linear elastodynamics problems. Journal of Computational Physics, 231(8):3143–3165, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000034.

Ivan:2013:MDF

[IDNG13] Lucian Ivan, Hans De Sterck, Scott A. Northrup, and Clinton P. T. Groth. Multi-dimensional finite-volume scheme for hyperbolic conservation laws on three-dimensional solutionadaptive cubed-sphere grids. Journal of Computational Physics, 255(??):205–227, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911300538X.

Izquierdo:2010:MTC

[IF10] Salvador Izquierdo and Norberto Fueyo. Momentum transfer correction for macroscopic-gradient boundary conditions in lattice Boltzmann methods. Journal of Computational Physics, 229(7):2497–2506, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006652.

Iadarola:2012:GAM

[IFD^{+}12] Giovanni Iadarola, Carlo Forestiere, Luca Dal Negro, Fabio Villone, and Giovanni Miano. GPU-accelerated T -matrix algorithm for light-scattering simulations. Journal of Computational Physics, 231(17):5640–5652, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001490.

Ivan:2014:HOS

[IG14] Lucian Ivan and Clinton P. T. Groth. High-order solutionadaptive central essentially non-oscillatory (CENO) method for viscous flows. Journal of Computational Physics, 257 (??):830–862, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006591.

Ingraham:2013:EVA

[IH13] Daniel Ingraham and Ray Hixon. External verification analysis: a code-independent verification technique for unsteady PDE codes. Journal of Computational Physics, 243(??):46–57, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001514.

Iga:2014:ISH

[iIT14] Shin ichi Iga and Hirofumi Tomita. Improved smoothness and homogeneity of icosahedral grids using the spring dynamics method. Journal of Computational Physics, 258(??):208–226, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006815.

Ito:2013:TSM

[IJZ13] Kazufumi Ito, Bangti Jin, and Jun Zou. A two-stage method for inverse medium scattering. Journal of Computational Physics, 237(??):211–223, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007358.

Ivanyshyn:2011:ISS

[IK11] Olha Ivanyshyn and Rainer Kress. Inverse scattering for surface impedance from phase-less far field data. Journal of Computational Physics, 230(9):3443–3452, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000635.

Ito:2014:HPC

[IKO^{+}14] Kei Ito, Tomoaki Kunugi, Shuji Ohno, Hideki Kamide, and Hiroyuki Ohshima. A high-precision calculation method for interface normal and curvature on an unstructured grid. Journal of Computational Physics, 273(??):38–53, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003386.

Iollo:2011:LSS

[IL11] Angelo Iollo and Damiano Lombardi. A Lagrangian scheme for the solution of the optimal mass transfer problem. Journal of Computational Physics, 230(9):3430–3442, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000623.

Innocenti:2013:MLM

[ILM^{+}13] M. E. Innocenti, G. Lapenta, S. Markidis, A. Beck, and A. Vapirev. A Multi Level Multi Domain Method for Particle In Cell plasma simulations. Journal of Computational Physics, 238(??):115–140, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007590.

Ionescu:2013:ALS

[Ion13] Ioan R. Ionescu. Augmented Lagrangian for shallow viscoplastic flow with topography. Journal of Computational Physics, 242(??):544–560, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001484.

Ii:2012:ICM

[IST^{+}12] Satoshi Ii, Kazuyasu Sugiyama, Shintaro Takeuchi, Shu Takagi, Yoichiro Matsumoto, and Feng Xiao. An interface capturing method with a continuous function: The THINC method with multi-dimensional reconstruction. Journal of Computational Physics, 231(5):2328–2358, March 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006942.

Ihme:2012:RRP

[ISZ12] Matthias Ihme, Lee Shunn, and Jian Zhang. Regularization of reaction progress variable for application to flamelet-based combustion models. Journal of Computational Physics, 231(23):7715–7721, October 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003440.

Tsubota:2014:SNB

[iT14] Ken ichi Tsubota. Short note on the bending models for a membrane in capsule mechanics: Comparison between continuum and discrete models. Journal of Computational Physics, 277(??):320–328, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005531.

Iwata:2010:MPE

[ITO^{+}10] Jun-Ichi Iwata, Daisuke Takahashi, Atsushi Oshiyama, Taisuke Boku, Kenji Shiraishi, Susumu Okada, and Kazuhiro Yabana. A massively-parallel electronic-structure calculations based on real-space density functional theory. Journal of Computational Physics, 229(6):2339–2363, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006676.

Ismail:2013:ABI

[IWG13] Mahmoud Ismail, Wolfgang A. Wall, and Michael W. Gee. Adjoint-based inverse analysis of windkessel parameters for patient-specific vascular models. Journal of Computational Physics, 244(??):113–130, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006298.

Ii:2010:GSW

[IX10] Satoshi Ii and Feng Xiao. A global shallow water model using high order multi-moment constrained finite volume method and icosahedral grid. Journal of Computational Physics, 229(5):1774–1796, March 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006275.

Ii:2014:ICM

[IXX14] Satoshi Ii, Bin Xie, and Feng Xiao. An interface capturing method with a continuous function: the THINC method on unstructured triangular and tetrahedral meshes. Journal of Computational Physics, 259(??):260–269, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007961.

Jardin:2012:RIM

[Jar12] S. C. Jardin. Review of implicit methods for the magnetohydrodynamic description of magnetically confined plasmas. Journal of Computational Physics, 231(3):822–838, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006996.

Jakeman:2011:CDH

[JAX11] John D. Jakeman, Richard Archibald, and Dongbin Xiu. Characterization of discontinuities in high-dimensional stochastic problems on adaptive sparse grids. Journal of Computational Physics, 230(10):3977–3997, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001112.

Jiang:2010:EBT

[JC10] Haiyan Jiang and Wei Cai. Effect of boundary treatments on quantum transport current in the Green’s function and Wigner distribution methods for a nano-scale DG–MOSFET. Journal of Computational Physics, 229(12):4461–4475, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000811.

Jamelot:2013:FNO

[JC13] Erell Jamelot and Patrick Ciarlet, Jr. Fast non-overlapping Schwarz domain decomposition methods for solving the neutron diffusion equation. Journal of Computational Physics, 241(??):445–463, May 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000703.

Jeng:2014:SCT

[JCC14] B.-W. Jeng, C.-S. Chien, and I.-L. Chern. Spectral collocation and a two-level continuation scheme for dipolar Bose–Einstein condensates. Journal of Computational Physics, 256(??):713–727, January 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006232.

Jiang:2011:AFI

[JCT11] Haiyan Jiang, Wei Cai, and Raphael Tsu. Accuracy of the Frensley inflow boundary condition for Wigner equations in simulating resonant tunneling diodes. Journal of Computational Physics, 230(5):2031–2044, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006662.

Jeschke:2011:EPS

[JEU11] Matthias Jeschke, Roland Ewald, and Adelinde M. Uhrmacher. Exploring the performance of spatial stochastic simulation algorithms. Journal of Computational Physics, 230 (7):2562–2574, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110007047.

Jakeman:2010:NAQ

[JEX10] John Jakeman, Michael Eldred, and Dongbin Xiu. Numerical approach for quantification of epistemic uncertainty. Journal of Computational Physics, 229(12):4648–4663, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001130.

Jia:2013:FPW

[JFC^{+}13] Weile Jia, Jiyun Fu, Zongyan Cao, Long Wang, Xuebin Chi, Weiguo Gao, and Lin-Wang Wang. Fast plane wave density functional theory molecular dynamics calculations on multi-GPU machines. Journal of Computational Physics, 251(??):102–115, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300329X.

Jouvet:2013:ANM

[JG13] Guillaume Jouvet and Carsten Gräser. An adaptive Newton multigrid method for a model of marine ice sheets. Journal of Computational Physics, 252(??):419–437, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004622.

Jia:2014:TNT

[JGZL14] Zupeng Jia, Xiangfei Gong, Shudao Zhang, and Jun Liu. Two new three-dimensional contact algorithms for staggered Lagrangian hydrodynamics. Journal of Computational Physics, 267(??):247–285, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001685.

Johnsen:2012:PNE

[JH12]  Eric Johnsen and Frank Ham.  Preventing numerical

errors generated by interface-capturing schemes in compressible multi-material flows. Journal of Computational Physics, 231(17):5705–5717, July 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002525.

Jing:2010:CSI

[JHDC10] Yan-Fei Jing, Ting-Zhu Huang, Yong Duan, and Bruno Carpentieri. A comparative study of iterative solutions to linear systems arising in quantum mechanics. Journal of Computational Physics, 229(22):8511–8520, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004274.

Jia:2010:FTA

[JHF10] Jun Jia, Robert Harrison, and George Fann. Fast transform from an adaptive multi-wavelet representation to a partial Fourier representation. Journal of Computational Physics, 229(17):5870–5878, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001828.

Jiang:2010:SME

[JHZ10] Kai Jiang, Yunqing Huang, and Pingwen Zhang. Spectral method for exploring patterns of diblock copolymers. Journal of Computational Physics, 229(20):7796–7805, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000361X.

Jin:2012:VBM

[Jin12] Bangti Jin. A variational Bayesian method to inverse problems with impulsive noise. Journal of Computational Physics, 231(2):423–435, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005298.

Jamali:2013:ABM

[JJM13] Yousef Jamali, Tahereh Jamali, and Mohammad R. K. Mofrad. An agent based model of integrin clustering: Exploring the role of ligand clustering, integrin homo-oligomerization, integrinligand affinity, membrane crowdedness and ligand mobility. Journal of Computational Physics, 244(??):264–278, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005414.

John:2010:VMM

[JK10] Volker John and Adela Kindl. A variational multiscale method for turbulent flow simulation with adaptive large scale space. Journal of Computational Physics, 229(2):301–312, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005178.

Jun:2011:AGF

[JK11] Sukky Jun and Do Wan Kim. Axial Green’s function method for steady Stokes flow in geometrically complex domains. Journal of Computational Physics, 230(5):2095–2124, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006716.

Jiang:2013:SKI

[JKQ13] Shidong Jiang, Mary Catherine A. Kropinski, and Bryan D. Quaife. Second kind integral equation formulation for the modified biharmonic equation and its applications. Journal of Computational Physics, 249(??):113–126, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003185.

Jennings:2014:WWPb

[JKR14] G. I. Jennings, S. Karni, and J. B. Rauch. Water wave propagation in unbounded domains. Part I: Nonreflecting boundaries. Journal of Computational Physics, 276(??):729–739, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001521.

Jebens:2011:PIP

[JKW11] Stefan Jebens, Oswald Knoth, and Rüdiger Weiner. Partially implicit peer methods for the compressible Euler equations. Journal of Computational Physics, 230(12):4955–4974, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001586.

Johnsen:2010:AHR

[JLB^{+}10] Eric Johnsen, Johan Larsson, Ankit V. Bhagatwala, William H. Cabot, Parviz Moin, Britton J. Olson, Pradeep S. Rawat, Santhosh K. Shankar, Björn Sjögreen, H. C. Yee, Xiaolin Zhong, and Sanjiva K. Lele. Assessment of high-resolution methods for numerical simulations of compressible turbulence with shock waves. Journal of Computational Physics, 229(4):1213–1237, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005804.

Jiang:2014:DAI

[JLC14] Haiyan Jiang, Tiao Lu, and Wei Cai. A device adaptive inflow boundary condition for Wigner equations of quantum transport. Journal of Computational Physics, 258(??):773–786, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007572.

Jie:2014:ASP

[JLS^{+}14] Liang Jie, KenLi Li, Lin Shi, RangSu Liu, and Jing Mei. Accelerating solidification process simulation for large-sized system of liquid metal atoms using GPU with CUDA. Journal of Computational Physics, 257(??):521–535, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006803.

Ji:2010:NAM

[JLY10] Hua Ji, Fue-Sang Lien, and Eugene Yee. A new adaptive mesh refinement data structure with an application to detonation. Journal of Computational Physics, 229(23):8981–8993, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004705.

Jia:2013:EIM

[JLZ13] Zupeng Jia, Jun Liu, and Shudao Zhang. An effective integration of methods for second-order three-dimensional multimaterial ALE method on unstructured hexahedral meshes using MOF interface reconstruction. Journal of Computational Physics, 236(??):513–562, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006638.

Jomaa:2010:SWE

[JM10] Z. Jomaa and C. Macaskill. The Shortley–Weller embedded finite-difference method for the 3D Poisson equation with mixed boundary conditions. Journal of Computational Physics, 229(10):3675–3690, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000446.

Jee:2012:CIF

[JM12] SolKeun Jee and Robert D. Moser. Conservative integral form of the incompressible Navier–Stokes equations for a rapidly pitching airfoil. Journal of Computational Physics, 231(19):6268–6289, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112001866.

Jagalur-Mohan:2013:GLS

[JMFO13] J. Jagalur-Mohan, G. Feijóo, and A. Oberai. A Galerkin least squares method for time harmonic Maxwell equations using Nédélec elements. Journal of Computational Physics, 235(??):67–81, February 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911200592X.

Juez:2013:SGF

[JMGN13] C. Juez, J. Murillo, and P. García-Navarro. 2D simulation of granular flow over irregular steep slopes using global and local coordinates. Journal of Computational Physics, 255(??):166–204, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005275.

Jordan:2012:FII

[JMKK12] Gerald Jordan, Martijn Marsman, Yoon-Suk Kim, and Georg Kresse. Fast iterative interior eigensolver for millions of atoms. Journal of Computational Physics, 231(14):4836–4847, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001829.

Jolliet:2012:PFG

[JMV^{+}12] S. Jolliet, B. F. McMillan, L. Villard, T. Vernay, P. Angelino, T. M. Tran, S. Brunner, A. Bottino, and Y. Idomura. Parallel filtering in global gyrokinetic simulations. Journal of Computational Physics, 231(3):745–758, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000544.

Ji:2012:NID

[JMW12] C. Ji, A. Munjiza, and J. J. R. Williams. A novel iterative direct-forcing immersed boundary method and its finite volume applications. Journal of Computational Physics, 231(4):1797–1821, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006528.

John:2012:ENO

[JN12] Volker John and Julia Novo. On (essentially) non-oscillatory discretizations of evolutionary convection-diffusion equations. Journal of Computational Physics, 231(4):1570–1586, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006255.

Janakiraman:2012:NVR

[JNM12] S. Janakiraman, Ravi S. Nanjundiah, and A. S. Vasudeva Murthy. A novel variable resolution global spectral method on the sphere. Journal of Computational Physics, 231(7):2794–2810, April 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111007352.

Jahangiri:2012:HOM

[JNSA12] Pouyan Jahangiri, Amir Nejat, Jila Samadi, and Ali Aboutalebi. A high-order Monte Carlo algorithm for the direct simulation of Boltzmann equation. Journal of Computational Physics, 231(14):4578–4596, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200143X.

Joggerst:2014:CCC

[JNW^{+}14] C. C. Joggerst, Anthony Nelson, Paul Woodward, Catherine Lovekin, Thomas Masser, Chris L. Fryer, P. Ramaprabhu, Marianne Francois, and Gabriel Rockefeller. Cross-code comparisons of mixing during the implosion of dense cylindrical and spherical shells. Journal of Computational Physics, 275(??):154–173, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114004446.

Jakeman:2013:MME

[JNX13] John D. Jakeman, Akil Narayan, and Dongbin Xiu. Minimal multi-element stochastic collocation for uncertainty quantification of discontinuous functions. Journal of Computational Physics, 242(??):790–808, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300154X.

Johnsen:2011:TCD

[Joh11] Eric Johnsen. On the treatment of contact discontinuities using WENO schemes. Journal of Computational Physics, 230(24):8665–8668, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005043.

Jennings:2014:WWPa

[JPC^{+}14] G. I. Jennings, D. Prigge, S. Carney, S. Karni, J. B. Rauch, and R. Abgrall. Water wave propagation in unbounded domains. Part II: Numerical methods for fractional PDEs. Journal of Computational Physics, 275(??):443–458, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004872.

Jin:2012:ISL

[JR12] Bangti Jin and William Rundell. An inverse Sturm–Liouville problem with a fractional derivative. Journal of Computational Physics, 231(14):4954–4966, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001763.

Jamshidian:2014:PFM

[JR14] M. Jamshidian and T. Rabczuk. Phase field modelling of stressed grain growth: Analytical study and the effect of microstructural length scale. Journal of Computational Physics, 261(??):23–35, March 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008218.

Johnen:2013:GVC

[JRG13] A. Johnen, J.-F. Remacle, and C. Geuzaine. Geometrical validity of curvilinear finite elements. Journal of Computational Physics, 233(??):359–372, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005219.

Jiang:2011:SKI

[JRTY11] Shidong Jiang, Bo Ren, Paul Tsuji, and Lexing Ying. Second kind integral equations for the first kind Dirichlet problem of the biharmonic equation in three dimensions. Journal of Computational Physics, 230(19):7488–7501, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003792.

Jaisankar:2013:DDR

[JS13a] S. Jaisankar and T. S. Sheshadri. Directional Diffusion Regulator (DDR) for some numerical solvers of hyperbolic conservation laws. Journal of Computational Physics, 233(??):83–99, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004159.

Ji:2013:MLM

[JS13b] Xia Ji and Jiguang Sun. A multi-level method for transmission eigenvalues of anisotropic media. Journal of Computational Physics, 255(??):422–435, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005688.

Jemison:2014:CMS

[JSA14] Matthew Jemison, Mark Sussman, and Marco Arienti. Compressible, multiphase semi-implicit method with moment of fluid interface representation. Journal of Computational Physics, 279(??):182–217, December 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006317.

Ji:2010:WFB

[JSX10] Hui Ji, Zuowei Shen, and Yuhong Xu. Wavelet frame based scene reconstruction from range data. Journal of Computational Physics, 229(6):2093–2108, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006536.

Jenny:2010:SAF

[JTH10] Patrick Jenny, Manuel Torrilhon, and Stefan Heinz. A solution algorithm for the fluid dynamic equations based on a stochastic model for molecular motion. Journal of Computational Physics, 229(4):1077–1098, February 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005531. See comment [Dad12].

Jauberteau:2014:MFS

[JTT14] F. Jauberteau, R. M. Temam, and J. Tribbia. Multiscale/ fractional step schemes for the numerical simulation of the rotating shallow water flows with complex periodic topography. Journal of Computational Physics, 270(??):506–531, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400223X.

Ju:2012:CUF

[JTXZ12] Lili Ju, Li Tian, Xiao Xiao, and Weidong Zhao. Covolumeupwind finite volume approximations for linear elliptic partial differential equations. Journal of Computational Physics, 231(18):6097–6120, July 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002288.

Jahnke:2010:SCM

[JU10] Tobias Jahnke and Tudor Udrescu. Solving chemical master equations by adaptive wavelet compression. Journal of Computational Physics, 229(16):5724–5741, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001919.

Jiang:2011:TRT

[JW11] Jianguo Jiang and Jichun Wu. A transition rate transformation method for solving advection-dispersion equation. Journal of Computational Physics, 230(14):5556–5563, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001859.

Jones:2012:DSV

[JW12] D. P. Jones and A. P. Watkins. Droplet size and velocity distributions for spray modelling. Journal of Computational Physics, 231(2):676–692, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005894.

Jacobson:2011:EAC

[JWNL11] M. Z. Jacobson, J. T. Wilkerson, A. D. Naiman, and S. K. Lele. The effects of aircraft on climate and pollution. Part I: Numerical methods for treating the subgrid evolution of discrete sizeand composition-resolved contrails from all commercial flights worldwide. Journal of Computational Physics, 230(12):5115– 5132, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001823.

Jackson:2012:NSC

[JWV12] R. H. Jackson, A. C. F. Wu, and J. P. Verboncoeur. Numerical solution of the cylindrical Poisson equation using the Local Taylor Polynomial technique. Journal of Computational Physics, 231(16):5421–5442, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002161.

Jin:2010:BDB

[JWYH10] Shi Jin, Hao Wu, Xu Yang, and Zhongyi Huang. Bloch decomposition-based Gaussian beam method for the Schrödinger equation with periodic potentials. Journal of Computational Physics, 229(13):4869–4883, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000483.

Jiang:2014:DRR

[JXD14] Jinfeng Jiang, Yan Xu, and Hui-Hui Dai. A dissipationrate reserving DG method for wave catching-up phenomena in a nonlinearly elastic composite bar. Journal of Computational Physics, 258(??):405–430, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007274.

Jin:2011:CAP

[JY11] Shi Jin and Bokai Yan. A class of asymptotic-preserving schemes for the Fokker–Planck–Landau equation. Journal of Computational Physics, 230(17):6420–6437, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002269.

Jeong:2010:IBT

[JYHT10] H. K. Jeong, H. S. Yoon, M. Y. Ha, and M. Tsutahara. An immersed boundary-thermal lattice Boltzmann method using an equilibrium internal energy density approach for the simulation of flows with heat transfer. Journal of Computational Physics, 229(7):2526–2543, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900672X.

Jin:2010:HBI

[JZ10] Bangti Jin and Jun Zou. Hierarchical Bayesian inference for ill-posed problems via variational method. Journal of Computational Physics, 229(19):7317–7343, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003311.

Jia:2011:NHO

[JZ11] Zupeng Jia and Shudao Zhang. A new high-order discontinuous Galerkin spectral finite element method for Lagrangian gas dynamics in two-dimensions. Journal of Computational Physics, 230(7):2496–2522, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000687X.

Jiang:2013:KII

[JZ13] Tian Jiang and Yong-Tao Zhang. Krylov implicit integration factor WENO methods for semilinear and fully nonlinear advection-diffusion-reaction equations. Journal of Computational Physics, 253(??):368–388, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004889.

Jiang:2014:NMQ

[JZ14] Kai Jiang and Pingwen Zhang. Numerical methods for quasicrystals. Journal of Computational Physics, 256(??):428–440, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300572X.

Kolobov:2012:TAK

[KA12] V. I. Kolobov and R. R. Arslanbekov. Towards adaptive kinetic-fluid simulations of weakly ionized plasmas. Journal of Computational Physics, 231(3):839–869, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003482.

Korzec:2013:TSM

[KA13] M. D. Korzec and T. Ahnert. Time-stepping methods for the simulation of the self-assembly of nano-crystals in Matlab on a GPU. Journal of Computational Physics, 251(??):396–413, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004063.

Kraus:2012:PML

[KAA12] C. Kraus, A. Adelmann, and P. Arbenz. Perfectly matched layers in a divergence preserving ADI scheme for electromagnetics. Journal of Computational Physics, 231(1):39–44, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005031.

Koren:2014:PCN

[KAB^{+}14] Barry Koren, Rémi Abgrall, Pavel Bochev, Jason Frank, and Blair Perot. Physics-compatible numerical methods. Journal of Computational Physics, 257(??):1039, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006918.

Kees:2011:CLS

[KAFB11] C. E. Kees, I. Akkerman, M. W. Farthing, and Y. Bazilevs. A conservative level set method suitable for variable-order approximations and unstructured meshes. Journal of Computational Physics, 230(12):4536–4558, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001197.

Kylasa:2014:PGR

[KAG14] S. B. Kylasa, H. M. Aktulga, and A. Y. Grama. PuReMD– GPU: a reactive molecular dynamics simulation package for GPUs. Journal of Computational Physics, 272(??):343–359, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002988.

Katsikadelis:2014:NSD

[Kat14] John T. Katsikadelis. Numerical solution of distributed order fractional differential equations. Journal of Computational Physics, 259(??):11–22, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007754.

Kaul:2010:TDE

[Kau10] Upender K. Kaul. Three-dimensional elliptic grid generation with fully automatic boundary constraints. Journal of Computational Physics, 229(17):5966–5979, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002044.

Kawai:2013:DFP

[Kaw13] Soshi Kawai. Divergence-free-preserving high-order schemes for magnetohydrodynamics: an artificial magnetic resistivity method. Journal of Computational Physics, 251(??):292–318, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003999.

Kazantsev:2010:IOD

[Kaz10] Eugene Kazantsev. Identification of an optimal derivatives approximation by variational data assimilation. Journal of Computational Physics, 229(2):256–275, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005105.

Kim:2014:SHR

[KB14] Jinho Kim and Dinshaw S. Balsara. A stable HLLC Riemann solver for relativistic magnetohydrodynamics. Journal of Computational Physics, 270(??):634–639, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002861.

Kwakkel:2013:ECM

[KBB13] Marcel Kwakkel, Wim-Paul Breugem, and Bendiks Jan Boersma. Extension of a CLSVOF method for droplet-laden flows with a coalescence/breakup model. Journal of Computational Physics, 253(??):166–188, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004762.

Klockner:2013:QEN

[KBGO13] Andreas Klöckner, Alexander Barnett, Leslie Greengard, and Michael O’Neil. Quadrature by expansion: a new method for the evaluation of layer potentials. Journal of Computational Physics, 252(??):332–349, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004579.

Kindelan:2010:ARM

[KBGRM10] Manuel Kindelan, Francisco Bernal, Pedro González-Rodríguez, and Miguel Moscoso. Application of the RBF meshless method to the solution of the radiative transport equation. Journal of Computational Physics, 229(5):1897–1908, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006330.

Kassebaum:2012:AGR

[KBRM12] P. G. Kassebaum, C. R. Boucher, and L. R. Ram-Mohan. Application of group representation theory to derive Hermite interpolation polynomials on a triangle. Journal of Computational Physics, 231(17):5747–5760, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002495.

Kenamond:2014:CTE

[KBS14] Mack Kenamond, Matthew Bement, and Mikhail Shashkov. Compatible, total energy conserving and symmetry preserving arbitrary Lagrangian–Eulerian hydrodynamics in 2D rz-cylindrical coordinates. Journal of Computational Physics, 268(??):154–185, July 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400165X.

Kramer:2014:ACE

[KBSV14] Richard M. J. Kramer, Pavel B. Bochev, Christopher M. Siefert, and Thomas E. Voth. Algebraically constrained extended edge element method (e XFEM-AC) for resolution of multi-material cells. Journal of Computational Physics, 276(??):596–612, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005142.

Karagadde:2012:CVI

[KBTD12] Shyamprasad Karagadde, Anirban Bhattacharya, Gaurav Tomar, and Pradip Dutta. A coupled VOF-IBM-enthalpy approach for modeling motion and growth of equiaxed dendrites in a solidifying melt. Journal of Computational Physics, 231(10):3987–4000, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000654.

K:2014:FVM

[KCG14] Sudarshan Kumar K., Praveen C., and G. D. Veerappa Gowda. A finite volume method for a two-phase multicomponent polymer flooding. Journal of Computational Physics, 275(??):667–695, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114004951.

Kluth:2010:DHU

[KD10] G. Kluth and B. Després. Discretization of hyperelasticity on unstructured mesh with a cell-centered Lagrangian scheme.

Journal of Computational Physics, 229(24):9092–9118, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004717.

Kazolea:2014:NTW

[KDS14] M. Kazolea, A. I. Delis, and C. E. Synolakis. Numerical treatment of wave breaking on unstructured finite volume approximations for extended Boussinesq-type equations. Journal of Computational Physics, 271(??):281–305, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000576.

Keaveny:2014:FFC

[Kea14] Eric E. Keaveny. Fluctuating force-coupling method for simulations of colloidal suspensions. Journal of Computational Physics, 269(??):61–79, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001867.

Kacem:2012:FMG

[KED^{+}12] S. Kacem, O. Eichwald, O. Ducasse, N. Renon, M. Yousfi, and K. Charrada. Full multi grid method for electric field computation in point-to-plane streamer discharge in air at atmospheric pressure. Journal of Computational Physics, 231(2):251–261, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100461X.

Koren:2012:CPP

[KEG^{+}12] Barry Koren, Ute Ebert, Tamas Gombosi, Hervé Guillard, Rony Keppens, and Dana Knoll. Computational plasma physics. Journal of Computational Physics, 231(3):717, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006589.

Komatitsch:2010:HOF

[KEGM10] Dimitri Komatitsch, Gordon Erlebacher, Dominik Göddeke, and David Michéa. High-order finite-element seismic wave propagation modeling with MPI on a large GPU cluster.

Journal of Computational Physics, 229(20):7692–7714, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003396.

Ketcheson:2010:RKM

[Ket10] David I. Ketcheson. Runge–Kutta methods with minimum storage implementations. Journal of Computational Physics, 229(5):1763–1773, March 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006251.

Kempe:2012:IIB

[KF12] Tobias Kempe and Jochen Fröhlich. An improved immersed boundary method with direct forcing for the simulation of particle laden flows. Journal of Computational Physics, 231(9):3663–3684, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000423.

Kast:2013:OBM

[KF13a] Steven M. Kast and Krzysztof J. Fidkowski. Outputbased mesh adaptation for high order Navier–Stokes simulations on deformable domains. Journal of Computational Physics, 252(??):468–494, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004300.

Koumoutsakos:2013:MSS

[KF13b] Petros Koumoutsakos and Justin Feigelman. Multiscale stochastic simulations of chemical reactions with regulated scale separation. Journal of Computational Physics, 244(??): 290–297, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112007103.

Katiyar:2014:PFP

[KFOS14] Amit Katiyar, John T. Foster, Hisanao Ouchi, and Mukul M. Sharma. A peridynamic formulation of pressure driven convective fluid transport in porous media. Journal of Computational Physics, 261(??):209–229, March 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008474.

Khayyer:2011:ESA

[KG11] Abbas Khayyer and Hitoshi Gotoh. Enhancement of stability and accuracy of the moving particle semi-implicit method. Journal of Computational Physics, 230(8):3093–3118, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000271.

Kelly:2012:CDG

[KG12a] James F. Kelly and Francis X. Giraldo. Continuous and discontinuous Galerkin methods for a scalable three-dimensional nonhydrostatic atmospheric model: Limited-area mode. Journal of Computational Physics, 231(24):7988–8008, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002252.

Kumar:2012:ABI

[KG12b] Amit Kumar and Michael D. Graham. Accelerated boundary integral method for multiphase flow in non-periodic geometries. Journal of Computational Physics, 231(20):6682–6713, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003051.

Khayyer:2013:EPS

[KG13a] Abbas Khayyer and Hitoshi Gotoh. Enhancement of performance and stability of MPS mesh-free particle method for multiphase flows characterized by high density ratios. Journal of Computational Physics, 242(??):211–233, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001010.

Ko:2013:MES

[KG13b] Jordan Ko and Josselin Garnier. Multi-element stochastic spectral projection for high quantile estimation. Journal of Computational Physics, 243(??):87–108, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000429.

Kreeft:2013:MMS

[KG13c] Jasper Kreeft and Marc Gerritsma. Mixed mimetic spectral element method for Stokes flow: a pointwise divergence-free solution. Journal of Computational Physics, 240(??):284–309, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200647X.

Kopera:2014:AAM

[KG14] Michal A. Kopera and Francis X. Giraldo. Analysis of adaptive mesh refinement for IMEX discontinuous Galerkin solutions of the compressible Euler equations with application to atmospheric simulations. Journal of Computational Physics, 275(??):92–117, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004331.

Kishor:2010:LSE

[KGG10] D. Krishna Kishor, S. Gopalakrishnan, and Ranjan Ganguli. A Legendre spectral element model for sloshing and acoustic analysis in nearly incompressible fluids. Journal of Computational Physics, 229(7):2605–2624, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006792.

Kumar:2013:WEG

[KGK13] G. Kumar, S. S. Girimaji, and J. Kerimo. WENOenhanced gas-kinetic scheme for direct simulations of compressible transition and turbulence. Journal of Computational Physics, 234(??):499–523, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005943.

Kempf:2013:NUY

[KGKS13] Andreas Kempf, Urs Ganse, Patrick Kilian, and Felix Spanier. Note on the use of Yee-lattices in (semi-) implicit particle-in-cell codes. Journal of Computational Physics, 237(??):56–60, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007279.

Kucharik:2010:CSI

[KGSS10] Milan Kucharik, Rao V. Garimella, Samuel P. Schofield, and Mikhail J. Shashkov. A comparative study of interface reconstruction methods for multi-material ALE simulations. Journal of Computational Physics, 229(7):2432–2452, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109003891.

Khan:2011:PDM

[KH11] Md. Ashfaquzzaman Khan and Martin C. Herbordt. Parallel discrete molecular dynamics simulation with speculation and in-order commitment. Journal of Computational Physics, 230(17):6563–6582, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002968.

Kang:2013:ELM

[KH13] Shin K. Kang and Yassin A. Hassan. The effect of lattice models within the lattice Boltzmann method in the simulation of wall-bounded turbulent flows. Journal of Computational Physics, 232(1):100–117, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003968.

Khader:2013:NTS

[Kha13] M. M. Khader. Numerical treatment for solving the perturbed fractional PDEs using hybrid techniques. Journal of Computational Physics, 250(??):565–573, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003987.

Kubrak:2013:LDS

[KHGW13] B. Kubrak, H. Herlina, F. Greve, and J. G. Wissink. Lowdiffusivity scalar transport using a WENO scheme and dual meshing. Journal of Computational Physics, 240(??):158–173, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000351.

Kupka:2012:TVD

[KHHK12] Friedrich Kupka, Natalie Happenhofer, Inmaculada Higueras, and Othmar Koch. Total-variation-diminishing implicitexplicit Runge–Kutta methods for the simulation of doublediffusive convection in astrophysics. Journal of Computational Physics, 231(9):3561–3586, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007522.

Korpilo:2013:NSM

[KHJ^{+}13] T. Korpilo, J. A. Heikkinen, S. J. Janhunen, T. P. Kiviniemi, S. Leerink, and F. Ogando. Numerically stable method for kinetic electrons in gyrokinetic particlein-cell simulation of toroidal plasmas. Journal of Computational Physics, 239(??):22–29, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000065.

Kong:2011:PCS

[KHRD11] Xianglong Kong, Michael C. Huang, Chuang Ren, and Viktor K. Decyk. Particle-in-cell simulations with chargeconserving current deposition on graphic processing units. Journal of Computational Physics, 230(4):1676–1685, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006479.

Kong:2010:SMI

[KHZ10] Linghua Kong, Jialin Hong, and Jingjing Zhang. Splitting multisymplectic integrators for Maxwell’s equations. Journal of Computational Physics, 229(11):4259–4278, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000835.

Kim:2014:DEB

[KIF^{+}14] Junghan Kim, Eldhose Iype, Arjan J. H. Frijns, Silvia V. Nedea, and Anton A. van Steenhoven. Development of EEM based silicon-water and silica-water wall potentials for non-reactive molecular dynamics simulations. Journal of Computational Physics, 268(??):51–62, July 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001752.

Keys:2011:CCP

[KIG11] Aaron S. Keys, Christopher R. Iacovella, and Sharon C. Glotzer. Characterizing complex particle morphologies through shape matching: Descriptors, applications, and algorithms. Journal of Computational Physics, 230(17):6438–6463, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002567.

Kim:2013:QDP

[Kim13] Jae Wook Kim. Quasi-disjoint pentadiagonal matrix systems for the parallelization of compact finite-difference schemes and filters. Journal of Computational Physics, 241(??):168–194, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000995.

Ketefian:2011:MEV

[KJ11] G. S. Ketefian and M. Z. Jacobson. A mass, energy, vorticity, and potential enstrophy conserving lateral boundary scheme for the shallow water equations using piecewise linear boundary approximations. Journal of Computational Physics, 230(8):2751–2793, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006108.

Kim:2013:ADS

[KJC13] Min-Geun Kim, Hong-Lae Jang, and Seonho Cho. Adjoint design sensitivity analysis of reduced atomic systems using generalized Langevin equation for lattice structures. Journal of Computational Physics, 240(??):1–19, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300051X.

Kent:2014:DERb

[KJWR14] James Kent, Christiane Jablonowski, Jared P. Whitehead, and Richard B. Rood. Determining the effective resolution of advection schemes. Part II: Numerical testing. Journal of Computational Physics, 278(??):497–508, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006214.

Kadioglu:2010:FSO

[KK10a] Samet Y. Kadioglu and Dana A. Knoll. A fully second order implicit/explicit time integration technique for hydrodynamics plus nonlinear heat conduction problems. Journal of Computational Physics, 229(9):3237–3249, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000069.

Kavouklis:2010:PAG

[KK10b] Christos Kavouklis and Yannis Kallinderis. Parallel adaptation of general three-dimensional hybrid meshes. Journal of Computational Physics, 229(9):3454–3473, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000264.

Kreeft:2010:NFK

[KK10c] Jasper J. Kreeft and Barry Koren. A new formulation of Kapila’s five-equation model for compressible two-fluid flow, and its numerical treatment. Journal of Computational Physics, 229(18):6220–6242, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002019.

Kaboudian:2014:GSM

[KK14] A. Kaboudian and B. C. Khoo. The ghost solid method for the elastic solid-solid interface. Journal of Computational Physics, 257(??):102–125, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006566.

Kadoch:2012:VPM

[KKAS12] Benjamin Kadoch, Dmitry Kolomenskiy, Philippe Angot, and Kai Schneider. A volume penalization method for incompressible flows and scalar advection-diffusion with moving obstacles. Journal of Computational Physics, 231(12):4365–4383, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000678.

Kuprat:2013:BCP

[KKC^{+}13] A. P. Kuprat, S. Kabilan, J. P. Carson, R. A. Corley, and D. R. Einstein. A bidirectional coupling procedure applied to multiscale respiratory modeling. Journal of Computational Physics, 244(??):148–167, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006225.

Kang:2010:NAL

[KKL10] Hyung-Min Kang, Kyu Hong Kim, and Dong-Ho Lee. A new approach of a limiting process for multi-dimensional flows. Journal of Computational Physics, 229(19):7102– 7128, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003153.

Kadioglu:2010:SOS

[KKLR10] Samet Y. Kadioglu, Dana A. Knoll, Robert B. Lowrie, and Rick M. Rauenzahn. A second order self-consistent IMEX method for radiation hydrodynamics. Journal of Computational Physics, 229(22):8313–8332, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004122.

Klein:2013:SBD

[KKO13] Benedikt Klein, Florian Kummer, and Martin Oberlack. A SIMPLE based discontinuous Galerkin solver for steady incompressible flows. Journal of Computational Physics, 237 (??):235–250, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112007334.

Karagiozis:2010:LND

[KKP10] K. Karagiozis, R. Kamakoti, and C. Pantano. A low numerical dissipation immersed interface method for the compressible Navier–Stokes equations. Journal of Computational Physics, 229(3):701–727, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005506.

Kalligiannaki:2012:MCG

[KKPV12] Evangelia Kalligiannaki, Markos A. Katsoulakis, Petr Plechác, and Dionisios G. Vlachos. Multilevel coarse graining and nano-pattern discovery in many particle stochastic systems. Journal of Computational Physics, 231(6):2599–2620, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007212.

Kaltenbacher:2013:MSV

[KKS13] Barbara Kaltenbacher, Manfred Kaltenbacher, and Imbo Sim. A modified and stable version of a perfectly matched layer technique for the 3-d second order wave equation in time domain with an application to aeroacoustics. Journal of Computational Physics, 235(??):407–422, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006055.

Kazemi-Kamyab:2014:AAH

[KKvZB14] V. Kazemi-Kamyab, A. H. van Zuijlen, and H. Bijl. Analysis and application of high order implicit Runge–Kutta schemes for unsteady conjugate heat transfer: a strongly-coupled approach. Journal of Computational Physics, 272(??):471–486, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002794.

Kim:2010:SDI

[KL10a] Yongsam Kim and Ming-Chih Lai. Simulating the dynamics of inextensible vesicles by the penalty immersed boundary method. Journal of Computational Physics, 229(12): 4840–4853, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001300.

Kokh:2010:ADN

[KL10b] S. Kokh and F. Lagoutière. An anti-diffusive numerical scheme for the simulation of interfaces between compressible fluids by means of a five-equation model. Journal of Computational Physics, 229(8):2773–2809, April 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006731.

Kress:2011:SDN

[KL11a] Rainer Kress and Kuo-Ming Lee. A second degree Newton method for an inverse obstacle scattering problem. Journal of Computational Physics, 230(20):7661–7669, August 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003846.

Kropinski:2011:ENM

[KL11b] Mary Catherine A. Kropinski and Enkeleida Lushi. Efficient numerical methods for multiple surfactant-coated bubbles in a two-dimensional Stokes flow. Journal of Computational Physics, 230(12):4466–4487, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001082.

Kalise:2012:MNA

[KL12a] Dante Kalise and Ivar Lie. Modeling and numerical approximation of a 2.5D set of equations for mesoscale atmospheric processes. Journal of Computational Physics, 231(21):7274–7298, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003609.

Kunze:2012:AMM

[KL12b] Rouven Künze and Ivan Lunati. An adaptive multiscale method for density-driven instabilities. Journal of Computational Physics, 231(17):5557–5570, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001349.

Kurganov:2012:NAA

[KL12c] Alexander Kurganov and Yu Liu. New adaptive artificial viscosity method for hyperbolic systems of conservation laws. Journal of Computational Physics, 231(24):8114–8132, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004251.

Karagiannis:2014:SPC

[KL14a] Georgios Karagiannis and Guang Lin. Selection of polynomial chaos bases via Bayesian model uncertainty methods with applications to sparse approximation of PDEs with stochastic inputs. Journal of Computational Physics, 259(??):114–134, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300778X.

Koopman:2014:ADP

[KL14b] E. A. Koopman and C. P. Lowe. An algorithm for detecting percolating structures in periodic systems — application to polymer networks. Journal of Computational Physics, 274(??):758–769, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004720.

Karamitros:2014:DCR

[KLB^{+}14] M. Karamitros, S. Luan, M. A. Bernal, J. Allison, G. Baldacchino, M. Davidkova, Z. Francis, W. Friedland, V. Ivantchenko, A. Ivantchenko, A. Mantero, P. Nieminem, G. Santin, H. N. Tran, V. Stepan, and S. Incerti. Diffusion-controlled reactions modeling in Geant4-DNA. Journal of Computational Physics, 274(??):841–882, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004185.

Kunze:2013:MMF

[KLL13] Rouven Künze, Ivan Lunati, and Seong H. Lee. A multilevel multiscale finite-volume method. Journal of Computational Physics, 255(??):502–520, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300586X.

Kah:2012:HOM

[KLMJ12] D. Kah, F. Laurent, M. Massot, and S. Jay. A high order moment method simulating evaporation and advection of a polydisperse liquid spray. Journal of Computational Physics, 231(2):394–422, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005262.

Kim:2010:NST

[KLP10] Yongsam Kim, Ming-Chih Lai, and Charles S. Peskin. Numerical simulations of two-dimensional foam by the immersed boundary method. Journal of Computational Physics, 229(13):5194–5207, July 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001543.

Kim:2014:NST

[KLPS14] Yongsam Kim, Ming-Chih Lai, Charles S. Peskin, and Yunchang Seol. Numerical simulations of three-dimensional foam by the immersed boundary method. Journal of Computational Physics, 269(??):1–21, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001971.

Katz:2011:NGF

[KM11] Ira Katz and Ioannis G. Mikellides. Neutral gas free molecular flow algorithm including ionization and walls for use in plasma simulations. Journal of Computational Physics, 230(4):1454–1464, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006212.

Keating:2013:FAD

[KM13] Johnwill Keating and Peter D. Minev. A fast algorithm for direct simulation of particulate flows using conforming grids. Journal of Computational Physics, 255(??):486–501, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005779.

Kappeli:2014:WBS

[KM14a] R. Käppeli and S. Mishra. Well-balanced schemes for the Euler equations with gravitation. Journal of Computational Physics, 259(??):199–219, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007900.

Kim:2014:HAC

[KM14b] Sun Ung Kim and Charles W. Monroe. High-accuracy calculations of sixteen collision integrals for Lennard–Jones (12–6) gases and their interpolation to parameterize neon, argon, and krypton. Journal of Computational Physics, 273(??):358–373, September 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003647.

Khazaeli:2013:AGF

[KMA13] Reza Khazaeli, Saeed Mortazavi, and Mahmud Ashrafizaadeh. Application of a ghost fluid approach for a thermal lattice Boltzmann method. Journal of Computational Physics, 250(??):126–140, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003380.

Kraczek:2010:ASM

[KMHJ10] B. Kraczek, S. T. Miller, R. B. Haber, and D. D. Johnson. Adaptive spacetime method using Riemann jump conditions for coupled atomistic–continuum dynamics. Journal of Computational Physics, 229(6):2061–2092, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006524.

Kuzmin:2010:FFL

[KMSS10] Dmitri Kuzmin, Matthias Möller, John N. Shadid, and Mikhail Shashkov. Failsafe flux limiting and constrained data projections for equations of gas dynamics. Journal of Computational Physics, 229(23):8766–8779, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004468.

Keppens:2012:PGA

[KMvM^{+}12] R. Keppens, Z. Meliani, A. J. van Marle, P. Delmont, A. Vlasis, and B. van der Holst. Parallel, grid-adaptive approaches for relativistic hydro and magnetohydrodynamics. Journal of Computational Physics, 231(3):718–744, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000386.

Karimi:2014:MTS

[KN14] S. Karimi and K. B. Nakshatrala. On multi-time-step monolithic coupling algorithms for elastodynamics. Journal of Computational Physics, 273(??):671–705, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003970.

Kurkcu:2011:IRG

[KNR11] Harun Kurkcu, Nilima Nigam, and Fernando Reitich. An integral representation of the Green function for a linear array of acoustic point sources. Journal of Computational Physics, 230(8):2838–2856, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007084.

Komura:2012:GBS

[KO12] Yukihiro Komura and Yutaka Okabe. GPU-based singlecluster algorithm for the simulation of the Ising model. Journal of Computational Physics, 231(4):1209–1215, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005882.

Kuzuoglu:2014:CPT

[KO14] Mustafa Kuzuoglu and Ozlem Ozgun. Combining perturbation theory and transformation electromagnetics for finite element solution of Helmholtz-type scattering problems. Journal of Computational Physics, 274(??):883–897, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004756.

Kolgan:2011:APM

[Kol11] V. P. Kolgan. Application of the principle of minimizing the derivative to the construction of finite-difference schemes for computing discontinuous solutions of gas dynamics. Journal of Computational Physics, 230(7):2384–2390, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110007072.

Kang:2013:RAM

[KPH13] Seongwon Kang, Heinz Pitsch, and Nahmkeon Hur. On a robust ALE method with discrete primary and secondary conservation. Journal of Computational Physics, 254(??):1–7, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004853.

Kropinski:2011:FIE

[KQ11] Mary Catherine A. Kropinski and Bryan D. Quaife. Fast integral equation methods for the modified Helmholtz equation. Journal of Computational Physics, 230(2):425–434, January 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005334.

Kanschat:2010:SCF

[KR10] G. Kanschat and B. Rivière. A strongly conservative finite element method for the coupling of Stokes and Darcy flow. Journal of Computational Physics, 229(17):5933–5943, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000197X.

Kim:2012:DEN

[KRF12] Tae-Yeon Kim, Leo G. Rebholz, and Eliot Fried. A deconvolution enhancement of the navier–stokes–αβ model. Journal of Computational Physics, 231(11):4015–4027, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007017.

Krivodonova:2010:ELT

[Kri10] Lilia Krivodonova. An efficient local time-stepping scheme for solution of nonlinear conservation laws. Journal of Computational Physics, 229(22):8537–8551, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004304.

Kriel:2014:FSM

[Kri14] A. J. Kriel. A flux splitting method for the Euler equations. Journal of Computational Physics, 278(??):326–347, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400607X.

Kroner:2014:LDG

[KRT14] Dietmar Kröner, Michael Ruzicka, and Ioannis Toulopoulos. Local discontinuous Galerkin numerical solutions of non-Newtonian incompressible flows modeled by p-Navier–Stokes equations. Journal of Computational Physics, 270(??):182–202, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002320.

Kanamori:2011:SWD

[KS11a] Masashi Kanamori and Kojiro Suzuki. Shock wave detection in two-dimensional flow based on the theory of characteristics from CFD data. Journal of Computational Physics, 230(8):3085–3092, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000258.

Katz:2011:MQE

[KS11b] Aaron Katz and Venkateswaran Sankaran. Mesh quality effects on the accuracy of CFD solutions on unstructured meshes. Journal of Computational Physics, 230(20):7670–7686, August 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003871.

Keaveny:2011:ASK

[KS11c] Eric E. Keaveny and Michael J. Shelley. Applying a secondkind boundary integral equation for surface tractions in Stokes flow. Journal of Computational Physics, 230(5):2141–2159, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006741. See com-

ment [PSL20].

Kucharik:2012:OSH

[KS12] Milan Kucharik and Mikhail Shashkov. One-step hybrid remapping algorithm for multi-material arbitrary Lagrangian– Eulerian methods. Journal of Computational Physics, 231 (7):2851–2864, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111007546.

Kitamura:2013:TSS

[KS13a] Keiichi Kitamura and Eiji Shima. Towards shock-stable and accurate hypersonic heating computations: a new pressure flux for AUSM-family schemes. Journal of Computational Physics, 245(??):62–83, July 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001769.

Kong:2013:TCE

[KS13b] Rong Kong and Jerome Spanier. Transport-constrained extensions of collision and track length estimators for solutions of radiative transport problems. Journal of Computational Physics, 242(??):682–695, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001423.

Kucharik:2014:CMM

[KS14] Milan Kucharik and Mikhail Shashkov. Conservative multimaterial remap for staggered multi-material arbitrary Lagrangian– Eulerian methods. Journal of Computational Physics, 258 (??):268–304, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007262.

Koal:2012:ASV

[KSB12] K. Koal, J. Stiller, and H. M. Blackburn. Adapting the spectral vanishing viscosity method for large-eddy simulations in cylindrical configurations. Journal of Computational Physics, 231(8):3389–3405, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000356.

Kunz:2014:PNH

[KSB14] Matthew W. Kunz, James M. Stone, and Xue-Ning Bai. Pegasus: a new hybrid-kinetic particle-in-cell code for astrophysical plasma dynamics. Journal of Computational Physics, 259(??):154–174, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113007973.

Kuhnlein:2012:MAF

[KSD12] Christian Kühnlein, Piotr K. Smolarkiewicz, and Andreas Dörnbrack. Modelling atmospheric flows with adaptive moving meshes. Journal of Computational Physics, 231(7):2741–2763, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111007224.

Kurbanmuradov:2013:RSF

[KSK13] Orazgeldi Kurbanmuradov, Karl Sabelfeld, and Peter R. Kramer. Randomized spectral and Fourier-wavelet methods for multidimensional Gaussian random vector fields. Journal of Computational Physics, 245(??):218–234, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300199X.

Kawai:2010:ALA

[KSL10] Soshi Kawai, Santhosh K. Shankar, and Sanjiva K. Lele. Assessment of localized artificial diffusivity scheme for large-eddy simulation of compressible turbulent flows. Journal of Computational Physics, 229(5):1739–1762, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006160.

Khajeh-Saeed:2013:DNS

[KSP13] Ali Khajeh-Saeed and J. Blair Perot. Direct numerical simulation of turbulence using GPU accelerated supercomputers. Journal of Computational Physics, 235(??):241–257, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006547.

Khajeh-Saeed:2010:ASW

[KSPP10] Ali Khajeh-Saeed, Stephen Poole, and J. Blair Perot. Acceleration of the Smith–Waterman algorithm using single and multiple graphics processors. Journal of Computational Physics, 229(11):4247–4258, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000823.

Kedia:2014:SOC

[KSR^{+}14] Kushal S. Kedia, Cosmin Safta, Jaideep Ray, Habib N. Najm, and Ahmed F. Ghoniem. A second-order coupled immersed boundary-SAMR construction for chemically reacting flow over a heat-conducting Cartesian grid-conforming solid. Journal of Computational Physics, 272(??):408–428, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002824.

Kadoura:2014:AMC

[KSS14] Ahmad Kadoura, Shuyu Sun, and Amgad Salama. Accelerating Monte Carlo molecular simulations by reweighting and reconstructing Markov chains: Extrapolation of canonical ensemble averages and second derivatives to different temperature and density conditions. Journal of Computational Physics, 270(??):70–85, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002253.

Kanevsky:2010:MSL

[KST10] Alex Kanevsky, Michael J. Shelley, and Anna-Karin Tornberg. Modeling simple locomotors in Stokes flow. Journal of Computational Physics, 229(4):958–977, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109002861.

Kansa:2013:DCP

[KST13] E. Kansa, U. Shumlak, and S. Tsynkov. Discrete Calderon’s projections on parallelepipeds and their application to computing exterior magnetic fields for FRC plasmas. Journal of Computational Physics, 234(??):172–198, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005785.

Kapahi:2013:TDS

[KSU13] A. Kapahi, S. Sambasivan, and H. S. Udaykumar. A threedimensional sharp interface Cartesian grid method for solving high speed multi-material impact, penetration and fragmentation problems. Journal of Computational Physics, 241(??):308– 332, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000375.

Khatri:2014:EBM

[KT14] Shilpa Khatri and Anna-Karin Tornberg. An embedded boundary method for soluble surfactants with interface tracking for two-phase flows. Journal of Computational Physics, 256(??):768–790, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006244.

Karasozen:2012:NCP

[KTT12] Bülent Karasözen, Anastasia V. Trofimova, and Vyacheslav G. Tsybulin. Natural convection in porous annular domains: Mimetic scheme and family of steady states. Journal of Computational Physics, 231(7):2995–3005, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000149.

Kublik:2013:IIB

[KTT13] Catherine Kublik, Nicolay M. Tanushev, and Richard Tsai. An implicit interface boundary integral method for Poisson’s equation on arbitrary domains. Journal of Computational Physics, 247(??):279–311, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002337.

Kuzmin:2014:HSL

[Kuz14] Dmitri Kuzmin. Hierarchical slope limiting in explicit and implicit discontinuous Galerkin methods. Journal of Computational Physics, 257(??):1140–1162, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003161.

Kritsikis:2014:BFO

[KVBP^{+}14] E. Kritsikis, A. Vaysset, L. D. Buda-Prejbeanu, F. Alouges, and J.-C. Toussaint. Beyond first-order finite element schemes in micromagnetics. Journal of Computational Physics, 256(??):357–366, January 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005731.

Krishnamurthy:2013:PSM

[KVC^{+}13] Adarsh Krishnamurthy, Christopher T. Villongco, Joyce Chuang, Lawrence R. Frank, Vishal Nigam, Ernest Belezzuoli, Paul Stark, David E. Krummen, Sanjiv Narayan, Jeffrey H. Omens, Andrew D. McCulloch, and Roy C. P. Kerckhoffs. Patient-specific models of cardiac biomechanics. Journal of Computational Physics, 244(??):4–21, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005463.

Kallinderis:2012:LOF

[KVM12] Yannis Kallinderis, Panagiotis A. Vitsas, and Penelope Menounou. A low order flow/acoustics interaction method for the prediction of sound propagation using 3D adaptive hybrid grids. Journal of Computational Physics, 231(18):6121–6138, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002884.

Kolesnikov:2010:UPC

[KWH10] R. A. Kolesnikov, W. X. Wang, and F. L. Hinton. Unlikeparticle collision operator for gyrokinetic particle simulations. Journal of Computational Physics, 229(15):5564–5572, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001944.

Krakos:2012:SAL

[KWHD12] Joshua A. Krakos, Qiqi Wang, Steven R. Hall, and David L. Darmofal. Sensitivity analysis of limit cycle oscillations. Journal of Computational Physics, 231(8):3228–3245, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000071.

Kent:2014:DERa

[KWJR14] James Kent, Jared P. Whitehead, Christiane Jablonowski, and Richard B. Rood. Determining the effective resolution of advection schemes. Part I: Dispersion analysis. Journal of Computational Physics, 278(??):485–496, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000825.

Kwon:2014:NAM

[Kwo14] Youngdon Kwon. Numerical aspects in modeling high Deborah number flow and elastic instability. Journal of Computational Physics, 265(??):128–144, May 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000990.

Katz:2011:ASM

[KWS^{+}11] Aaron Katz, Andrew M. Wissink, Venkateswaran Sankaran, Robert L. Meakin, and William M. Chan. Application of strand meshes to complex aerodynamic flow fields. Journal of Computational Physics, 230(17):6512–6530, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002841.

Kamkar:2011:FDC

[KWSJ11] S. J. Kamkar, A. M. Wissink, V. Sankaran, and A. Jameson. Feature-driven Cartesian adaptive mesh refinement for vortex-dominated flows. Journal of Computational Physics, 230(16):6271–6298, July 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002725.

Kotov:2014:CCS

[KYP^{+}14] Dmitry V. Kotov, H. C. Yee, Marco Panesi, Dinesh K. Prabhu, and Alan A. Wray. Computational challenges for simulations related to the NASA Electric Arc Shock Tube (EAST) experiments. Journal of Computational Physics, 269(??):215–233, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002022.

Lange:2010:USD

[LA10] Kyle J. Lange and W. Kyle Anderson. Using sensitivity derivatives for design and parameter estimation in an atmospheric plasma discharge simulation. Journal of Computational Physics, 229(17):6071–6083, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002251.

Langer:2014:AMM

[Lan14] Stefan Langer. Agglomeration multigrid methods with implicit Runge–Kutta smoothers applied to aerodynamic simulations on unstructured grids. Journal of Computational Physics, 277(??):72–100, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005439.

Lapenta:2011:DAP

[Lap11] Giovanni Lapenta. DEMOCRITUS: An adaptive particle in cell (PIC) code for object-plasma interactions. Journal of Computational Physics, 230(12):4679–4695, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001306.

Lapenta:2012:PSS

[Lap12] Giovanni Lapenta. Particle simulations of space weather. Journal of Computational Physics, 231(3):795–821, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001860.

Lyon:2010:HOU

[LB10] Mark Lyon and Oscar P. Bruno. High-order unconditionally stable FC–AD solvers for general smooth domains II. Elliptic, parabolic and hyperbolic PDEs; theoretical considerations. Journal of Computational Physics, 229(9):3358–3381, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000215.

LePenven:2012:SAF

[LB12] Lionel Le Penven and Marc Buffat. On the spectral accuracy of a fictitious domain method for elliptic operators in multi-dimensions. Journal of Computational Physics, 231(23):7893–7906, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004287.

Lee:2013:MMT

[LB13] Yongchang Lee and Cemal Basaran. A multiscale modeling technique for bridging molecular dynamics with finite element method. Journal of Computational Physics, 253 (??):64–85, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113004695.

Li:2014:SGR

[LB14] Shan Li and John P. Boyd. Symmetrizing grids, radial basis functions, and Chebyshev and Zernike polynomials for the D_{4} symmetry group; interpolation within a squircle, Part I. Journal of Computational Physics, 258(??):931–947, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007766.

Lagaert:2014:HSP

[LBC14] J.-B. Lagaert, G. Balarac, and G.-H. Cottet. Hybrid spectralparticle method for the turbulent transport of a passive scalar. Journal of Computational Physics, 260(??):127–142, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008334.

Leiderman:2013:RMN

[LBCL13] Karin Leiderman, Elizabeth L. Bouzarth, Ricardo Cortez, and Anita T. Layton. A regularization method for the numerical solution of periodic Stokes flow. Journal of Computational Physics, 236(??):187–202, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005803.

Luo:2010:MEF

[LBK10] Xian Luo, Ali Beskok, and George Em Karniadakis. Modeling electrokinetic flows by the smoothed profile method. Journal of Computational Physics, 229(10):3828–3847, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000537.

Lee:2010:MFC

[LBM10] Jin-Luen Lee, Rainer Bleck, and Alexander E. MacDonald. A multistep flux-corrected transport scheme. Journal of Computational Physics, 229(24):9284–9298, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000495X.

Liu:2014:TOB

[LBM14] Jinjie Liu, Moysey Brio, and Jerome V. Moloney. Transformation optics based local mesh refinement for solving Maxwell’s equations. Journal of Computational Physics, 258(??):359–370, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007249.

Luchtenburg:2014:LTU

[LBR14] Dirk M. Luchtenburg, Steven L. Brunton, and Clarence W. Rowley. Long-time uncertainty propagation using generalized polynomial chaos and flow map composition. Journal of Computational Physics, 274(??):783–802, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004367.

Lopez-Blanco:2013:ELM

[LBRA^{+}13] José R. López-Blanco, Ruymán Reyes, José I. Aliaga, Rosa M. Badia, Pablo Chacón, and Enrique S. Quintana-Ortí. Exploring large macromolecular functional motions on clusters of multicore processors. Journal of Computational Physics, 246(??):275–288, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002106.

Luu:2010:GRF

[LBS10] Thomas Luu, Eugene D. Brooks III, and Abraham Szőke. Generalized reference fields and source interpolation for the difference formulation of radiation transport. Journal of Computational Physics, 229(5):1626–1642, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900610X.

Lee:2011:FAQ

[LBT11] Wei-Koon Lee, Alistair G. L. Borthwick, and Paul H. Taylor. A fast adaptive quadtree scheme for a two-layer shallow water model. Journal of Computational Physics, 230(12):4848–4870, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001501.

Liu:2010:GFA

[LBZ^{+}10] Jinjie Liu, Moysey Brio, Yong Zeng, Armis R. Zakharian, Walter Hoyer, Stephan W. Koch, and Jerome V. Moloney. Generalization of the FDTD algorithm for simulations of hydrodynamic nonlinear Drude model. Journal of Computational Physics, 229(17):5921–5932, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001920.

Li:2013:PCS

[LC13] Jun Li and Victor M. Calo. Phase-coexistence simulations of fluid mixtures by the Markov Chain Monte Carlo method using single-particle models. Journal of Computational Physics, 249(??):233–248, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002659.

Liska:2014:PFM

[LC14] Sebastian Liska and Tim Colonius. A parallel fast multipole method for elliptic difference equations. Journal of Computational Physics, 278(??):76–91, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005415.

Luke:2012:FMD

[LCB12] Edward Luke, Eric Collins, and Eric Blades. A fast mesh deformation method using explicit interpolation. Journal of Computational Physics, 231(2):586–601, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005535.

Lin:2010:NMS

[LCE^{+}10] Ling Lin, Xiuyuan Cheng, Weinan E., An-Chang Shi, and Pingwen Zhang. A numerical method for the study of nucleation of ordered phases. Journal of Computational Physics, 229(5):1797–1809, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006287.

Li:2011:THC

[LCK11] Qing-Hua Li, Shen-Shen Chen, and Guang-Xiao Kou. Transient heat conduction analysis using the MLPG method and modified precise time step integration method. Journal of Computational Physics, 230(7):2736–2750, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000374.

Lee:2010:UMW

[LCNK10] Lie-Quan Lee, Arno Candel, Cho Ng, and Kwok Ko. On using moving windows in finite element time domain simulation for long accelerator structures. Journal of Computational Physics, 229(24):9235–9245, December 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004924.

Liang:2013:CCE

[LCP13] Chunlei Liang, Christopher Cox, and Michael Plesniak. A comparison of computational efficiencies of spectral difference method and correction procedure via reconstruction. Journal of Computational Physics, 239(??):138–146, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000259.

Lo:2012:REM

[LCWN12] Wing-Cheong Lo, Long Chen, Ming Wang, and Qing Nie. A robust and efficient method for steady state patterns in reaction-diffusion systems. Journal of Computational Physics, 231(15):5062–5077, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001775.

Li:2011:NAF

[LCY11] Changpin Li, An Chen, and Junjie Ye. Numerical approaches to fractional calculus and fractional ordinary differential equation. Journal of Computational Physics, 230(9):3352–3368, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000556.

Li:2012:EOF

[LD12] S. Li and E. Darve. Extension and optimization of the FIND algorithm: Computing Green’s and less-than Green’s functions. Journal of Computational Physics, 231(4):1121–1139, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100338X.

Liang:2014:EDS

[LD14] Dong Liang and Chuanbin Du. The efficient S-DDM scheme and its analysis for solving parabolic equations. Journal of Computational Physics, 272(??):46–69, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002782.

Loseille:2010:FAG

[LDA10] A. Loseille, A. Dervieux, and F. Alauzet. Fully anisotropic goal-oriented mesh adaptation for 3D steady Euler equations. Journal of Computational Physics, 229(8):2866–2897, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109007001.

Lockerby:2013:TSC

[LDDBR13] Duncan A. Lockerby, Carlos A. Duque-Daza, Matthew K. Borg, and Jason M. Reese. Time-step coupling for hybrid simulations of multiscale flows. Journal of Computational Physics, 237(??):344–365, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007127.

Liu:2013:IAA

[LDE^{+}13] Ju Liu, Luca Dedè, John A. Evans, Micheal J. Borden, and Thomas J. R. Hughes. Isogeometric analysis of the advective Cahn–Hilliard equation: Spinodal decomposition under shear flow. Journal of Computational Physics, 242(??):321– 350, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001186.

Lee:2014:NBG

[LDM14]  Curtis Lee, John Dolbow, and Peter J. Mucha.  A

narrow-band gradient-augmented level set method for multiphase incompressible flow. Journal of Computational Physics, 273(??):12–37, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003313.

Loris:2010:NRT

[LDN^{+}10] I. Loris, H. Douma, G. Nolet, I. Daubechies, and C. Regone. Nonlinear regularization techniques for seismic tomography. Journal of Computational Physics, 229(3):890–905, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005725.

Laraufie:2011:DFM

[LDS11] Romain Laraufie, Sébastien Deck, and Pierre Sagaut. A dynamic forcing method for unsteady turbulent inflow conditions. Journal of Computational Physics, 230(23):8647– 8663, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004840.

Lessig:2012:EAR

[LdWF12] C. Lessig, T. de Witt, and E. Fiume. Efficient and accurate rotation of finite spherical harmonics expansions. Journal of Computational Physics, 231(2):243–250, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005468.

LeRoux:2012:SIO

[Le 12]  Daniel Y. Le Roux.  Spurious inertial oscillations in

shallow-water models. Journal of Computational Physics, 231(24):7959–7987, October 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002872.

LeLouer:2014:HOS

[Le 14a] Frédérique Le Louër. A high order spectral algorithm for elastic obstacle scattering in three dimensions. Journal of Computational Physics, 279(??):1–17, December 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006238.

LeLouer:2014:SAN

[Le 14b] Frédérique Le Louër. Spectrally accurate numerical solution of hypersingular boundary integral equations for threedimensional electromagnetic wave scattering problems. Journal of Computational Physics, 275(??):662–666, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005154.

Luque:2012:DMS

[LE12] A. Luque and U. Ebert. Density models for streamer discharges: Beyond cylindrical symmetry and homogeneous media. Journal of Computational Physics, 231(3):904–918, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002580.

Landreman:2013:NVS

[LE13] Matt Landreman and Darin R. Ernst. New velocity-space discretization for continuum kinetic calculations and Fokker– Planck collisions. Journal of Computational Physics, 243 (??):130–150, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113001605.

Lee:2012:ECN

[Lee12a] Sang-Hyeon Lee. Effects of condition number on preconditioning for low Mach number flows. Journal of Computational Physics, 231(10):4001–4014, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000745.

Lee:2012:ASM

[Lee12b] Wei-Ming Lee. Acoustic scattering by multiple elliptical cylinders using collocation multipole method. Journal of Computational Physics, 231(14):4597–4612, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001489.

Lee:2013:SAE

[Lee13] Dongwook Lee. A solution accurate, efficient and stable unsplit staggered mesh scheme for three dimensional magnetohydrodynamics. Journal of Computational Physics, 243(??):269–292, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001836.

Lee:2014:ISP

[Lee14a] Kuo-Ming Lee. Inverse scattering problem from an impedance obstacle via two-steps method. Journal of Computational Physics, 274(??):182–190, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400415X.

Lee:2014:AAH

[Lee14b] Wonjung Lee. Adaptive approximation of higher order posterior statistics. Journal of Computational Physics, 258 (??):833–855, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007778.

Li:2010:SHC

[LEH10] Chao Li, Ute Ebert, and Willem Hundsdorfer. Spatially hybrid computations for streamer discharges with generic features of pulled fronts: I. Planar fronts. Journal of Computational Physics, 229(1):200–220, January 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005191.

Li:2012:SHC

[LEH12] Chao Li, Ute Ebert, and Willem Hundsdorfer. Spatially hybrid computations for streamer discharges: II. Fully 3D simulations. Journal of Computational Physics, 231(3):1020–1050, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004517.

Litvinov:2010:SSH

[LEHA10] S. Litvinov, M. Ellero, X. Y. Hu, and N. A. Adams. A splitting scheme for highly dissipative smoothed particle dynamics. Journal of Computational Physics, 229(15):5457–5464, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001592.

Lauritzen:2011:SIR

[LEM11] Peter H. Lauritzen, Christoph Erath, and Rashmi Mittal. On simplifying ‘incremental remap’-based transport schemes. Journal of Computational Physics, 230(22):7957– 7963, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004037.

Lerat:2013:SDS

[Ler13] Alain Lerat. Steady discrete shocks of high-order RBC schemes. Journal of Computational Physics, 252(??):350–364, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004609.

Lerat:2014:SDS

[Ler14] Alain Lerat. Steady discrete shocks of 5th and 7th-order RBC schemes and shock profiles of their equivalent differential equations. Journal of Computational Physics, 272(??):629–643, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003143.

Leung:2011:EAC

[Leu11] Shingyu Leung. An Eulerian approach for computing the finite time Lyapunov exponent. Journal of Computational Physics, 230(9):3500–3524, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000799.

Lei:2011:TDO

[LFK11] Huan Lei, Dmitry A. Fedosov, and George Em Karniadakis. Time-dependent and outflow boundary conditions for Dissipative Particle Dynamics. Journal of Computational Physics, 230(10):3765–3779, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000854.

Lamballais:2011:SHO

[LFL11] Eric Lamballais, Véronique Fortuné, and Sylvain Laizet. Straightforward high-order numerical dissipation via the viscous term for direct and large eddy simulation. Journal of Computational Physics, 230(9):3270–3275, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000659.

Losch:2014:PJF

[LFLV14] Martin Losch, Annika Fuchs, Jean-François Lemieux, and Anna Vanselow. A parallel jacobian-free Newton–Krylov solver for a coupled sea ice-ocean model. Journal of Computational Physics, 257(??):901–911, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006402.

Liovic:2010:ESS

[LFRM10] Petar Liovic, Marianne François, Murray Rudman, and Richard Manasseh. Efficient simulation of surface tensiondominated flows through enhanced interface geometry interrogation. Journal of Computational Physics, 229(19):7520– 7544, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003499.

Luzzatto-Fegiz:2011:EGN

[LFW11] Paolo Luzzatto-Fegiz and Charles H. K. Williamson. An efficient and general numerical method to compute steady uniform vortices. Journal of Computational Physics, 230 (17):6495–6511, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911100283X.

Li:2014:STA

[LG14] Tong Li and YuanTong Gu. A stochastic thermostat algorithm for coarse-grained thermomechanical modeling of largescale soft matters: Theory and application to microfilaments. Journal of Computational Physics, 263(??):177–184, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000485.

Lerat:2013:DHO

[LGC13] A. Lerat, K. Grimich, and P. Cinnella. On the design of high order residual-based dissipation for unsteady compressible flows. Journal of Computational Physics, 235(??):32–51, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006596.

Liu:2013:FEV

[LGE^{+}13] Ju Liu, Hector Gomez, John A. Evans, Thomas J. R. Hughes, and Chad M. Landis. Functional entropy variables: a new methodology for deriving thermodynamically consistent algorithms for complex fluids, with particular reference to the isothermal Navier–Stokes–Korteweg equations. Journal of Computational Physics, 248(??):47–86, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002544.

Lentine:2011:USF

[LGF11] Michael Lentine, Jón Tómas Grétarsson, and Ronald Fedkiw. An unconditionally stable fully conservative semi-Lagrangian method. Journal of Computational Physics, 230(8):2857–2879, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110007102.

Liang:2013:FMM

[LGG^{+}13] Zhi Liang, Zydrunas Gimbutas, Leslie Greengard, Jingfang Huang, and Shidong Jiang. A fast multipole method for the Rotne–Prager–Yamakawa tensor and its applications. Journal of Computational Physics, 234(??):133–139, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005529.

LeMetayer:2010:NSG

[LGH10a] O. Le Métayer, S. Gavrilyuk, and S. Hank. A numerical scheme for the Green–Naghdi model. Journal of Computational Physics, 229(6):2034–2045, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006408.

Lopez:2010:VFA

[LGH10b] J. López, P. Gómez, and J. Hernández. A volume of fluid approach for crystal growth simulation. Journal of Computational Physics, 229(19):6663–6672, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002871.

Li:2014:HOW

[LGLX14] Maojun Li, Philippe Guyenne, Fengyan Li, and Liwei Xu. High order well-balanced CDG-FE methods for shallow water waves by a Green–Naghdi model. Journal of Computational Physics, 257(??):169–192, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006645.

Lee:2010:IBM

[LGP10] Pilhwa Lee, Boyce E. Griffith, and Charles S. Peskin. The immersed boundary method for advection-electrodiffusion with implicit timestepping and local mesh refinement. Journal of Computational Physics, 229(13):5208–5227, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001555.

Long:2014:ITE

[LGP14] A. R. Long, N. A. Gentile, and T. S. Palmer. The iterative thermal emission method: a more implicit modification of IMC. Journal of Computational Physics, 277(??):228–247, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005713.

Lahivaara:2010:NUB

[LH10a] T. Lähivaara and T. Huttunen. A non-uniform basis order for the discontinuous Galerkin method of the 3D dissipative wave equation with perfectly matched layer. Journal of Computational Physics, 229(13):5144–5160, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000149X.

Leicht:2010:EEA

[LH10b] Tobias Leicht and Ralf Hartmann. Error estimation and anisotropic mesh refinement for 3d laminar aerodynamic flow simulations. Journal of Computational Physics, 229(19): 7344–7360, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003347.

Li:2010:AMA

[LH10c] Xianping Li and Weizhang Huang. An anisotropic mesh adaptation method for the finite element solution of heterogeneous anisotropic diffusion problems. Journal of Computational Physics, 229(21):8072–8094, October 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003785.

Lopez:2010:RIC

[LH10d] J. López and J. Hernández. On reducing interface curvature computation errors in the height function technique.

Journal of Computational Physics, 229(13):4855–4868, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001518.

Larriba:2013:FMC

[LH13] Carlos Larriba and Christopher J. Hogan, Jr. Free molecular collision cross section calculation methods for nanoparticles and complex ions with energy accommodation. Journal of Computational Physics, 251(??):344–363, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911300404X.

Li:2014:AAN

[LH14a] Jichun Li and Jan S. Hesthaven. Analysis and application of the nodal discontinuous Galerkin method for wave propagation in metamaterials. Journal of Computational Physics, 258(??):915–930, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007808.

Liu:2014:EIB

[LH14b] Cheng Liu and Changhong Hu. An efficient immersed boundary treatment for complex moving object. Journal of Computational Physics, 274(??):654–680, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004495.

Ling:2010:NSS

[LHB10] Y. Ling, A. Haselbacher, and S. Balachandar. A numerical source of small-scale number-density fluctuations in Eulerian– Lagrangian simulations of multiphase flows. Journal of Computational Physics, 229(5):1828–1851, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006305.

Leinonen:2014:ASG

[LHH14] Matti Leinonen, Harri Hakula, and Nuutti Hyvönen. Application of stochastic Galerkin FEM to the complete electrode model of electrical impedance tomography. Journal of Computational Physics, 269(??):181–200, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001843.

Leyendecker:2012:VCI

[LHK12] Sigrid Leyendecker, Carsten Hartmann, and Michael Koch. Variational collision integrator for polymer chains. Journal of Computational Physics, 231(10):3896–3911, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000381.

Luostari:2014:UWV

[LHM14] Teemu Luostari, Tomi Huttunen, and Peter Monk. The ultra weak variational formulation of thin clamped plate problems. Journal of Computational Physics, 260(??):85–106, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008358.

Lu:2010:PNP

[LHMZ10] Benzhuo Lu, Michael J. Holst, J. Andrew McCammon, and Y. C. Zhou. Poisson–Nernst–Planck equations for simulating biomolecular diffusion-reaction processes I: Finite element solutions. Journal of Computational Physics, 229(19):6979–6994, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002962.

Lopez-Herrera:2011:CCA

[LHPH11] J. M. López-Herrera, S. Popinet, and M. A. Herrada. A charge-conservative approach for simulating electrohydrodynamic two-phase flows using volume-of-fluid. Journal of Computational Physics, 230(5):1939–1955, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006571.

Lu:2013:CMN

[LHV13] Changna Lu, Weizhang Huang, and Erik S. Van Vleck. The cutoff method for the numerical computation of nonnegative solutions of parabolic PDEs with application to anisotropic diffusion and lubrication-type equations. Journal of Computational Physics, 242(??):24–36, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001307.

Li:2012:DTD

[LHY12] Jichun Li, Yunqing Huang, and Wei Yang. Developing a time-domain finite-element method for modeling of electromagnetic cylindrical cloaks. Journal of Computational Physics, 231(7):2880–2891, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007479.

Li:2013:AEF

[LHY13] Jichun Li, Yunqing Huang, and Wei Yang. An adaptive edge finite element method for electromagnetic cloaking simulation. Journal of Computational Physics, 249(??):216–232, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300291X.

Li:2010:CFE

[Li10a] Peijun Li. Coupling of finite element and boundary integral methods for electromagnetic scattering in a two-layered medium. Journal of Computational Physics, 229(2):481–497, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005324.

Li:2010:FOD

[Li10b] Shengtai Li. A fourth-order divergence-free method for MHD flows. Journal of Computational Physics, 229(20):7893–7910, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003682.

Li:2012:POS

[Li12] Jian-Guo Li. Propagation of ocean surface waves on a spherical multiple-cell grid. Journal of Computational Physics, 231(24):8262–8277, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911200441X.

Li:2013:ALE

[Li13] Jie Li. An arbitrary Lagrangian Eulerian method for threephase flows with triple junction points. Journal of Computational Physics, 251(??):1–16, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003951.

Lv:2014:DGM

[LI14] Yu Lv and Matthias Ihme. Discontinuous Galerkin method for multicomponent chemically reacting flows and combustion. Journal of Computational Physics, 270(??):105–137, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002101.

Lipatov:2012:MPM

[Lip12] Alexander S. Lipatov. Merging for Particle-Mesh Complex Particle Kinetic modeling of the multiple plasma beams. Journal of Computational Physics, 231(8):3101–3118, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007327.

Liu:2011:PMH

[Liu11a] Qianlong Liu. Physalis method for heterogeneous mixtures of dielectrics and conductors: Accurately simulating one million particles using a PC. Journal of Computational Physics, 230(22):8256–8274, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004529.

Liu:2011:EKS

[Liu11b] X. Liu. Extension of Kleiser and Schumann’s influence-matrix method for generalized velocity boundary conditions. Journal of Computational Physics, 230(22):7911–7916, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004359.

Liu:2013:NMP

[Liu13] Jinn-Liang Liu. Numerical methods for the Poisson– Fermi equation in electrolytes. Journal of Computational Physics, 247(??):88–99, August 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002428.

Livermore:2010:GOP

[Liv10] Philip W. Livermore. Galerkin orthogonal polynomials. Journal of Computational Physics, 229(6):2046–2060, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006512.

Li:2013:SCA

[LJ13a] Yanheng Li and Wei Ji. Stability and convergence analysis of a dynamics-based collective method for random sphere packing. Journal of Computational Physics, 250(??):373–387, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300363X.

Lieu:2013:CFR

[LJ13b] Binh K. Lieu and Mihailo R. Jovanović. Computation of frequency responses for linear time-invariant PDEs on a compact interval. Journal of Computational Physics, 250(??):246–269, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003410.

Liu:2014:SSO

[LJG14] Jie Liu, Rajeev K. Jaiman, and Pardha S. Gurugubelli. A stable second-order scheme for fluid-structure interaction with strong added-mass effects. Journal of Computational Physics, 270(??):687–710, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002836.

Leng:2014:FET

[LJX^{+}14] Wei Leng, Lili Ju, Yan Xie, Tao Cui, and Max Gunzburger. Finite element three-dimensional Stokes ice sheet dynamics model with enhanced local mass conservation. Journal of Computational Physics, 274(??):299–311, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004215.

Lee:2014:LAG

[LK14] Wanho Lee and Do Wan Kim. Localized axial Green’s function method for the convection-diffusion equations in arbitrary domains. Journal of Computational Physics, 275(??):390–414, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004641.

Lee:2011:SSF

[LKCY11] Jongho Lee, Jungwoo Kim, Haecheon Choi, and Kyung-Soo Yang. Sources of spurious force oscillations from an immersed boundary method for moving-body problems. Journal of Computational Physics, 230(7):2677–2695, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100012X.

Lemieux:2014:SOA

[LKLG14] Jean-François Lemieux, Dana A. Knoll, Martin Losch, and Claude Girard. A second-order accurate in time IMplicit-EXplicit (IMEX) integration scheme for sea ice dynamics. Journal of Computational Physics, 263(??):375–392, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400031X.

Larsen:2013:PIT

[LKM13] Edward W. Larsen, Akansha Kumar, and Jim E. Morel. Properties of the implicitly time-differenced equations of thermal radiation transport. Journal of Computational Physics, 238(??):82–96, April 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007140.

Latorre:2014:NMC

[LKP14] Juan C. Latorre, Peter R. Kramer, and Grigorios A. Pavliotis. Numerical methods for computing effective transport properties of flashing Brownian motors. Journal of Computational Physics, 257(??):57–82, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006116.

Lemieux:2012:CJF

[LKT^{+}12] Jean-François Lemieux, Dana A. Knoll, Bruno Tremblay, David M. Holland, and Martin Losch. A comparison of the Jacobian-free Newton–Krylov method and the EVP model for solving the sea ice momentum equation with a viscousplastic formulation: a serial algorithm study. Journal of Computational Physics, 231(17):5926–5944, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002665.

Liao:2010:ECE

[LKTL10] W. Liao, E. P. C. Koh, H. M. Tsai, and F. Liu. Euler calculations with embedded Cartesian grids and smallperturbation boundary conditions. Journal of Computational Physics, 229(9):3523–3542, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000029X.

Lee:2010:LBS

[LL10a] Taehun Lee and Lin Liu. Lattice Boltzmann simulations of micron-scale drop impact on dry surfaces. Journal of Computational Physics, 229(20):8045–8063, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003761.

Lutjens:2010:XFI

[LL10b] Hinrich Lütjens and Jean-François Luciani. XTOR–2F: a fully implicit Newton–Krylov solver applied to nonlinear 3D extended MHD in tokamaks. Journal of Computational Physics, 229(21):8130–8143, October 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003967.

Li:2012:ACS

[LL12a] Yi Li and Anita T. Layton. Accurate computation of Stokes flow driven by an open immersed interface. Journal of Computational Physics, 231(15):5195–5215, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200201X.

Lu:2012:WMS

[LL12b] Wangtao Lu and Ya Yan Lu. Waveguide mode solver based on Neumann-to-Dirichlet operators and boundary integral equations. Journal of Computational Physics, 231(4):1360–1371, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006164.

Lin:2014:HRR

[LL14a] Bo-Wen Lin and Pong-Jeu Lu. High-resolution Roe’s scheme and characteristic boundary conditions for solving complex wave reflection phenomena in a tree-like arterial structure. Journal of Computational Physics, 260(??):143–162, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008413.

Lu:2014:EHO

[LL14b] Wangtao Lu and Ya Yan Lu. Efficient high order waveguide mode solvers based on boundary integral equations. Journal of Computational Physics, 272(??):507–525, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002915.

Lage:2011:FSF

[LLdSC11] Marcos Lage, Hélio Lopes, and Marcio da S. Carvalho. Flows with suspended and floating particles. Journal of Computational Physics, 230(20):7736–7754, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004049.

Liu:2010:NTS

[LLH10] Li Liu, Xiaodong Li, and Fang Q. Hu. Nonuniform time-step Runge–Kutta discontinuous Galerkin method for Computational Aeroacoustics. Journal of Computational Physics, 229(19):6874–6897, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002895.

Li:2010:OGS

[LLJ10] Kuan Li, Philip W. Livermore, and Andrew Jackson. An optimal Galerkin scheme to solve the kinematic dynamo eigenvalue problem in a full sphere. Journal of Computational Physics, 229(23):8666–8683, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004365.

Li:2010:FEH

[LLL10] Shaojing Li, Boris Livshitz, and Vitaliy Lomakin. Fast evaluation of Helmholtz potential on graphics processing units (GPUs). Journal of Computational Physics, 229(22):8463–8483, November 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004225.

Lin:2014:WGF

[LLMY14] Guang Lin, Jiangguo Liu, Lin Mu, and Xiu Ye. Weak Galerkin finite element methods for Darcy flow: Anisotropy and heterogeneity. Journal of Computational Physics, 276(??):422–437, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004793.

Luo:2010:RDG

[LLN^{+}10] Hong Luo, Luqing Luo, Robert Nourgaliev, Vincent A. Mousseau, and Nam Dinh. A reconstructed discontinuous Galerkin method for the compressible Navier–Stokes equations on arbitrary grids. Journal of Computational Physics, 229(19):6961–6978, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002949.

Liu:2010:SAP

[LLP10] Jian-Guo Liu, Jie Liu, and Robert L. Pego. Stable and accurate pressure approximation for unsteady incompressible viscous flow. Journal of Computational Physics, 229(9):3428– 3453, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000252.

Li:2014:HDG

[LLP14] Liang Li, Stéphane Lanteri, and Ronan Perrussel. A hybridizable discontinuous Galerkin method combined to a Schwarz algorithm for the solution of 3D time-harmonic Maxwell’s equation. Journal of Computational Physics, 256(??):563–581, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006086.

Lowrie:2011:PMQ

[LLS11] W. Lowrie, V. S. Lukin, and U. Shumlak. A priori mesh quality metric error analysis applied to a highorder finite element method. Journal of Computational Physics, 230(14):5564–5586, June 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001872.

Li:2014:GDS

[LLSW14] Jingzhi Li, Hongyu Liu, Yuanchang Sun, and Qi Wang. Ground detection by a single electromagnetic far-field measurement.

Journal of Computational Physics, 273(??):472–487, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003738.

Li:2014:EML

[LLW14] Jingzhi Li, Hongyu Liu, and Qi Wang. Enhanced multilevel linear sampling methods for inverse scattering problems. Journal of Computational Physics, 257(??):554–571, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006621.

Li:2011:CGR

[LLX11a] Jiequan Li, Qibing Li, and Kun Xu. Comparison of the generalized Riemann solver and the gas-kinetic scheme for inviscid compressible flow simulations. Journal of Computational Physics, 230(12):5080–5099, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001793.

Li:2011:ESB

[LLX11b] Jing Li, Jinglai Li, and Dongbin Xiu. An efficient surrogatebased method for computing rare failure probability. Journal of Computational Physics, 230(24):8683–8697, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004803.

Lin:2011:FCH

[LLY11] Lin Lin, Jianfeng Lu, and Lexing Ying. Fast construction of hierarchical matrix representation from matrix-vector multiplication. Journal of Computational Physics, 230(10):4071–4087, May 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001227.

Lin:2012:ALB

[LLYE12a] Lin Lin, Jianfeng Lu, Lexing Ying, and Weinan E. Adaptive local basis set for Kohn–Sham density functional theory in a discontinuous Galerkin framework I: Total energy calculation.

Journal of Computational Physics, 231(4):2140–2154, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006875.

Lin:2012:OLB

[LLYE12b] Lin Lin, Jianfeng Lu, Lexing Ying, and Weinan E. Optimized local basis set for Kohn–Sham density functional theory. Journal of Computational Physics, 231(13):4515–4529, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001507.

Leung:2011:GBP

[LLZ11] Shingyu Leung, John Lowengrub, and Hongkai Zhao. A grid based particle method for solving partial differential equations on evolving surfaces and modeling high order geometrical motion. Journal of Computational Physics, 230(7):2540–2561, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110007035.

Li:2014:AAB

[LLZ14] Weixuan Li, Guang Lin, and Dongxiao Zhang. An adaptive ANOVA-based PCKF for high-dimensional nonlinear inverse modeling. Journal of Computational Physics, 258(??):752–772, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300781X.

Loire:2013:SFA

[LM13] Sophie Loire and Igor Mezić. Spatial filter averaging approach of probabilistic method to linear second-order partial differential equations of the parabolic type. Journal of Computational Physics, 233(??):175–191, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004925.

Lipnikov:2014:HOM

[LM14a] K. Lipnikov and G. Manzini. A high-order mimetic method on unstructured polyhedral meshes for the diffusion equation.

Journal of Computational Physics, 272(??):360–385, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002848.

Lombard:2014:NMN

[LM14b] Bruno Lombard and Jean-François Mercier. Numerical modeling of nonlinear acoustic waves in a tube connected with Helmholtz resonators. Journal of Computational Physics, 259(??):421–443, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113007985.

Lipnikov:2011:MFD

[LMBB11] K. Lipnikov, G. Manzini, F. Brezzi, and A. Buffa. The mimetic finite difference method for the 3D magnetostatic field problems on polyhedral meshes. Journal of Computational Physics, 230(2):305–328, January 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005012.

LeMetayer:2011:DMC

[LMFH11] O. Le Métayer, A. Massol, N. Favrie, and S. Hank. A discrete model for compressible flows in heterogeneous media. Journal of Computational Physics, 230(7):2470–2495, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006844.

Li:2013:NFT

[LMGS13] Ying Li, Liquan Mei, Jiatai Ge, and Feng Shi. A new fractional time-stepping method for variable density incompressible flows. Journal of Computational Physics, 242(??):124–137, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001204.

Li:2013:BCT

[LMK13] Like Li, Renwei Mei, and James F. Klausner. Boundary conditions for thermal lattice Boltzmann equation method. Journal of Computational Physics, 237(??):366–395, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007073.

Lagrava:2012:AMD

[LMLC12] D. Lagrava, O. Malaspinas, J. Latt, and B. Chopard. Advances in multi-domain lattice Boltzmann grid refinement. Journal of Computational Physics, 231(14):4808–4822, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200157X.

Ling:2014:SMD

[LMM14] You Ling, Joshua Mullins, and Sankaran Mahadevan. Selection of model discrepancy priors in Bayesian calibration. Journal of Computational Physics, 276(??):665–680, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005518.

Longoni:2011:ABN

[LMQ^{+}11] Matteo Longoni, A. Cristiano I. Malossi, Alfio Quarteroni, Andrea Villa, and Paolo Ruffo. An ALE-based numerical technique for modeling sedimentary basin evolution featuring layer deformations and faults. Journal of Computational Physics, 230(8):3230–3248, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000520.

Loubere:2010:RRB

[LMS^{+}10] Raphaël Loubère, Pierre-Henri Maire, Mikhail Shashkov, Jérôme Breil, and Stéphane Galera. ReALE: a reconnectionbased arbitrary-Lagrangian–Eulerian method. Journal of Computational Physics, 229(12):4724–4761, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000121X.

Lipnikov:2011:AMC

[LMS11] K. Lipnikov, G. Manzini, and D. Svyatskiy. Analysis of the monotonicity conditions in the mimetic finite difference method for elliptic problems. Journal of Computational Physics, 230(7):2620–2642, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007138.

Lipnikov:2014:MFD

[LMS14] Konstantin Lipnikov, Gianmarco Manzini, and Mikhail Shashkov. Mimetic finite difference method. Journal of Computational Physics, 257(??):1163–1227, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005135.

Liang:2014:ECP

[LMZ14] Chunlei Liang, Koji Miyaji, and Bin Zhang. An efficient correction procedure via reconstruction for simulation of viscous flow on moving and deforming domains. Journal of Computational Physics, 256(??):55–68, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005901.

Lantos:2010:PML

[LN10a] Nicolas Lantos and Frédéric Nataf. Perfectly matched layers for the heat and advection-diffusion equations. Journal of Computational Physics, 229(24):9042–9052, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004419.

Lindstrom:2010:SHO

[LN10b] Jens Lindström and Jan Nordström. A stable and high-order accurate conjugate heat transfer problem. Journal of Computational Physics, 229(14):5440–5456, July 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001865.

Liu:2010:CIF

[LN10c] Xinfeng Liu and Qing Nie. Compact integration factor methods for complex domains and adaptive mesh refinement. Journal of Computational Physics, 229(16):5692–5706, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001798.

Lundquist:2014:SST

[LN14] Tomas Lundquist and Jan Nordström. The SBP–SAT technique for initial value problems. Journal of Computational Physics, 270(??):86–104, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002356.

LeMartelot:2013:LLG

[LNS13] S. LeMartelot, B. Nkonga, and R. Saurel. Liquid and liquid-gas flows at all speeds. Journal of Computational Physics, 255(??):53–82, December 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005263.

Lauritzen:2010:CSL

[LNU10] Peter H. Lauritzen, Ramachandran D. Nair, and Paul A. Ullrich. A conservative semi-Lagrangian multi-tracer transport scheme (CSLAM) on the cubed-sphere grid. Journal of Computational Physics, 229(5):1401–1424, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900597X.

Lee:2012:MMM

[LOG12] Dongryeol Lee, Arkadas Ozakin, and Alexander G. Gray. Multibody multipole methods. Journal of Computational Physics, 231(20):6827–6845, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003427.

Lombard:2011:NMT

[LP11] Bruno Lombard and Joël Piraux. Numerical modeling of transient two-dimensional viscoelastic waves. Journal of Computational Physics, 230(15):6099–6114, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002476.

Lau:2012:SST

[LP12a] Stephen R. Lau and Richard H. Price. Sparse spectraltau method for the three-dimensional helically reduced wave equation on two-center domains. Journal of Computational Physics, 231(22):7695–7714, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003658.

Lejay:2012:SDP

[LP12b] Antoine Lejay and Géraldine Pichot. Simulating diffusion processes in discontinuous media: a numerical scheme with constant time steps. Journal of Computational Physics, 231 (21):7299–7314, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003713.

LeChenadec:2013:CUF

[LP13a] Vincent Le Chenadec and Heinz Pitsch. Combination of 3D unsplit forward and backward volume-of-fluid transport and coupling to the level set method. Journal of Computational Physics, 233(??):10–33, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003920.

LeChenadec:2013:MPC

[LP13b] Vincent Le Chenadec and Heinz Pitsch. A monotonicity preserving conservative sharp interface flow solver for high density ratio two-phase flows. Journal of Computational Physics, 249(??):185–203, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002921.

Li:2014:ERK

[LP14a] Qin Li and Lorenzo Pareschi. Exponential Runge–Kutta for the inhomogeneous Boltzmann equations with high order of accuracy. Journal of Computational Physics, 259(??):402–420, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007821.

Liu:2014:AED

[LP14b] Hailiang Liu and Michael Pollack. Alternating evolution discontinuous Galerkin methods for Hamilton–Jacobi equations. Journal of Computational Physics, 258(??):31–46, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006529.

Lemieux:2011:IJF

[LPE^{+}11] Jean-François Lemieux, Stephen F. Price, Katherine J. Evans, Dana Knoll, Andrew G. Salinger, David M. Holland, and Antony J. Payne. Implementation of the Jacobian-free Newton–Krylov method for solving the firstorder ice sheet momentum balance. Journal of Computational Physics, 230(17):6531–6545, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002853.

Lin:2013:P

[LPK13] Ching-Long Lin, Grace C. Y. Peng, and George Karniadakis. Preface. Journal of Computational Physics, 244(??):1–3, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001137.

Leung:2010:BPF

[LQ10a] Shingyu Leung and Jianliang Qian. The backward phase flow and FBI-transform-based Eulerian Gaussian beams for the Schrödinger equation. Journal of Computational Physics, 229(23):8888–8917, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004523.

Li:2010:HWE

[LQ10b] Gang Li and Jianxian Qiu. Hybrid weighted essentially non-oscillatory schemes with different indicators. Journal of Computational Physics, 229(21):8105–8129, October 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003955.

Luo:2011:FSH

[LQ11] Songting Luo and Jianliang Qian. Factored singularities and high-order Lax–Friedrichs sweeping schemes for pointsource traveltimes and amplitudes. Journal of Computational Physics, 230(12):4742–4755, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001331.

Luo:2014:FHS

[LQB14] Songting Luo, Jianliang Qian, and Robert Burridge. Fast Huygens sweeping methods for Helmholtz equations in inhomogeneous media in the high frequency regime. Journal of Computational Physics, 270(??):378–401, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400254X.

Labrosse:2011:ONP

[LR11] G. Labrosse and A. Redondo. The optimal 3-node preconditioner of the d^{2}/dx^{2} Fourier and Chebyshev spectral operators. Journal of Computational Physics, 230(1):147–158, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005103.

Li:2012:MDL

[LR12] Wanai Li and Yu-Xin Ren. The multi-dimensional limiters for solving hyperbolic conservation laws on unstructured grids II: Extension to high order finite volume schemes. Journal of Computational Physics, 231(11):4053–4077, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000599.

Liu:2013:HME

[LR13] Qianlong Liu and Kenneth L. Reifsnider. Heterogeneous mixtures of elliptical particles: Directly resolving local and global properties and responses. Journal of Computational Physics, 235(??):161–181, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005840.

Lin:2014:FVE

[LRBT14] Zhigui Lin, Ricardo Ruiz-Baier, and Canrong Tian. Finite volume element approximation of an inhomogeneous Brusselator model with cross-diffusion. Journal of Computational Physics, 256(??):806–823, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006141.

Liseikin:2011:ACG

[LRK11] V. D. Liseikin, A. D. Rychkov, and A. V. Kofanov. Applications of a comprehensive grid method to solution of threedimensional boundary value problems. Journal of Computational Physics, 230(21):7755–7774, September 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003512.

Lee:2010:HFB

[LRL10] Seung-Cheol Lee, Vineet Rawat, and Jin-Fa Lee. A hybrid finite/boundary element method for periodic structures on non-periodic meshes using an interior penalty formulation for Maxwell’s equations. Journal of Computational Physics, 229(13):4934–4951, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001245.

Li:2011:MDL

[LRLL11] Wanai Li, Yu-Xin Ren, Guodong Lei, and Hong Luo. The multi-dimensional limiters for solving hyperbolic conservation laws on unstructured grids. Journal of Computational Physics, 230(21):7775–7795, September 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003822.

Lipnikov:2013:MDT

[LRN13] Konstantin Lipnikov, James Reynolds, and Eric Nelson. Mimetic discretization of two-dimensional magnetic diffusion equations. Journal of Computational Physics, 247(??):1–16, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002349.

Lebiedz:2010:MCT

[LRS10] Dirk Lebiedz, Volkmar Reinhardt, and Jochen Siehr. Minimal curvature trajectories: Riemannian geometry concepts for slow manifold computation in chemical kinetics. Journal of Computational Physics, 229(18):6512–6533, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000269X.

Leclaire:2013:PIL

[LRT13] Sébastien Leclaire, Marcelo Reggio, and Jean-Yves Trépanier. Progress and investigation on lattice Boltzmann modeling of multiple immiscible fluids or components with variable density and viscosity ratios. Journal of Computational Physics, 246(??):318–342, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002179. See [dv23].

Lin:2010:TLS

[LS10a] Paul T. Lin and John N. Shadid. Towards large-scale multisocket, multicore parallel simulations: Performance of an MPIonly semiconductor device simulator. Journal of Computational Physics, 229(19):6804–6818, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002846.

Lipnikov:2010:FDM

[LS10b] K. Lipnikov and M. Shashkov. A framework for developing a mimetic tensor artificial viscosity for Lagrangian hydrocodes on arbitrary polygonal meshes. Journal of Computational Physics, 229(20):7911–7941, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003694.

Le:2013:FSI

[LS13a] Trung Bao Le and Fotis Sotiropoulos. Fluid-structure interaction of an aortic heart valve prosthesis driven by an animated anatomic left ventricle. Journal of Computational Physics, 244(??):41–62, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004937.

Lei:2013:CPF

[LS13b] Siu-Long Lei and Hai-Wei Sun. A circulant preconditioner for fractional diffusion equations. Journal of Computational Physics, 242(??):715–725, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001447.

Lindblom:2013:SPD

[LS13c] Lee Lindblom and Béla Szilágyi. Solving partial differential equations numerically on manifolds with arbitrary spatial topologies. Journal of Computational Physics, 243(??):151– 175, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001502.

Liu:2013:TSD

[LS13d] Yang Liu and Mrinal K. Sen. Time-space domain dispersionrelation-based finite-difference method with arbitrary evenorder accuracy for the 2D acoustic wave equation. Journal of Computational Physics, 232(1):327–345, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004718.

Lenaers:2014:NHO

[LSBJ14] Peter Lenaers, Philipp Schlatter, Geert Brethouwer, and Arne V. Johansson. A new high-order method for the simulation of incompressible wall-bounded turbulent flows. Journal of Computational Physics, 272(??):108–126, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002976.

Lin:2013:EIM

[LSE13] Lin Lin, Sihong Shao, and Weinan E. Efficient iterative method for solving the Dirac–Kohn–Sham density functional theory. Journal of Computational Physics, 245(??):205–217, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002088.

Levine:2011:FAM

[LSK11] Benjamin G. Levine, John E. Stone, and Axel Kohlmeyer. Fast analysis of molecular dynamics trajectories with graphics processing units-radial distribution function histogramming. Journal of Computational Physics, 230(9):3556–3569, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000829.

Laadhari:2014:CDB

[LSM14] Aymen Laadhari, Pierre Saramito, and Chaouqi Misbah. Computing the dynamics of biomembranes by combining conservative level set and adaptive finite element methods. Journal of Computational Physics, 263(??):328–352, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008395.

Lipnikov:2010:MFV

[LSV10] K. Lipnikov, D. Svyatskiy, and Y. Vassilevski. A monotone finite volume method for advection-diffusion equations on unstructured polygonal meshes. Journal of Computational Physics, 229(11):4017–4032, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000586.

Liska:2010:OBS

[LSVW10] Richard Liska, Mikhail Shashkov, Pavel Váchal, and Burton Wendroff. Optimization-based synchronized flux-corrected conservative interpolation (remapping) of mass and momentum for arbitrary Lagrangian–Eulerian methods. Journal of Computational Physics, 229(5):1467–1497, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006007.

Lilienthal:2014:NDS

[LSW14] M. Lilienthal, S. M. Schnepp, and T. Weiland. Nondissipative space–time hp-discontinuous Galerkin method for the time-dependent Maxwell equations. Journal of Computational Physics, 275(??):589–607, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005087.

Lin:2013:LFI

[LSZ13] Tao Lin, Dongwoo Sheen, and Xu Zhang. A locking-free immersed finite element method for planar elasticity interface problems. Journal of Computational Physics, 247(??):228–247, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002374.

Le:2010:LDL

[LT10a] Duc-Vinh Le and Zhijun Tan. Large deformation of liquid capsules enclosed by thin shells immersed in the fluid. Journal of Computational Physics, 229(11):4097–4116, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000732.

Lindbo:2010:SAF

[LT10b] Dag Lindbo and Anna-Karin Tornberg. Spectrally accurate fast summation for periodic Stokes potentials. Journal of Computational Physics, 229(23):8994–9010, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004730.

LeFloch:2011:GTM

[LT11a] Philippe G. LeFloch and Mai Duc Thanh. A Godunov-type method for the shallow water equations with discontinuous topography in the resonant regime. Journal of Computational Physics, 230(20):7631–7660, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003810.

Lindbo:2011:SAF

[LT11b] Dag Lindbo and Anna-Karin Tornberg. Spectral accuracy in fast Ewald-based methods for particle simulations. Journal of Computational Physics, 230(24):8744–8761, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005092.

Long:2011:NWP

[LT11c] D. Long and J. Thuburn. Numerical wave propagation on non-uniform one-dimensional staggered grids. Journal of Computational Physics, 230(7):2643–2659, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000027.

Li:2012:ECD

[LT12] Yan Li and Zhen F. Tian. An exponential compact difference scheme for solving 2D steady magnetohydrodynamic (MHD) duct flow problems. Journal of Computational Physics, 231(16):5443–5468, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002343.

Li:2014:HOR

[LT14] Yong Li and Catalin Trenchea. A higher-order Robert– Asselin type time filter. Journal of Computational Physics, 259(??):23–32, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007845.

Lalescu:2010:IHO

[LTC10] C. C. Lalescu, B. Teaca, and D. Carati. Implementation of high order spline interpolations for tracking test particles in discretized fields. Journal of Computational Physics, 229(17):5862–5869, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199910900607X.

Lalescu:2013:INS

[LTC13] C. C. Lalescu, B. Teaca, and D. Carati. Influence of numerical schemes on statistical properties of computed charged particle trajectories in turbulent electromagnetic fields. Journal of Computational Physics, 241(??):546–557, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006097.

Lin:2014:AEC

[LTC14] Huimin Lin, Huazhong Tang, and Wei Cai. Accuracy and efficiency in computing electrostatic potential for an ion channel model in layered dielectric/electrolyte media. Journal of Computational Physics, 259(??):488–512, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008164.

Lee:2013:AOT

[LTCN13] Bok Jik Lee, Eleuterio F. Toro, Cristóbal E. Castro, and Nikolaos Nikiforakis. Adaptive Osher-type scheme for the Euler equations with highly nonlinear equations of state. Journal of Computational Physics, 246(??):165–183, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002246.

Lunati:2011:IMF

[LTL11] Ivan Lunati, Manav Tyagi, and Seong H. Lee. An iterative multiscale finite volume algorithm converging to the exact solution. Journal of Computational Physics, 230(5):1849–1864, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006510.

Li:2014:ECD

[LTL^{+}14] Zhen Li, Yu-Hang Tang, Huan Lei, Bruce Caswell, and George Em Karniadakis. Energy-conserving dissipative particle dynamics with temperature-dependent properties. Journal of Computational Physics, 265(??):113–127, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000977.

Lemieux:2010:INC

[LTS^{+}10] Jean-François Lemieux, Bruno Tremblay, Jan Sedláček, Paul Tupper, Stephen Thomas, David Huard, and Jean-Pierre Auclair. Improving the numerical convergence of viscous-plastic sea ice models with the Jacobian-free Newton–Krylov method. Journal of Computational Physics, 229(8):2840–2852, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006822.

Li:2010:SDM

[LTSH10] Ben-Wen Li, Shuai Tian, Ya-Song Sun, and Zhang-Mao Hu. Schur-decomposition for 3D matrix equations and its application in solving radiative discrete ordinates equations discretized by Chebyshev collocation spectral method. Journal of Computational Physics, 229(4):1198–1212, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005774.

Lin:2010:UQR

[LTT10] G. Lin, A. M. Tartakovsky, and D. M. Tartakovsky. Uncertainty quantification via random domain decomposition and probabilistic collocation on sparse grids. Journal of Computational Physics, 229(19):6995–7012, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002974.

Luo:2013:USO

[Luo13] Songting Luo. A uniformly second order fast sweeping method for eikonal equations. Journal of Computational Physics, 241(??):104–117, May 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000958.

Leroy:2014:USA

[LVFK14] A. Leroy, D. Violeau, M. Ferrand, and C. Kassiotis. Unified semi-analytical wall boundary conditions applied to 2-D incompressible SPH. Journal of Computational Physics, 261(??):106–129, March 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113008437.

Lucas:2010:FUF

[LvZB10] Peter Lucas, Alexander H. van Zuijlen, and Hester Bijl. Fast unsteady flow computations with a Jacobian-free Newton– Krylov algorithm. Journal of Computational Physics, 229(24):9201–9215, December 10, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004808.

Liu:2014:LBP

[LVZK14] Haihu Liu, Albert J. Valocchi, Yonghao Zhang, and Qinjun Kang. Lattice Boltzmann phase-field modeling of thermocapillary flows in a confined microchannel. Journal of Computational Physics, 256(??):334–356, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005986.

Le:2011:FTM

[LW11] Duc-Vinh Le and Sum Thai Wong. A front-tracking method with Catmull–Clark subdivision surfaces for studying liquid capsules enclosed by thin shells in shear flow. Journal of Computational Physics, 230(9):3538–3555, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000817.

Li:2013:TDH

[LW13a] Peijun Li and Aihua Wood. A two-dimensional Helmholtz equation solution for the multiple cavity scattering problem. Journal of Computational Physics, 240(??):100–120, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200753X.

Liu:2013:FAN

[LW13b] Zhi-Feng Liu and Xiao-Hong Wang. Finite analytic numerical method for two-dimensional fluid flow in heterogeneous porous media. Journal of Computational Physics, 235(??):286–301, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006468.

Liu:2014:FES

[LW14] Hailiang Liu and Zhongming Wang. A free energy satisfying finite difference method for Poisson–Nernst–Planck equations. Journal of Computational Physics, 268(??):363–376, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001624.

Liu:2010:MSE

[LWCR10] Jin Liu, Moran Wang, Shiyi Chen, and Mark O. Robbins. Molecular simulations of electroosmotic flows in rough nanochannels. Journal of Computational Physics, 229(20): 7834–7847, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003657.

Li:2013:ESF

[LWD13a] S. Li, W. Wu, and E. Darve. Erratum to “A fast algorithm for sparse matrix computations related to inversion” [J. Comput. Phys. 242 (2013) 915–945]. Journal of Computational Physics, 251(??):535–536, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004087. See [LWD13b].

Li:2013:FAS

[LWD13b] S. Li, W. Wu, and E. Darve. A fast algorithm for sparse matrix computations related to inversion. Journal of Computational Physics, 242(??):915–945, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000892. See erratum [LWD13a].

Liu:2011:LSA

[LWF11] Hailiang Liu, Zhongming Wang, and Rodney O. Fox. A level set approach for dilute non-collisional fluid-particle flows. Journal of Computational Physics, 230(4):920–936, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004778.

Liu:2014:SNA

[LWS14] Wenjie Liu, Boying Wu, and Jiebao Sun. Some numerical algorithms for solving the highly oscillatory secondorder initial value problems. Journal of Computational Physics, 276(??):235–251, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005269.

Lin:2013:MCS

[LWZ^{+}13] Y. Lin, F. Wang, X. Zheng, H. Gao, and L. Zhang. Monte Carlo simulation of the Ising model on FPGA. Journal of Computational Physics, 237(??):224–234, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200736X.

Li:2010:EFP

[LX10] Jing Li and Dongbin Xiu. Evaluation of failure probability via surrogate models. Journal of Computational Physics, 229(23):8966–8980, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004699.

Li:2012:AOE

[LX12] Fengyan Li and Liwei Xu. Arbitrary order exactly divergencefree central discontinuous Galerkin methods for ideal MHD equations. Journal of Computational Physics, 231(6):2655– 2675, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111007285.

Li:2013:ADIa

[LX13a] Limei Li and Da Xu. Alternating direction implicit-Euler method for the two-dimensional fractional evolution equation. Journal of Computational Physics, 236(??):157–168, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200664X.

Liu:2013:BFQ

[LX13b] Yan Liu and H. L. Xing. A boundary focused quadrilateral mesh generation algorithm for multi-material structures. Journal of Computational Physics, 232(1):516–528, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005104.

Li:2014:TSF

[LX14] Xueyang Li and Aiguo Xiao. Time-splitting finite difference method with the wavelet-adaptive grids for semiclassical Gross–Pitaevskii equation in supercritical case. Journal of Computational Physics, 267(??):146–161, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001454.

Li:2010:HOG

[LXF10] Qibing Li, Kun Xu, and Song Fu. A high-order gaskinetic Navier–Stokes flow solver. Journal of Computational Physics, 229(19):6715–6731, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002809.

Luo:2012:HWR

[LXL^{+}12] Hong Luo, Yidong Xia, Shujie Li, Robert Nourgaliev, and Chunpei Cai. A Hermite WENO reconstruction-based discontinuous Galerkin method for the Euler equations on tetrahedral grids. Journal of Computational Physics, 231(16):5489–5503, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002355.

Li:2013:ADIb

[LXL13] Limei Li, Da Xu, and Man Luo. Alternating direction implicit Galerkin finite element method for the two-dimensional fractional diffusion-wave equation. Journal of Computational Physics, 255(??):471–485, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300569X.

Le:2014:TDG

[LXM14] N. T. P. Le, H. Xiao, and R. S. Myong. A triangular discontinuous Galerkin method for non-Newtonian implicit constitutive models of rarefied and microscale gases. Journal of Computational Physics, 273(??):160–184, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003593.

Luo:2013:RDG

[LXS^{+}13] Hong Luo, Yidong Xia, Seth Spiegel, Robert Nourgaliev, and Zonglin Jiang. A reconstructed discontinuous Galerkin method based on a Hierarchical WENO reconstruction for compressible flows on tetrahedral grids. Journal of Computational Physics, 236(??):477–492, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007061.

Lind:2012:ISP

[LXSR12] S. J. Lind, R. Xu, P. K. Stansby, and B. D. Rogers. Incompressible smoothed particle hydrodynamics for free-surface flows: a generalised diffusion-based algorithm for stability and validations for impulsive flows and propagating waves. Journal of Computational Physics, 231(4):1499–1523, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006279.

Li:2011:CDG

[LXY11] Fengyan Li, Liwei Xu, and Sergey Yakovlev. Central discontinuous Galerkin methods for ideal MHD equations with the exactly divergence-free magnetic field. Journal of Computational Physics, 230(12):4828–4847, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001495.

Lee:2013:IGC

[LY13a] Jinmo Lee and Donghyun You. An implicit ghost-cell immersed boundary method for simulations of moving body problems with control of spurious force oscillations. Journal of Computational Physics, 233(??):295–314, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005128.

Liang:2013:SFO

[LY13b] Dong Liang and Qiang Yuan. The spatial fourth-order energy-conserved S-FDTD scheme for Maxwell’s equations. Journal of Computational Physics, 243(??):344–364, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001599.

Liu:2014:EES

[LY14] Hailiang Liu and Hui Yu. Entropy/energy stable schemes for evolutionary dispersal models. Journal of Computational Physics, 256(??):656–677, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005706.

Lun:2010:SCF

[LYD10] Lisa Lun, Andrew Yeckel, and Jeffrey J. Derby. A Schur complement formulation for solving free-boundary, Stefan problems of phase change. Journal of Computational Physics, 229(20):7942–7955, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003700.

Li:2010:MCM

[LYE10] Xiantao Li, Jerry Z. Yang, and Weinan E. A multiscale coupling method for the modeling of dynamics of solids with application to brittle cracks. Journal of Computational Physics, 229(10):3970–3987, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000707.

Lin:2014:PIM

[LYJ14] Fu-Rong Lin, Shi-Wei Yang, and Xiao-Qing Jin. Preconditioned iterative methods for fractional diffusion equation. Journal of Computational Physics, 256(??):109–117, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005378.

Liu:2014:UGK

[LYXZ14] Sha Liu, Pubing Yu, Kun Xu, and Chengwen Zhong. Unified gas-kinetic scheme for diatomic molecular simulations in all flow regimes. Journal of Computational Physics, 259(??):96–113, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113007924.

Liu:2010:PFM

[LZ10] Haihu Liu and Yonghao Zhang. Phase-field modeling droplet dynamics with soluble surfactants. Journal of Computational Physics, 229(24):9166–9187, December 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000478X.

Li:2013:AGS

[LZ13] Jiequan Li and Yongjin Zhang. The adaptive GRP scheme for compressible fluid flows over unstructured meshes. Journal of Computational Physics, 242(??):367–386, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001022.

Liang:2013:SSA

[LZC^{+}13] Gongyou Liang, Zhong Zeng, Yu Chen, Junya Onishi, Hirotada Ohashi, and Shiyi Chen. Simulation of selfassemblies of colloidal particles on the substrate using a lattice Boltzmann pseudo-solid model. Journal of Computational Physics, 248(??):323–338, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002568.

Liu:2012:MST

[LZV12] Haihu Liu, Yonghao Zhang, and Albert J. Valocchi. Modeling and simulation of thermocapillary flows using lattice Boltzmann method. Journal of Computational Physics, 231 (12):4433–4453, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112000940.

Li:2014:CDG

[LZZC14] Xiang-Gui Li, Jiang Zhu, Rong-Pei Zhang, and Shengshan Cao. A combined discontinuous Galerkin method for the dipolar Bose–Einstein condensation. Journal of Computational Physics, 275(??):363–376, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004938.

Lv:2010:NCL

[LZZR10] Xin Lv, Qingping Zou, Yong Zhao, and Dominic Reeve. A novel coupled level set and volume of fluid method for sharp interface capturing on 3D tetrahedral grids. Journal of Computational Physics, 229(7):2573–2604, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006767.

Liu:2013:NCC

[LZZS13] Xuliang Liu, Shuhai Zhang, Hanxin Zhang, and Chi-Wang Shu. A new class of central compact schemes with spectrallike resolution I: Linear schemes. Journal of Computational Physics, 248(??):235–256, September 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002635.

Mattsson:2013:SSI

[MA13] Ken Mattsson and Martin Almquist. A solution to the stability issues with block norm summation by parts operators. Journal of Computational Physics, 253(??):418–442, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004865.

Macedo:2014:AFS

[MA14] Rodrigo Panosso Macedo and Marcus Ansorg. Axisymmetric fully spectral code for hyperbolic equations. Journal of Computational Physics, 276(??):357–379, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005336.

Maire:2013:NSO

[MAB^{+}13] Pierre-Henri Maire, Rémi Abgrall, Jérôme Breil, Raphaël Loubère, and Bernard Rebourcet. A nominally secondorder cell-centered Lagrangian scheme for simulating elasticplastic flows on two-dimensional unstructured grids. Journal of Computational Physics, 235(??):626–665, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006183.

Macrossan:2010:SNN

[Mac10] M. N. Macrossan. Searching for a near neighbor particle in DSMC cells using pseudo-subcells. Journal of Computational Physics, 229(17):5857–5861, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002299.

Mattsson:2014:ODN

[MAC14] Ken Mattsson, Martin Almquist, and Mark H. Carpenter. Optimal diagonal-norm SBP operators. Journal of Computational Physics, 264(??):91–111, May 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008498.

Mohebbi:2013:HOU

[MAD13] Akbar Mohebbi, Mostafa Abbaszadeh, and Mehdi Dehghan. A high-order and unconditionally stable scheme for the modified anomalous fractional sub-diffusion equation with a nonlinear source term. Journal of Computational Physics, 240(??):36– 48, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000430.

Menz:2014:SSP

[MAK14] William J. Menz, Jethro Akroyd, and Markus Kraft. Stochastic solution of population balance equations for reactor networks. Journal of Computational Physics, 256(??):615–629, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006268.

Mani:2012:AON

[Man12] Ali Mani. Analysis and optimization of numerical sponge layers as a nonreflective boundary treatment. Journal of Computational Physics, 231(2):704–716, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006176.

Magalhaes:2013:AMF

[MAPP13] João P. P. Magalhães, Duarte M. S. Albuquerque, José M. C. Pereira, and José C. F. Pereira. Adaptive mesh finitevolume calculation of 2D lid-cavity corner vortices. Journal of Computational Physics, 243(??):365–381, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001617.

Marksteiner:2010:UTI

[Mar10a] Q. R. Marksteiner. The use of tricubic interpolation with spectral derivatives to integrate particle trajectories in complicated electromagnetic fields. Journal of Computational Physics, 229(19):6688–6695, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000272X.

Marquina:2010:DFC

[Mar10b]  Antonio Marquina.  Diffusion front capturing schemes

for a class of Fokker–Planck equations: Application to the relativistic heat equation. Journal of Computational Physics, 229(7):2659–2674, April 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006858.

Martinsson:2013:DSV

[Mar13] P. G. Martinsson. A direct solver for variable coefficient elliptic PDEs discretized via a composite spectral collocation method. Journal of Computational Physics, 242(??):460–479, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001320.

Mattsson:2014:DNS

[Mat14] Ken Mattsson. Diagonal-norm summation by parts operators for finite difference approximations of third and fourth derivatives. Journal of Computational Physics, 274(??):432–454, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004343.

Maier:2010:LBA

[MB10a] R. S. Maier and R. S. Bernard. Lattice-Boltzmann accuracy in pore-scale flow simulation. Journal of Computational Physics, 229(2):233–255, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004987.

Monard:2010:ASF

[MB10b] François Monard and Guillaume Bal. An accurate solver for forward and inverse transport. Journal of Computational Physics, 229(13):4952–4979, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001269.

Mata:2011:NSP

[MB11] Matthew R. Mata and Andrea L. Bertozzi. A numerical scheme for particle-laden thin film flow in two dimensions. Journal of Computational Physics, 230(16):6334–6353, July 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002774.

Maire:2012:NSO

[MB12] Pierre-Henri Maire and Jérôme Breil. A nominally second-order accurate finite volume cell-centered scheme for anisotropic diffusion on two-dimensional unstructured grids. Journal of Computational Physics, 231(5):2259–2299, March 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100684X.

Ma:2013:NMT

[MB13a] C. Ma and D. Bothe. Numerical modeling of thermocapillary two-phase flows with evaporation using a twoscalar approach for heat transfer. Journal of Computational Physics, 233(??):552–573, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005426.

Mazzeo:2013:MDF

[MB13b] Brian A. Mazzeo and David D. Busath. From molecular dynamics to fluorescence anisotropy of fluorophores bound to oriented structures. Journal of Computational Physics, 232(1):482–497, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200513X.

Meyer:2014:SRK

[MBA14] Chad D. Meyer, Dinshaw S. Balsara, and Tariq D. Aslam. A stabilized runge–kutta-legendre method for explicit supertime-stepping of parabolic and mixed equations. Journal of Computational Physics, 257(??):594–626, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005597.

Moguen:2013:SIC

[MBD13] Yann Moguen, Pascal Bruel, and Erik Dick. Semi-implicit characteristic-based boundary treatment for acoustics in low Mach number flows. Journal of Computational Physics, 255 (??):339–361, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005573.

Muller:2013:CBA

[MBGW13] Andreas Müller, Jörn Behrens, Francis X. Giraldo, and Volkmar Wirth. Comparison between adaptive and uniform discontinuous Galerkin simulations in dry 2D bubble experiments. Journal of Computational Physics, 235(??):371–393, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006419.

Metivier:2014:RAL

[MBL^{+}14] L. Métivier, R. Brossier, S. Labbé, S. Operto, and J. Virieux. A robust absorbing layer method for anisotropic seismic wave modeling. Journal of Computational Physics, 279(??):218– 240, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114006330.

Meng:2010:MCN

[MBLM10] Jun Meng, Amir Boag, Vitaliy Lomakin, and Eric Michielssen. A multilevel Cartesian non-uniform grid time domain algorithm. Journal of Computational Physics, 229(22):8430–8444, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004195.

Mazzotti:2013:BEF

[MBMV13] M. Mazzotti, I. Bartoli, A. Marzani, and E. Viola. A 2.5D boundary element formulation for modeling damped waves in arbitrary cross-section waveguides and cavities. Journal of Computational Physics, 248(??):363–382, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002623.

Macdonald:2011:SEP

[MBR11] Colin B. Macdonald, Jeremy Brandman, and Steven J. Ruuth. Solving eigenvalue problems on curved surfaces using the Closest Point Method. Journal of Computational Physics, 230(22):7944–7956, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003858.

Mao:2013:CGO

[MBS13] Xuerui Mao, Hugh M. Blackburn, and Spencer J. Sherwin. Calculation of global optimal initial and boundary perturbations for the linearised incompressible Navier–Stokes equations. Journal of Computational Physics, 235(??):258–273, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006535.

Magni:2012:ANO

[MC12] Adrien Magni and Georges-Henri Cottet. Accurate, nonoscillatory, remeshing schemes for particle methods. Journal of Computational Physics, 231(1):152–172, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005237.

Marrone:2013:ASM

[MCA^{+}13] S. Marrone, A. Colagrossi, M. Antuono, G. Colicchio, and G. Graziani. An accurate SPH modeling of viscous flows around bodies at low and moderate Reynolds numbers. Journal of Computational Physics, 245(??):456–475, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001885.

McCaslin:2014:FMA

[MCD14a] Jeremy O. McCaslin, Émilien Courtine, and Olivier Desjardins. A fast marching approach to multidimensional extrapolation. Journal of Computational Physics, 274(??):393–412, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004306.

McDermott:2014:VDC

[McD14b] Randall J. McDermott. A velocity divergence constraint for large-eddy simulation of low-Mach flows. Journal of Computational Physics, 274(??):413–431, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004264.

Montecinos:2012:CSG

[MCDT12] G. Montecinos, C. E. Castro, M. Dumbser, and E. F. Toro. Comparison of solvers for the generalized Riemann problem for hyperbolic systems with source terms. Journal of Computational Physics, 231(19):6472–6494, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200318X.

Malecha:2014:MAS

[MCJ14] Ziemowit Malecha, Greg Chini, and Keith Julien. A multiscale algorithm for simulating spatially-extended Langmuir circulation dynamics. Journal of Computational Physics, 271(??):131–150, August 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113004749.

Marrone:2010:FFS

[MCLG10] S. Marrone, A. Colagrossi, D. Le Touzé, and G. Graziani. Fast free-surface detection and level-set function definition in SPH solvers. Journal of Computational Physics, 229(10):3652–3663, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000343.

Mackay:2013:CMP

[MD13] F. E. Mackay and C. Denniston. Coupling MD particles to a lattice-Boltzmann fluid through the use of conservative forces. Journal of Computational Physics, 237(??):289–298, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007188.

McCaslin:2014:LRI

[MD14] Jeremy O. McCaslin and Olivier Desjardins. A localized reinitialization equation for the conservative level set method. Journal of Computational Physics, 262(??):408–426, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000448.

Meyer:2010:CII

[MDH^{+}10] M. Meyer, A. Devesa, S. Hickel, X. Y. Hu, and N. A. Adams. A conservative immersed interface method for Large-Eddy simulation of incompressible flows. Journal of Computational Physics, 229(18):6300–6317, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002275.

Monasse:2012:CCA

[MDM^{+}12] L. Monasse, V. Daru, C. Mariotti, S. Piperno, and C. Tenaud. A conservative coupling algorithm between a compressible flow and a rigid body using an Embedded Boundary method. Journal of Computational Physics, 231(7):2977–2994, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000101.

Mai-Duy:2013:NSS

[MDPTK13] N. Mai-Duy, N. Phan-Thien, and B. C. Khoo. A numerical study of strongly overdamped Dissipative Particle Dynamics (DPD) systems. Journal of Computational Physics, 245(??):150–159, July 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001903.

Mai-Duy:2011:CLI

[MDTC11] N. Mai-Duy and T. Tran-Cong. Compact local integrated-RBF approximations for second-order elliptic differential problems. Journal of Computational Physics, 230(12):4772–4794, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001355.

Mai-Duy:2013:CFP

[MDTC13] N. Mai-Duy and T. Tran-Cong. A compact five-point stencil based on integrated RBFs for 2D second-order differential problems. Journal of Computational Physics, 235(??):302–321, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006523.

Mauger:2013:GAF

[MdVBS13] Sarah Mauger, Guillaume Colin de Verdière, Luc Bergé, and Stefan Skupin. GPU accelerated fully space and time resolved numerical simulations of self-focusing laser beams in SBS-active media. Journal of Computational Physics, 235(??):606–625, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006377.

Melchionna:2011:ISS

[Mel11] Simone Melchionna. Incorporation of smooth spherical bodies in the Lattice Boltzmann method. Journal of Computational Physics, 230(10):3966–3976, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001100.

Mendl:2012:EAT

[Men12] Christian B. Mendl. Efficient algorithm for two-center Coulomb and exchange integrals of electronic prolate spheroidal orbitals. Journal of Computational Physics, 231(15):5157–5175, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002033.

Michoski:2010:DGM

[MESV10] C. Michoski, J. A. Evans, P. G. Schmitz, and A. Vasseur. A discontinuous Galerkin method for viscous compressible multifluids. Journal of Computational Physics, 229(6):2249–2266, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006627.

Metivier:2011:IOF

[Mét11] Ludovic Métivier. Interlocked optimization and fast gradient algorithm for a seismic inverse problem. Journal of Computational Physics, 230(19):7502–7518, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003834.

Mikyska:2010:IHO

[MF10] Jirí Mikyska and Abbas Firoozabadi. Implementation of higher-order methods for robust and efficient compositional simulation. Journal of Computational Physics, 229(8):2898– 2913, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109007013.

Maddison:2012:DIU

[MF12] J. R. Maddison and P. E. Farrell. Directional integration on unstructured meshes via supermesh construction. Journal of Computational Physics, 231(12):4422–4432, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000885.

Moortgat:2013:HOC

[MF13] Joachim Moortgat and Abbas Firoozabadi. Higher-order compositional modeling of three-phase flow in 3D fractured porous media based on cross-flow equilibrium. Journal of Computational Physics, 250(??):425–445, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003409.

Main:2014:SOT

[MF14] Alex Main and Charbel Farhat. A second-order time-accurate implicit finite volume method with exact two-phase Riemann problems for compressible multi-phase fluid and fluid-structure problems. Journal of Computational Physics, 258(??):613–633, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300747X.

Markov:2012:SFM

[MFM12] Andrey A. Markov, Igor A. Filimonov, and Karen S. Martirosyan. Simulation of front motion in a reacting condensed two phase mixture. Journal of Computational Physics, 231(20):6714–6724, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003063.

Majda:2014:NPS

[MG14a] Andrew J. Majda and Ian Grooms. New perspectives on superparameterization for geophysical turbulence. Journal of Computational Physics, 271(??):60–77, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006190.

Mirzadeh:2014:CDP

[MG14b] Mohammad Mirzadeh and Frédéric Gibou. A conservative discretization of the Poisson–Nernst–Planck equations on adaptive Cartesian grids. Journal of Computational Physics, 274(??):633–653, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400446X.

Ma:2010:DCA

[MGJZ10] Xia Ma, Paul T. Giguere, Balaji Jayaraman, and Duan Z. Zhang. Distribution coefficient algorithm for small mass nodes in material point method. Journal of Computational Physics, 229(20):7819–7833, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003645.

Meier:2012:MOB

[MGLS12] E. T. Meier, A. H. Glasser, V. S. Lukin, and U. Shumlak. Modeling open boundaries in dissipative MHD simulation. Journal of Computational Physics, 231(7):2963–2976, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000137.

Murillo:2010:EBC

[MGN10a] J. Murillo and P. García-Navarro. An Exner-based coupled model for two-dimensional transient flow over erodible bed. Journal of Computational Physics, 229(23):8704–8732, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004432.

Murillo:2010:WSP

[MGN10b] J. Murillo and P. García-Navarro. Weak solutions for partial differential equations with source terms: Application to the shallow water equations. Journal of Computational Physics, 229(11):4327–4368, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000896.

Murillo:2011:IRS

[MGN11] J. Murillo and P. García-Navarro. Improved Riemann solvers for complex transport in two-dimensional unsteady shallow flow. Journal of Computational Physics, 230(19):7202–7239, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003330.

Murillo:2012:AVH

[MGN12a] J. Murillo and P. García-Navarro. Augmented versions of the HLL and HLLC Riemann solvers including source terms in one and two dimensions for shallow flow applications. Journal of Computational Physics, 231(20):6861–6906, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003464.

Murillo:2012:WRD

[MGN12b] J. Murillo and P. García-Navarro. Wave Riemann description of friction terms in unsteady shallow flows: Application to water and mud/debris floods. Journal of Computational Physics, 231(4):1963–2001, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006693.

Murillo:2013:EBN

[MGN13] J. Murillo and P. García-Navarro. Energy balance numerical schemes for shallow water equations with discontinuous topography. Journal of Computational Physics, 236(??):119–142, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006584.

Mendez:2014:USS

[MGN14a] S. Mendez, E. Gibaud, and F. Nicoud. An unstructured solver for simulations of deformable particles in flows at arbitrary Reynolds numbers. Journal of Computational Physics, 256(??):465–483, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006074.

Murillo:2014:ANM

[MGN14b] J. Murillo and P. García-Navarro. Accurate numerical modeling of 1D flow in channels with arbitrary shape. Application of the energy balanced property. Journal of Computational Physics, 260(??):222–248, March 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008486.

McClarren:2010:RAF

[MH10] Ryan G. McClarren and Cory D. Hauck. Robust and accurate filtered spherical harmonics expansions for radiative transfer. Journal of Computational Physics, 229(16):5597–5614, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001622.

Mascarenhas:2010:CMG

[MHA10] Brendan S. Mascarenhas, Brian T. Helenbrook, and Harold L. Atkins. Coupling p-multigrid to geometric multigrid for discontinuous Galerkin formulations of the convection-diffusion equation. Journal of Computational Physics, 229(10):3664–3674, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000434.

Munafo:2014:SLB

[MHGM14] Alessandro Munafò, Jeffrey R. Haack, Irene M. Gamba, and Thierry E. Magin. A spectral-Lagrangian Boltzmann solver for a multi-energy level gas. Journal of Computational Physics, 264(??):152–176, May 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000631.

Morales-Hernandez:2012:LTS

[MHGNM12] M. Morales-Hernandez, P. García-Navarro, and J. Murillo. A large time step 1D upwind explicit scheme (CFL > 1 ): Application to shallow water equations. Journal of Computational Physics, 231(19):6532–6557, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003324.

Morales-Hernandez:2014:ELT

[MHHGN14] M. Morales-Hernández, M. E. Hubbard, and P. García-Navarro. A 2D extension of a Large Time Step explicit scheme (CFL > 1 ) for unsteady problems with wet/dry boundaries. Journal of Computational Physics, 263(??):303–327, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000461.

Markidis:2014:FKP

[MHL^{+}14] Stefano Markidis, Pierre Henri, Giovanni Lapenta, Kjell Rönnmark, Maria Hamrin, Zakaria Meliani, and Erwin Laure. The Fluid-Kinetic Particle-in-Cell method for plasma simulations. Journal of Computational Physics, 271(??):415–429, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000965.

Mieussens:2013:APP

[Mie13] Luc Mieussens. On the asymptotic preserving property of the unified gas kinetic scheme for the diffusion limit of linear kinetic models. Journal of Computational Physics, 253 (??):138–156, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113004737.

Mignone:2014:HOC

[Mig14] A. Mignone. High-order conservative reconstruction schemes for finite volume methods in cylindrical and spherical coordinates. Journal of Computational Physics, 270(??):784–814, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002538.

May:2010:HMM

[MIJ10] Georg May, Francesca Iacono, and Antony Jameson. A hybrid multilevel method for high-order discretization of the Euler equations on unstructured meshes. Journal of Computational Physics, 229(10):3938–3956, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000598.

Motamarri:2012:HOA

[MIKG12] Phani Motamarri, Mrinal Iyer, Jaroslaw Knap, and Vikram Gavini. Higher-order adaptive finite-element methods for orbital-free density functional theory. Journal of Computational Physics, 231(20):6596–6621, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002185.

Min:2010:RLS

[Min10a] Chohong Min. On reinitializing level set functions. Journal of Computational Physics, 229(8):2764–2772, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109007189.

Miniati:2010:HSG

[Min10b] Francesco Miniati. A hybrid scheme for gas-dust systems stiffly coupled via viscous drag. Journal of Computational Physics, 229(10):3916–3937, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000574.

Minjeaud:2013:APC

[Min13]  Sebastian Minjeaud.  An adaptive pressure correction

method without spurious velocities for diffuse-interface models of incompressible flows. Journal of Computational Physics, 236(??):143–156, March 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007024.

Mitran:2013:CKM

[Mit13] Sorin Mitran. Continuum-kinetic-microscopic model of lung clearance due to core-annular fluid entrainment. Journal of Computational Physics, 244(??):193–211, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000909.

Meyer:2013:ACE

[MJ13a] Daniel W. Meyer and Patrick Jenny. Accurate and computationally efficient mixing models for the simulation of turbulent mixing with PDF methods. Journal of Computational Physics, 247(??):192–207, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300243X.

Movahed:2013:SAM

[MJ13b] Pooya Movahed and Eric Johnsen. A solution-adaptive method for efficient compressible multifluid simulations, with application to the Richtmyer–Meshkov instability. Journal of Computational Physics, 239(??):166–186, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000478.

Mennemann:2014:PML

[MJ14] Jan-Frederik Mennemann and Ansgar Jüngel. Perfectly matched layers versus discrete transparent boundary conditions in quantum device simulations. Journal of Computational Physics, 275(??):1–24, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400463X.

Mennemann:2013:TSP

[MJK13] Jan-Frederik Mennemann, Ansgar Jüngel, and Hans Kosina. Transient Schrödinger–Poisson simulations of a high-frequency resonant tunneling diode oscillator. Journal of Computational Physics, 239(??):187–205, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007401.

Muller:2013:MMC

[MJM13] Florian Müller, Patrick Jenny, and Daniel W. Meyer. Multilevel Monte Carlo for two phase flow and Buckley–Leverett transport in random heterogeneous porous media. Journal of Computational Physics, 250(??):685–702, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002015.

Minoli:2011:DGS

[MK11] Cesar A. Acosta Minoli and David A. Kopriva. Discontinuous Galerkin spectral element approximations on moving meshes. Journal of Computational Physics, 230(5):1876–1902, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006534.

Minoli:2012:BSR

[MK12] Cesar A. Acosta Minoli and David A. Kopriva. Boundary states at reflective moving boundaries. Journal of Computational Physics, 231(11):4160–4184, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000915.

Markish:2013:CFG

[MK13] Ofer Markish and Raphael Kastner. Cylindrical FDTD grid-compatible Green’s functions. Journal of Computational Physics, 240(??):198–210, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007425.

Markesteijn:2014:TAR

[MK14a] A. P. Markesteijn and S. A. Karabasov. Time asynchronous relative dimension in space method for multiscale problems in fluid dynamics. Journal of Computational Physics, 258(??):137–164, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007110.

Morinishi:2014:SSC

[MK14b] Yohei Morinishi and Kazuki Koga. Skew-symmetric convection form and secondary conservative finite difference methods for moving grids. Journal of Computational Physics, 257(??):1081–1112, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000934.

Moguen:2012:PVC

[MKB^{+}12] Yann Moguen, Tarik Kousksou, Pascal Bruel, Jan Vierendeels, and Erik Dick. Pressure-velocity coupling allowing acoustic calculation in low Mach number flow. Journal of Computational Physics, 231(16):5522–5541, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002240.

Morgan:2013:ATC

[MKB^{+}13] Nathaniel R. Morgan, Mark A. Kenamond, Donald E. Burton, Theodore C. Carney, and Daniel J. Ingraham. An approach for treating contact surfaces in Lagrangian cell-centered hydrodynamics. Journal of Computational Physics, 250(??):527–554, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300346X.

Mabuza:2014:CPP

[MKCB14] Sibusiso Mabuza, Dmitri Kuzmin, Suncica Canić, and Martina Bukac. A conservative, positivity preserving scheme for reactive solute transport problems in moving domains. Journal of Computational Physics, 276(??):563–595, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005427.

Marras:2012:VMS

[MKGV12] S. Marras, J. F. Kelly, F. X. Giraldo, and M. Vázquez. Variational multiscale stabilization of high-order spectral elements for the advection-diffusion equation. Journal of Computational Physics, 231(21):7187–7213, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003439.

Markidis:2011:ECP

[ML11a] Stefano Markidis and Giovanni Lapenta. The energy conserving particle-in-cell method. Journal of Computational Physics, 230(18):7037–7052, August 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003445.

Min:2011:SED

[ML11b] Misun Min and Taehun Lee. A spectral-element discontinuous Galerkin lattice Boltzmann method for nearly incompressible flows. Journal of Computational Physics, 230(1):245–259, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005279.

Misiats:2013:SOA

[ML13] Oleksandr Misiats and Konstantin Lipnikov. Second-order accurate monotone finite volume scheme for Richards’ equation. Journal of Computational Physics, 239(??):123–137, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005232.

Mata:2014:VID

[ML14a] Pablo Mata and Adrian J. Lew. Variational integrators for the dynamics of thermo-elastic solids with finite speed thermal waves. Journal of Computational Physics, 257(??):1423–1443, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300644X.

Moyner:2014:MTP

[ML14b] Olav Møyner and Knut-Andreas Lie. A multiscale twopoint flux-approximation method. Journal of Computational Physics, 275(??):273–293, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004835.

Morgan:2014:LSG

[MLBK14] Nathaniel R. Morgan, Konstantin N. Lipnikov, Donald E. Burton, and Mark A. Kenamond. A Lagrangian staggered grid Godunov-like approach for hydrodynamics. Journal of Computational Physics, 259(??):568–597, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008127.

Morales:2014:SCM

[MLBS14] Jorge A. Morales, Matthieu Leroy, Wouter J. T. Bos, and Kai Schneider. Simulation of confined magnetohydrodynamic flows with Dirichlet boundary conditions using a pseudospectral method with volume penalization. Journal of Computational Physics, 274(??):64–94, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400401X.

Murillo:2012:RSU

[MLGN12] J. Murillo, B. Latorre, and P. García-Navarro. A Riemann solver for unsteady computation of 2D shallow flows with variable density. Journal of Computational Physics, 231 (14):4775–4807, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112001581.

Moulla:2012:PSM

[MLM12] R. Moulla, L. Lefévre, and B. Maschke. Pseudo-spectral methods for the spatial symplectic reduction of open systems of conservation laws. Journal of Computational Physics, 231(4):1272–1292, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006012.

Mani:2010:EEA

[MM10] Karthik Mani and Dimitri J. Mavriplis. Error estimation and adaptation for functional outputs in time-dependent flow problems. Journal of Computational Physics, 229(2):415–440, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005269.

Mittal:2011:HDO

[MM11] Arpit Mittal and Sandip Mazumder. Hybrid discrete ordinatesspherical harmonics solution to the Boltzmann transport equation for phonons for non-equilibrium heat conduction. Journal of Computational Physics, 230(18):6977–7001, August 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003354.

Metivier:2012:SSI

[MM12] Ludovic Métivier and Philippe Montarnal. Strategies for solving index one DAE with non-negative constraints: Application to liquid-liquid extraction. Journal of Computational Physics, 231(7):2945–2962, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200006X.

Minoshima:2011:MMA

[MMA11] Takashi Minoshima, Yosuke Matsumoto, and Takanobu Amano. Multi-moment advection scheme for Vlasov simulations. Journal of Computational Physics, 230(17):6800–6823, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003147.

Minoshima:2013:MMA

[MMA13] Takashi Minoshima, Yosuke Matsumoto, and Takanobu Amano. Multi-moment advection scheme in three dimension for Vlasov simulations of magnetized plasma. Journal of Computational Physics, 236(??):81–95, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007048.

Michoski:2011:AHT

[MMD^{+}11] C. Michoski, C. Mirabito, C. Dawson, D. Wirasaet, E. J. Kubatko, and J. J. Westerink. Adaptive hierarchic transformations for dynamically p-enriched slope-limiting over discontinuous Galerkin systems of generalized equations. Journal of Computational Physics, 230(22):8028–8056, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004141.

Muddle:2012:EPM

[MMH12] Richard L. Muddle, Milan Mihajlović, and Matthias Heil. An efficient preconditioner for monolithically-coupled largedisplacement fluid-structure interaction problems with pseudosolid mesh updates. Journal of Computational Physics, 231 (21):7315–7334, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003579.

Mast:2012:MKL

[MMHA^{+}12] C. M. Mast, P. Mackenzie-Helnwein, P. Arduino, G. R. Miller, and W. Shin. Mitigating kinematic locking in the material point method. Journal of Computational Physics, 231(16):5351–5373, June 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002148.

Michoski:2014:DGM

[MMIW14] C. Michoski, D. Meyerson, T. Isaac, and F. Waelbroeck. Discontinuous Galerkin methods for plasma physics in the scrape-off layer of tokamaks. Journal of Computational Physics, 274(??):898–919, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004768.

Muller:2014:SBV

[MMJ14] Florian Müller, Daniel W. Meyer, and Patrick Jenny. Solverbased vs. grid-based multilevel Monte Carlo for two phase flow and transport in random heterogeneous porous media. Journal of Computational Physics, 268(??):39–50, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001776.

Martinez:2011:BAS

[MMM11] E. Martínez, P. R. Monasterio, and J. Marian. Billionatom synchronous parallel kinetic Monte Carlo simulations of critical 3D Ising systems. Journal of Computational Physics, 230(4):1359–1369, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000608X.

Malandain:2013:ODC

[MMM13] Mathias Malandain, Nicolas Maheu, and Vincent Moureau. Optimization of the deflated Conjugate Gradient algorithm for the solving of elliptic equations on massively parallel machines. Journal of Computational Physics, 238(??):32–47, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007280.

Mazzia:2011:BBG

[MMP11] Annamaria Mazzia, Gianmarco Manzini, and Mario Putti. Bad behavior of Godunov mixed methods for strongly anisotropic advection-dispersion equations. Journal of Computational Physics, 230(23):8410–8426, September 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004499.

Maginot:2012:NNM

[MMR12] Peter G. Maginot, Jim E. Morel, and Jean C. Ragusa. A non-negative moment-preserving spatial discretization scheme for the linearized Boltzmann transport equation in 1-D and 2-D Cartesian geometries. Journal of Computational Physics, 231(20):6801–6826, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003336.

Marxen:2013:MDN

[MMSI13] Olaf Marxen, Thierry E. Magin, Eric S. G. Shaqfeh, and Gianluca Iaccarino. A method for the direct numerical simulation of hypersonic boundary-layer instability with finite-rate chemistry. Journal of Computational Physics, 255(??):572–589, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005111.

Montecinos:2014:HRV

[MMT14] Gino I. Montecinos, Lucas O. Müller, and Eleuterio F. Toro. Hyperbolic reformulation of a 1D viscoelastic blood flow model and ADER finite volume schemes. Journal of Computational Physics, 266(??):101–123, June 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001338.

Marras:2013:SMC

[MMV^{+}13a] Simone Marras, Margarida Moragues, Mariano Vázquez, Oriol Jorba, and Guillaume Houzeaux. Simulations of moist convection by a variational multiscale stabilized finite element method. Journal of Computational Physics, 252(??):195–218, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004294.

Marras:2013:VMS

[MMV^{+}13b] Simone Marras, Margarida Moragues, Mariano Vázquez, Oriol Jorba, and Guillaume Houzeaux. A variational multiscale stabilized finite element method for the solution of the Euler equations of nonhydrostatic stratified flows. Journal of Computational Physics, 236(??):380–407, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006626.

Mavriplis:2011:GCL

[MN11] Dimitri J. Mavriplis and Cristian R. Nastase. On the geometric conservation law for high-order discontinuous Galerkin discretizations on dynamically deforming meshes. Journal of Computational Physics, 230(11):4285–4300, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000477.

Mentrelli:2012:APS

[MN12] Andrea Mentrelli and Claudia Negulescu. Asymptoticpreserving scheme for highly anisotropic non-linear diffusion equations. Journal of Computational Physics, 231(24): 8229–8245, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911200438X.

Mosahebi:2014:IAN

[MN14] A. Mosahebi and S. Nadarajah. An implicit and adaptive nonlinear frequency domain approach for periodic viscous flows. Journal of Computational Physics, 278(??):92–116, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005762.

Men:2010:BOT

[MNF^{+}10] H. Men, N. C. Nguyen, R. M. Freund, P. A. Parrilo, and J. Peraire. Bandgap optimization of two-dimensional photonic crystals using semidefinite programming and subspace methods. Journal of Computational Physics, 229(10):3706–3725, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000046X.

McCormick:2013:SLV

[MNKS13] M. McCormick, D. A. Nordsletten, D. Kay, and N. P. Smith. Simulating left ventricular fluid-solid mechanics through the cardiac cycle under LVAD support. Journal of Computational Physics, 244(??):80–96, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004512.

Motamarri:2013:HOA

[MNL^{+}13] P. Motamarri, M. R. Nowak, K. Leiter, J. Knap, and V. Gavini. Higher-order adaptive finite-element methods for Kohn–Sham density functional theory. Journal of Computational Physics, 253(??):308–343, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004774.

Marques:2011:CFM

[MNR11] Alexandre Noll Marques, Jean-Christophe Nave, and Rodolfo Ruben Rosales. A Correction Function Method for Poisson problems with interface jump conditions. Journal of Computational Physics, 230(20):7567–7597, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003780.

Moore:2013:CCL

[MNS13] Brian E. Moore, Laura Noreña, and Constance M. Schober. Conformal conservation laws and geometric integration for damped Hamiltonian PDEs. Journal of Computational Physics, 232(1):214–233, January 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004561.

Monkola:2013:OBA

[Mön13] Sanna Mönkölä. An optimization-based approach for solving a time-harmonic multiphysical wave problem with higher-order schemes. Journal of Computational Physics, 242(??):439–459, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001411.

Morinishi:2010:SSF

[Mor10] Yohei Morinishi. Skew-symmetric form of convective terms and fully conservative finite difference schemes for variable density low-Mach number flows. Journal of Computational Physics, 229(2):276–300, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005130.

Meister:2012:CDG

[MOSW12] Andreas Meister, Sigrun Ortleb, Thomas Sonar, and Martina Wirz. A comparison of the Discontinuous-Galerkin-and Spectral-Difference-Method on triangulations using PKD polynomials. Journal of Computational Physics, 231(23): 7722–7729, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112004093.

McGill:2012:EGE

[MOV12] Jacob A. McGill, Babatunde A. Ogunnaike, and Dionisios G. Vlachos. Efficient gradient estimation using finite differencing and likelihood ratios for kinetic Monte Carlo simulations. Journal of Computational Physics, 231(21):7170–7186, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003695.

Miller:2013:NNP

[MP13] Gregory H. Miller and Elbridge Gerry Puckett. A Neumann– Neumann preconditioned iterative substructuring approach for computing solutions to Poisson’s equation with prescribed jumps on an embedded boundary. Journal of Computational Physics, 235(??):683–700, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006249.

Miki:2012:PMU

[MPPP12a] K. Miki, M. Panesi, E. E. Prudencio, and S. Prudhomme. Probabilistic models and uncertainty quantification for the ionization reaction rate of atomic nitrogen. Journal of Computational Physics, 231(9):3871–3886, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999112000265. See corrigendum [MPPP12b].

Miki:2012:CPM

[MPPP12b] Kenji Miki, Marco Panesi, Ernesto E. Prudencio, and Serge Prudhomme. Corrigendum to “Probabilistic models and uncertainty quantification for the ionization reaction rate of atomic Nitrogen” [JCOMP 231(9) (2012) 3871–3886]. Journal of Computational Physics, 231(15):5216, June 1, 2012.

CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002124. See [MPPP12a].

Marchandise:2012:CMR

[MPR12] E. Marchandise, C. Piret, and J.-F. Remacle. CAD and mesh repair with Radial Basis Functions. Journal of Computational Physics, 231(5):2376–2387, March 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006899.

Madankan:2014:CPH

[MPS^{+}14] R. Madankan, S. Pouget, P. Singla, M. Bursik, J. Dehn, M. Jones, A. Patra, M. Pavolonis, E. B. Pitman, T. Singh, and P. Webley. Computation of probabilistic hazard maps and source parameter estimation for volcanic ash transport and dispersion. Journal of Computational Physics, 271(??):39–59, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007948.

Muller:2013:WBH

[MPT13] Lucas O. Müller, Carlos Parés, and Eleuterio F. Toro. Wellbalanced high-order numerical schemes for one-dimensional blood flow in vessels with varying mechanical properties. Journal of Computational Physics, 242(??):53–85, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001277.

Menz:2013:ASW

[MPWK13] William J. Menz, Robert I. A. Patterson, Wolfgang Wagner, and Markus Kraft. Application of stochastic weighted algorithms to a multidimensional silica particle model. Journal of Computational Physics, 248(??):221–234, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002593.

McGurn:2013:ELM

[MRD13] M. T. McGurn, K. P. Ruggirello, and P. E. DesJardin. An Eulerian–Lagrangian moving immersed interface method for simulating burning solids. Journal of Computational Physics, 241(??):364–387, May 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113000983.

Madaule:2014:ECD

[MRS14a] Éric Madaule, Marco Restelli, and Eric Sonnendrücker. Energy conserving discontinuous Galerkin spectral element method for the Vlasov–Poisson system. Journal of Computational Physics, 279(??):261–288, December 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006469.

Mettot:2014:CES

[MRS14b] Clément Mettot, Florent Renac, and Denis Sipp. Computation of eigenvalue sensitivity to base flow modifications in a discrete framework: Application to open-loop control. Journal of Computational Physics, 269(??):234–258, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002034.

Mack:2010:PKT

[MS10a] Christoph J. Mack and Peter J. Schmid. A preconditioned Krylov technique for global hydrodynamic stability analysis of large-scale compressible flows. Journal of Computational Physics, 229(3):541–560, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005117.

Majid:2010:ASG

[MS10b] Abdul Majid and Sultan Sial. Application of Sobolev gradient method to Poisson–Boltzmann system. Journal of Computational Physics, 229(16):5742–5754, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001932.

Majid:2011:ASP

[MS11] Abdul Majid and Sultan Sial. Approximate solutions to Poisson–Boltzmann systems with Sobolev gradients. Journal of Computational Physics, 230(14):5732–5738, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100218X.

Maroulas:2012:IPF

[MS12] Vasileios Maroulas and Panos Stinis. Improved particle filters for multi-target tracking. Journal of Computational Physics, 231(2):602–611, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005663.

Malaspinas:2014:WML

[MS14a] O. Malaspinas and P. Sagaut. Wall model for large-eddy simulation based on the lattice Boltzmann method. Journal of Computational Physics, 275(??):25–40, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004276.

Meldi:2014:ANM

[MS14b] M. Meldi and P. Sagaut. An adaptive numerical method for solving EDQNM equations for the analysis of longtime decay of isotropic turbulence. Journal of Computational Physics, 262(??):72–85, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000217.

Montefuscolo:2014:HOA

[MSB14] Felipe Montefuscolo, Fabricio S. Sousa, and Gustavo C. Buscaglia. High-order ALE schemes for incompressible capillary flows. Journal of Computational Physics, 278(??):133–147, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005841.

Messner:2012:FDM

[MSD12] Matthias Messner, Martin Schanz, and Eric Darve. Fast directional multilevel summation for oscillatory kernels based on Chebyshev interpolation. Journal of Computational Physics, 231(4):1175–1196, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005705.

Mishra:2012:MLM

[MSS12] S. Mishra, Ch. Schwab, and J. Sukys. Multi-level Monte Carlo finite volume methods for nonlinear systems of conservation laws in multi-dimensions. Journal of Computational Physics, 231(8):3365–3388, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000320.

Mauro:2014:FPK

[MSS^{+}14] Ava J. Mauro, Jon Karl Sigurdsson, Justin Shrake, Paul J. Atzberger, and Samuel A. Isaacson. A First-Passage Kinetic Monte Carlo method for reaction-drift-diffusion processes. Journal of Computational Physics, 259(??):536–567, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300822X.

Messner:2014:FGM

[MST14] Michael Messner, Martin Schanz, and Johannes Tausch. A fast Galerkin method for parabolic space–time boundary integral equations. Journal of Computational Physics, 258(??):15–30, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007055.

Mohamed:2013:TIH

[MSTL13] M. Shadi Mohamed, Mohammed Seaid, Jon Trevelyan, and Omar Laghrouche. Time-independent hybrid enrichment for finite element solution of transient conductionradiation in diffusive grey media. Journal of Computational Physics, 251(??):81–101, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003963.

Mohamed:2014:EFE

[MSTL14] M. Shadi Mohamed, Mohammed Seaid, Jon Trevelyan, and Omar Laghrouche. An enriched finite element model with q-refinement for radiative boundary layers in glass cooling. Journal of Computational Physics, 258(??):718–737, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007511.

Mignone:2010:SOU

[MT10] Andrea Mignone and Petros Tzeferacos. A second-order unsplit Godunov scheme for cell-centered MHD: The CTU–GLM scheme. Journal of Computational Physics, 229(6):2117–2138, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910900655X.

McDonald:2013:ARM

[MT13] James McDonald and Manuel Torrilhon. Affordable robust moment closures for CFD based on the maximum-entropy hierarchy. Journal of Computational Physics, 251(??):500–523, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004130.

Montecinos:2014:RGA

[MT14a] Gino I. Montecinos and Eleuterio F. Toro. Reformulations for general advection-diffusion-reaction equations and locally implicit ADER schemes. Journal of Computational Physics, 275(??):415–442, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004252.

Muradoglu:2014:SSS

[MT14b] Metin Muradoglu and Gretar Tryggvason. Simulations of soluble surfactants in 3D multiphase flow. Journal of Computational Physics, 274(??):737–757, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004318.

Morzfeld:2012:RMI

[MTAC12] Matthias Morzfeld, Xuemin Tu, Ethan Atkins, and Alexandre J. Chorin. A random map implementation of implicit filters. Journal of Computational Physics, 231(4):2049–2066, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006772.

Mignone:2010:HOC

[MTB10] Andrea Mignone, Petros Tzeferacos, and Gianluigi Bodo. Highorder conservative finite difference GLM–MHD schemes for cell-centered MHD. Journal of Computational Physics, 229 (17):5896–5920, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001890.

Moon:2014:SPC

[MTD14] Haksu Moon, Fernando L. Teixeira, and Burkay Donderici. Stable pseudoanalytical computation of electromagnetic fields from arbitrarily-oriented dipoles in cylindrically stratified media. Journal of Computational Physics, 273(??):118–142, September 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003428.

Mirzadeh:2011:SOD

[MTG11] Mohammad Mirzadeh, Maxime Theillard, and Frédéric Gibou. A second-order discretization of the nonlinear Poisson– Boltzmann equation over irregular geometries using nongraded adaptive Cartesian grids. Journal of Computational Physics, 230(5):2125–2140, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006728.

Mohan:2011:VOS

[MTS^{+}11] P. Surya Mohan, Tanja Tarvainen, Martin Schweiger, Aki Pulkkinen, and Simon R. Arridge. Variable order spherical harmonic expansion scheme for the radiative transport equation using finite elements. Journal of Computational Physics, 230(19):7364–7383, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003536.

Meng:2012:CSM

[MTSG12] Xing Meng, Gábor Tóth, Igor V. Sokolov, and Tamas I. Gombosi. Classical and semirelativistic magnetohydrodynamics with anisotropic ion pressure. Journal of Computational Physics, 231(9):3610–3622, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000113.

Medvinsky:2013:HON

[MTT13] M. Medvinsky, S. Tsynkov, and E. Turkel. High order numerical simulation of the transmission and scattering of waves using the method of difference potentials. Journal of Computational Physics, 243(??):305–322, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001927.

Ming:2014:ESM

[MTZ14] Ju Ming, Qinglin Tang, and Yanzhi Zhang. An efficient spectral method for computing dynamics of rotating twocomponent Bose–Einstein condensates via coordinate transformation. Journal of Computational Physics, 258(??):538–554, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007201.

Murman:2010:CUS

[Mur10] Scott M. Murman. Compact upwind schemes on adaptive octrees. Journal of Computational Physics, 229(4):1167–1180, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005713.

Mulet:2013:SLA

[MV13] Pep Mulet and Francesco Vecil. A semi-Lagrangian AMR scheme for 2D transport problems in conservation form. Journal of Computational Physics, 237(??):151–176, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200719X.

Moghadam:2013:MNM

[MVCFM13] Mahdi Esmaily Moghadam, Irene E. Vignon-Clementel, Richard Figliola, and Alison L. Marsden. A modular numerical method for implicit 0D/3D coupling in cardiovascular finite element simulations. Journal of Computational Physics, 244(??):63–79, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004202.

Morris:2011:MCS

[MVG11] A. B. Morris, P. L. Varghese, and D. B. Goldstein. Monte Carlo solution of the Boltzmann equation via a discrete velocity model. Journal of Computational Physics, 230(4):1265–1280, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006030.

Miron:2012:AMA

[MVG^{+}12] Philippe Miron, Jérôme Vétel, André Garon, Michel Delfour, and Mouhammad El Hassan. Anisotropic mesh adaptation on Lagrangian Coherent Structures. Journal of Computational Physics, 231(19):6419–6437, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003221.

Meldi:2013:ALE

[MVS13] M. Meldi, E. Vergnault, and P. Sagaut. An arbitrary Lagrangian–Eulerian approach for the simulation of immersed moving solids with Lattice Boltzmann Method. Journal of Computational Physics, 235(??):182–198, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006031.

Majhool:2012:SAI

[MW12] Ahmed Abed Al-Kadhem Majhool and A. P. Watkins. Spray algorithm without interface construction. Journal of Computational Physics, 231(9):3647–3662, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000332.

Montazeri:2014:BFA

[MW14a] Hanif Montazeri and C. A. Ward. A balanced-force algorithm for two-phase flows. Journal of Computational Physics, 257(??):645–669, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006736.

More:2014:DYT

[MW14b] Jorge J. Moré and Stefan M. Wild. Do you trust derivatives or differences? Journal of Computational Physics, 273(??):268–277, September 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003325.

Ma:2014:GCC

[MWP14] Z. H. Ma, H. Wang, and S. H. Pu. GPU computing of compressible flow problems by a meshless method with space-filling curves. Journal of Computational Physics, 263(??):113–135, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000503.

Meng:2013:AES

[MWRZ13] Jianping Meng, Lei Wu, Jason M. Reese, and Yonghao Zhang. Assessment of the ellipsoidal-statistical bhatnagargross-krook model for force-driven Poiseuille flows. Journal of Computational Physics, 251(??):383–395, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004129.

Mu:2013:WGM

[MWW^{+}13] Lin Mu, Junping Wang, Guowei Wei, Xiu Ye, and Shan Zhao. Weak Galerkin methods for second order elliptic interface problems. Journal of Computational Physics, 250(??):106–125, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003367.

Mu:2014:SNA

[MWY14] Lin Mu, Junping Wang, and Xiu Ye. A stable numerical algorithm for the Brinkman equations by weak Galerkin finite element methods. Journal of Computational Physics, 273(??):327–342, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002800.

Moroney:2013:EST

[MY13] Timothy Moroney and Qianqian Yang. Efficient solution of two-sided nonlinear space-fractional diffusion equations using fast Poisson preconditioners. Journal of Computational Physics, 246(??):304–317, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002076.

Mor-Yossef:2014:UST

[MY14]  Y. Mor-Yossef.  Unconditionally stable time marching

scheme for Reynolds stress models. Journal of Computational Physics, 276(??):635–664, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005403.

Miyatake:2012:NIO

[MYM12] Yuto Miyatake, Takaharu Yaguchi, and Takayasu Matsuo. Numerical integration of the Ostrovsky equation based on its geometric structures. Journal of Computational Physics, 231(14):4542–4559, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001386.

Ma:2010:AHD

[MZ10] Xiang Ma and Nicholas Zabaras. An adaptive high-dimensional stochastic model representation technique for the solution of stochastic partial differential equations. Journal of Computational Physics, 229(10):3884–3915, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000562.

Ma:2011:KPC

[MZ11a] Xiang Ma and Nicholas Zabaras. Kernel principal component analysis for stochastic input model generation. Journal of Computational Physics, 230(19):7311–7331, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003494.

Ma:2011:SMF

[MZ11b] Xiang Ma and Nicholas Zabaras. A stochastic mixed finite element heterogeneous multiscale method for flow in porous media. Journal of Computational Physics, 230(12):4696–4722, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001318.

Mencinger:2011:PVM

[MZ11c] Jure Mencinger and Iztok Zun. A PLIC–VOF method suited for adaptive moving grids. Journal of Computational Physics, 230(3):644–663, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005553.

Meng:2011:AAH

[MZ11d] Jianping Meng and Yonghao Zhang. Accuracy analysis of high-order lattice Boltzmann models for rarefied gas flows. Journal of Computational Physics, 230(3):835–849, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005796.

Meng:2014:DRB

[MZ14] Jianping Meng and Yonghao Zhang. Diffuse reflection boundary condition for high-order lattice Boltzmann models with streaming-collision mechanism. Journal of Computational Physics, 258(??):601–612, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300733X.

Momox:2012:SSE

[MZB12] E. Momox, N. Zakhleniuk, and N. Balkan. Solution of the 1D Schrödinger equation in semiconductor heterostructures using the immersed interface method. Journal of Computational Physics, 231(18):6173–6180, July 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002598.

Nagaosa:2014:NMG

[Nag14a] Ryuichi S. Nagaosa. A numerical modelling of gas exchange mechanisms between air and turbulent water with an aquarium chemical reaction. Journal of Computational Physics, 256(??):69–87, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005652.

Nagaosa:2014:RNM

[Nag14b] Ryuichi S. Nagaosa. Reprint of: a numerical modelling of gas exchange mechanisms between air and turbulent water with an aquarium chemical reaction. Journal of Computational Physics, 271(??):172–190, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002617.

Norman:2013:VNM

[NB13] Ryan B. Norman and Steve R. Blattnig. Validation of nuclear models used in space radiation shielding applications. Journal of Computational Physics, 233(??):464–479, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005359.

Noormohammadi:2014:FDM

[NB14] N. Noormohammadi and B. Boroomand. A fictitious domain method using equilibrated basis functions for harmonic and bi-harmonic problems in physics. Journal of Computational Physics, 272(??):189–217, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002745.

Nitsche:2010:HOQ

[NCKN10] M. Nitsche, H. D. Ceniceros, A. L. Karniala, and S. Naderi. High order quadratures for the evaluation of interfacial velocities in axi-symmetric Stokes flows. Journal of Computational Physics, 229(18):6318–6342, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002305.

Novak:2010:SMW

[NCV10] J. Novak, J.-L. Cornou, and N. Vasset. A spectral method for the wave equation of divergence-free vectors and symmetric tensors inside a sphere. Journal of Computational Physics, 229(2):399–414, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005257.

Niegemann:2012:ELS

[NDB12] Jens Niegemann, Richard Diehl, and Kurt Busch. Efficient low-storage Runge–Kutta schemes with optimized stability regions. Journal of Computational Physics, 231(2):364–372, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005213.

Nigro:2014:SOA

[NDBG14] Alessandra Nigro, Carmine De Bartolo, Francesco Bassi, and Antonio Ghidoni. Up to sixth-order accurate A-stable implicit schemes applied to the Discontinuous Galerkin discretized Navier–Stokes equations. Journal of Computational Physics, 276(??):136–162, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400521X.

Nickovic:2011:MEP

[NDV^{+}11] Slobodan Nickovic, Vladimir Djurdjevic, Mirjam Vujadinovic, Zavisa I. Janjic, Milan Curcic, and Borivoj Rajkovic. Method for efficient prevention of gravity wave decoupling on rectangular semi-staggered grids. Journal of Computational Physics, 230(5):1865–1875, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006522.

Nordstrom:2012:WSW

[NEE12] Jan Nordström, Sofia Eriksson, and Peter Eliasson. Weak and strong wall boundary procedures and convergence to steadystate of the Navier–Stokes equations. Journal of Computational Physics, 231(14):4867–4884, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001787.

Norman:2012:MMA

[NF12] M. R. Norman and H. Finkel. Multi-moment ADER-taylor methods for systems of conservation laws with source terms in one dimension. Journal of Computational Physics, 231 (20):6622–6642, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112002859.

Nasr:2014:LDA

[NGV14] N. Ben Nasr, G. A. Gerolymos, and I. Vallet. Low-diffusion approximate Riemann solvers for Reynolds-stress transport. Journal of Computational Physics, 268(??):186–235, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001302.

Najafi:2014:ASF

[NHE14] Mehdi Najafi, Kazem Hejranfar, and Vahid Esfahanian. Application of a shock-fitted spectral collocation method for computing transient high-speed inviscid flows over a blunt nose. Journal of Computational Physics, 257(??):954–980, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006517.

Numata:2010:AAG

[NHT^{+}10] Ryusuke Numata, Gregory G. Howes, Tomoya Tatsuno, Michael Barnes, and William Dorland. AstroGK: Astrophysical gyrokinetics code. Journal of Computational Physics, 229(24):9347–9372, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0021999110005000. See corrigendum [NHT^{+}13].

Numata:2013:CAA

[NHT^{+}13] Ryusuke Numata, Gregory G. Howes, Tomoya Tatsuno, Michael Barnes, and William Dorland. Corrigendum to “AstroGK: Astrophysical gyrokinetics code” [J. Comput. Phys. 229 (2010) 9347-9372]. Journal of Computational Physics, 245(??):493–494, July 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002490. See [NHT^{+}10].

Nicholas:2010:HAA

[Nic10] Mike Nicholas. A high accuracy algorithm for 3D periodic electromagnetic scattering. Journal of Computational Physics, 229(21):8211–8223, October 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004213.

Nicholls:2011:EEF

[Nic11] David P. Nicholls. Efficient enforcement of far-field boundary conditions in the Transformed Field Expansions method. Journal of Computational Physics, 230(22):8290–8303, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004578.

Noda:2014:MPE

[NIN^{+}14] Masashi Noda, Kazuya Ishimura, Katsuyuki Nobusada, Kazuhiro Yabana, and Taisuke Boku. Massively-parallel electron dynamics calculations in real-time and real-space: Toward applications to nanostructures of more than ten-nanometers in size. Journal of Computational Physics, 265(??):145–155, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001004.

Nishikawa:2010:FOS

[Nis10] Hiroaki Nishikawa. A first-order system approach for diffusion equation. II: Unification of advection and diffusion. Journal of Computational Physics, 229(11):3989–4016, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006019.

Nishikawa:2012:DFS

[Nis12] Hiroaki Nishikawa. Divergence formulation of source term. Journal of Computational Physics, 231(19):6393–6400, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003014.

Nishikawa:2014:FSTa

[Nis14a] Hiroaki Nishikawa. First-, second-, and third-order finitevolume schemes for diffusion. Journal of Computational Physics, 256(??):791–805, January 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006384.

Nishikawa:2014:FSTb

[Nis14b] Hiroaki Nishikawa. First, second, and third order finitevolume schemes for advection-diffusion. Journal of Computational Physics, 273(??):287–309, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003672.

Noetinger:2012:QSS

[NJ12] B. Noetinger and N. Jarrige. A quasi steady state method for solving transient Darcy flow in complex 3D fractured networks. Journal of Computational Physics, 231(1):23–38, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100502X.

Nonomura:2014:SIS

[NKF14] Taku Nonomura, Keiichi Kitamura, and Kozo Fujii. A simple interface sharpening technique with a hyperbolic tangent function applied to compressible two-fluid modeling. Journal of Computational Physics, 258(??):95–117, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006979.

Nordsletten:2010:NCM

[NKS10] D. Nordsletten, D. Kay, and N. Smith. A non-conforming monolithic finite element method for problems of coupled mechanics. Journal of Computational Physics, 229(20):7571–7593, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003189.

Nandipati:2012:LPR

[NKSR12] Giridhar Nandipati, Abdelkader Kara, Syed Islamuddin Shah, and Talat S. Rahman. Off-lattice pattern recognition scheme for kinetic Monte Carlo simulations. Journal of Computational Physics, 231(9):3548–3560, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007509.

Nair:2010:CDF

[NL10] Ramachandran D. Nair and Peter H. Lauritzen. A class of deformational flow test cases for linear transport problems on the sphere. Journal of Computational Physics, 229(23):8868–8887, November 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004511.

Nakata:2012:FSI

[NL12a] Toshiyuki Nakata and Hao Liu. A fluid-structure interaction model of insect flight with flexible wings. Journal of Computational Physics, 231(4):1822–1847, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006474.

Ni:2012:CCS

[NL12b] Ming-Jiu Ni and Jun-Feng Li. A consistent and conservative scheme for incompressible MHD flows at a low magnetic Reynolds number. Part III: On a staggered mesh. Journal of Computational Physics, 231(2):281–298, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004852.

Nordstrom:2013:SPT

[NL13] Jan Nordström and Tomas Lundquist. Summation-by-parts in time. Journal of Computational Physics, 251(??):487–499, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004099.

Nguyen:2014:FEM

[NLGL14] Dang Van Nguyen, Jing-Rebecca Li, Denis Grebenkov, and Denis Le Bihan. A finite elements method to solve the Bloch– Torrey equation applied to diffusion magnetic resonance imaging. Journal of Computational Physics, 263(??):283–302, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000308.

Nazari:2014:OHO

[NMCZ14] Farshid Nazari, Abdolmajid Mohammadian, Martin Charron, and Ayrton Zadra. Optimal high-order diagonallyimplicit Runge–Kutta schemes for nonlinear diffusive systems on atmospheric boundary layer. Journal of Computational Physics, 271(??):118–130, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000771.

Nonomura:2012:NEI

[NMT^{+}12] Taku Nonomura, Seiichiro Morizawa, Hiroshi Terashima, Shigeru Obayashi, and Kozo Fujii. Numerical (error) issues on compressible multicomponent flows using a high-order differencing scheme: Weighted compact nonlinear scheme. Journal of Computational Physics, 231(8):3181–3210, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000022.

Nakshatrala:2013:NFD

[NMV13] K. B. Nakshatrala, M. K. Mudunuru, and A. J. Valocchi. A numerical framework for diffusion-controlled bimolecular-reactive systems to enforce maximum principles and the non-negative constraint. Journal of Computational Physics, 253(??):278–307, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300483X.

Narayan:2012:SDA

[NMX12] Akil Narayan, Youssef Marzouk, and Dongbin Xiu. Sequential data assimilation with multiple models. Journal of Computational Physics, 231(19):6401–6418, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003038.

Nazari:2013:SAS

[NMZC13] Farshid Nazari, Abdolmajid Mohammadian, Ayrton Zadra, and Martin Charron. A stable and accurate scheme for nonlinear diffusion equations: Application to atmospheric boundary layer. Journal of Computational Physics, 236(??):271–288, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006420.

Niino:2012:PBC

[NN12] Kazuki Niino and Naoshi Nishimura. Preconditioning based on Calderon’s formulae for periodic fast multipole methods for Helmholtz’ equation. Journal of Computational Physics, 231(1):66–81, January 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111005067.

Norman:2011:LCL

[NNS11] Matthew R. Norman, Ramachandran D. Nair, and Fredrick H. M. Semazzi. A low communication and large time step explicit finite-volume solver for non-hydrostatic atmospheric dynamics. Journal of Computational Physics, 230(4):1567–1584, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006376.

Norman:2013:AIS

[Nor13] M. R. Norman. Algorithmic improvements for schemes using the ADER time discretization. Journal of Computational Physics, 243(??):176–178, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001757.

Norman:2014:WLA

[Nor14] Matthew R. Norman. A WENO-limited, ADER-DT, finitevolume scheme for efficient, robust, and communicationavoiding multi-dimensional transport. Journal of Computational Physics, 274(??):1–18, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004094.

Nickaeen:2014:NML

[NOT14] M. Nickaeen, A. Ouazzi, and S. Turek. Newton multigrid leastsquares FEM for the V-V-P formulation of the Navier–Stokes equations. Journal of Computational Physics, 256(??):416–427, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006165.

Nguyen:2012:HDG

[NP12] N. C. Nguyen and J. Peraire. Hybridizable discontinuous Galerkin methods for partial differential equations in continuum mechanics. Journal of Computational Physics, 231 (18):5955–5988, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112001544.

Nguyen:2011:HOI

[NPC11a] N. C. Nguyen, J. Peraire, and B. Cockburn. Highorder implicit hybridizable discontinuous Galerkin methods for acoustics and elastodynamics. Journal of Computational Physics, 230(10):3695–3718, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100060X.

Nguyen:2011:HDG

[NPC11b] N. C. Nguyen, J. Peraire, and B. Cockburn. Hybridizable discontinuous Galerkin methods for the time-harmonic Maxwell’s equations. Journal of Computational Physics, 230 (19):7151–7175, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111003226.

Nguyen:2011:IHO

[NPC11c] N. C. Nguyen, J. Peraire, and B. Cockburn. An implicit highorder hybridizable discontinuous Galerkin method for the incompressible Navier–Stokes equations. Journal of Computational Physics, 230(4):1147–1170, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005887.

Nobile:2014:IAP

[NPV14] Fabio Nobile, Matteo Pozzoli, and Christian Vergara. Inexact accurate partitioned algorithms for fluid-structure interaction problems with finite elasticity in haemodynamics. Journal of Computational Physics, 273(??):598–617, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003660.

Nave:2010:GAL

[NRS10] Jean-Christophe Nave, Rodolfo Ruben Rosales, and Benjamin Seibold. A gradient-augmented level set method with an optimally local, coherent advection scheme. Journal of Computational Physics, 229(10):3802–3827, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000525.

Natarajan:2011:INV

[NS11a] Ganesh Natarajan and Fotis Sotiropoulos. IDeC(k): a new velocity reconstruction algorithm on arbitrarily polygonal staggered meshes. Journal of Computational Physics, 230 (17):6583–6604, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911100297X.

Nishiura:2011:PVA

[NS11b] Daisuke Nishiura and Hide Sakaguchi. Parallel-vector algorithms for particle simulations on shared-memory multiprocessors. Journal of Computational Physics, 230(5):1923–1938, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006558.

Niemeyer:2014:AMS

[NS14a] Kyle E. Niemeyer and Chih-Jen Sung. Accelerating moderately stiff chemical kinetics in reactive-flow simulations using GPUs. Journal of Computational Physics, 256(??):854–871, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006396.

Novak:2014:CAP

[NS14b] Igor L. Novak and Boris M. Slepchenko. A conservative algorithm for parabolic problems in domains with moving boundaries. Journal of Computational Physics, 270(??):203–213, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001879.

Nabovati:2011:LBM

[NSA11] Aydin Nabovati, Daniel P. Sellan, and Cristina H. Amon. On the lattice Boltzmann method for phonon transport. Journal of Computational Physics, 230(15):5864–5876, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002233.

Norris:2014:ALR

[NT14] Scott A. Norris and Skyler Tweedie. Adaptive Localized Replay: an efficient integration scheme for accurate simulation of coarsening dynamical systems. Journal of Computational Physics, 273(??):1–11, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003350.

Nagendra:2014:NAC

[NTV14] Krishnamurthy Nagendra, Danesh K. Tafti, and Kamal Viswanath. A new approach for conjugate heat transfer problems using immersed boundary method for curvilinear grid based solvers. Journal of Computational Physics, 267(??):225– 246, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001739.

Najm:2014:EPI

[NV14] Habib N. Najm and Mauro Valorani. Enforcing positivity in intrusive PC–UQ methods for reactive ODE systems. Journal of Computational Physics, 270(??):544–569, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002484.

Nurijanyan:2013:HDG

[NvdVB13] S. Nurijanyan, J. J. W. van der Vegt, and O. Bokhove. Hamiltonian discontinuous Galerkin FEM for linear, rotating incompressible Euler equations: Inertial waves. Journal of Computational Physics, 241(??):502–525, May 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300048X.

Norman:2013:GAC

[NVS13] Paul Norman, Paolo Valentini, and Thomas Schwartzentruber. GPU-accelerated classical trajectory calculation direct simulation Monte Carlo applied to shock waves. Journal of Computational Physics, 247(??):153–167, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002441.

Nochetto:2010:HVF

[NW10] Ricardo H. Nochetto and Shawn W. Walker. A hybrid variational front tracking-level set mesh generator for problems exhibiting large deformations and topological changes. Journal of Computational Physics, 229(18):6243–6269, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002226.

Norris:2012:SKC

[NW12] Scott A. Norris and Stephen J. Watson. Simulating the kinematics of completely faceted surfaces. Journal of Computational Physics, 231(14):4560–4577, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001441.

Nedelcu:2012:GIB

[NWLS12] S. Nedelcu, M. Werner, M. Lang, and J.-U. Sommer. GPU implementations of the bond fluctuation model. Journal of Computational Physics, 231(7):2811–2824, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007339.

Najafi-Yazdi:2013:LDL

[NYM13] A. Najafi-Yazdi and L. Mongeau. A low-dispersion and lowdissipation implicit Runge–Kutta scheme. Journal of Computational Physics, 233(??):315–323, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005189.

Nguyen:2014:TDM

[NZ14] Duc Duy Nguyen and Shan Zhao. Time-domain matched interface and boundary (MIB) modeling of Debye dispersive media with curved interfaces. Journal of Computational Physics, 278(??):298–325, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006068.

Zhang:2011:ADI

[nZzS11] Ya nan Zhang and Zhi zhong Sun. Alternating direction implicit schemes for the two-dimensional fractional sub-diffusion equation. Journal of Computational Physics, 230(24):8713– 8728, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111005079.

Zhang:2014:FDM

[nZzSlL14] Ya nan Zhang, Zhi zhong Sun, and Hong lin Liao. Finite difference methods for the time fractional diffusion equation on non-uniform meshes. Journal of Computational Physics, 265(??):195–210, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001041.

Ohwada:2010:ACM

[OA10] Taku Ohwada and Pietro Asinari. Artificial compressibility method revisited: Asymptotic numerical method for incompressible Navier–Stokes equations. Journal of Computational Physics, 229(5):1698–1723, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006147.

Onodera:2011:LES

[OAK11] Naoyuki Onodera, Takayuki Aoki, and Hiromichi Kobayashi. Large-eddy simulation of turbulent channel flows with conservative IDO scheme. Journal of Computational Physics, 230 (14):5787–5805, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111002282.

Obrecht:2014:HPI

[OAKR14] Christian Obrecht, Pietro Asinari, Frédéric Kuznik, and Jean-Jacques Roux. High-performance implementations and large-scale validation of the link-wise artificial compressibility method. Journal of Computational Physics, 275(??):143–153, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004823.

Ohwada:2013:RAS

[OAXL13] Taku Ohwada, Ryo Adachi, Kun Xu, and Jun Luo. On the remedy against shock anomalies in kinetic schemes. Journal of Computational Physics, 255(??):106–129, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005329.

Osterloh:2013:RIM

[OBNN13] Lukas Osterloh, Christine Böckmann, Doina Nicolae, and Anca Nemuc. Regularized inversion of microphysical atmospheric particle parameters: Theory and application. Journal of Computational Physics, 237(??):79–94, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007206. See corrigendum

[OBNN14].

Osterloh:2014:CRI

[OBNN14] Lukas Osterloh, Christine Böckmann, Doina Nicolae, and Anca Nemuc. Corrigendum to “Regularized inversion of microphysical atmospheric particle parameters: Theory and application” [j. comput. phys. 237 (2013) 79–94]. Journal of Computational Physics, 275(??):696, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005348. See [OBNN13].

Onishi:2011:LSF

[OBT11] Ryo Onishi, Yuya Baba, and Keiko Takahashi. Large-scale forcing with less communication in finite-difference simulations of stationary isotropic turbulence. Journal of Computational Physics, 230(10):4088–4099, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001239.

Orszaghova:2012:PBB

[OBT12] Jana Orszaghova, Alistair G. L. Borthwick, and Paul H. Taylor. From the paddle to the beach — a Boussinesq shallow water numerical wave tank based on Madsen and Sørensen’s equations. Journal of Computational Physics, 231(2):328–344, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100516X.

Ober-Blobaum:2013:VIE

[OBTC^{+}13] Sina Ober-Blöbaum, Molei Tao, Mulin Cheng, Houman Owhadi, and Jerrold E. Marsden. Variational integrators for electric circuits. Journal of Computational Physics, 242 (??):498–530, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113001162.

Ogbonna:2012:DOG

[OD12] Nneoma Ogbonna and Dugald B. Duncan. Decoupled overlapping grids for the numerical modeling of oil wells. Journal of Computational Physics, 231(1):135–151, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005201.

Owkes:2013:DGC

[OD13] Mark Owkes and Olivier Desjardins. A discontinuous Galerkin conservative level set scheme for interface capturing in multiphase flows. Journal of Computational Physics, 249(??):275–302, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003215.

Owkes:2014:CFC

[OD14] Mark Owkes and Olivier Desjardins. A computational framework for conservative, three-dimensional, unsplit, geometric transport with application to the volume-of-fluid (VOF) method. Journal of Computational Physics, 270(??):587–612, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400285X.

Olbrant:2012:RPD

[OHF12] Edgar Olbrant, Cory D. Hauck, and Martin Frank. A realizability-preserving discontinuous Galerkin method for the M1 model of radiative transfer. Journal of Computational Physics, 231(17):5612–5639, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001362.

Oskooi:2011:DCI

[OJ11] Ardavan Oskooi and Steven G. Johnson. Distinguishing correct from incorrect PML proposals and a corrected unsplit PML for anisotropic, dispersive media. Journal of Computational Physics, 230(7):2369–2377, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000143.

Oh:2010:GPP

[OJW10] Hae-Soo Oh, Woo Jeong Jae, and Tak Hong Won. The generalized product partition of unity for the meshless methods. Journal of Computational Physics, 229(5):1600–1620, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006081.

Ozgun:2010:DCA

[OK10] Ozlem Ozgun and Mustafa Kuzuoglu. Domain compression via anisotropic metamaterials designed by coordinate transformations. Journal of Computational Physics, 229(3):921–932, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005750.

Omelchenko:2012:HUA

[OK12] Y. A. Omelchenko and H. Karimabadi. HYPERS: a unidimensional asynchronous framework for multiscale hybrid simulations. Journal of Computational Physics, 231(4):1766–1780, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006462.

Ozgun:2013:SMT

[OK13] Ozlem Ozgun and Mustafa Kuzuoglu. Software metamaterials: Transformation media based multiscale techniques for computational electromagnetics. Journal of Computational Physics, 236(??):203–219, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006924.

Oevermann:2014:ASA

[OK14] M. Oevermann and R. Klein. An asymptotic solution approach for elliptic equations with discontinuous coefficients. Journal of Computational Physics, 261(??):230–243, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008346.

Olson:2013:DAF

[OL13] Britton J. Olson and Sanjiva K. Lele. Directional artificial fluid properties for compressible large-eddy simulation. Journal of Computational Physics, 246(??):207–220, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002040.

Olson:2013:MDE

[OLC13] Sarah D. Olson, Sookkyung Lim, and Ricardo Cortez. Modeling the dynamics of an elastic rod with intrinsic curvature and twist using a regularized Stokes formulation. Journal of Computational Physics, 238(??):169–187, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007577.

Olbrant:2013:ADN

[OLFS13] E. Olbrant, E. W. Larsen, M. Frank, and B. Seibold. Asymptotic derivation and numerical investigation of timedependent simplified P_{N} equations. Journal of Computational Physics, 238(??):315–336, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006614.

Oliver:2014:MCS

[Oli14] Dean S. Oliver. Minimization for conditional simulation: Relationship to optimal transport. Journal of Computational Physics, 265(??):1–15, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400093X.

Ortega:2014:NSE

[OLPM14] A. López Ortega, M. Lombardini, D. I. Pullin, and D. I. Meiron. Numerical simulation of elastic-plastic solid mechanics using an Eulerian stretch tensor approach and HLLD Riemann solver. Journal of Computational Physics, 257(??):414–441, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300675X.

Olson:2011:SOT

[Ols11] Gordon L. Olson. Second order time evolution of the multigroup diffusion and P_{1} equations for radiation transport. Journal of Computational Physics, 230(20):7548–7566, August 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003470.

Olson:2012:ACR

[Ols12] Gordon L. Olson. Alternate closures for radiation transport using Legendre polynomials in 1D and spherical harmonics in 2D. Journal of Computational Physics, 231(7):2786–2793, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100725X.

Oxtoby:2012:MFI

[OM12] O. F. Oxtoby and A. G. Malan. A matrix-free, implicit, incompressible fractional-step algorithm for fluidstructure interaction applications. Journal of Computational Physics, 231(16):5389–5405, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002197.

Oterkus:2014:PTD

[OMA14] Selda Oterkus, Erdogan Madenci, and Abigail Agwai. Peridynamic thermal diffusion. Journal of Computational Physics, 265(??):71–96, May 15, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000540.

OKane:2014:ERL

[OMCO14] Terence J. O’Kane, Richard J. Matear, Matthew A. Chamberlain, and Peter R. Oke. ENSO regimes and the late 1970’s climate shift: the role of synoptic weather and South Pacific ocean spiciness. Journal of Computational Physics, 271(??):19–38, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007341.

Ovaysi:2010:DPL

[OP10] Saeed Ovaysi and Mohammad Piri. Direct pore-level modeling of incompressible fluid flow in porous media. Journal of Computational Physics, 229(19):7456–7476, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003438.

Oppelstrup:2013:MCC

[Opp13] Tomas Oppelstrup. Matrix compression by common subexpression elimination. Journal of Computational Physics, 247 (??):100–108, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113002209.

Olshanskii:2010:VVH

[OR10] Maxim A. Olshanskii and Leo G. Rebholz. Velocity-vorticity-helicity formulation and a solver for the Navier–Stokes equations. Journal of Computational Physics, 229(11):4291–4303, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000859.

Oliver:2013:SFE

[ORHH13] Robin C. Oliver, Daniel J. Read, Oliver G. Harlen, and Sarah A. Harris. A stochastic finite element model for the dynamics of globular macromolecules. Journal of Computational Physics, 239(??):147–165, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007589.

Olson:2011:SAM

[OS11a] Luke N. Olson and Jacob B. Schroder. Smoothed aggregation multigrid solvers for high-order discontinuous Galerkin methods for elliptic problems. Journal of Computational Physics, 230(18):6959–6976, August 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003135.

Ortega:2011:GCS

[OS11b] A. López Ortega and G. Scovazzi. A geometricallyconservative, synchronized, flux-corrected remap for arbitrary Lagrangian–Eulerian computations with nodal finite elements. Journal of Computational Physics, 230(17):6709–6741, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003020.

Ozcelikkale:2012:LSS

[OS12] Altug Ozcelikkale and Cuneyt Sert. Least-squares spectral element solution of incompressible Navier–Stokes equations with adaptive refinement. Journal of Computational Physics, 231(9):3755–3769, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000459.

Owen:2014:ALE

[OS14] J. Michael Owen and Mikhail Shashkov. Arbitrary Lagrangian Eulerian remap treatments consistent with staggered compatible total energy conserving Lagrangian methods. Journal of Computational Physics, 273(??):520–547, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003696.

Osting:2010:OSF

[Ost10] Braxton Osting. Optimization of spectral functions of Dirichlet–Laplacian eigenvalues. Journal of Computational Physics, 229(22):8578–8590, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004389.

Onishi:2013:EPS

[OTV13] Ryo Onishi, Keiko Takahashi, and J. C. Vassilicos. An efficient parallel simulation of interacting inertial particles in homogeneous isotropic turbulence. Journal of Computational Physics, 242(??):809–827, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001460.

Poursina:2012:LRF

[PA12] Mohammad Poursina and Kurt S. Anderson. Long-range force and moment calculations in multiresolution simulations of molecular systems. Journal of Computational Physics, 231(21):7237–7254, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003750.

Poursina:2014:IFM

[PA14] Mohammad Poursina and Kurt S. Anderson. An improved fast multipole method for electrostatic potential calculations in a class of coarse-grained molecular simulations. Journal of Computational Physics, 270(??):613–633, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002885.

Phillips:2011:PRN

[PAG11] Carolyn L. Phillips, Joshua A. Anderson, and Sharon C. Glotzer. Pseudo-random number generation for Brownian Dynamics and Dissipative Particle Dynamics simulations on GPU devices. Journal of Computational Physics, 230(19):7191–7201, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003329.

Puigt:2010:DDF

[PAM10] G. Puigt, V. Auffray, and J.-D. Müller. Discretisation of diffusive fluxes on hybrid grids. Journal of Computational Physics, 229(5):1425–1447, March 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005981.

Paravento:2010:RNM

[Par10] F. Paravento. A robust numerical model for premixed flames with high density ratios based on new pressure correction and IMEX schemes. Journal of Computational Physics, 229(12):4613–4647, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001051.

Parent:2012:PPH

[Par12a] Bernard Parent. Positivity-preserving high-resolution schemes for systems of conservation laws. Journal of Computational Physics, 231(1):173–189, January 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005249.

Park:2012:TDS

[Par12b] Won-Kwang Park. Topological derivative strategy for onestep iteration imaging of arbitrary shaped thin, curvelike electromagnetic inclusions. Journal of Computational Physics, 231(4):1426–1439, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006140.

Parent:2013:PPF

[Par13] Bernard Parent. Positivity-preserving flux difference splitting schemes. Journal of Computational Physics, 243(??):194–209, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001812.

Paciorri:2011:SIC

[PB11] Renato Paciorri and Aldo Bonfiglioli. Shock interaction computations on unstructured, two-dimensional grids using a shock-fitting technique. Journal of Computational Physics, 230(8):3155–3177, April 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111000362.

Peixoto:2013:AGI

[PB13a] Pedro S. Peixoto and Saulo R. M. Barros. Analysis of grid imprinting on geodesic spherical icosahedral grids. Journal of Computational Physics, 237(??):61–78, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007218.

Peron:2013:ABO

[PB13b] Stéphanie Péron and Christophe Benoit. Automatic off-body overset adaptive Cartesian mesh method based on an octree approach. Journal of Computational Physics, 232(1):153–173, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004135.

Peixoto:2014:VFR

[PB14a] Pedro S. Peixoto and Saulo R. M. Barros. On vector field reconstructions for semi-Lagrangian transport methods on geodesic staggered grids. Journal of Computational Physics, 273(??):185–211, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400312X.

Prignitz:2014:PFS

[PB14b] R. Prignitz and E. Bänsch. Particulate flows with the subspace projection method. Journal of Computational Physics, 260(??):249–272, March 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113008371.

Paster:2014:CDS

[PBB14] Amir Paster, Diogo Bolster, and David A. Benson. Connecting the dots: Semi-analytical and random walk numerical solutions of the diffusion-reaction equation with stochastic initial conditions. Journal of Computational Physics, 263(??):91–112, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000473.

Panarese:2011:MCM

[PBC^{+}11a] A. Panarese, D. Bruno, G. Colonna, P. Diomede, A. Laricchiuta, S. Longo, and M. Capitelli. A Monte Carlo model for determination of binary diffusion coefficients in gases. Journal of Computational Physics, 230(14):5716–5721, June 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002154.

Planas:2011:AIM

[PBC11b] Ramon Planas, Santiago Badia, and Ramon Codina. Approximation of the inductionless MHD problem using a stabilized finite element method. Journal of Computational Physics, 230(8):2977–2996, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000088.

Paredes:2014:PTM

[PBC^{+}14] A. Paredes, H. Bufferand, G. Ciraolo, F. Schwander, E. Serre, P. Ghendrih, and P. Tamain. A penalization technique to model plasma facing components in a tokamak with temperature variations. Journal of Computational Physics, 274(??):283–298, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003714.

Pierre:2014:NCH

[PBdGP14] Charles Pierre, Julien Bouyssier, Frédéric de Gournay, and Franck Plouraboué. Numerical computation of 3D heat transfer in complex parallel heat exchangers using generalized Graetz modes. Journal of Computational Physics, 268(??):84–105, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001636.

Pelanti:2011:RSS

[PBM11] Marica Pelanti, François Bouchut, and Anne Mangeney. A Riemann solver for single-phase and two-phase shallow flow models based on relaxation. Relations with Roe and VFRoe solvers. Journal of Computational Physics, 230(3):515–550, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005450.

Pan:2014:SPH

[PBT14] Wenxiao Pan, Jie Bao, and Alexandre M. Tartakovsky. Smoothed particle hydrodynamics continuous boundary force method for Navier–Stokes equations subject to a Robin boundary condition. Journal of Computational Physics, 259(??): 242–259, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113008139.

Park:2011:UTM

[PBWK11] C.-H. Park, N. Böttcher, W. Wang, and O. Kolditz. Are upwind techniques in multi-phase flow models necessary? Journal of Computational Physics, 230(22):8304–8312, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100458X.

Pirozzoli:2013:GCR

[PC13] Sergio Pirozzoli and Tim Colonius. Generalized characteristic relaxation boundary conditions for unsteady compressible flow simulations. Journal of Computational Physics, 248 (??):109–126, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113002866.

Pataki:2013:FHO

[PCF^{+}13] Andras Pataki, Antoine J. Cerfon, Jeffrey P. Freidberg, Leslie Greengard, and Michael O’Neil. A fast, high-order solver for the Grad–Shafranov equation. Journal of Computational Physics, 243(??):28–45, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001721.

Plante:2013:RSG

[PDC13] Ianik Plante, Luc Devroye, and Francis A. Cucinotta. Random sampling of the Green’s functions for reversible reactions with an intermediate state. Journal of Computational Physics, 242(??):531–543, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001009.

P:2010:CFV

[PDSS10] Arnaud Fosso P., Hugues Deniau, Frédéric Sicot, and Pierre Sagaut. Curvilinear finite-volume schemes using high-order compact interpolation. Journal of Computational Physics, 229(13):5090–5122, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001464.

Panda:2014:DGM

[PDZ^{+}14] Nishant Panda, Clint Dawson, Yao Zhang, Andrew B. Kennedy, Joannes J. Westerink, and Aaron S. Donahue. Discontinuous Galerkin methods for solving Boussinesq– Green–Naghdi equations in resolving non-linear and dispersive surface water waves. Journal of Computational Physics, 273(??):572–588, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003982.

Pearce:2010:PSA

[PE10] J. D. Pearce and J. G. Esler. A pseudo-spectral algorithm and test cases for the numerical solution of the twodimensional rotating Green–Naghdi shallow water equations. Journal of Computational Physics, 229(20):7594–7608, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003244.

Persson:2013:SHO

[Per13] Per-Olof Persson. A sparse and high-order accurate line-based discontinuous Galerkin method for unstructured meshes. Journal of Computational Physics, 233(??):414–429, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005384.

Pataki:2011:FES

[PG11] Andras Pataki and Leslie Greengard. Fast elliptic solvers in cylindrical coordinates and the Coulomb collision operator. Journal of Computational Physics, 230(21):7840–7852, September 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004098.

Poux:2011:IOT

[PGA11] A. Poux, S. Glockner, and M. Azaïez. Improvements on open and traction boundary conditions for Navier– Stokes time-splitting methods. Journal of Computational Physics, 230(10):4011–4027, May 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001136.

Parsani:2011:AIL

[PGL11] M. Parsani, G. Ghorbaniasl, and C. Lacor. Analysis of the implicit LU–SGS algorithm for 3rd- and 4th-order spectral volume scheme for solving the steady Navier–Stokes equations. Journal of Computational Physics, 230(19):7073–7085, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003378.

Perko:2014:GBA

[PGLK14] Zoltán Perkó, Luca Gilli, Danny Lathouwers, and Jan Leen Kloosterman. Grid and basis adaptive polynomial chaos techniques for sensitivity and uncertainty analysis. Journal of Computational Physics, 260(??):54–84, March 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008322.

Parsani:2010:IHO

[PGLT10] M. Parsani, G. Ghorbaniasl, C. Lacor, and E. Turkel. An implicit high-order spectral difference approach for large eddy simulation. Journal of Computational Physics, 229 (14):5373–5393, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001579.

Papac:2010:ESD

[PGR10] Joseph Papac, Frédéric Gibou, and Christian Ratsch. Efficient symmetric discretization for the Poisson, heat and Stefan-type problems with Robin boundary conditions. Journal of Computational Physics, 229(3):875–889, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005622.

Piret:2013:RBF

[PH13] Cécile Piret and Emmanuel Hanert. A radial basis functions method for fractional diffusion equations. Journal of Computational Physics, 238(??):71–81, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006444.

Peng:2014:WIM

[PHD14] Ji Peng, Jerrad Hampton, and Alireza Doostan. A weighted ℓ_{1}-minimization approach for sparse polynomial chaos expansions. Journal of Computational Physics, 267(??):92–111, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001442.

Phillips:2014:LHS

[Phi14] Carolyn L. Phillips. A learning heuristic for space mapping and searching self-organizing systems using adaptive mesh refinement. Journal of Computational Physics, 272(??):799–813, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003337.

Papac:2013:LSA

[PHRG13] Joseph Papac, Asdis Helgadottir, Christian Ratsch, and Frederic Gibou. A level set approach for diffusion and Stefan-type problems with Robin boundary conditions on quadtree/octree adaptive Cartesian grids. Journal of Computational Physics, 233(??):241–261, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005062.

Plunkett:2014:SAS

[PHSA14] Pat Plunkett, Jonathan Hu, Christopher Siefert, and Paul J. Atzberger. Spatially adaptive stochastic methods for fluidstructure interactions subject to thermal fluctuations in domains with complex geometries. Journal of Computational Physics, 277(??):121–137, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005440.

Pierre:2012:PBM

[Pie12] Charles Pierre. Preconditioning the bidomain model with almost linear complexity. Journal of Computational Physics, 231(1):82–97, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111005122.

Pettersson:2014:SGM

[PIN14] Per Pettersson, Gianluca Iaccarino, and Jan Nordström. A stochastic Galerkin method for the Euler equations with Roe variable transformation. Journal of Computational Physics, 257(??):481–500, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006797.

Pirozzoli:2010:GCA

[Pir10] Sergio Pirozzoli. Generalized conservative approximations of split convective derivative operators. Journal of Computational Physics, 229(19):7180–7190, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003219.

Pirozzoli:2011:SND

[Pir11] Sergio Pirozzoli. Stabilized non-dissipative approximations of Euler equations in generalized curvilinear coordinates. Journal of Computational Physics, 230(8):2997–3014, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100009X.

Piret:2012:OGM

[Pir12] Cécile Piret. The orthogonal gradients method: a radial basis functions method for solving partial differential equations on arbitrary surfaces. Journal of Computational Physics, 231(14):4662–4675, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001477.

Pletzer:2014:CCC

[PJCS14] Alexander Pletzer, Ben Jamroz, Robert Crockett, and Scott Sides. Compact cell-centered discretization stencils at finecoarse block structured grid interfaces. Journal of Computational Physics, 260(??):25–36, March 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300819X.

Peng:2010:NCD

[PL10] Zhen Peng and Jin-Fa Lee. Non-conformal domain decomposition method with second-order transmission conditions for time-harmonic electromagnetics. Journal of Computational Physics, 229(16):5615–5629, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001683.

Poette:2012:NII

[PL12] Gaël Poëtte and Didier Lucor. Non intrusive iterative stochastic spectral representation with application to compressible gas dynamics. Journal of Computational Physics, 231(9):3587– 3609, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000058.

Patronis:2014:MSN

[PL14] Alexander Patronis and Duncan A. Lockerby. Multiscale simulation of non-isothermal microchannel gas flows. Journal of Computational Physics, 270(??):532–543, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002587.

Patronis:2013:HCM

[PLBR13] Alexander Patronis, Duncan A. Lockerby, Matthew K. Borg, and Jason M. Reese. Hybrid continuum-molecular modelling of multiscale internal gas flows. Journal of Computational Physics, 255(??):558–571, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005718.

Potter:2011:FSF

[PLN11] M. E. Potter, M. Lamoureux, and M. D. Nauta. An FDTD scheme on a face-centered-cubic (FCC) grid for the solution of the wave equation. Journal of Computational Physics, 230(15):6169–6183, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002750.

Peigney:2014:FPK

[PLT14] B. E. Peigney, O. Larroche, and V. Tikhonchuk. Fokker–Planck kinetic modeling of suprathermal α-particles in a fusion plasma. Journal of Computational Physics, 278(??):416–444, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006019.

Polivka:2014:CMP

[PM14] Ondrej Polívka and Jirí Mikyska. Compositional modeling in porous media using constant volume flash and flux computation without the need for phase identification. Journal of Computational Physics, 272(??):149–169, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002927.

Pfeiffer:2013:GIP

[PMF13] M. Pfeiffer, A. Mirza, and S. Fasoulas. A grid-independent particle pairing strategy for DSMC. Journal of Computational Physics, 246(??):28–36, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001964.

Pardo:2011:SWS

[PMM^{+}11] David Pardo, Pawel Matuszyk, Ignacio Muga, Carlos Torres-Verdín, Angel Mora, and Victor M. Calo. Simulation of wireline sonic logging measurements acquired with Borehole– Eccentered tools using a high-order adaptive finite-element method. Journal of Computational Physics, 230(16):6320– 6333, July 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111002762.

Pethiyagoda:2014:JFN

[PMMB14] Ravindra Pethiyagoda, Scott W. McCue, Timothy J. Moroney, and Julian M. Back. Jacobian-free Newton–Krylov methods with GPU acceleration for computing nonlinear ship wave patterns. Journal of Computational Physics, 269(??):297–313, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002058.

Parent:2011:ADD

[PMS11] Bernard Parent, Sergey O. Macheret, and Mikhail N. Shneider. Ambipolar diffusion and drift in computational weakly-ionized plasmadynamics. Journal of Computational Physics, 230(22):8010–8027, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004116.

Parent:2014:EIT

[PMS14] Bernard Parent, Sergey O. Macheret, and Mikhail N. Shneider. Electron and ion transport equations in computational weakly-ionized plasmadynamics. Journal of Computational Physics, 259(??):51–69, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007912.

Peng:2010:NCG

[PMSH10] Yih-Ferng Peng, Rajat Mittal, Amalendu Sau, and Robert R. Hwang. Nested Cartesian grid method in incompressible viscous fluid flow. Journal of Computational Physics, 229(19):7072–7101, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000313X.

Pisarenco:2012:ESM

[PMSM12] M. Pisarenco, J. M. L. Maubach, I. D. Setija, and R. M. M. Mattheij. Efficient solution of Maxwell’s equations for geometries with repeating patterns by an exchange of discretization directions in the aperiodic Fourier modal method. Journal of Computational Physics, 231(24):8209–8228, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004378.

Pinelli:2010:IBM

[PNPF10] A. Pinelli, I. Z. Naqavi, U. Piomelli, and J. Favier. Immersedboundary methods for general finite-difference and finitevolume Navier–Stokes solvers. Journal of Computational Physics, 229(24):9073–9091, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004687.

Pieraccini:2012:MIM

[PP12] Sandra Pieraccini and Gabriella Puppo. Microscopically implicit-macroscopically explicit schemes for the BGK equation. Journal of Computational Physics, 231(2):299–327, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005146.

Popov:2014:IES

[PP14] Pavel P. Popov and Stephen B. Pope. Implicit and explicit schemes for mass consistency preservation in hybrid particle/ finite-volume algorithms for turbulent reactive flows. Journal of Computational Physics, 257(??):352–373, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006104.

Patil:2014:MLB

[PPB14] Dhiraj V. Patil, Kannan N. Premnath, and Sanjoy Banerjee. Multigrid lattice Boltzmann method for accelerated solution of elliptic equations. Journal of Computational Physics, 265(??):172–194, May 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000941.

Peres:2012:PMC

[PPS12] Noele Peres, Sébastien Poncet, and Eric Serre. A 3D pseudospectral method for cylindrical coordinates. application to the simulations of rotating cavity flows. Journal of Computational Physics, 231(19):6290–6305, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200215X.

Pan:2013:NIC

[PPTMDK13] Dingyi Pan, Nhan Phan-Thien, Nam Mai-Duy, and Boo Cheong Khoo. Numerical investigations on the compressibility of a DPD fluid. Journal of Computational Physics, 242(??):196– 210, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300123X.

Prakash:2011:HOS

[PPWZ11] A. Prakash, N. Parsons, X. Wang, and X. Zhong. High-order shock-fitting methods for direct numerical simulation of hypersonic flow with chemical and thermal nonequilibrium. Journal of Computational Physics, 230(23):8474–8507, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111004591.

Payette:2011:RML

[PR11] G. S. Payette and J. N. Reddy. On the roles of minimization and linearization in least-squares finite element models of nonlinear boundary-value problems. Journal of Computational Physics, 230(9):3589–3613, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000842.

Puri:2014:ARS

[PR14a] Kunal Puri and Prabhu Ramachandran. Approximate Riemann solvers for the Godunov SPH (GSPH). Journal of Computational Physics, 270(??):432–458, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002423.

Puri:2014:CSS

[PR14b] Kunal Puri and Prabhu Ramachandran. A comparison of SPH schemes for the compressible Euler equations. Journal of Computational Physics, 256(??):308–333, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006049.

Peco:2013:AMM

[PRA13] C. Peco, A. Rosolen, and M. Arroyo. An adaptive meshfree method for phase-field models of biomembranes. Part II: a Lagrangian approach for membranes in viscous fluids. Journal of Computational Physics, 249(??):320–336, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003239.

Palha:2014:PCD

[PRH^{+}14] Artur Palha, Pedro Pinto Rebelo, René Hiemstra, Jasper Kreeft, and Marc Gerritsma. Physics-compatible discretization techniques on single and dual grids, with application to the Poisson equation of volume forms. Journal of Computational Physics, 257(??):1394–1422, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005330.

Price:2012:SPH

[Pri12] Daniel J. Price. Smoothed particle hydrodynamics and magnetohydrodynamics. Journal of Computational Physics, 231 (3):759–794, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110006753.

Priimak:2014:FDN

[Pri14] Dmitri Priimak. Finite difference numerical method for the superlattice Boltzmann transport equation and case comparison of CPU(C) and GPU(CUDA) implementations. Journal of Computational Physics, 278(??):182–192, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005828.

Peng:2010:OWD

[PRL10] Zhen Peng, Vineet Rawat, and Jin-Fa Lee. One way domain decomposition method with second order transmission conditions for solving electromagnetic wave problems. Journal of Computational Physics, 229(4):1181–1197, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005762.

Pang:2012:MMF

[PS12] Hong-Kui Pang and Hai-Wei Sun. Multigrid method for fractional diffusion equations. Journal of Computational Physics, 231(2):693–703, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005985.

Pelanti:2014:MEC

[PS14a] Marica Pelanti and Keh-Ming Shyue. A mixture-energy-consistent six-equation two-phase numerical model for fluids with interfaces, cavitation and evaporation waves. Journal of Computational Physics, 259(??):331–357, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008024. See corrigendum

[PS15].

Pisarenco:2014:CAF

[PS14b] M. Pisarenco and I. D. Setija. On the complexity of aperiodic Fourier modal methods for finite periodic structures. Journal of Computational Physics, 261(??):130–144, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008590.

Pelanti:2015:CME

[PS15] Marica Pelanti and Keh-Ming Shyue. Corrigendum to “A mixture-energy-consistent six-equation two-phase numerical model for fluids with interfaces, cavitation and evaporation waves” [J. Comput. Phys. 259 (2014) 331–357]. Journal of Computational Physics, 288(??):196–197, May 1, 2015. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999115000406. See [PS14a].

Perrin:2013:KLE

[PSDF13] G. Perrin, C. Soize, D. Duhamel, and C. Funfschilling. Karhunen–Loève expansion revisited for vector-valued random fields: Scaling, errors and optimal basis. Journal of Computational Physics, 242(??):607–622, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001551.

Pinto:2014:NVP

[PSF^{+}14] Martin Campos Pinto, Eric Sonnendrücker, Alex Friedman, David P. Grote, and Steve M. Lund. Noiseless Vlasov– Poisson simulations with linearly transformed particles. Journal of Computational Physics, 275(??):236–256, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004392.

Padrino:2020:CAS

[PSL20] Juan C. Padrino, James E. Sprittles, and Duncan A. Lockerby. Comment on “Applying a second-kind boundary integral equation for surface tractions in Stokes flow”. Journal of Computational Physics, 401(??):Article 109007, January 15, 2020. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999119307132. See [KS11c].

Parent:2011:GOL

[PSM11] Bernard Parent, Mikhail N. Shneider, and Sergey O. Macheret. Generalized Ohm’s law and potential equation in computational weakly-ionized plasmadynamics. Journal of Computational Physics, 230(4):1439–1453, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006200.

Parent:2013:SGE

[PSM13] Bernard Parent, Mikhail N. Shneider, and Sergey O. Macheret. Sheath governing equations in computational weakly-ionized plasmadynamics. Journal of Computational Physics, 232(1):234–251, January 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004573.

Petropavlovsky:2012:NDA

[PT12] S. V. Petropavlovsky and S. V. Tsynkov. A non-deteriorating algorithm for computational electromagnetism based on quasi-lacunae of Maxwell’s equations. Journal of Computational Physics, 231(2):558–585, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005511.

Pan:2013:SPH

[PTM13] W. Pan, A. M. Tartakovsky, and J. J. Monaghan. Smoothed particle hydrodynamics non-Newtonian model for ice-sheet and ice-shelf dynamics. Journal of Computational Physics, 242(??):828–842, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112006286.

Pudykiewicz:2011:NSS

[Pud11] Janusz A. Pudykiewicz. On numerical solution of the shallow water equations with chemical reactions on icosahedral geodesic grid. Journal of Computational Physics, 230(5):1956–1991, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006601.

Peerenboom:2011:ACM

[PvDGvM11] K. S. C. Peerenboom, J. van Dijk, W. J. Goedheer, and J. J. A. M. vander Mullen. On the ambipolar constraint in multi-component diffusion problems. Journal of Computational Physics, 230(10):3651–3655, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001185.

Peerenboom:2011:MCF

[PvDtTB^{+}11] K. S. C. Peerenboom, J. van Dijk, J. H. M. ten Thije Boonkkamp, L. Liu, W. J. Goedheer, and J. J. A. M. van der Mullen. Mass conservative finite volume discretization of the continuity equations in multi-component mixtures. Journal of Computational Physics, 230(9):3525–3537, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000805.

Polimeridis:2014:SFJ

[PVDW14] A. G. Polimeridis, J. F. Villena, L. Daniel, and J. K. White. Stable FFT–JVIE solvers for fast analysis of highly inhomogeneous dielectric objects. Journal of Computational Physics, 269(??):280–296, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002071.

Parsani:2010:ILS

[PVLT10] M. Parsani, K. Van den Abeele, C. Lacor, and E. Turkel. Implicit LU–SGS algorithm for high-order methods on unstructured grid with p-multigrid strategy for solving the steady Navier–Stokes equations. Journal of Computational Physics, 229(3):828–850, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005592.

Philip:2014:DIA

[PWB^{+}14] B. Philip, Z. Wang, M. A. Berrill, M. Birke, and M. Pernice. Dynamic implicit 3D adaptive mesh refinement for non-equilibrium radiation diffusion. Journal of Computational Physics, 262(??):17–37, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000138.

Patterson:2011:SWP

[PWK11] Robert I. A. Patterson, Wolfgang Wagner, and Markus Kraft. Stochastic weighted particle methods for population balance equations. Journal of Computational Physics, 230(19):7456– 7472, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111003603.

Park:2010:MDL

[PYK10] Jin Seok Park, Sung-Hwan Yoon, and Chongam Kim. Multi-dimensional limiting process for hyperbolic conservation laws on unstructured grids. Journal of Computational Physics, 229(3):788–812, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005567.

Preisig:2010:TPF

[PZ10] Matthias Preisig and Thomas Zimmermann. Two-phase freesurface fluid dynamics on moving domains. Journal of Computational Physics, 229(7):2740–2758, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006998.

Perot:2014:DFS

[PZ14] J. Blair Perot and Christopher J. Zusi. Differential forms for scientists and engineers. Journal of Computational Physics, 257(??):1373–1393, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005354.

Paolucci:2014:WAWb

[PZG14] Samuel Paolucci, Zachary J. Zikoski, and Temistocle Grenga. WAMR: an adaptive wavelet method for the simulation of compressible reacting flow. Part II. The parallel algorithm. Journal of Computational Physics, 272(??):842–864, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002460.

Pan:2012:GKS

[PZTW12] Liang Pan, Guiping Zhao, Baolin Tian, and Shuanghu Wang. A gas kinetic scheme for the Baer–Nunziato twophase flow model. Journal of Computational Physics, 231(22):7518–7536, September 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002537.

Paolucci:2014:WAWa

[PZW14] Samuel Paolucci, Zachary J. Zikoski, and Damrongsak Wirasaet. WAMR: an adaptive wavelet method for the simulation of compressible reacting flow. Part I. Accuracy and efficiency of algorithm. Journal of Computational Physics, 272(??):814–841, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000527.

Quinn:2011:DAU

[QA11] John C. Quinn and Henry D. I. Abarbanel. Data assimilation using a GPU accelerated path integral Monte Carlo approach. Journal of Computational Physics, 230(22):8168– 8178, September 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004347.

Quaife:2014:HVF

[QB14] Bryan Quaife and George Biros. High-volume fraction simulations of two-dimensional vesicle suspensions. Journal of Computational Physics, 274(??):245–267, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004203.

Qiu:2010:CHO

[QC10] Jing-Mei Qiu and Andrew Christlieb. A conservative high order semi-Lagrangian WENO method for the Vlasov equation. Journal of Computational Physics, 229(4):1130–1149, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005610.

Qian:2011:CLT

[QL11] ZhanSen Qian and Chun-Hian Lee. A class of large time step Godunov schemes for hyperbolic conservation laws and applications. Journal of Computational Physics, 230(19):7418–7440, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003573.

Qian:2012:LTS

[QL12] ZhanSen Qian and Chun-Hian Lee. On large time step TVD scheme for hyperbolic conservation laws and its efficiency evaluation. Journal of Computational Physics, 231(21):7415–7430, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003889.

Qian:2014:GRP

[QLW14] Jianzhen Qian, Jiequan Li, and Shuanghu Wang. The generalized Riemann problems for compressible fluid flows: Towards high order. Journal of Computational Physics, 259 (??):358–389, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113008000.

Qiu:2011:CHO

[QS11a] Jing-Mei Qiu and Chi-Wang Shu. Conservative high order semi-Lagrangian finite difference WENO methods for advection in incompressible flow. Journal of Computational Physics, 230(4):863–889, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000224X.

Qiu:2011:PPS

[QS11b] Jing-Mei Qiu and Chi-Wang Shu. Positivity preserving semi-Lagrangian discontinuous Galerkin formulation: Theoretical analysis and application to the Vlasov–Poisson system. Journal of Computational Physics, 230(23):8386–8409, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111004372.

Quan:2011:SMF

[Qua11] Shaoping Quan. Simulations of multiphase flows with multiple length scales using moving mesh interface tracking with adaptive meshing. Journal of Computational Physics, 230(13):5430–5448, June 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002117.

Qian:2010:FGW

[QY10] Jianliang Qian and Lexing Ying. Fast Gaussian wavepacket transforms and Gaussian beams for the Schrödinger equation. Journal of Computational Physics, 229(20):7848–7873, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003669.

Qamar:2012:STC

[QY12] Shamsul Qamar and Muhammad Yousaf. The space–time CESE method for solving special relativistic hydrodynamic equations. Journal of Computational Physics, 231(10):3928– 3945, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112000708.

Qiu:2012:FLS

[QZM^{+}12] Zhouhua Qiu, Zhong Zeng, Huan Mei, Liang Li, Liping Yao, and Liangqi Zhang. A Fourier–Legendre spectral element method in polar coordinates. Journal of Computational Physics, 231(2):666–675, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005869.

Rendall:2010:RSP

[RA10a] T. C. S. Rendall and C. B. Allen. Reduced surface point selection options for efficient mesh deformation using radial basis functions. Journal of Computational Physics, 229(8):2810–2820, April 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006779.

Ricchiuto:2010:ERK

[RA10b]  M. Ricchiuto and R. Abgrall.  Explicit Runge–Kutta

residual distribution schemes for time dependent problems: Second order case. Journal of Computational Physics, 229(16):5653–5691, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110001786.

Rangan:2011:EMG

[Ran11] Aaditya V. Rangan. Efficient methods for grouping vectors into low-rank clusters. Journal of Computational Physics, 230(14):5684–5703, June 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002099.

Rangan:2012:DLR

[Ran12a] Aaditya V. Rangan. Detecting low-rank clusters via random sampling. Journal of Computational Physics, 231(1):215–222, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111005286.

Rangan:2012:SFD

[Ran12b] Aaditya V. Rangan. A simple filter for detecting low-rank submatrices. Journal of Computational Physics, 231(7):2682–2690, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111007534.

Radice:2013:NSH

[RARO13] David Radice, Ernazar Abdikamalov, Luciano Rezzolla, and Christian D. Ott. A new spherical harmonics scheme for multi-dimensional radiation transport I. Static matter configurations. Journal of Computational Physics, 242(??):648–669, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001125.

Rosatti:2010:RPO

[RB10] G. Rosatti and L. Begnudelli. The Riemann Problem for the one-dimensional, free-surface Shallow Water Equations with a bed step: Theoretical analysis and numerical simulations. Journal of Computational Physics, 229(3):760–787, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005555.

Rosatti:2013:CIG

[RB13] Giorgio Rosatti and Lorenzo Begnudelli. A closure-independent generalized Roe solver for free-surface, two-phase flows over mobile bed. Journal of Computational Physics, 255(??): 362–383, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005585.

Rocca:2014:LSR

[RB14] G. Della Rocca and G. Blanquart. Level set reinitialization at a contact line. Journal of Computational Physics, 265(??):34– 49, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000783.

Raeini:2012:MTP

[RBB12] Ali Q. Raeini, Martin J. Blunt, and Branko Bijeljic. Modelling two-phase flow in porous media at the pore scale using the volume-of-fluid method. Journal of Computational Physics, 231(17):5653–5668, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001830.

Rossinelli:2010:GAS

[RBCK10] Diego Rossinelli, Michael Bergdorf, Georges-Henri Cottet, and Petros Koumoutsakos. GPU accelerated simulations of bluff body flows using vortex particle methods. Journal of Computational Physics, 229(9):3316–3333, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000197.

Ruas:2013:HFE

[RBK13] Vitoriano Ruas, Diego Brandão, and Mauricio Kischinhevsky. Hermite finite elements for diffusion phenomena. Journal of Computational Physics, 235(??):542–564, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005815.

Reynolds-Barredo:2012:MCT

[RBNS^{+}12] J. M. Reynolds-Barredo, D. E. Newman, R. Sanchez, D. Samaddar, L. A. Berry, and W. R. Elwasif. Mechanisms for the convergence of time-parallelized, parareal turbulent plasma simulations. Journal of Computational Physics, 231(23):7851–7867, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004123.

Reynolds-Barredo:2013:AMC

[RBNS13] J. M. Reynolds-Barredo, D. E. Newman, and R. Sanchez. An analytic model for the convergence of turbulent simulations time-parallelized via the parareal algorithm. Journal of Computational Physics, 255(??):293–315, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005664.

Rajpoot:2012:SLR

[RBS12] Manoj K. Rajpoot, Swagata Bhaumik, and Tapan K. Sengupta. Solution of linearized rotating shallow water equations by compact schemes with different grid-staggering strategies. Journal of Computational Physics, 231(5):2300–2327, March 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006802.

Rhebergen:2012:STH

[RC12] Sander Rhebergen and Bernardo Cockburn. A space–time hybridizable discontinuous Galerkin method for incompressible flows on deforming domains. Journal of Computational Physics, 231(11):4185–4204, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000903.

Roger:2014:HTD

[RCCC14] M. Roger, C. Caliot, N. Crouseilles, and P. J. Coelho. A hybrid transport-diffusion model for radiative transfer in absorbing and scattering media. Journal of Computational Physics, 275(??):346–362, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400494X.

Richardson:2010:FEF

[RCCW10] G. A. Richardson, J. T. Cassibry, T. J. Chung, and S. T. Wu. Finite element form of FDV for widely varying flowfields. Journal of Computational Physics, 229(1):145–167, January 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005154.

Ran:2011:GAC

[RCQL11] Wei Ran, Wan Cheng, Fenghua Qin, and Xisheng Luo. GPU accelerated CESE method for 1D shock tube problems. Journal of Computational Physics, 230(24):8797–8812, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005134.

Rhebergen:2013:STD

[RCvdV13] Sander Rhebergen, Bernardo Cockburn, and Jaap J. W. van der Vegt. A space–time discontinuous Galerkin method for the incompressible Navier–Stokes equations. Journal of Computational Physics, 233(??):339–358, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005372.

Rasmussen:2011:MRV

[RCW11] Johannes Tophøj Rasmussen, Georges-Henri Cottet, and Jens Honoré Walther. A multiresolution remeshed Vortex-In-Cell algorithm using patches. Journal of Computational Physics, 230(17):6742–6755, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003032.

Rohlfs:2012:TPE

[RDHK12] Wilko Rohlfs, Georg F. Dietze, Herman D. Haustein, and Reinhold Kneer. Two-phase electrohydrodynamic simulations using a volume-of-fluid approach: a comment. Journal of Computational Physics, 231(12):4454–4463, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000733.

Redeker:2013:FAA

[RE13] Magnus Redeker and Christof Eck. A fast and accurate adaptive solution strategy for two-scale models with continuous inter-scale dependencies. Journal of Computational Physics, 240(??):268–283, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007565.

Remis:2011:RBF

[Rem11] Rob F. Remis. On the relation between FDTD and Fibonacci polynomials. Journal of Computational Physics, 230(4):1382–1386, February 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000611X.

Renac:2011:IRP

[Ren11] Florent Renac. Improvement of the recursive projection method for linear iterative scheme stabilization based on an approximate eigenvalue problem. Journal of Computational Physics, 230(14):5739–5752, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002191.

Ricchiuto:2014:URD

[RF14a] M. Ricchiuto and A. G. Filippini. Upwind residual discretization of enhanced Boussinesq equations for wave propagation over complex bathymetries. Journal of Computational Physics, 271(??):306–341, August 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008565.

Roberts:2014:MDP

[RF14b] Jeremy A. Roberts and Benoit Forget. Multigroup diffusion preconditioners for multiplying fixed-source transport problems. Journal of Computational Physics, 274(??):455–472, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004410.

Rycroft:2012:SSB

[RG12] Chris H. Rycroft and Frédéric Gibou. Simulations of a stretching bar using a plasticity model from the shear transformation zone theory. Journal of Computational Physics, 231(5):2155–2179, March 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006024.

Rasthofer:2013:MSS

[RG13] U. Rasthofer and V. Gravemeier. Multifractal subgridscale modeling within a variational multiscale method for large-eddy simulation of turbulent flow. Journal of Computational Physics, 234(??):79–107, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200544X.

Ragusa:2012:RDA

[RGK12] J. C. Ragusa, J.-L. Guermond, and G. Kanschat. A robust S_{N}-DG-approximation for radiation transport in optically thick and diffusive regimes. Journal of Computational Physics, 231(4):1947–1962, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006723.

Renac:2013:FTI

[RGMC13] Florent Renac, Sophie Gérald, Claude Marmignon, and Frédéric Coquel. Fast time implicit-explicit discontinuous Galerkin method for the compressible Navier–Stokes equations. Journal of Computational Physics, 251(??):272–291, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004105.

Razafindralandy:2013:TIA

[RH13] Dina Razafindralandy and Aziz Hamdouni. Time integration algorithm based on divergent series resummation, for ordinary and partial differential equations. Journal of Computational Physics, 236(??):56–73, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006237.

Reichert:2012:ESN

[RHB12] Adam Reichert, Michael T. Heath, and Daniel J. Bodony. Energy stable numerical methods for hyperbolic partial differential equations using overlapping domain decomposition. Journal of Computational Physics, 231(16):5243–5265, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001398.

Rivera:2010:PFE

[RHGT10] Christian A. Rivera, Mourad Heniche, Roland Glowinski, and Philippe A. Tanguy. Parallel finite element simulations of incompressible viscous fluid flow by domain decomposition with Lagrange multipliers. Journal of Computational Physics, 229(13):5123–5143, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001476.

Ren:2014:RKC

[RHXQ14] Tan Ren, Jun Hu, Tao Xiong, and Jing-Mei Qiu. Runge– Kutta central discontinuous Galerkin BGK method for the Navier–Stokes equations. Journal of Computational Physics, 274(??):592–610, October 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114004525.

Richter:2013:FEF

[Ric13] Thomas Richter. A fully Eulerian formulation for fluidstructure-interaction problems. Journal of Computational Physics, 233(??):227–240, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112005153.

Rieper:2010:DMU

[Rie10] Felix Rieper. On the dissipation mechanism of upwind-schemes in the low Mach number regime: a comparison between Roe and HLL. Journal of Computational Physics, 229(2):221–232, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199910900535X.

Rieper:2011:LMN

[Rie11] Felix Rieper. A low-Mach number fix for Roe’s approximate Riemann solver. Journal of Computational Physics, 230(13):5263–5287, June 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111001689.

Rehman:2013:CEM

[RJC13] Tafizur Rehman, M. Jaipal, and Abhijit Chatterjee. A cluster expansion model for predicting activation barrier of atomic processes. Journal of Computational Physics, 243(??):244–259, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001782.

Rupp:2010:MCS

[RJG10] K. Rupp, A. Jüngel, and T. Grasser. Matrix compression for spherical harmonics expansions of the Boltzmann transport equation for semiconductors. Journal of Computational Physics, 229(23):8750–8765, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004456.

Randles:2014:PTA

[RK14] Amanda Randles and Efthimios Kaxiras. Parallel in time approximation of the lattice Boltzmann method for laminar flows. Journal of Computational Physics, 270(??):577–586, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002605.

Ridgers:2014:MGR

[RKD^{+}14] C. P. Ridgers, J. G. Kirk, R. Duclous, T. G. Blackburn, C. S. Brady, K. Bennett, T. D. Arber, and A. R. Bell. Modelling gamma-ray photon emission and pair production in high-intensity laser-matter interactions. Journal of Computational Physics, 260(??):273–285, March 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008061.

Ren:2014:HBE

[RKGM14] Zhengyong Ren, Thomas Kalscheuer, Stewart Greenhalgh, and Hansruedi Maurer. A hybrid boundary element-finite element approach to modeling plane wave 3D electromagnetic induction responses in the Earth. Journal of Computational Physics, 258(??):705–717, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300750X.

Rauser:2011:PGE

[RKM11] Florian Rauser, Peter Korn, and Jochem Marotzke. Predicting goal error evolution from near-initial-information: a learning algorithm. Journal of Computational Physics, 230 (19):7284–7299, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111003408.

Ren:2014:APM

[RLJ14] Wei Ren, Hong Liu, and Shi Jin. An asymptotic-preserving Monte Carlo method for the Boltzmann equation. Journal of Computational Physics, 276(??):380–404, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005221.

Rosic:2012:SFL

[RLPM12] Bojana V. Rosić, Alexander Litvinenko, Oliver Pajonk, and Hermann G. Matthies. Sampling-free linear Bayesian update of polynomial chaos representations. Journal of Computational Physics, 231(17):5761–5787, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002483.

Rukavishnikov:2012:NNM

[RM12] Victor A. Rukavishnikov and Andrey O. Mosolapov. New numerical method for solving time-harmonic Maxwell equations with strong singularity. Journal of Computational Physics, 231(6):2438–2448, March 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006863.

Rybak:2014:MTS

[RM14] Iryna Rybak and Jim Magiera. A multiple-time-step technique for coupled free flow and porous medium systems. Journal of Computational Physics, 272(??):327–342, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400299X.

Rufat:2014:CCM

[RMMD14] Dzhelil Rufat, Gemma Mason, Patrick Mullen, and Mathieu Desbrun. The chain collocation method: a spectrally accurate calculus of forms. Journal of Computational Physics, 257(??):1352–1372, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005494.

Robinson-Mosher:2011:SPD

[RMSF11] Avi Robinson-Mosher, Craig Schroeder, and Ronald Fedkiw. A symmetric positive definite formulation for monolithic fluid structure interaction. Journal of Computational Physics, 230(4):1547–1566, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006364.

Ryser:2012:WPS

[RNT12] Marc D. Ryser, Nilima Nigam, and Paul F. Tupper. On the well-posedness of the stochastic Allen–Cahn equation in two dimensions. Journal of Computational Physics, 231 (6):2537–2550, March 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111007005.

Rodionov:2012:CKP

[Rod12] Alexander V. Rodionov. Complement to the “Kolgan project”. Journal of Computational Physics, 231(13):4465–4468, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001520.

Rosswog:2010:CSR

[Ros10] Stephan Rosswog. Conservative, special-relativistic smoothed particle hydrodynamics. Journal of Computational Physics, 229(22):8591–8612, November 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004390.

Rothkopf:2013:IME

[Rot13] Alexander Rothkopf. Improved maximum entropy analysis with an extended search space. Journal of Computational Physics, 238(??):106–114, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007541.

Ranjan:2013:CMI

[RP13] R. Ranjan and C. Pantano. A collocated method for the incompressible Navier–Stokes equations inspired by the Box scheme. Journal of Computational Physics, 232(1):346–382, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004676.

Rosolen:2013:AMM

[RPA13] A. Rosolen, C. Peco, and M. Arroyo. An adaptive meshfree method for phase-field models of biomembranes. Part I: Approximation with maximum-entropy basis functions. Journal of Computational Physics, 249(??):303–319, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003483.

Rinaldi:2014:EJI

[RPC14] Enrico Rinaldi, Rene Pecnik, and Piero Colonna. Exact Jacobians for implicit Navier–Stokes simulations of equilibrium real gas flows. Journal of Computational Physics, 270(??):459–477, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002459.

Ranjith:2013:NSB

[RPV13] S. Kumar Ranjith, B. S. V. Patnaik, and Srikanth Vedantam. No-slip boundary condition in finite-size dissipative particle dynamics. Journal of Computational Physics, 232(1):174–188, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004317.

Reuther:2014:SDS

[RR14] K. Reuther and M. Rettenmayr. Simulating dendritic solidification using an anisotropy-free meshless front-tracking method. Journal of Computational Physics, 279(??):63–66, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006299.

Ricketson:2014:EBT

[RRCD14] L. F. Ricketson, M. S. Rosin, R. E. Caflisch, and A. M. Dimits. An entropy based thermalization scheme for hybrid simulations of Coulomb collisions. Journal of Computational Physics, 273(??):77–99, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003398.

Rosin:2014:MMC

[RRD^{+}14] M. S. Rosin, L. F. Ricketson, A. M. Dimits, R. E. Caflisch, and B. I. Cohen. Multilevel Monte Carlo simulation of Coulomb collisions. Journal of Computational Physics, 274(??):140–157, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003933.

Rector:2010:SIL

[RS10a] David R. Rector and Mark L. Stewart. A semi-implicit lattice method for simulating flow. Journal of Computational Physics, 229(19):6732–6743, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002810.

Roussel:2010:CVS

[RS10b] Olivier Roussel and Kai Schneider. Coherent Vortex Simulation of weakly compressible turbulent mixing layers using adaptive multiresolution methods. Journal of Computational Physics, 229(6):2267–2286, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006639.

Rossmanith:2011:PPH

[RS11] James A. Rossmanith and David C. Seal. A positivitypreserving high-order semi-Lagrangian discontinuous Galerkin scheme for the Vlasov–Poisson equations. Journal of Computational Physics, 230(16):6203–6232, July 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002579.

Roget:2013:WDS

[RS13a] Beatrice Roget and Jayanarayanan Sitaraman. Wall distance search algorithm using voxelized marching spheres. Journal of Computational Physics, 241(??):76–94, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000880.

Russo:2013:GWP

[RS13b] Giovanni Russo and Peter Smereka. The Gaussian wave packet transform: Efficient computation of the semi-classical limit of the Schrödinger equation. Part 1: Formulation and the one dimensional case. Journal of Computational Physics, 233(??):192–209, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004640.

Roget:2014:REO

[RS14a] Beatrice Roget and Jayanarayanan Sitaraman. Robust and efficient overset grid assembly for partitioned unstructured meshes. Journal of Computational Physics, 260(??):1–24, March 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113008206.

Russo:2014:GWP

[RS14b] Giovanni Russo and Peter Smereka. The Gaussian wave packet transform: Efficient computation of the semi-classical limit of the Schrödinger equation. Part 2. Multidimensional case. Journal of Computational Physics, 257(??):1022–1038, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006281.

Rajpoot:2010:OTA

[RSD10] Manoj K. Rajpoot, Tapan K. Sengupta, and Pravir K. Dutt. Optimal time advancing dispersion relation preserving schemes. Journal of Computational Physics, 229(10):3623–3651, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000331.

Romano:2013:DDM

[RSFS13] Paul K. Romano, Andrew R. Siegel, Benoit Forget, and Kord Smith. Data decomposition of Monte Carlo particle transport simulations via tally servers. Journal of Computational Physics, 252(??):20–36, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300435X.

Racec:2013:EWE

[RSK13] Paul N. Racec, Stanley Schade, and Hans-Christoph Kaiser. Eigensolutions of the Wigner–Eisenbud problem for a cylindrical nanowire within finite volume method. Journal of Computational Physics, 252(??):52–64, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004348.

Raza:2010:ATE

[RSS10] Nauman Raza, Sultan Sial, and S. Siddiqi. Approximating time evolution related to Ginzburg–Landau functionals via Sobolev gradient methods in a finite-element setting. Journal of Computational Physics, 229(5):1621–1625, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006093.

Reboux:2012:SOL

[RSS12] Sylvain Reboux, Birte Schrader, and Ivo F. Sbalzarini. A self-organizing Lagrangian particle method for adaptiveresolution advection-diffusion simulations. Journal of Computational Physics, 231(9):3623–3646, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000472.

Reisner:2013:STS

[RSS13] J. Reisner, J. Serencsa, and S. Shkoller. A space–time smooth artificial viscosity method for nonlinear conservation laws. Journal of Computational Physics, 235(??):912–933, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004731.

Rainwater:2014:NCS

[RT14] G. Rainwater and M. Tokman. A new class of split exponential propagation iterative methods of Runge–Kutta type (sEPIRK) for semilinear systems of ODEs. Journal of Computational Physics, 269(??):40–60, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001855.

Ringler:2010:UAE

[RTKS10] T. D. Ringler, J. Thuburn, J. B. Klemp, and W. C. Skamarock. A unified approach to energy conservation and potential vorticity dynamics for arbitrarily-structured C-grids. Journal of Computational Physics, 229(9):3065–3090, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006780.

Rana:2013:RNM

[RTS13] Anirudh Rana, Manuel Torrilhon, and Henning Struchtrup. A robust numerical method for the R13 equations of rarefied gas dynamics: Application to lid driven cavity. Journal of Computational Physics, 236(??):169–186, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112007036.

Radhakrishnan:2013:TMA

[RUL^{+}13] R. Radhakrishnan, B. Uma, J. Liu, P. S. Ayyaswamy, and D. M. Eckmann. Temporal multiscale approach for nanocarrier motion with simultaneous adhesion and hydrodynamic interactions in targeted drug delivery. Journal of Computational Physics, 244(??):252–263, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006274.

Rapun:2010:ROM

[RV10] María-Luisa Rapún and José M. Vega. Reduced order models based on local POD plus Galerkin projection. Journal of Computational Physics, 229(8):3046–3063, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109007153.

Rahimian:2010:DSL

[RVB10] Abtin Rahimian, Shravan Kumar Veerapaneni, and George Biros. Dynamic simulation of locally inextensible vesicles suspended in an arbitrary two-dimensional domain, a boundary integral method. Journal of Computational Physics, 229(18):6466–6484, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002585.

Reps:2010:IHE

[RVbZ10] Bram Reps, Wim Vanroose, and Hisham bin Zubair. On the indefinite Helmholtz equation: Complex stretched absorbing boundary layers, iterative analysis, and preconditioning. Journal of Computational Physics, 229(22):8384–8405, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004158.

Reckinger:2014:AWC

[RVFK14] Shanon M. Reckinger, Oleg V. Vasilyev, and Baylor Fox-Kemper. Adaptive wavelet collocation method on the shallow water model. Journal of Computational Physics, 271 (??):342–359, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114002307.

Rycroft:2013:CTD

[RW13] Chris H. Rycroft and Jon Wilkening. Computation of threedimensional standing water waves. Journal of Computational Physics, 255(??):612–638, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005640.

Rosa:2011:AEM

[RWMG11] B. Rosa, L.-P. Wang, M. R. Maxey, and W. W. Grabowski. An accurate and efficient method for treating aerodynamic interactions of cloud droplets. Journal of Computational Physics, 230(22):8109–8133, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004311.

Ren:2011:DAW

[RWX11] Xiaoan Ren, Wenquan Wu, and Leonidas S. Xanthis. A dynamically adaptive wavelet approach to stochastic computations based on polynomial chaos — capturing all scales of random modes on independent grids. Journal of Computational Physics, 230(19):7332–7346, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003500.

Ren:2014:DAC

[RXLS14] Zhuyin Ren, Chao Xu, Tianfeng Lu, and Michael A. Singer. Dynamic adaptive chemistry with operator splitting schemes for reactive flow simulations. Journal of Computational Physics, 263(??):19–36, April 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000436.

Rawat:2010:HOS

[RZ10] Pradeep Singh Rawat and Xiaolin Zhong. On high-order shock-fitting and front-tracking schemes for numerical simulation of shock-disturbance interactions. Journal of Computational Physics, 229(19):6744–6780, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002822.

Ren:2013:CDS

[RzSZ13] Jincheng Ren, Zhi zhong Sun, and Xuan Zhao. Compact difference scheme for the fractional sub-diffusion equation with Neumann boundary conditions. Journal of Computational Physics, 232(1):456–467, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200472X.

Stehle:2014:HTD

[SAA14] Nicholas D. Stehle, Dmitriy Y. Anistratov, and Marvin L. Adams. A hybrid transport-diffusion method for 2D transport problems with diffusive subdomains. Journal of Computational Physics, 270(??):325–344, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002435.

Shahbazi:2011:MDF

[SABH11] Khosro Shahbazi, Nathan Albin, Oscar P. Bruno, and Jan S. Hesthaven. Multi-domain Fourier-continuation/WENO hybrid solver for conservation laws. Journal of Computational Physics, 230(24):8779–8796, October 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005110.

Shang:2014:CAH

[SAHS14] J. S. Shang, D. A. Andrienko, P. G. Huang, and S. T. Surzhikov. A computational approach for hypersonic nonequilibrium radiation utilizing space partition algorithm and Gauss quadrature. Journal of Computational Physics, 266(??):1–21, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400103X.

Salman:2014:TSP

[Sal14] Hayder Salman. A time-splitting pseudospectral method for the solution of the Gross–Pitaevskii equations using spherical harmonics with generalised-Laguerre basis functions. Journal of Computational Physics, 258(??):185–207, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006773.

Sankaran:2010:MSC

[SAM10] Sethuraman Sankaran, Charles Audet, and Alison L. Marsden. A method for stochastic constrained optimization using derivative-free surrogate pattern search and collocation. Journal of Computational Physics, 229(12):4664–4682, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001154.

Samuel:2014:LST

[Sam14] Henri Samuel. A level set two-way wave equation approach for Eulerian interface tracking. Journal of Computational Physics, 259(??):617–635, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007894.

Sanderse:2013:ECR

[San13] B. Sanderse. Energy-conserving Runge–Kutta methods for the incompressible Navier–Stokes equations. Journal of Computational Physics, 233(??):100–131, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200424X.

Sorek:2012:MEL

[SB12] Shaul Sorek and Vyacheslav Borisov. Modified Eulerian– Lagrangian formulation for hydrodynamic modeling. Journal of Computational Physics, 231(8):3083–3100, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007030.

Sengupta:2014:EDD

[SB14] Tapan K. Sengupta and Ashish Bhole. Error dynamics of diffusion equation: Effects of numerical diffusion and dispersive diffusion. Journal of Computational Physics, 266(??):240–251, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001417.

Sigurdsson:2013:HCP

[SBA13] Jon K. Sigurdsson, Frank L. H. Brown, and Paul J. Atzberger. Hybrid continuum-particle method for fluctuating lipid bilayer membranes with diffusing protein inclusions. Journal of Computational Physics, 252(??):65–85, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004403.

Slaba:2010:FMA

[SBB10] T. C. Slaba, S. R. Blattnig, and F. F. Badavi. Faster and more accurate transport procedures for HZETRN. Journal of Computational Physics, 229(24):9397–9417, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005048.

Salvadore:2013:GAF

[SBB13] Francesco Salvadore, Matteo Bernardini, and Michela Botti. GPU accelerated flow solver for direct numerical simulation of turbulent flows. Journal of Computational Physics, 235(??):129–142, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112006018.

Sagert:2014:HSW

[SBC^{+}14] Irina Sagert, Wolfgang Bauer, Dirk Colbry, Jim Howell, Rodney Pickett, Alec Staber, and Terrance Strother. Hydrodynamic shock wave studies within a kinetic Monte Carlo approach. Journal of Computational Physics, 266(??):191–213, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001399.

Sandve:2012:EMP

[SBN12] T. H. Sandve, I. Berre, and J. M. Nordbotten. An efficient multi-point flux approximation method for Discrete Fracture-Matrix simulations. Journal of Computational Physics, 231(9):3784–3800, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000447.

Suryanarayana:2011:MFC

[SBO11] Phanish Suryanarayana, Kaushik Bhattacharya, and Michael Ortiz. A mesh-free convex approximation scheme for Kohn– Sham density functional theory. Journal of Computational Physics, 230(13):5226–5238, June 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001616.

Sengupta:2011:NCD

[SBS11] T. K. Sengupta, S. Bhaumik, and U. Shameem. A new compact difference scheme for second derivative in non-uniform grid expressed in self-adjoint form. Journal of Computational Physics, 230(5):1822–1848, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006509.

Sushko:2013:SUR

[SBS^{+}13] Gennady B. Sushko, Victor G. Bezchastnov, Ilia A. Solov’yov, Andrei V. Korol, Walter Greiner, and Andrey V. Solov’yov. Simulation of ultra-relativistic electrons and positrons channeling in crystals with MBN Explorer. Journal of Computational Physics, 252(??):404–418, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004580.

Slaba:2013:RDE

[SBT13] Tony C. Slaba, Steve R. Blattnig, and John Tweed. Reduced discretization error in HZETRN. Journal of Computational Physics, 234(??):217–229, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005876.

Salas-Boni:2014:PCD

[SBT14]  Rebeca Salas-Boni and Esteban G. Tabak.  Principal

components: a descent algorithm. Journal of Computational Physics, 267(??):162–175, June 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001594.

Sollie:2011:STD

[SBvdV11] W. E. H. Sollie, O. Bokhove, and J. J. W. van der Vegt. Space– time discontinuous Galerkin finite element method for twofluid flows. Journal of Computational Physics, 230(3):789–817, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005759.

Sabo:2014:FET

[SBWF14] Doron Sabo, Oded Barzelay, Shlomo Weiss, and Miriam Furst. Fast evaluation of a time-domain non-linear cochlear model on GPUs. Journal of Computational Physics, 265(??):97–112, May 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000898.

Sijoy:2010:VFA

[SC10] C. D. Sijoy and Shashank Chaturvedi. Volume-of-fluid algorithm with different modified dynamic material ordering methods and their comparisons. Journal of Computational Physics, 229(10):3848–3863, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000549.

Simeonov:2011:PBI

[SC11] Julian A. Simeonov and Joseph Calantoni. A pressure boundary integral method for direct fluid-particle simulations on Cartesian grids. Journal of Computational Physics, 230(5):1749–1765, March 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000642X.

Steinhoff:2012:SSW

[SC12] John Steinhoff and Subhashini Chitta. Solution of the scalar wave equation over very long distances using nonlinear solitary waves: Relation to finite difference methods. Journal of Computational Physics, 231(19):6306–6322, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200232X.

Smolarkiewicz:2013:ECM

[SC13] Piotr K. Smolarkiewicz and Paul Charbonneau. EULAG, a computational model for multiscale flows: an MHD extension. Journal of Computational Physics, 236(??):608–623, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006845.

Smolentsev:2010:IEC

[SCB10] S. Smolentsev, S. Cuevas, and A. Beltrán. Induced electric current-based formulation in computations of low magnetic Reynolds number magnetohydrodynamic flows. Journal of Computational Physics, 229(5):1558–1572, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006056.

Shu:2014:AGA

[SCC14] Yu-Chen Shu, I-Liang Chern, and Chien C. Chang. Accurate gradient approximation for complex interface problems in 3D by an improved coupling interface method. Journal of Computational Physics, 275(??):642–661, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005105.

Servan-Camas:2013:AMB

[SCGE13] Borja Serván-Camas and Julio García-Espinosa. Accelerated 3D multi-body seakeeping simulations using unstructured finite elements. Journal of Computational Physics, 252(??):382–403, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004476.

Schram:2013:RDM

[Sch13a] R. D. Schram. Reaction-diffusion model Monte Carlo simulations on the GPU. Journal of Computational Physics, 241(??):95–103, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113000946.

Sidler:2013:PSM

[SCH13b] Rolf Sidler, José M. Carcione, and Klaus Holliger. A pseudospectral method for the simulation of poro-elastic seismic wave propagation in 2D polar coordinates using domain decomposition. Journal of Computational Physics, 235(??):846–864, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005918.

Schiavazzi:2014:MPA

[SCIE14] Daniele Schiavazzi, Filippo Coletti, Gianluca Iaccarino, and John K. Eaton. A matching pursuit approach to solenoidal filtering of three-dimensional velocity measurements. Journal of Computational Physics, 263(??):206–221, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113008577.

Scovazzi:2012:LSH

[Sco12] G. Scovazzi. Lagrangian shock hydrodynamics on tetrahedral meshes: a stable and accurate variational multiscale approach. Journal of Computational Physics, 231(24):8029–8069, October 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003488.

Spinneken:2014:FCA

[SCS14] Johannes Spinneken, Marios Christou, and Chris Swan. Forcecontrolled absorption in a fully-nonlinear numerical wave tank. Journal of Computational Physics, 272(??):127–148, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002812.

Sheu:2011:DHO

[SCY11] Tony W. H. Sheu, P. H. Chiu, and C. H. Yu. On the development of a high-order compact scheme for exhibiting the switching and dissipative solution natures in the Camassa–Holm equation. Journal of Computational Physics, 230(13):5399–5416, June 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100204X.

Sellier:2014:MBW

[SD14a] J. M. Sellier and I. Dimov. The many-body Wigner Monte Carlo method for time-dependent ab-initio quantum simulations. Journal of Computational Physics, 273(??):589– 597, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114004021.

Sellier:2014:WMC

[SD14b] J. M. Sellier and I. Dimov. A Wigner Monte Carlo approach to density functional theory. Journal of Computational Physics, 270(??):265–277, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002526.

Sun:2014:ASC

[SD14c] Huafei Sun and David L. Darmofal. An adaptive simplex cutcell method for high-order discontinuous Galerkin discretizations of elliptic interface problems and conjugate heat transfer problems. Journal of Computational Physics, 278(??):445–468, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006032.

Sudhakar:2014:ARE

[SdAW14] Y. Sudhakar, J. P. Moitinho de Almeida, and Wolfgang A. Wall. An accurate, robust, and easy-to-implement method for integration over arbitrary polyhedra: Application to embedded interface methods. Journal of Computational Physics, 273(??):393–415, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003659.

Sun:2012:ITS

[SDH12] Huafei Sun, David L. Darmofal, and Robert Haimes. On the impact of triangle shapes for boundary layer problems using high-order finite element discretization. Journal of Computational Physics, 231(2):541–557, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100550X.

Syrakos:2012:NEE

[SEBG12] Alexandros Syrakos, Georgios Efthimiou, John G. Bartzis, and Apostolos Goulas. Numerical experiments on the efficiency of local grid refinement based on truncation error estimates. Journal of Computational Physics, 231(20):6725–6753, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003385.

Slaybaugh:2013:MEP

[SEDW13] R. N. Slaybaugh, T. M. Evans, G. G. Davidson, and P. P. H. Wilson. Multigrid in energy preconditioner for Krylov solvers. Journal of Computational Physics, 242(??):405–419, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001228.

Sen:2013:NFC

[Sen13] Shuvam Sen. A new family of (5,5)CC–4OC schemes applicable for unsteady Navier–Stokes equations. Journal of Computational Physics, 251(??):251–271, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004142.

Sauerland:2011:EFE

[SF11] Henning Sauerland and Thomas-Peter Fries. The extended finite element method for two-phase and free-surface flows: a systematic study. Journal of Computational Physics, 230(9):3369– 3390, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000581.

Shin:2012:CPM

[SF12] Wonseok Shin and Shanhui Fan. Choice of the perfectly matched layer boundary condition for frequencydomain Maxwell’s equations solvers. Journal of Computational Physics, 231(8):3406–3431, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000344.

Shao:2014:HPC

[SF14] Yan-Lin Shao and Odd M. Faltinsen. A harmonic polynomial cell (HPC) method for 3D Laplace equation with application in marine hydrodynamics. Journal of Computational Physics, 274(??):312–332, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004288.

Shetty:2010:HOI

[SFCF10] Dinesh A. Shetty, Travis C. Fisher, Aditya R. Chunekar, and Steven H. Frankel. High-order incompressible large-eddy simulation of fully inhomogeneous turbulent flows. Journal of Computational Physics, 229(23):8802–8822, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004481.

Sitaraman:2010:PDC

[SFWP10] Jayanarayanan Sitaraman, Matthew Floros, Andrew Wissink, and Mark Potsdam. Parallel domain connectivity algorithm for unsteady flow computations using overlapping and adaptive grids. Journal of Computational Physics, 229(12):4703–4723, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000118X.

Savadatti:2010:ABCb

[SG10a] Siddharth Savadatti and Murthy N. Guddati. Absorbing boundary conditions for scalar waves in anisotropic media. Part 1: Time harmonic modeling. Journal of Computational Physics, 229(19):6696–6714, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002792.

Savadatti:2010:ABCa

[SG10b] Siddharth Savadatti and Murthy N. Guddati. Absorbing boundary conditions for scalar waves in anisotropic media. Part 2: Time-dependent modeling. Journal of Computational Physics, 229(18):6644–6662, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002780.

Savadatti:2012:AABa

[SG12a] Siddharth Savadatti and Murthy N. Guddati. Accurate absorbing boundary conditions for anisotropic elastic media. Part 1: Elliptic anisotropy. Journal of Computational Physics, 231(22):7584–7607, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003026.

Savadatti:2012:AABb

[SG12b] Siddharth Savadatti and Murthy N. Guddati. Accurate absorbing boundary conditions for anisotropic elastic media. Part 2: Untilted non-elliptic anisotropy. Journal of Computational Physics, 231(22):7608–7625, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003142.

Strand:2013:GSA

[SG13] James S. Strand and David B. Goldstein. Global sensitivity analysis for DSMC simulations of hypersonic shocks. Journal of Computational Physics, 246(??):184–206, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002131.

Shukla:2014:IFV

[SG14] Ratnesh K. Shukla and Pritam Giri. Isotropic finite volume discretization. Journal of Computational Physics, 276 (??):252–290, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911400518X.

Scovazzi:2013:DGM

[SGC13] G. Scovazzi, A. Gerstenberger, and S. S. Collis. A discontinuous Galerkin method for gravity-driven viscous fingering instabilities in porous media. Journal of Computational Physics, 233(??):373–399, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005220.

Sreekanth:2014:ALC

[SGL14] R. Sreekanth, J. B. Greenberg, and Y. Levy. An approximate linearized characteristic Riemann solver based on blending of Riemann invariants. Journal of Computational Physics, 278(??):469–484, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006196.

Schmid:2013:HOF

[SGS13] K. S. Schmid, S. Geiger, and K. S. Sorbie. Higher order FE-FV method on unstructured grids for transport and twophase flow with variable viscosity in heterogeneous porous media. Journal of Computational Physics, 241(??):416–444, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007486.

Stephani:2013:NES

[SGV13] K. A. Stephani, D. B. Goldstein, and P. L. Varghese. A non-equilibrium surface reservoir approach for hybrid DSMC/ Navier–Stokes particle generation. Journal of Computational Physics, 232(1):468–481, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004639.

Schnepp:2010:HFI

[SGW10] S. Schnepp, E. Gjonaj, and T. Weiland. A hybrid Finite Integration–Finite Volume Scheme. Journal of Computational Physics, 229(11):4075–4096, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000720.

Sharma:2011:FSI

[SH11] Prateek Sharma and Gregory W. Hammett. A fast semiimplicit method for anisotropic diffusion. Journal of Computational Physics, 230(12):4899–4909, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001525.

Schoenawa:2014:DGD

[SH14] Stefan Schoenawa and Ralf Hartmann. Discontinuous Galerkin discretization of the Reynolds-averaged Navier–Stokes equations with the shear-stress transport model. Journal of Computational Physics, 262(??):194–216, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000242.

So:2011:ADM

[SHA11] K. K. So, X. Y. Hu, and N. A. Adams. Anti-diffusion method for interface steepening in two-phase incompressible flow. Journal of Computational Physics, 230(13):5155–5177, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001549.

So:2012:ADI

[SHA12] K. K. So, X. Y. Hu, and N. A. Adams. Anti-diffusion interface sharpening technique for two-phase compressible flow simulations. Journal of Computational Physics, 231(11):4304–4323, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000927.

Shang:2013:TLS

[Sha13] Yueqiang Shang. A two-level subgrid stabilized Oseen iterative method for the steady Navier–Stokes equations. Journal of Computational Physics, 233(??):210–226, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004706.

Scovazzi:2013:FCU

[SHCY13] G. Scovazzi, H. Huang, S. S. Collis, and J. Yin. A fullycoupled upwind discontinuous Galerkin method for incompressible porous media flows: High-order computations of viscous fingering instabilities in complex geometry. Journal of Computational Physics, 252(??):86–108, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004361.

Shestakov:2013:MRD

[She13] Aleksei I. Shestakov. Multifrequency radiation diffusion equations for homogeneous, refractive, lossy media and their interface conditions. Journal of Computational Physics, 243 (??):293–304, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113001940.

Sengupta:2012:EAC

[SHFB12] Ankush Sengupta, Paul S. Hammond, Daan Frenkel, and Edo S. Boek. Error analysis and correction for Lattice Boltzmann simulated flow conductance in capillaries of different shapes and alignments. Journal of Computational Physics, 231(6):2634–2640, March 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007029.

Shunn:2012:VVD

[SHM12] Lee Shunn, Frank Ham, and Parviz Moin. Verification of variable-density flow solvers using manufactured solutions. Journal of Computational Physics, 231(9):3801–3827, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000484.

Schneiders:2013:AMB

[SHMS13] Lennart Schneiders, Daniel Hartmann, Matthias Meinke, and Wolfgang Schröder. An accurate moving boundary formulation in cut-cell methods. Journal of Computational Physics, 235(??):786–809, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005839.

Sarmany:2013:USS

[SHR13] D. Sármány, M. E. Hubbard, and M. Ricchiuto. Unconditionally stable space–time discontinuous residual distribution for shallow-water flows. Journal of Computational Physics, 253(??):86–113, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004816.

Shukla:2014:NPE

[Shu14] Ratnesh K. Shukla. Nonlinear preconditioning for efficient and accurate interface capturing in simulation of multicomponent compressible flows. Journal of Computational Physics, 276(??):508–540, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005270.

Shyue:2010:HRM

[Shy10] K.-M. Shyue. A high-resolution mapped grid algorithm for compressible multiphase flow problems. Journal of Computational Physics, 229(23):8780–8801, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000447X.

Shahbazi:2013:MDH

[SHZ13] Khosro Shahbazi, Jan S. Hesthaven, and Xueyu Zhu. Multidimensional hybrid Fourier continuation-WENO solvers for conservation laws. Journal of Computational Physics, 253 (??):209–225, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113004828.

Sogabe:2012:SGS

[SHZF12] Tomohiro Sogabe, Takeo Hoshi, Shao-Liang Zhang, and Takeo Fujiwara. Solution of generalized shifted linear systems with complex symmetric matrices. Journal of Computational Physics, 231(17):5669–5684, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002501.

Susanto:2013:HOC

[SIDG13] A. Susanto, L. Ivan, H. De Sterck, and C. P. T. Groth. High-order central ENO finite-volume scheme for ideal MHD. Journal of Computational Physics, 250(??):141–164, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003343.

Sugiyama:2011:FEF

[SIT^{+}11] Kazuyasu Sugiyama, Satoshi Ii, Shintaro Takeuchi, Shu Takagi, and Yoichiro Matsumoto. A full Eulerian finite difference approach for solving fluid-structure coupling problems. Journal of Computational Physics, 230(3):596–627, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000553X.

Stafford:2010:ULS

[SJ10] D. Shane Stafford and Thomas L. Jackson. Using level sets for creating virtual random packs of non-spherical convex shapes. Journal of Computational Physics, 229(9):3295–3315, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000185.

Schnabel:2011:AMM

[SJB11] Stefan Schnabel, Wolfhard Janke, and Michael Bachmann. Advanced multicanonical Monte Carlo methods for efficient simulations of nucleation processes of polymers. Journal of Computational Physics, 230(12):4454–4465, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001070.

Spencer:2012:FEF

[SJH12] J. Andrew Spencer, Jeong-Young Ji, and Eric D. Held. A finite element/Fourier treatment of the Fokker–Planck equation. Journal of Computational Physics, 231(18):6192–6206, July 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003075.

Shin:2011:CNM

[SJK11] Jaemin Shin, Darae Jeong, and Junseok Kim. A conservative numerical method for the Cahn–Hilliard equation in complex domains. Journal of Computational Physics, 230 (19):7441–7455, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111003585.

Sanderse:2012:AAE

[SK12a] B. Sanderse and B. Koren. Accuracy analysis of explicit Runge–Kutta methods applied to the incompressible Navier–Stokes equations. Journal of Computational Physics, 231(8):3041–3063, April 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111006838.

Solin:2012:AHO

[SK12b] Pavel Solin and Lukas Korous. Adaptive higher-order finite element methods for transient PDE problems based on embedded higher-order implicit Runge–Kutta methods. Journal of Computational Physics, 231(4):1635–1649, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006231.

Semiletov:2013:CSC

[SK13] V. A. Semiletov and S. A. Karabasov. CABARET scheme with conservation-flux asynchronous time-stepping for nonlinear aeroacoustics problems. Journal of Computational Physics, 253(??):157–165, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113004804.

Shu:2010:ACI

[SKCC10] Yu-Chen Shu, Chiu-Yen Kao, I-Liang Chern, and Chien C. Chang. Augmented coupling interface method for solving eigenvalue problems with sign-changed coefficients. Journal of Computational Physics, 229(24):9246–9268, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004936.

Sandvin:2013:AVM

[SKN13a] Andreas Sandvin, Eirik Keilegavlen, and Jan M. Nordbotten. Auxiliary variables for 3D multiscale simulations in heterogeneous porous media. Journal of Computational Physics, 238(??):141–153, April 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007474.

Skogestad:2013:DDS

[SKN13b] Jan Ole Skogestad, Eirik Keilegavlen, and Jan M. Nordbotten. Domain decomposition strategies for nonlinear flow problems in porous media. Journal of Computational Physics, 234(??):439–451, February 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200589X.

Starinshak:2014:NLS

[SKR14] David P. Starinshak, Smadar Karni, and Philip L. Roe. A new level set model for multimaterial flows. Journal of Computational Physics, 262(??):1–16, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008449.

Sovinec:2010:AMS

[SKt10] C. R. Sovinec, J. R. King, and the NIMROD Team. Analysis of a mixed semi-implicit/implicit algorithm for lowfrequency two-fluid plasma modeling. Journal of Computational Physics, 229(16):5803–5819, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001981.

Sambasivan:2013:SHS

[SKU13] S. Sambasivan, A. Kapahi, and H. S. Udaykumar. Simulation of high speed impact, penetration and fragmentation problems on locally refined Cartesian grids. Journal of Computational Physics, 235(??):334–370, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006341.

Smolarkiewicz:2014:CFD

[SKW14] Piotr K. Smolarkiewicz, Christian Kühnlein, and Nils P. Wedi. A consistent framework for discrete integrations of soundproof and compressible PDEs of atmospheric dynamics. Journal of Computational Physics, 263(??):185–205, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000588.

Schauer:2013:AEK

[SL13a] Volker Schauer and Christian Linder. All-electron Kohn– Sham density functional theory on hierarchic finite element spaces. Journal of Computational Physics, 250(??):644–664, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002696.

Suh:2013:NMC

[SL13b] Youngho Suh and Changhoon Lee. A numerical method for the calculation of drag and lift of a deformable droplet in shear flow. Journal of Computational Physics, 241(??):35–57, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000879.

Sohn:2011:FEA

[SLC^{+}11] Dongwoo Sohn, Jae Hyuk Lim, Young-Sam Cho, Jeong Ho Kim, and Seyoung Im. Finite element analysis of quasistatic crack propagation in brittle media with voids or inclusions. Journal of Computational Physics, 230(17):6866–6899, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003202.

Serre:2012:RRO

[SLGB12] Gilles Serre, Philippe Lafon, Xavier Gloerfelt, and Christophe Bailly. Reliable reduced-order models for time-dependent linearized Euler equations. Journal of Computational Physics, 231(15):5176–5194, June 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002008.

Smith:2011:DIS

[SLH^{+}11] Matthew R. Smith, K.-M. Lin, C.-T. Hung, Y.-S. Chen, and J.-S. Wu. Development of an improved spatial reconstruction technique for the HLL method and its applications. Journal of Computational Physics, 230(3):477–493, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005267.

Shashank:2010:CLI

[SLI10] Shashank, Johan Larsson, and Gianluca Iaccarino. A colocated incompressible Navier–Stokes solver with exact mass, momentum and kinetic energy conservation in the inviscid limit. Journal of Computational Physics, 229(12):4425–4430, June 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001208.

Seny:2014:EPI

[SLT^{+}14] Bruno Seny, Jonathan Lambrechts, Thomas Toulorge, Vincent Legat, and Jean-François Remacle. An efficient parallel implementation of explicit multirate Runge–Kutta schemes for discontinuous Galerkin computations. Journal of Computational Physics, 256(??):135–160, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005408.

Li:2010:MIM

[sLwG10] Xue song Li and Chun wei Gu. The momentum interpolation method based on the time-marching algorithm for All-Speed flows. Journal of Computational Physics, 229(20):7806–7818, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003621.

Srinath:2010:AMS

[SM10] D. N. Srinath and Sanjay Mittal. An adjoint method for shape optimization in unsteady viscous flows. Journal of Computational Physics, 229(6):1994–2008, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900638X.

Salac:2011:LSP

[SM11a] D. Salac and M. Miksis. A level set projection model of lipid vesicles in general flows. Journal of Computational Physics, 230(22):8192–8215, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004384.

Seo:2011:HOI

[SM11b] Jung Hee Seo and Rajat Mittal. A high-order immersed boundary method for acoustic wave scattering and low-Mach number flow-induced sound in complex geometries. Journal of Computational Physics, 230(4):1000–1019, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005711.

Seo:2011:SII

[SM11c] Jung Hee Seo and Rajat Mittal. A sharp-interface immersed boundary method with improved mass conservation and reduced spurious pressure oscillations. Journal of Computational Physics, 230(19):7347–7363, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003524.

Schaerf:2012:CCCa

[SM12a] T. M. Schaerf and C. Macaskill. On contour crossings in contour-advective simulations — part 1 — algorithm for detection and quantification. Journal of Computational Physics, 231(2):465–480, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100547X.

Schaerf:2012:CCCb

[SM12b] T. M. Schaerf and C. Macaskill. On contour crossings in contour-advective simulations — part 2 — analysis of crossing errors and methods for their prevention. Journal of Computational Physics, 231(2):481–504, January 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005456.

Sriram:2012:IMM

[SM12c] V. Sriram and Q. W. Ma. Improved MLPG R method for simulating 2D interaction between violent waves and elastic structures. Journal of Computational Physics, 231(22): 7650–7670, September 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003610.

Schutz:2013:HMM

[SM13] Jochen Schütz and Georg May. A hybrid mixed method for the compressible Navier–Stokes equations. Journal of Computational Physics, 240(??):58–75, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000508.

Scheff:2013:MMA

[SMF^{+}13] Jeremy D. Scheff, Panteleimon D. Mavroudis, Panagiota T. Foteinou, Gary An, Steve E. Calvano, John Doyle, Thomas E. Dick, Stephen F. Lowry, Yoram Vodovotz, and Ioannis P. Androulakis. A multiscale modeling approach to inflammation: a case study in human endotoxemia. Journal of Computational Physics, 244(??):279–289, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005554.

Sheu:2013:DWP

[SML13] Tony W. H. Sheu, C. M. Mei, and J. H. Li. Development of a wavenumber-preserving scheme for solving Maxwell’s equations in curvilinear non-staggered grids. Journal of Computational Physics, 247(??):208–227, August 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002465.

Scarbolo:2013:UFS

[SMP^{+}13] Luca Scarbolo, Dafne Molin, Prasad Perlekar, Mauro Sbragaglia, Alfredo Soldati, and Federico Toschi. Unified framework for a side-by-side comparison of different multicomponent algorithms: Lattice Boltzmann vs. phase field model. Journal of Computational Physics, 234(??):263–279, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005748.

Sriram:2014:HMM

[SMS14] V. Sriram, Q. W. Ma, and T. Schlurmann. A hybrid method for modelling two dimensional non-breaking and breaking waves. Journal of Computational Physics, 272(??):429–454, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002939.

Sato:2012:NCL

[SN12] Yohei Sato and Bojan Niceno. A new contact line treatment for a conservative level set method. Journal of Computational Physics, 231(10):3887–3895, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000642.

Sato:2013:SIP

[SN13a] Yohei Sato and Bojan Niceno. A sharp-interface phase change model for a mass-conservative interface tracking method. Journal of Computational Physics, 249(??):127–161, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003197.

Stefanescu:2013:PDN

[SN13b] R. Stefanescu and I. M. Navon. POD/DEIM nonlinear model order reduction of an ADI implicit shallow water equations model. Journal of Computational Physics, 237(??):95–114, March 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007152.

Svard:2014:RSP

[SN14] Magnus Svärd and Jan Nordström. Review of summationby-parts schemes for initial-boundary-value problems. Journal of Computational Physics, 268(??):17–38, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400151X.

Schoch:2013:EAC

[SNBN13] Stefan Schoch, Kevin Nordin-Bates, and Nikolaos Nikiforakis. An Eulerian algorithm for coupled simulations of elastoplasticsolids and condensed-phase explosives. Journal of Computational Physics, 252(??):163–194, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004440.

Samaddar:2010:PTN

[SNS10] D. Samaddar, D. E. Newman, and R. Sánchez. Parallelization in time of numerical simulations of fully-developed plasma turbulence using the parareal algorithm. Journal of Computational Physics, 229(18):6558–6573, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002731.

Sho:2013:QET

[SO13] Shohiro Sho and Shinji Odanaka. A quantum energy transport model for semiconductor device simulation. Journal of Computational Physics, 235(??):486–496, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006559.

Soares:2013:TDE

[Soa13] Delfim Soares, Jr. Time-domain electromagnetic wave propagation analysis by edge-based smoothed point interpolation methods. Journal of Computational Physics, 234(??):472–486, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005980.

Sabelnikov:2014:MLS

[SOG14a] Vladimir Sabelnikov, Andrey Yu. Ovsyannikov, and Mikhael Gorokhovski. Modified level set equation and its numerical assessment. Journal of Computational Physics, 278(??):1–30, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005725.

Soghrati:2014:HIE

[Sog14b] Soheil Soghrati. Hierarchical interface-enriched finite element method: an automated technique for mesh-independent simulations. Journal of Computational Physics, 275(??):41–52, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004239.

Son:2011:VCF

[Son11] Sang-Kil Son. Voronoi-cell finite difference method for accurate electronic structure calculation of polyatomic molecules on unstructured grids. Journal of Computational Physics, 230(5):2160–2173, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110006765.

Sousa:2014:EHO

[Sou14] Ercília Sousa. An explicit high order method for fractional advection diffusion equations. Journal of Computational Physics, 278(??):257–274, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006044.

Stupfel:2011:SUC

[SP11] Bruno Stupfel and Dorcas Poget. Sufficient uniqueness conditions for the solution of the time harmonic Maxwell’s equations associated with surface impedance boundary conditions. Journal of Computational Physics, 230(12):4571–4587, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001215.

Sahai:2013:UQH

[SP13] Tuhin Sahai and José Miguel Pasini. Uncertainty quantification in hybrid dynamical systems. Journal of Computational Physics, 237(??):411–427, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006316.

Shamir:2014:HBS

[SP14a] Ofer Shamir and Nathan Paldor. A Hermite-based Shallow Water solver for a thin “ocean” over a rotating sphere. Journal of Computational Physics, 269(??):80–97, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001880.

Suryanarayana:2014:ALF

[SP14b] Phanish Suryanarayana and Deepa Phanish. Augmented Lagrangian formulation of orbital-free density functional theory. Journal of Computational Physics, 275(??):524–538, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004860.

Shadid:2010:TSF

[SPB^{+}10] J. N. Shadid, R. P. Pawlowski, J. W. Banks, L. Chacón, P. T. Lin, and R. S. Tuminaro. Towards a scalable fullyimplicit fully-coupled resistive MHD formulation with stabilized FE methods. Journal of Computational Physics, 229 (20):7649–7671, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003335.

Shukla:2010:ICM

[SPF10] Ratnesh K. Shukla, Carlos Pantano, and Jonathan B. Freund. An interface capturing method for the simulation of multi-phase compressible flows. Journal of Computational Physics, 229(19):7411–7439, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003402.

Stevens:2013:ALC

[SPM^{+}13] D. Stevens, H. Power, C. Y. Meng, D. Howard, and K. A. Cliffe. An alternative local collocation strategy for highconvergence meshless PDE solutions, using radial basis functions. Journal of Computational Physics, 254(??):52–75, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005081.

Schmicker:2014:IGM

[SPS14] D. Schmicker, P.-O. Persson, and J. Strackeljan. Implicit Geometry Meshing for the simulation of Rotary Friction Welding. Journal of Computational Physics, 270(??):478–489, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002770.

Shen:2011:AMR

[SQC11] Chaopeng Shen, Jing-Mei Qiu, and Andrew Christlieb. Adaptive mesh refinement based on high order finite difference WENO scheme for multi-scale simulations. Journal of Computational Physics, 230(10):3780–3802, May 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000908.

Shirokoff:2011:EMI

[SR11] D. Shirokoff and R. R. Rosales. An efficient method for the incompressible Navier–Stokes equations on irregular domains with no-slip boundary conditions, high order up to the boundary. Journal of Computational Physics, 230(23):8619–8646, September 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004839.

Schwarz:2013:EKM

[SR13a] Karsten Schwarz and Heiko Rieger. Efficient kinetic Monte Carlo method for reaction-diffusion problems with spatially varying annihilation rates. Journal of Computational Physics, 237(??):396–410, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007164.

Sharma:2013:ASB

[SR13b] Gulshan B. Sharma and Douglas D. Robertson. Adaptive scapula bone remodeling computational simulation: Relevance to regenerative medicine. Journal of Computational Physics, 244(??):312–320, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112005736.

Sitaraman:2013:MFI

[SR13c] H. Sitaraman and L. L. Raja. A matrix free implicit scheme for solution of resistive magneto-hydrodynamics equations on unstructured grids. Journal of Computational Physics, 251(??):364–382, October 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004178.

Stuck:2013:ACV

[SR13d] Arthur Stück and Thomas Rung. Adjoint complement to viscous finite-volume pressure-correction methods. Journal of Computational Physics, 248(??):402–419, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000260.

Sun:2011:CFD

[SRL^{+}11] Zhen-Sheng Sun, Yu-Xin Ren, Cédric Larricq, Shi ying Zhang, and Yue cheng Yang. A class of finite difference schemes with low dispersion and controllable dissipation for DNS of compressible turbulence. Journal of Computational Physics, 230(12):4616–4635, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001276.

Safta:2010:HOL

[SRN10] Cosmin Safta, Jaideep Ray, and Habib N. Najm. A high-order low-Mach number AMR construction for chemically reacting flows. Journal of Computational Physics, 229(24):9299–9322, December 10, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004961.

Stracquadanio:2013:SDD

[SRN13] Giovanni Stracquadanio, Vittorio Romano, and Giuseppe Nicosia. Semiconductor device design using the Bi MADS algorithm. Journal of Computational Physics, 242(??):304–320, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000569.

Schrader:2010:DCG

[SRS10] Birte Schrader, Sylvain Reboux, and Ivo F. Sbalzarini. Discretization correction of general integral PSE Operators for particle methods. Journal of Computational Physics, 229(11):4159–4182, June 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000077X.

Soucasse:2014:SSM

[SRS14] L. Soucasse, Ph. Rivière, and A. Soufiani. Subgrid-scale model for radiative transfer in turbulent participating media. Journal of Computational Physics, 257(??):442–459, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006748.

Sengupta:2011:AAN

[SRSV11] Tapan K. Sengupta, Manoj K. Rajpoot, Shakti Saurabh, and V. V. S. N. Vijay. Analysis of anisotropy of numerical wave solutions by high accuracy finite difference methods. Journal of Computational Physics, 230(1):27–60, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004973.

Sani:2010:LIS

[SS10a] M. Sani and M. S. Saidi. A lagged implicit segregated data reconstruction procedure to treat open boundaries. Journal of Computational Physics, 229(14):5418–5431, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001816.

Sekora:2010:HGM

[SS10b] Michael D. Sekora and James M. Stone. A hybrid Godunov method for radiation hydrodynamics. Journal of Computational Physics, 229(19):6819–6852, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002858.

Shterev:2010:PBF

[SS10c] Kiril S. Shterev and Stefan K. Stefanov. Pressure based finite volume method for calculation of compressible viscous gas flows. Journal of Computational Physics, 229(2):461–480, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005348.

Suzuki:2010:CSR

[SS10d] Akihiro Suzuki and Toshikazu Shigeyama. A conservative scheme for the relativistic Vlasov–Maxwell system. Journal of Computational Physics, 229(5):1643–1660, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006111.

Szmelter:2010:EBU

[SS10e] Joanna Szmelter and Piotr K. Smolarkiewicz. An edge-based unstructured mesh discretisation in geospherical framework. Journal of Computational Physics, 229(13):4980–4995, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001270.

Santilli:2011:EMS

[SS11] Edward Santilli and Alberto Scotti. An efficient method for solving highly anisotropic elliptic equations. Journal of Computational Physics, 230(23):8342–8359, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100386X.

Saye:2012:AAV

[SS12] R. I. Saye and J. A. Sethian. Analysis and applications of the Voronoi Implicit Interface Method. Journal of Computational Physics, 231(18):6051–6085, July 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001751.

San:2013:CGP

[SS13a] Omer San and Anne E. Staples. A coarse-grid projection method for accelerating incompressible flow computations. Journal of Computational Physics, 233(??):480–508, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005347.

Sprittles:2013:FES

[SS13b] J. E. Sprittles and Y. D. Shikhmurzaev. Finite element simulation of dynamic wetting flows as an interface formation process. Journal of Computational Physics, 233(??):34–65, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003919. See corri-

gendum [SS14b].

Sui:2013:ECM

[SS13c] Y. Sui and Peter D. M. Spelt. An efficient computational model for macroscale simulations of moving contact lines. Journal of Computational Physics, 242(??):37–52, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001150.

Silva:2014:TER

[SS14a] Goncalo Silva and Viriato Semiao. Truncation errors and the rotational invariance of three-dimensional lattice models in the lattice Boltzmann method. Journal of Computational Physics, 269(??):259–279, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002083.

Sprittles:2014:CFE

[SS14b] J. E. Sprittles and Y. D. Shikhmurzaev. Corrigendum to “Finite element simulation of dynamic wetting flows as an interface formation process” [J. Comput. Phys. 233 (2013) 34– 65]. Journal of Computational Physics, 274(??):936, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004483. See [SS13b].

Subber:2014:DDM

[SS14c] Waad Subber and Abhijit Sarkar. A domain decomposition method of stochastic PDEs: an iterative solution techniques using a two-level scalable preconditioner. Journal of Computational Physics, 257(??):298–317, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006025.

Shojaaee:2012:AHD

[SSB^{+}12] Zahra Shojaaee, M. Reza Shaebani, Lothar Brendel, János Török, and Dietrich E. Wolf. An adaptive hierarchical domain decomposition method for parallel contact dynamics simulations of granular materials. Journal of Computational Physics, 231(2):612–628, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005675.

Sambasivan:2013:CCL

[SSB13] Shiv Kumar Sambasivan, Mikhail J. Shashkov, and Donald E. Burton. A cell-centered Lagrangian finite volume approach for computing elasto-plastic response of solids in cylindrical axisymmetric geometries. Journal of Computational Physics, 237(??):251–288, March 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007267.

Shao:2013:DIB

[SSC13] J. Y. Shao, C. Shu, and Y. T. Chew. Development of an immersed boundary-phase field-lattice Boltzmann method for Neumann boundary condition to study contact line dynamics. Journal of Computational Physics, 234(??):8–32, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005086.

Siegel:2012:ACC

[SSFM12] A. Siegel, K. Smith, P. Fischer, and V. Mahadevan. Analysis of communication costs for domain decomposed Monte Carlo methods in nuclear reactor analysis. Journal of Computational Physics, 231(8):3119–3125, April 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007261.

Shitrit:2011:AMS

[SSG11] Shlomy Shitrit, David Sidilkover, and Alexander Gelfgat. An algebraic multigrid solver for transonic flow problems. Journal of Computational Physics, 230(4):1707–1729, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006492.

Shoeybi:2010:AIE

[SSHM10] Mohammad Shoeybi, Magnus Svärd, Frank E. Ham, and Parviz Moin. An adaptive implicit-explicit scheme for the DNS and LES of compressible flows on unstructured grids. Journal of Computational Physics, 229(17):5944–5965, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002032.

Schroeder:2014:SOV

[SSHT14] Craig Schroeder, Alexey Stomakhin, Russell Howes, and Joseph M. Teran. A second order virtual node algorithm for Navier–Stokes flow problems with interfacial forces and discontinuous material properties. Journal of Computational Physics, 265(??):221–245, May 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001028.

Su:2012:LSS

[SSK^{+}12] C.-C. Su, M. R. Smith, F.-A. Kuo, J.-S. Wu, C.-W. Hsieh, and K.-C. Tseng. Large-scale simulations on multiple Graphics Processing Units (GPUs) for the direct simulation Monte Carlo method. Journal of Computational Physics, 231(23):7932–7958, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004238.

Steinmoeller:2013:SND

[SSL13] D. T. Steinmoeller, M. Stastna, and K. G. Lamb. A short note on the discontinuous Galerkin discretization of the pressure projection operator in incompressible flow. Journal of Computational Physics, 251(??):480–486, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004026.

Sun:2014:SOH

[sSLxRyZ14] Zhen sheng Sun, Lei Luo, Yu xin Ren, and Shi ying Zhang. A sixth order hybrid finite difference scheme based on the minimized dispersion and controllable dissipation technique. Journal of Computational Physics, 270(??):238–254, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002393.

Scrinzi:2014:NEP

[SSM14] A. Scrinzi, H. P. Stimming, and N. J. Mauser. On the nonequivalence of perfectly matched layers and exterior complex scaling. Journal of Computational Physics, 269(??):98–107, July 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001806.

Siegel:2013:ELI

[SSR^{+}13] A. R. Siegel, K. Smith, P. K. Romano, B. Forget, and K. Felker. The effect of load imbalances on the performance of Monte Carlo algorithms in LWR analysis. Journal of Computational Physics, 235(??):901–911, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003191.

Shvetsov-Shilovski:2014:SEM

[SSR14] N. I. Shvetsov-Shilovski and E. Räsänen. Stable and efficient momentum-space solutions of the time-dependent Schrödinger equation for one-dimensional atoms in strong laser fields. Journal of Computational Physics, 279(??):174–181, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006329.

Sengupta:2010:SNS

[SSS10] Tapan K. Sengupta, V. K. Suman, and Neelu Singh. Solving Navier–Stokes equation for flow past cylinders using single-block structured and overset grids. Journal of Computational Physics, 229(1):178–199, January 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199910900518X.

Schulze:2011:IOB

[SSS11] J. Schulze, P. Schmid, and J. Sesterhenn. Iterative optimization based on an objective functional in frequency-space with application to jet-noise cancellation. Journal of Computational Physics, 230(15):6075–6098, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002464.

Stecca:2010:UBF

[SST10] Guglielmo Stecca, Annunziato Siviglia, and Eleuterio F. Toro. Upwind-biased FORCE schemes with applications to free-surface shallow flows. Journal of Computational Physics, 229(18):6362–6380, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002500.

Smolarkiewicz:2013:UMA

[SSW13] Piotr K. Smolarkiewicz, Joanna Szmelter, and Andrzej A. Wyszogrodzki. An unstructured-mesh atmospheric model for nonhydrostatic dynamics. Journal of Computational Physics, 254(??):184–199, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005093.

Shishkova:2013:SBE

[SSX13] I. N. Shishkova, S. S. Sazhin, and J.-F. Xie. A solution of the Boltzmann equation in the presence of inelastic collisions. Journal of Computational Physics, 232(1):87–99, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200366X.

Sun:2013:TDS

[SSY13] Xiaosong Sun, Mikio Sakai, and Yoshinori Yamada. Threedimensional simulation of a solid-liquid flow by the DEM-SPH method. Journal of Computational Physics, 248(??):147–176, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002684.

Specogna:2011:DGA

[ST11a] Ruben Specogna and Francesco Trevisan. A discrete geometric approach to solving time independent Schrödinger equation. Journal of Computational Physics, 230(4):1370–1381, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006091.

Sun:2011:SMN

[ST11b] Y. Sun and P. S. P. Tse. Symplectic and multisymplectic numerical methods for Maxwell’s equations. Journal of Computational Physics, 230(5):2076–2094, March 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006704.

Shih:2013:TPC

[ST13] Yin-Tzer Shih and Chih-Ching Tsai. A two-parameter continuation algorithm using radial basis function collocation method for rotating Bose–Einstein condensates. Journal of Computational Physics, 252(??):37–51, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004427.

Song:2010:MWW

[STC10a] Hao Song, Longbin Tao, and Subrata Chakrabarti. Modelling of water wave interaction with multiple cylinders of arbitrary shape. Journal of Computational Physics, 229(5):1498–1513, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006020.

Sunarso:2010:GAM

[STC10b] Alfeus Sunarso, Tomohiro Tsuji, and Shigeomi Chono. GPU-accelerated molecular dynamics simulation for study of liquid crystalline flows. Journal of Computational Physics, 229(15):5486–5497, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000166X.

Srinivasan:2010:RWP

[STD^{+}10] G. Srinivasan, D. M. Tartakovsky, M. Dentz, H. Viswanathan, B. Berkowitz, and B. A. Robinson. Random walk particle tracking simulations of non-Fickian transport in heterogeneous media. Journal of Computational Physics, 229(11):4304–4314, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000872.

Sainath:2014:CPG

[STD14] Kamalesh Sainath, Fernando L. Teixeira, and Burkay Donderici. Complex-plane generalization of scalar Levin transforms: a robust, rapidly convergent method to compute potentials and fields in multi-layered media. Journal of Computational Physics, 269(??):403–422, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002277.

Stechmann:2014:MES

[Ste14] Samuel N. Stechmann. Multiscale eddy simulation for moist atmospheric convection: Preliminary investigation. Journal of Computational Physics, 271(??):99–117, August 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001053.

Sheng:2011:ENM

[STH11] Zhiqiang Sheng, Marc Thiriet, and Frederic Hecht. An efficient numerical method for the equations of steady and unsteady flows of homogeneous incompressible Newtonian fluid. Journal of Computational Physics, 230(3):551–571, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005462.

Stinis:2012:SGO

[Sti12] Panos Stinis. Stochastic global optimization as a filtering problem. Journal of Computational Physics, 231(4):2002–2014, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006747.

Sohn:2010:DMV

[STL^{+}10] Jin Sun Sohn, Yu-Hau Tseng, Shuwang Li, Axel Voigt, and John S. Lowengrub. Dynamics of multicomponent vesicles in a viscous fluid. Journal of Computational Physics, 229(1):119–144, January 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109005099.

Stolk:2013:RCD

[Sto13] Christiaan C. Stolk. A rapidly converging domain decomposition method for the Helmholtz equation. Journal of Computational Physics, 241(??):240–252, May 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000922.

Stupfel:2010:ITC

[Stu10] Bruno Stupfel. Improved transmission conditions for a onedimensional domain decomposition method applied to the solution of the Helmholtz equation. Journal of Computational Physics, 229(3):851–874, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005609.

Szalmas:2010:FIM

[SV10] Lajos Szalmás and Dimitris Valougeorgis. A fast iterative model for discrete velocity calculations on triangular grids. Journal of Computational Physics, 229(11):4315–4326, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110000884.

Schlomer:2013:OLS

[SV13] N. Schlömer and W. Vanroose. An optimal linear solver for the Jacobian system of the extreme type-II Ginzburg–Landau problem. Journal of Computational Physics, 234(??):560–572, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200602X.

Sanderse:2014:BTF

[SVK14] B. Sanderse, R. W. C. P. Verstappen, and B. Koren. Boundary treatment for fourth-order staggered mesh discretizations of the incompressible Navier–Stokes equations. Journal of Computational Physics, 257(??):1472–1505, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006670.

Shestakov:2011:DSM

[SVS11] A. I. Shestakov, R. M. Vignes, and J. S. Stölken. Derivation and solution of multifrequency radiation diffusion equations for homogeneous refractive lossy media. Journal of Computational Physics, 230(4):984–999, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005528.

Smolarkiewicz:2010:PRS

[SW10] Piotr K. Smolarkiewicz and C. Larrabee Winter. Pores resolving simulation of Darcy flows. Journal of Computational Physics, 229(9):3121–3133, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109007177.

Sandberg:2011:EMN

[SW11] Kristian Sandberg and Keith J. Wojciechowski. The EPS method: a new method for constructing pseudospectral derivative operators. Journal of Computational Physics, 230(15): 5836–5863, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111002208.

Stoll:2013:AOS

[SW13] Martin Stoll and Andy Wathen. All-at-once solution of time-dependent Stokes control. Journal of Computational Physics, 232(1):498–515, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005074.

Sulman:2011:OMT

[SWR11] Mohamed Sulman, J. F. Williams, and R. D. Russell. Optimal mass transport for higher dimensional adaptive grid generation. Journal of Computational Physics, 230(9):3302–3330, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000507.

Shi:2012:EMS

[SWX12] Wei Shi, Xinyuan Wu, and Jianlin Xia. Explicit multisymplectic extended leap-frog methods for Hamiltonian wave equations. Journal of Computational Physics, 231(22):7671– 7694, September 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003634.

Shyue:2014:EIS

[SX14] Keh-Ming Shyue and Feng Xiao. An Eulerian interface sharpening algorithm for compressible two-phase flow: the algebraic THINC approach. Journal of Computational Physics, 268(??):326–354, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114001831.

Sheng:2011:FVS

[SY11a] Zhiqiang Sheng and Guangwei Yuan. The finite volume scheme preserving extremum principle for diffusion equations on polygonal meshes. Journal of Computational Physics, 230(7):2588–2604, April 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110007114.

Shigeta:2011:MNS

[SY11b] Takemi Shigeta and D. L. Young. Mathematical and numerical studies on meshless methods for exterior unbounded domain problems. Journal of Computational Physics, 230(17):6900–6915, July 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111003214.

Singh:2011:MBA

[SY11c] Satbir Singh and Donghyun You. A multi-block ADI finite-volume method for incompressible Navier–Stokes equations in complex geometries. Journal of Computational Physics, 230(19):7400–7417, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100355X.

Schmitz:2012:FDS

[SY12a] Phillip G. Schmitz and Lexing Ying. A fast direct solver for elliptic problems on general meshes in 2D. Journal of Computational Physics, 231(4):1314–1338, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006115.

Sheng:2012:IMF

[SY12b] Zhiqiang Sheng and Guangwei Yuan. An improved monotone finite volume scheme for diffusion equation on polygonal meshes. Journal of Computational Physics, 231(9):3739–3754, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000368.

Sirotkin:2012:NPM

[SY12c] Fedir V. Sirotkin and Jack J. Yoh. A new particle method for simulating breakup of liquid jets. Journal of Computational Physics, 231(4):1650–1674, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006206.

Schmitz:2014:FND

[SY14a] Phillip G. Schmitz and Lexing Ying. A fast nested dissection solver for Cartesian 3D elliptic problems using hierarchical matrices. Journal of Computational Physics, 258(??):227–245, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007067.

Sundar:2014:HOM

[SY14b] D. Shyam Sundar and K. S. Yeo. A high order meshless method with compact support. Journal of Computational Physics, 272(??):70–87, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002733.

Strickland:2010:PAS

[SYE10] Michael Strickland and David Yager-Elorriaga. A parallel algorithm for solving the 3 d Schrödinger equation. Journal of Computational Physics, 229(17):6015–6026, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002196.

Shin:2011:LFR

[SYJ11] Seungwon Shin, Ikroh Yoon, and Damir Juric. The Local Front Reconstruction method for direct simulation of twoand three-dimensional multiphase flows. Journal of Computational Physics, 230(17):6605–6646, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002981.

Shen:2010:BVS

[SYL10] Zhijun Shen, Wei Yan, and Guixia Lv. Behavior of viscous solutions in Lagrangian formulation. Journal of Computational Physics, 229(12):4522–4543, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001002.

Shigeta:2012:AMM

[SYL12] Takemi Shigeta, D. L. Young, and Chein-Shan Liu. Adaptive multilayer method of fundamental solutions using a weighted greedy QR decomposition for the Laplace equation. Journal of Computational Physics, 231(21):7118–7132, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003099.

Sjogreen:2014:HOF

[SYV14] Björn Sjögreen, H. C. Yee, and Marcel Vinokur. On high order finite-difference metric discretizations satisfying GCL on moving and deforming grids. Journal of Computational Physics, 265(??):211–220, May 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000904. See corrigendum [SYV17].

Sjogreen:2017:CHO

[SYV17] Björn Sjögreen, H. C. Yee, and Marcel Vinokur. Corrigendum to “On High Order Finite-Difference Metric Discretizations Satisfying GCL on Moving and Deforming Grids” [J. Comput. Phys. 265 (2014) 211–220]. Journal of Computational Physics, 338(??):650, June 1, 2017. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999117301997. See [SYV14].

Shen:2014:RCRa

[SYY14a] Zhijun Shen, Wei Yan, and Guangwei Yuan. A robust and contact resolving Riemann solver on unstructured mesh, Part I, Euler method. Journal of Computational Physics, 268(??):432– 455, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001405.

Shen:2014:RCRb

[SYY14b] Zhijun Shen, Wei Yan, and Guangwei Yuan. A robust and contact resolving Riemann solver on unstructured mesh, Part II, ALE method. Journal of Computational Physics, 268(??):456– 484, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001740.

Shen:2010:LES

[SZ10] Yiqing Shen and Gecheng Zha. Large eddy simulation using a new set of sixth order schemes for compressible viscous terms. Journal of Computational Physics, 229(22):8296–8312, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110004006.

Shen:2011:GFC

[SZ11] Yiqing Shen and Gecheng Zha. Generalized finite compact difference scheme for shock/complex flowfield interaction. Journal of Computational Physics, 230(12):4419–4436, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000647.

Szalmas:2012:VRD

[Sza12] Lajos Szalmas. Variance-reduced DSMC for binary gas flows as defined by the McCormack kinetic model. Journal of Computational Physics, 231(9):3723–3738, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200037X.

Schroeder:2012:SIS

[SZF12] Craig Schroeder, Wen Zheng, and Ronald Fedkiw. Semiimplicit surface tension formulation with a Lagrangian surface mesh on an Eulerian simulation grid. Journal of Computational Physics, 231(4):2092–2115, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006760.

Shen:2012:CSE

[SZH12] Yiqing Shen, Gecheng Zha, and Manuel A. Huerta. E-CUSP scheme for the equations of ideal magnetohydrodynamics with high order WENO scheme. Journal of Computational Physics, 231(19):6233–6247, August 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001878.

Thalhammer:2012:NSA

[TA12] Mechthild Thalhammer and Jochen Abhau. A numerical study of adaptive space and time discretisations for Gross– Pitaevskii equations. Journal of Computational Physics, 231 (20):6665–6681, August 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112002999.

Tsuji:2013:MBP

[TA13] Tetsuro Tsuji and Kazuo Aoki. Moving boundary problems for a rarefied gas: Spatially one-dimensional case. Journal of Computational Physics, 250(??):574–600, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003495.

Thampi:2013:IDL

[TAAS13] Sumesh P. Thampi, Santosh Ansumali, R. Adhikari, and Sauro Succi. Isotropic discrete Laplacian operators from lattice hydrodynamics. Journal of Computational Physics, 234(??):1–7, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112004226.

Toggweiler:2014:NAT

[TAAY14] Matthias Toggweiler, Andreas Adelmann, Peter Arbenz, and Jianjun Yang. A novel adaptive time stepping variant of the Boris–Buneman integrator for the simulation of particle accelerators with space charge. Journal of Computational Physics, 273(??):255–267, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003441.

Tsoutsanis:2014:WSA

[TAD14] Panagiotis Tsoutsanis, Antonios Foivos Antoniadis, and Dimitris Drikakis. WENO schemes on arbitrary unstructured meshes for laminar, transitional and turbulent flows. Journal of Computational Physics, 256(??):254–276, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006062.

Telib:2011:ESS

[TAI11] H. Telib, E. Arian, and A. Iollo. The effect of shocks on second order sensitivities for the quasi-one-dimensional Euler equations. Journal of Computational Physics, 230(23): 8603–8618, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004827.

Takahashi:2014:IBF

[Tak14]  Toru Takahashi.  An interpolation-based fast-multipole

accelerated boundary integral equation method for the three-dimensional wave equation. Journal of Computational Physics, 258(??):809–832, February 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113007584.

Talamo:2013:NST

[Tal13] Alberto Talamo. Numerical solution of the time dependent neutron transport equation by the method of the characteristics. Journal of Computational Physics, 240(??):248–267, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007516.

Tang:2011:NCH

[Tan11] X. Z. Tang. Numerical computation of the helical Chandrasekhar– Kendall modes. Journal of Computational Physics, 230(4): 907–919, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110004262.

Taverniers:2014:NPH

[TAT14] Søren Taverniers, Francis J. Alexander, and Daniel M. Tartakovsky. Noise propagation in hybrid models of nonlinear systems: the Ginzburg–Landau equation. Journal of Computational Physics, 262(??):313–324, April 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000424.

Tautz:2012:SNT

[Tau12] R. C. Tautz. On simplified numerical turbulence models in test-particle simulations. Journal of Computational Physics, 231(14):4537–4541, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001222.

Thornber:2011:APC

[TBMH11] B. Thornber, R. W. Bilger, A. R. Masri, and E. R. Hawkes. An algorithm for LES of premixed compressible flows using the Conditional Moment Closure model. Journal of Computational Physics, 230(20):7687–7705, August 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003974.

Tzoufras:2011:VFP

[TBNT11] M. Tzoufras, A. R. Bell, P. A. Norreys, and F. S. Tsung. A Vlasov–Fokker–Planck code for high energy density physics. Journal of Computational Physics, 230(17):6475–6494, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002828.

Tumolo:2013:SIS

[TBR13] G. Tumolo, L. Bonaventura, and M. Restelli. A semiimplicit, semi-Lagrangian, p-adaptive discontinuous Galerkin method for the shallow water equations. Journal of Computational Physics, 232(1):46–67, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003105.

Tobon:2011:SSM

[TCL11] Luis Tobón, Jiefu Chen, and Qing Huo Liu. Spurious solutions in mixed finite element method for Maxwell’s equations: Dispersion analysis and new basis functions. Journal of Computational Physics, 230(19):7300–7310, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003469.

Thum:2010:EAM

[TCW^{+}10] P. Thum, T. Clees, G. Weyns, G. Nelissen, and J. Deconinck. Efficient algebraic multigrid for migration-diffusion-convection-reaction systems arising in electrochemical simulations. Journal of Computational Physics, 229(19):7260–7276, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003268.

Tierens:2011:BFS

[TD11a] Wouter Tierens and Daniël De Zutter. BOR–FDTD subgridding based on finite element principles. Journal of Computational Physics, 230(12):4519–4535, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001173.

Toulorge:2011:CCR

[TD11b] T. Toulorge and W. Desmet. CFL conditions for Runge– Kutta discontinuous Galerkin methods on triangular grids. Journal of Computational Physics, 230(12):4657–4678, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100129X.

Tierens:2012:UST

[TD12a] Wouter Tierens and Daniël De Zutter. An unconditionally stable time-domain discretization on cartesian meshes for the simulation of nonuniform magnetized cold plasma. Journal of Computational Physics, 231(15):5144–5156, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002094.

Toulorge:2012:ORK

[TD12b]  T. Toulorge and W. Desmet.  Optimal Runge–Kutta

schemes for discontinuous Galerkin space discretizations applied to wave propagation problems. Journal of Computational Physics, 231(4):2067–2091, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006796.

Trouette:2010:AEC

[TDL10] B. Trouette, C. Delcarte, and G. Labrosse. About the ellipticity of the Chebyshev–Gauss–Radau discrete Laplacian with Neumann condition. Journal of Computational Physics, 229(19):7277–7286, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003281.

Tian:2014:FSI

[TDL^{+}14] Fang-Bao Tian, Hu Dai, Haoxiang Luo, James F. Doyle, and Bernard Rousseau. Fluid-structure interaction involving large deformations: 3D simulations and applications to biological systems. Journal of Computational Physics, 258(??):451–469, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007237.

Titarev:2014:CCP

[TDU14] Vladimir Titarev, Michael Dumbser, and Sergey Utyuzhnikov. Construction and comparison of parallel implicit kinetic solvers in three spatial dimensions. Journal of Computational Physics, 256(??):17–33, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005950.

Theillard:2013:SOS

[TDVG13] Maxime Theillard, Landry Fokoua Djodom, Jean-Léopold Vié, and Frédéric Gibou. A second-order sharp numerical method for solving the linear elasticity equations on irregular domains and adaptive grids — application to shape optimization. Journal of Computational Physics, 233(??):430–448, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005207.

Toulopoulos:2011:ABC

[TE11] Ioannis Toulopoulos and John A. Ekaterinaris. Artificial boundary conditions for the numerical solution of the Euler equations by the discontinuous Galerkin method. Journal of Computational Physics, 230(15):5974–5995, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002324.

Teunissen:2014:CWS

[TE14] Jannis Teunissen and Ute Ebert. Controlling the weights of simulation particles: adaptive particle management using k– d trees. Journal of Computational Physics, 259(??):318–330, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113008048.

Tsuji:2012:SPT

[TEY12] Paul Tsuji, Bjorn Engquist, and Lexing Ying. A sweeping preconditioner for time-harmonic Maxwell’s equations with finite elements. Journal of Computational Physics, 231(9):3770– 3783, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000460.

Taylor:2010:CCS

[TF10] Mark A. Taylor and Aimé Fournier. A compatible and conservative spectral element method on unstructured grids. Journal of Computational Physics, 229(17):5879–5895, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110001841.

Tiwari:2013:DIM

[TFP13] Arpit Tiwari, Jonathan B. Freund, and Carlos Pantano. A diffuse interface model with immiscibility preservation. Journal of Computational Physics, 252(??):290–309, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004452.

Tatari:2014:GBN

[TG14a] Mehdi Tatari and Fatemeh Ghasemi. The Galerkin boundary node method for magneto-hydrodynamic (MHD) equation. Journal of Computational Physics, 258(??):634–649, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007328.

Tipireddy:2014:BAH

[TG14b] Ramakrishna Tipireddy and Roger Ghanem. Basis adaptation in homogeneous chaos spaces. Journal of Computational Physics, 259(??):304–317, February 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008085.

Turkel:2013:CSO

[TGGT13] Eli Turkel, Dan Gordon, Rachel Gordon, and Semyon Tsynkov. Compact 2D and 3D sixth order schemes for the Helmholtz equation with variable wave number. Journal of Computational Physics, 232(1):272–287, January 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004627.

Trias:2013:SAD

[TGO13] F. X. Trias, A. Gorobets, and A. Oliva. A simple approach to discretize the viscous term with spatially varying (eddy-)viscosity. Journal of Computational Physics, 253 (??):405–417, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113004944.

Toulorge:2013:RUC

[TGRL13] Thomas Toulorge, Christophe Geuzaine, Jean-François Remacle, and Jonathan Lambrechts. Robust untangling of curvilinear meshes. Journal of Computational Physics, 254(??):8–26, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004956.

Tamain:2010:TFC

[TGT^{+}10] P. Tamain, Ph. Ghendrih, E. Tsitrone, V. Grandgirard, X. Garbet, Y. Sarazin, E. Serre, G. Ciraolo, and G. Chiavassa. TOKAM-3D: a 3D fluid code for transport and turbulence in the edge plasma of Tokamaks. Journal of Computational Physics, 229(2):361–378, January 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005233.

Tan:2014:SGF

[TH14a] Sirui Tan and Lianjie Huang. A staggered-grid finite-difference scheme optimized in the time-space domain for modeling scalar-wave propagation in geophysical problems. Journal of Computational Physics, 276(??):613–634, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005373.

Tseng:2014:IBM

[TH14b] Yu-Hau Tseng and Huaxiong Huang. An immersed boundary method for endocytosis. Journal of Computational Physics, 273(??):143–159, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114003556.

Takamoto:2011:FNS

[TiI11] Makoto Takamoto and Shu ichiro Inutsuka. A fast numerical scheme for causal relativistic hydrodynamics with dissipation. Journal of Computational Physics, 230(18):7002–7017, August 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100341X.

Titarev:2012:IHO

[Tit12] V. A. Titarev. Implicit high-order method for calculating rarefied gas flow in a planar microchannel. Journal of Computational Physics, 231(1):109–134, January 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005183.

Terashima:2012:ASG

[TK12] Hiroshi Terashima and Mitsuo Koshi. Approach for simulating gas-liquid-like flows under supercritical pressures using a high-order central differencing scheme. Journal of Computational Physics, 231(20):6907–6923, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003361. See corrigendum [TK15].

Takezawa:2014:PFM

[TK14] Akihiro Takezawa and Mitsuru Kitamura. Phase field method to optimize dielectric devices for electromagnetic wave propagation. Journal of Computational Physics, 257(??):216–240, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006657.

Terashima:2015:CAS

[TK15] Hiroshi Terashima and Mitsuo Koshi. Corrigendum to “Approach for simulating gas-liquid-like flows under supercritical pressures using a high-order central differencing scheme” [J. Comput. Phys. 231 (20) (2012) 6907–6923]. Journal of Computational Physics, 283(??):609–610, February 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114007657. See [TK12].

Tomin:2013:HMF

[TL13] Pavel Tomin and Ivan Lunati. Hybrid Multiscale Finite Volume method for two-phase flow in porous media. Journal of Computational Physics, 250(??):293–307, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003513.

Tambue:2010:EIA

[TLG10] A. Tambue, G. J. Lord, and S. Geiger. An exponential integrator for advection-dominated reactive transport in heterogeneous porous media. Journal of Computational Physics, 229(10):3957–3969, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000604.

Tryoen:2010:IGM

[TLNE10] J. Tryoen, O. Le Maître, M. Ndjinga, and A. Ern. Intrusive Galerkin methods with upwinding for uncertain nonlinear hyperbolic systems. Journal of Computational Physics, 229(18):6485–6511, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110002688.

Trias:2014:SPD

[TLO^{+}14] F. X. Trias, O. Lehmkuhl, A. Oliva, C. D. Pérez-Segarra, and R. W. C. P. Verstappen. Symmetry-preserving discretization of Navier–Stokes equations on collocated unstructured grids. Journal of Computational Physics, 258(??):246–267, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007079.

Tian:2011:EIB

[TLZ^{+}11] Fang-Bao Tian, Haoxiang Luo, Luoding Zhu, James C. Liao, and Xi-Yun Lu. An efficient immersed boundary-lattice Boltzmann method for the hydrodynamic interaction of elastic filaments. Journal of Computational Physics, 230(19):7266–7283, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003391.

Tanaka:2010:SSP

[TM10] Masayuki Tanaka and Takayuki Masunaga. Stabilization and smoothing of pressure in MPS method by Quasi-Compressibility. Journal of Computational Physics, 229(11): 4279–4290, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110000847.

Takezawa:2010:STO

[TNK10] Akihiro Takezawa, Shinji Nishiwaki, and Mitsuru Kitamura. Shape and topology optimization based on the phase field method and sensitivity analysis. Journal of Computational Physics, 229(7):2697–2718, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006962.

Tokar:2010:NMT

[Tok10] Mikhail Z. Tokar. Numerical modeling of transport barrier formation. Journal of Computational Physics, 229(7):2625–2633, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006809.

Tokar:2011:NSM

[Tok11a] Mikhail Z. Tokar. Numerical solution of momentum balance equations for plasmas with two ion species. Journal of Computational Physics, 230(7):2696–2705, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000313.

Tokman:2011:NCE

[Tok11b] M. Tokman. A new class of exponential propagation iterative methods of Runge–Kutta type (EPIRK). Journal of Computational Physics, 230(24):8762–8778, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005109.

Tonti:2014:WSD

[Ton14] Enzo Tonti. Why starting from differential equations for computational physics? Journal of Computational Physics, 257(??):1260–1290, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005548.

Tricco:2012:CHD

[TP12] Terrence S. Tricco and Daniel J. Price. Constrained hyperbolic divergence cleaning for smoothed particle magnetohydrodynamics. Journal of Computational Physics, 231(21):7214–7236, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003737.

Tuckmantel:2014:HVT

[TP14] T. Tückmantel and A. Pukhov. H-VLPL: a three-dimensional relativistic PIC/fluid hybrid code. Journal of Computational Physics, 269(??):168–180, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002009.

Torres:2013:API

[TPA^{+}13] C. E. Torres, H. Parishani, O. Ayala, L. F. Rossi, and L.-P. Wang. Analysis and parallel implementation of a forced Nbody problem. Journal of Computational Physics, 245(??):235– 258, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001848.

Tartakovsky:2011:DRM

[TPF11] Alexandre M. Tartakovsky, Alexander Panchenko, and Kim F. Ferris. Dimension reduction method for ODE fluid models. Journal of Computational Physics, 230(23):8554–8572, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004633.

Tang:2014:OGM

[TQW14] H. S. Tang, K. Qu, and X. G. Wu. An overset grid method for integration of fully 3D fluid dynamics and geophysics fluid dynamics models to simulate multiphysics coastal ocean flows. Journal of Computational Physics, 273(??):548–571, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003568.

Traasdahl:2011:HON

[TR11] Øystein Tråsdahl and Einar M. Rønquist. High order numerical approximation of minimal surfaces. Journal of Computational Physics, 230(12):4795–4810, June 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001367.

Tu:2012:IAB

[TR12] Jonathan H. Tu and Clarence W. Rowley. An improved algorithm for balanced POD through an analytic treatment of impulse response tails. Journal of Computational Physics, 231(16):5317–5333, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002045.

Turcksin:2014:DDS

[TR14] Bruno Turcksin and Jean C. Ragusa. Discontinuous diffusion synthetic acceleration for s_{n} transport on 2D arbitrary polygonal meshes. Journal of Computational Physics, 274 (??):356–369, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114004100.

Touil:2014:DLE

[TRL14] Hatem Touil, Denis Ricot, and Emmanuel Lévêque. Direct and large-eddy simulation of turbulent flows on composite multi-resolution grids by the lattice Boltzmann method. Journal of Computational Physics, 256(??):220–233, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005299.

Toledo-Redondo:2013:PTA

[TRSMM^{+}13] Sergio Toledo-Redondo, Alfonso Salinas, Juan Antonio Morente-Molinera, Antonio Méndez, Jesús Fornieles, Jorge Portí, and Juan Antonio Morente. Parallel 3D-TLM algorithm for simulation of the Earth–ionosphere cavity. Journal of Computational Physics, 236(??):367–379, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006511.

Tryggvason:2012:NEC

[Try12] Gretar Tryggvason. Note: Enriching the content of the Journal of Computational Physics. Journal of Computational Physics, 231(22):7475, September 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003415.

Tan:2010:ILW

[TS10] Sirui Tan and Chi-Wang Shu. Inverse Lax–Wendroff procedure for numerical boundary conditions of conservation laws.

Journal of Computational Physics, 229(21):8144–8166, October 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003979.

Tan:2011:HOM

[TS11] Sirui Tan and Chi-Wang Shu. A high order moving boundary treatment for compressible inviscid flows. Journal of Computational Physics, 230(15):6023–6036, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911100235X.

Tolstykh:2012:VDM

[TS12] Mikhail A. Tolstykh and Vladimir V. Shashkin. Vorticitydivergence mass-conserving semi-Lagrangian shallow-water model using the reduced grid on the sphere. Journal of Computational Physics, 231(11):4205–4233, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001052.

Tranquilli:2014:EKM

[TS14] Paul Tranquilli and Adrian Sandu. Exponential-Krylov methods for ordinary differential equations. Journal of Computational Physics, 278(??):31–46, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005592.

Tanguy:2014:BNM

[TSL^{+}14] Sébastien Tanguy, Michaël Sagan, Benjamin Lalanne, Frédéric Couderc, and Catherine Colin. Benchmarks and numerical methods for the simulation of boiling flows. Journal of Computational Physics, 264(??):1–22, May 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000412.

Teigen:2011:DIM

[TSLV11] Knut Erik Teigen, Peng Song, John Lowengrub, and Axel Voigt. A diffuse-interface method for two-phase flows with soluble surfactants. Journal of Computational Physics, 230(2):375–393, January 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005231.

Turrell:2013:MCA

[TSR13] A. E. Turrell, M. Sherlock, and S. J. Rose. A Monte Carlo algorithm for degenerate plasmas. Journal of Computational Physics, 249(??):13–21, September 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113002362.

Tsukerman:2010:TDS

[Tsu10] Igor Tsukerman. Trefftz difference schemes on irregular stencils. Journal of Computational Physics, 229(8):2948–2963, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109007049.

Tokareva:2010:HTR

[TT10] S. A. Tokareva and E. F. Toro. HLLC-type Riemann solver for the Baer–Nunziato equations of compressible two-phase flow. Journal of Computational Physics, 229(10):3573–3604, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000318.

Tsoutsanis:2011:WSA

[TTD11] P. Tsoutsanis, V. A. Titarev, and D. Drikakis. WENO schemes on arbitrary mixed-element unstructured meshes in three space dimensions. Journal of Computational Physics, 230(4):1585–1601, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006388.

Thomas:2012:RVF

[TTR^{+}12] A. G. R. Thomas, M. Tzoufras, A. P. L. Robinson, R. J. Kingham, C. P. Ridgers, M. Sherlock, and A. R. Bell. A review of Vlasov–Fokker–Planck numerical modeling of inertial confinement fusion plasma. Journal of Computational Physics, 231(3):1051–1079, February 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005717.

Toth:2012:ANA

[TvdHS^{+}12] Gábor Tóth, Bart van der Holst, Igor V. Sokolov, Darren L. De Zeeuw, Tamas I. Gombosi, Fang Fang, Ward B. Manchester, Xing Meng, Dalal Najib, Kenneth G. Powell, Quentin F. Stout, Alex Glocer, Ying-Juan Ma, and Merav Opher. Adaptive numerical algorithms in space weather modeling. Journal of Computational Physics, 231(3):870–903, February 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100088X.

Thorgilsson:2014:RGF

[TVE14] G. Thorgilsson, G. Viktorsson, and S. I. Erlingsson. Recursive Green’s function method for multi-terminal nanostructures. Journal of Computational Physics, 261(??):256–266, March 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000096.

Trontin:2012:SCC

[TVEC12] P. Trontin, S. Vincent, J. L. Estivalezes, and J. P. Caltagirone. A subgrid computation of the curvature by a particle/level-set method. Application to a front-tracking/ghost-fluid method for incompressible flows. Journal of Computational Physics, 231(20):6990–7010, August 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003580.

Thompson:2014:SCR

[TWMM14] Richard J. Thompson, Andrew Wilson, Trevor Moeller, and Charles L. Merkle. A strong conservative Riemann solver for the solution of the coupled Maxwell and Navier–Stokes equations. Journal of Computational Physics, 258(??):431–450, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007171.

Tan:2012:EIH

[TWSN12] Sirui Tan, Cheng Wang, Chi-Wang Shu, and Jianguo Ning. Efficient implementation of high order inverse Lax–Wendroff boundary treatment for conservation laws. Journal of Computational Physics, 231(6):2510–2527, March 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006930.

Tian:2011:ECD

[TY11a] Zhen F. Tian and P. X. Yu. An efficient compact difference scheme for solving the streamfunction formulation of the incompressible Navier–Stokes equations. Journal of Computational Physics, 230(17):6404–6419, July 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007059.

Tsuji:2011:FDA

[TY11b] Paul Tsuji and Lexing Ying. A fast directional algorithm for high-frequency electromagnetic scattering. Journal of Computational Physics, 230(14):5471–5487, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001021.

Tygert:2010:FAS

[Tyg10] Mark Tygert. Fast algorithms for spherical harmonic expansions, III. Journal of Computational Physics, 229(18): 6181–6192, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110002561.

Tyliszczak:2014:HOC

[Tyl14] Artur Tyliszczak. A high-order compact difference algorithm for half-staggered grids for laminar and turbulent incompressible flows. Journal of Computational Physics, 276(??):438–467, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005361.

Urquiza:2014:WIS

[UGF14] José M. Urquiza, André Garon, and Marie-Isabelle Farinas. Weak imposition of the slip boundary condition on curved boundaries for Stokes flow. Journal of Computational Physics, 256(??):748–767, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005895.

Ullah:2013:TFT

[UGkM13] Mohammed Aman Ullah, Wenbin Gao, and De kang Mao. Towards front-tracking based on conservation in two space dimensions III, tracking interfaces. Journal of Computational Physics, 242(??):268–303, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001459.

Ullrich:2011:AFV

[UJ11] Paul A. Ullrich and Christiane Jablonowski. An analysis of 1D finite-volume methods for geophysical problems on refined grids. Journal of Computational Physics, 230(3):706–725, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005590.

Ullrich:2012:MNH

[UJ12] Paul A. Ullrich and Christiane Jablonowski. MCore: a non-hydrostatic atmospheric dynamical core utilizing highorder finite-volume methods. Journal of Computational Physics, 231(15):5078–5108, June 1, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112002057.

Ullrich:2010:HOF

[UJvL10] Paul A. Ullrich, Christiane Jablonowski, and Bram van Leer. High-order finite-volume methods for the shallowwater equations on the sphere. Journal of Computational Physics, 229(17):6104–6134, August 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002317.

Uddin:2014:CBE

[UKP14] H. Uddin, R. M. J. Kramer, and C. Pantano. A Cartesianbased embedded geometry technique with adaptive high-order finite differences for compressible flow around complex geometries. Journal of Computational Physics, 262(??):379–407, April 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000254.

Ueckermann:2013:NSD

[ULS13] M. P. Ueckermann, P. F. J. Lermusiaux, and T. P. Sapsis. Numerical schemes for dynamically orthogonal equations of stochastic fluid and ocean flows. Journal of Computational Physics, 233(??):272–294, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005098.

Urbano:2013:ARS

[UN13] Annafederica Urbano and Francesco Nasuti. An approximate Riemann solver for real gas parabolized Navier–Stokes equations. Journal of Computational Physics, 233(??):574–591, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005505.

Unfer:2013:AFS

[Unf13] Thomas Unfer. An asynchronous framework for the simulation of the plasma/flow interaction. Journal of Computational Physics, 236(??):229–246, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006961.

Usabiaga:2013:ICP

[UPDB13] F. Balboa Usabiaga, I. Pagonabarraga, and R. Delgado-Buscalioni. Inertial coupling for point particle fluctuating hydrodynamics. Journal of Computational Physics, 235(??): 701–722, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112006493.

Uminsky:2012:MMV

[UWB12] David Uminsky, C. Eugene Wayne, and Alethea Barbaro. A multi-moment vortex method for 2D viscous fluids. Journal of Computational Physics, 231(4):1705–1727, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111006437.

Vaibhav:2011:ABC

[Vai11] V. Vaibhav. Artificial boundary conditions for certain evolution PDEs with cubic nonlinearity for non-compactly supported initial data. Journal of Computational Physics, 230(8):3205–3229, April 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000490.

Vaibhav:2014:MAT

[Vai14] V. Vaibhav. Microlocal approach towards construction of nonreflecting boundary conditions. Journal of Computational Physics, 272(??):588–607, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400326X.

Validi:2014:LRS

[Val14] AbdoulAhad Validi. Low-rank separated representation surrogates of high-dimensional stochastic functions: Application in Bayesian inference. Journal of Computational Physics, 260(??):37–53, March 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113008310.

Vartdal:2013:OPL

[VB13] Magnus Vartdal and Arne Bøckmann. An oriented particle level set method based on surface coordinates. Journal of Computational Physics, 251(??):237–250, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004117.

Vincent:2010:ELM

[VBCM10] Stéphane Vincent, Guillaume Balmigère, Jean-Paul Caltagirone, and Erick Meillot. Eulerian–Lagrangian multiscale methods for solving scalar equations — application to incompressible two-phase flows. Journal of Computational Physics, 229(1):73–106, January 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109004926.

Villa:2013:APS

[VBGM13] Andrea Villa, Luca Barbieri, Marco Gondola, and Roberto Malgesini. An asymptotic preserving scheme for the streamer simulation. Journal of Computational Physics, 242(??):86–102, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001265.

Vikram:2010:ACE

[VBSK10] M. Vikram, A. Baczewski, B. Shanker, and L. Kempel. Accelerated Cartesian expansion (ACE) based framework for the rapid evaluation of diffusion, lossy wave, and Klein–Gordon potentials. Journal of Computational Physics, 229(24):9119–9134, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004729.

VanGenechten:2010:TBN

[VBVD10] Bert Van Genechten, Bart Bergen, Dirk Vandepitte, and Wim Desmet. A Trefftz-based numerical modelling framework for Helmholtz problems with complex multiple-scatterer configurations. Journal of Computational Physics, 229(18):6623–6643, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002779.

Vincent:2011:INA

[VCJ11] P. E. Vincent, P. Castonguay, and A. Jameson. Insights from von Neumann analysis of high-order flux reconstruction schemes. Journal of Computational Physics, 230(22):8134–8154, September 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004323.

Vogman:2014:DGH

[VCS14] G. V. Vogman, P. Colella, and U. Shumlak. Dory–Guest– Harris instability as a benchmark for continuum kinetic Vlasov–Poisson simulations of magnetized plasmas. Journal of Computational Physics, 277(??):101–120, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005609.

VanPoppel:2010:GFL

[VDD10a] B. P. Van Poppel, O. Desjardins, and J. W. Daily. A ghost fluid, level set methodology for simulating multiphase electrohydrodynamic flows with application to liquid fuel injection. Journal of Computational Physics, 229(20):7977–7996, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003724.

Vaughan:2010:RSS

[VDD10b] T. G. Vaughan, P. D. Drummond, and A. J. Drummond. A retrodictive stochastic simulation algorithm. Journal of Computational Physics, 229(10):3777–3791, May 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000501.

Vidovic:2013:CCD

[VDD^{+}13] D. Vidović, M. Dotlić, M. Dimkić, M. Pusić, and B. Pokorni. Convex combinations for diffusion schemes. Journal of Computational Physics, 246(??):11–27, August 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300212X.

Vincent:2014:LVT

[VdMS^{+}14] Stéphane Vincent, Jorge César Brändle de Motta, Arthur Sarthou, Jean-Luc Estivalezes, Olivier Simonin, and Eric Climent. A Lagrangian VOF tensorial penalty method for the DNS of resolved particle-laden flows. Journal of Computational Physics, 256(??):582–614, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005615.

Vidovic:2011:ANL

[VDP11] D. Vidović, M. Dimkić, and M. Pusić. Accelerated nonlinear finite volume method for diffusion. Journal of Computational Physics, 230(7):2722–2735, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000349.

vanDijk:2012:HDM

[vDS12] P. J. van Dijk and D. C. Swailes. Hermite–DG methods for pdf equations modelling particle transport and deposition in turbulent boundary layers. Journal of Computational Physics, 231(14):4904–4920, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911200188X.

vanderVegt:2012:HMSa

[vdVR12a] J. J. W. van der Vegt and S. Rhebergen. hp-multigrid as smoother algorithm for higher order discontinuous Galerkin discretizations of advection dominated flows: Part I. multilevel analysis. Journal of Computational Physics, 231(22): 7537–7563, September 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999112003129.

vanderVegt:2012:HMSb

[vdVR12b] J. J. W. van der Vegt and S. Rhebergen. HP-multigrid as smoother algorithm for higher order discontinuous Galerkin discretizations of advection dominated flows. Part II: Optimization of the Runge–Kutta smoother. Journal of Computational Physics, 231(22):7564–7583, September 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003117.

Viner:2013:SSS

[VED13] Kevin C. Viner, Craig C. Epifanio, and James D. Doyle. A steady-state solver and stability calculator for nonlinear internal wave flows. Journal of Computational Physics, 251(??):432–444, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001174.

vanEs:2014:FDS

[vEKdB14] Bram van Es, Barry Koren, and Hugo J. de Blank. Finitedifference schemes for anisotropic diffusion. Journal of Computational Physics, 272(??):526–549, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003155.

Villa:2010:ITM

[VF10] A. Villa and L. Formaggia. Implicit tracking for multifluid simulations. Journal of Computational Physics, 229 (16):5788–5802, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001968.

Vion:2014:DSP

[VG14] A. Vion and C. Geuzaine. Double sweep preconditioner for optimized Schwarz methods applied to the Helmholtz problem. Journal of Computational Physics, 266(??):171–190, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114001351.

Vay:2011:NMI

[VGCMG11] J.-L. Vay, C. G. R. Geddes, E. Cormier-Michel, and D. P. Grote. Numerical methods for instability mitigation in the modeling of laser wakefield accelerators in a Lorentz-boosted frame. Journal of Computational Physics, 230(15):5908–5929, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002270.

Vay:2013:DDM

[VHG13] Jean-Luc Vay, Irving Haber, and Brendan B. Godfrey. A domain decomposition method for pseudo-spectral electromagnetic simulations of plasmas. Journal of Computational Physics, 243(??):260–268, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001873.

vanHinsberg:2011:ESO

[vHtTBC11] M. A. T. van Hinsberg, J. H. M. ten Thije Boonkkamp, and H. J. H. Clercx. An efficient, second order method for the approximation of the Basset history force. Journal of Computational Physics, 230(4):1465–1478, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006224.

Vikas:2013:RTM

[VHWF13] V. Vikas, C. D. Hauck, Z. J. Wang, and R. O. Fox. Radiation transport modeling using extended quadrature method of moments. Journal of Computational Physics, 246(??):221–241, August 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002064.

Venturi:2012:NEE

[VK12] D. Venturi and G. E. Karniadakis. New evolution equations for the joint response-excitation probability density function of stochastic solutions to first-order nonlinear PDEs. Journal of Computational Physics, 231(21):7450–7474, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003865.

Vire:2011:MDE

[VKBK11] A. Viré, D. Krasnov, T. Boeck, and B. Knaepen. Modeling and discretization errors in large eddy simulations of hydrodynamic and magnetohydrodynamic channel flows. Journal of Computational Physics, 230(5):1903–1922, March 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006546.

Velechovsky:2013:SEB

[VKL^{+}13] J. Velechovský, M. Kucharík, R. Liska, M. Shashkov, and P. Váchal. Symmetry- and essentially-bound-preserving fluxcorrected remapping of momentum in staggered ALE hydrodynamics. Journal of Computational Physics, 255(??): 590–611, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005755.

vanLeer:2011:HOV

[vL11] Bram van Leer. A historical oversight: Vladimir P. Kolgan and his high-resolution scheme. Journal of Computational Physics, 230(7):2378–2383, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007060.

Violeau:2014:MTS

[VL14] Damien Violeau and Agnès Leroy. On the maximum time step in weakly compressible SPH. Journal of Computational Physics, 256(??):388–415, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006050.

Vie:2013:SVC

[VLM13] Aymeric Vié, Frédérique Laurent, and Marc Massot. Sizevelocity correlations in hybrid high order moment/multifluid methods for polydisperse evaporating sprays: Modeling and numerical issues. Journal of Computational Physics, 237(??):177–210, March 15, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007231.

Verschaeve:2010:CNS

[VM10] Joris C. G. Verschaeve and Bernhard Müller. A curved no-slip boundary condition for the lattice Boltzmann method. Journal of Computational Physics, 229(19):6781–6803, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002834.

Vilar:2014:DGD

[VMA14] François Vilar, Pierre-Henri Maire, and Rémi Abgrall. A discontinuous Galerkin discretization for solving the twodimensional gas dynamics equations written under total Lagrangian formulation on general unstructured grids. Journal of Computational Physics, 276(??):188–234, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114005233.

Vergnault:2011:TRL

[VMS11] E. Vergnault, O. Malaspinas, and P. Sagaut. A timereversal lattice Boltzmann method. Journal of Computational Physics, 230(22):8155–8167, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004335.

Vergnault:2012:LBM

[VMS12] E. Vergnault, O. Malaspinas, and P. Sagaut. A lattice Boltzmann method for nonlinear disturbances around an arbitrary base flow. Journal of Computational Physics, 231(24):8070–8082, October 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112003944.

Vollebregt:2014:NSE

[Vol14] E. A. H. Vollebregt. A new solver for the elastic normal contact problem using conjugate gradients, deflation, and an FFT-based preconditioner. Journal of Computational Physics, 257(??):333–351, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006724.

Varsakelis:2010:ELC

[VP10] C. Varsakelis and M. V. Papalexandris. The equilibrium limit of a constitutive model for two-phase granular mixtures and its numerical approximation. Journal of Computational Physics, 229(11):4183–4207, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000781.

Varsakelis:2014:NMT

[VP14] Christos Varsakelis and Miltiadis V. Papalexandris. A numerical method for two-phase flows of dense granular mixtures. Journal of Computational Physics, 257(??):737–756, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006992.

Villedieu:2011:TOR

[VQRD11] N. Villedieu, T. Quintino, M. Ricchiuto, and H. Deconinck. Third order residual distribution schemes for the Navier–Stokes equations. Journal of Computational Physics, 230(11):4301–4315, May 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000700X.

Vymazal:2011:HOU

[VQVD11] Martin Vymazal, Tiago Quintino, Nadège Villedieu, and Herman Deconinck. High-order upwind residual distribution schemes on isoparametric curved elements. Journal of Computational Physics, 230(4):890–906, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002883.

Vuik:2014:MTC

[VR14] Mathea J. Vuik and Jennifer K. Ryan. Multiwavelet troubled-cell indicator for discontinuity detection of discontinuous Galerkin schemes. Journal of Computational Physics, 270(??):138–160, August 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114002344.

Vanella:2010:DFE

[VRB10] Marcos Vanella, Patrick Rabenold, and Elias Balaras. A direct-forcing embedded-boundary method with adaptive mesh refinement for fluid-structure interaction problems. Journal of Computational Physics, 229(18):6427–6449, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000255X.

Veerapaneni:2011:FAS

[VRBZ11] Shravan K. Veerapaneni, Abtin Rahimian, George Biros, and Denis Zorin. A fast algorithm for simulating vesicle flows in three dimensions. Journal of Computational Physics, 230(14):5610–5634, June 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002063.

Vreman:2011:SEM

[Vre11] A. W. Vreman. Stabilization of the Eulerian model for incompressible multiphase flow by artificial diffusion. Journal of Computational Physics, 230(4):1639–1651, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006406.

Vreman:2014:PMI

[Vre14] A. W. Vreman. The projection method for the incompressible Navier–Stokes equations: the pressure near a no-slip wall. Journal of Computational Physics, 263(??):353–374, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400062X.

vanRees:2011:CVP

[vRLPK11] Wim M. van Rees, Anthony Leonard, D. I. Pullin, and Petros Koumoutsakos. A comparison of vortex and pseudospectral methods for the simulation of periodic vortical flows at high Reynolds numbers. Journal of Computational Physics, 230(8):2794–2805, April 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006467.

Vergnault:2011:ALB

[VS11] E. Vergnault and P. Sagaut. Application of Lattice Boltzmann method to sensitivity analysis via complex differentiation. Journal of Computational Physics, 230(13):5417–5429, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002051.

Vergnault:2014:ABL

[VS14] E. Vergnault and P. Sagaut. An adjoint-based lattice Boltzmann method for noise control problems. Journal of Computational Physics, 276(??):39–61, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005208.

Vincent:2011:ALP

[VSC^{+}11] Stéphane Vincent, Arthur Sarthou, Jean-Paul Caltagirone, Fabien Sonilhac, Pierre Février, Christian Mignot, and Grégoire Pianet. Augmented Lagrangian and penalty methods for the simulation of two-phase flows interacting with moving solids. Application to hydroplaning flows interacting with real tire tread patterns. Journal of Computational Physics, 230(4):956–983, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005504.

Vos:2010:EIF

[VSK10] Peter E. J. Vos, Spencer J. Sherwin, and Robert M. Kirby. From h to p efficiently: Implementing finite and spectral/hp element methods to achieve optimal performance for low- and high-order discretisations. Journal of Computational Physics, 229(13):5161–5181, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001506.

Viswanath:2013:NSS

[VT13] Divakar Viswanath and Ian Tobasco. Navier–Stokes solver using Green’s functions I: Channel flow and plane Couette flow. Journal of Computational Physics, 251(??):414–431, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300418X.

vanThang:2010:SLB

[vTCL^{+}10] Pham van Thang, Bastien Chopard, Laurent Lefèvre, Diemer Anda Ondo, and Eduardo Mendes. Study of the 1D lattice Boltzmann shallow water equation and its coupling to build a canal network. Journal of Computational Physics, 229(19):7373–7400, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003372.

Venturi:2013:EPE

[VTTK13] D. Venturi, D. M. Tartakovsky, A. M. Tartakovsky, and G. E. Karniadakis. Exact PDF equations and closure approximations for advective-reactive transport. Journal of Computational Physics, 243(??):323–343, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001733.

vantHof:2012:MME

[vV12] Bas van’t Hof and Arthur E. P. Veldman. Mass, momentum and energy conserving (MaMEC) discretizations on general grids for the compressible Euler and shallow water equations. Journal of Computational Physics, 231(14):4723–4744, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001428.

Veneziani:2013:AAS

[VV13] Alessandro Veneziani and Umberto Villa. ALADINS: an ALgebraic splitting time ADaptive solver for the Incompressible Navier–Stokes equations. Journal of Computational Physics, 238(??):359–375, April 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112007310.

vanVeldhuizen:2010:PNK

[vVVK10] S. van Veldhuizen, C. Vuik, and C. R. Kleijn. On projected Newton–Krylov solvers for instationary laminar reacting gas flows. Journal of Computational Physics, 229(5):1724–1738, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006159.

Vachal:2014:SPD

[VW14a] Pavel Váchal and Burton Wendroff. A symmetry preserving dissipative artificial viscosity in an r–z staggered Lagrangian discretization. Journal of Computational Physics, 258(??):118–136, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113007122.

Vong:2014:CDS

[VW14b] Seakweng Vong and Zhibo Wang. A compact difference scheme for a two dimensional fractional Klein–Gordon equation with Neumann boundary conditions. Journal of Computational Physics, 274(??):268–282, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400429X.

Vikas:2013:RHO

[VWF13] V. Vikas, Z. J. Wang, and R. O. Fox. Realizable high-order finite-volume schemes for quadrature-based moment methods applied to diffusion population balance equations. Journal of Computational Physics, 249(??):162–179, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003264.

Viswanathan:2011:NIM

[VWP11] Sharadha Viswanathan, Haifeng Wang, and Stephen B. Pope. Numerical implementation of mixing and molecular transport in LES/PDF studies of turbulent reacting flows. Journal of Computational Physics, 230(17):6916–6957, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100324X.

Vikas:2011:RHO

[VWPF11] V. Vikas, Z. J. Wang, A. Passalacqua, and R. O. Fox. Realizable high-order finite-volume schemes for quadrature-based moment methods. Journal of Computational Physics, 230 (13):5328–5352, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111001896.

Verma:2014:IBS

[VXB14] Siddhartha Verma, Y. Xuan, and G. Blanquart. An improved bounded semi-Lagrangian scheme for the turbulent transport of passive scalars. Journal of Computational Physics, 272(??):1–22, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002496.

vanyen:2010:WBD

[vydCNS^{+}10] Romain Nguyen van yen, Diego del Castillo-Negrete, Kai Schneider, Marie Farge, and Guangye Chen. Waveletbased density estimation for noise reduction in plasma simulations using particles. Journal of Computational Physics, 229(8):2821–2839, April 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109006810.

Wajid:2012:OBF

[WA12] Hafiz Abdul Wajid and Sobia Ayub. An optimally blended finite-spectral element scheme with minimal dispersion for Maxwell equations. Journal of Computational Physics, 231(24):8176–8187, October 15, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004329.

Wang:2011:TLD

[WABI11] Z. Wang, I. Akhtar, J. Borggaard, and T. Iliescu. Twolevel discretizations of nonlinear closure models for proper orthogonal decomposition. Journal of Computational Physics, 230(1):126–146, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110005097.

Waltz:2013:OST

[Wal13] Jacob Waltz. Operator splitting and time accuracy in Lagrange plus remap solution methods. Journal of Computational Physics, 253(??):247–258, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004890.

Wieselquist:2014:CLF

[WAM14] William A. Wieselquist, Dmitriy Y. Anistratov, and Jim E. Morel. A cell-local finite difference discretization of the loworder quasidiffusion equations for neutral particle transport on unstructured quadrilateral meshes. Journal of Computational Physics, 273(??):343–357, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400357X.

Wang:2010:MMR

[Wan10] Ping Wang. Modeling material responses by arbitrary Lagrangian Eulerian formulation and adaptive mesh refinement method. Journal of Computational Physics, 229(5):1573–1599, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006068.

Wan:2011:AHO

[Wan11a] Xiaoliang Wan. An adaptive high-order minimum action method. Journal of Computational Physics, 230(24):8669– 8682, October 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111004785.

Wang:2011:ENM

[Wan11b] Hanquan Wang. An efficient numerical method for computing dynamics of spin F = 2 Bose–Einstein condensates. Journal of Computational Physics, 230(15):6155–6168, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002609.

Wan:2013:MAM

[Wan13a] Xiaoliang Wan. A minimum action method for small random perturbations of two-dimensional parallel shear flows. Journal of Computational Physics, 235(??):497–514, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112005955.

Wang:2013:FAS

[Wan13b] Qiqi Wang. Forward and adjoint sensitivity computation of chaotic dynamical systems. Journal of Computational Physics, 235(??):1–13, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005360.

Wang:2014:PGM

[Wan14] Hanquan Wang. A projection gradient method for computing ground state of spin-2 Bose–Einstein condensates. Journal of Computational Physics, 274(??):473–488, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004227.

Werner:2013:MAS

[WBC13] Gregory R. Werner, Carl A. Bauer, and John R. Cary. A more accurate, stable, FDTD algorithm for electromagnetics in anisotropic dielectrics. Journal of Computational Physics, 255(??):436–455, December 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005391.

White:2011:RIT

[WC11] Alexander Thomas White and Chuh Khiun Chong. Rotational invariance in the three-dimensional lattice Boltzmann method is dependent on the choice of lattice. Journal of Computational Physics, 230(16):6367–6378, July 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002798.

Wang:2014:TPC

[WC14a] Y.-S. Wang and C.-S. Chien. A two-parameter continuation method for computing numerical solutions of spin-1 Bose–Einstein condensates. Journal of Computational Physics, 256(??):198–213, January 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006001.

Wu:2014:FID

[WC14b] Yuqi Wu and Xiao-Chuan Cai. A fully implicit domain decomposition based ALE framework for three-dimensional fluidstructure interaction with application in blood flow computation. Journal of Computational Physics, 258(??):524–537, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007225.

Waltz:2014:MST

[WCM^{+}14] J. Waltz, T. R. Canfield, N. R. Morgan, L. D. Risinger, and J. G. Wohlbier. Manufactured solutions for the threedimensional Euler equations with relevance to Inertial Confinement Fusion. Journal of Computational Physics, 267(??):196– 209, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001661.

Williams:2013:ESF

[WCVJ13] D. M. Williams, P. Castonguay, P. E. Vincent, and A. Jameson. Energy stable flux reconstruction schemes for advectiondiffusion problems on triangles. Journal of Computational Physics, 250(??):53–76, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113003318.

White:2011:HPH

[WD11] J. B. White III and J. J. Dongarra. High-performance highresolution semi-Lagrangian tracer transport on a sphere. Journal of Computational Physics, 230(17):6778–6799, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003123.

Wang:2013:FFD

[WD13a] Hong Wang and Ning Du. A fast finite difference method for three-dimensional time-dependent space-fractional diffusion equations and its efficient implementation. Journal of Computational Physics, 253(??):50–63, November 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004701.

Wang:2013:SPI

[WD13b] Hong Wang and Ning Du. A superfast-preconditioned iterative method for steady-state space-fractional diffusion equations. Journal of Computational Physics, 240(??):49–57, May 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004305.

Wang:2014:FAD

[WD14] Hong Wang and Ning Du. Fast alternating-direction finite difference methods for three-dimensional space-fractional diffusion equations. Journal of Computational Physics, 258 (??):305–318, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S002199911300716X.

Wu:2010:LPN

[WDGY10] Jiming Wu, Zihuan Dai, Zhiming Gao, and Guangwei Yuan. Linearity preserving nine-point schemes for diffusion equation on distorted quadrilateral meshes. Journal of Computational Physics, 229(9):3382–3401, May 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110000227.

Wang:2010:AIS

[WdHX10] Shen Wang, Maarten V. de Hoop, and Jianlin Xia. Acoustic inverse scattering via Helmholtz operator factorization and optimization. Journal of Computational Physics, 229(22):8445–8462, November 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004201.

Wang:2014:OGM

[WDP^{+}14] Gaofeng Wang, Florent Duchaine, Dimitrios Papadogiannis, Ignacio Duran, Stéphane Moreau, and Laurent Y. M. Gicquel. An overset grid method for large eddy simulation of turbomachinery stages. Journal of Computational Physics, 274(??):333–355, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004136.

Webb:2014:SIM

[Web14] Stephen D. Webb. Symplectic integration of magnetic systems. Journal of Computational Physics, 270(??):570–576, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114002368. See com-

ment [ZJS15].

Weigel:2012:PPS

[Wei12] Martin Weigel. Performance potential for simulating spin models on GPU. Journal of Computational Physics, 231 (8):3064–3082, April 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111007066.

Wendroff:2010:CAV

[Wen10] Burton Wendroff. A compact artificial viscosity equivalent to a tensor viscosity. Journal of Computational Physics, 229(19):6673–6675, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002950.

Waagan:2011:RNS

[WFK11] K. Waagan, C. Federrath, and C. Klingenberg. A robust numerical scheme for highly compressible magnetohydrodynamics: Nonlinear stability, implementation and tests. Journal of Computational Physics, 230(9):3331–3351, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111000519.

Wang:2010:HOC

[WFL10] Limin Wang, Yilin Fan, and Lingai Luo. Heuristic optimality criterion algorithm for shape design of fluid flow. Journal of Computational Physics, 229(20):8031–8044, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000375X.

Wodo:2011:CES

[WG11] Olga Wodo and Baskar Ganapathysubramanian. Computationally efficient solution to the Cahn–Hilliard equation:

Adaptive implicit time schemes, mesh sensitivity analysis and the 3D isoperimetric problem. Journal of Computational Physics, 230(15):6037–6060, July 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002361.

Wu:2014:IBS

[WG14] Jiming Wu and Zhiming Gao. Interpolation-based secondorder monotone finite volume schemes for anisotropic diffusion equations on general grids. Journal of Computational Physics, 275(??):569–588, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004914.

Wu:2012:SLP

[WGD12] Jiming Wu, Zhiming Gao, and Zihuan Dai. A stabilized linearity-preserving scheme for the heterogeneous and anisotropic diffusion problems on polygonal meshes. Journal of Computational Physics, 231(21):7152–7169, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003762.

Wang:2013:FOC

[WGX13] Tingchun Wang, Boling Guo, and Qiubin Xu. Fourthorder compact and energy conservative difference schemes for the nonlinear Schrödinger equation in two dimensions. Journal of Computational Physics, 243(??):382–399, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001824.

Weisse:2011:ECG

[WH11] Andrea Y. Weiße and Wilhelm Huisinga. Error-controlled global sensitivity analysis of ordinary differential equations. Journal of Computational Physics, 230(17):6824–6842, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003159.

Wang:2014:LSS

[WHB14] Qiqi Wang, Rui Hu, and Patrick Blonigan. Least Squares Shadowing sensitivity analysis of chaotic limit cycle oscillations. Journal of Computational Physics, 267(??):210–224, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001715.

White:2010:HOP

[Whi10] Michael D. White. High-order parabolic beam approximation for aero-optics. Journal of Computational Physics, 229 (15):5465–5485, August 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001658.

Wang:2014:AMS

[WHT14] Yixuan Wang, Hadi Hajibeygi, and Hamdi A. Tchelepi. Algebraic multiscale solver for flow in heterogeneous porous media. Journal of Computational Physics, 259(??):284–303, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007869.

Witteveen:2013:SSC

[WI13a] Jeroen A. S. Witteveen and Gianluca Iaccarino. Simplex stochastic collocation with ENO-type stencil selection for robust uncertainty quantification. Journal of Computational Physics, 239(??):1–21, April 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300003X.

Witteveen:2013:SRS

[WI13b] Jeroen A. S. Witteveen and Gianluca Iaccarino. Subcell resolution in simplex stochastic collocation for spatial discontinuities. Journal of Computational Physics, 251(??):17–52, October 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004014.

Waugh:2013:MFC

[WIJ13] Iain Waugh, Simon Illingworth, and Matthew Juniper. Matrix-free continuation of limit cycles for bifurcation analysis of large thermoacoustic systems. Journal of Computational Physics, 240(??):225–247, May 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000077.

Wilson:2013:SFP

[Wil13] Helen J. Wilson. Stokes flow past three spheres. Journal of Computational Physics, 245(??):302–316, July 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001988.

Witteveen:2010:SOF

[Wit10] Jeroen A. S. Witteveen. Second order front tracking for the Euler equations. Journal of Computational Physics, 229 (7):2719–2739, April 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109006986.

Wittek:2013:HPD

[Wit13] Peter Wittek. High-performance dynamic quantum clustering on graphics processors. Journal of Computational Physics, 233(??):262–271, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112005165.

Wu:2013:CHO

[WJC13] Yu Mao Wu, Li Jun Jiang, and Weng Cho Chew. Computing highly oscillatory physical optics integral on the polygonal domain by an efficient numerical steepest descent path method. Journal of Computational Physics, 236(??):408–425, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006560.

Wang:2013:CET

[WJCG13] Duo Wang, Xiangmin Jiao, Rebecca Conley, and James Glimm. On the curvature effect of thin membranes. Journal of Computational Physics, 233(??):449–463, January 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005190.

Wei:2010:LSM

[WJTY10] Dongming Wei, Shi Jin, Richard Tsai, and Xu Yang. A level set method for the semiclassical limit of the Schrödinger equation with discontinuous potentials. Journal of Computational Physics, 229(19):7440–7455, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003414.

Wang:2010:MEF

[WK10] Moran Wang and Qinjun Kang. Modeling electrokinetic flows in microchannels using coupled lattice Boltzmann methods. Journal of Computational Physics, 229(3):728–744, February 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005518.

Wang:2013:HDG

[WK13a] Bo Wang and B. C. Khoo. Hybridizable discontinuous Galerkin method (HDG) for Stokes interface flow. Journal of Computational Physics, 247(??):262–278, August 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002489.

Wei:2013:GAM

[WK13b] J. Wei and F. E. Kruis. GPU-accelerated Monte Carlo simulation of particle coagulation based on the inverse method. Journal of Computational Physics, 249(??):67–79, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002957.

Winters:2014:ADA

[WK14] Andrew R. Winters and David A. Kopriva. ALE-DGSEM approximation of wave reflection and transmission from a moving medium. Journal of Computational Physics, 263(??):233–267, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000497.

Wheatley:2010:RRS

[WKH10] V. Wheatley, H. Kumar, and P. Huguenot. On the role of Riemann solvers in Discontinuous Galerkin methods for magnetohydrodynamics. Journal of Computational Physics, 229(3):660–680, February 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005488.

Wang:2014:CGA

[WKL^{+}14] S. Wang, B. C. Khoo, G. R. Liu, G. X. Xu, and L. Chen. Coupling GSM/ALE with ES–FEM–T3 for fluiddeformable structure interactions. Journal of Computational Physics, 276(??):315–340, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005099.

Wang:2013:IES

[WKMJ13] Dafang Wang, Robert M. Kirby, Rob S. MacLeod, and Chris R. Johnson. Inverse electrocardiographic source localization of ischemia: an optimization framework and finite element solution. Journal of Computational Physics, 250(??):403–424, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003677.

Wollaber:2011:LSA

[WL11] Allan B. Wollaber and Edward W. Larsen. A linear stability analysis for nonlinear, grey, thermal radiative transfer problems. Journal of Computational Physics, 230(4):1528–1546, February 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110006340.

Wang:2013:INU

[WLL13] Yong-Lei Wang, Aatto Laaksonen, and Zhong-Yuan Lu. Implementation of non-uniform FFT based Ewald summation in dissipative particle dynamics method. Journal of Computational Physics, 235(??):666–682, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005542.

Wu:2011:SMR

[WLPG11] Jinlong Wu, Tiejun Li, Xiang Peng, and Hong Guo. Statistical method for resolving the photon-photoelectron-counting inversion problem. Journal of Computational Physics, 230(3):726–743, February 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110005693.

Wang:2013:FTA

[WLW13a] Bin Wang, Kai Liu, and Xinyuan Wu. A filon-type asymptotic approach to solving highly oscillatory second-order initial value problems. Journal of Computational Physics, 243(??):210–223, June 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001861.

Weller:2013:RKI

[WLW13b] Hilary Weller, Sarah-Jane Lock, and Nigel Wood. Runge– Kutta IMEX schemes for the Horizontally Explicit/Vertically Implicit (HEVI) solution of wave equations. Journal of Computational Physics, 252(??):365–381, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004555.

Wang:2014:FAN

[WLW14] Yan-Feng Wang, Zhi-Feng Liu, and Xiao-Hong Wang. Finite analytic numerical method for three-dimensional fluid flow in heterogeneous porous media. Journal of Computational Physics, 278(??):169–181, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005804.

Wang:2010:HOM

[WMI10] Qiqi Wang, Parviz Moin, and Gianluca Iaccarino. A high order multivariate approximation scheme for scattered data sets. Journal of Computational Physics, 229(18):6343–6361, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911000241X.

Wang:2013:MMP

[WMK13] Weiwei Wang, Nipa A. Mody, and Michael R. King. Multiscale model of platelet translocation and collision. Journal of Computational Physics, 244(??):223–235, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004603.

Wasserman:2010:RIM

[WMYYG10] M. Wasserman, Y. Mor-Yossef, I. Yavneh, and J. B. Greenberg. A robust implicit multigrid method for RANS equations with two-equation turbulence models. Journal of Computational Physics, 229(16):5820–5842, August 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001993.

Wollman:2012:NAV

[Wol12] Stephen Wollman. Numerical approximation of the Vlasov– Maxwell–Fokker–Planck system in two dimensions. Journal of Computational Physics, 231(9):3483–3517, May 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007303.

Ward:2010:HCD

[WP10] G. M. Ward and D. I. Pullin. A hybrid, center-difference, limiter method for simulations of compressible multicomponent flows with Mie–Grüneisen equation of state. Journal of Computational Physics, 229(8):2999–3018, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900713X.

Willert:2014:RMC

[WP14] Jeffrey Willert and H. Park. Residual Monte Carlo highorder solver for Moment-Based Accelerated Thermal Radiative Transfer equations. Journal of Computational Physics, 276(??):405–421, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005324.

Willert:2014:CAM

[WPK14] Jeffrey Willert, H. Park, and D. A. Knoll. A comparison of acceleration methods for solving the neutron transport k-eigenvalue problem. Journal of Computational Physics, 274(??):681–694, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004513.

Walther:2012:MSW

[WPKK12] Jens H. Walther, Matej Praprotnik, Evangelos M. Kotsalis, and Petros Koumoutsakos. Multiscale simulation of water flow past a C_{540} fullerene. Journal of Computational Physics, 231(7):2677–2681, April 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111007273.

Wang:2010:WSO

[WPP10] Haifeng Wang, Pavel P. Popov, and Stephen B. Pope. Weak second-order splitting schemes for Lagrangian Monte Carlo particle methods for the composition PDF/FDF transport equations. Journal of Computational Physics, 229(5):1852– 1878, March 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999109006317.

Wang:2013:ZLM

[WQCS13] Yibin Wang, Ning Qin, Greg Carnie, and Shahrokh Shahpar. Zipper layer method for linking two dissimilar structured meshes. Journal of Computational Physics, 255(??): 130–148, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113005500.

Wang:2011:SGO

[WR11] Yaqi Wang and Jean C. Ragusa. Standard and goal-oriented adaptive mesh refinement applied to radiation transport on 2D unstructured triangular meshes. Journal of Computational Physics, 230(3):763–788, February 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005747.

Weggler:2010:NNM

[WRH10] S. Weggler, V. Rutka, and A. Hildebrandt. A new numerical method for nonlocal electrostatics in biomolecular simulations. Journal of Computational Physics, 229(11):4059–4074, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000719.

Walker:2010:SOP

[WS10a] Shawn W. Walker and Michael J. Shelley. Shape optimization of peristaltic pumping. Journal of Computational Physics, 229(4):1260–1291, February 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109005828.

Wang:2010:ITT

[WS10b] Chunwu Wang and Chi-Wang Shu. An interface treating technique for compressible multi-medium flow with Runge– Kutta discontinuous Galerkin method. Journal of Computational Physics, 229(23):8823–8843, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004493.

Wu:2010:IIB

[WS10c] J. Wu and C. Shu. An improved immersed boundary-lattice Boltzmann method for simulating three-dimensional incompressible flows. Journal of Computational Physics, 229(13): 5022–5042, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110001348.

Wu:2011:SAL

[WS11] J. Wu and C. Shu. A solution-adaptive lattice Boltzmann method for two-dimensional incompressible viscous flows. Journal of Computational Physics, 230(6):2246–2269, March 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110006777.

Woodroffe:2014:HSW

[WS14] J. R. Woodroffe and A. V. Streltsov. Hybrid simulation of whistler excitation by electron beams in twodimensional non-periodic domains. Journal of Computational Physics, 276(??):468–478, November 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004732.

Wilcox:2010:HOD

[WSBG10] Lucas C. Wilcox, Georg Stadler, Carsten Burstedde, and Omar Ghattas. A high-order discontinuous Galerkin method for wave propagation through coupled elastic-acoustic media. Journal of Computational Physics, 229(24):9373–9396, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005024.

Wang:2012:HLS

[WSS12] Y. Wang, S. Simakhina, and M. Sussman. A hybrid level setvolume constraint method for incompressible two-phase flow. Journal of Computational Physics, 231(19):6438–6471, August 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200321X.

Wiens:2013:RSK

[WSW13] Jeffrey K. Wiens, John M. Stockie, and J. F. Williams. Riemann solver for a kinematic wave traffic model with discontinuous flux. Journal of Computational Physics, 242(??):1–23, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001435.

Wang:2012:HOF

[WSYS12] Wei Wang, Chi-Wang Shu, H. C. Yee, and Björn Sjögreen. High order finite difference methods with subcell resolution for advection equations with stiff source terms. Journal of Computational Physics, 231(1):190–214, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005250.

Wang:2012:FGM

[WT12a] Hong Wang and Hao Tian. A fast Galerkin method with efficient matrix assembly and storage for a peridynamic model. Journal of Computational Physics, 231(23):7730–7738, October 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003166.

Wang:2012:UQK

[WT12b] Peng Wang and Daniel M. Tartakovsky. Uncertainty quantification in kinematic-wave models. Journal of Computational Physics, 231(23):7868–7880, October 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004147.

Wang:2013:TRB

[WT13] Xiaochen Wang and Hamdi A. Tchelepi. Trust-region based solver for nonlinear transport in heterogeneous porous media. Journal of Computational Physics, 253(??):114–137, November 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004725.

Wu:2014:FVL

[WT14] Kailiang Wu and Huazhong Tang. Finite volume local evolution Galerkin method for two-dimensional relativistic hydrodynamics. Journal of Computational Physics, 256(??):277–307, January 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006013.

Willert:2014:LAA

[WTK14] Jeffrey Willert, William T. Taitano, and Dana Knoll. Leveraging Anderson Acceleration for improved convergence of iterative solutions to transport systems. Journal of Computational Physics, 273(??):278–286, September 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003611.

Wu:2010:SMM

[WTSZ10] Xue-Hong Wu, Wen-Quan Tao, Sheng-Ping Shen, and Xing-Wang Zhu. A stabilized MLPG method for steady state incompressible fluid flow simulation. Journal of Computational Physics, 229(22):8564–8577, November 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004377.

Wu:2013:OGG

[WTYC13] Nan-Jing Wu, Ting-Kuei Tsay, Tun-Chi Yang, and Hung-Yuan Chang. Orthogonal grid generation of an irregular region using a local polynomial collocation method. Journal of Computational Physics, 243(??):58–73, June 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000296.

Wang:2014:CDS

[WV14] Zhibo Wang and Seakweng Vong. Compact difference schemes for the modified anomalous fractional sub-diffusion equation and the fractional diffusion-wave equation. Journal of Computational Physics, 277(??):1–15, November 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114005580.

Wang:2011:ADF

[WW11] Hong Wang and Kaixin Wang. An O(N log^{2} N) alternatingdirection finite difference method for two-dimensional fractional diffusion equations. Journal of Computational Physics, 230(21):7830–7839, September 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004074.

Wang:2013:ECM

[WW13] Kun Wang and Yau Shu Wong. Error correction method for Navier–Stokes equations at high Reynolds numbers. Journal of Computational Physics, 255(??):245–265, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005421.

Wang:2014:IFT

[WW14] Bin Wang and Xinyuan Wu. Improved Filon-type asymptotic methods for highly oscillatory differential equations with multiple time scales. Journal of Computational Physics, 276(??):62–73, November 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005282.

Wang:2010:DFD

[WWS10] Hong Wang, Kaixin Wang, and Treena Sircar. A direct O(N log^{2} N) finite difference method for fractional diffusion equations. Journal of Computational Physics, 229(21):8095–8104, October 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003943.

Wu:2013:DNS

[WWS^{+}13a] Lei Wu, Craig White, Thomas J. Scanlon, Jason M. Reese, and Yonghao Zhang. Deterministic numerical solutions of the Boltzmann equation using the fast spectral method. Journal of Computational Physics, 250(??):27–52, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113003276.

Wu:2013:EEP

[WWS13b] Xinyuan Wu, Bin Wang, and Wei Shi. Efficient energypreserving integrators for oscillatory Hamiltonian systems. Journal of Computational Physics, 235(??):587–605, February 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112006043.

Wang:2010:HNS

[WWX^{+}10] J. Wang, L.-P. Wang, Z. Xiao, Y. Shi, and S. Chen. A hybrid numerical simulation of isotropic compressible turbulence. Journal of Computational Physics, 229(13):5257–5279, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001610.

Wang:2013:CFF

[WWZ13] Chu Wang, Xingshi Wang, and Lucy Zhang. Connectivityfree front tracking method for multiphase flows with free surfaces. Journal of Computational Physics, 241(??):58–75, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000545.

Wang:2014:SMB

[WX14] Weiwei Wang and Chuanju Xu. Spectral methods based on new formulations for coupled Stokes and Darcy equations. Journal of Computational Physics, 257(??):126–142, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113006505.

Wang:2013:CND

[WXY13] Dongling Wang, Aiguo Xiao, and Wei Yang. Crank–Nicolson difference scheme for the coupled nonlinear Schrödinger equations with the Riesz space fractional derivative. Journal of Computational Physics, 242(??):670–681, June 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001563.

Wang:2014:LIC

[WXY14] Dongling Wang, Aiguo Xiao, and Wei Yang. A linearly implicit conservative difference scheme for the space fractional coupled nonlinear Schrödinger equations. Journal of Computational Physics, 272(??):644–655, September 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114003167.

Wang:2010:EAD

[WXZ10] Bo Wang, Ziqing Xie, and Zhimin Zhang. Error analysis of a discontinuous Galerkin method for Maxwell equations in dispersive media. Journal of Computational Physics, 229(22):8552–8563, November 1, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004316.

Weymouth:2010:CVF

[WY10] G. D. Weymouth and Dick K.-P. Yue. Conservative Volumeof-Fluid method for free-surface simulations on Cartesiangrids. Journal of Computational Physics, 229(8):2853–2865, April 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006974.

Weymouth:2011:BDI

[WY11] G. D. Weymouth and Dick K. P. Yue. Boundary data immersion method for Cartesian-grid simulations of fluidbody interaction problems. Journal of Computational Physics, 230(16):6233–6247, July 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111002610.

Wu:2013:SWS

[WYC13] C. S. Wu, D. L. Young, and C. L. Chiu. Simulation of wavestructure interaction by hybrid cartesian/immersed boundary and arbitrary Lagrangian–Eulerian finite-element method. Journal of Computational Physics, 254(??):155–183, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004877.

Wang:2011:CLD

[WYS^{+}11] Wei Wang, H. C. Yee, Björn Sjögreen, Thierry Magin, and Chi-Wang Shu. Construction of low dissipative high-order well-balanced filter schemes for non-equilibrium flows. Journal of Computational Physics, 230(11):4316–4335, May 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002202.

Wang:2012:NVF

[WYS12a] Zhaoyuan Wang, Jianming Yang, and Frederick Stern. A new volume-of-fluid method with a constructed distance function on general structured grids. Journal of Computational Physics, 231(9):3703–3722, May 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112000435.

Wang:2012:SCO

[WYS12b] Zhaoyuan Wang, Jianming Yang, and Frederick Stern. A simple and conservative operator-splitting semi-Lagrangian volume-of-fluid advection scheme. Journal of Computational Physics, 231(15):4981–4992, June 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002100.

Wu:2014:TOA

[WYT14] Kailiang Wu, Zhicheng Yang, and Huazhong Tang. A third-order accurate direct Eulerian GRP scheme for the Euler equations in gas dynamics. Journal of Computational Physics, 264(??):177–208, May 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114000795.

Wang:2010:SGS

[WYYK10] X. Y. Wang, P. Yu, K. S. Yeo, and B. C. Khoo. SVD– GFD scheme to simulate complex moving body problems in 3D space. Journal of Computational Physics, 229(6):2314–2338, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006664.

Wolgemuth:2010:MBN

[WZ10] Charles W. Wolgemuth and Mark Zajac. The moving boundary node method: a level set-based, finite volume algorithm with applications to cell motility. Journal of Computational Physics, 229(19):7287–7308, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003293.

Wang:2011:IBM

[WZ11] Shizhao Wang and Xing Zhang. An immersed boundary method based on discrete stream function formulation for two- and three-dimensional incompressible flows. Journal of Computational Physics, 230(9):3479–3499, May 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111000787.

Wan:2013:PGM

[WZ13] Jiang Wan and Nicholas Zabaras. A probabilistic graphical model approach to stochastic multiscale partial differential equations. Journal of Computational Physics, 250(??):477–510, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113003471.

Wan:2014:PGM

[WZ14] Jiang Wan and Nicholas Zabaras. A probabilistic graphical model based stochastic input model construction. Journal of Computational Physics, 272(??):664–685, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003349.

Wang:2013:LBM

[WZG^{+}13] Haoming Wang, Haibo Zhao, Zhaoli Guo, Yongxiang He, and Chuguang Zheng. Lattice Boltzmann method for simulations of gas-particle flows over a backward-facing step. Journal of Computational Physics, 239(??):57–71, April 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000053.

Wang:2014:ARI

[WZN14] Dongyong Wang, Lei Zhang, and Qing Nie. Arrayrepresentation integration factor method for high-dimensional systems. Journal of Computational Physics, 258(??):585–600, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007481.

Wang:2012:RHO

[WZSN12] Cheng Wang, Xiangxiong Zhang, Chi-Wang Shu, and Jianguo Ning. Robust high order discontinuous Galerkin schemes for two-dimensional gaseous detonations. Journal of Computational Physics, 231(2):653–665, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005857.

Wen:2014:GIF

[WZT^{+}14] Binghai Wen, Chaoying Zhang, Yusong Tu, Chunlei Wang, and Haiping Fang. Galilean invariant fluid-solid interfacial dynamics in lattice Boltzmann simulations. Journal of Computational Physics, 266(??):161–170, June 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001387.

Wang:2014:HCP

[WZZ14] Li-Lian Wang, Jing Zhang, and Zhimin Zhang. On hpconvergence of prolate spheroidal wave functions and a new well-conditioned prolate-collocation scheme. Journal of Computational Physics, 268(??):377–398, July 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001788.

Wang:2014:EPF

[WZZZ14] Quanxiang Wang, Zhiyue Zhang, Xinhua Zhang, and Quanyong Zhu. Energy-preserving finite volume element method for the improved Boussinesq equation. Journal of Computational Physics, 270(??):58–69, August 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911400240X.

Xu:2014:NRK

[XCL14] Zhiliang Xu, Xu-Yan Chen, and Yingjie Liu. A new Runge– Kutta discontinuous Galerkin method with conservation constraint to improve CFL condition for solving conservation laws. Journal of Computational Physics, 278(??):348–377, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114006184.

Xu:2014:CCG

[XDZ^{+}14] Chuanfu Xu, Xiaogang Deng, Lilun Zhang, Jianbin Fang, Guangxue Wang, Yi Jiang, Wei Cao, Yonggang Che, Yongxian Wang, Zhenghua Wang, Wei Liu, and Xinghua Cheng. Collaborating CPU and GPU for large-scale high-order CFD simulations with complex grids on the TianHe-1A supercomputer. Journal of Computational Physics, 278(??):275–297, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114005786.

Xiao:2014:NLM

[XFB^{+}14] D. Xiao, F. Fang, A. G. Buchan, C. C. Pain, I. M. Navon, J. Du, and G. Hu. Non-linear model reduction for the Navier– Stokes equations using residual DEIM method. Journal of Computational Physics, 263(??):1–18, April 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114000321.

Xia:2014:MGM

[XFC^{+}14] Kelin Xia, Xin Feng, Zhan Chen, Yiying Tong, and Guo-Wei Wei. Multiscale geometric modeling of macromolecules I: Cartesian representation. Journal of Computational Physics, 257(??):912–936, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006487.

Xu:2010:UGK

[XH10] Kun Xu and Juan-Chen Huang. A unified gas-kinetic scheme for continuum and rarefied flows. Journal of Computational Physics, 229(20):7747–7764, October 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110003475.

Xu:2013:SIF

[XH13] Hui Xu and Yinnian He. Some iterative finite element methods for steady Navier–Stokes equations with different viscosities. Journal of Computational Physics, 232(1):136–152, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003932.

Xu:2014:SMD

[XH14] Qinwu Xu and Jan S. Hesthaven. Stable multi-domain spectral penalty methods for fractional partial differential equations. Journal of Computational Physics, 257(??):241–258, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113006554.

Xiao:2013:MSC

[XHF13] Nan Xiao, Jay D. Humphrey, and C. Alberto Figueroa. Multiscale computational model of three-dimensional hemodynamics within a deformable full-body arterial network. Journal of Computational Physics, 244(??):22–40, July 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005475.

Xiao:2011:RTS

[XIC11] Feng Xiao, Satoshi Ii, and Chungang Chen. Revisit to the THINC scheme: a simple algebraic VOF algorithm. Journal of Computational Physics, 230(19):7086–7092, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003615.

Xie:2014:NSD

[Xie14a] Dexuan Xie. New solution decomposition and minimization schemes for Poisson–Boltzmann equation in calculation of biomolecular electrostatics. Journal of Computational Physics, 275(??):294–309, October 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004926.

Xie:2014:MME

[Xie14b] Hehu Xie. A multigrid method for eigenvalue problem. Journal of Computational Physics, 274(??):550–561, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004379.

Xie:2014:MMF

[XIIX14] Bin Xie, Satoshi Ii, Akio Ikebata, and Feng Xiao. A multimoment finite volume method for incompressible Navier–Stokes equations on unstructured grids: Volume-average/point-value formulation. Journal of Computational Physics, 277(??):

138–162, November 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114005579.

Xing:2014:EWB

[Xin14] Yulong Xing. Exactly well-balanced discontinuous Galerkin methods for the shallow water equations with moving water equilibrium. Journal of Computational Physics, 257 (??):536–553, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006785.

Xiao:2012:CDM

[XJ12] Heng Xiao and Patrick Jenny. A consistent dual-mesh framework for hybrid LES/RANS modeling. Journal of Computational Physics, 231(4):1848–1865, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006516.

Xue:2014:PDP

[XJ14] Ming-Feng Xue and Jian-Ming Jin. A preconditioned dualprimal finite element tearing and interconnecting method for solving three-dimensional time-harmonic Maxwell’s equations. Journal of Computational Physics, 274(??):920–935, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004471.

Xu:2011:ACE

[XL11] Liang Xu and Tiegang Liu. Accuracies and conservation errors of various ghost fluid methods for multi-medium Riemann problem. Journal of Computational Physics, 230(12):4975– 4990, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111001641.

Xu:2011:PWH

[XLD^{+}11] Zhiliang Xu, Yingjie Liu, Huijing Du, Guang Lin, and Chi-Wang Shu. Point-wise hierarchical reconstruction for discontinuous Galerkin and finite volume methods for solving conservation laws. Journal of Computational Physics, 230 (17):6843–6865, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111003184.

Li:2013:PMC

[xLqW13] Zhao xiang Li and Zhong qing Wang. Pseudospectral methods for computing the multiple solutions of the Lane–Emden equation. Journal of Computational Physics, 255(??):407–421, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113005676.

Xu:2014:SFE

[XM14] Zhenli Xu and A. C. Maggs. Solving fluctuation-enhanced Poisson–Boltzmann equations. Journal of Computational Physics, 275(??):310–322, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114004847.

Xu:2013:CAS

[XMDG13] Gang Xu, Bernard Mourrain, Régis Duvigneau, and André Galligo. Constructing analysis-suitable parameterization of computational domain from CAD boundary by variational harmonic method. Journal of Computational Physics, 252 (??):275–289, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113004592.

See correction [XMDG14].

Xu:2014:CSA

[XMDG14] Gang Xu, Bernard Mourrain, Régis Duvigneau, and André Galligo. Corrigendum to “Constructing analysis-suitable parameterization of computational domain from CAD boundary by variational harmonic method” [J. Comput. Phys. 252 (2013) 275–289]. Journal of Computational Physics, 272(??):865, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911400309X. See [XMDG13].

Xu:2012:SAD

[XMS12] Hui Xu, Orestis Malaspinas, and Pierre Sagaut. Sensitivity analysis and determination of free relaxation parameters for the weakly-compressible MRT-LBM schemes. Journal of Computational Physics, 231(21):7335–7367, August 30, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112003646.

Xisto:2014:PBH

[XPO14] Carlos M. Xisto, José C. Páscoa, and Paulo J. Oliveira. A pressure-based high resolution numerical method for resistive MHD. Journal of Computational Physics, 275(??):323–345, October 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004896.

Xiong:2013:PMP

[XQX13] Tao Xiong, Jing-Mei Qiu, and Zhengfu Xu. A parametrized maximum principle preserving flux limiter for finite difference RK-WENO schemes with applications in incompressible flows. Journal of Computational Physics, 252(??):310–331, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004567.

Xiong:2014:HOM

[XQXC14] Tao Xiong, Jing-Mei Qiu, Zhengfu Xu, and Andrew Christlieb. High order maximum principle preserving semi-Lagrangian finite difference WENO schemes for the Vlasov equation. Journal of Computational Physics, 273(??):618–639, September 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114003969.

Xu:2014:LSM

[XR14] Jian-Jun Xu and Weiqing Ren. A level-set method for twophase flows with moving contact line and insoluble surfactant. Journal of Computational Physics, 263(??):71–90, April 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114000394.

Xu:2011:OLD

[XS11] Hui Xu and Pierre Sagaut. Optimal low-dispersion lowdissipation LBM schemes for computational aeroacoustics. Journal of Computational Physics, 230(13):5353–5382, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002014.

Xu:2013:AAL

[XS13] Hui Xu and Pierre Sagaut. Analysis of the absorbing layers for the weakly-compressible lattice Boltzmann methods. Journal of Computational Physics, 245(??):14–42, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113001915.

Xing:2015:CPE

[XS15] Tao Xing and Frederick Stern. Comment on “A procedure for the estimation of the numerical uncertainty of CFD calculations based on grid refinement studies” (L. Eça and M. Hoekstra, Journal of Computational Physics 262 (2014) 104–130). Journal of Computational Physics, 301(??):484–486, November 15, 2015. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999115005756. See [EH14]

and reply [EH15].

Xu:2013:NMN

[XST13] Jian Xu, Sihong Shao, and Huazhong Tang. Numerical methods for nonlinear Dirac equation. Journal of Computational Physics, 245(??):131–149, July 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300209X.

Xu:2011:BCC

[Xu11] Sheng Xu. A boundary condition capturing immersed interface method for 3D rigid objects in a flow. Journal of Computational Physics, 230(19):7176–7190, August 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003238.

Xuan:2010:WIB

[XW10] Li-Jun Xuan and Jie-Zhi Wu. A weighted-integral based scheme. Journal of Computational Physics, 229(17):5999–6014, August 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002184.

Xu:2014:HOA

[XW14] Li Xu and Peifen Weng. High order accurate and low dissipation method for unsteady compressible viscous flow computation on helicopter rotor in forward flight. Journal of Computational Physics, 258(??):470–488, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007092.

Xiong:2012:MCS

[XX12] Ye Xiong and Shi-Jie Xiong. Monte Carle simulation of quantum transport through nanostructures. Journal of Computational Physics, 231(4):1197–1208, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005651.

Xuan:2013:NGK

[XX13] Li-Jun Xuan and Kun Xu. A new gas-kinetic scheme based on analytical solutions of the BGK equation. Journal of Computational Physics, 234(??):524–539, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005967.

Xia:2010:LDG

[XXS10] Yinhua Xia, Yan Xu, and Chi-Wang Shu. Local discontinuous Galerkin methods for the generalized Zakharov system. Journal of Computational Physics, 229(4):1238–1259, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005816.

Xie:2014:MTI

[XY14] Hehu Xie and Xiaobo Yin. Metric tensors for the interpolation error and its gradient in L^{p} norm. Journal of Computational Physics, 256(??):543–562, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911300613X.

Xu:2012:LSC

[XYL12] Jian-Jun Xu, Yin Yang, and John Lowengrub. A level-set continuum method for two-phase flows with insoluble surfactant. Journal of Computational Physics, 231(17):5897–5909, July 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112002562.

Xia:2011:MME

[XZW11] Kelin Xia, Meng Zhan, and Guo-Wei Wei. MIB method for elliptic equations with multi-material interfaces. Journal of Computational Physics, 230(12):4588–4615, June 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111001264.

Xia:2012:ADM

[XZWW12] Kelin Xia, Meng Zhan, Decheng Wan, and Guo-Wei Wei. Adaptively deformed mesh based interface method for elliptic equations with discontinuous coefficients. Journal of Computational Physics, 231(4):1440–1461, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006267.

Yao:2014:EMA

[Yao14] Jin Yao. An easy method to accelerate an iterative algebraic equation solver. Journal of Computational Physics, 267(??):139–145, June 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999114001478.

Yu:2014:FDF

[YB14] Rixin Yu and Xue-Song Bai. A fully divergence-free method for generation of inhomogeneous and anisotropic turbulence with large spatial variation. Journal of Computational Physics, 256(??):234–253, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005998.

Yu:2013:GFM

[YBK13] Yue Yu, Hyoungsu Baek, and George Em Karniadakis. Generalized fictitious methods for fluid-structure interactions: Analysis and simulations. Journal of Computational Physics, 245(??):317–346, July 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113002039.

Yang:2011:PMM

[YC11] Chao Yang and Xiao-Chuan Cai. Parallel multilevel methods for implicit solution of shallow water equations with nonsmooth topography on the cubed-sphere. Journal of Computational Physics, 230(7):2523–2539, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110007011.

Yin:2013:MMI

[YCH^{+}13] Youbing Yin, Jiwoong Choi, Eric A. Hoffman, Merryn H. Tawhai, and Ching-Long Lin. A multiscale MDCT imagebased breathing lung model with time-varying regional ventilation. Journal of Computational Physics, 244(??):168–192, July 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007383.

Yang:2012:AAD

[YCLK12] Xiu Yang, Minseok Choi, Guang Lin, and George Em Karniadakis. Adaptive ANOVA decomposition of stochastic incompressible and compressible flows. Journal of Computational Physics, 231(4):1587–1614, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111006280.

Yang:2011:NMR

[YCY11] F.-L. Yang, C. H. Chen, and D. L. Young. A novel mesh regeneration algorithm for 2D FEM simulations of flows with moving boundary. Journal of Computational Physics, 230(9):3276– 3301, May 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911100026X.

Yang:2011:CIM

[YD11] Lun Yang and Kaushik Dayal. A completely iterative method for the infinite domain electrostatic problem with nonlinear dielectric media. Journal of Computational Physics, 230(21):7821–7829, September 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004050.

Yano:2012:OBF

[YD12] Masayuki Yano and David L. Darmofal. An optimization-based framework for anisotropic simplex mesh adaptation. Journal of Computational Physics, 231(22):7626–7649, September 15, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112003749.

Yamaleev:2010:LTA

[YDN10] Nail K. Yamaleev, Boris Diskin, and Eric J. Nielsen. Local-in-time adjoint-based method for design optimization of unsteady flows. Journal of Computational Physics, 229(14):5394–5407, July 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110001646.

Yuan:2011:CQM

[YF11] C. Yuan and R. O. Fox. Conditional quadrature method of moments for kinetic equations. Journal of Computational Physics, 230(22):8216–8246, September 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004396.

Yang:2014:CNC

[YF14] Chang Yang and Francis Filbet. Conservative and nonconservative methods based on Hermite weighted essentially non-oscillatory reconstruction for Vlasov equations. Journal of Computational Physics, 279(??):18–36, December 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911400624X.

Yang:2013:MSD

[YFL^{+}13] Xiaofeng Yang, M. Gregory Forest, Huiyuan Li, Chun Liu, Jie Shen, Qi Wang, and Falai Chen. Modeling and simulations of drop pinch-off from liquid crystal filaments and the leaky liquid crystal faucet immersed in viscous fluids. Journal of Computational Physics, 236(??):1–14, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112006456.

Yi:2013:DDG

[YHL13] Nianyu Yi, Yunqing Huang, and Hailiang Liu. A direct discontinuous Galerkin method for the generalized Korteweg– de Vries equation: Energy conservation and boundary effect. Journal of Computational Physics, 242(??):351–366, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113000752.

Young:2012:HON

[YHM12] P. Young, S. Hao, and P. G. Martinsson. A high-order Nyström discretization scheme for boundary integral equations defined on rotationally symmetric surfaces. Journal of Computational Physics, 231(11):4142–4159, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112000873.

Yang:2011:DEG

[YHT11] Zhicheng Yang, Peng He, and Huazhong Tang. A direct Eulerian GRP scheme for relativistic hydrodynamics: One-dimensional case. Journal of Computational Physics, 230(22):7964–7987, September 10, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999111004086.

Yin:2014:HPS

[Yin14] Zhaohua Yin. A Hermite pseudospectral solver for twodimensional incompressible flows on infinite domains. Journal of Computational Physics, 258(??):371–380, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007158.

Yuan:2012:NSF

[YJK12] Xuefei Yuan, Stephen C. Jardin, and David E. Keyes. Numerical simulation of four-field extended magnetohydrodynamics in dynamically adaptive curvilinear coordinates via Newton–Krylov–Schwarz. Journal of Computational Physics, 231(17):5822–5853, July 1, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002331.

Yildirim:2012:HSD

[YK12] B. Yildirim and George Em Karniadakis. A hybrid spectral/DG method for solving the phase-averaged ocean wave equation: Algorithm and validation. Journal of Computational Physics, 231(14):4921–4953, May 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112001854.

Yang:2013:RMM

[YK13] Xiu Yang and George Em Karniadakis. Reweighted ℓ_{1} minimization method for stochastic elliptic differential equations. Journal of Computational Physics, 248(??):87–108, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113002520.

Yee:2013:CSB

[YKWS13a] H. C. Yee, D. V. Kotov, Wei Wang, and Chi-Wang Shu. Corrigendum to “Spurious behavior of shock-capturing methods by the fractional step approach: Problems containing stiff source terms and discontinuities” [J. Comput. Phys. 241 (2013) 266–291]. Journal of Computational Physics, 250(??):703–712, October 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113003616. See [YKWS13b].

Yee:2013:SBS

[YKWS13b] H. C. Yee, D. V. Kotov, Wei Wang, and Chi-Wang Shu. Spurious behavior of shock-capturing methods by the fractional step approach: Problems containing stiff source terms and discontinuities. Journal of Computational Physics, 241(??):266–291, May 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000727. See corri-

gendum [YKWS13a].

Yu:2010:NSM

[YL10] D. Yu and A. J. C. Ladd. A numerical simulation method for dissolution in porous and fractured media. Journal of Computational Physics, 229(18):6450–6465, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002573.

Yan:2011:EHO

[YL11] Hongmei Yan and Yuming Liu. An efficient high-order boundary element method for nonlinear wave-wave and wave-body interactions. Journal of Computational Physics, 230(2):402–424, January 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005322.

You:2014:EMC

[YL14a] Guoqiao You and Shingyu Leung. An Eulerian method for computing the coherent ergodic partition of continuous dynamical systems. Journal of Computational Physics, 264(??):112–132, May 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114000618.

You:2014:VET

[YL14b] Guoqiao You and Shingyu Leung. VIALS: an Eulerian tool based on total variation and the level set method for studying dynamical systems. Journal of Computational Physics, 266(??):139–160, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911400134X.

Yang:2010:AGT

[YLW10] F. L. Yang, L. Ling, and T. Wei. An adaptive greedy technique for inverse boundary determination problem. Journal of Computational Physics, 229(22):8484–8496, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004249.

Yeo:2010:SCS

[YM10] Kyongmin Yeo and Martin R. Maxey. Simulation of concentrated suspensions using the force-coupling method. Journal of Computational Physics, 229(6):2401–2421, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109006706.

Yang:2014:AJI

[YM14] Xiyang I. A. Yang and Rajat Mittal. Acceleration of the Jacobi iterative method by factors exceeding 100 using scheduled relaxation. Journal of Computational Physics, 274(??):695–708, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114004173.

Yelash:2014:ADE

[YMLM^{+}14] L. Yelash, A. Müller, M. Lukácová-Medvid’ová, F. X. Giraldo, and V. Wirth. Adaptive discontinuous evolution Galerkin method for dry atmospheric flow. Journal of Computational Physics, 268(??):106–133, July 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114001600.

Yaguchi:2010:EDV

[YMS10] Takaharu Yaguchi, Takayasu Matsuo, and Masaaki Sugihara. An extension of the discrete variational method to nonuniform grids. Journal of Computational Physics, 229(11):4382–4423, June 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110000914.

Yaguchi:2012:DVD

[YMS12] Takaharu Yaguchi, Takayasu Matsuo, and Masaaki Sugihara. The discrete variational derivative method based on discrete differential forms. Journal of Computational Physics, 231(10):3963–3986, May 20, 2012. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112000666.

Yoshida:2010:MRT

[YN10] Hiroaki Yoshida and Makoto Nagaoka. Multiple-relaxation-time lattice Boltzmann model for the convection and anisotropic diffusion equation. Journal of Computational Physics, 229 (20):7774–7795, October 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110003529.

Yoshida:2014:LBM

[YN14] Hiroaki Yoshida and Makoto Nagaoka. Lattice Boltzmann method for the convection-diffusion equation in curvilinear coordinate systems. Journal of Computational Physics, 257 (??):884–900, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006499.

Yan:2011:LDG

[YO11] Jue Yan and Stanley Osher. A local discontinuous Galerkin method for directly solving Hamilton–Jacobi equations. Journal of Computational Physics, 230(1):232–244, January 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110005255.

Yokoi:2013:PNF

[Yok13] Kensuke Yokoi. A practical numerical framework for free surface flows based on CLSVOF method, multi-moment methods and density-scaled CSF model: Numerical simulations of droplet splashing. Journal of Computational Physics, 232(1):252–271, January 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004913.

Yokoi:2014:DSC

[Yok14] Kensuke Yokoi. A density-scaled continuum surface force model within a balanced force formulation. Journal of Computational Physics, 278(??):221–228, December 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114006020.

Yamaleev:2013:NMR

[YP13] Nail K. Yamaleev and Kedar A. Pathak. Nonlinear model reduction for unsteady discontinuous flows. Journal of Computational Physics, 245(??):1–13, July 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113001745.

Yang:2013:TLS

[YR13] Yushu Yang and Muruhan Rathinam. Tau leaping of stiff stochastic chemical systems via local central limit approximation. Journal of Computational Physics, 242(??):581–606, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113001216.

Yang:2011:SVFa

[YS11a] Di Yang and Lian Shen. Simulation of viscous flows with undulatory boundaries. Part I: Basic solver. Journal of Computational Physics, 230(14):5488–5509, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001252.

Yang:2011:SVFb

[YS11b] Di Yang and Lian Shen. Simulation of viscous flows with undulatory boundaries: Part II. Coupling with other solvers for two-fluid computations. Journal of Computational Physics, 230(14):5510–5531, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001240.

Yang:2012:SED

[YS12a] Jianming Yang and Frederick Stern. A simple and efficient direct forcing immersed boundary framework for fluid-structure interactions. Journal of Computational Physics, 231(15):5029– 5061, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112001842.

Yuffa:2012:OOE

[YS12b] Alex J. Yuffa and John A. Scales. Object-oriented electrodynamic S-matrix code with modern applications. Journal of Computational Physics, 231(14):4823–4835, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001799.

Yu:2013:DAC

[YS13] C. H. Yu and Tony W. H. Sheu. A dispersively accurate compact finite difference method for the Degasperis–Procesi equation. Journal of Computational Physics, 236(??):493–512, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911200650X.

Yang:2013:CFB

[YSW^{+}13] L. M. Yang, C. Shu, J. Wu, N. Zhao, and Z. L. Lu. Circular function-based gas-kinetic scheme for simulation of inviscid compressible flows. Journal of Computational Physics, 255(??):540–557, December 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113005639.

Yang:2014:SDF

[YSW14] L. M. Yang, C. Shu, and J. Wu. A simple distribution function-based gas-kinetic scheme for simulation of viscous incompressible and compressible flows. Journal of Computational Physics, 274(??):611–632, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004409.

Yang:2012:DEG

[YT12] Zhicheng Yang and Huazhong Tang. A direct Eulerian GRP scheme for relativistic hydrodynamics: Two-dimensional case. Journal of Computational Physics, 231(4):2116–2139, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111006814.

Yen:2012:SGF

[YTYJ12] Chien-Chang Yen, Ronald E. Taam, Ken Huai-Che Yeh, and Kang C. Jea. Self-gravitational force calculation of infinitesimally thin gaseous disks. Journal of Computational Physics, 231(24):8246–8261, October 15, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112004366.

Yalim:2011:FVI

[YVL^{+}11] M. S. Yalim, D. Vanden Abeele, A. Lani, T. Quintino, and H. Deconinck. A finite volume implicit time integration method for solving the equations of ideal magnetohydrodynamics for the hyperbolic divergence cleaning approach. Journal of Computational Physics, 230(15):6136–6154, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002592.

Ying:2013:KFB

[YW13] Wenjun Ying and Wei-Cheng Wang. A kernel-free boundary integral method for implicitly defined surfaces. Journal of Computational Physics, 252(??):606–624, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113004439.

Yu:2014:AED

[YWL14] Meilin Yu, Z. J. Wang, and Yen Liu. On the accuracy and efficiency of discontinuous Galerkin, spectral difference and correction procedure via reconstruction methods. Journal of Computational Physics, 259(??):70–95, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007857.

Yang:2013:DGM

[YWS13] Yang Yang, Dongming Wei, and Chi-Wang Shu. Discontinuous Galerkin method for Krause’s consensus models and pressureless Euler equations. Journal of Computational Physics, 252(??):109–127, November 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113004397.

Yin:2010:CFM

[YWY10] Li Yin, Jiming Wu, and Yanzhong Yao. A cell functional minimization scheme for parabolic problem. Journal of Computational Physics, 229(23):8935–8951, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004559.

Yu:2014:MLW

[YXD^{+}14] Peicheng Yu, Xinlu Xu, Viktor K. Decyk, Weiming An, Jorge Vieira, Frank S. Tsung, Ricardo A. Fonseca, Wei Lu, Luis O. Silva, and Warren B. Mori. Modeling of laser wakefield acceleration in Lorentz boosted frame using EM–PIC code with spectral solver. Journal of Computational Physics, 266(??):124– 138, June 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001363.

Yang:2013:CNQ

[YXZ13] Xuehua Yang, Da Xu, and Haixiang Zhang. Crank–Nicolson/ quasi-wavelets method for solving fourth order partial integrodifferential equation with a weakly singular kernel. Journal of Computational Physics, 234(??):317–329, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005827.

Yu:2012:IHO

[YYB12] Rixin Yu, Jiangfei Yu, and Xue-Song Bai. An improved high-order scheme for DNS of low Mach number turbulent reacting flows based on stiff chemistry solver. Journal of Computational Physics, 231(16):5504–5521, June 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112002306.

Yaji:2014:TOU

[YYY^{+}14] Kentaro Yaji, Takayuki Yamada, Masato Yoshino, Toshiro Matsumoto, Kazuhiro Izui, and Shinji Nishiwaki. Topology optimization using the lattice Boltzmann method incorporating level set boundary expressions. Journal of Computational Physics, 274(??):158–181, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004112.

Yin:2012:IBB

[YZ12] Xuewen Yin and Junfeng Zhang. An improved bounce-back scheme for complex boundary conditions in lattice Boltzmann method. Journal of Computational Physics, 231(11):4295– 4303, June 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112000939.

Yang:2014:OSC

[YZX14] Xuehua Yang, Haixiang Zhang, and Da Xu. Orthogonal spline collocation method for the two-dimensional fractional sub-diffusion equation. Journal of Computational Physics, 256(??):824–837, January 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113006219.

Zadehgol:2014:INK

[ZA14] Abed Zadehgol and Mahmud Ashrafizaadeh. Introducing a new kinetic model which admits an H-theorem for simulating the nearly incompressible fluid flows. Journal of Computational Physics, 274(??):803–825, October 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114004719.

Zadeh:2011:MCS

[Zad11] Kouroush Sadegh Zadeh. A mass-conservative switching algorithm for modeling fluid flow in variably saturated porous media. Journal of Computational Physics, 230(3):664–679, February 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005565. See com-

ments [HSK^{+}15].

Zandi:2012:EBFa

[ZBS12a] S. M. Zandi, B. Boroomand, and S. Soghrati. Exponential basis functions in solution of incompressible fluid problems with moving free surfaces. Journal of Computational Physics, 231(2):505–527, January 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111005481.

Zandi:2012:EBFb

[ZBS12b] S. M. Zandi, B. Boroomand, and S. Soghrati. Exponential basis functions in solution of problems with fully incompressible materials: a mesh-free method. Journal of Computational Physics, 231(21):7255–7273, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112003622.

Zhang:2010:DIA

[ZCD10] Lei Zhang, Long-Qing Chen, and Qiang Du. Diffuse-interface approach to predicting morphologies of critical nucleus and equilibrium structure for cubic to tetragonal transformations. Journal of Computational Physics, 229(18):6574–6584, September 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002743.

Zhou:2014:CFS

[ZCS14] Yunkai Zhou, James R. Chelikowsky, and Yousef Saad. Chebyshev-filtered subspace iteration method free of sparse diagonalization for solving the Kohn–Sham equation. Journal of Computational Physics, 274(??):770–782, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004744.

Zheng:2011:SOP

[ZCW11] Qiong Zheng, Duan Chen, and Guo-Wei Wei. Secondorder Poisson–Nernst–Planck solver for ion transport. Journal of Computational Physics, 230(13):5239–5262, June 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911100163X.

Zaslavsky:2010:STC

[ZD10] Mikhail Zaslavsky and Vladimir Druskin. Solution of time-convolutionary Maxwell’s equations using parameterdependent Krylov subspace reduction. Journal of Computational Physics, 229(12):4831–4839, June 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001294.

Zhang:2011:CGS

[ZD11a] Yong Zhang and Xuanchun Dong. On the computation of ground state and dynamics of Schrödinger–Poisson–Slater system. Journal of Computational Physics, 230(7):2660–2676, April 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111000076.

Zheng:2011:EBH

[ZD11b] X. Zheng and S. Dong. An eigen-based high-order expansion basis for structured spectral elements. Journal of Computational Physics, 230(23):8573–8602, September 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111004815.

Zhang:2012:CSD

[ZD12] Jingyan Zhang and Qiang Du. Constrained shrinking dimer dynamics for saddle point search with constraints. Journal of Computational Physics, 231(14):4745–4758, May 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112001465.

Ziegler:2011:AHO

[ZDSP11] Jack L. Ziegler, Ralf Deiterding, Joseph E. Shepherd, and D. I. Pullin. An adaptive high-order hybrid scheme for compressive, viscous flows with detailed chemistry. Journal of Computational Physics, 230(20):7598–7630, August 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111003809.

Zenginoglu:2011:HLH

[Zen11] Anil Zenginoglu. Hyperboloidal layers for hyperbolic equations on unbounded domains. Journal of Computational Physics, 230(6):2286–2302, March 20, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110006807.

Zerroukat:2010:SMC

[Zer10] M. Zerroukat. A simple mass conserving semi-Lagrangian scheme for transport problems. Journal of Computational Physics, 229(24):9011–9019, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110004547.

Zeng:2012:SAC

[ZF12] Xianyi Zeng and Charbel Farhat. A systematic approach for constructing higher-order immersed boundary and ghost fluid methods for fluid-structure interaction problems. Journal of Computational Physics, 231(7):2892–2923, April 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111007480.

Zhou:2014:SOA

[ZF14] Qiang Zhou and Liang-Shih Fan. A second-order accurate immersed boundary-lattice Boltzmann method for particle-laden flows. Journal of Computational Physics, 268(??):269–301, July 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114001648.

Zhai:2014:USC

[ZFH14] Shuying Zhai, Xinlong Feng, and Yinnian He. An unconditionally stable compact ADI method for three-dimensional time-fractional convection-diffusion equation. Journal of Computational Physics, 269(??):138–155, July 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999114002010.

Zhang:2011:MTM

[ZFJ11] Jun Zhang, Jing Fan, and Jianzheng Jiang. Multiple temperature model for the information preservation method and its application to nonequilibrium gas flows. Journal of Computational Physics, 230(19):7250–7265, August 10, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111003366.

Zheng:2010:KTB

[ZGSZ10] Lin Zheng, Zhaoli Guo, Baochang Shi, and Chuguang Zheng. Kinetic theory based lattice Boltzmann equation with viscous dissipation and pressure work for axisymmetric thermal flows. Journal of Computational Physics, 229(16):5843–5856, August 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110002020.

Zhang:2010:SST

[Zha10a] Mingyu Zhang. Simulation of surface tension in 2D and 3D with smoothed particle hydrodynamics method. Journal of Computational Physics, 229(19):7238–7259, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003256.

Zhao:2010:HOM

[Zha10b] Shan Zhao. High order matched interface and boundary methods for the Helmholtz equation in media with arbitrarily curved interfaces. Journal of Computational Physics, 229(9):3155– 3170, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109007207.

Zhao:2014:OSA

[Zha14] Shan Zhao. Operator splitting ADI schemes for pseudo-time coupled nonlinear solvation simulations. Journal of Computational Physics, 257(??):1000–1021, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006578.

Zhou:2010:BCD

[Zho10] Yunkai Zhou. A block Chebyshev–Davidson method with inner-outer restart for large eigenvalue problems. Journal of Computational Physics, 229(24):9188–9200, December 10, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004791.

Zhou:2011:ELS

[Zho11] Jian Guo Zhou. Enhancement of the LABSWE for shallow water flows. Journal of Computational Physics, 230(2):394–401, January 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110005309.

Zhou:2014:LBM

[Zho14a] Jian Guo Zhou. Lattice Boltzmann morphodynamic model. Journal of Computational Physics, 270(??):255–264, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002599.

Zhou:2014:NAS

[Zho14b] Zhennan Zhou. Numerical approximation of the Schrödinger equation with the electromagnetic field by the Hagedorn wave packets. Journal of Computational Physics, 272(??):386–407, September 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999114003106.

Zhang:2011:FSB

[ZHPS11] Bo Zhang, Jingfang Huang, Nikos P. Pitsianis, and Xiaobai Sun. A Fourier-series-based kernel-independent fast multipole method. Journal of Computational Physics, 230(15):5807– 5821, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002105.

Zheng:2010:AVM

[ZHS10] Haibiao Zheng, Yanren Hou, and Feng Shi. Adaptive variational multiscale methods for incompressible flow based on two local Gauss integrations. Journal of Computational Physics, 229(19):7030–7041, September 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002998.

Zein:2010:MPT

[ZHW10] Ali Zein, Maren Hantke, and Gerald Warnecke. Modeling phase transition for compressible two-phase flows applied to metastable liquids. Journal of Computational Physics, 229(8):2964–2998, April 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999109007050.

Zhang:2013:LCG

[ZHY13] Na Zhang, Zhaoqin Huang, and Jun Yao. Locally conservative Galerkin and finite volume methods for twophase flow in porous media. Journal of Computational Physics, 254(??):39–51, December 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911300507X.

Ziegler:2011:SDC

[Zie11] U. Ziegler. A semi-discrete central scheme for magnetohydrodynamics on orthogonal-curvilinear grids. Journal of Computational Physics, 230(4):1035–1063, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110005784.

Zhao:2010:SBI

[ZIOF10] Hong Zhao, Amir H. G. Isfahani, Luke N. Olson, and Jonathan B. Freund. A spectral boundary integral method for flowing blood cells. Journal of Computational Physics, 229(10):3726–3744, May 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110000471.

Zhang:2014:RVF

[ZJL^{+}14] Di Zhang, Chunbo Jiang, Dongfang Liang, Zhengbing Chen, Yan Yang, and Ying Shi. A refined volume-of-fluid algorithm for capturing sharp fluid interfaces on arbitrary meshes. Journal of Computational Physics, 274(??):709–736, October 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114004501.

Zhang:2015:CSI

[ZJS15] Shuangxi Zhang, Yuesong Jia, and Qizhi Sun. Comment on “Symplectic integration of magnetic systems” by Stephen D. Webb [J. Comput. Phys. 270 (2014) 570–576]. Journal of Computational Physics, 282(??):43–46, February 1, 2015. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114007505. See [Web14].

Zimmerman:2010:MFA

[ZJT10] Jonathan A. Zimmerman, Reese E. Jones, and Jeremy A. Templeton. A material frame approach for evaluating continuum variables in atomistic simulations. Journal of Computational Physics, 229(6):2364–2389, March 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006688.

Zhang:2012:INO

[ZJW12] Yaoxin Zhang, Yafei Jia, and Sam S. Y. Wang. An improved nearly-orthogonal structured mesh generation system with smoothness control functions. Journal of Computational Physics, 231(16):5289–5305, June 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002471.

Zayernouri:2013:FSL

[ZK13] Mohsen Zayernouri and George Em Karniadakis. Fractional Sturm–Liouville eigen-problems: Theory and numerical approximation. Journal of Computational Physics, 252(??): 495–517, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113004610.

Zayernouri:2014:EAS

[ZK14] Mohsen Zayernouri and George Em Karniadakis. Exponentially accurate spectral and spectral element methods for fractional ODEs. Journal of Computational Physics, 257 (??):460–480, January 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999113006530.

Zou:2010:CGC

[ZKKF10] Yu Zou, Michail E. Kavousanakis, Ioannis G. Kevrekidis, and Rodney O. Fox. Coarse-grained computation for particle coagulation and sintering processes by linking Quadrature Method of Moments with Monte-Carlo. Journal of Computational Physics, 229(14):5299–5314, July 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001178.

Ziemys:2011:HMD

[ZKM^{+}11] A. Ziemys, M. Kojic, M. Milosevic, N. Kojic, F. Hussain, M. Ferrari, and A. Grattoni. Hierarchical modeling of diffusive transport through nanochannels by coupling molecular dynamics with finite element method. Journal of Computational Physics, 230(14):5722–5731, June 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111002166.

Zheng:2010:AKC

[ZKV10] L. J. Zheng, M. T. Kotschenreuther, and J. W. Van Dam. AEGIS–K code for linear kinetic analysis of toroidally axisymmetric plasma stability. Journal of Computational Physics, 229(10):3605–3622, May 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911000032X.

Zhao:2010:DWM

[ZKZ10] Haibo Zhao, F. Einar Kruis, and Chuguang Zheng. A differentially weighted Monte Carlo method for two-component coagulation. Journal of Computational Physics, 229(19): 6931–6945, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999110002925.

Zhang:2010:HSF

[ZL10] Qinghai Zhang and Philip L.-F. Liu. Handling solidfluid interfaces for viscous flows: Explicit jump approximation vs. ghost cell approaches. Journal of Computational Physics, 229(11):4225–4246, June 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S002199911000080X.

Zhou:2013:NMD

[ZL13] Xuan Zhou and Shuixiang Li. A new mesh deformation method based on disk relaxation algorithm with predisplacement and post-smoothing. Journal of Computational Physics, 235(??):199–215, February 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112006250.

Zhang:2014:ESM

[ZLCW14] Bin Zhang, Hong Liu, Fang Chen, and Jian Hang Wang. The equilibrium state method for hyperbolic conservation laws with stiff reaction terms. Journal of Computational Physics, 263(??):151–176, April 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113008516.

Zhou:2012:MIB

[ZLH12] Y. C. Zhou, Jiangguo Liu, and Dennis L. Harry. A matched interface and boundary method for solving multi-flow Navier– Stokes equations with applications to geodynamics. Journal of Computational Physics, 231(1):223–242, January 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111005304.

Zhou:2010:LSB

[ZLL10] Shiwei Zhou, Wei Li, and Qing Li. Level-set based topology optimization for electromagnetic dipole antenna design. Journal of Computational Physics, 229(19):6915–6930, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110002913.

Zaeem:2010:FEM

[ZM10] Mohsen Asle Zaeem and Sinisa Dj. Mesarovic. Finite element method for conserved phase fields: Stress-mediated diffusional phase transformation. Journal of Computational Physics, 229(24):9135–9149, December 10, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004742.

Zadeh:2014:PFP

[ZM14] Kouroush Sadegh Zadeh and Hubert J. Montas. Parametrization of flow processes in porous media by multiobjective inverse modeling. Journal of Computational Physics, 259(??):390–401, February 15, 2014. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113007997.

Zitelli:2011:CDP

[ZMD^{+}11] J. Zitelli, I. Muga, L. Demkowicz, J. Gopalakrishnan, D. Pardo, and V. M. Calo. A class of discontinuous Petrov– Galerkin methods. Part IV: The optimal test norm and time-harmonic wave propagation in 1D. Journal of Computational Physics, 230(7):2406–2432, April 1, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006649.

Zheng:2014:ISM

[ZMD14] X. Zheng, Q. W. Ma, and W. Y. Duan. Incompressible SPH method based on Rankine source solution for violent water wave simulation. Journal of Computational Physics, 276(??):291–314, November 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114005294.

Zhang:2011:MPM

[ZMG11] Duan Z. Zhang, Xia Ma, and Paul T. Giguere. Material point method enhanced by modified gradient of shape function. Journal of Computational Physics, 230(16):6379–6398, July 10, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002804.

Zhang:2013:ATS

[ZMQ13] Zhengru Zhang, Yuan Ma, and Zhonghua Qiao. An adaptive time-stepping strategy for solving the phase field crystal model. Journal of Computational Physics, 249(??):204–215, September 15, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S002199911300315X.

Zakeri:2010:SCA

[ZN10] Gholam-Ali Zakeri and Mitra Navab. Sinc collocation approximation of non-smooth solution of a nonlinear weakly singular Volterra integral equation. Journal of Computational Physics, 229(18):6548–6557, September 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110002718.

Zhang:2014:CCS

[ZN14a] Jie Zhang and Ming-Jiu Ni. A consistent and conservative scheme for MHD flows with complex boundaries on an unstructured Cartesian adaptive system. Journal of Computational Physics, 256(??):520–542, January 1, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113005317.

Zhang:2014:DSM

[ZN14b] Jie Zhang and Ming-Jiu Ni. Direct simulation of multi-phase MHD flows on an unstructured Cartesian adaptive system. Journal of Computational Physics, 270(??):345–365, August 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114002174.

Zohdi:2013:NSI

[Zoh13] T. I. Zohdi. Numerical simulation of the impact and deposition of charged particulate droplets. Journal of Computational Physics, 233(??):509–526, January 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112005438.

Zhang:2011:VME

[ZOJR11] Lin Zhang, Jie Ouyang, Tao Jiang, and Chunlei Ruan. Variational multiscale element free Galerkin method for the water wave problems. Journal of Computational Physics, 230(12):5045–5060, June 1, 2011. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S002199911100177X.

Zhao:2011:OSI

[ZOL^{+}11] Su Zhao, Jeremy Ovadia, Xinfeng Liu, Yong-Tao Zhang, and Qing Nie. Operator splitting implicit integration factor methods for stiff reaction-diffusion-advection systems. Journal of Computational Physics, 230(15):5996–6009, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999111002336.

Zhang:2010:VME

[ZOWZ10] Lin Zhang, Jie Ouyang, Xiaoxia Wang, and Xiaohua Zhang. Variational multiscale element-free Galerkin method for 2D Burgers’ equation. Journal of Computational Physics, 229(19):7147–7161, September 20, 2010. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999110003190.

Zheligovsky:2010:MAE

[ZPF10] V. Zheligovsky, O. Podvigina, and U. Frisch. The Monge– Ampère equation: Various forms and numerical solution. Journal of Computational Physics, 229(13):5043–5061, July 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911000135X.

Zymaris:2010:AWF

[ZPGO10] A. S. Zymaris, D. I. Papadimitriou, K. C. Giannakoglou, and C. Othmer. Adjoint wall functions: a new concept for use in aerodynamic shape optimization. Journal of Computational Physics, 229(13):5228–5245, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001567.

Zhu:2011:LDM

[ZQ11] Jun Zhu and Jianxian Qiu. Local DG method using WENO type limiters for convection-diffusion problems. Journal of Computational Physics, 230(11):4353–4375, May 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001336.

Zhu:2013:HWS

[ZQ13] Jun Zhu and Jianxian Qiu. Hermite WENO schemes for Hamilton–Jacobi equations on unstructured meshes. Journal of Computational Physics, 254(??):76–92, December 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113005123.

Zhang:2013:IPV

[ZRM13] Yunhuang Zhang, Jean C. Ragusa, and Jim E. Morel. Iterative performance of various formulations of the SP_{N} equations. Journal of Computational Physics, 252(??):558–572, November 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113004336.

Zhang:2010:MPS

[ZS10a] Xiangxiong Zhang and Chi-Wang Shu. On maximum-principle-satisfying high order schemes for scalar conservation laws. Journal of Computational Physics, 229(9):3091–3120, May 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109007165.

Zhang:2010:PPH

[ZS10b] Xiangxiong Zhang and Chi-Wang Shu. On positivitypreserving high order discontinuous Galerkin schemes for compressible Euler equations on rectangular meshes. Journal of Computational Physics, 229(23):8918–8934, November 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110004535.

Zhang:2011:PPH

[ZS11] Xiangxiong Zhang and Chi-Wang Shu. Positivity-preserving high order discontinuous Galerkin schemes for compressible Euler equations with source terms. Journal of Computational Physics, 230(4):1238–1248, February 20, 2011. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110006017.

Zhang:2012:PPH

[ZS12] Xiangxiong Zhang and Chi-Wang Shu. Positivity-preserving high order finite difference WENO schemes for compressible Euler equations. Journal of Computational Physics, 231 (5):2245–2258, March 1, 2012. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0021999111006759.

Zhong:2013:SWE

[ZS13] Xinghui Zhong and Chi-Wang Shu. A simple weighted essentially nonoscillatory limiter for Runge–Kutta discontinuous Galerkin methods. Journal of Computational Physics, 232(1):397–415, January 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999112004743.

Zeng:2014:FIV

[ZS14] X. Zeng and G. Scovazzi. A frame-invariant vector limiter for flux corrected nodal remap in arbitrary Lagrangian– Eulerian flow computations. Journal of Computational Physics, 270(??):753–783, August 1, 2014. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999114002411.

Zerroukat:2010:MQS

[ZSW10] M. Zerroukat, A. Staniforth, and N. Wood. The monotonic Quartic Spline Method (QSM) for conservative transport problems. Journal of Computational Physics, 229(4):1150–1166, February 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109005634.

Zahedi:2010:DFA

[ZT10a] Sara Zahedi and Anna-Karin Tornberg. Delta function approximations in level set methods by distance function extension. Journal of Computational Physics, 229(6):2199–2219, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999109006597.

Zhou:2010:NCM

[ZT10b] Tao Zhou and Tao Tang. Note on coefficient matrices from stochastic Galerkin methods for random diffusion equations. Journal of Computational Physics, 229(22):8225–8230, November 1, 2010. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999110003992.

Zarmi:2013:GAH

[ZT13a] Asaf Zarmi and Eli Turkel. A general approach for high order absorbing boundary conditions for the Helmholtz equation. Journal of Computational Physics, 242(??):387–404, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999113000764.

Zhao:2013:RKD

[ZT13b] Jian Zhao and Huazhong Tang. Runge–Kutta discontinuous Galerkin methods with WENO limiter for the special relativistic hydrodynamics. Journal of Computational Physics, 242(??):138–168, June 1, 2013. CODEN JCT-PAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0021999113001319.

Zhang:2013:QCB

[ZTG^{+}13] Liangzhe Zhang, Michael R. Tonks, Derek Gaston, John W. Peterson, David Andrs, Paul C. Millett, and Bulent S. Biner. A quantitative comparison between C^{0} and C^{1} elements for solving the Cahn–Hilliard equation. Journal of Computational Physics, 236(??):74–80, March 1, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112007243.

Zhao:2013:SOR

[ZTL13] J. M. Zhao, J. Y. Tan, and L. H. Liu. A second order radiative transfer equation and its solution by meshless method with application to strongly inhomogeneous media. Journal of Computational Physics, 232(1):431–455, January 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112004664.

Zhu:2010:SWC

[ZTS^{+}10] Huajun Zhu, Lingyan Tang, Songhe Song, Yifa Tang, and Desheng Wang. Symplectic wavelet collocation method for Hamiltonian wave equations. Journal of Computational Physics, 229(7):2550–2572, April 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999109006755.

Zeman:2010:AFB

[ZVNM10] Jan Zeman, Jaroslav Vondrejc, Jan Novák, and Ivo Marek. Accelerating a FFT-based solver for numerical homogenization of periodic media by conjugate gradients. Journal of Computational Physics, 229(21):8065–8071, October 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003931.

Zhang:2010:NFE

[ZW10a] Shuo Zhang and Ming Wang. A nonconforming finite element method for the Cahn–Hilliard equation. Journal of Computational Physics, 229(19):7361–7372, September 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999110003359.

Zhou:2010:ABC

[ZW10b] Ying Zhou and Z. J. Wang. Absorbing boundary conditions for the Euler and Navier–Stokes equations with the spectral difference method. Journal of Computational Physics, 229(23):8733–8749, November 20, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110004444.

Zhu:2012:SOV

[ZWH^{+}12] Yongning Zhu, Yuting Wang, Jeffrey Hellrung, Alejandro Cantarero, Eftychios Sifakis, and Joseph M. Teran. A secondorder virtual node algorithm for nearly incompressible linear elasticity in irregular domains. Journal of Computational Physics, 231(21):7092–7117, August 30, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999112002574.

Zhu:2010:VPF

[ZWJW10] Liyong Zhu, Yu Wang, Lili Ju, and Desheng Wang. A variational phase field method for curve smoothing. Journal of Computational Physics, 229(6):2390–2400, March 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199910900669X.

Zhu:2010:VPC

[ZWL10] Shengfeng Zhu, Qingbiao Wu, and Chunxiao Liu. Variational piecewise constant level set methods for shape optimization of a two-density drum. Journal of Computational Physics, 229(13):5062–5089, July 1, 2010. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999110001452.

Zhang:2012:CHDa

[ZWL^{+}12a] Laiping Zhang, Liu Wei, He Lixin, Deng Xiaogang, and Zhang Hanxin. A class of hybrid DG/FV methods for conservation laws I: Basic formulation and one-dimensional systems. Journal of Computational Physics, 231(4):1081–1103, February 20, 2012. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111003597.

Zhang:2012:CHDb

[ZWL^{+}12b] Laiping Zhang, Liu Wei, He Lixin, Deng Xiaogang, and Zhang Hanxin. A class of hybrid DG/FV methods for conservation laws II: Two-dimensional cases. Journal of Computational Physics, 231(4):1104–1120, February 20, 2012. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999111001835.

Zhang:2013:OEF

[ZY13] Jin-Hai Zhang and Zhen-Xing Yao. Optimized explicit finitedifference schemes for spatial derivatives using maximum norm. Journal of Computational Physics, 250(??):511–526, October 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002945.

Zhou:2010:PFS

[ZYF^{+}10] Chunfeng Zhou, Pengtao Yue, James J. Feng, Carl F. Ollivier-Gooch, and Howard H. Hu. 3D phase-field simulations of interfacial dynamics in Newtonian and viscoelastic fluids. Journal of Computational Physics, 229(2):498–511, January 20, 2010. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999109005312.

Zhang:2013:AMF

[ZYHW13] Na Zhang, Jun Yao, Zhaoqin Huang, and Yueying Wang. Accurate multiscale finite element method for numerical simulation of two-phase flow in fractured media using discrete-fracture model. Journal of Computational Physics, 242(??):420–438, June 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999112007371.

Zhang:2013:NSS

[ZYLK13] Zhongqiang Zhang, Xiu Yang, Guang Lin, and George Em Karniadakis. Numerical solution of the Stratonovich- and Ito– Euler equations: Application to the stochastic piston problem. Journal of Computational Physics, 236(??):15–27, March 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S002199911200695X.

Zhao:2014:PPS

[ZYS14] Xuan Zhao, Yang Yang, and Charles E. Seyler. A positivitypreserving semi-implicit discontinuous Galerkin scheme for solving extended magnetohydrodynamics equations. Journal of Computational Physics, 278(??):400–415, December 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999114006202.

Zhang:2013:NEM

[ZYT13] Yong Zhang, Hong-Liang Yi, and He-Ping Tan. Natural element method for radiative heat transfer in a semitransparent medium with irregular geometries. Journal of Computational Physics, 241(??):18–34, May 15, 2013. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113000971.

Zhang:2014:GFR

[ZZB14] Sijun Zhang, Xiang Zhao, and Sami Bayyuk. Generalized formulations for the Rhie–Chow interpolation. Journal of Computational Physics, 258(??):880–914, February 1, 2014. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113007523.

Zhang:2014:PMP

[ZZK^{+}14] Na Zhang, Peng Zhang, Wei Kang, Danny Bluestein, and Yuefan Deng. Parameterizing the Morse potential for coarsegrained modeling of blood plasma. Journal of Computational Physics, 257(??):726–736, January 15, 2014. CO-DEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0021999113006542.

Zhao:2011:BTS

[ZzS11] Xuan Zhao and Zhi zhong Sun. A box-type scheme for fractional sub-diffusion equation with Neumann boundary conditions. Journal of Computational Physics, 230(15):6061–6074, July 1, 2011. CODEN JCTPAH. ISSN 0021-9991 (print),

1090-2716 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0021999111002373.

Zhang:2013:MPS

[ZZS13] Yifan Zhang, Xiangxiong Zhang, and Chi-Wang Shu. Maximumprinciple-satisfying second order discontinuous Galerkin schemes for convection-diffusion equations on triangular meshes. Journal of Computational Physics, 234(??):295–316, February 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999112005773.

Zhu:2013:RKD

[ZZSQ13] Jun Zhu, Xinghui Zhong, Chi-Wang Shu, and Jianxian Qiu. Runge–Kutta discontinuous Galerkin method using a new type of WENO limiters on unstructured meshes. Journal of Computational Physics, 248(??):200–220, September 1, 2013. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-

2716 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0021999113002611.